A Bibliography of Publications about the Fortran Programming Language: Part 3: 1990–date

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

22 August 2025 Version 2.188

Title word cross-reference

#55 [Och09]. #59 [Cha09].

+ [BMV03]. −1/2, 1/2, 3/2, 5/2 [Mac98]. 1 [WKM04]. 1/2 [PS08]. $14.95 [Ano90b]. $145.00 [Ano98a]. 2 [BKRG22, CMV09, RBS93a]. $22.50 [Ano99a, Ano99b]. $24.95 [Eme94, Ano96a]. 3 [BCE93, Fuj95, SC19]. $50.00 [Ano98b]. $65 [Ano03]. ^{2} [FGCG94]. ^{29}Si [SSLG91]. 40Ar [Xu93]. ^{40}Ar,^{39}Ar [Xu93]. ^{(}R) [LS04]. S [Lav91]. α [Jon92b]. AXB^{T} + CXD^{T} = E [Hop02, GWL^{+}92]. B [Lai92a, Lai92b]. BR[B → Xsy] [DGS08]. C^{1} [Ren04]. D [CHM91]. ℓ [KTMB02]. F [AS93, BKRG22]. f = 1 [BKRG22]. F_{−−} [NSJD98]. L_{1} [Dem03]. N [Hig93a]. p_{1/2+iτ}^{m} (x) [GST12]. π [KS12]. πps [Air04]. q [CHM91, MSP^{+}22]. R^{3} [MC96]. SU(3) [BW12]. t [Som98]. U(a, x) [GST06a, GST06b]. V (a, x) [GST06a, GST06b]. φ [Koi09]. W (a, x) [GST11].

-conjugated [KS12]. -D [WKM04, RBS93a]. -Dimensional [BCE93, CHM91]. -function [Jon92b]. -functions [Koi09]. -is [BN96]. -Kutta [GKKL19]. -Lattice [GAW96a, GAW96b]. -LSKUM [MSP^{+}22]. -nets [Lai92a, Lai92b]. -nodes [SG95]. -Percentiles [AS93]. -state [CHM91].

/Fortran [TBG^{+}02]. /Java [Och09]. /release [Dig90a].

0 [AB93, Ano90b, Ano91a, Gon01, Tay99]. 0-1 [BKK94]. 0-19-853772-7 [Ano91a].

0-262-61094-9 [Eme94]. 0-471-95596-5 [Gon01]. 0-521-43064-X [AB93]. 0-8493-2016-X [Tay99]. 0-87835-337-2 [Ano90b]. 007R1 [W^{+}95, ANS95].

1 [Car93, McB91, Sal92]. 1-D [Car93]. 1.1 [BDPW98, OA02]. 10 [Sun05]. 1003.9 [IEE92a]. 1003.9-1992 [IEE92a]. 1003.9-1993 [IEE93b]. 1005 [JSY^{+}20]. 100k [SC19]. 10th [Glo91b, IEE96]. 11i [TOML04]. 13th [HR92]. 14.9 [SMSY02]. 14/300 [SS93]. 1539-1 [Ame97b, IEC97, Int97a, Int97b, ISO04a, ISO04b, ISO10]. 1539-2 [IEC94, ISO94, Int00]. 1539-3 [IEC99, Int99]. 164 [Tou84]. 16th [FH90]. 186th [Ano95a]. 19 [IFI95]. 1989 [Olv91]. 198x [Ame87]. 1990 [STVS91]. 1990-12 [IEC90]. 1991 [Bar92]. 1992 [AC92, Ame92, HK93a, IEE92a]. 1993 [Hig94a]. 1994 [IEE94f, ISO94]. 1994-12 [IEC94]. 1995 [IFI95, ANS95]. 1996 [IEE96]. 1997 [Ame97b, Int97a, Int97b]. 1997-12 [IEC97]. 1998 [Int98a, Int98b]. 1998-12 [IEC98a, IEC98b]. 1999 [Int99, Met99c, Met99d]. 1999-02 [IEC99]. 199x [Ame90b]. 1st [Ano94d, Fer92, Kum94].

2 [Ano90b, BRH90, Car91b, Car92, CC95a, DV98, Int90e, Mös95, PAK^{+}90, PGH^{+}90, RBS93b, SS09, Sch97]. 2-D [RBS93b]. 2-dimensional [PT93]. 2.0 [AIS^{+}97, BKMC96, KK01]. 2.0.1 [Sun92a]. 2.1 [Wei94]. 2/3 [OC94]. 2/M [FK95]. 20 [Met99c]. 2000 [Coc03, Int00]. 2001 [ACM01]. 2002 [Coc03, Smi00]. 2003 [ACM03, AS14, FB12, MRC04, Mor15, Rei04, RMX05, RRX^{+}08, RAX10]. 2003/2008 [Mor15]. 2007 [Bjø08]. 2008 [Mor15, Sht19]. 2018 [MRC18]. 2023 [MRCB23]. 20th [Cip00]. 21st [ACM94b, Bra03]. 22nd [ACM95b]. 23rd [IEE92b]. 25-27 [Ano94p]. 25-28 [Ano93n]. 25th [Ano94l]. 26th [Ano95b]. 27 [SH07].

28 [SH09]. 29 [FC09]. 292pp [Gon01]. 2D [Azi23]. 2D-xtDCR-LTBEM [Azi23]. 2nd [HOP93, AB93, DT94, FK95, Yan94b]. 2PNO [SH07, SH09].

3 [HBG^{+}06, Hop97, KKS^{+}95, vH10]. 3-540-60529-0 [Hop97]. 3-540-60530-4 [Hop97]. 3.0 [Ano97c, Bra97c, KaM10, MMEH08]. 300/400 [Hew90b]. 3000 [Blu91]. 3090 [CK90, SSW91]. 3090/VF [CK90]. 32-bit [Ano92b, Ano93d]. 3772 [Cra93]. 3DModel4 [Bak91]. 3L [CA92, CA92]. 3rd [Rub93].

4 [Ein95, Hop97]. 4.0 [PS96]. 40 [PAK^{+}90]. 4th [CKMU94].

5 [Ano97d, Bra97d, DT93, Gon01, HP95b, KBKT94]. 524 [Bre78]. 528 [FHS78, GG99]. 5th [Ban93, Fri94, IEE94a, NBC92].

6.3/03 [Ing90a, Ing90b]. 6000

[Bel90a, Bel90b]. 600Js [WTW90]. 64 [AAC^{+}04]. 64-bit [YYX^{+}07]. 679 [DDHD90]. 689 [BB91]. 690 [BD91]. 692 [DGL91b]. 6th [BGNP94, HMPT94].

7 [Ano91a]. 703 [CC92a].

7040/FORTRAN [Anoxx]. 705 [Hop02]. 706 [BE92, Esp98]. 707 [NPB92]. 711 [NS92]. 717 [BGW93]. 720 [BCE93]. 724 [AS93]. 725 [Dre93]. 729 [HC94]. 730 [ARS94]. 734 [Hop98]. 751 [Ren96a, Ren99a]. 752 [Ren96b, Ren99b]. 755 [GJU96]. 757 [Mac96a]. 761 [Aki96, DV00, RB98]. 762 [BLL^{+}96]. 763 [Kea96b]. 769 [Hop03]. 77 [AL92, Ain90, Ain91, And90, BS91a, BBB00, BK06, BSV16, BKMC96, Bor91b, Bro90a, Bro92b, Bro92a, CM92, Cha94c, CC90, CS90a, dCH94, Dem03, Dem06, Dem07, ES93a, Ein95, EKB92, Ins91a, Ell81, Ell90, Ett90, Ett92, Ett93, Ett96, Ett97, FW90, GH18, GHN19, GG99, GP97, GST02a, Gil01, Her90, HB91a, Hop02, KF92d, Lah90, LM90b, Lig91a, Lig91b, Manxx, MC94, MC95a, Mir90, MA90, MN01, NL92, NL95a, NLN96, O’K93, OPE^{+}95, Pag95, Per93, Pre93a, PTV96, RZ94a, RZ94b, Rou90, Sil92a, Sil92b, Spe96a, SWBO93, SOP93, Sil01, SW91, SB92, Spe96b, SF93, Tor91, Tre95, WNO94, Wri91, Yip90, Zim07, ZB94b]. 77-Programmen [EMR93]. 77-programs [EMR93]. 77/386 [CSS90b, CSS91, Zei92]. 77/90 [Bro92a, Bro92b]. 778 [ZBLN97]. 779 [Mac98]. 77D [Cho92, CFH^{+}93]. 77to90 [Pre93b]. 77toHPF [Van94b]. 780 [Ham98]. 786 [Smi98]. 787 [AJ98, RFS98]. 79 [Jam96]. 792 [RB99]. 794 [Wie99]. 796 [DLM99b]. 797 [RR99]. 798 [BM99]. 7ET [Eme94]. 7th [PBG^{+}95].

800 [BBB00]. 801 [WSW00]. 806 [MS00a, MS00b]. 809 [MN01]. 811 [LV01]. 813 [BMR01]. 814 [Smi01]. 815 [FPR01, Has06]. 818 [DV02a]. 819 [GST02a]. 821 [HBG02]. 822 [GST02b]. 828 [Ren03]. 831 [GST04a]. 833 [Ren04]. 837 [ADD04]. 838 [Fab04]. 841 [HD05]. 842 [FGGL05]. 850 [GST06a]. 854 [BK06]. 857 [SMSW06]. 859 [AR06]. 860 [KR94, McB91, KR95]. 861 [Err06]. 863 [Dem07]. 865 [GRW07]. 867 [BB07]. 869 [ZBW07]. 877 [Kod08]. 879 [LS09]. 881 [FGG09]. 886 [CMV09]. 891 [RS09a]. 892 [Jon09]. 893 [Ren09]. 894 [Koi09]. 896 [LMV09]. 897 [HWS09]. 8th [Hua96]. 8X [Ano90c, AO90a, AO90b, AO90c, AKLS88, MR87, MR90a].

’90 [IEE90a, WN90, A^{+}92, AL92, ABW92, ABMS94, Aki99, AFAS99, AR06, And02, Ano92a, Ano93c, Ano93p, Ano94o, Ano95e, Ano97d, Ano99c, AAK01, Bai94, Bai95, Bai05a, Bai05b, Bak95, Ber91a, BDC^{+}96, BRdAHK04, Bla00, BGV94, BGA90, BGA94, BGA96, Bra97b, Bra97d, BG94, Bro92a, Bro97, Buc94a, Buc94c, Com91, CSC^{+}97, Cha95a, CCL01, CCL04, Cha94c, Cha97a, CC92b, CS95, Cou91, Cou97, DLLR96, DP96, DP99, DNS98, DG08, DL97c, Del93, DG99, DDH^{+}95, DDH^{+}96, DDHW96b, DS94, Cro92, Du 97, DV93, ES93a, Ein94, Ein95, Ein96, EPL94a, EPL94b, EPL95, Err06, EC13, FSPC^{+}02, For97, Fur93, Glo91a, Geh95, GK06, GST12, GBDB97, GOT03b, Hah94, Han92, HL94, Hen95, Hop98, Hud96, HLJ95, HLJ98, IFI93, KLM91, Kea95b, Kea96a, Kea96b, KMR96, Ker93a]. 90 [Ker93c, Ker93b, KLM00, Kir02, KS12, KZ94a, KKH10, KH13, KZ94b, Lig93, Manxx, MD97, Mai91, MKS^{+}96, MHT96, MC95b, McC96, Mei95, Mer92b, MR90b, MR91, MR92, Met92a, MRG^{+}93, MR93a, MR94, MR96a, MR99, Met99a, Mil04b, Mit02, MM98, MS93b, Mös95, MHdL12, NDS96, NSJD98, NLN96, NL96, NL97a, NL97b, Ola93, Ola95, Ort94a, PS08, Pre93a, Pre93c, PA94, PTM96, PTV96, Rat95, Red95, Rei92c, Rei92a, Rei92b, Rys95, SS09, Sat97, SS95, SSG^{+}10, SSG^{+}18, SM90, Sch93c, SKM94, Shi98, SM03, Smi95b, Smi01, Som98, SB01, SS10, SSS99, Taq16, Tay97, Tho97a, UM93, VCV97a, VCV97b, Wal93b, WD98, WAG98, WMMW97, Dub97, GMC96b, GMC96a, GMC96c, GMC96f]. 90-enhanced [And02]. 90/95 [Fah02, Cha97a, Ein94, MR96a, MR99, Tay97, DG08]. 90/95/HPF [Met99a]. 90/HPF [FSPC^{+}02]. 900 [Tor10]. 902 [RBD^{+}10]. 903 [CZ10]. 905 [TZW^{+}10]. 9076 [Mra94]. 90D [BCF^{+}94a, Ano94h, BCFH93, BCF^{+}93a, BCF^{+}93b, BCF^{+}93c, BCF^{+}94c, BCF^{+}94b, BCF^{+}94d, Cho92, CFH^{+}93, PHHF94a, PHHF94b, PH96, Pon94a, Pon94b, Rot93]. 90D/High [BCF^{+}94a]. 90D/HPF [BCFH93, BCF^{+}93b, BCF^{+}93c, BCF^{+}94c, BCF^{+}94b, BCF^{+}94d, Pon94a, Pon94b]. ’90s [Edg92]. ’91 [ACM91, IEE91]. 911 [Smi11]. 912 [Kod11]. 914 [GST11]. 916 [ZA11]. ’92 [IEE92b, IEE92d, KVK92]. 926 [Tho13]. ’93 [ACM93b, Ano93q, GGK^{+}93, IEE93a, IEE93d, KSW93]. 934 [EC13]. 936 [Kro14]. 937 [CS14]. ’94 [ACM94b, BLT94, BGG^{+}94, CGS94, DW94, Fri94, HMPT94, IEE94b, IEE94f]. 94-VAPP [BV94]. ’95 [ACM95b, HAM95b, Hua96, IEE95a, ABM^{+}97, Ano94o, Ano98c, Bee01a, Bee01c, Bro03, Cha97a, CS00, Cou97, DDF10, DV98, DVY00, DV01, DV02a, DV02b, Ein94, ECS96, FGJB19, Geh96, GT03, GT07, GRE99, GRW07, Jon09, KaM10, LS05, MR96a, MR99, MRC04, Moo95a, Moo95b, RMX05, RRX^{+}08, Sch99, Sch03, Sun05, Tay97, ANS95, vWAH^{+}02, vH06, vH10, Gen06, Hin06, Iha06, Sch07]. 95-007R1 [W^{+}95, ANS95]. 95/2003 [MRC04, RMX05, RRX^{+}08]. 9593-1 [IEC90, ISO90]. 9593-1-1990 [Ame97a]. ’96 [ACM96a, ACM96b, IEE96]. 961 [BSV16]. 9th [IEE95a].

= [Gom90b, RD91].

A. [Tsa01]. A.R [Gon01]. A1 [Bre78, Bre79]. AASHTO [Cro90]. ABBPACK [MKFB92]. ABD [AR06]. ABDPACK [MKFB92]. Abel [WJ94]. Aberth [Bin96]. Abilities [WR93]. ability [Cho91, TJ90]. Absoft [Ano96b]. Abstract [CS90b, SW94, SKP91, CM91, MKS^{+}96, SKM94]. Abstracting [Lor19]. Abstraction [Sug95, CS90b, RRX^{+}08]. abstracts [Sch93b]. accelerated [MSP^{+}22, iSYS12]. Accelerating [MVK25, SIOS02]. Acceleration [HJ97, HE13]. Accelerators [AC17]. Access [Ham93, KNS95b, LP92, LP93, BK89, BxCW01, KNS95a]. accesses [DSv94]. accessible [BDH^{+}05]. accommodate [SW91]. accompany [BS91b]. Accomplishments [SZAB98]. Accuracy [RB99, Aki96, DV00, Nak90]. accurate [PG10, Wal93b]. achieved [DPR94]. Achievement [MFI^{+}94]. acid [TRS91]. ACM [ACM97, Bee02, IEE02, PEP92, HOP93, ACM93c, PPP93, ACM93b, Ano95c, Kar95, RB99, Ham85, HM90a, RH84]. ACM/IEEE [ACM97, Kar95, ACM98]. acoustic [NJ94b]. Acoustics [KG99]. ACPC [Vol93]. ACPS [BH90]. ACRITH [Wal93a]. ACRITH-XSC [Wal93a]. Across [Bra97a, Coc03, Fri96]. ACSL [GOBG^{+}94]. ACSL-Model [GOBG^{+}94]. Activation [Ano90a]. AD [RP12]. Ada [Boo81, BH90, Cha09, FBC96, Gli96, Moo95a, Moo95b, Mor81, Och09, WBS97]. Adams [Ano98b, GMC96f]. adapted [Lav91]. Adapting [Fat94, Mer92b, GG99]. Adaptive [BE92, BCE93, DN09, KK94, Mit02, AES^{+}96, CC94, Esp98, GC03, HMS^{+}95, SPM^{+}94, WKM04]. adaptor [BV13, BZ94]. added [CA90]. addendum [Hew91b]. Adding [SZAB97]. Additions [HMT90]. Address [SSC00, TR96, SJ94]. Addresses [CGL^{+}95b, CGL^{+}93]. Adelaide [NBC92]. ADF95 [Str05]. ADIFOR [BCC^{+}92, BKMC96]. Adjoint [GK06, GRSS02]. Adjoints [NR06]. adjusted [ZMR^{+}91]. ADOL [GJU96]. ADOL-C [GJU96]. adoption [NDSG07]. Advanced [AMC01, Ben95, CZM94b, CZM94a, Don95, MKF95, MCAB^{+}02, Tem96, Wil95a, Wil95b, BLT94, Ben99b, CMZ94b, CMZ94a, FSPC^{+}02, PGH^{+}90, CMZ95, Ano96a]. Advances [FHP^{+}12, IEE97, Nic91]. advantage [VKB93]. Advantages [Rei92c, Rei92a, Rei92b]. advective [Car93]. advice [Uni2 ]. Aeroacoustic [NOL97]. aerodynamic [Con92]. AeroFcn [Con92]. aeronautical [Gro91]. aerospace [MZ00, MZ01]. Affine [SSC00]. after [Met92b]. against [BSPF01, BSB^{+}03, Ste91]. age [HK95]. ahead [Ano95d]. Aid [CT90, GV92, Gou93, Mil91]. aide [RD91]. Aided [IEE94g, Osy92, Bar92, HT91, IJCL96]. AIMS [Yan94a]. Airshed [SS00]. Airy [Fab04, GST02a]. AIX [Int90c, Int90d, Int90a, Int90b, Int90m, IBM93]. AIX/ [Int90m]. AIZ [GST02a]. al [Kon94, Ede90, Kon94, Tha93, Wu93]. Alan [Mil04a]. Albuquerque [IEE91, ACM93b]. Alexandria [Ano94d]. Algebra [DGL91b, DGL91c, DGL91a, DDHD90, DCHH88b, DCHH88a, DV98, DHP02, GGHvdG01, WD98, ACIK97, CWB92, CWB94, Coo95, GL10, Jon92a, Jon92b, Kea92, Lan90a, LFG00, Mal91, Mat90]. Algebraic [ACM94c, DDF10, Lev95a, Sen03, Ste95a, WN90, CC98, HBG^{+}05, KM99]. algebru [Mal91]. Algol [Wil93]. Algorithm [ARS92, ARS94, Amo90, AFS94, BS97, BGKZ91, Bai93a, BBCH95, BE92, BCE93, Bou97, BG97, Buc94a, Buc94b, CJL97, CP93, CV94, CT95, Cas89a, Cod93a, Cod93b, Cos97a, Dem95, DLS95, DGR92, DCHH88b, DV02b, FJS97, GGLM88, GWL^{+}92, GP97, GM97, GWDL08, IK96, JP95, KTMB02, KDDH94, Kea96a, KNS95b, Kra94, LS00, LP19, MKFB92, MR93b, MR95a, MN11, MGH81, Nat92, O’B93, PR91, PPR97, RG90a, Ren97a, Ren97b, RPL96, RH84, SF92, SH07, SM95, Shi93b, SW91, Smi91, SWH15, SHCP91, TZW^{+}10, TS25, Ves91, WAG98, WMMW97, Zag16, ADD04, ADH95, Coo94, Err06, EO94, Esp98, GC03, GST12, Gou93, GWDL10, Han92, HWS09, KNS95a, KY94, Koi09, KP93, MC96, MKC92, PW84, SG95, vV90, AS93, Aki96, ADD04, ARS94, AR06, AJ98]. Algorithm [BB07, BBB00, BK06, BSV16, BE92, BCE93, BM99, BD91, BMR01, BB91, Bre78, Bre79, BLL^{+}96, BGW93, CMV09, CC92a, CZ10, CS14, DLM99b, DV00, Dem07, DGL91b, DDHD90, Dre93, DV02a, Err06, EC13, Esp98, Fab04, FPR01, FHS78, Fox79, FGGL05, FGG09, GL90, GG99, GST02a, GST02b, GST04a, GST06a, GST11, GJU96, GRW07, Ham85, Ham98, HC94, HBG02, HH18, Has06, HWS09, Hig91, HM90a, Hop98, Hop02, Hop03, HD05, Jon09, JSY^{+}20, Kea96b, Kod08, Kod11, Koi09, Kro14, KHS17, LS09, LV01, LMV09, Mac98, Mac96a, MS00a, MS00b, MN01, NPB92, NS92, RBD^{+}10, RBD^{+}11, RS09a, Ren96b, Ren96a, RB98, RB99, Ren99a, Ren99b, Ren03, Ren04, Ren09, RFS98, RR99, Smi98, Smi01, Smi11, SMSW06, TZW^{+}10, Tho13, Tor10, Wie99, WSW00, ZA11, ZBLN97, ZBW07, SZG95]. Algorithmen [EMR93]. Algorithmic [FHP^{+}12, JSY^{+}20]. algorithmique [Robxx]. Algorithms [CFGG94, Cip00, DH92, EMU96, EMUP98, FGCG94, Ham85, HM90a, HHK94, Kea95b, MJR93, MT90, ONT95, RB99, RH84, SD92, Ste95a, TDMC97, WMMW97, dSZP92, Ano97a, BID95, Din99, EMR93, GJU96, Hop97, HMT90, LV01, Mal91, Num05, Rat95, SD93, Swa84, vPMF92, LMV09]. algoritmov [Mal91]. Alias [MHT96, NI03]. aliasing [LR91]. Align [HCLJ03]. Alignment [CGS93, HCLJ03, CZM93b, CMZ93a, Cha93, WI94]. alignments [vKS94]. Allocation [BCT94, MR93b, Tal91, CCK90, KH93, RMX05, RFRH96]. allow [Ano92b, OJ09]. alluvial [MB92]. Almost [MKFB92, Sch93c]. Alpha [Sha95, Jon92a, Sil01]. alpha-function [Jon92a]. Already [Sch93c]. Altarelli [KKK95, KL92]. Altarelli-Parisi [KKK95, KL92]. Alternate [Bro92a]. Alternative [Sha95, CZM93b, CMZ93a]. Alto [ACM01]. AM1 [HKŠ^{+}97]. AMAD [CA90]. AMBER [HKŠ^{+}97]. AMD [ADD04]. Ameko [McC95]. American [Ame90b, Ame92]. Among [SWM95, GLS93, SB91, SFB92, vV90]. AMPHAX [Cum90]. amplifier [MIN^{+}95]. AMT [CWB94]. Analyse [RD91]. Analyses [CI96, Ber92, CI98, Nar95, WCN92, YH93]. Analysis [AMC98, AM90, BH90, Cok95, CL94, DJ92, DFS95, EO91, FC09, Ger94b, GS97, HK92, Harxx, HHLS90, Kam00, KOM94, KKH10, LR94, LH92, Mas93b, MP93, Oku95, Pao99, RD92, SNK06, SVD96, SDv98, SAS90, SF93, Ueb97, vKK^{+}93, Agt94, AI90, AZ98, Bet97, BGH^{+}06, Bli90, Bra94a, CK86, CH96, CKT85, Cre90b, DSv94, Dot93, EKB92, GV92, HK90, HIK90, Hor23, IMS90b, IMS91c, IMS91g, IMS91h, KW94, KOM93, KH93, Kor99, KSM95, LPA95, Lef93, LN91, LYZ90, dLJEB95, LFG00, MH91, MSP^{+}22, Mas94, MHT96, Mir90, MDV07, Pao01, PBU95, PW93, RD91, SRH96, iSYS12, SZ90, SSG94, Tay99, Unixx, Zah92, vKK92, von92]. Analyst [Pap93]. analytic [AC16, BDH^{+}05]. analytical [Var97]. Analyzers [Dya95]. Analyzing [CHL94, HMW91, LW07, Sze90, HW95, HMW93]. Anasazi [BHLT09]. Anecdotes [Tom99]. Angles [Wal01, Wal02b]. anharmonic [TS06b]. animation [UHP91]. anisotropic [Azi23, KYSV^{+}15, MA09, YSVM^{+}16]. anisotropic-DCR [Azi23]. anisotropy [Mai90]. ’Annai [CEF^{+}95]. Anniversary [Ano93n]. announcement [SSG^{+}18, vH07]. Annual [ACM93c, IEE92b, van90a, ACM91]. anomaly [HKMC90]. ANOVA [WCN92]. ANOVA-based [WCN92]. anQCD [AC16]. ANSI [Ano98b, Ame87, AC92, A^{+}92, Ame96, Ame97b, Ame97a, ABW92, ABM^{+}97, Ein95]. ANSI-C [Ein95]. ANSI/ISO [A^{+}92, Ame97a, ABW92]. ANSI/ISO/IEC [Ame97b]. Any [See04, Kah01]. AP1000 [HDH^{+}94, HDH^{+}95, SIDH95]. AP87 [HM92]. APD [KP92]. API [Ins92, IEE92a, IEE93b]. APL [AP90]. apparent [CNP91, Dut94]. appendix [Mal91]. Application [AS92, AS91, AAK01, BCS00, BCS01, BC01, BGLP94, Fox94, Gar91a, Gar91b, GLPE97, Hem94, HIM91, Hun00, IEE92a, JBBH93, PHHF94a, YFH97, AFAS99, AH90, Ame90a, CN94, CWB94, GT92a, GT94, Ins92, Mal91, NG93, PSG03, Pel93, Sun93, GT92b]. Applications [ASS93, ASS95, Ano02, Ben99a, Bra94c, BCC^{+}96a, BCC^{+}96b, BCH^{+}06, BSPF01, CNBB96, CZM94b, CZM94a, CHKM93, DG94, FGRT00, Fer92, FK95, GS90b, GS01a, Gli96, HRW^{+}98, Irv91, JPE20, Jou95, KF92c, KSW93, LK93a, McD93, Nat00, Oku95, Pas95, RRM^{+}15, RZ94b, SN94, Yam95, AAS93, All93, BLT94, Ben99b, Ben00, Bra94d, BCC^{+}97a, BCC^{+}97b, BxCW01, BMV03, BSB^{+}03, Cen91, Cha94a, CMZ94b, CMZ94a, CMVZ94, CMZ95, DDcMR96, DSZ94, DKMS91, Don95, FGBN19, FG93, GBR15, GS90a, GB95, GR92, HZ99, IMS91b, IMS91f, IMS91d, IMS91e, KF93b, Law01, MM94, MZ00, MZ01, NBC92, PD96, Rap90, RBS93a, RBS93b, SRH96, SM02a, SFKL02, SIOS02, TMD13, YYX^{+}07]. applicative [OM90]. Applied [EK01, Glo91b, JSW93, KaM10, MA18, Mat90, Lev94]. Applns [KF92a]. Applying [CC93]. Appreciation [Rei96b]. Approach [ASS95, BCFH93, CS90a, CL93, HLJ98, Jéz93, Schxx, BC97, BCF^{+}94c, BHS92, CK86, EKC95, GBC92, HM92, KHJS94, SS09, Wag94, WW94, WTW90]. Approaches [CC93, SM02b, Rei97]. approche [LMG95, Lig93]. Approximants [CJL97]. Approximate [PPR97, RFS98, ADD04, FPR01, Has06, Hop03, RPL96, RR99]. Approximating [Gro90]. Approximation [BH92, Dem97, Dem07, MSA03, MKC92]. approximations [Mac96b]. apps [Ano92b]. APR [Wag94]. April [CKMU94, DR94a, Fri94, GH94a, GH94b, GH94c, IEE95a, IEE96, KSW93, Sie94a, Sie94b]. aquifer [MSZ90]. Arbitrary [Cap98, Ves91, Ple93, vH06, vH07, vH10]. arbitrary-order [vH06, vH07, vH10].

Arccosine [HFT97]. archetypal [HKM98]. Architectural [Ano94a, CHKM93, HDH^{+}94, HDH^{+}95]. Architecture [AAC^{+}04, Ano93b, MS94, AHJS90, BT01, CMVZ94, Par86, WMCU97, YYX^{+}07]. Architectures [BKP93, HHK94, Mer92b, Sab95, TLS91, BZ99, CGS94, HMPT94, Lan90a, TLS90, ZCP95, vPMF92]. Arcsine [HFT97]. area [BDH^{+}05, Deu90]. ARGON.f90 [BOPC05]. Argonne [BRH90, KLM91]. Argonne/GMD [BRH90]. Argument [Amo90, Kod08, Kod11, GST04a]. Arguments [NPB92, Tho13, TS25, GST04b]. ARIMA [Bel11]. Arising [MKFB92, WW90]. Arithmetic [Bon06, Bre78, Bre79, BHY80, CT90, Cse99, Kul95, Oku95, Sch99, Smi91, Smi98, SP91a, SP91b, Sun05, TOML04, VCV97b, AH92, AAK01, BBZ95, EP92, HM92, Sch03, Smi01, VCV97a, Vig93]. Arithmetics [FGG09, FGGL05]. Arles [Van95]. Arlington [IEE92c]. ARLOSS [Xu93]. Array [BBZ94, CCL01, CI96, HM96, HLJ95, HLJ98, HLJ01, KHS96, Mas93b, McC96, Mer92b, Rod90, SW94, Ste95b, SOG94, TCF94, Vio90, AW94, AKLS88, ARB94, ARB95, BBDR94, BBDR95, CCL04, Cha93, CD03, HK90, KHS95, KHRS95, Kir02, LPA95, Mas94, NRK98, OH90, PQ94, Rei02, TOC18, WW94, WW95, WI94, BCS00, BCS01, BC01, MDV07, NR98a, NR98b, Num05, Wal02a, CDMC06]. Arrays [Ber91a, CGS93, Mey01, Ros93, vDSP96, AH90, Bec91, BSCV95, CS90b, DSv94, RBS92, RN07]. arrival [Hor23]. arrival-time [Hor23]. arsenide [SMB90]. Art [Adl93, Kon94, Lev98, PTM96, PTV96, Yan94b, AB93, DeT90, Loz98, Nak90, PTVF92, Pre94b, Sch91a]. Artificial [BPG94, HR92]. Artistic [Mil93]. ARVO [BDH^{+}05]. ASCII [KP92]. Ascona [DR94a]. Asked [Ola93]. ASL [FGRT00].

ASME [KRB^{+}90]. Aspects

[PMBH93, Per94]. ASPLOS [Ano94a]. ASPLOS-VI [Ano94a]. Assessing [EP87, Nan93c]. Assessment [TAH^{+}01, Boi97]. Assignment [EB98, MR95a, PPR97, Ste95b, YKK96, Bar94, Hop03, KY98a, KY98b, RPL96]. assimilation [HBD^{+}93, ZZN94]. assist [Kik93]. assistant [CTS96]. Assisted [Nis95]. Assists [Tho90]. associated [Coo94, SSG^{+}10, SSG^{+}18]. Association [Hig94d]. associative [dSZP92]. AST [Cod90a]. astrophysical [RBS93a, RBS93b]. Asymmetric [CT95, NG93]. ATAN2 [Wal01, Wal02b]. Athens [HMPT94]. Atkinson [Ede90]. Atlantic [Coc03]. Atlas [Tho97a, Ano98a]. Atmospheric [HK93b, HK93a, NLVE15, PFS^{+}04]. Atomic [FGJB19, BF92, Hor09, RPG^{+}20]. atoms [MCA17]. Attempts [Gil91b, Gil91a]. Auckland [BG94]. AUGMENT [BHY80]. August [Agr95, Ban93, BGNP94, CGS94, HAM95b, Hua96, KRB^{+}90, PBG^{+}95, Sen03, Van95, WN90]. Austin [IEE94b]. Austria [BV94, Vol93]. Austrian [Fer92, FK95]. Austrian-Hungarian [Fer92, FK95]. auto deriv [SPF00, SF10]. autocorrelation [NVFNP93]. automata [SC19]. Automated [Che91, JPE20, Kin92, RFRH96, Van22, Yan94a]. Automates [Ano92c]. Automatic [AAN^{+}93, AK84, Bai92, Bai93b, BKMC96, BKK94, BEH^{+}94, BB96, BCH^{+}06, Cha93, CGSS94, CMKH03, CP94, DLS95, EJLC97, FGL01, FBZ92, Ger94a, GP92, Harxx, HZ99, Hor92, KK95a, KK98, LS90a, LS90b, LS00, LP98a, Maaxx, NL19, NVG94, Pre93a, RVV^{+}92, RD91, Sar97, SD01, SD03, UNF^{+}08, WI94, YH93, AK93, AGG^{+}97, BBB^{+}57, Bli90, CDGM96, EO94, GRSS02, GJU96, GB92, GKH^{+}93, LMJC96, LCD91, LP99, Lop90, Lov92, LP90, Mar92, O’K93, OPE^{+}95, SPF00, SF10, Str05, YB13, vH06, vH07, vH10, AJ98]. Automatically [Bla00]. automating [Cre90b]. Automation [Che95, LD90]. automatiques [RD91]. automatisée [Tro90]. Autonomous [NJ94c]. Autotasking [EO91, Nag90]. Autotransformation [AAN^{+}93]. auxiliary [HM92, Lin90]. Avenarius [Kro90]. Averaged [BD90]. Award [Lew94]. Aware [Wal91b, Zim07]. Away [Lor19]. axes [Cum90]. axi [Raj95]. axi-symmetric [Raj95]. AXIOM [JT94]. AXP [Ano97b]. azeotropic [Cra95].

B [Adl93, AB93, Eme94, MN11, UMM94, Yan94b, WKM04, ZBLN97]. B-spline [WKM04]. Babbage [BDH90]. BABDCR [AR06]. BACCII [CB94]. Backus [Bjø08, Aik07, Loh07, McJ17c]. BACOL [WKM04]. BADGER [HM12]. Balanced [KE93, vK94, vHK00]. balancing [Kin92]. Baltimore [IEE02]. banded [HD05]. Bangalore [Kum94]. Barbara [Ano95c, IEE95a]. Barcelona [ACM95a]. Barlow [Ano99a, Ano99b, Gon01]. Barnett [Ano99a, Ano99b, Gon01]. barotropic [KTP^{+}24]. Base [Ame97b, IEC97, Int97a, ISO04a, ISO04b, ISO10, Int97b]. Based [PEP92, Bai94, Bai95, BGLP94, DLM99b, For95, GGLM88, GL90, GLPE97, vHKS94a, HKS94, HIM91, HRW^{+}98, IGHG^{+}94, MSC96, PMBH93, Rotxx, SM03, TOML04, Ber92, CGL^{+}95a, Che90, CCJ93, DDH17, GV92, Ger98a, Ger98b, HW95, HZ94, Hun00, KO91, Koi09, MKS94, Nat00, Nai17, NIY^{+}94, NOL97, Num05, OP98b, SFKL02, SSG94, TS06b, WCN92, vKS94, vHKS94b, Dub97, Che91, KLV98, Sal92, WCN92]. bases [HKŠ^{+}97]. Basic [DGL91c, DGL91a, DDHD90, DCHH88b, DCHH88a, DV98, DHP02, HL94, HMKN91, WC92, SZG95, Dot93, FL91, RS92b, DGL91b, Jon93, RS92a, Sco93, nY90]. Basics [Cor92]. BASIN [HM93]. BASINMAT [Ano90a]. basins [HM93]. Basis [AAN^{+}93, MKFB92, TR96, PZY16, TS06b]. basis-set [TS06b]. batch [Phi91b, Phi92]. battle [MWM90]. Bayes [MHdL12]. Baym [KKY99]. Be [VJ97a, DPR94, FTPR04, VJ97b, Wal91b]. Beam [Mit93, Bec91, MKF95, QRH00]. Becomes [Rys95]. bed [Dut94]. Began [Mey00]. beginning [Gla92b]. Bell [DKMS91, STVS91]. Benchmark [McC95, Pre93c, PA94, SF02, Bak91, DS02, HJJ^{+}00, KLV98]. Benchmarking [BSPF01, BSB^{+}03, Nag95, PAK^{+}90, BGH^{+}06]. Benchmarks [AHOK02, MMY95b, BGH^{+}06, MMY95a, NNON02, VSH91, WYJ99, Cyb91]. Bending [Mit93, Dot93]. benefits [Wic89]. Berlin [Hop97]. Bessel [BBZ95, CRS90, GST04a, GST04b]. Best [Cip00, Dem03]. Better [BBCR98, CB94]. Between [Jéz93, Sil01, van90b, BID95, GRE99, MNZ90, MHT96, Nai17]. Beyond [AS91, HKŠ^{+}97, Sch93a, SC19]. BFGS [MN11, ZBLN97]. BHESS [HD05]. Bibliography [Bee96b, Bee96c, Bee96a, Bee97, Bee02]. BIEMS [MHdL12]. Bifurcation [Nis95]. BIGD.FOR [Nie92]. Bilevel [CV94]. Binary [Nan93c, Hig93a]. Binding [Ame96, DCR99a, Ins92, IEE92a, IEE93b, Par94, Coo95]. Bindings [Ame97a, JPE20, MFK09, Cha09, IEC90, IEE90b, ISO90]. Biochemistry [GDS94]. biocides [RKMJ92]. Biological [VBB18, CH96]. Bit [Kar96, Ano92b, Ano93d, Ano96b, YYX^{+}07]. Bivariate [CMV09]. BIZ [GST02a]. Blanch [Err06]. BLAS [Bee01b, DDP94, DD99, DH92, DVY00, DV01, DV02a, DV02b, DHP02, Hig90b, JSY^{+}20, KLV98, Lin93, MFK09, Per93, Phi91a, She92]. BLAST [Ano96b]. blend [Cra95]. Blinn [Bli94]. Blitz [AJJF14]. Block [ASS93, ASS95, DDP94, DH92, HMKN91, MKFB92, CH98, GRW07, HC08, Koi09, LW07, SZG95, VRT97, WO96]. block-cyclic [HC08, VRT97, WO96]. block-sparse [SZG95]. Blocked [DD99, vPMF92]. Blocks [BDK91, Que00, Deu90]. bodies [CA90, Raj95]. body [CZ10, MB95, MM02, ADB94]. Bondi [Rib02]. Book [AB93, Ano90b, Ano91a, Ano96a, Ano97a, Ano98a, Ano98b, Ano99a, Ano99b, Ano03, BCM99, Dub97, Eme94, EMUP98, GMC96b, GMC96a, GMC96c, GMC96d, GMC96f, GMC96e, Gen06, Gla92a, Hin06, Hop97, Iha06, KG99, Kon94, Kri86, Lev98, Loz98, Mar98, Olv91, Rag95, Sch07, Sch91b, Tay99, TDMC97, UMM94, V^{+}93, Wei94, Yan94b, dL12, Ano92e, VTP92, Vet93]. books [Met99c, Met99d]. Bookshelf [Rys95]. boost [Ano93j]. Bootstrap [Alo04]. bordered [AR06]. borehold [Dut94]. Borland [Hol94, Mit92]. Born [BG93]. Bose [BKRG22, KRG21, KLM^{+}19, TS06b]. Boson [HHCS95]. bosons [LSZ92]. Boston [Ano90b, KRB^{+}90]. Bouclettes [Bou96]. Boulder [Ano03, Sch93b]. Boulevard [ACM99]. Bound [ZBLN97, MN11, ZBW07]. Boundary [AJ98, Cos97a, Cos97b, CA90, EH07a, Gao05, Gao06, McG91, SSS99, KG99]. boundary-only [Gao05]. Boundary-Valued [Cos97a, Cos97b]. Bounding [HK93c]. Bounds [Nak95a]. Box [FJS97, Lai92a, Lai92b]. brackets [MBGK11]. Brain [Ano98b]. Brainerd [Ano98b, GMC96f]. Branch [BL93a, BL93b, PR91, Sht19, CGL^{+}95a]. Branch-and-Cut [PR91]. break [Mas92b]. Brent [BHY80]. Brest [IEE94c]. Brian [Gar93, Loz98]. Bridge [OPB^{+}20]. brief [Gri93, SSG93]. Brinch [Off98]. Bringing [FKKC96]. Bristol [Tay99]. BRKF45 [Cas89a, Hig91]. broadened [NG93]. broadening [NG93]. Bromhead [DeT90]. browser [Nai17]. browser-based [Nai17]. BTN [NS92]. Budapest [FK95]. buffers [MR96b]. bugs [DS02]. Builder [DL97c].

Building

[Cro92, Fos93, LJO05, Du 97, Fos95]. bulletin [Num90c]. BVSPIS [Cos97a]. Byte [Mas93a, Mas92a]. Bytecode [SD01, SD03].

C

[Ano96b, Ano98b, Ano99a, Ano99b, Ano02, Ano03, BBG^{+}93, BCM99, GMC96f, Gon01, JPE20, KSW93, Mit92, Tay99, TBG^{+}02, All90, AS91, Ano95g, Ano99c, BN93, BN96, Bee01b, Bee01d, Bee01g, Bee01e, Bel90a, Bel90b, BSS92, BV13, Bro90b, Bru96a, Bru96b, BSPF01, BSB^{+}03, CG96, CSC^{+}97, Cha09, CC94, Che95, SG93a, SG93b, SG93d, SG93c, DNS97, DFL92, Dot93, Ein95, EKC95, FGMS90c, FGMS90d, FGMS90a, FGMS90b, FGMS93, FGMS95, FCHE02, FH92, FES05, GH18, GHN19, GMF18, GPS99, Goo90a, Goo90c, Goo90e, Goo90f, Goo90d, Goo90b, GRE99, GJU96, Hol94, Hop97, Hor96, Int92, IBM93, Irv91, KM99, Kef92, Ker91a, Ker91b, KLS94b, KYSV^{+}15, KLM^{+}19, Lan01, Lar93, Lee90, LS04, Lem93a, Lem93b, Lem93d, Lem93c, Lev95b, Lev97, Liv91, Mac91b, Mac91a, Mac91c]. C [MSP^{+}22, MFK09, Mey00, Mey01, MMG98, Mös95, NLBB23, Phi91b, Phi92, PMM^{+}08, Poh97, PGH^{+}90, Pug94, RP93, Ros93, RAX10, RMX12, Sch91a, SD92, SD93, SAC^{+}92, Tee90, Tho86, Vel97, VJ97a, VJ97b, VCV97a, VCV97b, Wea94, YGS^{+}94, YSVM^{+}16, EMUP98, TDMC97]. C-Ghinsu [Liv91]. C-H-O-N-S-Ar [BSS92]. C-Language [SG93a, SG93b, SG93d, SG93c]. C-O-H [Lar93]. C-Shell [Phi91b, Phi92]. C-Tree [Ano96b]. C. [GMC96a, Sch91b]. C/C [GJU96, Lan01, VCV97a, VCV97b, MMG98]. C5 [MGH81]. CA [Ano95c, BBG^{+}95, IEE95a, Kar95]. Cache [PMBH93]. CADNA [JCL10]. CAF [GBR15]. Calculate [AS93, Air04, BS92a, BS92b, BS97, BD90, Cap98, CA90, Dan90, HMT90, Kir02, LHW01, Mai90, Nie92, Sar00, Sar17, SSLG91]. Calculates [Ano90a, AG95b, Cok93b, Hor09]. Calculating [BBCH95, Kod11, MHdL12, Coo94, Cum90, Car93, CB95, Joy92, KRY90, LZL11, Nar95, RPG^{+}20, STY15, STY18]. Calculation [KLA95, KDG99, Som07, AIS^{+}97, BSS92, MSB92, Var97]. Calculations [Kra94, BF92, FTD91, Kle93, RM90, Spe94, SH97, TYJ92]. calculus [RMX05]. Caldera [Ano97b]. calibration [LMK94, Neu01, Ple93]. Calif [ACM01]. California [ACM93a, ACM95b, ACM97, Ano94a, IEE93a]. Call [BDK91, HK92, FT03, YO95]. call-by-reference [FT03]. call-by-value [FT03]. callbacks [BV13]. called [AHZ90]. Calling [Ros93]. Cambridge [AB93, Ano98b]. CAMOS [Osy92]. Can [Mor81, TJ90, DPR94, FTPR04, NLBB23]. Canada [BGG^{+}94, CGS94, GGK^{+}93, Lev95a, BT01, HDR03]. Cancun [Sie94a, Sie94b]. CANPLOT [CM92]. capillary [Ude91]. CAPTools [IJCL96]. Carefully [HH18]. Caribbean [Pel93]. Carlo [BD93, CHM91, Heu90, MMV95]. Carolina [ACM93c]. cartographic [Wol91]. CASCON [BGG^{+}94, GGK^{+}93]. Case [BF01, Buc94c, EKC95, GLPE97, GKKL19, GS01b, McC96, MM98, RRM^{+}15, SN94, AJJF14, Bri00, CHT92, DS97, GMHC92, KNOR04, LSW92, NLVE15, PMHC92]. Castle [AFKL04]. catalyzed [RM90]. caught [Vel97]. caused [HM93]. CC [Fri94]. CD [Ano96b, Ano97a, Ano98a]. CD-ROM [Ano97a, Ano98a]. CDT [CT95]. Cedar [Eig90a, Eig90b, EHJ^{+}91, EHJ^{+}93, GPHL90]. ceiling [Coo94]. ceiling/floor [Coo94]. Celebrating [Lee97]. celestial [GL10]. Cell [ADHF96, CLiN^{+}02, Ves91, KSYE00, Smi93b, XHY^{+}24]. Cell-Structured [Ves91]. cellular [SC19]. Cenju [KKS^{+}95]. Cenju-3 [KKS^{+}95]. Center [ACM98, ACM99, IEE90a, IEE94b, Kar95, Wie94, ABB^{+}94, KT94]. Centers [KT94]. Centre [KSW93]. centrosymmetry [PSPE94]. Century [Bra03, Cip00, NRS92]. CERN [Van95]. Certification [Hop98]. Cetraro [Don95]. CF77 [Cra90, Cra91a, Cra93]. CFD [MFI^{+}94, SM17]. CFF95 [HKŠ^{+}97]. CFP [LSZ92]. CFPs [HHCS95, WHL95]. CG [SZG95]. CG-algorithm [SZG95]. CGA [Ame90a]. Challenge [BEH^{+}94, Sai95]. Challenges [Agr95, Ten93, BCM^{+}93]. Change [WW90]. changes [FR94, HCD^{+}98]. CHAPLIN [BD14]. CHAPLIN-Complex [BD14]. Chapter [AMC01, KK01]. Character [Cou97, Goo90a, Goo90b, IEC94, ISO94, Int00]. characteristics [Dan90]. Characterization [Vaj92, Ber92]. characterized [AF92]. charge [Spe94]. Charles [Eme94, Rag95, UMM94]. Charleston [ACM93c]. CHARMM [HKŠ^{+}97]. Chebyshev [BD91]. CHECK [LCC^{+}03]. checkerboard [BW12]. checking [LCC^{+}03]. Chemical [Cok95, EK01, HKŠ^{+}97, Lar93, SSLG91, WRL90]. Chemistry [GDS94, AFAS99]. Chemkin [Ano97c, Bra97c]. CHEMSHIFT [SSLG91]. Chenies [Eme94]. Chichester [Ano96a, Ano99a, Ano99b, Gon01]. CHICOM [GHN19]. Chief [Ano94c]. China [IEE97]. Chip [Kul95]. Chislennoe [AZ90]. Chivers [GMC96b]. CHIWEI [GH18]. chobo [nY90]. choice [AJJF14]. Cholesky [GRW07, GWDL08, GWDL10, JP95, RS09b, WAG98]. chonja [nY90]. Chonsan [KMmYsK92]. chromosomes [WTW90]. circle [Mil92, Mil92]. citation [CD03]. CL [LW95b, LW95a]. CL-PVM [LW95b]. Clamped [Cap98]. Class [Gre93, NCMF15, OT93, Hor96, PQ94]. Classes [GPS99, Que00, BB07]. Classic [Ano91b, App91, CT11]. clay [MSB92]. Clerman [Mol12]. Client [Ano93n, Ano96b, Sch93a]. Client/Server [Ano93n, Sch93a]. climate [NLVE15]. clinoamphibole [Cum90]. cloning [Das06]. Closed [TR96, MCA17]. closed-shell [MCA17]. Club [VSH91]. Cluster [CLiN^{+}02, BID95, CHM91, MCA17, SR95, Var97, WTW90]. clustered [KSZ90]. Clustering [SSW91]. Clusters [Del98, ADB94, BMV03, BL94, SPS^{+}91]. CM [CC95a, HP95b, KBKT94, MM94, NOL97, SMG91, Sab92, Sab94, Sai95, SOP93, Thi91]. CM-2 [CC95a]. CM-5 [HP95b, KBKT94]. CMFT [PRS99]. CMS [Int90h, Int90n, IBM91a, Int91f]. CNDO [SS09]. CNDO/2 [SS09]. CO [Ano03, NRK98, PTS92, Rei02, RN07, CD03, TOC18, Ano94l, BCS00, BCS01, BC01, CDMC06, MDV07, NR98a, NR98b, Num05, Wal02a]. Co-array [NRK98, Rei02, CD03, TOC18, BCS00, BCS01, BC01, MDV07, NR98a, NR98b, Num05, Wal02a, CDMC06]. Co-arrays [RN07]. Co-Current [PTS92]. Coarray [FGBN19, GBR15, MVK25, SM17, YBMCB14]. Coarrays [NLE^{+}20, RRM^{+}15, RLS20, SC19]. Coarse [BR96, HK91, KY98a, KY98b]. coarse-grain [KY98a, KY98b]. Coarse-Grained [BR96]. COBOL [Ing90a, Sal92, nY90]. Code [AC97a, AC97b, AMC01, BCC^{+}92, Buc94a, Buc94b, CLiN^{+}02, CG96, DCR99b, DLS95, DR93a, DR93b, FES05, GHN19, Gar91a, Gar91b, GM97, KaM10, LS90a, LS90b, LS00, MB95, Mit02, MWO95, MA18, NI03, NLBB23, Pau93, PS96, RD92, RA90, SGMS97, SM03, WW92, DR95b, vWAH^{+}02, AIS^{+}97, ACIK97, ADB94, Azi23, Bee01b, BW12, BGV94, BHS92, CHM91, CCJ93, Cro90, DET12, DGS08, DKM07, DR94b, DR95a, Duf04, Eli98, EH07a, EJLC97, EKC95, FTPR04, FBC96, GH18, GV92, Gao05, Gao06, GMF18, Gom90a, Gom90b, GS98, GAW96a, GAW96b, HIK90, Hop98, How91, KSYE00, KLM00, Kin92, KKY99, KDG99, LMJC96, LP90, Mar92, MDM05, NVC96, Ogi02, OT93, Paj90, PBU95, PG10, RBS93a, RBS93b, Rib02, SH91, Sil93, SPF00, SF10, Str05, SMSW06, SH97, TYJ92, Van22, WSW00]. code [XHY^{+}24, YK90, YB92, ZZSW19, ZT90, ZBW07, vK94, vHK00]. Codes [Adv98, ADHF96, BCC^{+}91a, BCC^{+}91b, DL97c, PAK^{+}90, SWH15, UNF^{+}08, WMMW97, AH90, dSZP92, BF92, BC97, BSCV95, Cah90, HWS09, IJCL96, Kir93, Kir98, MVK25, RBS92, Sai95, Sta94, SSS99, UZCZ96, YB13]. coding [BBB^{+}57, FKL94, FTPR04]. codon [Wri90b]. Coefficient [BH92, WS94]. Coefficients [Rhe93, Sil01, Azi23, Err06, NG93, Nie92]. coexisting [FT03]. Coffee [Cof93]. cognitive [Cho91]. Cohen [RR92]. Coherence [PMBH93]. cohort [ZMR^{+}91]. cokriging [GMHC92, PMHC92]. Collected [Ham85, HM90a, RH84, SS90]. Collection [BE92, BCE93, Mil04b, WNO94, DH84, Esp98, PBU95]. College [WR93]. Collier [DDH17]. collision [HMT90]. Collocation [BB91, MKFB92, WKM04]. Colorado [Ano94l, Sch93b]. Coloring [BCT94]. Columbus [Hua96]. Column [GP97, GP94]. COMAGMAT [AFBN93]. Combinatorial [Sug95]. combined [Mir90, Var97, MIN^{+}95]. combines [Ano93d]. Combining [DP94]. Coming [Coc03, HK95, HCD^{+}98]. Command [Mac90, Phi91b, Phi92]. Comment [GG99, McC95, Tsa01, Gho01]. Comments [BDH90]. Committee [W^{+}95, ANS95]. Committees [Ano94n, Ano94o]. Common [Rei96a, FBWR95, KA95]. Commun [Jam96]. Communication [BR98, BD96, CL97, CGL^{+}95b, CHKM93, KW94, Mac91b, Mac91c, MR96b, PSC93b, SOG94, TRV96, VRT97, BBDR94, BBDR95, CGL^{+}93, GKH^{+}92, KHJS94]. Communication-buffers [MR96b]. communication-efficient [KHJS94]. communication/computation [BBDR94, BBDR95]. Communications [Ano96b, Ano94d, BKT91, Cha94a, Coe94a, EJLC97]. Compact [MBGK11, PW84]. companion [Ing90a, Ing90b]. Compaq [Law01]. Comparative [GKKL19, LCD91]. compared [Gro91]. Comparing [Ano99c, BF01, FBC96, GGW96, GBR15, Ram90, GHN19]. Comparison [BGLP94, CSC^{+}97, CWB94, SG93b, SG93c, HKŠ^{+}97, Mös95, RS93, Wal02a, BDOS95a, BDOS95b, BID95, CF90, CWB92, GS95, HKM98, Las97, McB91, MMG98, NJ94a, SM02a, TOC18]. compartment [Coo94]. Compatibility [SM02a, BS13]. Compatible [BL90]. Compcon [IEE93a]. Competitive [Hil91]. Compilation [Adv98, BCFH93, BGS94b, CFH^{+}93, Coe94a, Coe94b, CA96, Eps94a, Eps94b, Fah94, FXAC94, HHKT92, HHKT96, Nik93, O’B93, PSC93b, TBC94a, UZCZ97, BCF^{+}94c, CGS94, Eps96, Hal91, IEC99, Int99, KY98a, KY98b, MCH96, PSC^{+}95]. Compile [ASS95, DCZ96, PH96, SPM^{+}94]. Compile-Time [ASS95, DCZ96, PH96, SPM^{+}94]. Compile-Time/Run-Time [DCZ96]. compiled [GMF18]. Compiler [ASS93, AS97, BGS94a, BBZ94, BL90, BCF^{+}93c, BMN^{+}97, BD96, Bra00, BHMS91a, CT11, DG99, DDHW96a, FW90, Fri94, GMS^{+}95, HKT91b, HKT91a, HKT91c, HKT92a, HKT92b, HKT93b, HKT94, IHKvW02, Jéz93, Ken94a, LFK^{+}93, McJ17b, NR06, Pad00, RVV^{+}92, STVS91, SZAB97, SZAB98, SZAB99, SIDH95, SAC^{+}92, TBC94b, TBG^{+}02, Tse97, WRL90, YYM93, ZCP95, vKK92, von92, vKK^{+}93, AFMP95, ABC^{+}96, Ano92b, Ano98c, AHJS90, ADG96, BCM^{+}93, Bee01a, Bee01c, BCF^{+}93a, BCF^{+}93b, BSCV95, BHMS91b, BGS82, CMT01, CD03, CSS90a, CSS90b, CSS91, DS97, Eig90a, Eig90b, HDH^{+}94, HKT91d, HKT93a, Int90a, Int90b, Int90c, Int90d, KKS^{+}95, LM94, Lov94, MCAB^{+}02, NR05, NVG94, Par86, PGH^{+}90, Sof93, Sab92, Sab94, SNK06, Spo94, Tou84, Tse93, VKB93, Ben99c]. compiler/6000 [Int90a, Int90b, Int90c, Int90d]. Compilers [Ano93m, Ano02, BB96, BCFH93, CT11, IK96, KLW93, LZ97, LHH^{+}91, Mar90, McC95, Nak95c, Pre93c, PA94, SF02, Sch93b, SS96, TT93, Ano93j, Ban93, BGNP94, BCF^{+}94c, CCKT86, CTS96, CC92b, Cre90a, DPR94, DFRR91, GB92, HDH^{+}95, Hua96, Int92, JH86, KW94, LCD91, LYZ90, LP92, Met99c, Met99d, Nak95b, Nic91, PBG^{+}95, Pon94a, Pon94b, SM02a, Sal92, SM92, Sar97, SNMC93, SLY90a, SLY90b, WFW^{+}94]. Compiling [AKLS88, BZ99, BCF^{+}93a, BCF^{+}94b, BCF^{+}94d, BMMN94, BMN^{+}95, Cho92, DT93, DD97, HBB^{+}95, HKT92c, JM94, KHS96, OE92, RMCKB97, SAC^{+}92, TIUG90, YYM93, Cra90, Cra91a, DDS99, HMS^{+}95, NNON02, WMCU97]. Complementary [Cod90b]. complet [Del93]. Complete [A^{+}92, ABW92, ABM^{+}97, Ano98b, Bee02]. Complex [Amo90, FGG09, GPS99, HFT94, HFT97, Kod08, Kod11, NPB92, Sht19, Smi98, AC16, DDH17, EC13, FGGL05, GST02a, GST02b, MSA03, BD14]. complex-step [MSA03]. complexity [BKT91]. Complying [Bak95]. Component [Bel11]. components [Av94]. Composition [HLJ98, MSB92, Nie92]. Compositional [KR94, KR95]. compression [SJ94]. Comput [Jam96]. Computation [ACM94c, Adl93, AMGM20, BF93a, BD96, Cre90b, Dre92, Dre93, IEE94a, JSW93, JB01a, Lev95a, SS94, SS95, Sch96a, Sen03, SB01, Sug95, Sun92b, TR96, Ueb97, Var95, WN90, YKK96, Bin96, BBDR94, BBDR95, BG94, CZ10, CN94, Eme94, GST02a, GST02b, Gro90, GDS94, HKM98, LP90, Mor81, PT93, SS90, Sre92, Vig93, Vol93, ABB^{+}94, KT94].

Computational [BFHH94, BLW02, Com91, DeV94, EKC95, Hun00, KM90, NBC92, Ric06, FGBN19, HF95, HP95b, Mal91, PSG03, VLLY92, WTW90, ZJEP95a, ZJEP95b, PRS99, Ano03]. Computations [Ano94p, Bra00, FB12, Fuj95, MFI^{+}94, MR95b, PCS98, ZMR^{+}91, CC94, GLS93, KNOR04, KO94, KB94, MR96b, Nak90, PDS^{+}93, PCS99, UZCZ95]. Compute [ABB^{+}91, JP95, McB06, RH94, Shi93b, BG93, Con92, EC13, FR94, KK90, Lar93, Mac96a, Sat97, SSG^{+}10, SSG^{+}18, ZZSW19]. Computer [Ame97a, Ano95b, AH92, Bon06, Cok93a, Cse99, EPL94b, FL91, IEC90, IEE94g, ISO90, IJCL96, JL93, Knu03, KZ94a, KZ94b, Lap96, MT90, Mra94, Nis95, Osy92, Rit90, Sab92, SNJ^{+}92, TIUG90, Ten93, vDSP96, AKLS88, Blu91, Car93, FCHE02, GL10, GR92, HCD^{+}98, HT91, Jon92a, Jon92b, Kea92, KSM95, LD87, Mat90, Mir90, SS93, Tou84, TJ90, Unixx, vV90, Bar92]. Computer-Aided [IEE94g, HT91]. Computers [BCF^{+}93c, BCF^{+}94d, Che92, Dec93, Don91, DV92, FYR99, FGG09, Hud91a, Hud91c, KRB^{+}90, MSC96, ONT95, PAK^{+}90, Schxx, SS96, Tho97b, TT93, YRF02, ALS91, All93, BCF^{+}93b, BCF^{+}94b, Don90, Duv92, FGGL05, Ger98a, Ger98b, Hew91a, Hew91b, Hud91b, KKS^{+}95, LP93, McB91, PW84, Sab94, SSW91, Swa84, Wie94, Wol92]. Computing [ACM97, ACM98, Ano93a, Ano93m, Ano97d, AH92, BGS94a, BBG^{+}95, BH92, BEH^{+}94, Bra97d, BKR^{+}91, CJL97, Cam13, CC95b, Cos97a, Cse99, DGR92, Dow93, For97, FJSD96, Fur93, GS01a, GST04b, GST06b, Glo91b, HH18, HR92, Hun00, IEE94d, IFI95, KNS95b, Kon00, Kon94, LP98b, LMR^{+}97, Mac91a, NR06, Ort94b, Ort94a, OPB^{+}20, PTM96, PTV96, Ric95, Sab95, Ste95a, Ten93, Tho97a, Van95, Vel97, VBA95, Wil93, Yan94b, ZA11, Zag16, AK93, AHZ90, AB93, Ano93q, Ano94d, Ano98a, Ban93, BGNP94, Bec91, BPG94, BDG^{+}94, BBB00, BK06, Bou95, BDH^{+}05, Cel96, Che90, CDF^{+}93, CNP91, Cyb91, DGR90, DT94, DW94, Don95, Dut94, Err06, EFP07, FPR01, GH94a, GH94b, GH94c, GST12, HH14, Has06, HL08, HS95, Hua96, IEE97, KSZ90, KT94, KNS95a, Kir98]. computing [LP05, Loz98, Mer92a, MMG98, MMG00, MM02, NDSG07, PG10, PBG^{+}95, PTVF92, Pre94b, Raj95, RBS93a, RBS93b, Sch93a, SMB90, TMD13, Wal93a, Wal93b, Zim07, Gon01, Lev98, Ano99a, Ano99b]. concave [Dem06]. concentrations [RKMJ92]. Concept [KaM10]. Concepts [Ano93h, DNS97, Fos95, MRG^{+}93, NDSG07]. conceptual [IJCL96]. concerning [MKS^{+}96]. Concerns [Off98]. concise [Yip90]. Concludes [Coc03]. Concurrent [BGMZ92, Bre92]. condensate [BKRG22, KRG21]. condensates [KLM^{+}19, RPG^{+}20]. Conditional [Air04, Eps94a, Eps94b, Eps96, IEC99, Int99]. conditions [EH07a]. conducting [PSPE94, WCN92]. conduction [iSYS12]. conductive [Car93]. Conference [HOP93, ACM93c, ACM93b, ACM94a, ACM94b, ACM95a, ACM95b, ACM96a, ACM96b, ACM97, ACM98, ACM01, Ano94a, Ano94l, BBG^{+}95, Boi97, BT01, BV94, CGS94, DSZ94, Ein91, ERS95, FH90, Fri94, GH94a, GH94b, Glo91b, HMPT94, HAM95b, HS95, HS94b, HS94a, IEE92b, IEE94d, IEE94e, IEE95b, IEE02, Kar95, KRB^{+}90, KSW93, MS94, NBC92, PRS99, RFC90, Vol93, Ano93q, BLT94, CKMU94, DR94a, GH94c]. Confidence [SB01, WS94]. confined [PS08]. Confirmatory [KKH10]. Confluent [NPB92]. Conformal [SS99, MKF95]. Congress [HR92, KSW93]. CONHYP [NPB92]. conical [GST12]. conjugate [LN91, MN01]. conjugate-gradient [LN91]. conjugated [KS12, SS10]. Connecticut [Ban93]. Connection [AKLS88, BL91, BHMS91a, BHMS91b, CC95a, DFL92, Sab92, Sab94]. connectivity [RTY90]. CONPAR [BV94]. Conquer [ARS92, ARS94]. consensus [TRS91]. conservation [Ste90]. considerations [KM99, LHHJ91]. constant [CCKT86, MS93a, SRH96]. Constants [GG99]. Constrained [FJS97, Kea95b, MHdL12, ZBLN97, CZ90, GOT03a, MN11, Ren96b, Ren96a, BMR01]. Constraints [FJ92, MP93, ZT90, ZBW07]. Construct [DP94, IFI93, Pug90, XH90, MC96, Tip91]. Constructing [Ano93b]. construction [Fri94, KLM00]. constructs [ABC^{+}96]. contained [AI90]. containing [BSCV95]. Contemporary [For97]. Content [Teo01, Coc03]. continue [Cof93]. continued [McG91]. continuum [PG10]. contouring [Gou93]. Contrasting [HFMS95]. Contribution [BBCR98]. Control [CFGG94, Enr95, FGCG94, FJ92, IEE94g, Kra94, AS92, BMO90, Bar92, CZ90, EH07b, EP92, RBD^{+}10, RBD^{+}11, RKMJ92]. control- [BMO90]. Controlled [NJ94c, Lie94a, Lie94b]. Convention [ACM98, ACM99, IEE94b, Kar95, FKL94]. Convergent [WMMW97]. Conversion [Buc94c, KP92, Manxx, BF92, Blu91, Che91, Nai17, SMH91]. Conversione [Anoxx]. Converter [FGMS90c, FGMS90d, FGMS90a, FGMS93, FGMS95, FGMS90b, FLQZ97]. Converting [AS91, FT03, Gli96, McD93]. Convex [Som98, Dem97, Dem06, BMR01]. Convex-Constrained [BMR01]. convex/concave [Dem06]. convexity [Ren04]. convexity-preserving [Ren04]. Convolution [BHMS91a, BHMS91b]. Cool [Sla96]. Cooperating [CF95]. Coordinate [BDOS95a, BDOS95b, OP98b, DRST03, DH95, SZ90]. coordinate-time [SZ90]. coordinates [NY91]. Coordination [OP98a]. Coprocessor [Kul95]. Copy [GS97]. CORBA [Nat00, SFKL02].

CORBA-based [Nat00, SFKL02]. Core [TBC94b, Av94, RS09b, TBC94a]. cores [SC19]. Corner [Bli94]. cornerturn [Hol90]. Corporation [AOL94a]. Corpus [CGL^{+}95a]. Corpus-based [CGL^{+}95a]. correction [Agt94, Lop90]. corrections [AIS^{+}97]. Correctness [CRDO16, Fos17]. correlated [PZY16]. Correlation [BH92, SD90, WS94, Ame90a]. Correlations [HFMS95, SB91, SFB92, SWM95, vV90]. Correspondence [Bru96a]. Corrigenda [Ano94b, FYR99]. Corrigendum [ARS94, Dre93, HC94, MS00b, RBD^{+}11, YRF02]. CORSURF [KSM95]. Cosine [BH92, SM95, Mac96b]. Cosine-Pi [BH92]. Cosmological [MB95]. Costs [GS97]. Coulomb [Sar00, Sar17]. Counter [PTS92]. Counter-Current [PTS92]. Counting [WS94]. County [ACM98, Bra94a]. Coupled [LJO05, BKRG22, KRG21, MCA17, Sil93]. coupled-cluster [MCA17]. Coupling [LJO05]. couplings [AC16]. cours [Ain90, Ain91]. Course [DeV94, Kam00, LE98, Manxx, Ric06, Smi94, Smi95b, LP05, LD87, ZJEP95a, ZJEP95b]. courses [Met99c, Met99d]. covariance [KSM95]. covariance-matrix [KSM95]. CPAR [CC94]. CPAR-language [CC94]. CRAFT [PMM94, SZG95]. Craftworks [Ano97b]. CRAY [Car91b, EO91, Nag90, PSG03, PAK^{+}90, SZG95, VSH91, WW92, Car92, HP95b, KLN90, KZ94a, KZ94b, Mac91b, Mac91c, MWO95, Oed93, Vaj92]. CRAY-2 [Car91b, PAK^{+}90, Car92]. CRAY-YMP [Car91b, HP95b]. CRC [Tay99]. Create [AG95a]. Creating [AC97a, AC97b, SWW90, Law01]. Creek [Neu01]. Criterion [KH13, van90b]. Critical [McB06, SB01, Som07, Las97]. CRop [van90a]. cropping [CNP91]. cross [Ame90a, Hor09]. cross-correlation [Ame90a]. CROSSOVER [McG91]. CRPC [Fox91a]. CRPC-TR92225 [Fox91a]. Crystal [Ger94a, KLA95, Nar95]. crystallographic [Cum90]. CSE [MM98]. CSMP [Lef93, Sto93, van90a]. CSMP-Spring [Sto93]. CSRFPACK [Ren04]. CSTRAIN [GF95a]. CTRAN [Che90]. Cubature [BE92, BCE93, Esp98, GC03]. cubic [Aki96, DV00]. CUDA [AC17, GML^{+}16, HE13, MVK25, NLVE15, iSYS12, XHY^{+}24]. Current [IFI95, PTS92]. Curses [Ano93b]. curve [MC96, Ren09]. curves [BG93, Dut94]. Custom [EFG^{+}05]. customization [IBM91a, IBM91b]. Cut [PR91]. CUTEr [GOT03a]. CUTOFF [MC92]. Cuts [Sht19]. CVFF [HKŠ^{+}97]. CXFTV2 [Dem97]. Cycle [KK95b]. Cycles [CL93, CL94]. cyclic [HC08, VRT97, WO96]. Cydra [DT93]. cylinder [GST06a, GST06b, GST11, YB92]. cylinders [YK90]. Cylindrical [Kod08, Kod11]. Cyprus [PRS99].

D [Ano90b, HKK^{+}91a, KR95, CMV09, Car93, FHK^{+}90b, FHK^{+}90a, Fox91b, Fuj95, Hal91, HHKT92, HHKT96, vHKS94a, HKS94, HKT91b, HKT91c, HKT91d, Hir91, HKK^{+}91b, HKT92c, HKT92a, HKT92b, HKK^{+}92, HKT93a, HKT93b, HKTW94, HKT94, Ken94b, Kon92, KR94, RBS93b, SC19, Tse93, Tse97, WKM04, Wie94, vKK92, von92, vKK^{+}93, vKS94, vHKS94b, RBS93a]. d1mach [GG95]. DaCapo [BGH^{+}06]. DAEs [CM94]. Dagstuhl [AFKL04]. Dallas [Ano94i, IEE93c]. damping [CA90]. dans [Cha94a, Gom90b]. DAP [SHCP91]. DAP510 [CWB94]. d’application [II90]. d’applications [Pic94]. DAPRE [SP91a, SP91b]. d’Arc [BLT94]. DaReL [KN95]. Data [ACG^{+}94, AMC98, ALS91, AZ98, Ano96b, Ben95, Bra00, CFK^{+}94, CMMZ93, CZM93a, CZM94b, CGS93, CGL^{+}95b, CHL94, Dem95, FC09, Fox94, FMW^{+}94, Guo01, GS97, HCLJ03, KP92, KY98a, Kea95a, Kea96a, Kea96b, KK95a, KNS95b, KK98, KP93, LR94, LH92, Mas93b, Meh93a, Meh93b, Nan93c, O’B93, PSC93b, PHD^{+}95, RSB97, Ren97b, RB99, SWW90, SSC00, SR04, Ste95b, TZW^{+}10, TR96, UZCZ96, Wal90, Wal92, Wal00, YKK96, ZCMM93, AW94, Aki96, AI90, Av94, AFMP95, ABC^{+}96, AGG^{+}97, BK89, BMO90, BG93, BKK94, BDOS95a, BDOS95b, Blu91, BID95, BxCW01, CM92, Cha94b, CMZ94b, CMVZ94, Cha93, CGL^{+}93, CS90b, CM91, CP94, DV00, Dem03, Dem06, Dem07, EKB92, Gep90, GB92, GKH^{+}92, GKH^{+}93, GHSJ94, GS95, HW95, HBD^{+}93, HC08, IEC98b, Int98b, KN95]. data [KY98b, KHJS94, KNS95a, KGV97, LYZ90, Lin90, MKS^{+}96, Mar93, Mas94, McG91, MBFC99, MC96, MR96b, NJ94a, Off94, OPP00, PPW94, PBU95, PW93, Per94, PD96, Phi91b, Phi92, PSC^{+}95, RBS92, Ren96b, Ren04, Ren09, SS90, SNK06, SZ90, SKM94, SSOG93, SV95, TBC94a, UZCZ95, WO96, WCN92, YO95, ZMR^{+}91, ZZN94, GG95, BCC^{+}97b]. data- [BMO90]. data-domain [RBS92]. Data-Flow [Mas93b, Mas94]. Data-Localization [YKK96, KY98a, KY98b]. Data-Parallel [ACG^{+}94, AMC98, CZM94b, CGS93, CGL^{+}95b, Guo01, GS97, KNS95b, PHD^{+}95, SSC00, Ste95b, UZCZ96, AFMP95, BDOS95a, BDOS95b, Cha93, CGL^{+}93, KNS95a, MR96b, UZCZ95]. data-parallelism [PPW94]. Data-structure [BCC^{+}97b]. Database [OC94, Bet97, Che91]. Dataflow [YYM93, YKK96, Cas14, SRH96, WMCU97]. date [Bee96c, Bee97, Din99]. dating [Xu93]. David [Ano96a, Eme94, Hin06, Iha06, Rag95, UMM94, Sch07]. Dawn [Ano03]. DC [IEE94f]. DCE [Sch93a, LK93b, RS93, Sch93a]. DCL [VKB93]. DCR [Azi23]. DCUTRI [BE92, Esp98]. DDE [TS06a]. DDFUN90 [Bai05a]. DDT [AGG^{+}97]. de-allocation [RMX05]. Debate [BDH90, Can92a, Can92b, Can91]. Debugger [But95, CH94, IGHG^{+}94, FSPC^{+}02]. debugging [BHS92, HKMC90, SSG94, SSG97]. Dec [Bjø08, Ano91d, Ano91c, Dig92, Dig93a, Dig93b, Dig93c, KLS94b, Lan93a, Lov94]. decay [Hor23]. decays [DET12, MDM05]. December [HHK94, IEE92c, IEE93c, Ing90b, Kar95, Kum94, Ing90a]. Decision [CFGG94, FGCG94, VBA95, DI90]. Deck [BP92, Mil91]. Decks [NOL97]. Decomposition [DDF10, GLPE97, RG90a, SWW90, NVG94, RG90b]. dedicated [GL10]. Deep [OPB^{+}20, AIS^{+}97, CNP91, Coc03]. defect [EH07b]. Defined [CMZ93b]. Defining [CM91]. definite [Duf04]. Definition [NSJD98]. deflection [HM93]. Deformation [FYR99, YRF02]. degree [ADD04]. Delaunay [CCW04, Ren96a, Ren97a]. Delft [DSZ94]. Delinearization [Mas92b]. demand [BMO90]. demand-driven [BMO90]. Demonstration [GB92, GMHC92, PMHC92]. d’emploi [Ber91b]. Denelcor [DH84]. Denmark [DW94]. Denotational [Guo01]. dense [RPL96]. Department [Bee01g, Bee01f, Bee01e]. Departments [Tom99]. departure [Dut94]. depend [Cof93]. Dependence [HHLS90, KK95b, Mas93b, MP93, OE92, SAS90, SF93, AZ98, BMO90, LYZ90, Mas92b, Mas94, MHT96]. Dependence-Analysis [HHLS90]. dependent [AFAS99, KYSV^{+}15, MA09, YSVM^{+}16, YSMA^{+}17, YSMBA23]. derivations [HW95]. Derivative [BCC^{+}91a, BCC^{+}91b, BCC^{+}92, GST11, MSA03]. Derivatives [AMGM20, Kub91a, Kub91b, Kub91c, BLL^{+}96, McG91, Met99c, Met99d]. Derived [PMM^{+}08, RMX05]. describe [GBC92]. Describing [Boo81]. Description [BDK91, IEE92a, AAK01, DFRR91]. descriptions [MKF95]. Design [ACM93b, BLLWW95, BDPW98, BCF^{+}93b, BCF^{+}93c, CTS96, Che95, Coe96, Cok95, EP92, GR92, HMR^{+}15, Her90, IEE94g, LM90b, MMT09, Mit02, PHHF94a, RAX10, SOG94, WBS97, Wri91, AM90, Bar92, Boo81, CKT85, DG08, Ell81, GT92a, GT94, ISKvW02, KM99, Ker90, LM90a, LFG00, QRH00, Ren09, Wri90a, ZE92, GT92b]. designed [DLW^{+}18, Str05]. Designing [Du 97, Fos95]. Designs [AC97a, AC97b, Air04, Cok93a]. Desk [Bra97a, Fri96]. Desktop [Ano97b, Tho97b]. Details [Cou97]. Detecting [Nan93b, RH94]. Detection [BEH^{+}94, HK91, McB06, van90b, CFMR95, HKMC90]. Determination [Gil91b, Gil91a, Gil01, LP90, RKMJ92]. determine [McG91]. Determining [HMW91, HMW93, WS94, Deu90]. Deterministic [CF95, CFMR95]. Developer [Lew94, Loh07, Sco93, Sun93]. Developers [HDR03]. Developing [Gen06, LS05, LM94, Nat00, CDF^{+}93, Sch07, Hin06, Iha06]. Development [AC17, Ana93a, Ana93b, Ano93i, Ano97b, BL90, Dan90, DG94, KG99, KKMP95b, MFI^{+}94, NJ94c, PHHF94a, Pel93, SFKL02, Tre97, XH90, BGH^{+}06, Che91, CKT85, DSZ94, FG93, KKMP95a, dLJEB95, Mic93b, MMRS92, RL91, Sal06, Wie94]. Developments [Cse99]. device [CM92]. DFN [RS93]. DFN-RPC [RS93]. dHPF [MCAB^{+}02]. DI-3000 [Blu91]. Diagnostic [HHLS90]. Diagonal [MKFB92, vH06, vH07]. Diagram [Ren97a, Tip91]. Diagrams [NCMF15, NL19]. Dialects [GPHL90, PCS98, CWB92]. Dialogsystems [Kru90a]. diatomic [PZY16]. Diego [ACM93a, Kar95]. Dies [Loh07, Mar07]. diesel [KRY90]. DIFALPHA [Sil01]. Difference [CC95a, Fuj95, Sil01, GV92, HE13, LD90]. Differences [Dem95, SB91, SWM95]. Different [EL97, Sil01]. Differential [BG97, Cas89a, CC92a, DPS02, EL97, Hig91, MD97, Nak95a, Nan93b, RH94, Shi93b, AZ90, BG94, GST04b, HBG^{+}05, HIS91, KM99, LS04, Sil93]. differential-algebraic [KM99]. differential/algebraic [HBG^{+}05]. Differentiation [BKMC96, BCH^{+}06, DLS95, FHP^{+}12, Gar91a, Gar91b, Hor92, JSY^{+}20, KN94, LS90a, LS90b, LS00, Maaxx, SP91a, SP91b, UNF^{+}08, AFBN93, CDGM96, GJU96, NR05, SPF00, SF10, Str05, YB13, vH06, vH07, vH10]. differentiation-enabled [NR05]. differentsial′ [AZ90]. Diffraction [BRdAHK04, MDD94]. diffusion [Tal94]. Digest [IEE93a]. Digit [Ves91, Kah01]. Digital [Ano91e, JSW93, Ple93, Tre97, AOL94a, BLLWW95, ED99]. digitised [SHCP91]. dike [CNP91]. dilute [TS06b]. Dimensional [BCE93, BM99, CLiN^{+}02, DM90, BSCV95, CRS90, CHM91, CA90, Eli98, Gao06, GF95a, Gou93, GMHC92, Heu90, KS12, Ogi02, PMHC92, PT93, Ren96a, SMSY02, SRM90, SWO92, VLLY92]. Dimensioned [Ros93]. Dimtest [SNJ^{+}92]. dip [McG91]. dipolar [KYSV^{+}15, YSMBA23]. dipping [FYR99, YRF02]. Dirac [Mac98]. Direct [DR93a, DR93b, Ham93, MMV95, DR95b, DR94b, DR95a, HWS09, KMJ^{+}23, SWH15]. directional [Lai92a, Lai92b]. Directions [Bod94, IFI95, Sav95]. directives [BCF^{+}93a, BC19]. directly [BG93]. Directory [PMBH93]. Directory-Based [PMBH93]. Discovering [CT90]. discovery [BD93]. Discrete [Ano90a, FJ92, SM95, Dem97, Kir93, Kir98, Nan93a, Ren03, Tor10]. Discrete-Time [FJ92]. discretization [Gao05]. discretization-theory [Gao05]. Discretized [BB91]. Discussing [Coe96]. discussion [BBF^{+}92]. disordered [LZL11]. Dispatcher [Mac90]. dispersions [ZZSW19]. displacement [FR94]. displacements [Ude91]. display [PBU95]. DISPMODULE [Jon09]. dissociation [TYJ92]. distance [MNZ90, ZBW07]. Distributed [AW94, BR96, BCF^{+}93c, BCF^{+}94d, BMMN94, BMN^{+}97, CL97, CMZ91, CZM93a, CH94, DCZ96, Ger94a, HM96, HHKT92, HBB^{+}95, HL08, HKT92b, HLJ01, IEE92c, IEE93c, KHS96, KMR96, KK98, LK93a, McD93, Mer92b, Nat00, O’B93, RSB97, RA90, Sch93b, vDSP96, BZ99, BCF^{+}93b, BCF^{+}94b, CN94, Cho92, CEF^{+}95, CK91, DSv94, DR94a, GHSJ94, Hal91, HM93, HKT91b, HKT91a, HKT91d, HKT92c, HMS^{+}95, IEE97, KN95, KMR^{+}97, KHS95, KGV97, PZA93, RBS92, SSH08, SM92, SNK06, Sch93a, TBC94a, Tse93, Wag94, WW95, WI94, YO95, Yu01, ZA93, vPMF92]. Distributed-Memory [Ger94a, HKT92b, KHS96, KMR96, KK98, HBB^{+}95, RA90, BZ99, Cho92, GHSJ94, Hal91, HKT91b, HKT91a, HKT91d, HKT92c, HMS^{+}95, KN95, KHS95, RBS92, SSH08, SNK06, Tse93]. Distribution [Ano90a, CGSS94, Gil94, Ham98, McB06, ACIK97, AGG^{+}97, BCF^{+}93a, CZM93b, CMZ93a, CP94, Gho01, LPA95, MZM94, Tsa01]. Distributions [CMMZ93, vHKS94a, HKS94, PHD^{+}95, ZCMM93, BSCV95, Cha94b, GKH^{+}92, GKH^{+}93, LHW01, Pon94a, Pon94b, PSC^{+}95, VRT97, WO96, vKS94, vHKS94b]. Divide [ARS92, ARS94]. Divided [Dem95]. divider [Kah01]. Division [FKL94, WBS97]. DL POLY [KSYE00]. dlja [ES93b]. DM [Ano97a]. DNA [HKŠ^{+}97]. DNAD [YB13]. DNSPLIN1 [Ren03]. Do [YWS^{+}94]. documentation [Kes92]. Doing [Koo90, Pif96]. Domain [DDF10, DRST03, GLPE97, Gao05, HE13, RBS92, Hew90a]. Domain-Decomposition [GLPE97]. Domains [CMV09]. Dominance [FC09]. Dominant [BS92a, BS92b, BS97]. dotCall64 [GMF18]. Double [FKKC96, LH92, Bai05a, Bai05b, Bai05c]. double-double [Bai05a, Bai05c]. Double-Precision [LH92]. double-single [Bai05b]. DQAINF [EO94]. Draft [Ame87, Ame90b, Fox91a, W^{+}95, ANS95, ISO00, ISO04a]. Drafted [Coc03]. dragon [Sal95]. Drexel [Sen03]. drive [Tea94]. Driven [BFKS93b, CMKH03, NJ94c, BMO90, BFKS93a]. Driver [Ano96b]. Drivers [Cod93a, Cod93b]. drop [Cok91]. DSblock [GOBG^{+}94]. DSblock-Model [GOBG^{+}94]. DSFUN90 [Bai05b]. Dt [Pas95]. Dt-FORTRAN [Pas95]. DTM [DH95]. dual [YB13]. Dublin [HR92]. due [How91, SH91]. d’un [Robxx, RD91]. DuPage [Bra94a]. Dusty [BP92, NOL97, Mil91]. dusty-deck [Mil91]. DVM [KKMP95a, KKMP95b]. Dybbuk [PSC^{+}93a, PDS^{+}93]. dyke [MM02]. dyke-like [MM02]. Dynamic [AAN^{+}93, AMKS02, CMMZ93, Cha94b, DS01, RMX05, SM90, SR04, Tal91, Teo01, Vio90, ZCMM93, AFAS99, CK86, CZ90, Kin92, KB94, Luc92]. dynamical [KLN90, Sat97]. Dynamics [BFHH94, BL91, DCR99a, DCR99b, EKC95, Hun00, Nis95, WBS97, BCS01, Cre90b, EFG^{+}05, HF95, KSYE00, NSWP90, QRH00, SZ90].

E4 [MGH81]. Earth [Fos93, FYR99, Par94, YRF02, FR94, Ogi02, SMSY02]. earthquake [Gep90]. Easy [Del98]. ECMWF [HK93a, HK93b, HK95]. ecological [Lef93]. ECPSSR [Hor09]. ed [AB93, KF92a, Rub93]. Ed. [Yan94b]. Eddie [DeT90]. Edge [Hil91, Agt94]. Edinburgh [Fri94]. Edited [Mal91]. Edition [Bro92a, GWE^{+}05, Wu93, Tha93, Loz98]. Editor [HKTW94, Ano94c, Bra97a, Fri96, HHK^{+}93, Hat94, KMT91]. Editor-in-Chief [Ano94c]. Editorial [Bra96]. Editors [Cip00]. EDSS [VBA95]. Education [Ano95b, KT94]. Effect [Cho91, Gil91b, Gil91a]. Effective [BS91a, BS91b, BC94, BC19, Wri90b]. Effectively [LR94]. Effectiveness [DP96, Fu90]. Effects [WR93, Agt94, CHT92, HK90, How91, SH91, Xu93, Yan95]. Efficiency [Zag16, XHY^{+}24]. Efficient [BB02, CCW04, DS01, EB98, GRSS02, HK92, JB01a, KHS96, KK94, Kru90b, LYZ90, SW94, SSC00, TR96, CCJ93, CFPS94, DLLR96, FTPR04, GMF18, GS98, KHJS94, LMJC96, Mas92b, Ove91, Paj90, PPW94]. Effort [Fah94]. EFGs [Spe94]. Eigenproblem [ARS92, ARS94]. Eigenproblems [BSV16, LS09]. EIGENTEST [LS09]. Eigenvalue [MR95a, BHLT09, LP90]. Eigenvalues [DGR92, LH92, MR93b, Shi93b, BBB00, BK06, DGR90]. Eigenvectors [DGR92, DGR90]. eighteenth [ACM91]. Eighth [ERS95, Sie94a, Sie94b]. Eingabe [Por90]. Einstein [BKRG22, KRG21, KLM^{+}19]. electric [HT91, Spe94]. Electrical [Oku95]. Electromagnetic [CLiN^{+}02]. Electromagnetics [Car90, PSG03]. electromechanical [GBC92]. Electron [BRdAHK04, LZL11]. electronic [KLM00, SMB90, SS10]. electronic-structure [KLM00]. Electronics [IEE92b]. electrons [Hor09]. ELEFUNT [Cod90a]. elegant [Sch91a]. Element [Ano94p, BGLP94, DFS95, Fen96, KG99, Nak95a, SM02b, Dot93, KBKT94, Nie92, OA02, Ste90, SSS99, TOC18, ZZN94]. element-by-element [OA02]. Elementary [HK93c, HFT94, KDKSH92, KDDH94]. elements [Dot93, Sar00, Sar17, Var97]. Elimination [BKT91, BC94, FTPR04, NH09, PW84]. ellipFor [TS25]. Elliptic [BD91, Nak95a, TS25, Car91a, STVS91].

Elliptic-Parabolic [BD91]. Elution [PTS92]. Embedded [KA95, RFC90, FH90]. emerging [CD03]. Empirical [SLY90b, HKŠ^{+}97, MS93a, SLY90a, VSH91]. Emulating [Mor15, Pug90]. Enable [IFI93]. enabled [NR05]. Enabling [SM94]. Enclosure [Kea95b, AH92]. encounter [Met92a]. end [Lov94]. Energies [Ano90a]. Energy [BPG94, BRdAHK04, BG93, BSS92, Ste90]. Engeln [Ano97a, Hop97, TDMC97]. Engeln-Müllges [TDMC97]. Engineering [Alt90, BPG94, BCM99, Bro90a, Cha94c, GS90b, Glo91b, KaM10, KM97, KF92c, Kri86, Lew94, LMR^{+}97, Pao99, Pao01, SS94, SS95, BN93, BN96, Ben99b, CC98, CKMU94, Eme94, EKC95, FG93, Fos95, GS90a, HGG93, HT91, IEE94c, KRB^{+}90, KF93b, Mat90, Pri93, SFKL02, ZJEP95a, ZJEP95b, KF92a, Tay99]. Engineers [BS91a, Bor91b, Bro95, Bro97, Cha97a, Edg92, Ett90, Ett92, Ett93, Ett96, Ett97, Fen96, For97, Hah94, HB91a, HB91b, Mar98, NL92, NL95a, NLN96, NL96, NL97a, RZ94b, Rub93, Smi94, Smi95b, Tor91, BS91b, Bor91a, Cha95b, CC95b, EK01, GR92, NL95b, NL97b, Ano90b]. England [Eme94]. Enhanced [HCLJ03, IEC98b, SZAB98, SZAB99, And02, Int98b]. enhancement [Boi97, HT91, LHHJ91]. enhances [Cok93a]. enhancing [BK89]. ensemble [Sha94]. Enterprise [SSW91]. entitled [Wie94]. entre [II90]. Entry [MD97, ZJEP95a, ZJEP95b]. entry-level [ZJEP95a, ZJEP95b]. Entwicklung [Ano93p, UM93]. Entwurf [Kru90a]. Enumeration [DS01, SVD96, SDv98]. Environment [Ana93a, Ana93b, BCC^{+}96b, DL97c, Don91, DV92, Ger94b, HRW^{+}98, JBBH93, Kea95b, SS00, VBA95, All93, BDG^{+}94, BCC^{+}97a, Cen91, CB94, CN94, CEF^{+}95, CKT85, Don90, GOT03a, HZ94, JA92, KHC92, Lav91, MMRS92, Pri93, Sch93a, Shi98, Sre92, Vee94, BLLWW95, GGHvdG01]. Environments [ACG^{+}94, DT94, Ein91, Ric95, Sch93b, TAH^{+}01, AES^{+}96, Che90, CWB94, CDH^{+}94, DR94a, IEE90b, SM92]. EPPP [NVG94]. EPROP [SMB90]. Equality [MHdL12, ZT90]. Equation [AJ98, BG97, Fuj95, GWL^{+}92, LZ97, Shi93b, Azi23, CRS90, CK86, CA90, GST04b, HM12, HBG^{+}05, HIS91, Hop02, KKK95, KYSV^{+}15, LS04, MA09, iSYS12, TS06b, YSVM^{+}16, YSMA^{+}17, YSMBA23]. Equations [BB91, Cas89a, CC92a, dCH94, Don91, DV92, DR93a, DR93b, EL97, Hig91, Kea95b, MD97, Nak95a, Nis95, AF92, AZ90, BKRG22, BG94, BDH^{+}05, CS14, Cro90, DH84, Don90, DR94b, KRG21, KM99, KKY99, KL92, Lie94a, Lie94b, Mas92b, Nat92, NY91, PS08, RRX^{+}08, Sil93, Ste90, SMSW06, WSW00, ZZN94]. EQUEL [Ing90a]. Equipment [AOL94a]. equipped [Hor09]. ERCS08 [Hor09]. Erratum [Jam96, KR95]. Error [Cod90b, Enr95, Kub91a, Kub91b, Kub91c, Nak95a, Bli90, JCL10]. Errors [Bel11, BDH90, Wal90, CBTL97]. Erstellen [Kru90a]. Erwin [NRS92]. ES/ [SPS^{+}91]. ES/3090 [WTW90]. ESPRIT [CDH^{+}94, Hey94]. ESQL [Ing90b]. Essays [NRS92]. ESSENTIALS [Smo94]. Establishing [MC92]. Estimate [WS94, dSZP92, Koi09, NVFNP93]. Estimates [Kub91a, Kub91b, Kub91c, How91, SH91]. Estimating [Rit90, Coh90, JCL10, Kay90]. Estimation [BGW93, Gar91a, Gar91b, EK01]. Estimator [Hig90a]. Etude [Robxx]. Euclidean [HH18]. EURO [HAM95b]. EURO-PAR [HAM95b]. Europe [Ano93n, Ano93q, HMPT94]. EUROSIM [DSZ94]. EVAL [KS90]. evaluate [BBZ95]. evaluates [Cok91]. Evaluating [BBDR94, BBDR95, Fu90, HKT94, Rhe93, Tho90, IMS90a]. Evaluation [PEP92, AAN^{+}93, AHOK02, BGZ94, BF93a, BFHH94, BB96, Bra94c, Bra94d, Cod90b, CL93, DJ92, Gao05, Han98, HKT91d, HKT92b, Hun00, KS90, LZ97, LHH^{+}91, NIY^{+}94, SF02, Schxx, SM02b, Som98, SOG94, YFH97, BF93b, Bli90, BR98, Bou95, Din99, Gao06, KLV98, KKS^{+}95, KHRS95, Kir93, KP93, MZM94, MAH^{+}02, NSU20, SM02a, SS93]. evaluation-interpolation [KP93]. Evaluator [NPB92]. Event [HMW91, IGHG^{+}94, HMW93, Nan93a]. Event-Based [IGHG^{+}94]. events [Hor23]. Ever’ [Cam13]. Everything [Cre03]. evolution [FHE95, Van84]. evolve [OJ09]. Evolving [ORO15]. Exact [CT95, Gil94, CZ10]. EXAFS [AI90]. examination [SMH91]. Example [Kon94, V^{+}93, Ano92e, DFRR91, VTP92, Vet93]. Examples [MFK09, Lef93]. Exception [HFT94, HFT97, IFI93, Rei95a, Rei95b, IEC98a, Int98a, ISO00, Rei95c, Rei97]. Exchange [PTS92, SNJ^{+}92]. Exchangers [PTS92]. excitable [FCHE02]. excitations [Taq16]. excluded [BDH^{+}05]. executables [Ano93j]. Executing [BMMN94]. Execution [Bai93a, KMS^{+}95, Mit97, RHH96, SWBO93, Sze90, Ano93d, KHS95, Rot93, SSW91, SZ91, VKB93]. Executive [Jus92]. exemple [LMG95]. exercices [Ain90, Ain91]. Exhibition [GH94a, GH94b, HS95, KSW93, Ano93q, GH94c]. Existing [SWW90, SF93]. exits [TLS90]. Expansion [Vio90, Err06]. expect [Ano93j]. Experience [HK90, Hig90a, Ola92, OE92, Sou91a, Sou91b, BC19]. Experiences [AS91, CNBB96, CDMC06, Ein96, HHK^{+}93, HHLS90, HKT93b, SM17, SAC^{+}92, HKT93a, Sai95]. experimental [FBC96, Hen94, SS93]. Experiments [CJL97, Coe94b, SZG95]. Expert [BPG94, HR92, AS92, LMK94]. EXPFIT4 [IDVV97]. Explained [Ano90c, Ano93b, Gla92a, MR87, MR90b, MR91, MR96a, MR99, MRC04, MRCB23, Rub93, MR90a, MR92, MR93a, MR94, MRC11, MRC18, GMC96d, Ano91a]. Explicit [CHKM93, NY91]. explicitly [PZY16]. exploitation [JA92]. Exploiting [HF95, Hig90b, JB01b, LE98, LR94, Nag90, OP98a, RSB97, SSOG93]. Exponential [Amo90, Ham98, Smi11, CRS90, PZY16]. exponents [Sat97]. Exposing [PMM^{+}08]. Exposition [KRB^{+}90]. Express [Bee01a, DNS97]. Expressing [MMV95, PCS98, CWB92]. expression [NH09, PCS99]. expressions [BBDR94, BBDR95, Mar93, MBGK11]. EXSHALL [NY91]. Extended [Ame90b, Ame92, DCHH88b, DCHH88a, GWE^{+}05, NL19, SKP91, AC92, DDH17, SJ94, Sch93b]. Extending [CMZ94b, CZM94b, Che95, DY99, MR95b, HM92]. Extensible [GWE^{+}05]. Extension [CMZ91, KA95, NSJD98, PH06, AAK01]. Extensions [AHOK02, DGL91c, DGL91a, UZCZ97, BCC^{+}97b, CC94, CA92, ISKvW02, SIOS02, Wea94]. extent [McG91]. External [BBB^{+}94, MFK09, Av94]. Extracting [Bar94, NCMF15]. extraction [GP92, OT93]. Extrapolation [EL97, Gro90]. Extreme [McB06]. Extrinsic [Hig94c].

F [Hop97, Sch91b, Ano96c, BEH97, Bra97b, BLL^{+}96, Geh97, HHCS95, Mei96, Rei96b, UNF^{+}08]. f2c [AS91, Lev95b, Lev97]. f2cl [BW96]. F90 [Das06, DP94, Gla92b, Ola93, Ola92]. f90ppr [Ola95]. F World [Sla96]. fa [yKxx]. facilitate [LMJC96]. Facilities [Moo95b, IEC98b, Int98b, Moo95a]. FACt [RPG^{+}20]. factor [Car91a, Cok93b]. Factorization [GWDL08, GRW07, GWDL10, vPMF92]. Factorizations [JP95]. Factors [MHdL12]. Faddeyeva [Zag16, ZA11]. fall [KKZ11]. Family [Hew01, GR92]. FAN’s [Ola93]. Farewell [Bli94]. Farms [AOL94a, AOL94b, KC94]. Fast [DH92, FBWR95, Hig90b, Lev92, ONT95, Ple93, RBS92, RHH96, SM95, Ves91, GHSJ94, Heu90, HS10, Kir02, LHW01, MH91, Rei96a]. Faster [VJ97a, VJ97b]. Father [Mar07]. fault [FYR99, Gep90, YRF02]. fault-STRGRV [YRF02]. fault/slickenside [Gep90]. Favor [Pif96]. FCRC [ACM96a]. FDCHQHP [WW14]. FDR [Som07]. Fe [USE94]. FE2DY [Ste90]. Feasibility [KR94, KR95, GT03, GT07]. feasible [Sav95, SH97, ZT90]. Features [ABMS94, Ano93c, Ano93k, SG93a, SG93d, Ham95a, KZ94a, KZ94b, AKLS88, LS00, UZCZ95, UZCZ96]. February [Ano93q, BBG^{+}95, IEE93a, IEE94a]. Feedback [MR95a, FPR01, Has06]. Feel [NLBB23, DFS95]. FEM [OA02]. Fermi [Mac98]. fermion [KKY99, WHL95]. fermions [KLN90, PS08]. fever [Cre90a]. FFT [Swa84]. FFTS [Car92, Car91b]. FIDE [LD90]. fidl [LK93b]. Field [Ano94i, Av94, BD93, Spe94]. fields [LN91]. Fifth [Ano95c, HK93b, IEE93c, Ano93q, HK93a]. File [BGMZ92, Met99b, Bre92, CFPS94, Met99c, Met99d, Phi91b, Phi92, iSYS12]. Files [KP92]. FILT [SRM90]. FILT-PC [SRM90]. filter [GT03, GT07, Tor10]. filter-trust-region [GT03, GT07]. filters [PT93]. FILTRANE [GT03, GT07]. Financial [Jou95]. Finding [DS02, TRS91]. Fine [CL97]. Fine-Grain [CL97]. Finite [Ano94p, BGLP94, CC95a, DFS95, Fen96, Nak95a, SM02b, Dot93, HE13, KBKT94, PSG03, Ste90, TOC18, ZZN94, LD90]. finite-difference [HE13]. finite-element [ZZN94]. finite-volume [PSG03]. fire [Coo94]. First [App91, Cas89a, CC92a, DeV94, DW94, HAM95b, Hig91, Kam00, Kum94, ONT95, SWM95, Smi94, Smi95b, TIUG90, LP05, LZL11, Met92a, Ove91, Shi98]. First-Order [Cas89a, CC92a, Hig91, ONT95, SWM95, TIUG90, Ove91]. first-principles [LZL11]. first-time [Shi98]. fit [GH18]. FITEVT [Hor23]. Fitting [Dem95, Aki96, CRS90, DV00, Gho01, Ren96b, Ren09, Tsa01]. Fitzroy [Eme94]. five [Leo91]. FL [ACM91]. FLAME [GGHvdG01]. FLAN [OC94]. Flannery [Adl93, AB93, Yan94b, Gar93, Loz98]. Flattening [GF95b]. Flavor [DGS08]. Flexible [FGG09, HC08, Lie94a, Lie94b]. flexural [HM93]. flight [How91, SH91, WBS97]. FLIPP [Kry94]. Floating [FBWR95, IEC98a, ISO00, Rei96a, Smi91, TOML04, Ume91, VCV97b, CA90, Smi01, VCV97a, Wic89, Int98a]. Floating-Point [Rei96a, Smi91, TOML04, VCV97b, IEC98a, ISO00, Ume91, Smi01, VCV97a, Wic89, Int98a]. Flood [Bra94a, RTY90]. floor [Coo94]. flop [MMG00]. Florida [ACM98]. Flow [LR94, Mas93b, Coo94, Dan90, Kle93, Mas94, TOC18]. flowfields [OM92]. flows [KBKT94]. fltgrh [FYR99]. fltgrv [FYR99]. fluctuations [RPG^{+}20]. Fluid [BFHH94, EKC95, Hun00, Nis95, SMSY02, San92, HF95, HIS91]. fluids [Cok91, Lar93]. fly [HKMC90]. FMSI [Gep90]. focal [Gep90]. Fock [HKŠ^{+}97, KS12, PS08, SS09]. fonctionnel [RD91]. fondamentaux [MRG^{+}93]. FORALL [ALS91]. Force [Jor90a, Jor90b]. Forecast [GK06, RHH96, Kor99]. forecasting [GS95]. foreland [HM93, HM93]. Forest [Gar91a, Gar91b]. Forever [Ano91e, Wal91a]. Form [Cap98, DGR92, Bar94, DGR90, HD05, Lie94a, Lie94b]. FORMAC [Gri93]. Formal [CS90b, Cho91, MKS^{+}96, SKM94, GGHvdG01]. Format [GOBG^{+}94, GWDL08, Pug90, GRW07, GWDL10, Bee90]. formation [HM93]. formations [Car93]. Formats [Bon06]. Forms [TR96]. Formula [AJJF14, Lor19, Nob90]. Formulation [WAG98]. Fortan [Loz98]. FORTDIFF [KN94]. ForTec [Bee02]. Forth [RFC90, FH90, Nob90, Rod90]. Fortieth [Lee97]. Fortnet [AH90, AHZ90, CA92]. FortPort [MMRS92]. FORTRAN [Adl93, Ano03, BH92, Dig90a, Ede90, EMR93, Gar93, Gen06, Gho01, IEE92a, Mic93b, Mic93a, Pas95, Rub93, Sch91b, Sun92a, Tay99, Tsa01, VTP92, Yan94b, YRF02, Ame87, AB93, Ano93b, Ano94o, Ano96b, Ano97b, Ano97c, Ano98a, Ano99a, Ano99b, AZ90, Bee01a, Bee01c, Bee02, Ben99c, Bra97c, Can92a, Cod90a, CGT92a, DG99, Dub97, ES93b, Eme94, Fah02, Fah94, GBR15, GMC96b, GMC96a, GMC96c, GMC96f, GMC96e, Ger94b, Gon01, Hew01, Hig92, HH18, Hin06, HKK^{+}91a, ISO04a, ISO04b, ISO10, Iha06, Jam96, KLM91, Kea96a, Kea96b, KK01, KF93a, KF97, Kon94, KR95, KHS17, Lev98, Loz98, MB95, Mai91, Mar98, McC95, Mol12, MN11, Rag95, Rei92a, Rys95, Sch07, Sch97, TDMC97, Tha93, UMM94, Wu93, dL12, AC92, Ame96, Ame97a, Abs91, AC97a, AC97b]. FORTRAN [Agt94, AIS^{+}97, AI90, Alg90, AH90, AHZ90, All93, All90, Av94, Ame90a, Amo90, Ano90a, Ano90c, Ano91b, Ano91f, Ano92d, Ano92e, Ano93j, Anoxx, App91, Are90, AFBN93, Ash81, AAK01, BBB^{+}57, BS92a, BS92b, BS97, BGKZ91, Bai93a, BKRG22, BD90, BG93, Bec91, BSS92, BL90, BRdAHK04, BD93, BRH90, BCM99, Bor91a, Bra90, Bra94a, BA95, BGMZ92, Bre92, BDH90, BH90, Con91, Cah90, CV94, CM92, Can91, CD92, CK90, Car90, Cas89a, CC92a, CMP02, Cha95b, Che90, Che91, CHM91, Che95, CC98, CFGG94, CNP91, Cod93a, Cod93b, Cof93, Coh90, CS90b, CJPA94, CA90, Con92, CHL94, Con90, Coo94, CS90c, CSS90a, CSS90b, CS91, CSS91, Cra95, Cre90a, Cro90, Cro91, Cum90, Dig90b, Car93, CB95, DeT90, Deu90, DGL91c, DGL91a, DGR92, DDS99]. FORTRAN [Dot93, DH95, DM90, DI90, Dut94, ES93b, Ell81, EMR93, EP87, EKC95, FJS97, FL91, FTD91, FGCG94, FR94, FYR99, FPR01, FC95, FHE95, FBC96, GGLM88, GL90, Mer91, GWL^{+}92, GS90a, Gep90, GF95a, Gil91b, Gil91a, Gil94, Gil01, Gom90a, Gom90b, GS98, GT92a, GT94, Goo90a, Goo90c, Goo90e, Goo90f, Goo90d, Goo90b, GMMM92, Gou93, Gro91, GMHC92, Hew90a, Hew90b, Hew91a, Hew91b, Hew92a, Hew92b, HW95, HHCS95, HC92, HC94, HGG93, Has06, HT91, Hig91, HW91, HM93, HP95a, HB91b, Hor09, Hor23, How91, HK93c, Hud91b, Hun00, Int90a, Int90b, Int90c, Int90d, Int90f, Int90g, Int90h, Int90l, Int90i, Int90j, Int90k, Int90m, Int90n, Int90e, Int91a, Int91b, IBM91a, IBM91b, IBM91e, IBM91c, IBM91d, Int91e, Int91c, Int91d, Int91f, Int92, Intxx, IEC90, IEE90b, Ins92, IEE93b]. FORTRAN [Lib90a, IMS90a, IMS90b, Lib90b, IMS91c, IMS91b, IMS91f, IMS91d, IMS91e, IMS91g, IMS91h, ISO90, ISO94, Ing90a, Ing90b, IDVV97, JC93, Jor90a, Jor90b, Joy92, KP92, KRG21, KDKSH92, Kes92, KSYE00, Kin92, Kir02, KS90, KKK95, KF90, KF92a, KF92d, KRY90, KVK92, KSM95, Kub91a, Kub91b, Kub91c, Kug92, KL92, yKxx, KDG99, Lan90e, LE98, LK93a, Lan93a, Lan01, Lar93, LD87, LM90a, Lef93, LMJC96, LN91, Lev95b, Lev97, LS90a, LS90b, dLJEB95, LSZ92, LH92, Lop90, Lou90, LHW01, LP90, Maaxx, Mac90, MB92, Mai90, MKFB92, Mal91, MCA17, Mar90, MJR93, Mas92a, MC91, McG91, McJ17a, MC92, Meh94, Mer92a, MSB92, Mey00, Mic91, Mil92, Mil04b, MR93b, MGH81, Mor81, Mös95, Num91b, Numxx, Nagxx, Nan93c, Nan93b, NY91, NK94, NJ94a, NJ94b, Neu01, NVC96, Nie92]. FORTRAN [NVFNP93, NLVE15, NL95b, Osy92, PMHC92, Pao99, Pao01, PT93, Par86, Par94, PBU95, PW93, Pau93, Pel93, Pon94a, Pon94b, PTVF92, Pre92a, Pre92b, Pre92c, Pre92d, Pre93e, Pre93f, Pre94b, Pug94, Raj95, Ram90, RS92b, Rap90, RBS92, RBS93a, RBS93b, RKMJ92, RR92, RG90a, RG90b, RVV^{+}92, Rhe93, Rib92, Rit90, Robxx, RH94, RPG^{+}20, RA90, Sci93, SPS^{+}91, Sal92, SH91, San92, Sar97, SS90, SSW91, SS94, Sv90, Sch90, SGMS97, SMB90, SSLG91, SD90, SB91, SFB92, Sil01, SW91, Smi91, Smi93a, Smi94, Smo94, SB01, SRM90, Spe94, Spe93, Spexx, SWO92, SP91a, SP91b, Ste90, Ste91, SHCP91, Sto93, Str05, Uni2 , Sun92b, Sun94, Sze90, SZ91, Tal91, Tho90, TYJ92, Tip91, TT92, TOC18, Tur93, Uni93, Ude91, Van84, VKB93, Vio90, WMCU97]. FORTRAN [Wal90, Wal91b, Wal92, Wal93b, Wam90a, Wam90b, WS94, WD98, WW92, WJ94, Wei91b, Wei93, WCN92, Wol91, WR93, Wri90a, XWK95, Xu93, Yan95, YYM93, YB92, Zah92, Zei92, ZE92, ZMR^{+}91, ZZN94, Zim07, ZB94a, GT92b, vV90, van90a, van90b, von92, Bee96b, Bee96c, Bee96a, Bee97, Ame90b, A^{+}92, Ame97b, Ame92, AL92, ABW92, ABMS94, ABM^{+}97, Adv98, AMC01, Ain90, Ain91, Ain93, Air04, ADHF96, Aki99, AS97, AKLS88, AG95a, AFAS99, Alo04, Alt90, AC17, dSZP92, AR06, Ana93a, Ana93b, ACIK97, And90, And92b, And02, AGS92, AS91, AH91, AH94, AOL94a, AOL94b, Ano91d, Ano91c, Ano91e, Ano92a, Ano92b, Ano93c, Ano93d, Ano93e, Ano93g, Ano93h, Ano93f, Ano93k, Ano93o, Ano93p, Ano94e, Ano94f, Ano94h, Ano94m, Ano94o, Ano95d, Ano95e, Ano95g, Ano97d]. Fortran [Ano98c, Ano99c, Ano02, Ano07, AJJF14, AHJS90, AO90a, AO90b, AO90c, ADG96, AS14, AC16, AGG^{+}97, Azi23, BCM^{+}93, Bac98, BGZ94, Bai92, Bai93b, Bai94, Bai95, Bai05a, Bai05b, Bak91, Bak95, BCS00, BCS01, BC01, BS13, BS91a, BT94, Bar94, BBCH95, BP92, Bee90, Bee91, Bee01b, Bee01d, Bee01g, Bee01f, Bee01e, Bel90a, Bel11, Bel90b, BBZ94, Ben95, BZ99, Ben99b, BB02, BBB00, BK06, BSV16, BW12, Ber91a, Ber91b, BGNP93, BCC^{+}91a, BCC^{+}91b, BCC^{+}92, BKMC96, BDC^{+}96, BF93a, BF93b, Bla00, Bli94, Bli90, Blu91, BKP93, BFHH94, BL91, Bor91b, BGV94, Bos19, BDGxx, Bou96, BCFH93, BCF^{+}93a, BCF^{+}93b, BCF^{+}93c, BCF^{+}94c, BCF^{+}94b, BCF^{+}94d, BCF^{+}94a, BMN^{+}95, BMN^{+}97, Bra91, BGA90, Bra94b, BGA94, BGA96, Bra97b, Bra03, Bra97d, Bra94c, Bra94d, BCC^{+}96a, BCC^{+}96b, BCC^{+}97a, BCC^{+}97b]. Fortran [BG94, BG97, Bre78, Bre79, BGS94b, BV13, BHMS91a, BHMS91b, Bro90a, Bro92b, Bro92a, Bro95, Bro03, Bro97, BW96, Bro90b, Bru96a, Bru96b, Buc94a, Buc94b, Buc94c, BC19, BD14, BDK91, BSPF01, BSB^{+}03, BDH^{+}05, Com91, CLiN^{+}02, CF90, Can92b, CG96, CSC^{+}97, CRS90, CZ10, Cel96, Cha94a, Cha95a, CTS96, CCL01, CCL04, CMZ91, CMZ92a, CMZ92b, CMMZ93, CZM93a, CZM93b, CMZ93a, CMZ93b, Cha94b, Cha94c, CZM94a, CMZ94a, CMZ95, Cha97a, Cha97b, CC95b, CC92b, CCW04, Che92, CZ90, CC90, CS90a, CS95, CS00, Cho91, Cho92, CKZ93, CFH^{+}93, CWB92, CWB94, CD03, CDMC06, Coc03, Cok95, CM91, CGT92b, CRDO16, CA92, CT11, CC93, Cou91, Cou97, CT90, Cra91b, Cra92, Cra93, Cre03, CI98, Cro07, dCH94, Cur94, DDF10, DLM99b, Dig92, Dig93a, Dig93b, Dig93c, Dan90]. Fortran [DDcMR96, DET12, DLLR96, DP96, DP99, DL97a, DL97b, DS97, Dec93, DNS98, DNG07, DG08, DGS08, DL97c, Del93, DZ98, Dem95, Dem97, Dem03, Dem06, Dem07, DDH17, DG99, DCR99a, Rocxx, DOP^{+}92, Din99, DKM07, DLS95, DGR90, Don90, Don91, DV92, DDH^{+}95, DDH^{+}96, DDHW96b, DCHH88b, DCHH88a, DD97, DS94, Cro92, Du 97, DV91, DV93, DB93a, DB93b, DY99, DR93a, DR93b, DR94b, DR95a, DV98, DVY00, DV01, DV02a, DV02b, Edg92, Eig90a, Eig90b, EHJ^{+}91, EHJ^{+}93, ES93a, Ein94, Ein95, ECS96, Ein96, EKB92, Ins91a, Eli98, Ell90, EPL94a, EPL94b, EPL95, EH07a, EMU96, Err06, EC13, EFP07, Ett90, Ett92, Ett93, Ett96, Ett97, ED99, EB98, FBZ92, FSPC^{+}02, FGBN19, FT03, FGMS90c, FGMS90d, FGMS90a, FGMS90b, FW90, FGMS93, FGMS95]. Fortran [FC09, FCHE02, FB12, FG93, FKL94, For97, Hig93b, FC92, Fos93, FXAC94, Fos17, FHK^{+}90b, FHK^{+}90a, Fox91a, Fox91b, FLQZ97, FES05, FGJB19, Fu90, Fur93, Glo91a, GH18, GHN19, Gal91, Gao05, Gao06, Gar91a, Gar91b, GMC96d, GS90b, GG99, Geh95, Geh96, GMF18, GB95, GP94, GP97, GKKL19, GK06, GST02a, GST02b, GST12, Gla92a, GBDB97, Gli96, God93, GML^{+}16, GU90, GT03, GOT03b, GT07, GRE99, Gre93, GOS94, GOBG^{+}94, GWE^{+}05, GRW07, GPHL90, Hig94a, Hig94b, Hig94c, Hig94d, HM96, HE13, Hah94, Hal91, HHKT92, HHKT96, Ham95a, Han98, Han92, HL94, HBG01, HBG02, HH14, vHKS94a, HKS94, HBB^{+}95, Hat94, HMR^{+}15, HF95, HM12, HZ99, Hen95, Her90, Heu90, HKT91b, HKT91c, HKT91d, Hir91, HKK^{+}91b, HKT92c, HKT92a, HKT92b, HKK^{+}92, HKT93a, HKT93b, HKT94]. Fortran [HFMS95, HIK90, HB91a, HK91, HMKN91, Hop98, Hop02, Hop03, Hor91a, Hor91b, Hor92, HJT97, HJJ^{+}00, Hud91c, Hud96, Hug96, HMT90, HLJ95, HLJ98, IBM93, IEC94, IEC97, IEC98a, IEC98b, IEC99, Ins91b, IFI93, II91, Int97a, Int97b, Int98a, Int98b, Int99, ISO00, Int00, Irv91, ISKvW02, Jam94, JCL10, JPE20, Jon93, Jon09, JSY^{+}20, JP95, Jus92, Ken91, Ken92a, Kah01, KMR^{+}97, KTP^{+}24, KaM10, KMJ^{+}23, KY98a, KY98b, KTMB02, Kay90, Kea92, KDDH94, KN94, Kea95b, Kea95a, KAČ^{+}22, Kef92, Ken94b, KK95a, KMR96, KS02, KKZ11, KT00, Ker90, Ker91a, Ker91b, Ker93a, Ker93c, Ker93b, KMBK96, KLM00, KMS^{+}95, KO91, Kir93, Kir98, KG99, KKZG94, KKZG95, KC94, KOM93, KOM94, KLW93, Koe92, KLS^{+}94a, KF92b, KF92c, KF93b, KF96, KKY99, Kon92, KS12, KKMP95a, KKMP95b, KLA95, KM90].

Fortran [KGV97, KK94, KK90, Kra94, KR94, Kro14, Kru90b, Kry94, KZ94a, KLS94b, KKH10, KH13, KP91, KYSV^{+}15, KLM^{+}19, KZ94b, Lan90b, Lan90c, Lan90d, Lah90, Lai92a, Lai92b, LK93b, LP05, Lan93b, Lan90a, LMG95, Las97, Law01, LM90b, Lee90, Lee97, LS04, Lem93a, Lem93b, Lem93d, Lem93c, LS05, Leo91, LW89, Lew94, LZL11, LS00, Lig91a, Lig91b, Lig93, LP98a, LP99, Lin93, Lin90, Loh07, LMMW96, Lor19, Lov92, Lov93, Lov94, LHH^{+}91, LCC^{+}03, LMK94, Manxx, Mac91b, Mac91c, MH91, MD97, MKS^{+}96, MSP^{+}22, MWM90, Mar92, Mar07, MDD94, Mas93a, MHT96, MC94, MC95a, MC95b, McC96, McD93, McJ17b, MVK25, MM94, MMY95a, Meh93a, Meh93b, MVZ98b, MVZ98a, MZ00, MZ01, MMEH08, Mei95, Mer92b, MH95, MCH96, MR87, MR90b, MR90a, MR91, MR92, Met92a, MRG^{+}93, MR93a]. Fortran [MR94, Met95, MR96a, MR99, Met99a, Met99c, Met99d, Met99b, MRC04, MRC11, MRC18, MRCB23, MBGK11, MMV95, MMY95b, Mil91, Mil04a, MMRS92, MR95a, Mir90, Mit02, Mit93, MDV07, MA90, MN01, MMG98, MM98, MSZ90, MS93b, Mor15, MWO95, MR95b, MDM05, MHdL12, MM02, MA09, Num90a, Num90b, Num90c, Num91a, Num91c, Num93a, Num93b, Nat00, Num92, Nag95, Nag01, Nag02, Nai17, NLE^{+}20, NCMF15, NL19, Nar95, NSU20, NR05, NR06, NS11, Ngu91, NI03, NSWP90, NLBB23, NDS96, NDSG07, NOL97, NR98a, NSJD98, NR98b, NRK98, Num05, NL92, NL95a, NLN96, NL96, NL97a, NL97b, O’K93, OPE^{+}95, Off98, Ola93, Ola95, Ola96, OC94, ORO15, Ort94b, Ort94a, OPB^{+}20, OE92, PZY16, Pad00, Pag95, PFS^{+}04, Paj90, PS08, PHHF94a, PHHF94b, PH96, PPR97]. Fortran [PSPE94, PH06, PMM93, PMM94, Paz96, PG10, Per93, Phi91b, Phi92, Pic94, Pif96, PMM^{+}08, Poh97, PGH^{+}90, Pra90, Pre93a, Pre93c, PA94, PS96, Pre93d, PTM96, PTV96, Pre93g, Pre99, PCS98, PCS99, Pug90, QRH00, RRM^{+}15, RP12, RM90, RTY90, RS92a, Rat95, RZ94a, RZ94b, Red95, RL91, Rei93, Rei92c, Rei92a, Rei92b, Rei95a, Rei95b, Rei97, Rei02, Rei03, Rei04, RN07, RS09a, RLS20, RPL96, RFS98, RR93, RR99, Rib02, Ric06, RP93, RD91, Rot93, RMCKB97, Rou90, RMX05, RRX^{+}08, RAX10, RMX12, Rub93, Rys95, Sil92a, Sil92b, Spe96a, Sof93, Scixxa, Scixxb, SMG91, Sab92, SWBO93, Sab94, SS09, Sai95, SFKL02, Sar00, Sar17, SZM98, iSYS12, Sat97, Sav95, SWW90, SOP93, SS95, SZ90, SSG^{+}10, SSG^{+}18, SM90, Sch93c, Sch99, Sch03, Sch96a, SZAB97, SZAB98]. Fortran [SZAB99, SNMC93, SKM94, See04, SKP91, SIOS02, SD01, SD03, STY15, STY18, Sha94, SM17, SLY90a, SLY90b, SH07, SH09, She91, Shi98, SM03, SC19, Sht19, Sil93, SD99, SWM95, SB92, Smi92, Smi93b, Smi95b, Smi00, Smi01, Sny07, Som98, SS10, Sou91a, Sou91b, Spe96b, SPF00, SF10, SD92, SD93, Ste93, SF93, SAC^{+}92, SSS99, SH97, Sun05, SSG97, Szy07, Thi91, Taq16, Tay97, Tee90, Tem96, Tho93, Tho97a, Tho97b, TS06a, TBG^{+}02, TMD13, TS06b, TT93, Tom99, Tor91, Tou84, Tre97, Tre91, Tre95, TS25, Tro90, Tse93, Tse97, Unixx, U.S01a, U.S01b, U.S01c, UM93, UNF^{+}08, UHP91, Utt90, Vai93, Van22, Var97, Vel97, VJ97a, VJ97b, VCV97a, VCV97b, Vet93, WRL90, Wag94, W^{+}95, Wal02a, Wal91a, Wal00, Wal01, Wal02b]. Fortran [Wal93a, WW14, WHL95, WAG98, WNO94, WMMW97, Wea94, WHL92a, WHL92b, Wei91c, Wei91a, Wes96, Wie99, Wil93, Wil95a, Wil95b, Wri91, ANS95, XHY^{+}24, YGS^{+}94, YBMCB14, YFH97, nY90, Yip90, YKK96, YSVM^{+}16, YSMA^{+}17, YSMBA23, YK90, Yu01, YB13, Zag16, ZZSW19, ZT90, ZBLN97, ZBC^{+}92, Zim92, ZCMM93, ZBC94, Zim02, ZB94b, Zos93, dSL98, DR95b, van94a, vWAH^{+}02, vKK92, vKK^{+}93, vKS94, vHKS94b, vH06, vH07, vH10, Olv91, Ano96a, Ano97a, EMUP98, Hop97, Kri86, Ano90b, Ano91a, Ano98b]. Fortran-77 [ES93b, CM92, GWL^{+}92, GH18, GHN19, Hop02, SF93, KDDH94]. FORTRAN-90 [BRdAHK04, ES93b, Bai94, Bai95, Bai05a, Bai05b, CC92b, DS94]. FORTRAN-based [Che91, Hun00, DDH17, NOL97]. Fortran-like [Wal93a, KGV97]. FORTRAN-Linda [Sci93]. Fortran-P [DOP^{+}92, O’K93, OPE^{+}95]. Fortran-S [BKP93]. FORTRAN-Scientific [Kri86]. Fortran-Style [SKP91]. FORTRAN-The [Yan94b]. Fortran-to-C [FGMS90c, FGMS90d, FGMS90a, FGMS93, FGMS95]. Fortran-to-Fortran [KN94]. Fortran-to-Java [FLQZ97]. FORTRAN-XSC [Wal93b]. FORTRAN/ [Hew90b, Hew91a, Hew91b, Hew92a, Hew92b, Wri90a]. Fortran/HPF [UZCZ97]. Fortran/PVM [MWO95]. Fortran2003 [DLW^{+}18]. FORTRAN77 [But95, BS91b, Rotxx]. Fortran90 [DNS97, DPS02, DDHW96a, LJO05, RY99]. Fortran95 [DN04]. FortranD [HKT91a]. FORTRANe [Mal91, AZ90]. Fortranning [Mol12]. Fortrans [Ham93]. FORTRESS [BKRG22, KRG21]. FORTREX [Jus92]. Forum [Bee02, Fox91a, DHP02, Bee02]. ForUML [NCMF15, NL19]. foundations [PD96]. Four [DM90, KNOR04, Lai92a, Lai92b, MBGK11]. Four-Dimensional [DM90]. four-directional [Lai92a, Lai92b]. four-particle [MBGK11]. Fourier [AI90, DLM99b, DLM99a, GHSJ94, Kam00, Kir02, MH91, Mat90, SRM90]. Fourth [PPP93, ACM93a, Cas89b, GM97, IEE92c]. Fourth-Order [GM97]. FPS [SAC^{+}92, Tou84]. FPS-164 [Tou84]. FR14C [Num90c]. fractional [Deu90]. Framework [FHS78, Fox79, GOS94, KK95b, MMT09, MA18, RSB97, vDSP96, ACIK97, CH98, MDV07, Sal06, vK94, vHK00]. France [BLT94, Van95, IEE94c]. Francisco [ACM95b, BBG^{+}95, IEE93a]. Frank [Ano97a, TDMC97]. FRB [CZ10]. Free [Ano97d, BL93a, Bra97d, BL93b, BSS92, CZ10, RRX^{+}08]. freezer [Cra95]. Frequently [Ola93]. Friend [BDH90]. friendly [CFPS94, TS06a]. Fritz [Coc03]. front [Lov94]. Frontiers [IEE94a]. Frontiers’95 [IEE94a]. fronts [EN96]. FSQP [ZT90]. Fthreads [Nag01]. fuel [KRY90]. Fujitsu [AHOK02, Bee01a]. Full [GWDL08, Rei02, GWDL10]. fully [MA09]. Function [BBCH95, Cod93a, Cod93b, HLJ98, KDKSH92, KDDH94, NPB92, Olv91, Wal01, Wal02b, Das06, GST11, GST12, Jon92a, Jon92b, MZM94, NVFNP93, PRS99]. Function-Composition [HLJ98]. Functional [SWW90, SKP91, GP92, MC96, RD91]. functionality [Gro91]. Functioning [Nan93b, RH94]. Functions [AAC^{+}04, Are90, CCL01, Cod90b, Fab04, HK93c, HFT94, HFT97, Kod08, Kod11, Maaxx, Mac98, MKFB92, MJR93, McB06, Nag95, Ros93, SGMS97, Sht19, Smi98, Smi11, Tho97a, Tho13, TS25, Zag16, Ano98a, BBZ95, BB07, CDGM96, EFG^{+}05, Err06, EC13, GST02a, GST02b, GST04a, GST06a, GST06b, Gro90, HIS91, IMS90a, IMS91f, Koi09, KVK92, Las97, Mac96a, PG10, Sar00, Sar17, SSG^{+}10, SSG^{+}18, Smi01, XWK95, ZA11, vWAH^{+}02]. fundamental [HCD^{+}98]. Fundamentals [BCM99]. Fusion [SMSY02, RM90]. Future [Ken92b, Rei03, SZAB98, Zim02, HCD^{+}98, MVZ98b].

G [Ano90b, Dub97, Hop97]. G4 [Hun00]. GALAHAD [GOT03b]. gallium [SMB90]. GAMM [AH92]. gamma [HIK90, Smi01, Tho13]. gamma-spectrum [HIK90]. Gas [Ano90a, Cok93a, Cok93b, Lar93, Tea94]. gas-liquid [Cok93a]. gases [TS06b].

Gateway [RVV^{+}92]. Gauge [GAW96a, GAW96b, BW12, Cah90]. Gauss [RBD^{+}11, BB07, RBD^{+}10]. Gauss-related [BB07]. Gaussian [BKT91, HD05, KDG99, MKFB92, PW84, SGMS97]. Gaussian-Type [SGMS97]. GCC [Bro03, HDR03]. GCG [CH96]. gdb [But95]. Gehrke [GMC96e]. Gem [Cha09, Och09]. GEMM [KLV98]. GEMM-based [KLV98]. gene [Wri90b]. General [DGR92, FJ92, HC92, HC94, Uni2 , CM94, DGR90, Gro90, Int90i, Int90j, Int91c, Ove91, SS10, SMSW06, ZT90, FGJB19]. generalised [NSU20]. generalization [Wol91]. Generalized [KH13, TS25]. generate [GBC92, NJ94a]. generated [FTPR04]. générateur [Gom90b]. Generating [BN97, BCC^{+}91a, BCC^{+}91b, Bla00, CV94, CGL^{+}93, CGL^{+}95b, DI90, FJS97, SOG94, Wal92, BB07, Cha09, Gho01, GS98, GKH^{+}92, Tsa01, ZT90, BCC^{+}92]. Generation [AMC01, Ano90a, CG96, CMKH03, FGL01, FES05, KHS96, NL19, Paj90, SGMS97, SSC00, BS13, BGV94, Cas89b, EFP07, EJLC97, GV92, GKH^{+}93, Hen95, KHS95, LMJC96, LP90, MS00a, MS00b, TRV96, VRT97, YH93]. Generator [HD93, Lev92, Wal91b, WW92, FSV90, Gom90a, Gom90b, Hen94, Jam94, Jam96, LS09, MZT90, STVS91, SOP93, Wol92, vWAH^{+}02]. generators [BS13, CBTL97, DW03, Jam90]. Generic [She92, BxCW01, CM91, Cur94, SSS99]. Genesis [Hey94]. genetic [FHE95]. Genotypic [CHL94]. GENTRAN [BGV94]. GENTRANS [Kea92]. geoacoustics [Ame90a]. geodesic [AF92, Rib02]. GEOFLUID [Lar93]. Geological [FKL94, Coh90, Car93, CB95, McG91]. geologischen [Por90]. Geometric [Sug95, Raj95]. geometries [AF92]. geophone [Bec91]. geophysics [Dut94]. Geosciences [FYR99, YRF02]. geothermal [Kut92]. German [EMR93]. Germany [Ano97a, AH92, BPG94, Ein91, GH94a, GH94b, GH94c, AFKL04, Sch93a, KSW93]. get [Ano95d, Wic89, HDH^{+}94]. Getting [SB92, Thi91]. GF [PW84]. Ghinsu [Liv91]. Ghosh [Tsa01]. Giants [Ola96]. Gibbs [BSS92, SH07, SH09]. Gigaflops [BHMS91a, BHMS91b]. Gisela [Ano97a, TDMC97]. Give [DZ98, vK94]. GIVE-N-TAKE [vK94]. given [Gil01]. GIZ [GST02b]. GKS [Ame96, LMG95, San92]. GLIM [ZMR^{+}91]. Global [GK06, PWD93, SWH15, ADH95, GBC92, HWS09, KH93, Ogi02]. Globally [WMMW97]. glossary [IBM91c]. GMAO [GK06]. GMD [BRH90]. GMRES [FGGL05, FGG09]. GMS [She91]. Gmunden [Vol93]. GNU [AS97, But95, SZAB97, SZAB98, SZAB99, YSMA^{+}17]. Gnuplot [AG95a]. Goals [Ano93e]. Goddard [WBS97]. Godunov [Mal91]. gold [Coh90]. Goldberg [GMC96f]. GOLDCALC [Coh90]. goodness [GH18]. GOR [Tea94]. Gordon [STVS91, DKMS91]. GPGPU [MA18]. GPIC [XHY^{+}24]. GPOPS [RBD^{+}11, RBD^{+}10]. GPSS [Chi91]. GPSS/PC [Chi91]. GPU [MSP^{+}22, iSYS12, XHY^{+}24]. GPUs [BC19, GML^{+}16, KMJ^{+}23]. gradient [Kut92, LN91, MN01, Sav95]. gradients [Spe94]. grafting [BN93]. Grain [CL97, HK91, LE98, KY98a, KY98b]. Grained [BR96]. Grand [BEH^{+}94, Ten93]. Grande [ACM01]. Grande/ISCOPE [ACM01]. Graph [BCT94, HK92, RR99]. Graphical [Ame96, HT91, UHP91, Cur94, LP05, Par94, She91]. Graphics [Ame97a, Lan93a, Ame90a, AP90, Blu91, HW91, Hor09, IEC90, Lib90a, Lib90b, ISO90]. graphique [II90]. Graphs [OE92, NH09]. GRASP [FPR01, Has06, Hop03, PPR97, RPL96, RFS98, RR99]. GRASP2018 [FGJB19]. gravitational [FYR99, YRF02]. gravity [FR94, Lop90, McG91]. Gray [Dub97]. Great [Lap96]. Greece [HMPT94]. GRESS [Hor91a, Hor91b]. grid [Deu90, RRX^{+}08, STVS91, SR04]. grid-free [RRX^{+}08]. gridded [MC96, Phi91b, Phi92]. Grids [BLW02, CN94, Gou93]. GROMOS96 [BCS01]. Gross [BKRG22, KRG21, KYSV^{+}15, MA09, TS06b, YSVM^{+}16, YSMA^{+}17, YSMBA23]. ground [HW95, Joy92]. ground-based [HW95]. Group [Zei92, Sch94]. Growth [Gar91a, Gar91b, Ger94a, Szy07]. GSAP [HIK90]. GSL [Rap94]. Guaranteed [Nak95a, Nak90]. Guest [Hat94]. GUI [Ana93a, Ana93b, Sal95]. Guide [Air04, And92a, And92b, ABB^{+}95, BDPW98, BGA90, BA95, BGA96, Cha95a, dCH94, ED99, Fah94, GMC96f, Geh95, Geh96, Geh97, HKS91, IBM93, Scixxb, Sco93, Sun92b, Sun92a, ZB94b, Ano98a, Bak91, BGA94, Con91, CZ90, Con90, CSS90a, CSS90b, CSS91, Cro90, Dig90a, Dig90b, Dig93a, Dig93b, Del93, FG93, Hew92a, Int90b, Int90d, Int90h, Int90m, Int90n, Int90e, Int91b, Int91f, Int92, Ing90a, Ing90b, Law01, Num90a, Num91b, Num93a, Pag95, Sil92b, Sof93, Spexx, Sun93, WHL92a, WHL92b, Yip90, ZT90, GMC96e]. Guidebook [LW89, Tho97a]. Guidelines [PWD93, Cok93a]. GUIs [Bra91]. GUT [EH07a]. Guy [Eme94, Rag95, UMM94].

H [Adl93, AB93, BSS92, Eme94, Gar93, GMC96f, Kon94, Loz98, Tha93, UMM94, Wu93, Yan94b, Anoxx, CT11, Lar93]. H. [Gho01]. H2SOLV [PZY16]. hadroproduction [WW14]. Haenszel [Nan93b, RH94]. Hague [Ano93n]. haksup [nY90]. Halos [Ben99a, Bra00, Ben00]. Halstead [Sch91b]. Hamiltonian [BBB00, BK06, BSV16, GBDB97]. Hamiltonian/Hamiltonian [BSV16]. Hampton [Wie94]. Hand [Fuj95, FTPR04]. hand-coding [FTPR04]. Handbook [A^{+}92, ABW92, ABM^{+}97, Num91a, Rag95, Ano98b, KLS^{+}94a, UMM94, Eme94]. handing [ISO00]. Handling [BBCR98, CZM93a, HFT94, HFT97, IFI93, Rei95b, Wei95, ABC^{+}96, IEC98a, Int98a, Rei95a, Rei95c, Rei97]. Hands [CS90a]. Hands-On [CS90a]. Handwritten [Dya95]. handyG [NSU20]. Hankel [Wie99]. Hansen [Off98]. hardware [SJ94]. Hare [Wei94]. Harmonic [BD14, MBGK11, PS08, TS06b]. harmonic-oscillator [MBGK11]. Harness [Gli96, AH90, AHZ90, CA92]. Harray [YYM93]. Hartree [HKŠ^{+}97, KS12, PS08, SS09]. Harvard [Par86]. Haven [Ban93]. having [MIN^{+}95]. Hawaii [ERS95, HS94b, HS94a, MS94]. headers [Cha09]. Healing [GWE^{+}05]. heat [Car93, iSYS12, Car93]. heavy [SH97, WW14]. Hector [RFRH96]. Heidelberg [Ano97a]. held [NBC92]. Helmholtz [Kir93, Kir98]. hemisphere [Cum90]. HeNCE [BDG^{+}94]. HEP [DH84]. here [JH86]. Hermitian [CS14]. Hessenberg [HD05]. heterogeneous [ADB94, BDG^{+}94]. HI [HS94a, IEE96, HS94b]. Hiebert [NRS92]. Hierarchical [Ame97a, IEC90, ISO90, BMV03, JC93]. Higgs [DKM07, EH07a]. High [ACM97, ACM98, AMGM20, AOL94b, Ano94d, BGS94a, BPG94, Bee96a, BBZ94, BM99, BEH^{+}94, BCF^{+}94a, BCC^{+}96b, CC95a, CMZ93a, CMZ94a, Cre90a, Don95, Dow93, Ein91, FJSD96, Fos94, Fox91a, FGG09, Fuj95, GS01a, GH94a, GH94b, Ger98a, Ger98b, God93, HMR^{+}15, HS95, IEE94d, IFI95, Lin93, Lov93, Lov94, MCH96, MA18, Per93, Rag95, Sab95, Ten93, USE94, UMM94, WD98, Wea94, Zos93, Ano93q, BCM^{+}93, BID95, Bre92, Car91b, CCJ93, CDF^{+}93, DLLR96, DH95, Duv92, Eme94, EN96, FGGL05, Jam94, Jam96, KLV98, KT00, KO94, KC94, Lee90, MKF95, OM92, Sar97, SSG97, XHY^{+}24, Zim07, Adv98, AMC01, ADHF96, ACIK97, AOL94a, Ano93c, Ano93e, Ano93g, Ano93h, Ano93f, Ano93k, Ano94e, Ano94f, Ano94m, Ano94o, AGG^{+}97, BZ99, Ben99b, BB02, BFHH94, BMN^{+}95]. High [BMN^{+}97, Bra94c, Bra94d, BCC^{+}96a, BCC^{+}97a, BCC^{+}97b, BGMZ92, CLiN^{+}02, CZM93b, CZM94a, CMZ95, CCW04, CKZ93, Cou97, DDcMR96, DL97a, DL97b, DS97, DZ98, DCR99a, Din99, Eli98, FXAC94, GH94c, GOS94, Hig92, HM96, Han98, HBB^{+}95, Hat94, HF95, HJT97, HJJ^{+}00, KMR^{+}97, Ken94b, KK95a, KK01, KS02, KKZ11, KMBK96, KMS^{+}95, KOM93, KOM94, Koe92, KLS^{+}94a, KGV97, KK94, LMMW96, MB95, MMY95a, Meh93a, Meh93b, Meh94, MVZ98b, MVZ98a, MZ00, MZ01, MH95, Met95, MMV95, MMY95b, MR95b, NOL97, Off98, PFS^{+}04, Paz96, RMCKB97, SSH08, SZM98, Sch96a, Sch97, SNMC93, SIOS02, Ste93, Tho93, Wag94, YGS^{+}94, YFH97, Zim02, dSL98, van94a]. High-Dimensional [BM99]. high-efficiency [XHY^{+}24]. High-level [Ger98a, Ger98b, Wea94, DLLR96]. High-Order [CC95a, Fuj95, AMGM20, Sar97]. High-Performance [BGS94a, Bee96a, BEH^{+}94, FJSD96, Fos94, FGG09, GH94a, HMR^{+}15, IEE94d, Lin93, Per93, GH94b, HS95, Ano93q, KLV98, BFHH94, Bra94c, CZM94a, GH94c, SSH08]. High-powered [Cre90a]. high-quality [Jam94, Jam96]. high-resolution [DH95]. high-speed [OM92]. higher [CM94, KHC92]. Highly [AAC^{+}04, HJT97, KSYE00, PW84]. Hilton [IEE90a]. HiPPI [JA92]. HIRLAM [GS95]. histograms [GH18, GHN19]. History [HOP93, Kin93, Lor19, McJ17a, RLS20, Zim02, Bac98, Nan93a, MVZ98b]. hits [Ano95g]. HiWEP [Zim02]. HIZ [GST02b]. hole [LZL11, Taq16]. Holland [Nan93c]. homogeneous [KKY99]. Homotopy [WMMW97, SMSW06, WSW00].

HOMPACK90 [WMMW97]. Honolulu [IEE96]. Honor [NRS92]. HOPL [HOP93]. HOPL-II [HOP93]. horizontal [Coo94, McG91]. House [Eme94]. HP [GMMM92, Hew90b, Hew91a, Hew91b, Hew92a, Hew92b, Hew01, TOML04]. HP-UX [TOML04]. HPC [Fox94, Loh10]. HPF [ABC^{+}96, Ano94g, Ano94h, AMKS02, AHOK02, Ben99b, Ben99a, Ben00, BF01, BDPW98, Bod94, BB96, BR98, Bou95, BCFH93, BCF^{+}93b, BCF^{+}93c, BCF^{+}94c, BCF^{+}94b, BCF^{+}94d, BMMN94, BID95, BZ94, BD96, BG96, BCC^{+}97b, Bra00, BSCV95, BxCW01, BLW02, Bri00, BMV03, CNBB96, CMT01, CL97, CMZ94b, CZM94b, CM98, CGSS94, Coe94a, Coe94b, Coe96, CA96, Cou97, DL97c, Del98, DS01, DS02, DCR99b, DRST03, EGKU99, EGKU02, FGL01, FGRT00, FSPC^{+}02, FKKC96, Fox94, GLPE97, GS01b, Guo01, GMS^{+}95, HKM98, HLJ01, HCLJ03, IK96, IHKvW02, ISKvW02, JB01a, JB01b, Jou95, KKS^{+}95, KHS96, KMS^{+}95, LZ97, MM94, MBFC99, Met99a, MAH^{+}02, Nak95c, Nak95b, NJ94c, NNON02, Ogi02, OA02, OP98a, OP98b, OPP00, PSG03, PHHF94a, PHHF94b, PH96, PD96, Pon94a, Pon94b, Sai95, SM02a, SF02]. HPF [SMSY02, SNK06, Sch96b, SZG95, SIOS02, SIDH95, SM02b, SVD96, SDv98, Smi95a, Spo94, SS00, SN94, TBC94b, TCF94, TRV96, UZCZ97, Van94b, Vee94, WSL94, Zim99, vDSP96, vWAH^{+}02]. HPF-Builder [DL97c]. HPF-combined [MIN^{+}95]. HPF-Like [Guo01, CMT01]. HPF/Fortran [Ano94h, PHHF94a, PHHF94b, PH96]. HPF/JA [AHOK02, ISKvW02, Ogi02, SIOS02]. HPF/SX [MAH^{+}02]. HPF2 [BBCR98]. HPFBench [HJJ^{+}00]. HPFIT [BCC^{+}96a, BCC^{+}96b, BCC^{+}97a, BCC^{+}97b]. HPO [Dig90a]. HSPEXP [LMK94]. HSPF [Neu01]. HTML [Nai17]. hui [yKxx]. Human [Gal91]. Hungarian [Fer92, FK95].

Hungary [Fer92, FK95, Cse99]. hybrid [GRW07, LW07, Sre92]. hydrodynamic [RBS93a, RBS93b]. hydrogenic [PG10, Sar00, Sar17]. hydrologic [Pel93]. Hydrological [Neu01, Uni93, Bra94a, LMK94]. Hyper [TBG^{+}02]. Hyper-Threading [TBG^{+}02]. Hypercube [BF92]. hypercubic [CHM91]. Hypergeometric [NPB92].

I. [GMC96b]. I/O

[BLW02, LG93, LHHJ91, SW94, Coe94a]. I/Os [CFPS94]. IA [AAC^{+}04]. IA-64 [AAC^{+}04]. IBM [BBB^{+}94, Bel90a, Bel90b, CK90, CT11, GR92, GMS^{+}95, Int90c, Int90d, Lin93, Mra94, Per93, Pet91, SPS^{+}91, Sai95, Sar91, Sar97, SSW91, WTW90]. ICFP [Coc03]. ICIP [IEE94b]. ICONIC [CB94]. ICPP [Agr95]. Id [Nik93]. Idaho [Neu01]. ideal [Loh10]. identical [LSZ92, WHL95]. IEC [IEC98a, IEC98b, ISO04a, AC92, Ame97b, ISO90, ISO94, Int97a, Int97b, Int98a, Int98b, Int99, Int00, ISO04b, ISO10]. IEEE [ACM97, IEE92a, ACM98, HM92, Ins92, IEE94g, Kar95]. IF1 [LE98]. IFAC [Bar92]. IFIP [Boi97, BT01, CGS94, DR94a]. Igniting [ACM03]. II [HS94a, YRF02, HOP93, Ano94e, Bac98, BK06, BCC^{+}97b, CM94, Goo90e, Goo90f, Hig94b, Hig94c, Hig94d, Mar92, McJ17a]. III [Ano94f, Bac98, BPG94, Hig94a, VKB93]. Illustrated [Tho97a, Ano98a]. IMACS [AH92, HR92]. IMACS-GAMM [AH92]. image [Lan90a, MKS94]. imaginary [GST04a, GST04b]. imhan [nY90]. IML [SB01]. immun [nY90]. Impact [BKT91, Hat94, WBS97, CKT85]. imperative [BMO90, OM90, OT93]. Implementation [ACM93b, ARS92, ARS94, BCF^{+}93c, CP93, CZM94a, CA92, DLM99a, DDP94, DD99, DGL91b, DDHD90, DCHH88b, HH18, KBKT94, KK94, LM90b, LZ97, MAH^{+}02, RP12, Sht19, SOG94, TBG^{+}02, Wal00, vDSP96, BRH90, BCF^{+}93b, BMV03, CTS96, CMZ94a, CMZ95, CFMR95, CDGM96, DLLR96, DCR99a, DV01, DV02a, DV02b, GRSS02, Jam94, Jam96, KKS^{+}95, LM90a, MKS^{+}96, Mic97, PSG03, PCS99, QRH00, RBS92, SS09, SKM94, VKB93]. Implementations [BCH^{+}06, MT90, CCW04, CDMC06, GML^{+}16, HKM98, KLV98, KM99, Phi91b, Phi92, Sul91]. Implemented [Lin93, Per93, ARB94, ARB95, Bos19, PW84]. Implementierung [Kru90a]. Implementing [AS97, BBG^{+}93, BD96, But95, DL97a, DPZ97, DVY00, GHSJ94, HFT94, HFT97, Rei93, SOP93, SH07, SD99, CM91, DN04, NNON02]. Implementor [CKZ93]. Implications [AH94, AH91]. implicit [KTP^{+}24, KBKT94]. IMPLO [GT92a, GT92b]. Importance [Bra03]. improve [TJ90]. Improved [JP95, NG93, GST12, Nag90]. Improvements [BCT94, Zag16]. Improving [CCK90, Lev95b, Sal92]. Inc. [Zei92]. incidence [YK90, YB92]. included [Ame96, Ano97a, Ano98a, Ude91]. Includes [Rub93]. Including [Cou97]. Inclusion [Air04, NVC96]. incommensurate [Smi93b]. Incomplete [JP95, Tho13, PW93]. incompressible [KBKT94]. Incorporating [MRCB23, MRC18]. incorrect [BBF^{+}92]. Increased [CP93]. Incremental [KHS95, SAS90, EKC95]. increments [How91, SH91]. Indefinite [DR95b, DR94b, DR95a, Duf04]. Independent [FC09, HKT92a, RFS98, SB91, SFB92, SWM95, Sil01, HKT91c, Ken94a, Str05]. Index [Ano07, KHS96, CM94, IBM91c, KHS95]. India [Kum94]. Indices [MC92]. indirect [DSv94]. INDO [SS09]. induced [How91, LR91, SH91]. Industrial [Kon00, BLT94]. industry [Ano95d]. inelastic [AIS^{+}97]. Inequality [MHdL12, ZT90]. Infer [VBB18]. inferred [ORO15]. infinite [EO94]. infinitely [CNP91, Dut94, YB92]. Influence [KZ94a, KZ94b]. InfoDock [Ano97b]. infographie [II90]. informatics [Pri93]. Information [Ame97b, Ame97a, Ano94n, Ano94o, Ins91a, IEC90, IEC94, IEC97, IEC98a, IEC98b, IEC99, IEE92a, IEE93b, ISO90, Int97a, ISO00, Int00, ISO04a, ISO04b, ISO10, JL93, KH13, Met99c, Met99d, Met99b, Int90i, Int90j, Int91c, Ins91b, Ins92, II90, II91, LMJC96, ISO94, Int97b, Int98a, Int98b, Int99]. infrastructure [WFW^{+}94]. infusion [WHL92a, WHL92b]. INGRES/EQUEL [Ing90a]. INGRES/ESQL [Ing90b]. Inheritance [Mor15, DNS98]. Initial [BG97, Cas89a, CC92a, EP87, Hig91, BG94, IDVV97, Xu93]. initial-value [IDVV97]. Initiation [BDGxx, AL92]. Initiative [KLM91]. initio [HKŠ^{+}97]. injection [PBU95]. inline [CHT92]. Innovation [ACM03]. Input [And90, And92b, Are90, MR93b, Pra90, TS25, BN96, Tho86]. Input/Output [And90, And92b]. INRIA [Glo91b]. Insight [IEE02]. Inspection [NJ94c]. Installation [BDPW98, Dig90a, Dig93a, IBM91a, IBM91b]. instantiation [DV98]. Institute [Ano94p]. Instruction [SS93, Vaj92, Cho91, HT91, KE93]. Instruction-level [SS93]. Instructional [Schxx]. Instructor [BS91b, Spexx]. Instrumentation [Bli90, Yan94a]. intake [Lin90]. Integer [AMC98, Shi93b, BKK94, Hig93a]. Integers [MP93, Ric06]. Integral [AJ98, BB91, Dre92, Dre93, Smi11, CA90, Jon92a, Jon92b, Kir93, Kir98, LP90]. Integrals [Amo90, SGMS97, Som98, TS25, Car91a, Gao05, Gao06, HMT90, Mac96b]. Integrated [ASS95, BGG^{+}94, BCC^{+}96a, BCC^{+}96b, CFK^{+}94, DCZ96, JL93, BCC^{+}97a, BCC^{+}97b]. Integrates [FXAC94]. Integrating [AP90, CM98, LMJC96, CMVZ94, MVK25, YWS^{+}94]. Integration [HIM91, CRS90, EO94]. Intel [Ano02, BRH90, GAW96a, GAW96b, KR94, KR95, McB91, SZG95, YSMA^{+}17]. Intelligence [BPG94, HR92]. Intel(R) [TBG^{+}02]. Intensive [Bel90a, Bel90b, GR92]. interacting [PS08]. interaction [DRST03, Eli98, HKŠ^{+}97, Sar00, Sar17]. interactions [GLS93, MMEH08]. Interactive [Ame97a, CC90, CS90a, HKTW94, IEC90, ISO90, KMT91, Kry94, Pao99, SAS90, Tay99, Ame90a, Coo95, HM92, HHK^{+}93, KNOR04, MKF95, Pao01, SMB90, Sil93, SWO92]. InterCall [Wei94]. interchange [AK84]. Interface [AG95a, Ano94j, Ano96b, BDC^{+}96, BKP93, BG96, BHY80, BDK91, DFL92, DDH^{+}95, DDHW96b, FKKC96, GRE99, HBG01, IEE92a, MJR93, NLE^{+}20, SW94, YGS^{+}94, Ano94k, BDPW98, BxCW01, CH96, Cur94, GMF18, HBG02, HDH^{+}94, Hen95, Hor09, IEE90b, Ins92, KTP^{+}24, Lan93a, Nag90, Par94, She91, Sil93, BBB^{+}94]. Interfaces [BBZ94, BFKS93b, Ins91a, IEE92a, IEE93b, BFKS93a, Hem94, Ins91b, Ins92, II90]. Interfacing [All90, LMMW96, Och09]. interim [MSZ90, Ngu91]. interior [GT92a, GT94, GT92b]. Interlanguage [Mac91b, Mac91c]. intermediate [Nie92]. intermolecular [HMT90]. Internal [FWH^{+}94]. Internation [MS94]. International [ACM94a, ACM94c, ACM95a, ACM96a, Ano93m, Ano94a, Ano94i, AH92, BPG94, BV94, CKMU94, Cse99, DW94, ERS95, EPL94b, Fri94, GH94a, GH94b, Glo91b, HMPT94, HAM95b, HS95, HS94b, HS94a, HHK94, IEE95a, IEE96, II91, ISO04a, KRB^{+}90, Kum94, KSW93, Lev95a, NBC92, PBG^{+}95, AFKL04, Sen03, Sie94a, Sie94b, Vol93, WN90, Ano93q, Ban93, BGNP94, BLT94, GH94c, Hua96, Sch93a]. Internetworking [Ano93b]. Interoperability [NLE^{+}20]. interoperable [YBMCB14]. Interpolating [Cos97a, Cos97b, MC96]. Interpolation [BM99, CMV09, Ren97b, RB99, TZW^{+}10, KP93, MKC92, Ren03, Ren04, Yu01]. interpretation [BMO90, Ude91]. Interpretations [Ano94o]. interpreter [Rap94]. Interpreting [Ano94h, PHHF94b]. Interprocedural [CCKT86, CI96, CI98, DJ92, Hal91, HHKT92, HHKT96, Harxx, MS93a, CKT85, HK90, LM94, SRH96, YH93]. Intersections [NRS92]. intersegmental [SZ90]. interstage [MIN^{+}95]. Interval [AS97, Kea95a, Kea96a, Kea96b, KVK92, Oku95, SZAB97, SZAB98, SZAB99, Sun05, Wal00, Wal01, Wal02b, WS94, BBZ95, PBU95, KVK92]. Interval-Enhanced [SZAB99]. INTERVAL ARITHMETIC [Kea96a, Kea96b]. Intervals [Alo04, SB01]. interview [Tay86]. INTLIB [KDKSH92, KDDH94]. INTLIB ARITHMETIC [Kea95a]. Intrinsic [CCL01, Hig94b, Nag95, Sht19, Wal01, Wal02b]. Intrinsics [AS97, CCL04]. Introducing [CS95, CS00, LS90c, GMC96b]. Introduction [Cha94c, Cha97b, Chi91, DW08, Ell90, FJSD96, Hat94, Hud91a, Hud91c, Hud91b, Jon93, Kug92, LP98b, Mal91, MH95, NLN96, NL96, NL97b, Ort94b, Ort94a, Rou90, Sch96b, Var95, Wri91, Zim99, Cha95b, Ede90, Gri93, Och09, SSG93, Vil94, Wri90a, Mar98]. Introductory [Ric06, LD87, Num90a, Num91b, Num93a]. intrusion [FR94]. invaded [Dut94]. Invariant [BS92a, BS92b, BS97]. Invention [NRS92, Pri10]. inverse [Koi09, MZM94, PT93]. Inversion [DLM99b, DLM99a, GGLM88, GL90, GWDL08, GWDL10, KP93, NJ94a, NJ94b, Ple93, Tur93]. inverting [Gep90]. Invited [Meh93b, Zim02]. involving [Sar00, Sar17]. IO [Sni92]. Ion [PTS92]. ions [MCA17]. IP [Ano93b, JA92]. IPMLO [GT94]. IPPS [IEE96, IEE95a]. iPSC [BRH90, KR94, McB91, KR95]. iPSC/2 [BRH90]. iPSC/860 [KR94, McB91, KR95]. Ireland [HR92]. Irregular [Ben99a, BBCR98, Bra00, DL97a, DL97b, GLPE97, HJT97, LP98a, PHD^{+}95, PCS98, PCS99, Sch96a, UZCZ97, vKK^{+}93, Ben00, BCC^{+}97b, BSCV95, CC94, GF95b, HMS^{+}95, LP99, Luc92, MR96b, Pon94a, Pon94b, PSC^{+}95, SSG^{+}10, SSG^{+}18, SPM^{+}94, dSL98, vKK92, von92]. irregularly [Yu01]. IRT [SH07]. ISBN [AB93, Ano90b, Ano91a, Eme94, Gon01, Hop97]. ISCOPE [ACM01]. Ising [DM90, Heu90]. Isn’t [RMX12]. ISO [Ame97b, Ano98b, ISO04a, AC92, A^{+}92, Ame97a, ABW92, ABM^{+}97, IEC90, IEC94, IEC97, IEC98a, IEC98b, IEC99, Int97a, Int97b, Int98a, Int98b, Int99, Int00, Wic89]. ISO-Pascal [Wic89]. ISO/ANSI [Ano98b, ABM^{+}97]. ISO/IEC [ISO04a, AC92, IEC90, IEC94, IEC97, IEC99, ISO90, ISO94, Int97a, Int97b, Int98a, Int98b, Int99, Int00, ISO04b, ISO10]. ISO/IEC/TR2 [IEC98a, IEC98b]. Isopycnic [BDOS95a, BDOS95b]. ISPAN [HHK94]. ISSAC [Sen03, WN90]. ISSAC’94 [ACM94c]. ISSAC’95 [Lev95a]. Issue [Ano94m, BDH90, Hig94b, Hig94c, Hig94d, KS02, W^{+}95, ANS95]. Issues [Coe96, FGL01, Goo90c, Goo90d, McC96, Nak95c, CMT01, MKS^{+}96, Nak95b]. Italy [DR94a, Don95, HS95]. Itanium [Hew01, TOML04]. Itanium-Based [TOML04]. Item [Nan93b, RH94, SH09]. Iterated [BB91]. iterates [ZT90]. Iterative [dCH94, DL97a, WW90, ARB94, ARB95, dSL98]. Ithaca [PBG^{+}95]. ITPACKV [KO90]. IV [Anoxx, BH92, CD92, Rocxx, KSM95, SD90, Spexx]. IWPP [Kum94]. IWPP-94 [Kum94]. IWWP [Kum94].

J [Ano91a, Ano98a, Ano99a, Ano99b, GMC96b, GMC96d, GMC96f, Gla92a]. J. [GMC96c]. JA [AHOK02, ISKvW02, Ogi02, SIOS02]. Jacobi [TS25]. Jacobian [FTPR04]. January [ACM91, ACM94b, ACM95b, Eme94, HS94b, HS94a, AFKL04, ACM93c]. Japan [CKMU94, HHK94, WN90, IFI95]. Japanese [SM02a, SF02]. Java [ACM01, Ano97b, Ano97c, BGH^{+}06, Bra97c, BSPF01, BSB^{+}03, DD97, DDS99, FCHE02, FLQZ97, LP05, LS04, MMG98, MMG00, Och09]. Jeanne [Ano98b]. Jerrold [Ano98b]. Jim [Bli94]. Jindivik [Lin90]. JLAPACK [DD97, DDS99]. JMASM [FC09, SH07, SH09]. JMASM1 [Fah02]. JMASM10 [Alo04]. John [Ano96a, Ano98a, Ano99a, Ano99b, Rub93, Aik07, Bjø08, Loh07, McJ17c]. Joint [BV94, KSW93]. Jose [ACM97, Ano94a]. Joseph [Hin06, Iha06, Sch07]. Journal [Ano93i]. Jr [ACM99, Eme94, Rag95, UMM94]. Julia [GKKL19, MSP^{+}22]. July [ACM95a, Ano95c, Bar92, Boi97, HMPT94, HR92, IEE92b, Lev95a, NBC92]. June [ACM01, Ano94l, Ano95a, DSZ94, DW94, Don95, FH90, IEE92b, Wie94, ACM93b]. JVM [SD01, SD03].

K17 [MDD94]. Kadanoff [KKY99].

Kadanoff-Baym [KKY99]. kaeron [nY90]. Kalman [Tor10]. Kanazawa [HHK94]. KAP [KLS94b]. kappa [RR92]. Karlsruhe [Ein91, KSW93, Sch93a]. KB0013 [Sal95]. KeLP [MBFC99]. Kemari [KMR^{+}97]. Kent [Ano90b]. Keras [OPB^{+}20]. Kergen [MSZ90]. Kergen-Kersim [MSZ90]. Kernel [Ame96, NLBB23, NLVE15]. Kernels [YFH97, WMCU97]. Kerrigan [GMC96c]. Kersim [MSZ90]. Keulen [Sto93]. Keulen-Seligman [Sto93]. Key [ABMS94]. Keyword [Tho86]. Keywords [Ham85, HM90a, RH84]. KFKI [FK95].

KFKI-1995-2 [FK95]. KFKI-1995-2/M [FK95]. Kind [BB91]. King [ACM99]. Kingdom [Boi97]. Kirchhoff [Cap98]. Kirchhoff-Plates [Cap98]. Kit [Ano96b]. Knapsack [MT90]. Kniga [ES93b]. knowledge [KT94, LMJC96]. known [Ste91]. Knoxville [IEE94d]. Koelbel [Eme94, Rag95, UMM94]. Koonin [Ano03]. KPP [AC17]. Kramnik [Coc03]. Krommes [Kro90]. KSR [Ken91, Ken92a]. KSR1 [Pap93]. Kutta [EH07b, GKKL19]. kyesan [nY90]. kyojae [nY90]. Kyoto [IFI95, IFI95].

L [Ano98b, Ede90, Eme94, Hin06, Iha06, MN11, UMM94, ZBLN97]. L-BFGS-B [MN11, ZBLN97]. L1PMA [Dem03]. L2CXCV [Dem06]. L2CXFT [Dem95]. L2WPMA [Dem07]. Lab [FH92, WC92]. Laboratories [Olv91]. Laboratory [CFGG94, FGCG94]. LAD [KTMB02]. LADFEUDX [ZZN94]. Lagrange [CMV09]. Laguerre [KL92]. Lahey [Ano92b, Bee01a]. Lahey/Fujitsu [Bee01a]. Lamb [STY15, STY18]. Lamb-shift [STY15, STY18]. lamina [Raj95]. laminates [ZE92]. Lancelot [CGT92b, CGT92a]. langage [BDGxx, DV91, II90, Robxx, RD91]. Langley [Wie94]. Language [ACM93b, Ame90b, Ame97a, Ame92, Ano91b, Ano93g, Ano94e, Ano94f, Ano94m, Ano94o, App91, Bee96b, Bee96c, Bee97, Ben95, BFHH94, BEH97, BDK91, CMZ91, SG93a, SG93b, SG93d, SG93c, DFS95, Ein95, Ins91a, Hig93b, Fos93, FC95, FHK^{+}90b, Fox91a, FMW^{+}94, GMC96e, Geh95, Geh96, Geh97, Guo01, Hig92, Hig94a, Hig94b, Hig94c, Hig94d, IEE92a, IEE93b, Irv91, KA95, KO91, KKMP95b, Lan90d, Lor19, Nik93, OC94, Rys95, See04, XH90, AC92, Ame97b, Alg90, Ano91d, Ano91e, AJJF14, Con91, Cel96, CS90c, CS91, Cra93, Dig92, Rocxx, EN96, FCHE02, FC92, FHK^{+}90a, Fox91b, Hew90a, HIS91, HMS^{+}95, Int90a, Int90c, Int90f, Int90g, Int90l, Int90k, Int91a, Int91e, Int91d, IEC90, IEC97, Ins91b, Ins92, ISO90, Int97a, Int97b, KKMP95a, Lan90e, Loh10, Mar07, Met99c, Met99d, Nor91, OJ09, RD91, Sil92a]. language [Szy07, UZCZ95, UZCZ96, Wal93a, WHL92a, WHL92b, ZBC^{+}92, Zim92, ISO04a, CC94, ISO04b, ISO10]. Languages [HOP93, ACM93c, ACM94b, ACM95b, Ano93m, Ano94a, Ban93, BGNP94, CZM94a, CMZ95, CMKH03, Fos94, Fox94, GKKL19, Hua96, Ker93c, Kin93, Knu03, Mar93, PZA93, PBG^{+}95, PMM^{+}08, PHD^{+}95, SM92, Sch93b, SS96, USE94, VCV97b, Wil93, ZA93, ACM91, Ame97b, ASM^{+}94, BMO90, BBF^{+}92, Cas14, CMZ94a, Dot93, HMPT94, IEC94, IEC97, IEC98a, IEC98b, IEC99, II91, ISO94, Int97a, Int97b, Int98a, Int98b, Int99, ISO00, Int00, ISO04a, ISO04b, ISO10, KGV97, Nan93a, Nic91, OJ09, ST95, Tay97, VCV97a, Zim07]. Lanthanum [KLA95]. LAPACK [And92a, ABB^{+}95, And02, Bee01b, BDC^{+}96, DDH^{+}95, DDH^{+}96, DDHW96a, DDHW96b, DD97, DDS99, MFK09, Phi91a, She92]. LAPACK3E [And02]. Laplace [AJ98, DLM99b, DLM99a, GGLM88, GL90]. LARC [Wie94]. Large [BC01, Bou97, CT95, CGT92b, JPE20, KSZ90, KMJ^{+}23, MA18, NPB92, PR91, Ric06, SF92, SM03, TT92, VBA95, BHLT09, BxCW01, CGT92a, GOT03b, LS09, LN91, LMV09, MN11, NY91, Str05, Tor10, TOC18, ZBLN97]. Large-Scale [BC01, CT95, CGT92b, JPE20, KMJ^{+}23, PR91, SF92, SM03, TT92, VBA95, BHLT09, CGT92a, GOT03b, LS09, LN91, LMV09, MN11, ZBLN97]. Lärobok [Ein94]. Larry [Rub93]. laser [Eli98]. lasers [Ano94l]. latency [KE93]. Latent [Gre93]. Lattice [Ano94i, GAW96a, GAW96b, BW12, Mai90]. lattices [CHM91]. Launcher [NLBB23]. Layer [OP98a, PFS^{+}04]. layered [CB95, FYR99, Par94, YRF02]. Layout [KK95a, KK98, BKK94]. Lazy [Mas93b, Mas94]. LCPC [Hua96]. Learn [Pif96]. Learned [MWO95]. Learning [OPB^{+}20, CB94]. Least [Dem95, Sou91a, Sou91b, CS14, Dem97, Dem06, Dem07, GT07]. least-squares [CS14, Dem07, GT07]. Lecture [Meh93b]. lectures [DKMS91]. LEED90 [BRdAHK04]. Leestma [Rub93]. left [VLLY92]. Legacy [AC97a, AC97b, Gli96, McD93, Nat00, RRM^{+}15, EKC95, Pre99, SFKL02, TMD13, Van22]. Legendre [SSG^{+}10, SSG^{+}18, TS25]. Lemaître [Rib02]. Lemmon [Hin06, Iha06, Sch07]. Length [Mey01, Ves91, CK86, IEC94, ISO94, Int00, Cou97]. Lengths [Sil01]. Lesson [GWE^{+}05]. Lessons [MWO95]. lethal [RKMJ92]. Letters [GWE^{+}05]. Level [DDP94, DH92, Ein91, Hig90b, Lin93, Per93, USE94, Vaj92, CCJ93, DLLR96, EN96, Ger98a, Ger98b, KLV98, MKF95, Mit02, SNK06, SS93, Wea94, ZJEP95a, ZJEP95b, DD99, DDHD90, DV98]. Level-3 [DDP94, DH92]. Lexical [Dya95]. lexically [BGS82]. Libm [TOML04]. Libraries [Cro92, EGKU99, Ham95a, IEE94e, IEE95b, KNOR04, MFK09, Bee01b, Bee01g, Bee01f, Bee01e, BCF^{+}94a, BV13, Hor96, LHW01]. Library [AMGM20, Bra00, BKR^{+}91, CMKH03, For95, FHS78, Fox79, Hig94b, Jéz93, KDKSH92, KDDH94, Kry94, MJR93, Olv91, SD99, TOML04, TS25, Wri99, BS13, CHHW94, Coo95, Cra91b, Cra92, DDH17, DLW^{+}18, Du 97, GT92a, GT94, GOT03b, HKM98, HM12, HW91, Int90f, Int90g, Int90l, Int90k, Int91a, Int91e, Int91d, JCL10, KN95, KVK92, Mar92, MS00a, MS00b, Num90a, Num90b, Num90c, Num91a, Num91b, Num91c, Num93a, Num93b, Numxx, PQ94, Rap90, Sv90, Sch90, Ste91, Wal93b, GT92b, IMS90a, IMS90b, IMS91f, IMS91d, IMS91e, IMS91g, IMS91h, DG99, Vil94, vWAH^{+}02, Kri86]. Life [NOL97]. Lifetime [Huf93, ZMR^{+}91]. Lifetime-sensitive [Huf93]. lift [How91, SH91]. Lightweight [IHKvW02]. Like [Guo01, NLBB23, CMT01, KGV97, MM02, Yam95, Wal93a]. Likelihood [BGW93, BLL^{+}96, WCN92]. likelihood-based [WCN92]. limb [SZ90]. Limitations [Meh93b, Meh93a]. Linda [Sci93]. Line [LP19, NG93]. line-broadening [NG93]. Linear [ACIK97, dCH94, DGL91b, DGL91c, DGL91a, DDHD90, Don91, DV92, DCHH88b, DCHH88a, DR93a, DR93b, DV98, DHP02, GK06, GGHvdG01, HL94, KNS95b, LZ97, MKFB92, ONT95, TIUG90, WD98, DR95b, AR06, ARB94, ARB95, CS14, CWB92, CWB94, DH84, Don90, DR94b, DR95a, GT92a, GT94, HS10, KNS95a, LFG00, Mal91, OH90, Ove91, SG95, SMSW06, WSW00, ZT90, GT92b, vV90]. Linear-Time [KNS95b, KNS95a]. lines [FBC96, NG93]. lineĭnuyu [Mal91]. link [KSYE00]. link-cell [KSYE00]. Linking [Bee01b, GPS99, RVV^{+}92]. Links [BKR^{+}91]. LINPACK [She92]. Linux [Ano96b, Ano97b, Ano97b, Ano02, Bee01a, Del98, Hug96, KM97]. Linux/AXP [Ano97b]. Linz [BV94]. Liouville [BGKZ91, GM97, Pry99]. liquid [Cok93a]. LISP [BGS82, Gro91, Ume91, BW96, FBWR95, FT03, KA95, Rap94, Rei96a, SSG93]. listing [WRL90]. Literate [AO90a, AO90b, AO90c, AAK01]. literate-programming [AAK01]. lithofacies [KDG99]. lithosphere [HM93]. lithospheric [Av94]. Little [Ola96]. Livermore [CF90]. Lives [GWE^{+}05]. LLDRLF [BLL^{+}96]. Load [KH93, HM93, Kin92]. Load/store [KH93]. loading [Mir90]. Local [CGL^{+}95b, DS01, EGKU02, SVD96, CGL^{+}93, SDv98]. locality [ADH95, LW07, Zim07]. locality-aware [Zim07]. Localization [YKK96, HC08, KY98a, KY98b]. Localizing [CT90]. Locating [Sal92, Wal90]. Loft [Lov92]. log [BLL^{+}96, BLL^{+}96]. log-F [BLL^{+}96]. log-likelihood [BLL^{+}96]. logarithmic [BB07]. Logic [Jon93, Pri10, Kah01]. Logical [van90b]. London [Eme94]. long [Cah90, FSV90, GMF18, TRS91, YB92]. long-period [FSV90]. Look [NLBB23]. Loop [Bou96, FGL01, IK96, SF93, AK84, Ber92, DDH17, GF95b, LPA95, LP92, LP93, WMCU97]. loop-based [Ber92]. Loops [KK95b, TLS91, CK90, DFRR91, RP95, TLS90, YWS^{+}94, CF90]. losses [Xu93]. lossy [MIN^{+}95]. Lost [CL93, CL94]. Loveman [Eme94, Rag95, UMM94]. Low [BRdAHK04, Cra95, Lin90]. Low-Energy [BRdAHK04]. Löwdin [Jon92a, Jon92b]. lower [Cum90]. LPAR [BK95]. LPF [MIN^{+}95]. LPF/HPF [MIN^{+}95]. LPF/HPF-combined [MIN^{+}95]. LPI [Lan90b, Lan90c, Lan90d, Lan90e]. LPI-FORTRAN [Lan90e, Lan90b, Lan90c, Lan90d]. LRPD [RP95]. LSA [LMV09]. LSKUM [MSP^{+}22]. LSNNO [TT92]. LSODE [Rei93]. LTBEM [Azi23]. LU [vPMF92]. Lüscher [Jam96, Jam94]. Luther [ACM99]. Lyap [Sat97]. Lyapunov [Sat97]. Lyngby [DW94].

M [Ano91a, Dub97, FK95, Gla92a, KSW93, EKB92, FC92, Fos93, FXAC94, FC95, HKŠ^{+}97]. M-prep [EKB92]. M. [GMC96d]. M.I.T [Eme94]. M/Fortran [U.S01a, U.S01b, U.S01c]. MA [Ano98b]. MA47 [DR94b, DR95a, DR95b]. MA48 [DR93a, DR93b]. MA57 [Duf04]. Machine [CC95a, GG99, HKT92a, Lor19, HKT91c, Ken94a, AKLS88, BL91, BHMS91a, BHMS91b, DFL92, Sab92, Sab94]. Machine-Independent [HKT92a, HKT91c, Ken94a]. Machines [BR96, BMN^{+}97, Fat94, HHKT92, HKT92b, HLJ01, KHS96, KK98, LG93, Nik93, Pri10, YYM93, BBDR94, BBDR95, CC92b, Cho92, GS98, Hal91, HKT91b, HKT91a, HKT91d, HKT92c, HMS^{+}95, KN95, KHS95, Sal92, SM92, SNK06, TBC94a, Tse93, WSL94]. Macintosh [Blu91]. Macro [YKK96]. Macro-Dataflow [YKK96]. MacroC [CG96]. MACROFORT [Gom90a, Gom90b, CG96]. macros [BRH90, van90a]. MACS80.VERS.4.1 [WRL90]. MACSYMA [Cah90, Kea92, GBC92]. mafic [Nie92]. Magazine [FH92, WC92]. magma [AFBN93, FR94]. magmas [Nie92]. magnetic [PT93]. magnetohydrodynamic [KT00]. magnetosphere [Ogi02]. Magnitudes [NPB92]. Magsat [Av94]. Mainframes [Sha95]. maintenance [Num90c]. Make [MP93, JT94]. Makefile [Wes96]. Maker [Wes96]. Making [Pif96, Van22, VBA95, KT94]. Management [BL90, JL93, Rotxx, AW94, Che91, YH93]. Managing [RMX12, Off94]. mangrove [Pel93]. Manipulation [PEP92, Goo90a, Goo90b, Wea94, AP90]. Mantel [Nan93b, RH94]. Manual [Lan90d, Sci93, Scixxb, Sun05, U.S01a, U.S01b, U.S01c, Ano91d, Ano91c, BS91b, Con91, Con92, Cra91b, Cra92, Cra93, Dig92, Dig93c, Intxx, IMS90a, IMS90b, Lib90b, IMS91f, IMS91d, IMS91e, IMS91g, IMS91h, Jor90a, Jor90b, Lan90e, Lah90, LMK94, MSZ90, Num90b, Num91c, Num93b, Numxx, Nagxx, Ngu91, Rap90, Sil92a, Sv90, Sch90, Sun94, Uni93]. Manufacturing [JL93]. Many [Maaxx]. MAPLE [Gom90a, Gom90b, CG96, GS98, LP05, LS04]. Mapping [EB98, LE98, HC08, MKF95, SNK06, SV95, WTW90]. Mappings [CMZ93b]. maps [SS99]. March [Ano94d, Bjø08, IEE94g, IEE97]. marine [Ame90a, LHW01]. Mark [Num90a, Num90c, Num91b, Num93a, Num90b, Num91c, Num93b, Dub97]. markers [CB95]. Markov [BBZ95].

Martin [ACM99, Ano98b]. Mary [Eme94, Rag95, UMM94]. MAS [SSLG91]. Maskenorientierte [Por90]. mass [CA90]. Massachusetts [KRB^{+}90]. masse [MM02]. masses [EH07a]. Massively [DSZ94, FBZ92, IEE94a, Oed93, ASM^{+}94, BBDR94, BBDR95, CC92b, DR94a, Ger98a, Ger98b, LSW92, O’K93, OPE^{+}95, Sta94]. master [IBM91c]. Mastering [And90]. Masters [Mit92]. match [MIN^{+}95]. matching [Wea94]. MATFIT [Ram90]. Math [Ano92c, Kor99, EFG^{+}05, IMS91f, IMS91d, IMS91e]. MATH/LIBRARY [IMS91f, IMS91d, IMS91e]. MathCode [FES05]. Mathematica [Tay99, Fre92, FES05, LP05, Pao99, Pao01, Pri93, Tam95, Tho97a, Var97, Ano98a]. Mathematical [AAC^{+}04, KSW93, Mil04b, Tho97a, WNO94, Ano98a, IMS91b, IMS91f, IMS91d, IMS91e, XWK95, Olv91]. Mathematics [Bee01g, Bee01f, Bee01e, DG99, NRS92, Bos19, Mat90]. Mathieu [Err06, EC13, Shi93b]. Mating [Rit90]. MATLAB [RBD^{+}11, Tay99, CFGG94, DP96, DP99, DH12, LHW01, Pao99, Pao01, RBD^{+}10, RVV^{+}92, Rei93, Ano97d, Ano97c, Bra97d, Bra97c, FGCG94, JM94, LS04]. Matrices [GP97, Ram90, Rei02, SD90, BBB00, BK06, DV98, GP94, Jon09, LW07, dSL98]. Matrix [BS92a, BS92b, BS97, BD90, DL97a, DL97b, DGR92, DS94, FB12, GWL^{+}92, Han92, Hig90b, Hig90a, JSY^{+}20, MSC96, CZ90, DN09, DGR90, Hop02, KP93, KSM95, LS09, Pet91, Sar00, Sar17, UZCZ95, Var97]. Matrix-Matrix [DS94]. Matrix-Vector [MSC96]. Maximizing [MSC96]. Maximum [BGW93, RFS98]. May [ACM93a, ACM96a, DT94, Hig94a, Hig94b, Hig94c, Hig94d, HS95, HDR03, IEE94d, Met99c, Met99d, SS96, W^{+}95, ANS95]. Maze [Gil94]. mazev1 [DG08]. mazev2 [DG08]. mazev3 [DG08]. McFarland [Mar90]. McLean [IEE94a]. MD [IEE02].

Means [KLA95, Bin96, Gro90]. Measure [CRDO16, ORO15]. Measurement [Mit97]. measurements [Cum90, Lop90, SZG95]. measuring [Fu90]. MEBDF [CC92a]. MECCA [AC17]. mechanical [AM90, ZE92]. mechanics [FTD91, GL10, San92]. mechanism [Gep90]. media [CB95, FCHE02, Ano93b]. MEDINA [AC17]. Meeting [Ano93n, Ano95a]. meets [Tam95]. megaflops [MMG00]. Meiko [RBS93a, RBS93b]. Memoriam [Aik07]. Memory [BR96, BP92, BGLP94, BKP93, BCF^{+}93c, BCF^{+}94d, BMMN94, BMN^{+}97, CL97, CMZ91, DCZ96, Ger94a, GS97, HHKT92, Ham93, HKT92b, HLJ01, KHS96, KNS95b, KMR96, KK98, Mer92b, O’B93, PMBH93, PWD93, RSB97, Sch93b, Tal91, dSZP92, BZ99, BB02, Bod94, BCF^{+}93b, BCF^{+}94b, Cho92, CK91, DPZ97, Ger98a, Ger98b, GHSJ94, Hal91, HBB^{+}95, HKT91b, HKT91a, HKT91d, HKT92c, HMS^{+}95, JC93, KN95, KMR^{+}97, KHS95, KNS95a, KE93, OH90, PZA93, Phi91a, RBS92, RS09a, RMX05, RA90, SSH08, SM92, SNK06, TBC94a, Tse93, Wag94, WYJ99, WW95, WI94, ZA93, vPMF92]. Meredith [Ano03]. Merge [YWS^{+}94]. mesh [GBR15, IJCL96]. Meshes [SM17, Lai92a, Lai92b]. meshfree [MSP^{+}22]. Mesoscale [Mic97]. Message [Ano94j, BGLP94, FKKC96, KHS17, vDSP96, Ano94k, BDOS95a, BDOS95b, CHHW94, CA92, DOSW96, GS95, Kro14]. Message-Passing [Ano94j, vDSP96, Ano94k, BDOS95a, BDOS95b, CHHW94, CA92]. messages [BL94]. Metcalf [Ano91a, GMC96d, Gla92a, Rub93]. Meteorology [HK93b, HK93a, HK95]. Method [BM99, Cap98, DLM99b, DLM99a, GGLM88, GL90, HD93, KG99, Rhe93, WS94, YYM93, BBB00, Bin96, CM94, CC98, CA90, Dan90, DN04, Dot93, GRSS02, Gro90, HE13, HKŠ^{+}97, HP95b, KL92, LD90, Luc92, LP90, MVK25, MN01, OM92, PW93, RBD^{+}10, TRS91, RBD^{+}11]. Methodology [Nan93c, CDF^{+}93, GKH^{+}92, Tre91]. Methods [Bor91b, Bou97, CMK00, CC95a, dCH94, EL97, Enr95, EP87, Ett92, Fen96, GKKL19, Glo91b, GGHvdG01, JSW93, KSW93, MMV95, NL95a, Pao99, RS92a, Ueb97, Yam95, AH92, Bor91a, Bos19, CRS90, DLW^{+}18, Don95, Ede90, EH07b, GT92a, GT94, HKŠ^{+}97, KM99, KBKT94, KHC92, Mac96b, NL95b, Pao01, PRS99, PSC^{+}95, RS92b, Shi93a, SS99, Tay99, GT92b, NSWP90]. Metrics [HIM91]. Mexico [IEE91, Sie94a, Sie94b, USE94, ACM93b]. MH [RH94]. MHD [Kle93, Ogi02]. Miami [BDOS95a, BDOS95b]. Mica [Neu01]. Michael [Rub93, Tay86]. MICOM [BDOS95a, BDOS95b]. Micro [EO91]. Micro- [EO91]. Microcomputer [Dot93, FTD91, RKMJ92]. Microcomputers [Mar90]. microscopic [Var97]. Microsoft [All90, Ano91f, Ano92d, Ano93j, BL90, HFMS95, KLA95, Mar90, Pas95, SWM95, WRL90, Wal91b, WS94, Olv91]. Microsoft-FORTRAN [Olv91]. Middleman [Mil91]. Migrating [Ano95e, Ker93a, Ker93c, Ker93b, MM94, GMC96c, Rys95]. migration [HZ99, IBM91d]. Milan [HS95]. millennium [Met99c, Met99d]. Miller [Mil04a]. MIMD [BCF^{+}93b, BCF^{+}93c, BCF^{+}94b, BCF^{+}94d, Cho92, DDP94, GGW96, Hal91, HHKT92, HKT91b, HKT91a, HKT91d, HKT92c, HKT92b, LG93, MSC96, Tse93]. MIMDizer [SWW90]. Mine [Wri91, Wri90a]. minerals [Mai90, MSB92, SSLG91]. mini [Nagxx]. Minimal [DET12, DGS08]. minimax [ZT90]. Minimization [Buc94a, Buc94b, SF92, Hop98, LN91]. Minimizing [EDA96]. minimum [ADD04].

MININEC [Car90]. mining [WW94]. Minis [Bra90]. MINIX [HBG^{+}06]. Minneapolis [Ano94p, IEE92d]. Minnesota [IEE92d]. MINRES [CS14]. MINRES-QLP [CS14]. MISCFUN [Mac96a]. mise [LMG95, MM02]. mise-a-la-masse [MM02]. Mississippi [IEE94e, IEE95b]. Mixed [Ein95, Nor91, OM90, OPP00, HS10, Kir02, MWM90, Wic89]. Mixed-language [Nor91]. mixed-precision [HS10]. mixed-radix [Kir02]. Mixing [Ein95]. mixtures [BSS92]. MK [Mar92]. MLD2P4 [DDF10]. MM5 [Mic97]. MM90 [Mic97]. MN [Ano94p]. MNDO [HKŠ^{+}97]. MNDO/M [HKŠ^{+}97]. Mobile [CGS93]. mobility [LZL11]. Mode [JSY^{+}20, Ber91b, KTP^{+}24]. Model [BDOS95a, BDOS95b, DGL91b, DDHD90, DM90, GK06, GOBG^{+}94, Guo01, HCLJ03, MKS94, NOL97, PMM93, PWD93, SWBO93, AFBN93, Bra94a, BMV03, CK90, CZM93b, CMZ93a, CHM91, Coo94, DV02a, DV02b, FR94, FYR99, GS95, HBD^{+}93, HM93, KLV98, KB94, KS12, PFS^{+}04, PMM94, Pel93, PD96, SMG91, Sch93a, STY15, STY18, SS10, Sto93, SH97, Var97, VLLY92, Yan95, YRF02, ZZN94, DET12, DCHH88b, LJO05, Mic97, SH97]. Model-based [MKS94]. Modeling [Car90, CC95a, Chi91, FGRT00, FMW^{+}94, SS00, CMP02, KHRS95, Par94, RAX10, She91, Xu93]. Models [Bel11, BGW93, CMK00, Fos93, Gar91a, Gar91b, Gre93, LJO05, MHdL12, RHH96, SH07, SH09, Wri91, AC16, BLL^{+}96, CDF^{+}93, Duv92, GF95a, Kay90, SS09, Wri90a, YO95]. Modern [Bro92b, Bro92a, Cel96, HH18, HMR^{+}15, MRC11, MRC18, MRCB23, Mol12, NLE^{+}20, PCS98, RMX12, AJJF14, HH14, KTP^{+}24, KMJ^{+}23, NDSG07, dL12]. Modes [GGW96]. Modification [Fuj95, SW91]. Modified [BM99, TZW^{+}10, GST04b, Par86, GST04a]. Modular [Bro90a, FC95, GBC92, HBG^{+}06, Hor92, UNF^{+}08, FC92]. Module [Cou97, Kea95a, Kea96a, Kea96b, Kod11, Mit97, Sch99, Tre97, BCS01, GST12, How91, Jon09, Rei95c, SH91, Sch03, Wal93b]. Modules [Kra94, DLLR96, van90a]. Modulo [EDA96, Huf93]. Molecular [BL91, DCR99a, DCR99b, KLA95, NSWP90, SGMS97, BCS01, EFG^{+}05, FTD91, KSYE00, Nar95, SWO92]. Molecule [XH90]. molecules [NG93, PZY16]. momenta [AC16]. Monitoring [Yan94a]. Monitors [BL94]. monohull [Mil92]. monolithic [MIN^{+}95]. Monomial [MKFB92]. monotonic [Dem03, Dem07]. Monte [BD93, CHM91, Heu90, MMV95]. Montreal [CGS94, Lev95a]. Morton [LW07]. Morton-hybrid [LW07]. Morton-order [LW07]. Moscow [KVK92]. mosquito [RKMJ92]. mostly [JH86]. motifs [Ste91]. motion [Lie94a, Lie94b]. motions [CZ10]. Moving [Pug94]. MP [Bre78, Bre79, Car92, Nag90, PAK^{+}90, VSH91, Vaj92]. MPAS [KTP^{+}24]. MPAS-ocean [KTP^{+}24]. MPI [Ano94j, Ano94k, BW12, BF01, CFMR95, Coo95, DZ98, GBR15, HZ99, LZ97, LCC^{+}03, OP98b, RFRH96, SM17, SM02b, SM03, SC19, TAH^{+}01, TOC18, WO96, YBMCB14]. MPI-Based [SM03, OP98b]. MPI-CHECK [LCC^{+}03]. MPI-interoperable [YBMCB14]. MPP [AGS92, DOSW96, PMM93]. MS [HT91, WRL90]. MS-DOS [HT91, WRL90]. MSSM [DGS08, DKM07, MDM05]. MTIEU1 [Shi93b]. MTIEU2 [Shi93b]. Müllges [Ano97a, TDMC97, Hop97]. mullihak [KMmYsK92]. Multi [KHRS95, Mit02, SH09, BSCV95, PFS^{+}04]. multi-dimensional [BSCV95]. multi-layer [PFS^{+}04]. Multi-level [Mit02]. Multi-phase [KHRS95]. Multi-unidimensional [SH09]. multibody [Lie94a, Lie94b, Sre92]. Multibox [Dya94, Dya95]. multicell [Mir90].

Multicomplex [AMGM20]. Multicomponent [PTS92]. Multicomputer [HRW^{+}98, KW94, SSOG93]. Multicomputers [KMR96, RSB97, GB92, WW95, WI94, Yan94a]. Multiconference [Ten93]. multicore [HL08]. multicriterion [Osy92]. Multidimensional [SW94, RBS92]. Multidisciplinary [MMT09, CMVZ94]. Multidual [AMGM20]. Multiflow [LFK^{+}93, SS93]. Multigrid [NRK98, KLM00]. Multiinput [MR95a]. multilayer [AG95b]. multilayered [FR94]. Multilevel [DDF10, Sal06]. Multilocus [CHL94]. multiloop [Mas92b]. MULTINOR [Tho90]. multiphoton [TYJ92]. Multiphysics [LJO05, RAX10]. Multiple [Bre78, Bre79, BHY80, CC93, MBFC99, Mor15, Smi91, Smi98, Smi11, SB01, vV90, Las97, RBD^{+}10, RBD^{+}11, Smi01]. multiple-phase [RBD^{+}10, RBD^{+}11]. Multiple-Precision [Bre78, Bre79, Smi91, Smi98, Smi11]. Multiplication [DS94, Hig90b, Han92, Pet91]. multiplications [DN09]. Multiprecision [Bai92, Bai93a, Bai93b, Bai94, Bai95]. Multiprocessing [NV94]. Multiprocessor [BP92, PWD93, SR04, EO91, KHC92, KLN90, Phi91a, Wag94]. MultiProcessors [BMV03, CMZ91, PMBH93, Sch93b, AW94, GHSJ94, KSZ90, SPS^{+}91, SMH91, TMD13, WYJ99]. Multirate [EL97]. Multitasking [Vai93, Nag90]. Multithreading [Nag01]. Multivariate [Dre92, Dre93, HP95a, KTMB02, Som98, TZW^{+}10, Tho90, EKB92, vH10]. MultiZ [AMGM20]. MUMPS [MFK09]. Munich [GH94a, GH94b, GH94c]. muon [RM90]. muon-catalyzed [RM90]. Mutation [KO91]. Mutation-based [KO91]. mutual [Szy07]. MVS [Int90h, Int90n, IBM91b, Int91f, LHHJ91].

N [BSS92, FK95, NRS92, vK94, DCR99a, ADB94, MB95]. N-body [ADB94, MB95]. N.A.Software [Bee01c]. N.S [Mol12]. N1122 [W^{+}95, ANS95]. NAG [KLM91, Bee01g, Bee01f, Bee01e, BKR^{+}91, For95, Mai91, Num90a, Num90b, Num90c, Num91a, Num91b, Num91c, Num93a, Num93b, Numxx]. NAGWare [NR06, Ola92]. NAL [MFI^{+}94]. Name [Cip00]. NAMELIST [Nai17]. NAS [AHOK02, NNON02, Sai95, WYJ99]. NASA [GK06, WBS97]. NASA/ [GK06]. National [Ame90b, Ame92]. Natural [NRS92, Cok93b, Nie92]. Nature [Gal91]. Navier [Fat94, RRX^{+}08]. NBI [FTD91]. NC [Agr95]. NCAR [Mic97]. NDA [LV01]. NDP [Ano97b]. near [CCW04]. Nearly [Dec93]. Need [NLE^{+}20, VCV97b, VCV97a]. Negative [Tho13]. Nekbone [GML^{+}16]. nested [PPW94]. Nestor [SD99]. nests [GF95b]. Netherlands [Ano93n, Ano93q, DSZ94]. nets [Lai92a, Lai92b]. Network [AAN^{+}93, Coe94b, Oku95, TT92, BDG^{+}94, BID95, DLW^{+}18, MIN^{+}95, MC96]. Networking [ACM97, ACM98, GH94a, GH94b, GH94c, HS95]. Networks [HHK94, VBB18, Fre92]. Neural [Fre92]. Neutral [GOBG^{+}94]. News [Ano93a, Ano93l, Ano97d, Ano97c, Bra97d, Bra97c, Coc03, Mar98, Ola93]. Newton [OM92, SF92]. next [RN07, DET12]. Next-to-Minimal [DET12]. NeXTSTEP [Sil93]. Nickel [KLA95]. Nicosia [PRS99]. Nineties [Rys95]. Ninety [ABMS94, Gla92b]. nitrogen [NG93]. NLEdit [Cur94]. NMR [SSLG91]. NMSDECAY [DET12]. NMSPEC [EH07a]. NMSSM [EH07a]. No [Dya95, GG95]. nodes [CK91, SG95]. non [Cra95, KB94, MC96]. non-azeotropic [Cra95]. non-gridded [MC96]. non-uniform [KB94]. noncompact [Cah90]. nonconcentric [NVC96]. nondifferentiable [LV01]. nonempirical [HKŠ^{+}97]. Noninteger [Shi93b]. Nonlinear [BB91, Buc94a, Buc94b, BGW93, CGT92b, Kea95b, TT92, WW90, Yam95, CM94, CGT92a, GT03, GOT03b, GT07, HBG^{+}05, Hop98, IDVV97, Joy92, Nat92, Ren03]. NONMEM [VKB93]. Nonnegative [Dem95, Kod08]. Nonnormal [Rhe93, Wal92]. nonperturbative [NJ94b]. nonsinglet [KKK95]. Nonstiff [Cas89a, Hig91]. Nonsymmetric [BS92a, BS92b, BS97]. nonuniform [Gou93]. Norm [Hig90a, HH18]. Normal [Dre92, Dre93, Lev92, Som98, KK90, KDG99, MZM94, YK90]. normalisé [DV91]. Normality [HP95a, Tho90]. normalization [LP92, LP93]. normative [MSB92]. NORMUL [HP95a]. Notation [CF95, McC96, Num05]. Note [GS01b, Hew01, KRY90, KK90, WMCU97]. Notes [Ano93l, Rocxx, FH92, WC92, WW93, Dig90a, SMG91]. November [ACM96b, ACM97, ACM98, ACM99, ACM03, BGG^{+}94, Fox91a, HK93a, HK95, IEE90a, IEE91, IEE92d, IEE93d, IEE94b, IEE94f, IEE02]. NSWC [DG99]. NT [Mic93b, TAH^{+}01, Vai93]. Nuclear [BPG94, KSW93, Hor23]. nuclear-decay [Hor23]. nuclei [SH97, Taq16]. nucleic [TRS91]. Null [Gil94, Rib02]. NUMA [AW94, LP92, LP93]. Number [Aki99, HD93, Lev92, Ric06, Wal91b, BS13, Coh90, CBTL97, DW03, FSV90, Hen94, Hen95, Jam90, Jam94, Jam96, MZT90, MS00a, MS00b, Ume91, Wol92, Wri90b]. Numbers [AMGM20, BS13, Str05, YB13]. Numeric [BKR^{+}91, Mer92a, GDS94, LP05]. Numerical [Ano92e, AZ90, ABB^{+}94, BK95, BGZ94, BFKS93b, Bin96, Bor91b, Bos19, Cap98, DLM99b, DLM99a, EGKU99, EMR93, EMU96, Ett92, For95, Gao06, GGLM88, GL90, Goo90c, Goo90d, HH14, IFI95, Irv91, JSW93, Kle93, Kon94, Lev98, Loz98, Mac91a, MFI^{+}94, Num92, NPB92, NL95a, Pap93, PTS92, PTVF92, Pre92a, Pre92b, Pre92c, Pre92d, Pre93e, Pre93f, Pre93d, Pre94b, PTM96, PTV96, RS92a, RS92b, TDMC97, Ueb97, VTP92, V^{+}93, Vet93, VBA95, Wie99, AP90, Azi23, BHLT09, BFKS93a, Boi97, Bor91a, CM94, CRS90, Du 97, GC03, Hor96, IDVV97, KMJ^{+}23, KM99, KKK95, KL92, LFG00, MMG98, Mor81, Nak90, NSU20, Num05, NL95b, Shi93a, Adl93, Ede90, EMUP98, Gar93, Hop97, Kon94, AFKL04, Spe93, Sun92b, TDMC97, Tha93, Wu93, AB93, Ano97a, EMUP98, Hop97, Yan94b]. Numerically [Bel90a, Bel90b, GR92]. Numerics [Ano97b, Cse99, AAS93]. Numerik [EMR93]. Numerik-Algorithmen [EMR93]. NumPy [AJJF14]. NUMVEC [Sv90, Sch90]. nuyu [Mal91]. NWT [MFI^{+}94]. NY [PBG^{+}95, SS96]. Nyhoff [Rub93]. nykh [AZ90].

O [BSS92, Lar93, BLW02, Coe94a, LG93, LHHJ91, SW94]. Oberammergau [BPG94]. Oberon [Mös95]. Oberon-2 [Mös95]. Object [AC97a, AC97b, Aki99, DG08, FB12, GRE99, KLM00, MD97, MMT09, Moo95a, Moo95b, NCMF15, NL19, RL91, Rotxx, Sch93c, Shi93a, Smi92, Smi93a, WBS97, Wil93, Abs91, Bod94, CSC^{+}97, CH98, LFG00, NDS96, OT93, QRH00, Sal06, She91, Wam90a, Wam90b, Dub97]. Object-Based [Dub97]. Object-Oriented [AC97a, AC97b, FB12, MD97, MMT09, Moo95b, NCMF15, NL19, WBS97, Wil93, DG08, KLM00, Moo95a, Shi93a, Smi92, Abs91, CSC^{+}97, CH98, QRH00, Sal06, She91, Wam90a, Wam90b]. objective [LN91]. Objects [RMX12, AFAS99]. oblique [DH95, YB92]. Observations [Mei96, PWD93]. obtain [Gep90]. Obtaining [LH92]. obyknovennykh [AZ90]. occam [AHZ90, Mar92, SAC^{+}92]. occupational [ZMR^{+}91]. Ocean [BDOS95a, BDOS95b, PWD93, KTP^{+}24, NJ94b]. Oceans [IEE94c, IEE94c]. October [Ano93n, Ano94a, Ano94i, Ano94p, AH92, BPG94, BT01, BGG^{+}94, Fer92, FK95, GGK^{+}93, IEE94e, IEE95b, IFI95, PRS99, Sch93a, Sch93b, USE94, Vol93]. ODBC [Ano96b]. ODE [Ano97c, Bra97c, BG94, Enr95, Rei93]. ODRPACK95 [ZBW07]. oeuvre [LMG95]. off [JCL10]. offer [Shi98]. offering [Int90e]. Offshore [CKMU94]. Ohio [Hua96]. Oil [Ano90a, KR94, KR95, BD93]. oil-field [BD93]. old [CC93]. Oldenburg [AH92]. Older [Pra90, Sal92]. O’Leary [Tay86]. on-the-fly [HKMC90]. ONC [RS93]. one [CRS90, DDH17, KS12, MCA17, SRM90, Car91a]. one-dimensional [CRS90, KS12, SRM90]. one-loop [DDH17]. one-valence [MCA17]. only [CD03, Gao05]. Ontario [BGG^{+}94, GGK^{+}93, HDR03]. onto [BN93, PWD93, WMCU97]. OoLALA [LFG00]. OOPL [dVdVI97]. OOPP [Wam90a, Wam90b]. Open [UNF^{+}08, Wri90a, Wri91, Dig93a, Dig93b, Dig93c]. Open-Source [UNF^{+}08]. OpenACC [GML^{+}16, NLVE15]. OpenAD [UNF^{+}08]. OpenAD/F [UNF^{+}08]. OpenMP [Ano97c, BF01, Bra97c, Bri00, BMV03, BC19, CM98, KaM10, KLM^{+}19, MVK25, TBG^{+}02, Wal02a, YSVM^{+}16, YSMA^{+}17, YSMBA23]. OpenSHMEM [NLE^{+}20]. Operating [Ano94a, Ing90a, Ing90b, LHHJ91]. Operation [HLJ95, HLJ01, Blu91]. Operational [RHH96, Cho91]. Operations [AS97, BGMZ92, HLJ98, JSY^{+}20, KMR96, Pri10, Bre92, KVK92, Off94, XWK95]. operator [STY15, STY18]. Operators [HL94, Kir93, Kir98]. OPLS [HKŠ^{+}97]. Oppenheimer [BG93]. Oppermann [Kro90]. OPS5 [Gro91]. optical [SS90]. Optimal [CA96, FJ92, Kra94, NH09, SV95, dSZP92, AM90, NSWP90, RBD^{+}10, RBD^{+}11]. optimisation [KHC92]. Optimization [AMC98, AMKS02, ABB^{+}94, Bou97, CGT92b, FJS97, Fah94, GPS99, Int92, IBM93, KLW93, LP19, Maaxx, MMT09, McC96, MGH81, NS92, NI03, OA02, Osy92, PAK^{+}90, SM03, SWH15, Sou91a, Sou91b, TT92, WW95, ZBLN97, ADH95, CCL04, CGT92a, Con90, CKT85, GT92a, GT94, GOT03b, HWS09, KM99, LV01, LMV09, MCAB^{+}02, MN11, RTY90, SMG91, Sav95, TRV96, ZT90, GT92b, BMR01]. Optimizations [HKT92b, HKT94, iSYS12, HKT91b, HKT91a, HKT91d]. Optimize [HLJ95, HLJ01, GKH^{+}92]. Optimized [AAC^{+}04, Sab92, SGMS97, BHS92]. Optimizing [Ano02, Ben99a, Ben00, BMN^{+}97, CL97, CT11, Das06, EGKU02, GS01a, JH86, JC93, KMR96, Rot93, Sab94, Sch91b, BGS82, CSS90a, CSS90b, CSS91, GMMM92, LM94, Tse93, WFW^{+}94]. Optimum [EDA96]. option [LD87]. optional [Das06]. Orange [ACM98]. orbit [BKRG22, KRG21]. Orbital [KLA95]. Orchestrating [GLS93]. Order [Cas89a, CC92a, CC95a, Fuj95, GM97, Hig91, Kod08, Kod11, KH13, Mac98, ONT95, Shi93b, SWM95, TIUG90, AMGM20, GST04a, IDVV97, KHC92, LW07, Ove91, Sar97, vH06, vH07, vH10]. Order-Restricted [KH13]. ordering [ADD04]. Orders [HMW91, GST04b, HMW93]. Ordinary [BG97, Cas89a, CC92a, Hig91, LS04, AZ90, GP92]. Oregon [ACM94b, ACM99, BGNP94, IEE93d]. Organization [Ano94n, Ano94o]. organizations [Cof93]. orientation [Mai90]. Oriented [AC97a, AC97b, Cam13, FB12, MD97, MMT09, Moo95b, NCMF15, NL19, Sch93c, WBS97, Wil93, Abs91, Aki99, CSC^{+}97, CH98, DG08, KLM00, LFG00, Moo95a, NDS96, QRH00, RL91, Sal06, She91, Shi93a, Smi92, Smi93a, Wam90a, Wam90b].

Origin2000 [Bri00]. Original [McJ17b]. Origins [Cro07]. Orlando [ACM91, ACM98]. Orthogonal [DPS02, Gou93, ZBW07]. Osaka [CKMU94]. OSCAR [HMKN91]. oscillating [EO94]. oscillator [MBGK11]. OSF [Sch93a]. OSIPE [CJPA94]. Other [BOPC05, GPHL90, PMM^{+}08, CB95]. Ottawa [BT01, HDR03]. Our [Pif96, BN93]. out-cropping [CNP91]. Out-of-Core [TBC94b, RS09b, TBC94a]. Outer [SF93]. Outlier [McB06]. Outline [MC94, MC95b, MC95a]. Output [And90, And92b, Lev95b, Lev97, Ngu91]. overlap [BBDR94, BBDR95, Jon92a, Jon92b]. overlapped [EJLC97]. overlapping [BDH^{+}05, CN94]. Overview [HKK^{+}91a, Hir91, HKK^{+}91b, HKK^{+}92, Koe92, KK94, DHP02, GR92, Hey94, IJCL96, Rei04, ZCP95, Zos93]. own [Cre03]. Ownership [JB01a, JB01b]. Oxford [Ano91a, Boi97, Gla92a]. Oxide [KLA95]. ozone [NG93].

P [Adl93, AB93, Gar93, Loz98, Yan94b, DOP^{+}92, O’K93, OPE^{+}95, PQ94]. P.C. [WNO94]. P03T [Fah94]. p4 [BL94]. PA [Sen03, ACM96b, Ano95a]. Pacific [Pre93b, Van94b]. Pacific-Sierra [Pre93b]. PACK [BR96]. PACK/UNPACK [BR96]. Package [BGKZ91, Bou97, Bre78, Bre79, BHY80, Cod93a, Cod93b, CGT92b, Cos97a, DDF10, DLM99b, DGL91b, GGLM88, GL90, GWL^{+}92, GP97, Kod08, LW95b, Maaxx, MSZ90, Pry99, SF92, Smi91, WD98, AI90, AR06, Bai05a, Bai05b, Bai05c, BB07, BGV94, BDH^{+}05, CGT92a, dCH94, Dem03, Dem06, Dem07, GMF18, GT03, GT07, GJU96, Hen94, Hop02, KSYE00, LW95a, LD90, Mac96a, NS11, Osy92, Ren96a, Ren09, SZ90, Tal94, Tor10, Tre91, Var97, WW14, Yan95, ZMR^{+}91, vH06, vH07, vH10, FGJB19]. Packages [Ano97d, Bra97d, EP87, GOT03b]. Packed [GWDL08, GWDL10]. PACKER [PBU95]. PACT [CGS94]. Padé [CJL97, MKC92]. PADRE2 [Kub91a, Kub91b, Kub91c]. Padua [CMV09]. pages [Ano96a, Ano97a, Ano98b, Ano99a, Ano99b, Ano03]. Palo [ACM01]. Pandore [AS92, AFMP95]. Panel [HP95b, BBF^{+}92, HCD^{+}98]. Pao [Tay99]. Paperback [Eme94]. Papers [Cse99, HR92, Knu03, Lap96, ACM93c, ACM94b, ACM95b, Ano93q, Bar92, IEE93a, AFKL04]. PARA [DW94]. Parabolic [BD91, GST06a, GST06b, WKM04, GST11]. Paradigm [Sug95, Wam90a, Wam90b]. Paradigms [CM98]. Parafrase [PGH^{+}90]. Parafrase-2 [PGH^{+}90]. Paragon [GAW96a, GAW96b, SZG95]. Parallel [PPP93, ACM93a, ACG^{+}94, AMC98, Agr95, AHZ90, AH94, Ano93m, Ano94p, Ano95c, AHOK02, ABB^{+}94, BK95, BR96, BBG^{+}95, Ben95, BBG^{+}93, BMMN94, Bra00, BLW02, BV94, CCL01, CZM94b, CGS93, CGL^{+}95b, CH94, CA92, CGS94, CL93, CL94, dCH94, Cyb91, CHKM93, DDF10, DDP94, DL97b, Dec93, DG94, DFRR91, FBZ92, Fat94, For95, FC95, Fox94, FMW^{+}94, Fuj95, Guo01, GS97, GPHL90, HHLS90, HKT92a, HKTW94, HK93b, HK93a, HK95, HIM91, HMKN91, HKMC90, HHK94, HCLJ03, IEE92c, IEE93c, IEE94a, IEE94e, IEE95b, IEE96, Ken94b, KNS95b, Kon00, KKMP95b, KP91, Kum94, LG93, LJO05, Lev94, Lev98, LZ97, LH92, LMR^{+}97, Meh93b, Mit02, Mra94, NS92, Nik93, Num05, Oed93, PHD^{+}95, PTM96, PTV96, RA90, SWBO93, Sab95, Sar91, SZM98, SWW90, SSC00, SM03, Sie94a, Sie94b, SWH15]. Parallel [Ste95b, Sze90, TR96, Vol93, XH90, YGS^{+}94, YYM93, dSL98, vDSP96, AES^{+}96, ALS91, AH90, All93, ASM^{+}94, AFMP95, ABC^{+}96, AH91, ADB94, BBB^{+}94, BKT91, Ban93, BGNP94, BS13, BB02, BDOS95a, BDOS95b, Bod94, BRH90, BBDR94, BBDR95, BID95, BxCW01, BL94, Cel96, CCL04, CMZ94b, Cha93, CGL^{+}93, CC92b, CC94, CCW04, CN94, CEF^{+}95, CWB92, CWB94, CDGM96, Coo95, CFPS94, CDH^{+}94, CK91, DR94a, DSZ94, DH84, DT94, Duv92, FC92, Fos95, Ger98a, Ger98b, GLS93, GS95, HMPT94, HHK^{+}93, HAM95b, HGG93, HWS09, HZ94, HZ99, HKT91c, Hua96, IBM91d, IEE95a, IEE97, JC93, Jor90a, Jor90b, KKS^{+}95, KMR^{+}97, Kas93, KY98a, KY98b, KMT91, KT94, Ken94a, KNS95a, KY94, KB94, KKMP95a, KP93, LPA95, LMJC96, Loz98, LSW92, Luc92, MCA17, McB91, Meh93a, Mic97]. parallel [MMRS92, MKS94, MC96, MR96b, Nic91, NNON02, NDS96, NR98a, NR98b, O’K93, OPE^{+}95, Off94, OH90, PSG03, PW84, PQ94, Per94, PD96, PBG^{+}95, QRH00, RBS92, RBS93a, RBS93b, RL91, Sal06, SSW91, ST95, Sta94, SMSW06, SV95, Sul91, SZ91, TBC94a, TMD13, UZCZ95, UZCZ96, UHP91, Utt90, WTW90, Wol92, YWS^{+}94, YO95, ZCP95, ABB^{+}91, BLLWW95, BBG^{+}93, DW94, HK93a, NLE^{+}20]. Parallelisation [HBD^{+}93, Jéz93, EJLC97, IJCL96, LMJC96, Lov92]. parallelise [PFS^{+}04]. Parallelism [BEH^{+}94, CFK^{+}94, CT90, Fos94, FKKC96, GOS94, HK91, LE98, LR94, OP98a, RSB97, ADH95, BK89, Bar94, Bod94, CMVZ94, Fu90, GP92, Mar93, MBFC99, OPP00, PPW94, PQ94, RLS20, SSOG93]. Parallelisme [Cha94a]. Parallelization [BC01, BB96, BCC^{+}96a, BCC^{+}96b, Che92, DCR99b, Eli98, EB98, FBZ92, Fah94, Ger94a, Harxx, IK96, Kik93, KO90, LP98a, RA90, SAS90, SF93, TLS90, TLS91, WYJ99, BCS01, BW12, BCC^{+}97a, BCC^{+}97b, DDcMR96, Kin92, LP99, RP95, STVS91, vPMF92]. Parallelize [RRM^{+}15]. Parallelizer [Bou96, KLS94b]. Parallelizing [AAN^{+}93, Adv98, ASS95, DPR94, GB95, KLN90, RHH96, RY99, SR95, ADG96, BC97, DFRR91, GF95b, GB92, HDH^{+}94, HDH^{+}95, LYZ90, Mil91, PGH^{+}90, SLY90a, SLY90b, WFW^{+}94]. Parameter [Gar91a, Gar91b, Que00, EK01, EC13]. Parameters [BGW93, Con92, Das06, Nar95, PG10, Smi93b, YO95]. Parametric [EFP07]. ParaScope [HHK^{+}93, KMT91]. Parent [RMX12]. Pareto [Air04]. Pariser [KS12, SS10]. Parisi [KKK95, KL92]. PARLE [HMPT94]. parler [Ain90, Ain91, Ain93]. Parlett [Koi09]. PARMACS [CHHW94, HZ99]. Parr [KS12, SS10]. Parsers [Dya94, Dya95]. Parsing [Bee90]. PARST [Nar95]. PARST95 [Nar95]. Part [Ame97b, Ame97a, Ano91b, Ano93g, Ano94e, Ano94f, App91, Bee96b, Bee96c, Bee97, SG93a, SG93b, SG93d, SG93c, FYR99, Goo90e, Hig94a, HFMS95, IEC90, IEC94, IEC97, IEC99, Ins92, IEE92a, IEE93b, ISO90, Int97a, Int00, ISO04a, ISO04b, ISO10, YRF02, BCC^{+}97b, BCC^{+}96b, BCC^{+}97a, Ein95, Goo90f, Hig94b, Hig94c, Hig94d, ISO94, Int97b, Int99]. Partial [PEP92, BGZ94, BF93a, BF93b, BC94, DJ92, MD97, Nak95a, SWM95, YKK96, HIS91, KY98a, KY98b, LS04]. partial-static [KY98a, KY98b]. partially [CDGM96, OH90]. Particle [ADHF96, CLiN^{+}02, DET12, DKM07, MBGK11, Taq16, XHY^{+}24]. Particle-in-Cell [ADHF96, CLiN^{+}02, XHY^{+}24]. particles [AIS^{+}97, Coh90, MDM05]. Partioning [O’B93]. partir [AL92]. partition [Nie92, PW93]. partitioned [WSW00]. Partitioning [PSC93b, GB92, LPA95]. PASCAL [Dot93, Cho91, Lem93a, Lem93b, Lem93d, Lem93c, MHT96, RS92a, RS92b, RR92, Wic89, WR93]. pass [KY94]. Passing [AFAS99, Ano94j, BGLP94, FKKC96, vDSP96, Ano94k, BDOS95a, BDOS95b, CHHW94, CA92, DOSW96, GS95]. Pattern [RD92, Wea94]. pattern-matching [Wea94]. Patterns [HMR^{+}15, MDD94, Agt94, Bec91, DG08, DRST03, RAX10, TRS91]. PC [DeT90, Ame90a, Bau93, CLiN^{+}02, Chi91, EN96, FH92, Ham93, Ham95a, KP92, Kul95, Lav91, PS96, RM90, Sal92, SRM90, WC92, Bod94]. PC-APD [KP92]. PC-based [Sal92]. PC-Fortran [RM90]. PCOMP [DLS95, LS90a, LS90b, LS00]. PCs [Bra90, FL91, NV94]. PCTE [HZ94]. PDE [GBR15, HRW^{+}98, OH90, LD90]. PDEQSOL [HIS91]. PDEs [BD91, WKM04]. PDS [HKM98]. Pedagogic [And92b]. Peek [Smi00]. Peerless [Kri86]. pelitic [MSB92]. PELLPACK [HRW^{+}98]. Penn [Mic97]. Pennsylvania [ACM96a]. penultimate [Met99c, Met99d]. Percentiles [AS93]. Perfect [Cyb91, VSH91]. Performance [ACM97, ACM98, Adv98, AMC01, ADHF96, ACIK97, AH91, AH94, AOL94a, Ano93b, Ano93c, Ano93e, Ano93g, Ano93h, Ano93f, Ano93k, Ano94e, Ano94f, Ano94h, Ano94m, Ano94o, AGG^{+}97, BGS94a, Bee96a, Bel90a, Bel90b, BBZ94, BZ99, Ben99b, BB02, BEH^{+}94, Bou95, BCF^{+}93c, BCF^{+}94a, BMN^{+}95, BMN^{+}97, Bra94d, BCC^{+}96a, BCC^{+}96b, BCC^{+}97a, BCC^{+}97b, BGMZ92, BH90, CLiN^{+}02, CMT01, CC95a, CZM93b, CMZ95, CCW04, CKZ93, Cod90b, Cou97, CL93, CL94, DDcMR96, DL97a, DL97b, DS97, DZ98, DCR99a, Din99, Don90, Don91, DV92, Dow93, EGKU02, Eli98, FBZ92, FGRT00, FJSD96, FXAC94, Fos94, Fox91a, FGG09, GS01a, GH94a, Ger94b, GKKL19, GOS94, Hig92, HM96, Han98, HBB^{+}95, Hat94, HMR^{+}15, HF95, HKŠ^{+}97, HIM91, HJT97, HJJ^{+}00, IEE94d, IFI95, KMR^{+}97, KaM10, Ken94b, KK95a, KK01]. Performance [KS02, KKZ11, KMBK96, KMS^{+}95, KOM93, KOM94, Koe92, KLS^{+}94a, KGV97, KK94, KZ94a, KZ94b, Lin93, LMMW96, Lov94, MB95, MSP^{+}22, McC95, MSC96, MMY95a, Meh93a, Meh93b, Meh94, MVZ98b, MVZ98a, MZ00, MZ01, MH95, MCH96, Met95, MMV95, MMY95b, MR95b, MA18, NOL97, Off98, PFS^{+}04, PHHF94b, PH96, Paz96, Per93, PMBH93, Pre93c, PA94, Rag95, RMCKB97, Sab95, SF02, SZM98, Sch96a, Sch97, SNMC93, SIOS02, SM02b, SM03, Ste93, SSG94, Ten93, Tho93, TBG^{+}02, Tse97, UMM94, Wag94, WD98, Yan94a, YGS^{+}94, YFH97, Zim02, dSL98, van94a, AOL94b, Ano93q, Ano94d, AHJS90, BCM^{+}93, Bli90, BCF^{+}93b, Bre92, BMV03, Car91b, Car92, CK90, CMZ93a, CMZ94a, CZ90, CDF^{+}93, Dig90b, Dig93b, DS02, Don95, Duv92, Eme94, FGGL05, GH94b, God93, GML^{+}16, HS95]. performance [HP95b, IJCL96, KLV98, KKS^{+}95, KT00, KC94, Lov93, LSW92, LHHJ91, MDV07, Phi91a, SM02a, Sal92, SZG95, SSG97, Zos93, BFHH94, Bra94c, CZM94a, GH94c, SSH08]. performance-prediction [BMV03]. Performances [DCR99b]. performing [NJ94b, WJ94]. PerfVisS [KC94]. period [FSV90]. periodic [KS12]. Perl [Wes96]. PERM [LH92]. permeability [Ude91]. Permuted [LH92]. Persistent [Kry94]. Personal [Mei96, TT93, Blu91, Lah90]. Perspective [Fox94, Pap93, Smi95a]. Perspectives [Wil93]. perturbation [GL10]. PESC [IEE92b]. peta [Zim07]. peta-scale [Zim07]. Peter [Cof93]. PETSc [HKM98]. Pfortran [BCS00, BCS01, BC01]. PGHPF [BMN^{+}97, Sch94]. pH [LHW01]. Phase [DD97, WW90, Cok91, KHRS95, LP90, RBD^{+}10, RBD^{+}11]. Phase-Change [WW90]. phi [SG95]. PHIGS [IEC90, ISO90, Ame97a, Lan93a]. Philadelphia [ACM96a, Sen03]. Philosophy [NRS92]. Phoenix [ACM03]. Phonetics [Ano96b]. phonon [ZZSW19]. photograph [DH95]. phq [ZZSW19]. Phys [Jam96]. Physical [GDS94, NRS92, WTW90]. Physics [DeV94, Gal91, KM90, Ano03, BPG94, XHY^{+}24]. Pi [BH92]. PIC [BMV03]. pictures [SHCP91]. piecewise [Dem03, Dem07]. pien [yKxx].

PIM [dCH94]. Pipeline [Che92, SR04]. Pipelined [BD96, TLS91, TLS90]. Pit [Wri91, Wri90a]. Pitaevskii [BKRG22, KRG21, KYSV^{+}15, MA09, TS06b, YSVM^{+}16, YSMA^{+}17, YSMBA23]. Pittsburgh [ACM96b, Ano95a]. PL [Sal92]. PL/1 [Sal92]. placement [vK94, vHK00]. placing [SG95]. Planar [AJ98, SZ90, Tip91]. planarization [RR99]. Plane [RB99, Raj95, YK90, YB92]. planning [Gro91, MKF95]. Plans [Ano94o, SZAB98, WSL94]. Plant [Rit90]. plasma [Eli98]. Plasmadynamics [Ano94l]. plate [Dot93]. Plates [Cap98]. Platforms [HRW^{+}98]. plots [GF95a]. Plotting [Pau93, Blu91, CM92, Ngu91]. pluribus [dVdVI97]. Plus [Ano96b, Cen91, MJR93, Las97, RD92, Yan95]. PM [MB95]. PM3 [HKŠ^{+}97]. Point [FBWR95, Mra94, Rei96a, RD92, Smi91, TOML04, VCV97b, Agt94, GT92a, GT94, IEC98a, ISO00, Smi01, Spe94, Ume91, VCV97a, Wic89, GT92b, Int98a]. point-charge [Spe94]. Pointer [LR91, MHT96]. Pointer-induced [LR91]. pointers [AZ98, MKS^{+}96, MHT96]. Points [CMV09, MKFB92, Las97, MNZ90, Yu01]. Poisson [Air04, Fuj95]. polar [CKMU94]. Polaris [FWH^{+}94, Wea94]. polarized [AIS^{+}97]. Politically [BBF^{+}92]. Pollution [SS00]. POLRAD [AIS^{+}97]. POLSYS GLP [SMSW06]. POLSYS PLP [WSW00]. POLTEV [HKŠ^{+}97]. polygon [Deu90]. polylogarithms [NSU20, BD14]. polymers [NSWP90]. polymorphism [DNS98, DN04]. Polynomial [BD91, GP97, MP93, Aki96, Bin96, DV00, GP94, KP93, SMSW06, WSW00, XWK95]. Polytomous [Gre93]. POPL [ACM91, ACM94b, ACM95b]. Pople [KS12, SS10]. Population [CHL94, WS94, FHE95]. porosity [Tur93]. Porous [PTS92]. Portability [BEH97, DB93a, KaM10, She92]. Portable [Amo90, ADG96, BK95, Bru96a, CHHW94, CH94, Cod93a, Cod93b, CDH^{+}94, Dec93, DB93b, DW03, FW90, FHS78, Fox79, HD93, IEE90b, KDKSH92, KDDH94, KKMP95b, KP91, dLJEB95, Mit97, NLBB23, RHH96, Sta94, WW92, Wei95, YBMCB14, All93, AFMP95, BRH90, Bru96b, CEF^{+}95, HZ94, Jor90a, Jor90b, KN95, KMR^{+}97, KKMP95a, Mar92, Rap94, RL91, Wal93b, Wol92]. Portage [Pic94]. portfolio [AS92]. Porting [Bau93, BP92, Bra90, KM97, MWO95, NOL97, PWD93, SN94, Wri99, AH90, Sai95]. Portland [ACM94b, ACM99, BGNP94, IEE93d, Sch94]. Ports [PS96]. positive [GST04a, GST04b]. POSIX [Ins91b, IEE92a, Ins91a, HBG01, HBG02, Ins92, IEE93b]. Possible [HMW91, HMW93]. possibly [ZT90]. poster [Sch93b]. PostScript [Ngu91]. potential [BG93, FR94]. potentials [HKŠ^{+}97, HMT90, PS08]. Potts [CHM91]. Powder [MDD94]. Power [BEH97, IEE92b, Rhe93, Sai95, Ash81, HT91, How91, Lib90a, Lib90b, MIN^{+}95, SH91]. powered [Cre90a]. Powerful [DB93b, Pri93]. PowerMac [Hun00]. Powerstation [Ano93d, PS96, Mic93b, Mic93a]. pp [AB93, Ano90b, Ano91a]. PPOPP [ACM93a, Ano95c]. PPPE [CDH^{+}94]. PPRPA [Taq16]. PPTran [KMBK96]. Practical [KOM94, KK95b, Din99, Ede90, KOM93]. Practice [PPP93, ACM93a, Ano95c, KVK92]. Practitioners [Tho97a, Ano98a]. PRASCH [Gre93]. pratique [Lig91b, Lig93]. Pre [AC17, BR98, Ola95]. Pre-evaluation [BR98]. Pre-processor [AC17, Ola95]. Precise [MP93, SRH96]. Precision [Bre78, Bre79, BHY80, LH92, Ric06, Sch99, Smi91, Smi98, Smi11, HS10, KO94, Mer92a, Sch03, Smi01, Wic89]. Precompiler [Kub91a, Kub91b, Kub91c].

Precompilers [Sou91a, Sou91b]. Preconditioners [DDF10, Sal06]. preconditioning [CH98, MN01]. predict [CK90]. Predictable [Ano93b, VCV97b, VCV97a]. Prediction [Alo04, BL93a, CL94, FBZ92, MA18, PH96, SWBO93, BL93b, BMV03, CGL^{+}95a, CDF^{+}93, Mil92, RTY90]. predictor [vV90]. Predictors [van90b]. predicts [Kut92]. preface [Mal91]. preferred [Mai90]. Preliminary [BFHH94, HKT93a, HKT93b]. PREMER [VBB18]. Prentice [McC95]. prep [EKB92]. preparatory [EKB92]. PREPARE [BBZ94, BSCV95, Vee94]. Prepreprocessor [RP12]. Preprints [HOP93]. Preprocessor [Hor91a, Hor91b, RY99, GMMM92]. Preprocessors [LHH^{+}91]. PREQN [MN01]. presentations [Sch93b]. presented [ACM93c, ACM94b, ACM95b, Cse99]. preservation [IEE94c]. Preserving [Cos97a, Cos97b, Ren04]. Press [Adl93, AB93, Ano91a, Ano98b, Ano03, Eme94, Gar93, Kon94, Loz98, Sch91b, Tay99, Tha93, Wu93, Yan94b]. pressure [Cok91, PBU95, Ude91]. Presto [Tal94]. pRETS [HGG93]. Pretty [Ola95, Jon09]. Pretty-printer [Ola95]. pretty-printing [Jon09]. Prettyprinting [Ash81]. Preview [Rub93]. previous [Gil01]. Price [Eme94, Gon01]. prilozheniem [Mal91]. primitive [CCJ93]. primitive-based [CCJ93]. Principles [ACM91, ACM93c, PPP93, ACM93a, ACM94b, ACM95b, Ano93b, Ano95c, Ano99a, Ano99b, AAK01, Gon01, LZL11]. printer [Ola95]. printing [Jon09]. prior [Kir02]. privatization [RP95]. Prize [DKMS91, STVS91]. Pro [Ano97b]. Probabilistic [dSZP92]. Probabilities [Air04]. Problem [Ano92c, Bro97, Edg92, Ein91, HRW^{+}98, KF90, KF92d, MR93b, MR95a, Ric95, Sab95, WR93, Cho91, CWB92, GDS94, LP05, LR91, TJ90]. Problem-Solving [WR93, Cho91, GDS94, TJ90]. Problems [BK95, BGKZ91, BBCR98, BG97, CV94, CT95, Cas89a, CC92a, DL97a, DL97b, DPS02, FJS97, FJ92, GM97, Hig91, HJT97, MT90, MC94, MC95b, Mit93, Nis95, PR91, PPR97, RFS98, SF92, Sou91a, Sou91b, TT92, UZCZ97, WW90, vKK^{+}93, AS92, Ano95d, BHLT09, BCC^{+}97b, Cas89b, CS14, Cre90b, FPR01, GT03, GT07, Has06, HIS91, Hop03, IDVV97, MC95a, Pri93, RBD^{+}10, RBD^{+}11, RPL96, RR99, SPM^{+}94, Tor10, ZT90, vKK92, von92]. Procedure [BBZ94, BDK91, Wal92, YO95, Phi91b, Phi92, VKB93]. Procedures [CZM93a, Hig94b, Hig94c, KS90, Rei95c, Som07]. Proceedings [ACM91, PEP92, ACM93b, ACM93a, ACM94a, ACM94c, ACM96b, ACM97, ACM98, ACM01, Agr95, Ano93m, Ano93n, BBG^{+}95, BGG^{+}94, ERS95, Fer92, FK95, GGK^{+}93, Glo91b, HS94b, HK93b, HDR03, IEE90a, IEE91, IEE92c, IEE92d, IEE93c, IEE93d, IEE94g, IEE94d, IEE94e, IEE95b, IEE02, Kar95, MS94, Sen03, Sie94a, Sie94b, Ten93, USE94, ACM95a, ACM96a, Ano94a, Ano94i, AH92, Ban93, BGNP94, BLT94, BPG94, Boi97, BV94, CGS94, DSZ94, Fri94, GH94a, GH94b, HMPT94, HAM95b, HS95, HK93a, HK95, HHK94, Hua96, IEE94a, IEE94b, IEE94c, IEE97, KRB^{+}90, Kum94, Lev95a, NBC92, PRS99, PBG^{+}95, Van95, Vol93, WN90, HS94a, IEE94f, IEE96, KSW93, DW94, GH94c, Sch93a]. Process [Cok95, Schxx, Av94, Ker90]. Processes [CF95, AFBN93, Lef93, Tal94]. Processing [Ame97a, Agr95, ABB^{+}91, BBG^{+}95, BBZ94, BSCV95, FBWR95, HMKN91, IEE92c, IEE93c, IEE96, Rei96a, Sie94a, Sie94b, SD92, YYM93, ASM^{+}94, BV94, DSZ94, HAM95b, IEC90, IEE95a, ISO90, JC93, Kas93, KY98a, KY98b, Kum94, Lin90, MKS94, Nic91, ST95, SD93, Wic89].

Processor [Hew01, Oed93, Rod90, Vaj92, AC17, HM92, HC08, Kro14, NIY^{+}94, Ola95, RBS92, VSH91, YYX^{+}07, KHS17]. Processors [DDP94, DD99, HK93b, HK93a, HK95, RA90, CK91, KMR^{+}97, LSW92, O’K93, OPE^{+}95, OH90, Sta94]. produced [FYR99, Kea92, YRF02]. producing [CCJ93]. Product [MSC96, SMSW06, WSW00]. Production [MA18]. Productivity [CP93, KaM10, Zim07]. Products [Ano96b, Ano97b, Ano97d, Bra97d, Ano97c, Bra97c]. professional [Pag95]. Professor [Tay86]. Profiler [Sze90]. profiles [CB95]. Program [PEP92, AS93, AMC98, AG95b, Ano90a, BS13, BD90, BP92, BH92, Bel11, Car90, Cok91, Cok93b, CL93, DM90, FL91, Ger94a, Gil91b, Gil91a, Gil94, Gre93, HFMS95, HP95a, HIM91, HK91, HMKN91, IEE92a, KP92, KTMB02, KKZG94, KKZG95, KS90, KKMP95b, KKH10, KH13, Kut92, LMK94, MDD94, MC92, Mit93, MHdL12, Nan93c, Nan93b, OE92, RH94, SD90, SB91, SFB92, SWM95, Sil01, Smi93b, Som98, SNJ^{+}92, Tea94, Tho90, Wal90, Wal92, WS94, Wea94, van90b, Agt94, AI90, Ame90a, AFBN93, BMO90, Bec91, BSS92, BRdAHK04, Blu91, BD93, Bra94a, BOPC05, CM92, CRS90, CNP91, Coh90, Cok93a, CA90, Con92, Cum90, Dan90, Car93, CB95, Dut94, EKB92, EFP07, FTD91, FR94, FHE95, Gep90, GF95a, Gho01, Gil01, GMHC92, HW95, HHCS95]. program [Heu90, HM93, HKMC90, Hor09, Hor23, Int90e, Ins92, IDVV97, Joy92, Kah01, KKK95, KS12, KKMP95a, KRY90, KK90, KSM95, KL92, Lar93, LN91, LZL11, dLJEB95, Lin90, LSZ92, Lop90, MH91, MB92, Mai90, MCA17, McG91, MSB92, MBGK11, Mil92, Mir90, MM02, NY91, NJ94a, NJ94b, Neu01, Nie92, PS08, PMHC92, PT93, PW93, Raj95, RKMJ92, Sar00, Sar17, Sat97, Sav95, SSG^{+}10, SSG^{+}18, SMB90, STY15, STY18, SSLG91, SRM90, SS10, Spe94, SWO92, Ste90, Ste91, Taq16, TS06b, Tsa01, Tur93, Unixx, Uni93, Ude91, Van22, Var97, WRL90, WHL95, WJ94, WHL92a, WHL92b, Wie99, WCN92, Xu93, Yan95, YH93, Yu01, ZE92, ZMR^{+}91, ZZN94]. program-package [AI90]. Programacion [Mer91]. Programmable [RY99]. programmation [Ain90, Ain91]. programme [II90]. Programmen [EMR93, Por90]. Programmer [Ame97a, BGA90, BGA94, BA95, BGA96, Del93, ES93a, ED99, ISO90, U.S01a, Con91, Hew92a, Hew92b, Sil92b, GMC96f, IEC90]. Programmers [Ano93b, Bro90b, Lem93a, Lem93b, Lem93c, Lou90, Manxx, Poh97, Ano95g, Gla92b, Lan93a, Lem93d, Pag95, RP93, Uni2 ]. programmes [RD91, Tro90]. Programmieren [Ano93o, Lan93b]. Programming [HOP93, ACM93c, PPP93, ACM93b, ACM93a, ACM94b, ACM95b, Ame90b, Ame97b, Ame92, ACG^{+}94, Ano94a, Ano95c, AO90a, AO90b, AO90c, BK95, Bee96b, Bee96c, Bee97, BGLP94, BF01, BCM99, BEH97, CMK00, CV94, CMZ92a, CMZ92b, Che92, Coc03, Cro91, SG93a, SG93b, SG93d, SG93c, DR94a, Del98, DFL92, DB93c, Dub97, DY99, Ein91, Ein95, Ell90, EPL94a, EPL94b, FC95, GGW96, GRE99, GWE^{+}05, HHLS90, HBG^{+}06, HKK^{+}91a, HKK^{+}91b, HKT92a, HKK^{+}92, HKTW94, Hol94, Hug96, IEC94, IEC97, IEC98a, IEC98b, IEC99, Int97a, ISO00, Int00, ISO04a, ISO04b, ISO10, Jon93, Ken91, Ken94b, Ker93c, Kin93, Kru90b, Kry94, KP91, Lor19, MD97, Mas92a, Mas93a, MC94, MC95b, Meh93b, Mil93, MS93b, Mös95, Nik93, PMM93, Pri10, Rib92, RR93, Rod90, SZM98, ST95, Smi94, Smi95b, ST90]. Programming [Sun05, Tal91, Tem96, Tre95, VCV97b, Wal02a, WR93, ACM91, AC92, AES^{+}96, Aki99, Alt90, ASM^{+}94, Ano99a, Ano99b, AAK01, BN93, BB02, BKK94, Bod94, BRH90, BL94, CHHW94, CB94, CSC^{+}97, Cas14, Cho91, CEF^{+}95, CJPA94, CDH^{+}94, DS97, DeT90, Duv92, EPL95, EN96, FG93, FC92, Ger98a, Ger98b, Gon01, Gro91, GS95, HHK^{+}93, HCD^{+}98, HZ94, Hem94, HKT91c, Hir91, Hor96, Int90h, Int90m, Int90n, Int91b, Int91f, II91, Ken92a, KMT91, Ken94a, KF90, KF92d, KB94, Kon92, KVK92, Kug92, dLJEB95, Loh10, Mar07, MC95a, Meh93a, Met92b, MKS94, Nan93a, NSWP90, NDS96, NDSG07, Nor91, NR98a, NR98b, OJ09, PMM94, Per94, PD96, PZA93, iSYS12, Sch91a, Shi93a, Sny07, Szy07, TJ90, Vee94, VCV97a, Wam90a, Wam90b, ZA93, ISO94, Int97b, Int98a, Int98b, Int99]. programming-are [HCD^{+}98]. programmiruyushchikh [ES93b]. Programs [Air04, AG95a, AH94, Ano96c, AMKS02, AJ98, BGZ94, Bai92, Bai93a, Bai93b, Bel90a, Bel90b, BCC^{+}91a, BCC^{+}91b, BCC^{+}92, BKMC96, BF93a, BMMN94, Bra90, Bra97b, BZ94, Bru96a, CL97, CGS93, CGL^{+}95b, CH94, Che92, Cod90b, Cok95, CHL94, CRDO16, CC93, CT90, Dec93, DDHD90, DCHH88b, DB93b, EHJ^{+}91, EHJ^{+}93, EB98, FGL01, FBZ92, Fos17, Ger94b, GS97, Harxx, Hor91a, Hor91b, Hor92, HLJ95, HLJ01, HCLJ03, KN94, KM97, KNS95b, KMS^{+}95, KLS94b, LM90b, LP98a, MKFB92, McB06, Mer92b, Nat00, Pao99, Pra90, RS92a, Rit90, SWW90, SKP91, SSC00, SB01, SF93, Sze90, TBC94b, TCF94, TR96, Tho97a, WNO94, Wri91, XH90, Av94, AZ98, AH91, Ano98a, AC16, Bak91, BKRG22, BF93b, Bli90, Bru96b, CMP02, CCL04, Cha93, CGL^{+}93, Che90, CC98, CCJ93, CZ90, CI98]. programs [Cur94, DP99, DH95, DI90, EMR93, EO91, FSPC^{+}02, FCHE02, FYR99, Fos95, GP92, GS98, HCD^{+}98, HMT90, HMS^{+}95, JCL10, JC93, KRG21, KNS95a, KY94, KC94, KYSV^{+}15, KLM^{+}19, LM90a, LP99, Lov92, Luc92, LCC^{+}03, MMRS92, MSZ90, MA09, NJ94a, O’K93, OPE^{+}95, OM90, Pao01, PSPE94, Phi91b, Phi92, RM90, RS92b, RL91, RD91, Rot93, SFKL02, SS90, SSW91, SLY90a, SLY90b, SR95, SSG97, SZ91, Tay99, TBC94a, Tip91, UHP91, Utt90, Wam90a, Wam90b, Wri90a, YO95, YSVM^{+}16, YSMA^{+}17, YSMBA23, YRF02, Zah92, Sch91b]. Progress [CM94, Fei94]. progressif [Ain90, Ain91]. Project [Zim99, CDH^{+}94, Hey94, Sha94]. Projected [SH97]. projections [Cum90]. projects [Tea94]. Prolog [MWM90]. Prolog/Fortran [MWM90]. Promise [Ano93n]. Promises [Meh93b, Meh93a]. PROMOT [Ste91]. Proof [Nis95]. propagation [CCKT86, JCL10, MS93a, SRH96, TYJ92]. properties [KRY90, Lar93, Raj95, SS90, SMB90, ZE92, ZZSW19]. Proposal [DDH^{+}95, DDHW96b, She92, Sni92, Wic89]. Proposed [Ame87, Ame90b]. prospecting [PT93]. Prospects [MVZ98a]. protein [MMEH08, Ste91]. proteins [MMEH08]. Protocols [Ano93b]. Prototype [Tse97, YYM93, FLQZ97]. Prototyping [CC92b, Kea95b]. providing [GMF18]. Proximity [BD90]. PRPQ [IBM91d]. Pseudo [BT94, NIY^{+}94]. Pseudo-recursive [BT94]. pseudocode [Tro90]. Pseudorandom [Ham98, Jam90, Jam94, Jam96, MS00a, MS00b]. pseudoskin [AG95b]. pseudospectral [RBD^{+}10, RBD^{+}11]. pseudostress [LN91]. ptchg [Spe94]. Pthreads [AS14]. PTOOL [HHLS90]. PTRAN [Sar91]. PTX [iSYS12]. Publications [Bee96b, Bee96c, Bee96a, Bee97]. purpose [Lan90a, SS10]. PUT [HDH^{+}94]. PUT/GET [HDH^{+}94]. PVM [BID95, CDGM96, DP94, DPZ97, For95, LW95b, LW95a, MWO95, OPP00, RS93]. PWS [Ano90b]. PWS-Kent [Ano90b]. pyMDO [MMT09]. Python [DY99, GH18, GHN19, MSP^{+}22, MMEH08, SSH08]. PyTrilinos [SSH08].

QCD [AC16, DLLR96, KLN90, Sta94]. QCDF90 [DLLR96]. QCDMAPT [NS11]. QCDMAPT F [NS11]. QD [Bai05c]. QDMC [CZ90]. QEDMOD [STY15, STY18]. QLP [CS14]. QR [RG90a, RG90b]. quad [Bai05c]. QUADLOG [BB07]. Quadratic [CV94, PPR97, Car91a, CZ90, Hop03, RPL96]. quadrature [BB07]. Quality [Boi97, Sht19, Jam94, Jam96]. Quantity [Fos17]. quantum [AFAS99, HKŠ^{+}97, HMT90]. quarkonium [WW14]. Quasi [BGW93]. Quasi-Likelihood [BGW93]. quasiparticle [ZZSW19]. quaternion [AAK01]. Queries [BDH90]. Query [HM96]. QUERYFLEX [Ano97b]. question [BC97]. Questionnaire [Wal90]. Queues [CC93]. Quick [Lib90a, IMS91a, IMS91c, IMS91b]. QuickBasic [Tay99, CD92, Pao99, Pao01]. quotient [Kah01]. quotient-digit [Kah01].

R [Ano96a, Ano99a, Ano99b, Hin06, Iha06, GH18, GHN19, GMF18, LHW01, Wie94]. R.J [Gon01]. R1996 [Ame96]. R1997 [Ame97a]. R1998 [IEE92a]. R22 [KK90]. race [CFMR95]. radar [HW95]. Radial [Gil94]. radiation [SS90, Unixx]. radiative [AIS^{+}97]. radix [Kir02]. raising [Lib90a, Lib90b]. Raleigh [Agr95]. Random [HD93, Lev92, Wal91b, Yan95, BS13, CBTL97, DW03, FSV90, Hen94, Hen95, MNZ90, MZT90, Wol92]. Random-effects [Yan95]. Randy [Dub97]. RANEXP [Hen94]. Range [HK93c, KH93, PG10]. RANGEN [Fah02]. ranking [Cas14]. RANLUX [HJ97, Jam94, Jam96]. Rapid [MZM94, NSU20, RKMJ92]. Rasch [Gre93]. RATCHET [SJ94]. rates [ZMR^{+}91]. Rational [Mac96b, MKC92]. Ratios [McB06]. Ray [DP94]. Ray-Tracing [DP94]. RCCPAC [MCA17]. re [EKC95]. re-engineering [EKC95]. reactivation [RM90]. Reading [HK93a, HK95, Que00]. Ready [DZ98]. Real [Ano93b, BBCH95, Bra94c, FCHE02, FGG09, Bra94d, FGGL05, GST06b, SJ94, SIOS02, Ano94o, GST06a]. Real-Time [Ano93b, FCHE02, SJ94, Ano94o]. real-world [SIOS02]. Realistic [SN94, HMT90, Tea94]. realities [LSW92]. Reality [Ano93n]. Realization [BG96, PD96]. Reasoning [Ste95b]. Receives [Lew94]. Recipes [Adl93, Gar93, Kon94, Lev98, Pre93d, PTM96, PTV96, Spe93, Tha93, V^{+}93, Wu93, Yan94b, AB93, Ano92e, Loz98, PTVF92, Pre92a, Pre92b, Pre92c, Pre92d, Pre93e, Pre93f, Pre94b, VTP92, Vet93]. recognition [PQ94]. Recompilation [AMKS02]. Record [IEE92b, ACM93c, ACM94b, ACM95b]. rectangles [MNZ90]. Rectangular [GWDL08, GWDL10, Deu90, Dot93, FYR99, YRF02]. recurrence [Ove91]. Recurrences [ONT95, TIUG90]. recurrent [GF95b]. Recursive [WAG98, BT94]. Red [Ano96b]. Redistribution [TCF94, WO96, HC08, KN95, KHJS94, KHRS95, WW94, WW95]. Redistributions [BG96, GHSJ94]. Reduce [GS97, BGV94, GV92, Kea92, LD90]. reduced [BBB00]. Reducing [Mra94]. Reduction [DGR92, GP97, DGR90, GP94, HD05, RP95]. reductions [YWS^{+}94]. Redundancy [BC94]. Redwine [GMC96a]. refactoring [OJ09]. Reference [A^{+}92, ABW92, ABM^{+}97, Cha95a, HH18, Lan90d, Rap90, Sci93, Spe96a, Scixxb, Spe96b, Sun05, ZB94b, AFAS99, Ano91d, Ano91f, Con91, CS90c, CS91, Cra91b, Cra92, Cra93, Dig92, DV02a, DV02b, FT03, Hew90a, Hew90b, Hew91a, Hew91b, Hew92b, HW91, Int90a, Int90c, Int90f, Int90g, Int90l, Int90k, Int91a, IBM91e, Int91e, Int91d, Lib90a, IMS91a, IMS91c, IMS91b, Lan90e, Lah90, Sil92a, She91, Sun94, Ano98b].

References

[Ham85, HM90a, RH84, MHT96]. refine [Smi93b]. reflection [CB95]. Refrigerant [KK90, Cra95]. REGCMPNT [Bel11]. Region [CI96, GT03, GT07]. regional [Gep90]. Regions [Pau93, Som98]. Register [BCT94, EDA96, CCK90, FSV90, KH93]. registers [NIY^{+}94]. Regression [Bel11, BGW93, KTMB02, ZBW07, vV90]. Regrowing [OJ09]. regular [Gao05, SSG^{+}10, SSG^{+}18]. regularizations [DDH17]. Reid [Ano91a, GMC96d, Gla92a, Rub93]. Rejection [HD93]. Rekindled [Can92b, Can91, Can92a]. Related [HFMS95, Smi11, BB07, Smi01]. relation [Ove91]. Relations [van90b]. relative [Ude91]. Relativistic [FGJB19, MCA17]. Relayed [Coc03]. Release [CGT92b, CGT92a, Hew01, Int90l, IBM91a, IBM91b, IBM91e, Int91e, Dig90a, Int90i, Int90j, IBM91d, Ing90a, Ing90b, Uni93]. reliability [EP92, Kle93, Zah92]. Reliable [Cse99, Enr95, EH07b, Vig93, Wal93b]. Remappings [CA96]. Remark [AFS94, Bre79, DV00, Esp98, Fox79, GL90, Ham85, HH18, Has06, Hig91, HM90a, Hop02, Hop03, KHS17, LS00, MN11, Rei96a, RB98, Ren99a, Ren99b, SWH15, Zag16]. Remembering [McJ17c]. REML [Yan95]. Remote [BDK91, GS97, Ano96b]. Removal [KK95b, Hor09]. renewed [Ano91e]. Repeat [Pug90]. Replace [Kef92, Mor81]. replay [CFMR95]. Replicated [CGS93]. Replication [Gil91b, Gil91a]. Reply [Gho01]. Report [Ano97b, DZ98, NR98b, Ste93, MSZ90, Ngu91]. Reports [Ano93a, Ano94g, Cof93]. Representation [FWH^{+}94, Ric06, BMO90]. required [ZE92]. Requirements [ACG^{+}94, CHKM93, EDA96]. resampling [KHC92]. RESDYK [CNP91]. Research [ASM^{+}94, ABB^{+}94, GS01b, HM90b, Kea95b, SZAB98, AGG^{+}97, BLT94, BPG94, KT94, Lav91, Sha94, Wie94, WFW^{+}94, Lev94, Oed93]. resection [DH95]. Reservoir [KR94, KR95, AG95b, Kut92]. reshenie [AZ90]. residual [Av94]. Resilient [FGBN19]. resin [WHL92a, WHL92b]. resistivity [CNP91, Dut94]. Resolution [PR91, DH95]. Resources [FKL94, Met99b, Ano07]. Response [Gre93, SH09, Joy92, MM02, Tre91]. responses [Ple93]. Restricted [KH13]. Restructuration [Tro90]. Restructuring [EHJ^{+}91, EHJ^{+}93, TMD13, DDcMR96, Eig90b, LP92, LP93]. result [AK93, AFKL04]. Resultant [GV92]. resulting [FR94, HM93]. Results [BCF^{+}93c, Cod90a, McC95, MMY95b, NOL97, PA94, BCF^{+}93b, DFRR91, FBC96, HKŠ^{+}97, MMY95a, Nar95]. Retargetable [BCM^{+}93, IGHG^{+}94, SNMC93]. retargeting [Lan90a]. Retire [Can91, Can92b, NK94, Can92a]. Retraining [JL93]. Reuse [Jéz93, PSC93b]. Reversal [Kar96, Ves91]. Reverse [Hor92, JSY^{+}20, CC98, HGG93]. reverse-engineering [HGG93]. Review [AB93, Ano90b, Ano91a, Ano96a, Ano97a, Ano98a, Ano98b, Ano99a, Ano99b, Ano03, Dub97, Eme94, GMC96b, GMC96a, GMC96c, GMC96d, GMC96f, GMC96e, Gen06, Gla92a, Hin06, Hop97, Iha06, Kri86, Lev98, Mai91, Olv91, Rag95, Sch07, Spe93, Tay99, TDMC97, UMM94, Wei94, Yan94b, dL12, Jam90, Mol12, Rys95, Sch91b]. Reviews [Ano97c, BCM99, Bra97c, EMUP98, KG99, Loz98, Mar98]. Revised [HR92, AFKL04, MB92]. revisions [HMT90]. revisited [GG95, GOT03a]. revived [Cel96]. revolution [HL08]. Riccati [BBZ95]. Rice [MCAB^{+}02]. Ricicle [Smi93b]. RIDGE [vV90]. Right [Fuj95]. Right-Hand [Fuj95]. rigid [CZ10]. RISC [Bel90a, Bel90b, Ano91c, DD99, IBM93, MSC96]. rise [KKZ11]. risks [ZMR^{+}91]. RKSUITE [BG94]. rksuite 90 [BG97]. RNGSSELIB [BS13]. Robert [Eme94, Rag95, UMM94]. Roberts [Dub97]. ROBETH [MJR93]. Robot [NJ94c]. Robots [Coc03]. Robust [EH07b, LHW01, MJR93, Sug95, HS10, KB94]. Rochester [FH90, RFC90]. Rockefeller [IEE90a]. rocks [MSB92]. Role [JL93, Wie94]. ROM [Ano97a, Ano98a]. root [Hig93a]. roots [BN93]. Rosenbaum [Nan93c]. rotating [KLM^{+}19]. rotor [NG93]. round [JCL10]. round-off [JCL10]. Rounding [CBTL97]. roundoff [Bli90]. route [Gro91]. Routine [Alo04, BBCH95, DPS02, LH92, BG93, NVFNP93]. Routines [Cod93a, Cod93b, FGG09, Lan01, MJR93, Mil04b, RVV^{+}92, BB07, CZ10, DH84, FGGL05, GST02a, GST02b, LS04, Nar95, Ngu91, Par94]. Royale [BLT94]. RPC [RS93]. RSSP [Tre91]. Run [FL91, OP98b, PQ94, Sch93b, SPM^{+}94, SS96, DNS98, DN04, RP95, SM92]. Run-Time [DCZ96, OP98b, Sch93b, SS96, PQ94, SPM^{+}94, DNS98, DN04, RP95, SM92]. Runge [EH07b, GKKL19]. Runtime [ASS93, ASS95, AES^{+}96, BBG^{+}93, HMS^{+}95, Jus92, PSC93b, PSC^{+}95, TBC94b, TCF94, AFMP95, PSC^{+}93a, PDS^{+}93]. Russian [AZ90, Mal91]. Ryan [Mar90]. Ryan-McFarland [Mar90].

S [Adl93, AB93, Ano98b, Eme94, GMC96f, UMM94, Yan94b, Mal91, BSS92, BKP93, FGCG94, Las97, MJR93, RD92, Yan95]. S-PLUS [MJR93, Las97, RD92, Yan95]. S. [Mal91]. S02YSCODE [CFGG94]. S12 [HKS91]. SAC [GS01b]. safe [GOT03b, Van22]. Sale [SW91]. Salesman [CT95, PR91]. Saline [BLT94]. Sample [Ano96c, Gil91b, Gil91a, Gil01]. Sample-Size [Gil91b, Gil91a]. samples [Coh90, Gho01, Tsa01]. Sampling [Air04, SH07, SH09]. San [ACM93a, ACM95b, ACM97, Ano94a, BBG^{+}95, IEE93a, Kar95]. Sanford [Rub93].

Santa [Ano95c, IEE95a, USE94]. SAS [SB01]. SAS-IML [SB01]. saturated [TOC18]. saturation [EN96]. Saul [Gar93, Loz98]. Savez [Ain90, Ain91, Ain93]. Savez-vous [Ain90, Ain91, Ain93]. SC2002 [IEE02]. SC2003 [ACM03]. SC22 [W^{+}95]. SC22/WG56 [W^{+}95]. SC22/WG56-N1122 [W^{+}95]. SC33 [ANS95]. SC33/WG5 [ANS95]. SC33/WG5-N1122 [ANS95]. SC97 [ACM97, ACM97]. SC98 [ACM98, ACM98]. SC’99 [ACM99]. Scalability [PMBH93, SSG94]. Scalable [BBG^{+}93, BCF^{+}94a, Fox94, IEE94d, IEE94e, IEE95b, SS96, ARB94, ARB95, BBB^{+}94, BB02, MS00a, MS00b, Mic97, PSG03, Sal06, ZCP95]. ScaLAPACK [BDPW98, BG96, LMMW96]. scalar [Phi91a, SSS99]. Scale [BC01, CT95, CGT92b, JPE20, KMJ^{+}23, PR91, SF92, SM03, TT92, VBA95, BHLT09, CDF^{+}93, CGT92a, EH07a, GOT03b, KSZ90, LS09, LN91, LMV09, MN11, Tor10, Zim07, ZBLN97]. Scales [EL97]. SCAN [Cse99, Ste91, AH92]. SCAN-91 [AH92]. SCAN-98 [Cse99]. Scattered [Ren97b, RB99, TZW^{+}10, Aki96, DV00, Ren96b, Ren04]. scattered-data [Aki96, DV00]. scattering [AIS^{+}97, NVC96, YK90, YB92]. Schafer [Sch07, Hin06, Iha06]. Schaum [MC95a, MC95b]. Schaum’s [MC94]. Schedule [EDA96, PSC93b]. Scheduling [EDA96, KK94, Huf93, KE93, KY94, Luc92, LFK^{+}93]. schématique [Tro90]. Scheme [HK91, HMKN91, HLJ95, HBD^{+}93, KY98a, KY98b, Harxx]. Schemes [Fuj95, SVD96, GV92, SDv98]. Schofield [Sch91b]. School [Van95]. Schreiber [Eme94, Rag95, UMM94]. Schrödinger [CRS90]. Schur [Koi09]. Science [Ano93a, Ano95b, Bro90a, Cha94c, EPL94b, Gla92a, HK93b, Lap96, NRS92, Pri10, SMSY02, HCD^{+}98, HK93a, Kor99, LD87, ZJEP95a, ZJEP95b]. Sciences [ERS95, Glo91b, HS94b, HS94a, MS94]. Scientific [AK93, Adl93, Adv98, AH92, BBG^{+}95, BC01, BN93, BN96, Ber92, BSPF01, CNBB96, Cse99, CHKM93, DS97, DB93c, DY99, Ein91, Eme94, FJSD96, FMW^{+}94, GRE99, Hun00, JPE20, KM97, Kon94, Lev98, LP98b, LMR^{+}97, Mös95, Nat00, Ort94b, Ort94a, OPB^{+}20, PAK^{+}90, PTM96, PTV96, Ric95, Sny07, Ste95a, Vel97, Wic89, Wil95a, Yan94b, AHZ90, AB93, Ano96a, Ben99b, BC97, BT01, Bou95, BSB^{+}03, CSC^{+}97, Cel96, Che90, CJPA94, Din99, DT94, DW94, EFP07, FCHE02, KB94, Kug92, LP05, Lav91, Loz98, NDSG07, PD96, PTVF92, Pre94b, SM02a, Szy07, Tou84, Vig93, Wal93a, Wal93b, Wil95b, YYX^{+}07, Ano97c, Bra97c, Kri86]. scientifiques [Cha94a]. Scientists [BS91a, Bro95, Bro97, Cha97a, Edg92, Ett90, Ett92, Ett93, Ett96, Ett97, For97, Hah94, HB91a, HB91b, NL92, NL95a, NLN96, NL96, NL97a, RZ94b, Rub93, Smi94, Smi95b, Ano90b, Ano99a, Ano99b, BS91b, Cha95b, CC95b, GR92, Gon01, NL95b, NL97b, Mar98]. Scope [Ano93e]. scoped [BGS82]. Scorer [GST02b]. scores [KDG99]. Scratch [SWW90]. Scratchpad [JT94]. Screen [BL90, Wei95]. Screening [MC92]. SCS [PAK^{+}90]. SCS-40 [PAK^{+}90]. SDECAY [MDM05]. sea [SS90]. Search [CL93, Lan01, LP19, McJ17b, Hig93a, Sav95]. Second [BB91, Dem95, Kub91a, Kub91b, Kub91c, Tha93, Vol93, Wu93, IDVV97, Loz98]. second-order [IDVV97]. Secrets [Mit92]. section [Mir90]. sections [Hor09]. SEEK [Sav95]. Segmented [HCLJ03]. segodnia [GU90]. seismic [CB95, Joy92, Mai90]. Seismograms [DP94]. seismograph [Ple93]. Selected [HR92, Knu03, McC95, PA94, WW93, Bar92, SMB90]. Selection [KS90, AS92, Kah01, Sar97, dVdVI97]. Self [GG99, GWE^{+}05, AI90]. Self-adapting [GG99]. self-contained [AI90]. Self-Healing [GWE^{+}05]. Seligman [Sto93]. Semantical [DJ92]. Semantics [PEP92, EB98, Guo01]. Semantics-Based [PEP92]. semiconductors [LZL11]. semiempirical [HKŠ^{+}97]. seminar [AFKL04]. Sempa [LMR^{+}97]. Senans [BLT94]. sensitive [Huf93]. Sensitivity [Hor91a, Hor91b]. sep [Koi09]. sep-inverse [Koi09]. separable [CDGM96, GBDB97]. Separated [Lie94a, Lie94b, OM92]. Separated-form [Lie94a, Lie94b]. separator [Cok93a]. September [BLT94, BV94, FK95, IEE94c, Sch93b, Van95]. Sequence [Hig94d, KNS95b, NL19, KNS95a]. Sequences [TR96, BD93, CH96, Ste91, SV95, TRS91]. Sequent [Cod90a]. Sequential [HMW91, HMW93, SR95]. Serial [SWH15, BF92, GS98, HWS09]. Series [DLM99b, DLM99a, EPL94b, Rit90, SAC^{+}92, App91, Eme94, GL10, GMMM92, Hew90b, Hew91a, Hew91b, Kay90, Mat90, PW93, Sat97]. Server [Ano93n, Ano96b, Sch93a, ABB^{+}91]. Servers [Teo01]. Service [Kri86]. Set [BCC^{+}96a, BCC^{+}96b, DDHD90, DCHH88b, DCHH88a, FGG09, KHS96, KN94, Lin93, Per93, Pre93d, RFS98, van90b, Ano95g, BCC^{+}97a, BCC^{+}97b, CZ90, DLLR96, FPR01, FGGL05, Has06, TS06b, XHY^{+}24]. Sets [AMC98, CGL^{+}95b, JB01a, JB01b, Wal92, BxCW01, CGL^{+}93, KHS95, PW93]. Seventeenth [NRS92]. Seventh [BBG^{+}95, HS94b, HS94a, MS94]. Several [MMY95b, GBR15, MMY95a]. Severe [Wic99]. sFr [Ano97a]. SFUN [IMS90a]. SFUN/LIBRARY [IMS90a]. SGI [Sai95]. Shadow [GRE99]. Shadow-Object [GRE99]. shallow [NY91, Ste90, ZZN94]. shallow-water [NY91, Ste90, ZZN94]. Shanghai [IEE97]. Shape [Cos97a, Cos97b]. Shape-Preserving [Cos97a, Cos97b].

SHARE [Ano93n]. Shared [BP92,

BGLP94, BKP93, BMMN94, CL97, DCZ96, PMBH93, PWD93, BB02, Bod94, DPZ97, Ger98a, Ger98b, OH90, Phi91a, WYJ99]. Shared-Memory [BP92, PMBH93, PWD93]. Shell [Phi91b, Phi92, SH97, MCA17]. Shepard [BM99, TZW^{+}10]. shielding [Unixx]. Shift [KMR96, FSV90, STY15, STY18, SSLG91]. shift-register [FSV90]. ship [Mil92]. shock [KK90]. short [ZCP95]. SHPF [MCH96]. SIAM [BBG^{+}95]. sic [RBD^{+}10]. Side [Fuj95, CHT92, HK90]. Sierra [Pre93b, Van94b]. Sieve [Alo04]. SifDec [GOT03a]. SIGACT [ACM93c, ACM94b, ACM95b]. SIGCSE [Ano95b]. Signal [SD92, SD93]. signals [Ame90a]. signatures [Pre99]. Significance [SD90]. Significant [GKKL19]. SIGPLAN [HOP93, HOP93, ACM93c, PPP93, ACM93b, ACM93a, ACM94b, ACM95b, Ano95c]. SIGPLAN-SIGACT [ACM93c, ACM94b, ACM95b]. silicate [SSLG91]. silicon [SMB90]. SIMD [GGW96, KLW93, Rot93]. similar [HD05]. Simple [Wal92, BC19, Ngu91, YB13]. Simplices [BCE93, GC03]. simplification [Nat92]. simplified [CK90, DN04, Shi98]. Simplifying [MP93]. simulate [FHE95, MB92]. simulated [GF95a]. Simulating [MMEH08, Wic99]. Simulation [ADHF96, Chi91, Cok95, DFS95, Ger94a, Hun00, KR94, KR95, LMK94, MDD94, MMV95, PTS92, SMSY02, Ten93, BD93, Bra94a, CZ90, Cra95, DCR99a, GBC92, GAW96a, GAW96b, Heu90, KSYE00, KDG99, Lef93, MWM90, MSZ90, Nan93a, Neu01, Ogi02, Sre92, Tal94, Tre91, Uni93, WHL92a, WHL92b, XHY^{+}24, van90a]. Simulations [GPS99, MB95, SM02b, Cah90, DLLR96, FCHE02, KMJ^{+}23, KT00, NSWP90, QRH00, TOC18]. Simulator [OC94, SMSY02]. Simultaneous [CJL97, SB01]. Sinc [SS99]. Since [NRS92]. sine [Mac96b]. Single [EB98, MR93b, Bai05b, Bar94, Cok91, NH09, VKB93]. single-expression-use [NH09]. Single-Input [MR93b]. single-step [VKB93]. singular [Gao05, Gao06]. SISAL [CF90]. situ [SS90]. Sixth [Ano94a, HK95]. Size [Gil91b, Gil91a, Coh90, Gil01]. sizes [Kir02]. Skew [BSV16]. Skew-Hamiltonian [BSV16]. Skew-Hamiltonian/Hamiltonian [BSV16]. SL-GMS [She91]. Sleightholme [GMC96b]. SLEUTH [GM97]. SLHPF [BDPW98]. slicewise [SMG91]. Slicing [DSv94]. slickenside [Gep90]. slide [NIY^{+}94]. slide-windowed [NIY^{+}94]. slip [YRF02]. slower [Sal92]. small [FHE95]. Smith [Ano98b]. SMMP [MMEH08]. smoothing [Dem03, Dem06]. Smoothly [PS96]. SNA [KSW93]. Sneak [Smi00]. Society [IEE94g]. SoftBench [Bet97]. SofTech [Spo94]. Software [Ano92c, Ano95f, Ano96b, Ano97b, Ano97d, BGKZ91, BPG94, BLLWW95, BD91, BMR01, Bou97, Bra97d, BG97, CFGG94, DLM99b, Don91, DV92, DCZ96, FGCG94, GGLM88, GL90, GWL^{+}92, Gen06, Her90, HS94b, HS94a, Hin06, IFI95, Iha06, Ken92b, KO91, LMR^{+}97, MFK09, Mil04a, MGH81, NS92, Sch07, TS25, Ueb97, WNO94, Ano97c, AP90, BHLT09, Boi97, BT01, Boo81, Bra97c, CMVZ94, CH96, CKT85, Don90, DPZ97, FG93, Fos95, GBC92, Hop02, Kas93, Kor99, Lav91, LS05, Mac96b, Mac96a, Mil91, Nag90, Osy92, Paj90, RBD^{+}10, RBD^{+}11, AFKL04, Ren96b, Smi01, SMH91, Tal94, WKM04, Ano93p, UM93]. Software-Entwicklung [Ano93p, UM93]. Solaris [Sun93]. Solution [BSV16, CT95, DL97b, DPS02, DR93a, DR93b, GWDL08, HIS91, MD97, MKFB92, PPR97, RFS98, Rhe93, WW90, DR95b, AR06, AZ90, BHLT09, DR94b, DR95a, Duf04, GWDL10, HS10, Hop03, IDVV97, KKK95, KL92, OH90, Pri93, RPL96, RR99, Ste90, Tea94]. solution-gas-drive [Tea94]. Solutions [BGG^{+}94, Nak95a, Shi93b, AF92, Azi23, CWB92, FPR01, GST04b, Has06, Rib02, Sch91a]. Solve [MR93b, MR95a, BKRG22, PS08, TS06b]. solvent [BDH^{+}05]. Solver [Fat94, LZ97, NRK98, SM17, CK91, GBDB97, HIS91, HS10, KTP^{+}24, OM92, PZY16, Rei93, RS09b, TS06a]. solver.Scientific [BG94]. Solvers [DL97a, Pry99, ARB94, ARB95, HBG^{+}05, MSP^{+}22, SSH08, dSL98]. Solving [Ano92c, Bro97, Cas89a, CC92a, Edg92, Ein91, GWL^{+}92, GM97, Hig91, HRW^{+}98, Ric95, Sab95, TT92, WR93, Cho91, GT03, GT07, GDS94, Hop02, KRG21, KM99, KF90, KF92d, KKY99, LP05, LD90, NY91, RBD^{+}10, RBD^{+}11, Sil93, SMSW06, TJ90, WSW00, YSVM^{+}16, YSMA^{+}17, YSMBA23, ZT90]. Some [Bra94c, BKR^{+}91, Che92, SG93a, SG93d, Ein96, FBC96, HK93c, McC96, Per94, Tay97, Bra94d, BLL^{+}96, Cof93]. Somerville [Som07]. son [II90]. SONGS [OC94]. Sons [Ano96a, Ano98a, Ano99a, Ano99b]. Sopron [Fer92]. Soputuweo [nY90]. SOR [Yam95]. SOR-like [Yam95]. SORCERER [SOP93]. Source [KMBK96, SD99, UNF^{+}08, AC17, Che91, FTPR04, SOP93]. source-to-source [AC17, SOP93]. sous [Pic94]. South [ACM93c]. SP2 [GMS^{+}95, Sai95]. Space [AF92, CMP02, XHY^{+}24]. Space-time [AF92, CMP02]. spacetime [Rib02]. Spain [ACM95a, IEE92b]. SPARCompiler [Sun92b, Sun92a]. Sparse [Bou97, CCL01, DL97a, DL97b, DGL91b, DGL91c, DGL91a, DR93a, DR93b, FB12, LP98a, MSC96, PPR97, Pet91, Rei02, UZCZ97, DR95b, CCL04, DR94b, DR95a, DV98, Duf04, HS10, Hop03, LP99, RS09b, SZG95, UZCZ95, UZCZ96, dSL98, DVY00, DV01, DV02a, DV02b, DHP02]. sparticle [EH07a]. Spatial [RD92, AM90, SZ90]. spatio [Azi23]. spatio-temporal [Azi23].

SPECFUN [Cod93a, Cod93b]. Special [Ano94m, Cod93a, Cod93b, Hig94b, Hig94c, Hig94d, KS02, NR98b, SF02, W^{+}95, ANS95, IMS90a, IMS91f, Lan90a, Mac96a]. special-purpose [Lan90a]. Specialists [IEE92b]. Specialization [Bla00, KKZG94, KKZG95]. speciation [WRL90]. species [Lar93]. Specification [Ano93g, Ano94e, Ano94f, Ano94m, Ano94o, Hig93b, FHK^{+}90b, Fox91a, Hig94a, Hig94b, Hig94c, Hig94d, CS90b, FHK^{+}90a, Fox91b, MKS^{+}96, SKM94, ZBC^{+}92, Zim92, Hig92]. specifications [CC98, Sha94]. specificités [DV93]. specified [PSC^{+}95]. Specifying [Bla00]. Spector [Mol12]. spectral [Eli98, GS95, MH91]. spectrometers [SS90]. spectroscopic [BG93]. spectrum [DKM07, HIK90, Cas14]. speculative [RP95]. Speed [ARB94, ARB95, Ano93j, BID95, Lee90, Lin90, OM92]. Speed-up [ARB94, ARB95]. Speeding [CC93]. SPG [BMR01]. Sphere [Ren97a, Ren97b, NVC96]. spheres [BDH^{+}05]. spherical [NY91, NVC96]. SPICE [Wri99]. SPiDER [FSPC^{+}02]. spin [BKRG22, HHCS95, KRG21, PS08]. spin- [BKRG22, PS08]. spin-1 [KRG21]. spin-orbit [BKRG22, KRG21]. Splancs [RD92]. SPLASH [Hol90]. Spline [MKFB92, Ren03, Ren09, WKM04]. Splines [Cos97a, Cos97b, Lai92a, Lai92b, Yu01]. SPMD [RLS20, Wal02a]. Spotlight [Ano95f]. spring [IEE93a, Sto93]. Springer [Ano97a, Hop97]. Springer-Verlag [Ano97a]. Springs [Ano94l]. SPRNG [MS00a, MS00b]. Spyglass [Ano96b]. SQL [Che91]. square [BBB00, Hig93a]. square-reduced [BBB00]. Squares [Dem95, Sou91a, Sou91b, CS14, Dem97, Dem06, Dem07, GT07]. SR [Cra93]. SR-3772 [Cra93]. SRFPACK [Ren96b]. SRRIT [BS92a, BS92b, BS97]. SRT [Kah01]. SRTEST [Kah01]. SSRFPACK [Ren97b]. Stability [DH92, Pau93, CZ90, GV92]. stability/performance [CZ90]. Stable [CJL97]. Stage [EDA96, Cho91, Cum90]. Standard [Ano94j, DET12, Don91, DV92, DB93a, Ins91a, Ins91b, IEE92a, ISO04a, Ano94k, Ano95d, BN96, Don90, DOSW96, DHP02, Ins92, II91, Met99c, Met99d, Rap94, Ame90b, AC92, Ame92, Ano94o, Ano97c, Bra97c, Ell90, Nag02, RN07, Vil94]. Standards [Fei94, FKKC96, Ano94n, Ano94o]. STAR [Coo95]. STAR/MPI [Coo95]. Starbase [LS90c]. started [SB92, Thi91]. STAT [IMS90b, IMS91g, IMS91h]. STAT/LIBRARY [IMS90b, IMS91g, IMS91h]. STATCN [PSPE94]. STATCW [PSPE94]. State [IEE94e, IEE95b, KAČ^{+}22, Mic97, MR95a, Nak90, CHM91, DLW^{+}18, GRSS02, HM12, Ude91]. State/NCAR [Mic97]. statement [ALS91, KHS95]. Statements [Bee90, BBZ94, KHS96, SOG94, GG95]. Static [GS97, YKK96, ACIK97, Bet97, CGL^{+}95a, KY98a, KY98b]. Stations [WNO94]. Statistic [Nan93b, RH94]. Statistical [Gen06, Hin06, Iha06, Mil04b, Sch07, IMS90b, IMS91c, IMS91g, IMS91h, LS05, PSPE94, PW93]. Statistically [GKKL19]. Statistics [Gil94, HFMS95, MJR93]. Status [DZ98, MVZ98a, Nak95c, Nak95b, Ste93, Zim02, MVZ98b]. stdio.h [Lev97]. Steele [Eme94, Rag95, UMM94]. Steering [JBBH93]. Stencils [RMCKB97]. STENMIN [Bou97]. Step [She92, MSA03, NY91, VKB93]. stereographic [Cum90]. Steuerung [Por90]. Steven [Ano03]. Stiff [CC92a]. Still [Met92b]. Stochastic [CFGG94, FGCG94, AS92, Vig93]. Stockholm [HAM95b]. Stokes [Fat94, RRX^{+}08]. STOL [How91, SH91]. StopWatch [Mit97]. Storage [Hig94d, Rotxx, SVD96, Cra95, SDv98]. store [KH93]. storm [CDF^{+}93]. storm-scale [CDF^{+}93]. Strassen [Han92]. Strategies [BB96, DCR99b, MCAB^{+}02, PCS99]. Strategy [RRM^{+}15, CCJ93]. stratigraphic [CM92, CB95]. stratigraphy [MB92]. stream [MSZ90, YYX^{+}07]. stream-aquifer [MSZ90]. streams [BS13]. Street [Eme94]. Strength [Kon00]. stress [Gep90]. STRGRH [YRF02]. STRGRV [YRF02]. strike [YRF02]. strike-slip [YRF02]. Strings [Cou97, SM90, IEC94, ISO94, Int00]. strip [WW94]. STRIPACK [Ren97a]. structural [Kay90, dLJEB95]. Structure [CHL94, FGJB19, BF92, BCC^{+}97b, KLM00, Nar95, Off94, PGH^{+}90, SS10, SSS99, Tur93]. Structured [ASS93, ASS95, Ett90, Ett93, Ett96, Ett97, Lin93, Mas93a, Per93, Ves91, Alt90, DR94b, Ell81, GBR15, IJCL96, KF90, KF92d, Mas92a]. structurée [Ain90, Ain91]. structures [KGV97, Smi93b, Unixx]. STS1 [Kay90]. Student [Cam13]. Student-Oriented [Cam13]. Students [Pif96, WR93, Cho91]. Studies [Hor91a, Hor91b, KNOR04]. Studio [Sun05]. Study [BF01, Buc94c, GLPE97, GS01b, KR94, KR95, McC96, MM98, Pre93c, RRM^{+}15, SN94, AJJF14, Bri00, CHM91, CZ90, CHT92, DS97, EKC95, GF95a, Gil01, LCD91, LSW92, MS93a, NLVE15, RTY90, SLY90a, SLY90b, VSH91]. Sturm [BGKZ91, GM97, Pry99]. Style [Mol12, SKP91, SS00, Wal90, dL12, Kes92]. suan [yKxx]. Sub [Tay97]. Sub-languages [Tay97]. Subprograms [DGL91a, DCHH88b, DCHH88a, DDHD90, DV98, DHP02]. Subroutine [Amo90, BS97, CT95, Cas89a, CC92a, Dem95, FJS97, Hig91, Kod08, Kri86, MR93b, MR95a, Ram90, TT92, BS92a, BS92b, Coo94, Dem97, Deu90, GP94, Kay90, Kir02, SHCP91, Wol91]. Subroutines [BSV16, BFKS93b, BL90, BGW93, CV94, DGR92, HC92, HC94, JSY^{+}20, JP95, KN94, MGH81, PPR97, RG90a, RFS98, Shi93b, ZBLN97, BBB00, BK06, BFKS93a, Das06, DGR90, Err06, EC13, FPR01, GRW07, Has06, Hop03, IMS90a, IMS90b, IMS91c, IMS91b, IMS91f, IMS91d, IMS91e, IMS91g, IMS91h, Kea92, Ker90, Lai92a, Lai92b, Las97, MN01, MN11, Pre99, RG90b, RPL96, RR99, XWK95, DGL91b, DGL91c]. subscripted [CCK90]. Subscripts [SSC00]. Subset [Ano93c, Ola96, Gla92b, Par86, MCH96]. subsets [Shi98]. Subspace [BS92a, BS92b, BS97, Ram90]. substitution [CHT92]. subsurface [Tur93]. SUIF [WFW^{+}94]. Suite [DG94, SF02, WMMW97, DS02, DFRR91, HJJ^{+}00, HBG^{+}05]. Suited [HD93]. summarizing [BK89]. Summary [Bee91, SZAB98, IBM91e]. Summit [HDR03, BC19]. SUNDIALS [HBG^{+}05]. SunSoft [Ano95g]. SUPERB [ZBC94]. Supercomputer [Ano94p, DKMS91, GAW96a, GAW96b, ST90]. Supercomputers [Ken92b, LW89, Car91b, Car92]. Supercomputing [ACM94a, ACM95a, ACM96a, ACM96b, Ano93q, HK93b, IEE90a, IEE91, IEE92d, IEE93d, IEE94f, Kar95, KSW93, BBF^{+}92, HK93a, KT94]. SUPERFLUID [BSS92]. supernode [Mar92]. supersonic [Dan90]. Supersymmetric [DKM07, DET12, MDM05]. Support [ASS93, AH94, Ano94a, Bra00, BGS94b, BLW02, But95, CFK^{+}94, CCL04, FBZ92, HKT92a, Ken94b, MR95b, OP98b, Sch96a, SZAB97, TBC94b, AES^{+}96, AH91, Bro03, DNS98, HDH^{+}94, HDH^{+}95, HKT91c, HMS^{+}95, II90, PSC^{+}95, SPM^{+}94]. supported [San92]. Supporting [Pon94a, Pon94b, PHD^{+}95, BMO90, GMF18]. Supports [CCL01]. SUPRENUM [Hem94, AHJS90, McB91, ST90]. Suprenum-1 [McB91]. Surface [Ren97a, Ren97b, Tre91, Yu01, Aki96, BDH^{+}05, DV00, RBS93a, RBS93b, Ren96b, Ren04]. Survey [FKL94, Paz96]. survivors [App91]. SuSpect [DKM07]. SusyBSG [DGS08]. SVM [BGNP93, BGS94b, Ger94b, GB95]. SVM-Fortran [GB95, Ger94b]. Swansea [Bar92]. Sweden [HAM95b]. Sweep3D [CDMC06]. SX [MAH^{+}02]. Sylvester [GWL^{+}92, Hop02]. symbol [AP90]. Symbolic [ACM94c, Ano97d, BF93a, Bra97d, BKR^{+}91, Cre90b, DJ92, GDS94, HR92, Lev95a, LP90, Sen03, Var95, WN90, Coo95, FSPC^{+}02, Gro90, LP05, Mat90, Paj90]. Symbolic-numeric [GDS94]. Symmetric [BMV03, NV94, PR91, DR95b, CS14, DR94b, DR95a, Duf04, HS10, Raj95]. Symmetry [Cod90a]. Symposium [ACM93c, PPP93, ACM93a, ACM94b, ACM94c, ACM95b, Ano94d, Ano94i, Ano94p, Ano95c, Ano95b, AH92, Cse99, HHK94, IEE92c, IEE93c, IEE94a, IEE94g, IEE96, Lev95a, Sen03, Sie94a, Sie94b, Ten93, USE94, WN90, ACM91, Bar92, IEE95a]. Synchronization [AH94, GS97, AH91]. syntax [Num05]. Synthesis [HLJ95, HLJ01, Per94]. Synthesize [HLJ98]. Synthetic [DP94]. System [Ame96, Ame97a, Ano93b, AO90a, AO90b, AO90c, BK95, Bai94, Bai95, BBG^{+}93, BGS94b, DCZ96, ERS95, Fos93, FXAC94, FES05, Gar91a, Gar91b, HBG^{+}06, HS94b, HS94a, HKK^{+}91b, HKK^{+}92, HIM91, HM90b, IBM93, IEC90, IEE92a, IEE93b, IEE94g, ISO90, Kas93, KO91, Koo90, MS94, Oed93, Osy92, Sar91, SP91a, SP91b, WW90, YYM93, AS92, AKLS88, BBB^{+}57, BL94, Che91, CFPS94, CK91, Cra90, Cra91a, GV92, GL10, GBDB97, Gro91, HHCS95, Heu90, Hir91, IEE90b, Ins92, Ing90a, Ing90b, KMR^{+}97, Kik93, KLN90, KKY99, KVK92, Lev94, LSZ92, LHHJ91, LMK94, MCH96, Mic93b, MSZ90, Nar95, PSG03, PSC^{+}93a, PDS^{+}93, RS09a, Sof93, Sat97, She91, Utt90, WHL95, Bel90a, Bel90b, Fah94, GR92, HKK^{+}91a, SSW91, Yan94a]. System-Harray [YYM93]. System/3090 [SSW91]. System/390 [GR92]. System/6000 [IBM93, Bel90a, Bel90b]. Systematic [KK95b]. Systematical [NJ94c]. Systemes [II90]. Systems [Ame97a, Ano94a, BPG94, BD91, BBG^{+}93, BMMN94, Cas89a, CC92a, Che92, CFGG94, dCH94, DR93a, DR93b, FBZ92, FGCG94, Ger94a, HC92, HC94, HBB^{+}95, Hig91, HR92, Hun00, IEE94g, JL93, KZ94a, KZ94b, MKFB92, MR93b, MR95a, RFC90, Rit90, SM03, SS96, TOML04, Tho97b, DR95b, vDSP96, AR06, Ano91c, AM90, BBB^{+}94, Bar92, BB02, Coo95, Dig93a, Dig93b, Dig93c, DR94a, DR94b, DR95a, Duf04, EO91, FH90, GBC92, HS10, IEC90, ISO90, IDVV97, JC93, Kea92, KNOR04, KS12, Lef93, Lie94a, Lie94b, LHW01, Nat92, PZA93, Ple93, SS10, SMSW06, Wag94, WSW00, ZA93]. systems-using [GBC92]. Systolic [MKC92]. Szeged [Cse99].

T [Adl93, AB93, Ano98b, Gar93, Kon94, Loz98, Yan94b, Gil01, SAC^{+}92]. T-Series [SAC^{+}92]. T3D [MWO95, Oed93, SZG95]. T3E [PSG03]. table [Car91a]. tables [DI90]. TAE [Cen91]. tails [EO94]. TAKE [vK94]. Talk [Zim02]. talks [Sch93b]. Taming [DH12, Sal95]. Tangent [GK06]. tangye [nY90]. TAPENADE [PH06]. Targeting [BC19]. Task [CFK^{+}94, Fos94, FKKC96, Fox94, GOS94, OP98a, RSB97, YKK96, CMVZ94, KY98a, KY98b, OPP00, PQ94, RFRH96, SSOG93]. Tasking [KaM10]. Tasks [OP98b, DRST03, SV95]. taxonomy [LR91]. TaylUR [vH06, vH07, vH10]. TC2 [BT01, Boi97]. TC2/WG2.5 [BT01, Boi97]. Tcl [AG95a]. Tcl/Tk [AG95a]. TCP [Ano93b, JA92]. TCP/IP [Ano93b, JA92]. teach [Mat90]. Teaching [Ein96, Fur93, Mei96, Tre91]. Technical [Ano95b, Bru96a, DHP02, KRY90, KK90, Hew91b, MMG00]. Technique [AMKS02, SR04, BK89, HC08]. Techniques [Adv98, BGLP94, BMMN94, Cro91, DP99, FB12, GS01a, Jon93, KLW93, NNON02, PSC93b, Tal91, TIUG90, BPG94, CGS94, GDS94, GB92, MKF95, NBC92, Pet91]. Technology [Ano96b, Ano97d, Ano97c, Bra97d, Bra97c, Ins91a, HS94b, HS94a, IEE92a, IEE93b, Sci92, TBG^{+}02, Ame97b, ABC^{+}96, Don95, IEC94, IEC97, IEC98a, IEC98b, IEC99, Ins91b, Ins92, IEE94c, II91, ISO94, Int97a, Int97b, Int98a, Int98b, Int99, ISO00, Int00, ISO04a, ISO04b, ISO10, Ken94a, Kor99, ZCP95]. Telescoping [CMKH03]. temperature [Cra95, Kut92]. Template [BFKS93b, BFKS93a, Vil94]. Template-Driven [BFKS93b, BFKS93a]. templates [CZM93b, CMZ93a]. temporal [Azi23]. Ten [BHMS91a, BHMS91b]. Tennessee [IEE94d]. Tension [Ren97b, Ren96b, Ren09]. Tensor [Bou97, DLW^{+}18, Gep90, Num05]. TenXpert [Ano96b]. Terabytes [IEE02]. Terms [Ano93h]. TERRACE [Phi91b, Phi92]. terracing [Phi91b, Phi92]. terrain [Lop90]. Test [CV94, Cod90a, Cod93a, Cod93b, DGL91b, DDHD90, DCHH88b, Pry99, Sil01, DFRR91, Gil01, Kah01, LS09, Lin90, Mac96b, NJ94a, PBU95, RP95]. Testing [AS97, DG94, DG99, GKKL19, HP95a, KO91, MGH81, SD90, SB91, SFB92, SWM95, Sil01, SB01, GOT03a]. Tests [RB99, GH18, PSPE94]. TETRA [BH92]. Tetrachoric [BH92]. Teukolsky [Adl93, AB93, Gar93, Loz98, Yan94b]. Texas [Ano94i, IEE92c, IEE93c, IEE94b]. text [Ano07]. textual [CB94]. TFLOPS [SMSY02]. Their [CZM94a, UZCZ97, Ano93j, BRH90, CMZ94a, CMZ95, LK93a, MKS^{+}96, SKM94, Yam95]. Them [Pif96]. Theology [NRS92]. theories [Cah90]. Theory [Ano94i, BCH^{+}06, Gao06, KDG99, MC94, MC95b, SH09, U.S01b, vV90, AAS93, BW12, Gao05, MC95a, PRS99, GAW96a, GAW96b]. therapy [MKF95]. thermal [EN96]. thermodynamic [KRY90]. these [Met92b]. thick [Dut94]. thin [Mir90, VLLY92]. thin-walled [VLLY92]. Thinking [WSL94]. thinning [SHCP91]. Third [BPG94, PRS99, AI90, AH92, BV94]. Thompson [Ano98a]. Thoughts [Tay97]. thread [GOT03b]. thread-safe [GOT03b]. Threading [TBG^{+}02]. Threads [HBG01, HBG02]. Three [CLiN^{+}02, Fat94, Ogi02, SMSY02, Tho97b, Eli98, GMHC92, Heu90, Lai92a, Lai92b, PMHC92, SWO92, VLLY92]. three- [Lai92a, Lai92b]. Three-Dimensional [CLiN^{+}02, Ogi02, SMSY02, Eli98, GMHC92, Heu90, PMHC92, SWO92, VLLY92]. Thresholds [MC92]. Thrust [FYR99]. Tight [DCR99a]. Tight-Binding [DCR99a]. Tim [DeT90]. Time [ASS95, Ano93b, DCZ96, EL97, FJ92, KNS95b, Mit97, OP98b, PH96, Sch93b, SS96, AFAS99, AF92, CMP02, CB95, DNS98, DN04, FCHE02, HE13, HM93, Hor23, Kay90, KNS95a, KYSV^{+}15, MA09, NY91, NK94, PQ94, PW93, RP95, SM92, Sat97, SJ94, SZ90, SPM^{+}94, Shi98, SG95, Tor10, YSVM^{+}16, YSMA^{+}17, YSMBA23, Ano94o]. time-dependent [AFAS99, KYSV^{+}15, MA09, YSVM^{+}16, YSMA^{+}17, YSMBA23]. time-domain [HE13]. time-step [NY91]. time-varying [HM93]. Time/Run [DCZ96]. Timings [Bra97b]. Tiny [Gla92b]. Tiny-Ninety [Gla92b]. Tk [AG95a]. TN [DT94]. TNO [DS02]. TNPACK [SF92]. TNSPackage [DLW^{+}18]. today [IEE94c, Pre93g]. Toeplitz [HC92, HC94]. Together [Bru96a, Bru96b]. Tokyo [WN90]. Toledo [IEE92b]. Tolman [Rib02]. tomographic [Tur93]. tomography [NJ94a, NJ94b]. tomorrow [IEE94c, Pre93g]. TOMP [Kra94]. Tool [Bla00, BZ94, Bru96a, DG94, HKTW94, HIM91, LK93b, Liv91, RVV^{+}92, SD99, SPF00, SF10, Str05, UNF^{+}08, VBB18, Ano95g, AGG^{+}97, Bru96b, CTS96, CJPA94, DDcMR96, EJLC97, HHK^{+}93, Kon92, Lov92, LCC^{+}03, Mil91, Nai17, SSG97, YB13]. Toolbox [Ano97c, Bra97c, EP92, RPG^{+}20]. Toolkit [AG95a, Ano96b, LJO05, PHHF94a, Sar94, LJO05]. Tools [BC01, BCC^{+}96a, BCC^{+}96b, CT90, HHLS90, Hug96, KP91, Paz96, BCC^{+}97a, BCC^{+}97b, CSS90a, CSS90b, CSS91, DT94, EO91, Fos95, IJCL96, KNOR04, LMJC96, dLJEB95, Met99c, Met99d, OJ09, ST90]. Toolset [Ano97b, HGG93]. Top [Cip00, Lew94, ABMS94, Cas14, DG94]. TOPOVEL [Tur93]. Toronto [BGG^{+}94, GGK^{+}93]. Tortoise [Wei94]. total [Fu90]. Touch [Coc03]. Townsend [DT94]. TR [Int98a, Int98b]. TR2 [IEC98a, IEC98b]. TR92225 [Fox91a]. Trac [U.S01a, U.S01b, U.S01c]. Trac-M [U.S01a, U.S01b, U.S01c]. Trac-M/Fortran [U.S01a, U.S01b, U.S01c]. TRACE [SS93, LFK^{+}93, Nie92, SJ94]. trace-element [Nie92]. Traces [HMW91, HMW93, SJ94]. Tracing [DP94]. Tracking [EN96]. tradeoffs [AJJF14]. training [dSZP92]. traitement [II90]. TRANS4 [Dut94]. TransactNet [Ano96b]. Transfer [SR04, KT94]. Transfers [Mra94]. Transform [DLM99b, DLM99a, DL97c, GGLM88, GL90, SM95, Kir02, MH91, SRM90, Sul91, Wie99]. Transformation [BZ94, MA18, Rhe93, FTPR04, RD91, Van22]. transformational [vWAH^{+}02]. Transformations [BGS94a, SM94, SKP91, SD99, BK89, Sar97]. Transforming [SWW90, BN97, Che90]. transforms [GHSJ94, WJ94]. transient [Ple93]. transition [NDSG07]. transitional [Dut94]. Translating [Tee90]. Translation [Bai92, Bai93a, Bai93b, BW96, Pre93a, Sar91, SD01, Zag16, AJJF14, Bar94, DP99, Mar92, O’K93, OPE^{+}95, SD03, Lor19].

Translator

[DP96, Goo90e, Goo90f, GOBG^{+}94, KMBK96, Lev97, O’K93, CD92, Lee90, Lev95b, Mai91, OPE^{+}95, SOP93, Nob90]. Transparent [Jéz93]. transport [Car93, KRY90, PFS^{+}04]. transportable [Cen91]. Transputer [Fer92, FK95, AH90]. Transputers [BLT94, ARB94, ARB95, CA92]. TransTOOL [BCC^{+}97a, DDcMR96, BCC^{+}96b]. trap [KYSV^{+}15, MA09, YSVM^{+}16]. traps [TS06b]. travel [CB95]. Traveling [PR91]. Travelling [CT95]. Treatment [Ric06]. Tree [Ano96b, ADB94]. trend [KSM95]. trend-analysis [KSM95]. Trends [Duv92, SFB92]. Triangles [BE92, Esp98]. Triangulation [Ren97a, CCW04, Ren96a]. Tridiagonal [DGR92, CK91, DGR90]. TRIP [GL10]. TRIPACK [Ren96a]. Troy [SS96]. truly [KT94]. Truncated [SF92, KDG99]. trust [GT03, GT07]. Tsai [Gho01]. TSPACK [Ren09]. TTUTIL [Rap90]. Tucson [IEE94g]. Tuned [Lin93, Per93]. Tuning [Ano93b, Bel90a, Bel90b, IBM93, Int92, Yan94a]. tunnel [Lin90, MFI^{+}94]. Turbo [RR92]. turbulence [KMJ^{+}23]. Turkel [NY91]. Turkel-Zwas [NY91]. turning [Mil92]. Tutorial [ECS96, Pas95, WW93, Smi92, Smi93a]. tutorials [Met99c, Met99d, San92]. Twentieth [ACM93c]. Twenty [ERS95, HS94b, HS94a, MS94]. Twenty-Eighth [ERS95]. Twenty-Seventh [HS94b, HS94a, MS94]. twin [WCN92]. TWINAN90 [WCN92]. Two [CM98, EP87, Mar90, Ram90, Rei97, Shi93b, BB07, CWB92, CA90, CB95, Gao06, GF95a, GST02a, GST02b, Gou93, KY94, Ngu91, NVFNP93, Ren96a, SNK06, Taq16, YK90]. two-dimensional [CA90, Gao06, GF95a, Gou93, Ren96a]. two-hole [Taq16]. two-level [SNK06]. two-particle [Taq16]. two-pass [KY94]. two-way [CB95]. TWYTT [CB95]. Type [CMKH03, Kea95a, Kea96a, Kea96b, SGMS97, Coo94, IEC98b, Int98b, Pre99, RD91, RMX05, Van22]. Type-Driven [CMKH03]. typed [RD91]. Types [PMM^{+}08, Wal00, CM91, MKS^{+}96, ORO15, SKM94].

U.S. [FKL94]. Uhlig [Ano97a, Hop97, TDMC97]. UK [Bar92, HK93a, Fri94, HK95]. Ultra [Car91b, KO94]. Ultra-high [KO94]. Ultrahigh [Car92]. Ultrahigh-performance [Car92]. ULTRIX [Ano91c]. UMFPACK [MFK09]. UML [NCMF15, NL19]. Uncertain [Gil91b, Gil91a, KE93]. uncommenting [GG95]. uncommon [Mac96a]. Unconstrained [Bou97, Buc94a, Buc94b, Kea95b, MGH81, NS92, GOT03a, Hop98]. Understand [Scixxa, Scixxb]. Understanding [BF93a, ZB94a, ZB94b, BF93b]. Undocumented [Ham95a]. Unexpected [CHT92]. UNICOS [Cra91b, Cra92]. unidimensional [SH09]. Unidimensionality [Nan93c]. Unified [CFH^{+}93, HBD^{+}93]. uniform [KB94]. Unimodular [SM94]. Uniprocessors [Kar96]. unit [Smi93b]. Unitary [ARS92, ARS94]. United [Boi97, Olv91]. Units [CRDO16, ORO15]. Units-of-Measure [CRDO16, ORO15]. Univ [AB93, Ano91a]. univariate [Kay90]. universal [Cum90, MZT90]. University [FH90, IEE95b, NBC92, Sen03, Bee01g, Bee01f, Bee01e, Cam13]. Unix [Ano93b, Che90, HM90b, Lou90, Phi91b, Phi92]. Unknown [Pre94a]. UNPACK [BR96]. unrealized [VKB93]. unsteady [Azi23, Ude91]. unsteady-state [Ude91]. Unstructured [MR95b, PDS^{+}93, SM17, SM02b, SM03]. Unsymmetric [DR93a, DR93b]. unum [dVdVI97]. unweighted [GH18, GHN19]. up-to-date [Din99]. Update [Car90, Nar95]. Updated [TOML04, MBGK11]. updates [Ano92b]. Updating [RG90a, RG90b]. Upgrade [Ano96b]. Upgrading [Red95, GMC96a]. upward [McG91]. upward-continued [McG91]. uravneniĭna [AZ90]. US$49.95 [AB93, Gon01]. USA [IEE96, ACM96a, ACM97, ACM98, Agr95, Ano94i, Ano94p, BBG^{+}95, Ban93, BGNP94, HS94b, HS94a, Hua96, IEE94g, IEE94d, IEE02, Kar95, PBG^{+}95, Sen03, SS96, USE94]. usable [KT94]. Usage [SF92, HW95, Mol12, dL12]. Use [Bru96a, HHLS90, HK93b, HK93a, HK95, LK93a, Schxx, Ste95a, Wal00, BK89, Bru96b, Cah90, Cre90b, FKL94, MKS^{+}96, MWM90, NH09, Tre91]. used [KDG99, Wri90b]. Useful [SG93a, SG93d]. User [And92a, ABB^{+}95, BBB^{+}94, CMZ93b, CZ90, Con92, CFPS94, dCH94, HKS91, IMS90a, Lib90b, IMS91f, IMS91d, IMS91e, IMS91g, IMS91h, MSZ90, Ngu91, Scixxb, Smi95a, Sou91a, Sou91b, U.S01c, WHL92a, WHL92b, ZT90, Ano91c, Bak91, CSS90a, CSS90b, CSS91, Cur94, Dig93c, Hor09, Int90b, Int90d, Int90e, Jor90a, Jor90b, LMJC96, Par94, PSC^{+}95, Sof93, Uni93]. User-friendly [CFPS94]. user-specified [PSC^{+}95]. Users [CKZ93, IMS90b, LMK94, Shi98, Sun92a]. uses [BOPC05]. Using [AMC98, AMGM20, AG95a, Ano90a, AHOK02, BBZ95, Bee01d, Bee01g, Bee01f, Bee01e, Ben99a, BM99, Bou97, BCC^{+}96a, BCC^{+}96b, BH90, CLiN^{+}02, Chi91, CL94, DL97b, DG99, Don91, DV92, Fah94, For97, HBG01, Her90, HFT94, HFT97, KT00, LK93b, Lev97, LZ97, Mat90, MR95a, NLE^{+}20, Nan93c, NRK98, PFS^{+}04, PPR97, PHD^{+}95, Pre93a, RRM^{+}15, RFS98, SZM98, SD92, TR96, Vio90, YKK96, Ben00, BKK94, Blu91, BL91, Bra94a, BID95, BCC^{+}97a, BCC^{+}97b, BW96, CF90, CRS90, CK86, CC98, CDGM96, CA92, CFPS94, Dan90, DDcMR96, DS97, Don90, Dot93, DH95, Eli98, Err06, FGBN19, FPR01, GBC92, Gou93, GHSJ94, HHK^{+}93, Han92, Has06, HHLS90, Hop03, KTP^{+}24, KMJ^{+}23, KY98a, KY98b, Kea92, KMT91, KS12, KVK92, LP05, LN91, MH91, McG91, Ogi02]. using [RBD^{+}10, RBD^{+}11, Rei93, RPL96, RR99, RD91, SM02a, Sav95, SOP93, SS10, SD93, SSS99, VSH91, WO96, WTW90, XHY^{+}24, Yan95, Yu01, YB13, ZMR^{+}91]. Utah [Bee01g, Bee01f, Bee01e]. Utility [OC94, Pra90, Rap90]. utilizing [Cra95]. Utrecht [Ano93q]. UX [TOML04].

V [Ede90, Mal91, MMEH08, Zei92, How91, SH91]. V/STOL [How91, SH91]. v1.0 [CA92, HM12]. v1.1 [BRdAHK04]. V1.8.0 [Cod90a]. V2 [MAH^{+}02, TOML04]. v2.5 [Hew01]. v5.5 [Bee01a]. VA [Ano94d, Wie94]. Valarrays [Ano99c]. valence [MCA17]. Validated [Cse99]. Validation [AAS93, BMV03, Yan95]. Value [BG97, Cas89a, CC92a, EP87, vHKS94a, HKS94, Hig91, McB06, vKS94, vHKS94b, BG94, FT03, Gil01, IDVV97]. Value-Based [vHKS94a, HKS94, vKS94, vHKS94b]. Valued [Cos97a, Cos97b]. Values [BBCH95, BD90, McB06, SB01, Som07, EC13]. VAPP [BV94]. variability [FHE95]. Variable [Mey01, Sch99, Sch03, van90b, Cou97]. Variable-Length [Cou97]. Variables [Maaxx, CCK90, NVFNP93, Str05, vV90]. Variably [Ros93, TOC18]. Variance [KKH10, Mra94]. Variants [DS94]. variational [ZZN94]. VARIATM [LN91]. variogram [KDG99]. Various [Don91, DV92, AC16, Don90]. Varying [IEC94, Int00, HM93, ISO94]. VASE [JBBH93]. VAST [Int90e, Pre93a, Pre93b, Van94b]. VAST-2 [Int90e]. VAST-90 [Pre93a]. VAST-HPF [Van94b]. VAST/77to90 [Pre93b]. VAST/77toHPF [Van94b]. VAX [She91, Dig90a, Dig90b, Dig93a, Dig93b, Dig93c, Mac90, Phi91b, Phi92, VKB93, Vio90, Wei91a, Wei91b, Wei93, Wei91c]. Vector [BV94, Che92, DDP94, GPHL90, KZ94a, Kul95, KZ94b, LP19, LHH^{+}91, MSC96, ONT95, PAK^{+}90, Sab95, SAC^{+}92, Sul91, TIUG90, CK90, CTS96, CK91, KSZ90, NIY^{+}94, Pet91, SSS99, Swa84]. Vector-Pipeline [Che92]. VECTORFORTH [Rod90]. Vectorial [MDD94]. vectorised [GS98, KSYE00]. Vectorizable [TYJ92]. Vectorization [Che92, KO90, Ove91]. vectorized [FSV90, Heu90]. vectorizers [Fu90]. vectorizing [LCD91, VKB93]. Vectors [TR96, GMF18]. velocity [Tur93]. vent [Coo94]. VENTCF2 [Coo94]. ventricle [VLLY92]. Verification [NI03, AK93, AFKL04]. verified [KNOR04, Wal93a]. Verlag [Ano97a]. Version [Hud91c, IBM91a, IBM91b, IBM91e, Int91e, IBM93, KM90, Num91a, Pas95, Scixxb, Sch99, Sch97, U.S01a, U.S01b, U.S01c, AI90, And02, BG94, CZ90, dCH94, FGJB19, Hud91b, Int90f, Int90g, Int90h, Int90l, Int90i, Int90j, Int90k, Int90m, Int90n, Int91a, Int91b, IBM91c, IBM91d, Int91c, Int91d, Int91f, JCL10, NS11, SSG^{+}18, She91, Sto93, VKB93, WRL90, ZT90, ZBC^{+}92, van90a, vH07, Hig92, Hig94a, Hig94b, Hig94c, Hig94d, Met99c, Met99d, Ano03]. Versions [CFGG94, FGCG94, GK06, BDOS95a, BDOS95b]. Versus [BH90, Vel97, CF90, Coc03]. vertex [FTPR04, NH09]. vertical [CNP91]. Very [USE94, CCJ93]. Vesta [CFPS94]. Vetterling [Adl93, AB93, Gar93, Kon94, Loz98, Yan94b]. VF [CK90, Lin93, Per93, Pet91]. VFC [Ben99c]. VHLL [USE94]. VI [Ano94a, BV94]. Via [FKKC96, Aki99, BDH^{+}05, Das06, EDA96, Hig93a, KK94, Paj90, RP12, SS99]. viable [LD87]. Vibrations [Cap98]. Vienna [Ben95, Ben99c, BGS94b, CMZ91, CMZ92a, CMZ92b, CMMZ93, CZM93a, CMZ93b, Cha94b, Fah94, UZCZ97, ZBC^{+}92, Zim92, ZCMM93, ZBC94]. Vienna-Fortran [UZCZ97]. Vienna-Fortran/HPF [UZCZ97]. Views [Coc03]. Violation [DGS08]. Virginia [IEE94a]. Virtual [BGLP94, BKP93, Bod94, DPZ97, Ger98a, Ger98b, RS09a]. viscoelastic [FYR99, YRF02]. viscoelastic-gravitational [FYR99, YRF02]. viscous [OM92]. Visionaries [Tay86]. Visions [BCWWB94]. Visual [DL97c, ED99, Law01, Tre97, Kon92, Taq16, Nag01, Sco93]. Visualization [Ano97b, HM96, JBBH93, KGV97, BCC^{+}97b, SWO92, Utt90]. visualizer [KC94]. Visualizing [KMS^{+}95, SZ91]. VLIW [SS93]. VMS [Dig93a, Dig93b, Dig93c, Ing90a, Ing90b, KK90]. VMS/Fortran [KK90]. Voigt [Zag16, ZA11]. Vol [HS94a, MS94]. Vol.II [HS94b]. Volterra [BB91]. Volume [Ano93b, Loz98, BSS92, BDH^{+}05, PSG03]. volumetric [MKF95]. Voronoi [Ren97a, Tip91]. Vorzuge [Rei92a]. Vote [WS94]. Vote-Counting [WS94]. vous [Ain90, Ain91, Ain93]. VPP [AHOK02, ISKvW02]. vs [CB94, GS01b, Irv91, Kro90, SC19, Int90f, Int90g, Int90h, Int90l, Int90i, Int90j, Int90k, Int90m, Int90n, Int90e, Int91a, Int91b, IBM91a, IBM91b, IBM91e, IBM91c, IBM91d, Int91e, Int91c, Int91d, Int91f]. VTDIRECT95 [HWS09, SWH15]. Vvedenie [Mal91]. vychislitel′ [Mal91].

W [Adl93, Aik07, AB93, Kon94, Loh07, Mol12, Yan94b, BBCH95, KF92a]. W-Function [BBCH95]. W. [GMC96f, GMC96e]. W/Engineering [KF92a]. Wabi [Ano97b].

Wagenauslauflisten [Kru90a]. Wagener [Ano98b]. Wailea [HS94b, HS94a]. walled [Mir90, VLLY92]. Walter [Ano98b]. WAPR [BBCH95]. Ward [DeT90]. Warner [Bjø08]. Warp [Tam95]. Washington [IEE94f]. ‘Wat [Cam13]. WATCOM [CS90c, CSS90a, CSS90b, CS91, CSS91, HW91, Ano93j, Zei92]. Water [FKL94, NY91, Ste90, ZZN94]. Waterloo [Cam13]. wave [PG10, Sar00, Sar17, TYJ92]. wavelet [Sul91]. waves [KK90, NVC96, YK90, YB92]. Way [Del98, Cre03, CB95, Mas92b]. WC1E [Eme94]. Weakly [CJL97]. Weather [GK06, MA18, RHH96, SWBO93, Wic99, CDF^{+}93, GS95]. Web [Ano96b, BMO90, Ano97d, Bra97d, CH96, Teo01]. WebThreads [Ano97b]. Weeks [GGLM88, GL90]. Weibull [Gho01, Tsa01]. weight [BB07]. weighted [Dem07, GH18, GHN19, RR92, ZBW07]. Weinman [Ano90b]. Well [HD93, Sch93c]. Westview [Ano03]. WetC3D [Bak91]. wetland [Pel93]. WG10.3 [CGS94, DR94a]. WG2.5 [BT01, Boi97]. WG5-N1122 [ANS95]. WG56-N1122 [W^{+}95]. wheat [Sto93]. Which [BC97, CB94, How91, SH91]. WHILE [TLS91]. Who [Koo90]. wide [PG10]. wideband [MIN^{+}95]. Wiley [Ano96a, Ano98a, Ano99a, Ano99b, Gon01, Olv91]. Will [Kef92, VJ97a, VJ97b]. Willé [Ano96a]. William [Ano98a, Gar93, Loz98, Tha93, Wu93, Loz98]. Wind [MFI^{+}94, LN91, Lin90]. Window [AG95a]. windowed [NIY^{+}94]. Windows [Ano93d, Ano02, Hol94, Hor09, Law01, Mic93b, Nor91, Par94, Pic94, Rib92, RR93, TAH^{+}01, Tem96, Vai93]. Winograd [DN09]. Within [Hig90b, CG96, Deu90, PQ94, SS09, Tre97]. without [BW12, CZM93b, CMZ93a]. Wizard [Tre97]. Wolfe [LP19]. woman [Ano95d]. Work [PPW94, WNO94]. Work-efficient [PPW94]. workbook [Lem93d, MC91]. Working [Boi97, BT01, CGS94, Ein91, DR94a]. workload [Ber92]. WorkPlace [Ano97c, Bra97c]. Works [Pas95]. Workshop [PEP92, Agr95, Ano93m, BPG94, CKZ93, DT94, DW94, Fer92, FK95, HK93b, HK93a, HK95, IFI95, Kum94, PBG^{+}95, Sch93a, Sch93b, Smi95a, Wie94, Ano95g, Ban93, BGNP94, Don95, Hua96]. Workstation [AOL94a, AOL94b, KC94, Num91a]. Workstations [Bau93, Coe94b, BID95, DOSW96, Lan93a, SR95]. World [HR92, SIOS02]. WRAPGEN [Bru96a, Bru96b]. Wrapi [Sar94]. Wrapper [AS14, FCHE02]. Write [Dec93, See04, Cah90]. Writer [Ano97b]. Writing [NRK98, Que00, Wes96, Ano92b]. Written [KaM10, MDD94, GJU96]. WWW2GCG [CH96].

X [AG95a, AB93, PAK^{+}90, Tay99]. X-MP [PAK^{+}90]. X-Window [AG95a]. X3.124 [Ame96]. X3.124-1985 [Ame96]. X3.124.1 [Ame96]. X3.124.1-1985 [Ame96]. X3.198 [AC92, Ame92, Ame90b]. X3.198-1992 [AC92]. X3.9 [Ame87]. X3J3 [W^{+}95]. X3J3/95 [W^{+}95, ANS95]. X3J3/95-007R1 [W^{+}95, ANS95]. XANES [AI90]. xHPF [DS97, Lev94]. XL [Int90a, Int90b, Int90c, Int90d, Sar97, Int92]. XML [Nai17]. XML/HTML [Nai17]. XSC [Wal93a, Wal93b]. xtDCR [Azi23]. xxx [DV02b, Buc94b].

Y-MP [Car92, Nag90, VSH91, Vaj92]. Y. [Tay99]. YAKL [NLBB23]. yazyke [ES93b]. year [Met99c, Met99d]. Years [Szy07, Met92b]. Yielding [Kub91a, Kub91b, Kub91c]. YMP [Car91b, HP95b]. York [Ano98a, IEE90a, Olv91]. YSCODE [FGCG94].

Z [Cok93b, FHS78, Fox79]. Z-factor [Cok93b]. zavtra [GU90]. ZERO [McG91]. zeros [Bin96]. zone [Coo94, Dut94]. zonetype [Coo94]. Zosel [Eme94, UMM94, Rag95]. zur [Por90]. Zwas [NY91].

References

ANSI:1992:FHC

[A^{+}92] Jeanne C. Adams et al. Fortran 90 Handbook: Complete ANSI/ ISO Reference. McGraw-Hill, New York, NY, USA, 1992. ISBN 0-07-000406-4. xi + 740 pp. LCCN QA76.73.F28 F67 1992.

Akutin:2004:HOM

[AAC^{+}04] Yuri Akutin, Cristina Anderson, Marius Cornea, Alexey Ershov, Eugeny Gladkov, Evgeny Gvozdev, Bob Hanek, John Harrison, Alexander Isaev, Andrey Kolesov, Alexey Kovalev, Elena Luneva, Sergey Maidanov, Andrey Naraikin, Bob Norin, Pavel Shelepugin, Vladimir Sorokin, Shane Story, and Ping Tak Peter Tang. Highly optimized mathematical functions for the IA-64 architecture. Application note 245410-011, Intel Corporation, Santa Clara, CA, USA, December 16, 2004. 14 pp.

                                                    URL ftp://download.intel.
                                                    com/software/opensource/
                                                    numerics/libm.pdf;   http:
                                                    //www.intel.com/cd/software/
                                                    products/asmo-na/eng/219868.
                                                    htm;   http://www.intel.
                                                    com/cd/software/products/

asmo-na/eng/219871.htm?prn=
y.

Atlamazoglou:2001:ALP

[AAK01] P. E. Atlamazoglou, H. T.

Anastassiu, and D. I. Kaklamani. Application of literateprogramming principles for the description of a FORTRAN 90 extension to quaternion arithmetic. IEEE Antennas and Propagation Magazine, 43(4): 104–114, August 2001. CO-DEN IAPMEZ. ISSN 1045-9243.

Abramov:1993:AEN

[AAN^{+}93] S. M. Abramov, A. I. Adamowitch, I. A. Nesterov, S. P. Pimenov, and Yu.V. Shevchuck. Autotransformation of evaluation network as a basis for automatic dynamic parallelizing. In S. Atkins and A. S. Wagner, editors, Transputer Research and Applications, NATUG-6; Proceedings of the Sixth Conference of the North American Transputer Users Group, pages 333–344. IOS Press, Amsterdam, The Netherlands, 1993.

Albrecht:1993:VNT

[AAS93] R. Albrecht, G. Alefeld, and H. J. Stetter, editors. Validation numerics: theory and applications, volume 9 of Computing. Supplementum. Springer, Wien / New York, 1993. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Ames:1993:BRNb

[AB93] W. F. Ames and C. Brezinski. Book review: Numerical recipes in Fortran (The art of scientific computing): W. H. Press, S. A. Teukolsky, W. T. Vetterling and B. P. Flannery, Cambridge Univ. Press, Cambridge, 2nd ed., 1992. 963 pp., US$49.95, ISBN 0-521-43064-X. Mathematics and Computers in Simulation, 35(5):451–452, November 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

                                        //www.sciencedirect.com/
                                        science/article/pii/037847549390043T.

Ammann:1991:PPC

[ABB^{+}91] E. M. Ammann, R. R. Berbec, G. Bozman, M. Faix, G. A. Goldrian, J. A. Pershing, Jr., J. Ruvolo-Chong, and F. Scholz. The Parallel Processing Compute Server. IBM Journal of Research and Development, 35(5/6):653–666, September/November 1991. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Averick:1994:NOA

[ABB^{+}94] B. Averick, C. Bischof, B. Bixby, A. Carle, J. Dennis, M. El-Alem, A. El-Bakry, A. Griewank, G. Johnson, R. Lewis, J. Moré, R. Tapia, V. Torczon, and K. Williamson. Numerical optimization at the Center for Research on Parallel Computation. The International Journal of Supercomputer Applications and High Performance Computing, 8(2):143– 153, Summer 1994. CODEN IJSAE9. ISSN 0890-2720.

Anderson:1995:LUG

[ABB^{+}95] E. Anderson, Z. Bai, C. Bischof, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, A. McKenney, S. Ostrouchov, and D. Sorensen. LAPACK Users’ Guide. SIAM Press, Philadelphia, PA, USA, second edition, 1995. ISBN 0-89871-345-5 (paperback). xix + 325 pp. LCCN QA76.73.F25 L36 1995.

Andre:1996:NCT

[ABC^{+}96] F. Andre, P. Brezany, O. Cheron, W. Denissen, J.-L. Pazat, and K. Sanjari. A new compiler technology for handling HPF data parallel constructs. In Szymanski and Sinharoy [SS96], pages 279–282. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Adams:1997:FHC

[ABM^{+}97] Jeanne C. Adams, Walter S.

Brainerd, Jeanne T. Martin, Brian T. Smith, and Jerrold L. Wagener, editors. Fortran 95 Handbook: Complete ISO/ANSI Reference. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, November 1997.

ISBN 0-262-51096-0. xii + 711 pp. LCCN QA76.73.F25 F6 1997. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=0262510960;   http:
                                        //www.mitpress.com/book-
                                        home.tcl?isbn=0262510960.

Adams:1994:FTN

[ABMS94] Jeanne Adams, Walt Brainerd, Jeanne Martin, and Brian Smith. Fortran top 90: Ninety Key Features of Fortran 90. Unicomp, Albuquerque, NM, USA, December 1, 1994. ISBN 0-9640135-0-9 (paperback). 252 pp. LCCN QA76.73.F25F679 1994. US$20.00.

Absoft:1991:FOF

[Abs91] Absoft Corp. FORTRAN 77 an object-oriented FORTRAN, 1991. 1 computer optical disk cartridge.

Adams:1992:FHC

[ABW92] Jeanne C. Adams, Walter S. Brainerd, and Jerrold L. Wagener. Fortran 90 Handbook: Complete ANSI/ISO Reference. McGraw-Hill, New York, NY, USA, 1992. ISBN 0-07-000406-4. xi + 740 pp. LCCN QA76.73.F28 F67 1992. US$79.50. See [Ame92].

ANSI:1992:ANSc

[AC92] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for programming language, FORTRAN — extended: ANSI X3.198-1992: ISO/IEC 1539: 1991 (E). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, September 21, 1992. xix + 369 pp.

Achee:1997:COD

[AC97a] B. L. Achee and Doris L.

Carver. Creating objectoriented designs from legacy FORTRAN code. The Journal of Systems and Software, 39(2):179–??, ???? 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Achee:1997:COO

[AC97b] B. L. Achee and Doris L.

Carver. Creating objectoriented designs from legacy FORTRAN code. The Journal of Systems and Software, 39(2):179–194, November 1997. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Ayala:2016:AFP

[AC16] César Ayala and Gorazd Cvetic. anQCD: Fortran programs for couplings at complex momenta in various analytic QCD models. Computer Physics Communications, 199(??):114–117, February 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0010465515003719.

Alvanos:2017:PMM

[AC17] Michail Alvanos and Theodoros Christoudias. MEDINA: MECCA development in accelerators — KPP Fortran to CUDA sourceto-source pre-processor. Journal of Open Research Software, 5(1):13–??, April 28, 2017. CODEN ???? ISSN 2049-9647. URL https://

                                        openresearchsoftware.metajnl.
                                        com/articles/10.5334/jors.
                                        158/.

Adve:1994:RDP

[ACG^{+}94] Vikram Adve, Alan Carle, Elana Granston, Seema Hiranandani, and Ken Kennedy. Requirements for data-parallel programming environments. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):48, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Ancourt:1997:LAF

[ACIK97] Corinne Ancourt, Fabien Coelho, François Irigoin, and Ronan Keryell. Linear algebra framework for static High Performance Fortran code distribution. Scientific Programming, 6(1):3–27, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

ACM:1991:PPE

[ACM91] ACM, editor. POPL ’91.

Proceedings of the eighteenth annual ACM symposium on Principles of programming languages, January 21–23, 1991, Orlando, FL. ACM Press, New York, NY 10036, USA, 1991. ISBN ???? LCCN ????

ACM:1993:PFA

[ACM93a] ACM, editor. Proceedings of the Fourth ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming, PPOPP: San Diego, California, May 19–22, 1993, volume 28(7) of ACM SIG-PLAN Notices. ACM Press, New York, NY 10036, USA, July 1993. CODEN SINODQ. ISBN 0-89791-589-5. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7 .S54 v.28:7.

ACM:1993:PAS

[ACM93b] Proceedings of the ACM SIG-PLAN ’93 Conference on Programming Language Design and Implementation: Albuquerque, New Mexico, June 23– 25, 1993, SIGPLAN notices; ISSN: 0362-1340; v. 28, no. 6 (June 1993). ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-598-4. LCCN QA76.7.A152 v. 28, no. 6.

ACM:1993:CRT

[ACM93c] Conference record of the Twentieth Annual ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the symposium, Charleston, South Carolina, January 10–13, 1993. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993. ACM order number 549930.

ACM:1994:CPI

[ACM94a] ACM, editor. Conference Proceedings. 1994 International Conference on Supercomputing. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-665-4. LCCN ????

ACM:1994:CRP

[ACM94b]  ACM, editor.  Conference

record of POPL ’94, 21st ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the Symposium: Portland, Oregon, January 17– 21, 1994. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-636-0. LCCN QA76.7 .A15 1994. URL

http://www.acm.org/pubs/
contents/proceedings/plan/
174675/index.html.

ACM:1994:IPI

[ACM94c] ACM, editor. ISSAC’94. Proceedings of the International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-638-7. LCCN ????

ACM:1995:CPI

[ACM95a] ACM, editor. Conference proceedings of the 1995 International Conference on Supercomputing, Barcelona, Spain, July 3–7, 1995, Conference Proceedings of the International Conference on Supercomputing. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995.

ACM:1995:CRP

[ACM95b]  ACM, editor.  Conference

record of POPL ’95, 22nd ACM SIGPLAN-SIGACT Symposium on Principles of Programming Languages: papers presented at the Symposium: San Francisco, California, January 22–25, 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-692-1. LCCN QA 76.7 A11 1995.

URL http://www.acm.org/
pubs/contents/proceedings/
plan/199448/index.html. ACM

order number: 549950.

ACM:1996:FCP

[ACM96a] ACM, editor. FCRC ’96:

Conference proceedings of the 1996 International Conference on Supercomputing: Philadelphia, Pennsylvania, USA, May 25–28, 1996. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

ACM:1996:SCP

[ACM96b] ACM, editor. Supercomputing ’96 Conference Proceedings: November 17–22, Pittsburgh, PA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-89791-854-1. LCCN QA 76.88 S8573 1996. URL

http://www.supercomp.org/
sc96/proceedings/. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

ACM:1997:SHP

[ACM97] ACM, editor. SC’97: High Performance Networking and Computing: Proceedings of the 1997 ACM/IEEE SC97 Conference: November 15– 21, 1997, San Jose, California, USA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-89791-985-8. LCCN ???? URL

http://www.supercomp.org/
sc97/proceedings/. ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

ACM:1998:SHP

[ACM98] ACM, editor. SC’98: High Performance Networking and Computing: Proceedings of the 1998 ACM/IEEE SC98 Conference: Orange County Convention Center, Orlando, Florida, USA, November 7–13, 1998. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN ???? LCCN ????

URL http://www.supercomp.
org/sc98/papers/.

ACM:1999:SPO

[ACM99] ACM, editor. SC’99: Oregon Convention Center 777 NE Martin Luther King Jr. Boulevard, Portland, Oregon, November 11–18, 1999. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999.

ACM:2001:PAJ

[ACM01]  ACM, editor.  Proceed-

ings of the ACM 2001 Java Grande/ISCOPE Conference: Palo Alto, Calif., June 2– 4, 2001. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-359-6. LCCN QA76.9.O35 A26 2001.

ACM:2003:SII

[ACM03] ACM, editor. SC2003: Igniting Innovation. Phoenix, AZ, November 15–21, 2003. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 1-58113-695-1. LCCN ????

Antonuccio-Delogu:1994:PTN

[ADB94] V. Antonuccio-Delogu and U. Becciani. A parallel tree N-body code for heterogeneous clusters. In Dongarra and Wasniewski [DW94], pages 17–32. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

Amestoy:2004:AAA

[ADD04] Patrick R. Amestoy, Timothy A. Davis, and Iain S. Duff. Algorithm 837: AMD, an approximate minimum degree ordering algorithm. ACM Transactions on Mathematical Software, 30(3):381–388, September 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Averbuch:1996:PPF

[ADG96] A. Averbuch, R. Dekel, and E. Gabber. Portable parallelizing Fortran compiler. Concurrency, practice and experience, 8(2):91–123, March 1996. CODEN CPEXEI. ISSN 1040-3108. URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        23268.

Appelbe:1995:NAG

[ADH95] B. Appelbe, S. Doddapaneni, and C. Hardnett. A new algorithm for global optimization for parallelism and locality. In Pingali et al. [PBG^{+}95], pages 125–140. ISBN 3-540-58868-X. LCCN QA76.58 .W656 1994.

Akarsu:1996:PCS

[ADHF96] Erol Akarsu, Kivanc Dincer, Tomasz Haupt, and Geoffrey C. Fox. Particle-in-cell simulation codes in High Performance Fortran. In ACM [ACM96b], page ?? ISBN 0-89791-854-1. LCCN QA 76.88 S8573 1996. URL http:

//www.supercomp.org/sc96/
proceedings/SC96PROC/AKARSU/

INDEX.HTM. ACM Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Adler:1993:NRF

[Adl93] Fred Adler. Numerical Recipes in FORTRAN: The Art of Scientific Computation (W. H. Press, S. A. Teukolsky, W. T. Vetterling, and B. P. Flannery). Mathematical Biosciences, 118(1):123– ??, November 1993. CODEN MABIAR. ISSN 0025-5564.

Adve:1998:HPF

[Adv98] Vikram S. Adve. High Performance Fortran compilation techniques for parallelizing scientific codes. In ACM [ACM98], page ?? ISBN ???? LCCN ???? URL http:

//www.supercomp.org/sc98/
papers/.

Agrawal:1996:RSP

[AES^{+}96] G. Agrawal, G. Edjlali, A. Sussman, J. Humphries, and J. Saltz. Runtime support for programming in adaptive parallel environments. In Szymanski and Sinharoy [SS96], pages 241–252. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Andrew:1992:SGC

[AF92] K. Andrew and C. G. Fleming. Space-time geometries characterized by solutions to the geodesic equations. Computers in physics, 6(5):498– 505, September–October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Alexandrov:1999:PDO

[AFAS99] V. Alexandrov, I. Filippov, L. Adamowicz, and W. R. Salzman. Passing dynamic objects by reference in Fortran 90: application to time-dependent quantum chemistry. Computers and Chemistry, 23(1):25– ??, ???? 1999. CODEN COCHDK. ISSN 0097-8485.

Ariskin:1993:CFP

[AFBN93] Alexei A. Ariskin, Mikhail Ya. Frenkel, Galina S. Barmina, and Roger L. Nielsen. CO-MAGMAT: a FORTRAN program to model magma differentiation processes. Computers and Geosciences, 19(8):1155– ??, September 1993. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Rene:2004:NSR

[AFKL04] René Alt, Andreas Frommer, R. Baker Kearfott, and Wolfram Luther, editors. Numerical software with result verification. International Dagstuhl seminar, Dagstuhl Castle, Germany, January 19–24, 2003. Revised papers. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-21260-4 (paperback).

Andre:1995:PDC

[AFMP95] F. Andre, M. Le Fur, Y. Maheo, and J.-L. Pazat. The Pandore data-parallel compiler and its portable runtime. In Hertzberger and Serazzi [HS95], pages 176–183. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Averbukh:1994:RA

[AFS94] Victoria Z. Averbukh, Samuel Figueroa, and Tamar Schlick. Remark on algorithm 566. ACM Transactions on Mathematical Software, 20(3):282– 285, September 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [MGH81].

Alexander:1995:HCX

[AG95a] P. Alexander and L. F. Gladden. How to create an Xwindow interface to Gnuplot and Fortran programs using the Tcl/Tk toolkit. Computers in physics, 9(1):57–??, 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Ambastha:1995:PCP

[AG95b] Anil K. Ambastha and Edmond Gomes. Program calculates pseudoskin in a multilayer reservoir. The Oil and Gas Journal, 93(24):121–??, June 1995. CODEN OIGJAV. ISSN 0030-1388.

Ayguade:1997:DRT

[AGG^{+}97] Eduard Ayguade, Jordi Garcia, Merce Girones, M. Luz Grande, and Jesus Labarta. DDT: a research tool for automatic data distribution in High Performance Fortran. Scientific Programming, 6(1):73– 94, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Agrawal:1995:PIW

[Agr95] D. P. Agrawal, editor. Proceedings of the 1995 ICPP Workshop on Challenges for Parallel Processing, August 14, 1995, Raleigh, NC, USA. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1995. ISBN 0-8493-2618-4. LCCN QA76.58.I34 1995.

Andreev:1992:FM

[AGS92] V. M. Andreev, I. S. Golosov, and S. V. Sprogis. Fortran for MPP. Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 17(5):281–289, July 1992. CODEN PCSODA. ISSN 0361-7688 (print), 1608-3261 (electronic).

Agterberg:1994:FPA

[Agt94] Frederik Agterberg. FOR-TRAN program for the analysis of point patterns with correction for edge effects. Computers and Geosciences, 20(2): 229–??, March 1994. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Allan:1990:FAP

[AH90] R. J. Allan and L. Heck. Fortnet: a parallel FORTRAN harness for porting application codes to transputer arrays. In L. Freeman and C. Phillips, editors, Applications of Transputers, pages 82–89. IOS Press, Amsterdam, The Netherlands, 1990.

Anik:1991:PIS

[AH91] S. Anik and W. W. Hwu. Performance implications of synchronization support for parallel Fortran programs. Technical Report CRHC-91-21, University of Illinois, Urbana-Champaign, Center for Reliable and High-Performance Computing, Urbana, IL, USA, 1991.

Atanassova:1992:CAE

[AH92] Lidiya Atanassova and Jurgen Herzberger, editors. Computer arithmetic and enclosure methods: proceedings of the third International IMACS-GAMM Symposium on Computer Arithmetic and Scientific Computing (SCAN-91), Oldenburg, Germany, 1–4 October 1991. North-Holland Publishing Co., Amsterdam, The Netherlands, 1992. ISBN 0-444-89834-4. LCCN QA76.9.C62I559 1992.

Anik:1994:PIS

[AH94] S. Anik and W.-M. W. Hwu. Performance implications of synchronization support for parallel Fortran programs. Journal of Parallel and Distributed Computing, 22(2):202– 215, August 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1994.1081/production;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1994.1081/production/pdf.

Ashauer:1990:SFC

[AHJS90] R. Ashauer, T. Hoppe, G. Jost, and K. Solchenbach. The Suprenum Fortran compiler

— architecture and performance. Supercomputer, 7(3): 19–25, May 1990. CODEN SP-COEL. ISSN 0168-7875.

Asaoka:2002:EHJ

[AHOK02] Kae Asaoka, Akio Hirano, Yasuo Okabe, and Masanori Kanazawa. Evaluation of the HPF/JA extensions on Fujitsu VPP using the NAS parallel benchmarks. Lecture Notes in Computer Science, 2327:503– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2327/23270503.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270503.
pdf.

Allan:1990:PFS

[AHZ90] R. J. Allan, L. Heck, and S. Zurek. Parallel FORTRAN in scientific computing: a new occam harness called Fortnet. Computer Physics Communications, 59(2):325–344, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
001046559090182Z.

Aldea:1990:FAE

[AI90] Nicolae Aldea and Emil Indrea. Fourier analysis of EXAFS and XANES data — a self-contained FOR-TRAN program-package — the third version. Computer Physics Communications, 60(1):145–154, August 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
001046559090083D.

Aiken:2007:MJW

[Aik07] Alex Aiken. In memoriam:

John W. Backus, 1924–2007. ;login: the USENIX Association newsletter, 32(3):68– 69, June 2007. CODEN LOGNEM. ISSN 1044-6397.

Ain:1990:SPF

[Ain90] Maryse Ain. Savez-vous parler Fortran?: cours progressif de programmation structurée en Fortran 77, cours et exercices. Publications universitaires scientifiques. Orsay Plus, Orsay, France, 1990. ISBN 2-87800-000-5. 260 pp.

Ain:1991:SPF

[Ain91] Maryse Ain. Savez-vous parler Fortran?: cours progressif de programmation structurée en Fortran 77, cours et exercices. Publications universitaires scientifiques. Orsay Plus, Orsay, France, 1991. ISBN 2-87800-000-5. 280 pp.

Ain:1993:SPF

[Ain93] Maryse Ain. Savez-vous parler Fortran? Bibliothèque des universités. Informatique. Paris Sud. De Boeck-Wesmael, Bruxelles, Belgium, 1993. ISBN 2-8041-1755-3. 501 pp.

Aires:2004:GFP

[Air04] Nibia Aires. A guide to the Fortran programs to calculate inclusion probabilities for conditional Poisson sampling and Pareto πps sampling designs. Computational Statistics, 19 (3):337–345, September 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/BF03372100.

Akushevich:1997:PFC

[AIS^{+}97] I. Akushevich, A. Ilyichev, N. Shumeiko, A. Soroko, and A. Tolkachev. POLRAD 2.0. FORTRAN code for the radiative corrections calculation to deep inelastic scattering of polarized particles. Computer Physics Communications, 104(1–3):201–244, August 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465597000623.

Atkinson:1998:AAB

[AJ98] Kendall Atkinson and Youngmok Jeon. Algorithm 787: Automatic boundary integral equation programs for the planar Laplace equation. ACM Transactions on Mathematical Software, 24(4):395–417, December 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Arabas:2014:FTB

[AJJF14] Sylwester Arabas, Dorota Jarecka, Anna Jaruga, and Maciej Fijalkowski. Formula translation in Blitz++, NumPy and modern Fortran: A case study of the language choice tradeoffs. Scientific Programming, 22(3):201–222, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Allen:1984:ALI

[AK84] John R. Allen and Ken Kennedy. Automatic loop interchange. ACM SIGPLAN Notices, 19(6):233–246, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adams:1993:SCA

[AK93] E. Adams and U. Kulisch, editors. Scientific computing with automatic result verification, volume 189 of Mathematics in science and engineering. Academic Press, New York, NY, USA, 1993. ISBN 0-12-044210-8. LCCN QA76.S368 1993.

Akima:1996:ASS

[Aki96] Hiroshi Akima. Algorithm 761: scattered-data surface fitting that has the accuracy of a cubic polynomial. ACM Transactions on Mathematical Software, 22(3):362–371, September 1996. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/toms/
                                        1996-22-3/p362-akima/. See

remarks [RB98, DV00].

Akin:1999:NOO

[Aki99] J. E. Akin. Number 1: Object oriented programming via Fortran 90. Engineering Computations, 16(1):26–??, ???? 1999. CODEN ENCOEN. ISSN 0264-4401.

Albert:1988:CFA

[AKLS88] Eugene Albert, Kathleen Knobe, Joan D. Lukas, and Guy L. Steele, Jr. Compiling Fortran 8x array features for the Connection Machine computer system. ACM SIGPLAN Notices, 23 (9):42–56, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/62115/p42-albert/.

Aberti:1992:FIP

[AL92] C. Aberti and Patrice Lignelet.

Fortran 90: initiation à partir du Fortran 77. Serie informatique. Studio image, Menton, France, 1992. ISBN ???? 144 pp. LCCN ????

Algonquin:1990:FL

[Alg90] Algonquin College.Continuing Education.Business Dept. FOR-TRAN language. Media Algonquin, Nepean, Ontario, Canada, 1990. ISBN 0-88620-847-5. 13 + 36 + [2] pp.

Allison:1990:IMC

[All90] Bob Allison. Interfacing Microsoft C and FORTRAN. The Journal of C Language Translation, 1(4):300–304, March 1990. ISSN 1042-5721.

Allan:1993:TPE

[All93]  R. J. Allan.  Towards a

portable environment for FOR-TRAN applications on parallel computers. Theoretica Chimica Acta, 84(4):257–??, January 1993. CODEN TCHAAM. ISSN 0040-5744.

Alonso:2004:JFR

[Alo04] Andrés M. Alonso. JMASM10: a Fortran routine for sieve bootstrap prediction intervals. Journal of Modern Applied Statistical Methods, 3(1):239–249, May 2004. CODEN ???? ISSN 1538-9472.

Albert:1991:DPC

[ALS91] Eugene Albert, Joan D. Lukas, and Guy L. Steele, Jr. Data parallel computers and the FORALL statement. Journal of Parallel and Distributed Computing, 13(2):185–192, October 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Altin:1990:EPS

[Alt90]  Vural Altin.  Engineering

programming in structured Fortran. Technical report, Bogazici University, Istanbul, Turkey, 1990. 394 pp.

Ashrafiuon:1990:AOD

[AM90] H. Ashrafiuon and N. K. Mani. Analysis and optimal design of spatial mechanical systems. Journal of Mechanisms, Transmissions, and Automation in Design, 112(2):200–207, June 1990. CODEN JMTDDK. ISSN 0738-0666.

Adve:1998:UIS

[AMC98] Vikram Adve and John Mellor-Crummey. Using integer sets for data-parallel program analysis and optimization. ACM SIGPLAN Notices, 33(5):186–198, May 1998. CODEN SINODQ. ISBN 0-89791-987-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/277650/p186-adve/.

Adve:2001:CAC

[AMC01] Vikram Adve and John Mellor-Crummey. Chapter 16. advanced code generation for High Performance Fortran. Lecture Notes in Computer Science, 1808:553–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1808/18080553.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1808/18080553.
pdf.

ANSI:1987:DPA

[Ame87] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed ANSI Fortran X3.9– 198x, September 18, 1987. See also [MR87].

Amenta:1990:IFP

[Ame90a] Roddy V. Amenta. An interactive FORTRAN program for cross-correlation of signals on a PC with CGA graphics: an application in marine geoacoustics. Computers and Geosciences, 16(1):75– ??, 1990. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

ANSI:1990:DPA

[Ame90b] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed American National Standard Programming Language Fortran Extended X3.198–199x, September 24, 1990. See also [MR87].

ANSI:ftn92

[Ame92] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. American National Standard Programming Language Fortran Extended X3.198–1992, 1992. This US Standard is identical to the international standard, ISO 1539:1991. See also [ABW92, BGA94, Cou91, MR92].

ANSI:1996:AXR

[Ame96] American National Standards Institute. ANSI X3.124.1-1985 (R1996): Graphical Kernel System (GKS) FORTRAN Binding (included in ANSI X3.124-1985). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1996. ISBN ???? ???? pp. LCCN ???? US$32.

URL http://www.ansi.org/.

ANSI:1997:AIR

[Ame97a] American National Standards Institute. ANSI/ISO 9593-1-1990 (R1997): Information Processing Systems — Computer Graphics — Programmer’s Hierarchical Interactive Graphics System (PHIGS) Language Bindings (Part 1: FORTRAN). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1997. ISBN ???? ???? pp. LCCN ???? US$80.

URL http://www.ansi.org/.

ANSI:1997:AII

[Ame97b] American National Standards Institute. ANSI/ISO/IEC 1539-1:1997: Information technology — Programming languages — Fortran — Part 1: Base language. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1997. ISBN ???? ???? pp. LCCN ???? US$18.00.

URL http://webstore.ansi.
org/ansidocstore/product.
asp?sku=ANSI%2FISO%2FIEC+
1539%2D1%3A1997; http://
www.fortran.com/fortran/
iso1539.html.

Aguirre-Mesa:2020:MLC

[AMGM20] Andres M. Aguirre-Mesa, Manuel J. Garcia, and Harry Millwater. MultiZ: a library for computation of high-order derivatives using multicomplex or multidual numbers. ACM Transactions on Mathematical Software, 46(3):23:1–23:30, September 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL https://dl.acm.org/
doi/10.1145/3378538.

Araki:2002:OHP

[AMKS02] Takuya Araki, Hitoshi Murai, Tsunehiko Kamachi, and Yoshiki Seo. Optimization of HPF programs with dynamic recompilation technique. Lecture Notes in Computer Science, 2327:551–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2327/23270551.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270551.
pdf.

Amos:1990:APF

[Amo90] Donald E. Amos. Algorithm 683: a portable FORTRAN subroutine for exponential integrals of a complex argument. ACM Transactions on Mathematical Software, 16(2): 178–182, June 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0098-
3500/78934.html.

Anantbaraman:1993:GEF

[Ana93a] Vinod Anantbaraman. A GUI environment for Fortran development. Dr. Dobb’s Journal of Software Tools, 18(6):104–??, June 1993. CODEN DDJOEB. ISSN 1044-789X. Anantharaman:1993:GEF

[Ana93b] Vinod Anantharaman. A GUI environment for Fortran development. Dr. Dobb’s Journal of Software Tools, 18(6):104, 106– 108, 110–111, June 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Anderson:1990:MIO

[And90] Oliver D. Anderson. Mastering input/output in Fortran 77. Interface (Santa Cruz), 12(??): 53–??, Winter 1990. CODEN INFCDB. ISSN 0163-6626.

Anderson:1992:LUG

[And92a] E. Anderson. LAPACK Users’ Guide. SIAM Press, Philadelphia, PA, USA, May 1992. ISBN 0-89871-294-7. xv + 235 pp. LCCN QA76.73.F25 L36 1992. Includes a Quick Reference Guide.

Anderson:1992:PGI

[And92b] Oliver D. Anderson. A pedagogic guide to input/output in Fortran 77. Interface (Santa Cruz), 14(??):27–34, Spring 1992. CODEN INFCDB. ISSN 0163-6626.

Anderson:2002:LFE

[And02]  Edward Anderson.  LA-

PACK3E — a Fortran 90enhanced version of LA-PACK. LAPACK Working Note 158, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, December 2002. URL http:

//www.netlib.org/lapack/
lawns/lawn158.ps;   http:
//www.netlib.org/lapack/
lawnspdf/lawn158.pdf. UT-

CS-02-497, December 2002.

Anonymous:1990:BFP

[Ano90a]  Anonymous.  BASINMAT:

FORTRAN program calculates oil and gas generation using a distribution of discrete activation energies. Geobyte, 5(2): 37–??, April 1990. ISSN 0885-6362.

Anonymous:1990:BRFb

[Ano90b] Anonymous. Book review: Fortran for scientists and engineers: D. G. Weinman, PWS-Kent, Boston, 1989. 531 pp., US $14.95, ISBN 0-87835-337-

2. Mathematics and Computers in Simulation, 32(4):423, October 1990. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/037847549090160K.

Anonymous:1990:FE

[Ano90c] Anonymous. FORTRAN 8X explained. Computer applications in the biosciences: CABIOS, 6(4):417–??, 1990. CODEN COABER. ISSN 0266-7061.

Anonymous:1991:BRFe

[Ano91a] Anonymous. Book review:

Fortran 90 explained: M. Metcalf and J. Reid, Oxford Univ. Press, Oxford, 1990. 294 pp., £14.95, ISBN 0-19-853772-7. Mathematics and Computers in Simulation, 33(1):83–84, July 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/037847549190068E.

Anonymous:1991:CLP

[Ano91b]  Anonymous.  Classic lan-

guage, part 1: FORTRAN. BYTE Magazine, 16(9):147– 152, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:DFU

[Ano91c] Anonymous. DEC Fortran for ULTRIX RISC systems: user manual. Digital Equipment Corp., Maynard, MA, USA, 1991. various pp.

Anonymous:1991:DFL

[Ano91d] Anonymous. DEC Fortran:

language reference manual. Digital Equipment Corp., Maynard, MA, USA, 1991. various pp.

Anonymous:1991:FFD

[Ano91e] Anonymous. Forever Fortran:

Digital language renewed. Science News, 140(2):23, July 1991. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Anonymous:1991:MFR

[Ano91f] Anonymous. Microsoft FOR-TRAN reference. Microsoft Press, Bellevue, WA, USA, version 5.1 edition, 1991. xxv + 534 pp.

Anonymous:1992:F

[Ano92a]  Anonymous.  Fortran 90.

Chemical engineer, ??(527):25– ??, September 1992. CODEN CMERA9. ISSN 0302-0797.

Anonymous:1992:LUF

[Ano92b] Anonymous. Lahey updates Fortran compiler to allow writing 32-bit DOS apps. InfoWorld, 14(12):20–21, March 1992. CODEN INWODU. ISSN 0199-6649.

Anonymous:1992:MSA

[Ano92c] Anonymous. Math software automates problem solving. Mechanical Engineering, 114(5): 58–63, May 1992. CODEN MEENAH. ISSN 0025-6501.

Anonymous:1992:MF

[Ano92d] Anonymous. Microsoft FOR-TRAN, 1992.

Anonymous:1992:NRE

[Ano92e] Anonymous. Numerical recipes: example book (FORTRAN). Cambridge University Press, Cambridge, UK, second edition, 1992. ISBN 0-521-43721-0. viii + 245 pp. LCCN QA76.73.F25N84 1992.

Anonymous:1993:CSN

[Ano93a] Anonymous. Computing in science: News reports. Science, 261(5123):841–??, August 1993. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1993:CPR

[Ano93b] Anonymous. Constructing Predictable Real-Time System; Unix for Fortran Programmers; Unix Curses Explained; Internetworking with TCP/IP, Volume I: Principles, Protocols, and Architecture; System Performance Tuning; New Media. IEEE Software, 10(3):106–111, May 1993. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1993:FFS

[Ano93c] Anonymous. Fortran 90 features in subset High Performance Fortran. ACM SIG-PLAN FORTRAN Forum, 12 (4):151–??, December 1, 1993. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:FPC

[Ano93d] Anonymous. Fortran powerstation combines Windows with 32-bit DOS execution. EDN, 38 (6):136–??, March 1993. CO-DEN EDNSBH. ISSN 0012-7515, 0364-6637.

Anonymous:1993:GSH

[Ano93e] Anonymous. Goals and scope of High Performance Fortran.

ACM SIGPLAN FORTRAN Forum, 12(4):1–??, December 1, 1993. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:HPFc

[Ano93f]  Anonymous.  High Perfor-

mance Fortran. Scientific Programming, 2(1–2):1–170, Spring–Summer 1993. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:1993:HPFa

[Ano93g] Anonymous. High Performance Fortran language specification (part I). ACM SIGPLAN FORTRAN Forum, 12(4):1–86, December 1, 1993. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:HPFb

[Ano93h] Anonymous. High Performance Fortran terms and concepts. ACM SIGPLAN FORTRAN Forum, 12(4):169–??, December 1, 1993. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:JD

[Ano93i] Anonymous. Journal of development. Scientific Programming, 2(1-2):A1–44, Spring– Summer 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:1993:MWE

[Ano93j] Anonymous. Microsoft and Watcom expect their FOR-TRAN compilers to boost the speed of executables. PC Week, 10(10):67–??, March 1993. ISSN 0740-1604.

Anonymous:1993:NFH

[Ano93k] Anonymous. New features in High Performance Fortran. ACM SIGPLAN FORTRAN Forum, 12(4):3–??, December 1, 1993. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1993:NN

[Ano93l] Anonymous. News and notes.

Journal of Scientific and Industrial Research, 52(12):837– ??, December 1993. CODEN JSIRAC. ISSN 0022-4456.

Anonymous:1993:PIW

[Ano93m] Anonymous, editor. Proceedings of International Workshop on Languages and Compilers for Parallel Computing. Min. Sci. and Technol, Jerusalem, Israel, 1993. ISBN ???? LCCN ????

Anonymous:1993:PSE

[Ano93n]  Anonymous, editor.  Pro-

ceedings. SHARE Europe Anniversary Meeting Client/ Server — the Promise and the Reality: October 25-28, 1993, the Hague, the Netherlands. SHARE Europe, Carouge/Geneva, Switzerland, 1993. ISBN ???? ISSN 0254-6213. LCCN ????

Anonymous:1993:PF

[Ano93o] Anonymous. Programmieren in Fortran. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., March 1993. ISBN 3-211-82446-4. US$88.50. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=3211824464.

Anonymous:1993:SEF

[Ano93p] Anonymous. Software-Entwicklung in Fortran 90. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., March 1993. ISBN 3-211-82450-2. US$25.01. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=3211824502.

Anonymous:1993:SEC

[Ano93q] Anonymous, editor. Supercomputing Europe ’93 conference papers: fifth international exhibition and conference highperformance computing, 22– 24 February, 1993, Utrecht, The Netherlands. Royal Dutch Fairs, Utrecht, Netherlands, 1993. ISBN ???? LCCN ????

Anonymous:1994:AVP

[Ano94a] Anonymous, editor. ASPLOS-VI proceedings / Sixth International Conference on Architectural Support for Programming Languages and Operating Systems, San Jose, California, October 4–7, 1994, volume 29(11) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, November 1994. CODEN SINODQ. ISBN 0-89791-660-3. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.9.A73 I558 1994.

Anonymous:1994:C

[Ano94b] Anonymous. Corrigenda. ACM Transactions on Mathematical Software, 20(4):553, December 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [HFT94].

Anonymous:1994:EC

[Ano94c] Anonymous. From the Editorin-Chief. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3): 2–??, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1994:HPC

[Ano94d] Anonymous, editor. High performance computing and communications 1st Symposium High performance computing and communications symposium: – March 1994, Alexandria, VA. ARPA-CSTO, Arlington, VA, USA, 1994. ISBN ???? LCCN ????

Anonymous:1994:HPFa

[Ano94e] Anonymous. High Performance Fortran language specification (part II). ACM SIGPLAN FORTRAN Forum, 13(2):87– 150, June 1, 1994. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:HPFb

[Ano94f] Anonymous. High Performance Fortran language specification (part III). ACM SIGPLAN FORTRAN Forum, 13(3):22– 55, September 1, 1994. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:HR

[Ano94g] Anonymous. HPF reports. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):71, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1994:IPH

[Ano94h] Anonymous. Interpreting the performance of HPF/Fortran 90D. In IEEE [IEE94f], pages 743–752. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:ISL

[Ano94i] Anonymous, editor. Lattice 93: proceedings of the International Symposium on Lattice Field Theory, Dallas Texas, USA, 12–16 October 1993, volume 34 of Nuclear Physics B, Proceedings Supplements. North-Holland Publishing Co., Amsterdam, The Netherlands, April 1994. CODEN NPBSE7. ISBN ???? ISSN 0920-5632 (print), 1873-3832 (electronic). LCCN ????

Anonymous:1994:MMI

[Ano94j] Anonymous. MPI: a messagepassing interface standard. The International Journal of Supercomputer Applications and High Performance Computing, 8(3/4):159–416, Fall/Winter 1994. CODEN IJSAE9. ISSN 0890-2720.

Anonymous:1994:MMP

[Ano94k] Anonymous. MPI: a messagepassing interface standard. The International Journal of Supercomputer Applications and High Performance Computing, 8(3/4):159–416, Fall-Winter 1994. CODEN IJSAE9. ISSN 0890-2720.

Anonymous:1994:PLC

[Ano94l] Anonymous, editor. Plasmadynamics and lasers: 25th Conference — June 1994, Colorado Springs, CO, number 94 in Papers — American Institute of Aeronautics and Astronautics 1994; 2262. American Institute of Aeronautics and Astronautics, 370 L’Enfant Promenade SW, Washington, DC 20024– 2518, 1994. ISBN ???? LCCN ????

Anonymous:1994:SIH

[Ano94m] Anonymous. Special issue:

High Performance Fortran language specification. ACM SIG-PLAN FORTRAN Forum, 13 (2), June 1, 1994. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIO

[Ano94n] Anonymous. Standards information. organization of Standards Committees. ACM SIG-PLAN FORTRAN Forum, 13 (1):4–??, March 1, 1994. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SIOa

[Ano94o] Anonymous. Standards information. organization of Standards Committees; Real-Time Fortran Standard; Fortran 90 interpretations; plans for Fortran 95; High Performance Fortran language specification. ACM SIGPLAN FORTRAN Forum, 13(1):4–6, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:SPF

[Ano94p] Anonymous, editor. Symposium on Parallel Finite Element Computations: October 25-27, 1993, Supercomputer Institute, Minneapolis, MN, USA, volume 119(1–2) of Computer Methods in Applied Mechanics and Engineering. Elsevier, Amsterdam, The Netherlands, November 1994. CODEN CMMECC. ISBN ???? ISSN 0045-7825, 0374-2830. LCCN ????

Anonymous:1995:MJP

[Ano95a] Anonymous, editor. 186th Meeting: — June 1995, Pittsburgh, PA, volume 27(2) of Bulletin — American Astronomical Society. American Institute of Physics, Woodbury, NY, USA, 1995. ISBN ???? ISSN 0002-7537. LCCN ????

Anonymous:1995:STS

[Ano95b]  Anonymous, editor.  26th

SIGCSE Technical Symposium on Computer Science Education, volume 27(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, March 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:1995:FAS

[Ano95c] Anonymous, editor. Fifth ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming (PPoPP), Santa Barbara, CA, July 19– 21, 1995, volume 30(8) of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, August 1995. CODEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN ????

Anonymous:1995:HGA

[Ano95d]  Anonymous.  How to get

ahead as a woman in industry, standard problems in Fortran. Physics world, 8(7):17– ??, July 1995. CODEN PH-WOEW. ISSN 0953-8585.

Anonymous:1995:MF

[Ano95e] Anonymous. Migrating to Fortran 90. IEEE Software, 12 (4):109–??, 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1995:SS

[Ano95f] Anonymous. Software in the spotlight. ACM SIGPLAN FORTRAN Forum, 14(4):4–??, December 1, 1995. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:SHW

[Ano95g] Anonymous. SunSoft hits the workshop with new tool set for C++ and Fortran programmers. Network World, 12(19): 38–??, May 1995. ISSN 0887-7661 (print), 1944-7655 (electronic).

Anonymous:1996:BRAa

[Ano96a] Anonymous. Book review: Advanced scientific Fortran: By David R. Willé. John Wiley & Sons, Chichester. (1995). 234 pages. $24.95. Computers and Mathematics with Applications, 31(2):128, January 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122196900772.

Anonymous:1996:NPA

[Ano96b] Anonymous. New products: Absoft Fortran 77 for Linux; Spyglass Client Web Technology Kit in Red Hat Linux; BLAST Communications Software for Linux; TenXpert CD Server Upgrade; TransactNet Web Interface Toolkit; 32 bit ODBC Driver for C-Tree Plus; Phonetics Data Remote. Linux Journal, 31:??, November 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1996:SFP

[Ano96c] Anonymous. Sample F programs. Fortran Journal, 8(6): ??, November/December 1996. ISSN 1060-0221. URL http:/

/www.fortran.com/fortran/
FJ/9611/#samples.

Anonymous:1997:BRNc

[Ano97a] Anonymous. Book review: Numerical algorithms with Fortran: By Gisela Engeln– Müllges and Frank Uhlig. Springer-Verlag, Heidelberg, Germany. (1996). 602 pages. DM 68.00, sFr 60.00 (CD-ROM included). Computers and Mathematics with Applications, 33(4):132, February 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122197900390.

Anonymous:1997:NPW

[Ano97b]  Anonymous.  New prod-

ucts: WebThreads 1.0.1; QUERYFLEX Report Writer; Linux Pro Desktop 1.0; NDP Fortran for Linux; Numerics and Visualization for Java; Craftworks Linux/AXP 2.2; InfoDock Linux Software Development Toolset; Caldera Wabi 2.2 for Linux. Linux Journal, 34:??, February 1997. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1997:TNR

[Ano97c] Anonymous. Technology news & reviews: Chemkin software; OpenMP Fortran Standard; ODE toolbox for Matlab; Java products; Scientific WorkPlace 3.0. IEEE Computational Science & Engineering, 4(4):75–??, October/December 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1997/pdf/
c4075.pdf.

Anonymous:1997:TNF

[Ano97d] Anonymous. Technology news: Fortran 90 news; free software; symbolic computing packages; Matlab 5; Web products. IEEE Computational Science & Engineering, 4(1):87–??, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1997/pdf/
c1087.pdf.

Anonymous:1998:BRAk

[Ano98a] Anonymous. Book review: Atlas for computing mathematical functions: an illustrated guide for practitioners, with programs in Fortran 90 and Mathematica: By William J. Thompson. John Wiley & Sons, Inc., New York. (1997). $145.00 (CD-ROM included). Computers and Mathematics with Applications, 36(8):125, October 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122198911671.

Anonymous:1998:BRFb

[Ano98b] Anonymous. Book review:

Fortran 95 handbook: Complete ISO/ANSI reference: By Jeanne C. Adams, Walter S. Brainerd, Jeanne T. Martin, Brain T. Smith and Jerrold L. Wagener. MIT Press, Cambridge, MA. (1997). 711 pages. $50.00. Computers and Mathematics with Applications, 35(9):142, May 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122198907052.

Anonymous:1998:FC

[Ano98c] Anonymous. Fortran 95 compiler. Computers in physics, 12(4):355–??, July 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.168695.

Anonymous:1999:BRCb

[Ano99a] Anonymous. Book review:

Computing for scientists: Principles of programming with Fortran 90 and C++: By R. J. Barlow and A. R. Barnett. John Wiley & Sons, Inc., Chichester. (1998). 276 pages. $22.50. Computers and Mathematics with Applications, 37(3):136, February 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122199904376.

Anonymous:1999:BRCd

[Ano99b] Anonymous. Book review: Computing for scientists: Principles of programming with Fortran 90 and C++: By R. J. Barlow and A. R. Barnett. John Wiley & Sons, Inc., Chichester. (1998). 276 pages. $22.50. Computers and Mathematics with Applications, 37(6):136, March 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0898122199903012.

Anonymous:1999:CFC

[Ano99c] Anonymous. Comparing Fortran 90 and C++ valarrays. C/C++ Users Journal, 17(3): ??, March 1999. CODEN CCU-JEX. ISSN 1075-2838.

Anonymous:19xx:CFI

[Anoxx]  Anonymous.  Conversione

FORTRAN IV 7040/FORTRAN H 360. ????, 19xx. 16 pp.

Anonymous:2002:OAI

[Ano02] Anonymous. Optimizing applications with the Intel C++ and Fortran compilers for Windows and Linux. Technical report, Hewlett–Packard Corporation, Palo Alto, CA, USA, 2002. URL http:

//www.intel.com/software/
products/compilers/c60/techtopics/
Compiler_Optimization_6.
pdf.

Anonymous:2003:BRCf

[Ano03] Anonymous. Book review:

Computational physics: FOR-TRAN version: By Steven E. Koonin and Dawn C. Meredith. Westview Press, Boulder, CO. (1990). 639 pages. $65. Computers and Mathematics with Applications, 46 (2–3):502, July/August 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122103900435.

Anonymous:2007:IRT

[Ano07] Anonymous. Index of resources: text: Fortran. Computer History Museum Web archive., 2007. URL https:/

/archive.computerhistory.
org/resources/text/Fortran/

.

X3J3:1995:SIF

[ANS95] ANSI Accredited Technical Subcommittee X3J3. Special issue: Fortran 95, committee draft, May 1995 (X3J3/ 95-007r1, SC33/WG5-N1122). ACM SIGPLAN FORTRAN Forum, 14(2):xviii + 354, June 1995. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Avenarius:1990:ALP

[AO90a] Adrian Avenarius and Siegfried Oppermann. FWEB: a literate programming system for Fortran 8X. ACM SIGPLAN Notices, 25(1):52–58, January 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Avenarius:1990:FLP

[AO90b] Adrian Avenarius and Siegfried Oppermann. FWEB: a literate programming system for Fortran 8X. ACM SIGPLAN Notices, 25(1):52–58, January 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Avenarius:fortran-web

[AO90c] Adrian Avenarius and Siegfried Oppermann. FWEB: a literate programming system for Fortran 8X. ACM SIGPLAN Notices, 25(1):52–58, January 1990. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Annaratone:1994:DEC

[AOL94a] M. Annaratone, C. D. Offner, and D. B. Loveman. Digital Equipment Corporation — High Performance Fortran on workstation farms. In Siegel [Sie94b], pages 664–669. ISBN 0-8186-5602-6, 0-8186-5601-8. ISSN 1063-7133. LCCN QA 76.58 I56 1994.

Annaratone:1994:HPF

[AOL94b] M. Annaratone, C. D. Offner, and D. B. Loveman. High performance Fortran on workstation farms. In Siegal [Sie94a], pages 664–669. ISBN 0-8186-5602-6. LCCN QA76.58.I58 1994. IEEE catalog no. 94CH34819.

Apiola:1990:IAS

[AP90] H. Apiola and P. Peltola. Integrating APL with symbol manipulation, numerical software and graphics. APL Quote Quad, 20(4):10–17, July 1990. CODEN APLQD9. ISSN 0163-6006.

Appleby:1991:CLP

[App91] D. Appleby. Classic language, part 1: FORTRAN: First in a series on language survivors. BYTE Magazine, 16(9):147– 148, 150, September 1991. CO-DEN BYTEDJ. ISSN 0360-5280.

Amodio:2006:ABF

[AR06] Pierluigi Amodio and Giuseppe Romanazzi. Algorithm 859: BABDCR—a Fortran 90 package for the solution of bordered ABD linear systems. ACM Transactions on Mathematical Software, 32(4):597–608, December 2006. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Asenov:1994:SSI

[ARB94] A. Asenov, D. Reid, and J. R.

Barker. Speed-up of scalable iterative linear solvers implemented on an array of transputers. Parallel Computing, 20 (3):375–387, March 10, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1994&
volume=20&issue=3&aid=832.

Asenov:1995:SSI

[ARB95] A. Asenov, D. Reid, and J. R. Barker. Speed-up of scalable iterative linear solvers implemented on an array of transputers. Parallel Computing, 21 (4):669–682, April 1, 1995. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1995&
volume=21&issue=4&aid=931.

Arenius:1990:FIF

[Are90] Kristopher Arenius. FOR-TRAN input functions. The VAX professional, 12(6):29–??, December 1990. CODEN VAX-PEN. ISSN 8750-9628.

Ammar:1992:IDC

[ARS92] G. S. Ammar, L. Reichel, and D. C. Sorensen. An implementation of a divide and conquer algorithm for the unitary eigenproblem. ACM Transactions on Mathematical Software, 18(3):292–307, September 1992. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1992-18-3/p292-ammar/. See

also [ARS94].

Ammar:1994:CAI

[ARS94] G. S. Ammar, L. Reichel, and D. C. Sorensen. Corrigendum: “Algorithm 730: An implementation of a divide and conquer algorithm for the unitary eigenproblem”. ACM Transactions on Mathematical Software, 20(1):161, March 1994. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1994-20-1/p161-ammar/. See

[ARS92].

Angus:1991:ECA

[AS91] I. G. Angus and Janice L.

Stolzy. Experiences converting an application from Fortran to C++: Beyond f2c. In C++ at Work Conference. ????, ????, November 1991.

Akian:1992:APE

[AS92] M. Akian and A. Sulem. Application of ‘Pandore’, an expert system in stochastic control, to portfolio selection problems. In Houstis and Rice [HR92], pages 389–398. ISBN 0-444-89703-8. LCCN Q334 .I45 1991.

Abernathy:1993:APC

[AS93] Roger W. Abernathy and Robert P. Smith. Algorithm 724: Program to calculate F - percentiles. ACM Transactions on Mathematical Software, 19(4):481–483, December 1993. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1993-19-4/p481-abernathy/.

Akkas:1997:ITI

[AS97] A. Akkas and M. J. Schulte. Implementing and testing interval operations and intrinsics in the GNU Fortran compiler. Technical report, Lehigh University, Bethlehem, PA, USA, December 1997.

URL http://home.ku.edu.
tr/~ahakkas/publications/
Impl_Test_Intv_Op.pdf.

Awile:2014:PWF

[AS14] Omar Awile and Ivo F. Sbalzarini. A Pthreads wrapper for Fortran 2003. ACM Transactions on Mathematical Software, 40(3):19:1–19:15, April 2014. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Ashworth:1981:PP

[Ash81] Robert Ashworth. Prettyprinting for power (FORTRAN). ACM SIGPLAN Notices, 16 (2):16–17, February 1981. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Amamiya:1994:RPL

[ASM^{+}94] M. Amamiya, M. Satoh, A. Makinouchi, K. Hagiwara, T. Yuasa, H. Aida, K. Ueda, K. Araki, T. Ida, and T. Baba. Research on programming languages for massively parallel processing. In Horiguchi et al. [HHK94], pages 443–450. ISBN 0-8186-6507-6. LCCN QA76.58 .I5673 1994. IEEE catalog no. 94TH0697-3.

Agrawal:1993:CRS

[ASS93] Gagan Agrawal, Alan Sussman, and Joel Saltz. Compiler and runtime support for structured and block structured applications. In IEEE [IEE93d], pages 578–587. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Agrawal:1995:IRC

[ASS95] Gagan Agrawal, Alan Sussman, and Joel Saltz. An integrated runtime and compiletime approach for parallelizing structured and block structured applications. IEEE Transactions on Parallel and Distributed Systems, 6(7):747– 754, July 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td1995/l0747abs.
htm.

Alsdorf:1994:FPP

[Av94] Douglas E. Alsdorf and Ralph R. B. von Frese. FORTRAN programs to process Magsat data for lithospheric, external field, and residual core components. NASA technical memorandum 104612, National Aeronautics and Space Administration, Goddard Space Flight Center; National Technical Information Service, distributor, Greenbelt, MD, USA, 1994. ???? pp.

Abdelrahman:1994:DAD

[AW94] T. S. Abdelrahman and T. N. Wong. Distributed array data management on NUMA multiprocessors. In IEEE [IEE94d], pages 551–559. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994.

Arushanyan:1990:CRO

[AZ90] O. B. Arushanyan and S. F. Zalëtkin. Chislennoe reshenie obyknovennykh differentsial′nykh uravneniĭna Fortrane. (Russian) [Numerical solution of ordinary differential equations in Fortran]. Moskov. Gos. Univ., Moscow, Russia, 1990. ISBN 5-211-00957-6. 336 pp.

Amme:1998:DDA

[AZ98] Wolfram Amme and Eberhard Zehendner. Data dependence analysis in programs with pointers. Parallel Computing, 24(3–4):505– 525, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.elsevier.
com/cas/tree/store/parco/
sub/1998/24/3-4/1300.pdf.

Azis:2023:XLF

[Azi23] Mohammad Ivan Azis. 2D-xtDCR-LTBEM:

a Fortran code for numerical solutions to the unsteady anisotropic-DCR equation of spatio-temporal coefficients. SoftwareX, 23(??): ??, July 2023. CODEN ???? ISSN 2352-7110. URL

http://www.sciencedirect.
com/science/article/pii/
S2352711023002066.

Brainerd:1995:PGF

[BA95] W. S. Brainerd and J. C.

Adams. Programmer’s Guide to FORTRAN 90. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, September 1995. ISBN 0-387-94570-9. 445 pp. US$39.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0387945709.

Backus:1998:HFI

[Bac98] J. Backus. The history of Fortran I, II, and III. IEEE Annals of the History of Computing, 20 (4):68–78, October–December 1998. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/15706/00728232.pdf.

Bailey:1992:ATF

[Bai92] David H. Bailey. Automatic translation of Fortran programs to multiprecision. RNR Technical Report RNR-91-025, NAS Applied Research Branch, NASA Ames Research Center, Moffett Field, CA 94035, 17 April 1992.

Bailey:1993:AMT

[Bai93a] David H. Bailey. Algorithm 719: Multiprecision translation and execution of FOR-TRAN programs. ACM Transactions on Mathematical Software, 19(3):288–319, September 1993. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0098-3500/
                                        155767.html.

Bailey:1993:ATF

[Bai93b]  David H. Bailey.  Auto-

matic translation of Fortran programs to multiprecision. Technical Report RNR-91-025, NASA Ames Research Center, Moffett Field, CA, USA, May 1993. ???? pp.

Bailey:1994:FBM

[Bai94] David H. Bailey. A Fortran-90 based multiprecision system. Technical Report RNR-94-013, NASA Ames Research Center, Moffett Field, CA, USA, June 1994. 12 pp.

Bailey:1995:FBM

[Bai95] David H. Bailey. A Fortran-90 based multiprecision system. ACM Transactions on Mathematical Software, 21(4):379– 387, December 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
toms/1995-21-4/p379-bailey/

. See also extension to complex arithmetic [Smi98].

Bailey:2005:DFDa

[Bai05a] David H. Bailey. DDFUN90:

Fortran-90 double-double package. World-Wide Web site with software archives., March

11, 2005. URL http://crd.
lbl.gov/~dhbailey/mpdist/
;   http://crd.lbl.gov/~
dhbailey/mpdist/ddfun90.
tar.gz.

Bailey:2005:DFDb

[Bai05b] David H. Bailey. DSFUN90:

Fortran-90 double-single package. World-Wide Web site with software archives., March

11, 2005. URL http://crd.
lbl.gov/~dhbailey/mpdist/
;   http://crd.lbl.gov/~
dhbailey/mpdist/dsfun90.
tar.gz.

Bailey:2005:QDD

[Bai05c] David H. Bailey. QD: doubledouble and quad double package. World-Wide Web site with software archives., August

24, 2005. URL http://crd.

lbl.gov/~dhbailey/mpdist/
;   http://crd.lbl.gov/~
dhbailey/mpdist/qd.tar.gz.

Baker:1991:FBP

[Bak91] Lawrence M. Baker. Fortran benchmark programs WetC3D and 3DModel4 user’s guide. Open-file report 91-328, U.S. Geological Survey; Books and Open-File Reports Section distributor, Menlo Park, CA, USA, 1991. ???? pp.

Baker:1995:CF

[Bak95] Steven Baker. Complying with Fortran 90. Dr. Dobb’s Journal of Software Tools, 20(1):68, 70, 72, 74–76, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Banerjee:1993:LCP

[Ban93] Utpal Banerjee, editor. Languages and compilers for parallel computing: 5th international workshop, New Haven, Connecticut, USA, August 3– 5, 1992: proceedings, volume 757 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57502-2, 0-387-57502-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .L36 1993.

Barker:1992:SJS

[Bar92] H. A. Barker, editor. Computer aided design in control systems: selected papers from the IFAC symposium, Swansea, UK, 15– 17 July 1991, number 1 in IFAC Symposia Series. Pergamon Press, New York, NY, USA, 1992. ISBN 0-08-041269-6. LCCN TJ213 .C57 1992.

Barry:1994:EPF

[Bar94] Robert J. Barry. Extracting parallelism from Fortran by translation to a single assignment form. Thesis (M.S. in computer engineering), Southern Methodist University, Dallas, TX, USA, 1994. viii + 46 pp.

Bauer:1993:PWP

[Bau93] Barr E. Bauer. Porting from workstations to PC’s. Dr. Dobb’s Journal of Software Tools, 18(9):68, 70, 72–73, September 1993. CODEN DDJOEB. ISSN 1044-789X.

Blom:1991:ADC

[BB91] J. G. Blom and H. Brunner. Algorithm 689: Discretized collocation and iterated collocation for nonlinear Volterra integral equations of the second kind. ACM Transactions on Mathematical Software, 17(2): 167–177, June 1991. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
toms/1991-17-2/p167-blom/.

Boulet:1996:EAP

[BB96] P. Boulet and T. Brandes.

Evaluation of automatic parallelization strategies for HPF compilers. Lecture Notes in Computer Science, 1067:778– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benkner:2002:EPP

[BB02] Siegfried Benkner and Thomas Brandes. Efficient parallel programming on scalable shared memory systems with High Performance Fortran. Concurrency and Computation: Practice and Experience, 14(8–9): 789–803, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/95016125/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016125{\&}PLACEBO=IE.
pdf.

Beebe:2007:AQP

[BB07] Nelson H. F. Beebe and James S. Ball. Algorithm 867: QUADLOG—a package of routines for generating Gauss-related quadrature for two classes of logarithmic weight functions. ACM Transactions on Mathematical Software, 33(3):20:1–20:30, August 2007. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Backus:1957:FAC

[BBB^{+}57] J. W. Backus, R. J. Beeber, S. Best, R. Goldberg, L. M. Haibt, H. L. Herrick, R. A. Nelson, D. Sayre, P. B. Sheridan, H. Stern, I. Ziller, R. A. Hughes, and R. Nutt. The FORTRAN automatic coding system. In Proceedings of the Western Joint Computer Conference, February 26–28, 1957, Los Angeles, CA, USA, pages 188–198. Institute of Radio Engineers, 1 East 79th Street, New York 21, NY, USA, 1957. ISSN 0449-1173. LCCN TK7885.A1 J6. The online edition of the Oxford English Dictionary cites this as the second earliest mention of the name FORTRAN, with the extract “The programmer attended a one-day course on FORTRAN and … then programmed the job in four hours using 47 FOR-TRAN statements.”.

Bala:1994:IEU

[BBB^{+}94] Vasanth Bala, Jehoshua Bruck, Raymond Bryant, Robert Cypher, Peter de Jong, Pablo Elustondo, D. Frye, Alex Ho, Ching-Tien Ho, Gail Irwin, Shlomo Kipnis, Richard Lawrence, and Marc Snir. The IBM External User Interface for scalable parallel systems. Parallel Computing, 20(4):445– 462, March 31, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1994&
volume=20&issue=4&aid=862.

Benner:2000:AFS

[BBB00] Peter Benner, Ralph Byers, and Eric Barth. Algorithm 800: Fortran 77 subroutines for computing the eigenvalues of Hamiltonian matrices I: the square-reduced method. ACM Transactions on Mathematical Software, 26(1):49–77, March 2000. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/toms/
                                        2000-26-1/p49-benner/;
                                        http://www.acm.org/pubs/
                                        citations/journals/toms/
                                        2000-26-1/p49-benner/p49-
                                        benner.pdf.

Barry:1995:AWF

[BBCH95] D. A. Barry, S. J. Barry, and P. J. Culligan-Hensley. Algorithm 743: WAPR: A Fortran routine for calculating real values of the W-function. ACM Transactions on Mathematical Software, 21(2):172–181, June 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Brandes:1998:CBH

[BBCR98] T. Brandes, F. Bregier, M. C. Counilh, and J. Roman. Contribution to better handling of irregular problems in HPF2. Lecture Notes in Computer Science, 1470:639–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouchitte:1994:EAE

[BBDR94] V. Bouchitte, P. Boulet, A. Darte, and Y. Robert. Evaluating array expressions on massively parallel machines with communication/ computation overlap. In Buchberger and Volkert [BV94], pages 713–724. ISBN 3-540-58430-7 (Berlin), 0-387-58430-7 (New York). LCCN QA76.58 .J65 1994.

Bouchitte:1995:EAE

[BBDR95] V. Bouchitte, P. Boulet, A. Darte, and Y. Robert. Evaluating array expressions on massively parallel machines with communication/ computation overlap. The International Journal of Supercomputer Applications and High Performance Computing, 9(3):205–219, Fall 1995. CO-DEN IJSCFG. ISSN 1078-3482.

Bohm:1992:PIL

[BBF^{+}92] W. Bohm, J. C. Browne, D. W. Forslund, A. Goforth, McGraw, J. R., and K. S. Smith. Politically incorrect languages for supercomputing — a panel discussion. In IEEE [IEE92d], pages 704–706. ISBN 0-8186-2630-5. LCCN QA76.5 .S894 1992. IEEE catalog no. 92CH3216-9.

Bodin:1993:IPC

[BBG^{+}93] F. Bodin, P. Beckman, D. Gannon, S. Yang, S. Kesavan, A. Malony, and B. Mohr. Implementing a Parallel C++ runtime system for scalable parallel systems. In IEEE [IEE93d], pages 588–597. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Bailey:1995:PSS

[BBG^{+}95] D. H. Bailey, P. E. Bjorstad, J. R. Gilbert, M. V. Mascagni, R. S. Schreiber, H. D. Simon, V. J. Torczon, and L. T. Watson, editors. Proceedings of the Seventh SIAM Conference on Parallel Processing for Scientific Computing (15– 17 February, 1995, San Francisco, CA, USA). SIAM Press, Philadelphia, PA, USA, 1995. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Benkner:1994:PAS

[BBZ94] S. Benkner, P. Brezany, and H. Zima. Processing array statements and procedure interfaces in the PREPARE High Performance Fortran compiler. In Fritzson [Fri94], pages 324– 338. ISBN 0-387-57877-3 (New York), 3-540-57877-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 I52 1994.

Bartholomew-Biggs:1995:UMI

[BBZ95] M. C. Bartholomew-Biggs and S. Zakovic. Using Markov’s interval arithmetic to evaluate Bessel–Riccati functions. Numerical Algorithms, 10(3-4): 261–287, October 1995. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Briggs:1994:EPR

[BC94] Preston Briggs and Keith D.

Cooper. Effective partial redundancy elimination. ACM SIGPLAN Notices, 29(6):159– 170, June 1994. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/pldi/178243/
p159-briggs/.

Berthou:1997:WAP

[BC97] Jean-Yves Berthou and Laurent Colombet. Which approach to parallelizing scientific codes — that is the question. Parallel Computing, 23 (1–2):165–179, April 16, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1997&
volume=23&issue=1-2&aid=
1154.

Bala:2001:PCA

[BC01] Piotr Bała and Terry W. Clark.

Pfortran and Co-Array Fortran as tools for parallelization of a large-scale scientific application. Lecture Notes in Computer Science, 1900:511– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1900/19000511.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1900/19000511.
                                        pdf.

Budiardja:2019:TGO

[BC19] Reuben D. Budiardja and Christian Y. Cardall. Targeting GPUs with OpenMP directives on Summit: a simple and effective Fortran experience. Parallel Computing, 88(??):Article 102544, ???? 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167819119301358.

Bisc91f

[BCC^{+}91a] Christian Bischof, Alan Carle, George Corliss, Andreas Griewank, and Paul Hovland. Generating derivative codes from Fortran programs. Preprint MCS–P263–0991, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 S. Cass Ave., Argonne, IL 60439–4801, 1991. Also appeared as Technical Report 91185, Center for Research in Parallel Computation, Rice University, Houston, TX 77251.

Bischof:1991:GDC

[BCC^{+}91b] Christian Bischof, Alan Carle, George Corliss, Andreas Griewank, and Paul Hovland. Generating derivative codes from Fortran programs. Preprint MCS–P263–0991, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 S. Cass Ave., Argonne, IL 60439–4801, 1991. Also appeared as Technical Report 91185, Center for Research in Parallel Computation, Rice University, Houston, TX 77251.

Bischof:1992:AGD

[BCC^{+}92] Christian H. Bischof, Alan Carle, George F. Corliss, Andreas Griewank, and Paul Hovland. ADIFOR: Generating derivative code from Fortran programs. Scientific Programming, 1(1):11–29, 1992. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Brandes:1996:HSIa

[BCC^{+}96a] T. Brandes, S. Chaumette, M.-C. Counilh, A. Darte, J. Mignot, F. Desprez, and J. Roman. HPFIT: a set of integrated tools for the parallelization of applications using High Performance Fortran. Rapport de recherche 96-28, LIP ENS, Lyon, France, 1996.

Brandes:1996:HSIb

[BCC^{+}96b] T. Brandes, S. Chaumette, M.-C. Counilh, A. Darte, J. Mignot, F. Desprez, and J. Roman. HPFIT: a set of integrated tools for the parallelization of applications using high performance Fortran: Part I: HP-FIT and the Transtool environment’. In J. Dongarra and B. Tourancheau, editors, Third Workshop on Environments and Tools for Parallel Scientific Computing. SIAM Press, Philadelphia, PA, USA, 1996. ISBN ???? LCCN ????

Brandes:1997:HSIa

[BCC^{+}97a] T. Brandes, S. Chaumette, M. C. Counilh, J. Roman, A. Darte, F. Desprez, and J. C. Mignot. HPFIT: a set of integrated tools for the parallelization of applications using High Performance Fortran. Part I: HPFIT and the TransTOOL environment. Parallel Computing, 23(1–2): 71–87, April 16, 1997. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1997&
volume=23&issue=1-2&aid=
1148.

Brandes:1997:HSIb

[BCC^{+}97b] T. Brandes, S. Chaumette, M. C. Counilh, J. Roman, F. Desprez, and J. C. Mignot. HPFIT: a set of integrated tools for the parallelization of applications using High Performance Fortran. PART II: Datastructure visualization and HPF extensions for irregular problems. Parallel Computing, 23(1–2):89–105, April 16, 1997.

CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1997&
volume=23&issue=1-2&aid=
1149.

Berntsen:1993:AAA

[BCE93] Jarle Berntsen, Ronald Cools, and Terje O. Espelid. Algorithm 720: An algorithm for adaptive cubature over a collection of 3-dimensional simplices. ACM Transactions on Mathematical Software, 19(3): 320–332, September 1993. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bozkus:1993:CDD

[BCF^{+}93a] Z. Bozkus, A. Choudhary, G. Fox, T. Haupt, and S. Ranka. Compiling distribution directives in a Fortran 90D compiler. In IEEE [IEE93c], pages 617–620. ISBN 0-8186-4222-X. LCCN QA 76.58 I42 1993. IEEE catalog no. 93TH0584-3.

Bozkus:1993:FCD

[BCF^{+}93b] Z. Bozkus, A. Choudhary, G. Fox, T. Haupt, and S. Ranka. Fortran 90D/HPF compiler for distributed memory MIMD computers: Design, implementation, and performance results. In IEEE [IEE93d], pages 351–360. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Bozkus:1993:FHC

[BCF^{+}93c] Zeki Bozkus, Alok Choudhary, Geoffrey Fox, Tomasz Haupt, and Sanjay Ranka. Fortran 90D/HPF compiler for distributed memory MIMD computers: Design, implementation, and performance results. In IEEE [IEE93d], pages 351–360. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Bozkus:1994:SLF

[BCF^{+}94a] Z. Bozkus, A. Choudhary, G. Fox, T. Haupt, S. Ranka, R. Thakur, and Jhy-Chun Wang. Scalable libraries for Fortran 90D/High Performance Fortran. In IEEE [IEE94e], pages 67–76. ISBN 0-8186-4980-1. LCCN QA76.58.S34 1993.

Bozkus:1994:CFD

[BCF^{+}94b] Z. Bozkus, A. Choudhary, G. Fox, T. Haupt, S. Ranka, and Min-You Wu. Compiling Fortran 90D/HPF for distributed memory MIMD computers. Journal of Parallel and Distributed Computing, 21 (1):15–26, April 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bozkus:1994:CAF

[BCF^{+}94c] Zeki Bozkus, Alok Choudhary, Geoffrey Fox, Tomasz Haupt, and Sanjay Ranka. A compilation approach for Fortran 90D/HPF compilers. In Banerjee et al. [BGNP94], pages 200–215. ISBN 3-540-57659-2 (Berlin), 0-387-57659-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1993. DM122.00.

Bozkus:1994:CFH

[BCF^{+}94d] Zeki Bozkus, Alok Choudhary, Geoffrey Fox, Tomasz Haupt, Sanjay Ranka, and Min-You Wu. Compiling Fortran 90D/HPF for distributed memory MIMD computers. Journal of Parallel and Distributed Computing, 21 (1):15–26, April 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1994.1039/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1994.1039/production/pdf.

Bozkus:1993:CAF

[BCFH93] Z. Bozkus, A. Choudhary, G. Fox, and T. Haupt. A compilation approach for Fortran 90D/HPF compilers. In Banerjee [Ban93], pages 200–

215. ISBN 3-540-57502-2, 0-387-57502-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .L36 1993.

Bucker:2006:ADA

[BCH^{+}06] Martin Bücker, George Corliss, Paul Hovland, Uwe Naumann, and Boyana Norris, editors. Automatic Differentiation: Applications, Theory, and Implementations, volume 50 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNCSA6. ISBN 3-540-28403-6 (print), 3-540-28438-9 (ebook). ISSN 1439-7358. LCCN QA304 .I58 2006. URL http:

//link.springer.com/book/
10.1007/3-540-28438-9. Pro-

ceedings of the Fourth International Conference on Automatic Differentiation, July 20– 23, Chicago, Illinois.

Babb:1993:RHP

[BCM^{+}93] R. Babb II, A. Choudhary, L. Meadows, S. Nakamoto, and V. J. Schuster. Retargetable high performance Fortran compiler challenges. In IEEE [IEE93a], pages 137–146. ISBN 0-8186-3400-6. LCCN QA75.5.C58 1993. IEEE catalog no. 93CH3251-6.

Bonham-Carter:1999:BRF

[BCM99] G. F. Bonham-Carter and H. R. Myler. Book reviews: Fundamentals of Engineering Programming with C and FORTRAN. Computers and Geosciences, 25(1):99–??, ???? 1999. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Bala:2000:APC

[BCS00] P. Bala, T. Clark, and L. R.

Scott. Application of Pfortran and Co-Array Fortran in the . . .. Scientific Programming, 9(1):61–68, 2000. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bala:2001:APC

[BCS01] Piotr Bała, Terry Clark, and L. Ridgway Scott. Application of Pfortran and Co-Array Fortran in the parallelization of the GROMOS96 molecular dynamics module. Scientific Programming, 9(1):61–68, ???? 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.com/
app/home/contribution.asp%
3Fwasp=f2779jvvqg63jq64qwtm%
26referrer=parent%26backto=
issue%2C6%2C6%3Bjournal%
2C2%2C9%3Blinkingpublicationresults%
2C1%2C1.

Briggs:1994:IGC

[BCT94] Preston Briggs, Keith D.

Cooper, and Linda Torczon. Improvements to graph coloring register allocation. ACM Transactions on Programming Languages and Systems, 16(3): 428–455, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL http://www.acm.org/

                                        pubs/toc/Abstracts/0164-
                                        0925/177575.html.

Bell:1994:V

[BCWWB94] Gordon Bell, James Cownie, Steve Wallach, and Irving Wladawsky-Berger. Visions. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):5, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Bateson:1990:FPC

[BD90] Allan G. Bateson and Christy L. De Vader. A FORTRAN program to calculate proximity values and an averaged proximity matrix. Educational and Psychological Measurement, ?? (1):131–??, Spring 1990. CO-DEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Berzins:1991:ACP

[BD91] M. Berzins and P. M. Dew. Algorithm 690: Chebyshev polynomial software for ellipticparabolic systems of PDEs. ACM Transactions on Mathematical Software, 17(2):178– 206, June 1991. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/toms/
                                        1991-17-2/p178-berzins/.

Bohling:1993:FPM

[BD93] Geoffrey C. Bohling and John C. Davis. A FOR-TRAN program for Monte Carlo simulation of oil-field discovery sequences. Computers and Geosciences, 19(10):1529– ??, November 1993. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Brandes:1996:IPC

[BD96] T. Brandes and F. Desprez.

Implementing pipelined computation and communication in an HPF compiler. Lecture Notes in Computer Science, 1123:459–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buehler:2014:CCH

[BD14] Stephan Buehler and Claude Duhr. CHAPLIN-Complex Harmonic Polylogarithms in Fortran. Computer Physics Communications, 185(10):2703– 2713, October 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465514001969.

Blackford:1996:FIL

[BDC^{+}96] L. Susan Blackford, Jack J.

Dongarra, Jeremy Du Croz, Sven Hammarling, and Jerzy Wasniewski. A Fortran 90 interface for LAPACK. LAPACK Working Note 117, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, December 1996. URL http:

                                        //www.netlib.org/lapack/
                                        lawns/lawn117.ps;   http:
                                        //www.netlib.org/lapack/
                                        lawnspdf/lawn117.pdf. UT-

CS-96-341, December 1996.

Beguelin:1994:HHN

[BDG^{+}94] Adam Beguelin, Jack J. Dongarra, George Al Geist, Robert Manchek, and Keith Moore. HeNCE: a heterogeneous network computing environment. Scientific Programming, 3(1): 49–60, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL http://www.netlib.
                                        org/utk/people/JackDongarra/
                                        PAPERS/HeNCE-A-Heterogeneous-
                                        Network-Computing-Environment.
                                        pdf.

Boudoul:19xx:IAL

[BDGxx] J. Boudoul, A. Duyck, and J.- P Greiveldinger. Initiation au langage Fortran. Technical report, I.N.S.E.E., Paris, France, 19xx. various pp.

Buchholz:1990:CQDb

[BDH90] Werner Buchholz, Donald Watts Davies, and Mark Halpern. Comments, queries, and debate: Babbage’s friend; errors in the FORTRAN issue. Annals of the History of Computing, 12(2):147, April/June 1990. CODEN AHCOE5. ISSN 0164-1239.

                                        URL http://dlib.computer.
                                        org/an/books/an1990/pdf/
                                        a2147a.pdf;   http://www.

computer.org/annals/an1990/
a2147aabs.htm.

Busa:2005:AFP

[BDH^{+}05] Ján Busa, Jozef Dzurina, Edik Hayryan, Shura Hayryan, Chin-Kun Hu, Ján Plávka, Imrich Pokorný, Jaroslav Skrivánek, and Ming-Chya Wu. ARVO: a Fortran package for computing the solvent accessible surface area and the excluded volume of overlapping spheres via analytic equations. Computer Physics Communications, 165(1):59–96, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Buis:1991:FIB

[BDK91] P. Buis, W. Dyksen, and J. Korb. Fortran interface blocks as an interface description language for remote procedure call. In Einarsson [Ein91], page 116. ISBN ???? LCCN ????

Bleck:1995:CDM

[BDOS95a] R. Bleck, S. Dean, M. O’Keefe, and A. Sawdey. A comparison of data-parallel and messagepassing versions of the Miami Isopycnic Coordinate Ocean Model (MICOM). Parallel Computing, 21(10):1695–1720, October 1995. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bleck:1995:CDP

[BDOS95b] Rainer Bleck, Sumner Dean, Matthew O’Keefe, and Aaron Sawdey. A comparison of data-parallel and messagepassing versions of the Miami Isopycnic Coordinate Ocean Model (MICOM). Parallel Computing, 21(10):1695–1720, November 29, 1995. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1995&
volume=21&issue=10&aid=1022.

Blackford:1998:IGD

[BDPW98] L. S. Blackford, J. J. Dongarra, C. A. Papadopoulos, and R. C. Whaley. Installation guide and design of the HPF 1.1 interface to ScaLA-PACK, SLHPF. LAPACK Working Note 137, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August 1998. URL http:

//www.netlib.org/lapack/
lawns/lawn137.ps;   http:
//www.netlib.org/lapack/
lawnspdf/lawn137.pdf. UT-

CS-98-396, August 1998.

Berntsen:1992:ADA

[BE92] Jarle Berntsen and Terje O. Espelid. Algorithm 706: DCUTRI: An algorithm for adaptive cubature over a collection of triangles. ACM Transactions on Mathematical Software, 18(3):329–342, September 1992. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1992-18-3/p329-berntsen/.

See remark [Esp98].

Becklehimer:1991:FPC

[Bec91] Jeffrey L. Becklehimer. A FORTRAN program for computing beam patterns of geophone arrays. Computers and Geosciences, 17(5):633–??, 1991. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Beebe:1990:PFF

[Bee90] Nelson H. F. Beebe. Parsing Fortran FORMAT statements. Technical report, Center for Scientific Computing and Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 20 1990. 83 pp.

Beebe:1991:SF

[Bee91] Nelson H. F. Beebe. A summary of Fortran. Technical report, Center for Scientific Computing and Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 13 1991. 61 pp. Supplemental class notes prepared for Mathematics 118 and 119.

Beebe:1996:BPAk

[Bee96a]  Nelson H. F. Beebe.  A

bibliography of publications about High-Performance Fortran. Technical report, Center for Scientific Computing and Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 11, 1996. 16 pp. URL https:

//www.math.utah.edu/pub/
tex/bib/index-table-h.html#

hpfortran. This report is updated frequently.

Beebe:1996:BPAc

[Bee96b] Nelson H. F. Beebe. A bibliography of publications about the Fortran programming language: Part 1: 1956–1980. Technical report, Center for Scientific Computing and Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 3, 1996. 247 pp. URL https:

//www.math.utah.edu/pub/
tex/bib/index-table-f.html#
fortran1.

Beebe:1996:BPAd

[Bee96c] Nelson H. F. Beebe. A bibliography of publications about the Fortran programming language: Part 2: 1981–date. Technical report, Center for Scientific Computing and Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 3, 1996. 303 pp. URL https:

//www.math.utah.edu/pub/
tex/bib/index-table-f.html#
fortran2.

Beebe:1997:BPAh

[Bee97] Nelson H. F. Beebe. A bibliography of publications about the Fortran programming language: Part 3: 1990–date.

Technical report, Center for Scientific Computing and Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, October 12, 1997. 190 pp. URL https:

//www.math.utah.edu/pub/
tex/bib/index-table-f.html#

fortran3. This report is updated frequently.

Beebe:2001:LFF

[Bee01a] Nelson H. F. Beebe. Lahey/Fujitsu Fortran 95 compiler: Linux Express v5.5. World-Wide Web document.,

2001. URL https://www.
math.utah.edu/software/lahey.
html.

Beebe:2001:LLB

[Bee01b] Nelson H. F. Beebe. Linking LAPACK and BLAS libraries with Fortran and C code. World-Wide Web document.,

2001. URL https://www.
math.utah.edu/software/lapack.
html.

Beebe:2001:SFC

[Bee01c] Nelson H. F. Beebe. N.A.Software Fortran 95 compiler. World-Wide Web document., 2001.

URL https://www.math.utah.
edu/software/nasoftware.
html.

Beebe:2001:UCC

[Bee01d] Nelson H. F. Beebe. Using C and C++ with Fortran. Technical report, Center for Scientific Computing and Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, November 27, 2001. 23

                                        pp. URL https://www.math.
                                        utah.edu/software/c-with-
                                        fortran.html.

Beebe:2001:UNFc

[Bee01e] Nelson H. F. Beebe. Using the NAG Fortran and C libraries at the University of Utah Mathematics Department. World-Wide Web document., 2001. URL https://

                                        www.math.utah.edu/software/
                                        nag-libraries.html.

Beebe:2001:UNFb

[Bee01f] Nelson H. F. Beebe. Using the NAG Fortran libraries at the University of Utah Mathematics Department. World-Wide Web document., 2001. URL https://

                                        www.math.utah.edu/software/
                                        nag-libraries-fortran.html.

Beebe:2001:UNFa

[Bee01g] Nelson H. F. Beebe. Using the NAG Fortran libraries at the University of Utah Mathematics Department from C and C++. World-Wide Web document.,

                                        2001. URL https://www.
                                        math.utah.edu/software/nag-
                                        libraries-c-to-fortran.html.

Beebe:2002:CBAb

[Bee02] Nelson H. F. Beebe. A complete bibliography of ACM ForTec Forum and ACM Fortran Forum. Technical report, Center for Scientific Computing and Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 7, 2002. 58 pp. URL https:

//www.math.utah.edu/pub/
tex/bib/index-table-f.html#
fortran-forum.

Blume:1994:ADP

[BEH^{+}94] William Blume, Rudolf Eigenmann, Jay Hoeflinger, David Padua, and Paul Petersen. Automatic detection of parallelism: a grand challenge for high-performance computing. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):37–47, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Brainerd:1997:PPF

[BEH97] Walt Brainerd, David Epstein, and Richard Hendrickson. Portability and power with the F programming language. Linux Journal, 42:??, October 1997. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL

ftp://ftp.ssc.com/pub/lj/
listings/issue42/2401.tgz.

Bell:1990:IRS

[Bel90a] Ron Bell. IBM RISC System/6000 performance tuning for numerically intensive Fortran and C programs. Technical Report GG24-3611-00, IBM Corporation, New York, NY, USA, August 1990.

Bell:rs6000-tuning

[Bel90b] Ron Bell. IBM RISC System/6000 performance tuning for numerically intensive Fortran and C programs. Technical Report GG24-3611-00, IBM Corporation, August 1990.

Bell:2011:RFP

[Bel11] William R. Bell. REGCMPNT — a Fortran program for regression models with ARIMA component errors. Journal of Statistical Software, 41(7):??, May 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

                                        jstatsoft.org/v41/i07.

Benkner:1995:VFA

[Ben95] S. Benkner. Vienna Fortran 90 — an advanced data parallel language. Lecture Notes in Computer Science, 964:142– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benkner:1999:OIH

[Ben99a] S. Benkner. Optimizing irregular HPF applications using halos. Lecture Notes in Computer Science, 1586:1015– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benkner:1999:HHP

[Ben99b] Siegfried Benkner. HPF+: High Performance Fortran for advanced scientific and engineering applications. Future Generation Computer Systems, 15(3):381–??, ???? 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Benkner:1999:VVF

[Ben99c]  Siegfried Benkner.  VFC:

The Vienna Fortran Compiler. Scientific Programming, 7(1):67–81, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.com/
app/home/contribution.asp%
3Fwasp=64cr5a4mg33tuhcbdr02%
26referrer=parent%26backto=
issue%2C5%2C7%3Bjournal%
2C8%2C9%3Blinkingpublicationresults%
2C1%2C1.

Benkner:2000:OIH

[Ben00] Siegfried Benkner. Optimizing irregular HPF applications using halos. Concurrency, practice and experience, 12(2-3):137–155, February–March 2000. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72504938/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72504938&PLACEBO=IE.pdf.

Bernecky:1991:FA

[Ber91a] Robert Bernecky. Fortran 90 arrays. ACM SIGPLAN Notices, 26(2):83–98, February 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bernheim:1991:FMD

[Ber91b] Madeleine Bernheim. Fortran mode d’emploi. iia: Informatique Intelligence Artificielle. InterEditions, Paris, France, 1991. ISBN 2-7296-0303-4. 220 pp.

Berry:1992:SWC

[Ber92] Michael W. Berry. Scientific workload characterization by loop-based analyses. ACM SIGMETRICS Performance Evaluation Review, 19(3):17– 29, February 1992. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bethke:1997:SSA

[Bet97] R. C. Bethke. The SoftBench static analysis database. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 48(1):16–18, February 1997. CODEN HPJOAX. ISSN 0018-1153. URL http://www.

                                        hp.com/hpj/97feb/fe97a3.
                                        htm.

Bentley:1992:HCS

[BF92] M. Bentley and C. Froese Fischer. Hypercube conversion of serial codes for atomic structure calculations. Parallel Computing, 18(9):1023– 1031, September 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Blazy:1993:PES

[BF93a] S. Blazy and P. Facon. Partial evaluation and symbolic computation for the understanding of Fortran programs. Lecture Notes in Computer Science, 685:184–198, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blazy:1993:PEU

[BF93b] S. Blazy and P. Facon. Partial evaluation for the understanding of Fortran programs. In Software Engineering and Knowledge Engineering, San Francisco, California, June 1993, pages 517–525. ????, ????, 1993.

Berthou:2001:COH

[BF01] Jean-Yves Berthou and Eric Fayolle. Comparing OpenMP, HPF, and MPI programming: a study case. The International Journal of High Performance Computing Applications, 15(3):297–309, Fall 2001. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Bogucz:1994:PEH

[BFHH94] E. A. Bogucz, G. C. Fox, T. Haupt, and K. A. Hawick. Preliminary evaluation of High-Performance Fortran as a language for computational fluid dynamics. In Anonymous [Ano94l], page ALL. ISBN ???? LCCN ????

Bentley:1993:TIN

[BFKS93a] J. L. Bentley, M. F. Fernandez, B. W. Kernighan, and N. L. Schryer. Template-driven interfaces for numerical subroutines. ACM Transactions on Mathematical Software, 19(3): 265–287, September 1993. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bentley:1993:TDI

[BFKS93b] Jon L. Bentley, Mary F. Fernandez, Brian W. Kernighan, and Norman L. Schryer. Template-driven interfaces for numerical subroutines. ACM Transactions on Mathematical Software, 19(3):265–287, September 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Battaglia:1993:FRC

[BG93] Franco Battaglia and Emilio Gallicchio. FORTRAN routine to compute Born–Oppenheimer potential energy curves directly from spectroscopic data. Journal of Computational Chemistry, 14(5):579–586, May 1993. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Brankin:1994:FVR

[BG94] R. W. Brankin and I. Gladwell. A Fortran 90 version of RKSUITE: an ODE initial value solver.Scientific computation and differential equations (Auckland, 1993). Ann. Numer. Math., 1:363– 375, 1994.

Brandes:1996:RHI

[BG96] T. Brandes and D. Greco. Realization of an HPF interface to ScaLAPACK with redistributions. Lecture Notes in Computer Science, 1067:834– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brankin:1997:ARF

[BG97] R. W. Brankin and I. Gladwell.

Algorithm 771. rksuite 90: Fortran software for ordinary differential equation initial value problems. ACM Transactions on Mathematical Software, 23(3):402–415, September 1997. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1997-23-3/p402-brankin/.

Brainerd:1990:PGF

[BGA90] Walter S. Brainerd, Charles H.

Goldberg, and Jeanne C. Adams. Programmer’s Guide to Fortran 90. McGraw-Hill, New York, NY, USA, 1990. ISBN 0-07-000248-7. vii + 410 pp. LCCN QA76.73.F25 B735 1990. US$37.95. See [Ame92].

Brainerd:1994:PGF

[BGA94] Walter S. Brainerd, Charles H.

Goldberg, and Jeanne C. Adams. Programmer’s guide to Fortran 90. Unicomp, Albuquerque, NM, USA, second edition, 1994. ISBN ???? x + 456 pp. LCCN ????

Brainerd:1996:PGF

[BGA96] Walter S. Brainerd, Charles H. Goldberg, and Jeanne C. Adams. Programmer’s Guide to Fortran 90. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, June 1996. ISBN 0-387-94570-9 (paperback). xi + 445 pp. LCCN QA76.73.F25 B735 1996. US$39.95.

Botsford:1994:PCI

[BGG^{+}94] J. Botsford, A. Gawman, M. Gentleman, E. Kidd, K. Lyons, and J. Slonim, editors. Proceedings. CAS-CON ’94. Integrated Solutions: Toronto, Ontario, Canada, 31 October–3 November 1994. Nat. Res. Council Canada, Ottawa, Ont., Canada, 1994. ISBN ???? LCCN ????

Blackburn:2006:DBJa

[BGH^{+}06] Stephen M. Blackburn, Robin Garner, Chris Hoffmann, Asjad M. Khang, Kathryn S. McKinley, Rotem Bentzur, Amer Diwan, Daniel Feinberg, Daniel Frampton, Samuel Z. Guyer, Martin Hirzel, Antony Hosking, Maria Jump, Han Lee, J. Eliot B. Moss, B. Moss, Aashish Phansalkar, Darko Stefanović, Thomas Van-Drunen, Daniel von Dincklage, and Ben Wiedermann. The DaCapo benchmarks: Java benchmarking development and analysis. ACM SIGPLAN Notices, 41(10):169–190, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bailey:1991:AFS

[BGKZ91] Paul B. Bailey, Burton S. Garbow, Hans G. Kaper, and Anton Zettl. Algorithm 700: A FORTRAN software package for Sturm–Liouville problems. ACM Transactions on Mathematical Software, 17(4):500– 501, December 1991. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0098-
3500/210239.html.

Berrendorf:1994:CSV

[BGLP94] Rudolf Berrendorf, Michael Gerndt, Zakaria Lahjomri, and Thierry Priol. A comparison of shared virtual memory and message passing programming techniques based on a finite element application. Technical Report KFA-ZAM-IB-9410, KFA Research Centre, Jülich, Jülich, Germany, 1994. 12

pp. URL ftp://ftp.zam.kfa-
juelich.de/pub/zamdoc/ib/
ib-94/ib-9410.ps.

Brezany:1992:CFOa

[BGMZ92] Peter Brezany, Michael Gerndt, Piyush Mehrotra, and Hans Zima. Concurrent file operations in a High Performance FORTRAN. In IEEE [IEE92d], pages 230–237. ISBN 0-8186-2630-5. LCCN QA76.5 .S894 1992. IEEE catalog no. 92CH3216-9.

Berrendorf:1993:SF

[BGNP93] Rudolf Berrendorf, Michael Gerndt, Wolfgang E. Nagel, and Joachim Pruemmer. SVM Fortran. Technical Report KFA-ZAM-IB-9322, KFA Research Centre, Jülich, Jülich, Germany, 1993. 23 pp.

URL ftp://ftp.zam.kfa-
juelich.de/pub/zamdoc/ib/
ib-93/ib-9322.ps.

Banerjee:1994:LCP

[BGNP94] Uptal Banerjee, D. Gelernter, A. Nicolau, and D. Padua, editors. Languages and compilers for parallel computing: 6th international workshop, Portland, Oregon, USA, August 12– 14, 1993: proceedings, volume 768 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57659-2 (Berlin), 0-387-57659-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1993. DM122.00.

Brooks:1982:OCL

[BGS82] Rodney A. Brooks, Richard P. Gabriel, and Guy L. Steele, Jr. An optimizing compiler for lexically scoped LISP. ACM SIG-PLAN Notices, 17(6):261–275, June 1982. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bacon:1994:CTH

[BGS94a] David F. Bacon, Susan L. Graham, and Oliver J. Sharp. Compiler transformations for high-performance computing. ACM Computing Surveys, 26 (4):345–420, December 1994. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/
0360-0300/197406.html.

Brezany:1994:SSV

[BGS94b] Peter Brezany, Michael Gerndt, and Viera Sipkova. SVM support in the Vienna Fortran compilation system. Technical Report KFA-ZAM-IB-9401, KFA Research Centre, Jülich, Jülich, Germany, 1994. 10

pp. URL ftp://ftp.zam.kfa-
juelich.de/pub/zamdoc/ib/
ib-94/ib-9401.ps.

Borst:1994:GRP

[BGV94] W. N. Borst, V. V. Goldman, and J. A. Van Hulzen. GENTRAN 90: a REDUCE package for the generation of Fortran 90 code. In ACM [ACM94c], pages 45–51. ISBN 0-89791-638-7. LCCN ????

Bunch:1993:ASM

[BGW93] David S. Bunch, David M.

Gay, and Roy E. Welsch. Algorithm 717: Subroutines for maximum likelihood and quasilikelihood estimation of parameters in nonlinear regression models. ACM Transactions on Mathematical Software, 19(1):109–130, March 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1993-19-1/p109-bunch/.

Baier:1994:PEN

[BGZ94] Romana Baier, Robert Glück, and Robert Zöchling. Partial evaluation of numerical programs in Fortran. In ACM SIGPLAN Workshop on Partial Evaluation and Semantics-Based Program Manipulation, pages 119–132. ????, ????, 1994.

Byrne:1990:AVF

[BH90] Dan J. Byrne and Richard C. Ham. Ada versus FORTRAN: Performance analysis using the ACPS. ACM SIGADA Ada Letters, 10(3):139–145, Winter 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Becker:1992:TFI

[BH92] Gilbert Becker and John Hilton. TETRA: A FORTRAN IV program for computing the cosine-pi approximation to the tetrachoric correlation coefficient. Educational and Psychological Measurement, 52(1): 81–86, Spring 1992. CODEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Baker:2009:ASN

[BHLT09] C. G. Baker, U. L. Hetmaniuk, R. B. Lehoucq, and H. K. Thornquist. Anasazi software for the numerical solution of large-scale eigenvalue problems. ACM Transactions on Mathematical Software, 36 (3):13:1–13:23, July 2009. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bromley:1991:FAT

[BHMS91a] Mark Bromley, Steven Heller, Tim McNerney, and Guy L. Steele, Jr. Fortran at ten gigaflops: The Connection Machine convolution compiler. ACM SIGPLAN Notices, 26 (6):145–156, June 1991. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Published as SIGPLAN Notices, volume 26, number 6.

Bromley:1991:FTG

[BHMS91b] Mark Bromley, Steven Heller, Tim McNerney, and Guy L. Steele, Jr. Fortran at ten gigaflops: the Connection Machine convolution compiler. ACM SIGPLAN Notices, 26 (6):145–156, June 1991. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brooks:1992:NAD

[BHS92] Gary Brooks, Gilbert J.

Hansen, and Steve Simmons.

A new approach to debugging optimized code. ACM SIGPLAN Notices, 27(7):1–11, July 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/143095/p1-brooks/.

Brent:1980:AIB

[BHY80] Richard P. Brent, Judith A. Hooper, and J. Michael Yohe. An AUGMENT interface for Brent’s multiple precision arithmetic package. ACM Transactions on Mathematical Software, 6(2):146–149, June 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Bre78, Bre79, Smi98].

Branca:1995:CBH

[BID95] A. Branca, M. Ianigro, and A. Distante. A comparison between HPF and PVM for data parallel algorithms on a cluster of workstations using a high speed network. In Hertzberger and Serazzi [HS95], pages 930–

931. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Bini:1996:NCP

[Bin96] Dario Andrea Bini. Numerical computation of polynomial zeros by means of Aberth’s method. Numerical Algorithms, 13(3-4):179–200, ????

1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bjorner:2008:JWB

[Bjø08] Dines Bjørner. John Warner Backus: 3 Dec 1924–17 March 2007. Formal Aspects of Computing, 20(3): 239–240, May 2008. CO-DEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s00165-008-
0077-4;   http://link.springer.
com/article/10.1007/s00165-
008-0077-4.

Balasundaram:1989:TSD

[BK89] Vasanth Balasundaram and Ken Kennedy. A technique for summarizing data access and its use in parallelism enhancing transformations. ACM SIGPLAN Notices, 24(7):41–53, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/73141/p41-balasundaram/

.

Baden:1995:PPP

[BK95] Scott B. Baden and Scott R.

Kohn. Portable parallel programming of numerical problems under the LPAR system. Journal of Parallel and Distributed Computing, 27 (1):38–55, May 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jpdc.1995.1070/production;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1995.1070/production/pdf.

Benner:2006:AFS

[BK06] Peter Benner and Daniel Kressner. Algorithm 854: Fortran 77 subroutines for computing the eigenvalues of Hamiltonian matrices II. ACM Transactions on Mathematical Software, 32 (2):352–373, June 2006. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bixby:1994:ADL

[BKK94] R. Bixby, K. Kennedy, and U. Kremer. Automatic data layout using 0-1 integer programming. In Cosnard et al. [CGS94], pages 111–122. CO-DEN ITATEC. ISBN 0-444-81926-6. ISSN 0926-5473. LCCN QA76.58 .I46 1994.

Bischof:1996:AAD

[BKMC96] Christian Bischof, Peyvand Khademi, Andrew Mauer, and Alan Carle. Adifor 2.0: Automatic differentiation of Fortran 77 programs. IEEE Computational Science & Engineering, 3(3):18–32, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c3018abs.
                                        htm.

Bodin:1993:FFI

[BKP93] F. Bodin, L. Kervella, and T. Priol. Fortran-S: A Fortran interface for shared virtual memory architectures. In IEEE [IEE93d], pages 274–283. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Broughan:1991:SSC

[BKR^{+}91] K. A. Broughan, G. Keady, T. D. Robb, M. G. Richardson, and M. C. Dewar. Some symbolic computing links to the NAG numeric library. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(3): 28–37, July 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Banger:2022:FFP

[BKRG22] Paramjeet Banger, Pardeep Kaur, Arko Roy, and Sandeep Gautam. FORTRESS: FOR-TRAN programs to solve coupled Gross–Pitaevskii equations for spin-orbit coupled spin-f Bose–Einstein condensate with spin f = 1 or

2. Computer Physics Communications, 279(??):Article 108442, October 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465522001618.

Bampis:1991:ICC

[BKT91] E. Bampis, J. C. Konig, and D. Trystram. Impact of communications on the complexity of the parallel Gaussian elimination. Parallel Computing, 17(1):55–61, April 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bilodeau:1990:DSM

[BL90] M. Bilodeau and D. Laguitton. Development of screen management subroutines compatible with Microsoft FOR-TRAN compiler. The Computer Journal, 33(5):444–452, October 1990. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Boltjes:1991:MDC

[BL91] B. Boltjes and S. W. De Leeuw. Molecular dynamics on the Connection Machine using Fortran. Molecular simulation, 7(??):1–??, 1991. CODEN MOSIEA. ISSN 0892-7022.

Ball:1993:BPFa

[BL93a] Thomas Ball and James R. Larus. Branch prediction for free. Technical Report TR 1137, Computer Science Department, University of Wisconsin — Madison, Madison, WI, USA, February 1993.

Ball:1993:BPFb

[BL93b] Thomas Ball and James R. Larus. Branch prediction for free. ACM SIGPLAN Notices, 28(6):300–313, June 1993. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/155090/p300-ball/.

Butler:1994:MMC

[BL94] Ralph M. Butler and Ewing L. Lusk. Monitors, messages, and clusters: The p4 parallel programming system. Parallel Computing, 20(4):547– 564, March 31, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1994&
volume=20&issue=4&aid=864.

Blazy:2000:SAG

[Bla00] S. Blazy. Specifying and automatically generating a specialization tool for Fortran 90. Automated Software Engineering, 7(4):345–376, December 2000. CODEN ???? ISSN 0928-8910.

Bliss:1990:IFP

[Bli90] Brian Eugene Bliss. Instrumentation of Fortran programs for automatic roundoff error analysis and performance evaluation. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana, IL, USA, 1990. xiii + 132 pp.

Blinn:1994:JBC

[Bli94] Jim Blinn. Jim Blinn’s corner: Farewell to Fortran. IEEE Computer Graphics and Applications, 14(2):86–89, March 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Brown:1996:ALL

[BLL^{+}96] Barry W. Brown, Lawrence B. Levy, James Lovato, Kathy Russell, and Floyd M. Spears. Algorithm 762: LLDRLF, loglikelihood and some derivatives for log-F models. ACM Transactions on Mathematical Software, 22(3):372–382, September 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Benson:1995:DDP

[BLLWW95] Edward G. Benson, David C. P. LaFrance-Linden, Richard A. Warren, and Santa Wiryaman. Design of Digital’s Parallel Software Environment. Digital Technical Journal, 7(3): 24–38, Fall 1995. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n3/Design_of_Digitals_
Parallel_S_02jan1996DTJJ02P8.
ps;   http://www.digital.
com:80/info/DTJJ02/DTJJ02AH.
HTM;   http://www.digital.
com:80/info/DTJJ02/DTJJ02P8.
PS;   http://www.digital.
com:80/info/DTJJ02/DTJJ02PF.
PDF;   http://www.digital.

com:80/info/DTJJ02/DTJJ02SC.
TXT.

Becker:1994:TPI

[BLT94] M. Becker, L. Litzler, and M. Tehel, editors. Transputers ’94: [advanced research and industrial applications]: proceedings of the international conference, September 21–23, 1994, Saline Royale d’Arc et Senans, France. IOS Press, Amsterdam, The Netherlands, 1994. ISBN 90-5199-179-7 (IOS Press), 4-274-90003-7 (Ohmsha). LCCN ????

Blunt:1991:CFD

[Blu91] David Mark Blunt. The conversion of a Fortran data plotting program using DI-3000 graphics to operation on a Macintosh personal computer. Flight mechanics technical memorandum; 446, Dept. of Defence, Defence Science and Technology Organisation Aeronautical Research Laboratory, Melbourne, Victoria, Australia, 1991. ISBN 0-646-10146-3 (not printed on item). 34 pp.

Brezany:2002:PSH

[BLW02] Peter Brezany, Jonghyun Lee, and Marianne Winslett. Parallel I/O support for HPF on computational grids. Lecture Notes in Computer Science, 2327:539–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/

                                        0558/bibs/2327/23270539.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2327/23270539.
                                        pdf.

Berry:1999:AHD

[BM99] Michael W. Berry and Karen S. Minser. Algorithm 798: Highdimensional interpolation using the modified Shepard method. ACM Transactions on Mathematical Software, 25(3): 353–366, September 1999. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bozkus:1994:TCE

[BMMN94] Z. Bozkus, L. Meadows, D. Miles, and S. Nakamoto. Techniques for compiling and executing HPF programs on shared memory and distributed memory parallel systems. In Kumar [Kum94], pages 515– 520. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Bozkus:1995:CHP

[BMN^{+}95] Z. Bozkus, L. Meadows, S. Nakamoto, V. Schuster, and M. Young. Compiling High Performance Fortran. In Bailey et al. [BBG^{+}95], pages 704–709. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Bozkus:1997:POH

[BMN^{+}97] Zeki Bozkus, Larry Meadows, Steven Nakamoto, Vincent Schuster, and Mark Young. PGHPF — an optimizing High Performance Fortran compiler for distributed memory machines. Scientific Programming, 6(1):29–40, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ballance:1990:PDW

[BMO90] Robert A. Ballance, Arthur B.

Maccabe, and Karl J. Ottenstein. The program dependence web: a representation supporting control-, data- , and demand-driven interpretation of imperative languages. ACM SIGPLAN Notices, 25(6):257–271, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/93542/p257-ottenstein/

.

Birgin:2001:ASS

[BMR01] Ernesto G. Birgin, José Mario Martínez, and Marcos Raydan. Algorithm 813: SPG—software for Convex-Constrained Optimization. ACM Transactions on Mathematical Software, 27 (3):340–349, September 2001. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Briguglio:2003:PPM

[BMV03] Sergio Briguglio, Beniamino Di Martino, and Gregorio Vlad. A performance-prediction model for PIC applications on clusters of symmetric multiprocessors: Validation with hierarchical HPF + OpenMP implementation. Scientific Programming, 11(2):159–176, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Barton:1993:SEP

[BN93] J. J. Barton and L. R. Nackman. Scientific and engineering programming in C++: grafting onto our roots. C++ Report, 5(7):26–31, September 1993. CODEN CRPTE7. ISSN 1040-6042.

Barton:1996:SEC

[BN96] J. J. Barton and L. R. Nackman. Scientific and engineering C++-is this standard input? C++ Report, 8(9):66, 68–69, October 1996. CODEN CRPTE7. ISSN 1040-6042.

Barton:1997:GT

[BN97] J. J. Barton and L. R. Nackman. Generating and transforming. C++ Report, 9(2):62– 64, February 1997. CODEN CRPTE7. ISSN 1040-6042.

Bodin:1994:DPP

[Bod94] François Bodin. Directions in parallel programming HPF, shared virtual memory and object parallelism in pC++. NASA contractor report NASA CR-194943; ICASE report no. 94-54, Institute for Computer Applications in Science and Engineering, NASA Langley Research Center, Hampton, VA, USA, 1994.

Boisvert:1997:QNS

[Boi97] Ronald F. Boisvert, editor.

Quality of numerical software: assessment and enhancement / proceedings of the IFIP TC2/WG2.5 Working Conference on the Quality of Numerical Software, Assessment and Enhancement, Oxford, United Kingdom, 8–12 July 1996. Chapman and Hall, Ltd., London, UK, 1997. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Bonten:2006:ACF

[Bon06] J. H. M. Bonten. Arithmetic computer formats. Web document, October 5, 2006. URL

http://home.hetnet.nl/mr_
1/81/jhm.bonten/computers/
bitsandbytes/wordsizes/.

Booch:1981:DSD

[Boo81] Grady Booch. Describing software design in Ada. ACM SIGPLAN Notices, 16(9):42– 47, September 1981. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bradley:2005:OUP

[BOPC05] J. Bradley, S. F. C. O’Rourke, R. T. Pedlow, and D. S. F. Crothers. Other uses of the program ‘ARGON.f90’. Computer Physics Communications, 172(1):42–44, October 15, 2005. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465505003978.

Borse:1991:FNMb

[Bor91a] Garold J. Borse. FORTRAN 77 and numerical methods for engineers. PWS-KENT series in engineering and computer science. PWS-Kent Pub. Co., Boston, MA, USA, second edition, 1991. ISBN 0-534-98353-7. xiv + 722 pp.

Borse:1991:FNMa

[Bor91b] Garold J. (Garold J.) Borse. Fortran 77 and Numerical Methods for Engineers. PWS-Kent series in engineering and computer science. PWS-Kent Pub. Co., Boston, MA, USA, second edition, May 1991. ISBN 0-534-92562-6. xiv + 722 pp. LCCN TA 345 B67 1991. US$57.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0-534-92562-6.

Bose:2019:NMM

[Bos19] Sujit Kumar Bose. Numerical methods of mathematics implemented in Fortran. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2019. ISBN 981-13-7114-8. xvii + 467 pp. LCCN ????

Boulter:1995:PEH

[Bou95] B. Boulter. Performance evaluation of HPF for scientific computing. In Hertzberger and Serazzi [HS95], pages 652– 657. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Boulet:1996:BFL

[Bou96] P. Boulet. Bouclettes: a Fortran loop parallelizer. Lecture Notes in Computer Science, 1067:784–791, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouaricha:1997:ASS

[Bou97]  Ali Bouaricha.  Algorithm

765: STENMIN — a software package for large, sparse unconstrained optimization using tensor methods. ACM Transactions on Mathematical Software, 23(1):81–90, March 1997. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1997-23-1/p81-bouaricha/.

Beavis:1992:PDD

[BP92] Andy Beavis and Chris Phillips.

Porting a dusty deck Fortran program to a shared-memory multiprocessor. Concurrency, practice and experience, 4(8): 575–587, December 1992. CO-DEN CPEXEI. ISSN 1040-3108.

Becks:1994:NCT

[BPG94] K.-H. Becks and D. Perret-Gallix, editors. New computing techniques in physics research III: proceedings of the Third International Workshop on Software Engineering, Artificial Intelligence and Expert Systems for High Energy and Nuclear Physics: October 4– 8, 1993, Oberammergau, Germany. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1994. ISBN 981-02-1699-

8. LCCN QC793.47.E4I58 1993.

Bae:1996:PUC

[BR96] Seungjo Bae and Sanjay Ranka. PACK/UNPACK on coarse-grained distributed memory parallel machines. Journal of Parallel and Distributed Computing, 38(2):204– 216, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1996.0141/production;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0141/production/pdf.

Boulet:1998:CPH

[BR98] P. Boulet and X. Redon. Communication pre-evaluation in HPF. Lecture Notes in Computer Science, 1470:263–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradberry:1990:PFP

[Bra90] John L. Bradberry. Porting FORTRAN programs from minis to PCs. Dr. Dobb’s Journal of Software Tools, 15(9):26– 36, September 1990. CODEN DDJOEB. ISSN 1044-789X.

Bradberry:1991:FG

[Bra91] John L. Bradberry. Fortran and GUIs. Dr. Dobb’s Journal of Software Tools, 16(6):16–18, 20, 22, 25, 101–102, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Bradley:1994:FAD

[Bra94a] Allen Bradley. Flood analysis in DuPage County using hydrological simulation program, FORTRAN model. Transportation research record 1471, DuPage County, IL, USA, 1994. 41–46 pp.

Brainerd:1994:F

[Bra94b] W. S. Brainerd. Fortran.

Computer Standards and Interfaces, 16(5):459–??, 1994. CO-DEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Brandes:1994:EHF

[Bra94c] T. Brandes. Evaluation of High-Performance Fortran on some real applications. In Gentzsch and Harms [GH94b], pages 417–422. ISBN 3-540-57981-8 (Berlin: vol. 2: paperback), 0-387-57981-8 (New York: vol. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1– 2 (c1994). DM96.00. Two volumes.

Brandes:1994:EHP

[Bra94d] T. Brandes. Evaluation of High Performance Fortran on some real applications. In Gentzsch and Harms [GH94a], pages 417–422. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Brainerd:1996:E

[Bra96] Walt Brainerd. Editorial. Fortran Journal, 8(6):??, November/December 1996. ISSN 1060-0221. URL http:/

                                        /www.fortran.com/fortran/
                                        FJ/9611/#editorial.

Brainerd:1997:AED

[Bra97a] Walt Brainerd. Across the Editor’s desk. Fortran Journal, 9 (1):??, 1997. ISSN 1060-0221.

                                        URL http://www.fortran.
                                        com/fortran/FJ/9701/desk.
                                        html.

Brainerd:1997:TFP

[Bra97b] Walt Brainerd. Timings of F programs and Fortran 90 programs. Fortran Journal, 9(1):??, 1997. ISSN 1060-0221. URL http:/

                                        /www.fortran.com/fortran/
                                        FJ/9701/f_times.html.

Bramley:1997:TNRb

[Bra97c] Randall Bramley. Technology news & reviews: Chemkin software; OpenMP Fortran Standard; ODE toolbox for Matlab; Java products; Scientific WorkPlace 3.0. IEEE Computational Science & Engineering, 4(4):75–78, October/December 1997. CO-DEN ISCEE4. ISSN 1063-6552 (print), 1558-1861 (electronic).

URL http://dlib.computer.
org/cs/books/cs1997/pdf/
c4075.pdf.

Bramley:1997:TNF

[Bra97d] Randall Bramley. Technology news: Fortran 90 news; free software; symbolic computing packages; Matlab 5; Web products. IEEE Computational Science & Engineering, 4(1):87, 89–90, January/March 1997. CODEN ISCEE4. ISSN 1063-6552 (print), 1558-1861 (electronic).

URL http://dlib.computer.
org/cs/books/cs1997/pdf/
c1087.pdf;   http://www.
cs.berkeley.edu/~xiaoye;
http://www.cs.cornell.edu/
home/vavasis/qmg-home.html;
http://www.globus.org/sage;
http://www.macsyma.com/;
http://www.mathworks.com/
; http://www.netlib.org/
benchmark/linpackjava;
http://www.netlib.org/scalapack;
http://www.ucmp.berkeley.
edu/subway/phylogen.html;
http://www.vni.com/products/
wpd/jnl/jnl_1_0.html;   http:
//www.wolfram.com/look/cse.

Brandes:2000:HLC

[Bra00] T. Brandes. HPF library and compiler support for halos in data parallel irregular computations. Parallel Processing Letters, 10 (2/3):189–??, September 2000. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL http:

//ejournals.wspc.com.sg/
ppl/10/1002_03/S0129626400000196.
html.

Brainerd:2003:IFC

[Bra03] Walt Brainerd. The importance of Fortran in the 21st Century. Journal of Modern Applied Statistical Methods, 2 (1):14–15, May 2003. CO-DEN ???? ISSN 1538-9472.

URL http://tbf.coe.wayne.
edu/jmasm/.

Blanco-Rey:2004:FLE

[BRdAHK04] Maria Blanco-Rey, Pedro de Andres, Georg Held, and David A. King. A FORTRAN-90 low-energy electron diffraction program (LEED90 v1.1). Computer Physics Communications, 161(3):151–165, August 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465504002310.

Brent:1978:AMF

[Bre78] Richard P. Brent. Algorithm 524: MP, A Fortran multipleprecision arithmetic package [A1]. ACM Transactions on Mathematical Software, 4(1): 71–81, March 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [Bre79, BHY80, Smi98].

Brent:1979:RMF

[Bre79] R. P. Brent. Remark on “Algorithm 524: MP, A Fortran multiple-precision arithmetic package [A1]”. ACM Transactions on Mathematical Software, 5(4):518–519, December 1979. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Bre78, BHY80, Smi98].

Brezany:1992:CFOb

[Bre92] Peter Brezany. Concurrent file operations in a high performance FORTRAN. NASA contractor report CR-189711; ICASE report no. 92-46, National Aeronautics and Space Administration, Langley Research Center; National Technical Information Service, distributor, Hampton, VA, USA, 1992. i + 15 pp.

Bomans:1990:AGM

[BRH90] L. Bomans, D. Roose, and R. Hempel. The argonne/ GMD macros in FORTRAN for portable parallel programming and their implementation on the Intel iPSC/2. Parallel Computing, 15(1–3):119– 132, September 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Brieger:2000:HOO

[Bri00]  Leesa Brieger.  HPF to

OpenMP on the Origin2000: a case study. Concurrency, practice and experience, 12 (12):1147–1154, October 2000. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76500351/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500351&PLACEBO=IE.pdf.

Bronson:1990:MFS

[Bro90a]  Gary J. Bronson.  Modu-

lar Fortran 77 for Science and Engineering. Scott/ Jones, El Granada, CA, USA, February 1990. ISBN 0-9624230-0-9. various pp. LCCN QA76.73.F25 B75 1990. US$28.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0-9624230-0-9.  In-

cludes disk.

Brown:1990:CFP

[Bro90b] T. D. Brown. C for Fortran Programmers. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, April 1990. ISBN 0-929306-01-5, 0-13-120528-5 (paperback). xii + 206 pp. LCCN QA76.73.C15B774 1990. US$22.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0-929306-01-5.

Bronson:1992:MFA

[Bro92a] Gary Bronson. Modern Fortran 77/90: The Alternate Edition. Scott/Jones, El Granada, CA, USA, January 1992. ISBN 0-9624230-5-X. US$35.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=096242305X. Includes

disk.

Bronson:1992:MF

[Bro92b] Gary J. Bronson. Modern Fortran 77/90. Scott/Jones, El Granada, CA, USA, alternate edition, 1992. ISBN 0-9624230-5-X. various pp.

Bronson:1995:FSE

[Bro95]  Gary J. Bronson.  For-

tran for Scientists and Engineers. Scott/Jones, El Granada, CA, USA, second edition, April 1995. ISBN 1-881991-39-3. viii + 546 pp. US$35.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=1881991393.

Brooks:1997:PSF

[Bro97] David Brooks. Problem Solving With Fortran 90: For Scientists and Engineers. Undergraduate Texts in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-98229-9. xxxiv + 682 pp. US$49.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0387982299.

Brook:2003:FSG

[Bro03] Paul Brook. Fortran 95 support in GCC. In Hutton et al. [HDR03], pages 35–42. ISBN

????  LCCN ????  URL

                                        http://www.linux.org.uk/
                                        ~ajh/gcc/gccsummit-2003-
                                        proceedings.pdf.

Bruccoleri:1996:TCW

[Bru96a] Robert E. Bruccoleri. Technical correspondence: WRAP-GEN — a tool for the use of Fortran and C together in portable programs. ACM SIGPLAN Notices, 31(4):20– 27, April 1996. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruccoleri:1996:WTU

[Bru96b] Robert E. Bruccoleri. WRAP-GEN: a tool for the use of Fortran and C together in portable programs. ACM SIG-PLAN Notices, 31(4):20–27, April 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barnard:1991:EFE

[BS91a] David T. Barnard and David B. Skillicorn. Effective Fortran 77 for Engineers and Scientists. Wm. C. Brown Publishers, Dubuque, IA, USA, second edition, June 1991. ISBN 0-697-08546-5. xxii + 504 pp. LCCN QA76.73.F25B366 1992. US$63.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0-697-08546-5.

Barnard:1991:IMA

[BS91b] David T. Barnard and David B.

Skillicorn. Instructor’s manual to accompany Effective Fortran77 for engineers and scientists. Wm. C. Brown Publishers, Dubuque, IA, USA, second edition, 1991. ISBN 0-697-08547-3. v + 132 pp.

Bai:1992:SAF

[BS92a] Z. Bai and G. W. Stewart.

SRRIT — A FORTRAN subroutine to calculate the dominant invariant subspace of a nonsymmetric matrix. Technical Report 2908, Department of Computer Science, University of Maryland, College Park, MD, USA, 1992.

Bai:1992:SFS

[BS92b] Zhaojun Bai and G. W. Stewart. SRRIT — A FORTRAN subroutine to calculate the dominant invariant subspace of a nonsymmetric matrix. Computer science technical report series CS-TR-2908, UMIACS-TR-92-61, University of Maryland, College Park, MD, USA, May 1992. 25 pp.

Bai:1997:ASF

[BS97] Z. Bai and G. W. Stewart. Algorithm 776. SRRIT — a FOR-TRAN subroutine to calculate the dominant invariant subspace of a nonsymmetric matrix. ACM Transactions on Mathematical Software, 23(4): 494–513, December 1997. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Barash:2013:RPL

[BS13] L. Yu. Barash and L. N. Shchur. RNGSSELIB: Program library for random number generation. More generators, parallel streams of random numbers and Fortran compatibility. Computer Physics Communications, 184(10):2367–2369, October 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465513001422.

Bull:2003:BJA

[BSB^{+}03] J. M. Bull, L. A. Smith, C. Ball, L. Pottage, and R. Freeman. Benchmarking Java against C and Fortran for scientific applications. Concurrency and Computation: Practice and Experience, 15(3–5): 417–430, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brezany:1995:PIC

[BSCV95] P. Brezany, K. Sanjari, O. Cheron, and E. Van Konijnenburg. Processing irregular codes containing arrays with multidimensional distributions by the PREPARE HPF compiler.

In Hertzberger and Serazzi [HS95], pages 526–531. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Bull:2001:BJA

[BSPF01] Mark Bull, Lorna Smith, Lindsay Pottage, and Robin Freeman. Benchmarking Java against C and Fortran for scientific applications. In ACM [ACM01], pages 97–

105. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

2001.  URL http://www.
philippsen.com/JGI2001/camerareadyabstracts/
30.html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500097.pdf.

Belonoshki:1992:SFP

[BSS92] A. B. Belonoshki, Pingfang Shi, and S. K. Saxena. SUPER-FLUID: a FORTRAN 77 program for calculation of Gibbs free energy and volume of C-H-O-N-S-ar mixtures. Computers and Geosciences, 18(9): 1267–??, October 1992. CO-DEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Benner:2016:AFS

[BSV16] Peter Benner, Vasile Sima, and Matthias Voigt. Algorithm 961: Fortran 77 subroutines for the solution of skew-Hamiltonian/ Hamiltonian eigenproblems. ACM Transactions on Mathematical Software, 42(3):24:1– 24:26, May 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Barrett:1994:PF

[BT94] K. E. Barrett and R. J. Tiddy. Pseudo-recursive Fortran. Advances in engineering software, 21(1):17–??, 1994. CODEN AESODT. ISSN 0965-9978 (print), 0141-1195 (electronic).

Boisvert:2001:ASS

[BT01] Ronald F. Boisvert and Ping Tak Peter Tang, editors. The architecture of scientific software: IFIP TC2/WG2.5 Working Conference on the Architecture of Scientific Software, October 2–4, 2000, Ottawa, Canada, volume 60 of IFIP. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-7339-1. LCCN QA76.758 .I345 2000.

Buckley:1994:AFC

[Buc94a] A. G. Buckley. Algorithm 734: a Fortran 90 code for unconstrained nonlinear minimization. ACM Transactions on Mathematical Software, 20(3): 354–372, September 1994. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Buckley:1994:AXF

[Buc94b] A. G. Buckley. Algorithm XXX: A Fortran 90 code for unconstrained nonlinear minimization. ACM Transactions on Mathematical Software, 1994. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). To appear, 1994.

Buckley:1994:CFC

[Buc94c]  A. G. Buckley.  Conver-

sion to Fortran 90: a case study. ACM Transactions on Mathematical Software, 20(3):308–353, September 1994. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1994-20-3/p308-buckley/.

Butt:1995:IFS

[But95] Farooq Butt. Implementing FORTRAN77 support in the GNU gdb debugger. ACM SIGPLAN Notices, 30(5):29– 36, May 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buchberger:1994:PPC

[BV94] Bruno Buchberger and Jens Volkert, editors. Parallel processing: CONPAR 94-VAPP VI: third Joint International Conference on Vector and Parallel Processing, Linz, Austria, September 6–8, 1994: proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58430-7 (Berlin), 0-387-58430-7 (New York). LCCN QA76.58 .J65 1994.

Broeckhove:2013:ACC

[BV13] J. Broeckhove and K. Vanmechelen. An adaptor for C++ callbacks with C and Fortran libraries. Computer Physics Communications, 184(3):824–832, March 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465512003591.

Broughan:1996:FLT

[BW96] K. A. Broughan and D. M. K. Willcock. Fortran to Lisp translation using f2cl. Software—Practice and Experience, 26(10):1127–1139, October 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
16765.

Berg:2012:FCL

[BW12] Bernd A. Berg and Hao Wu. Fortran code for SU(3) lattice gauge theory with and without MPI checkerboard parallelization. Computer Physics Communications, 183(10):2145–2157, October 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465512001269.

Brezany:2001:GIP

[BxCW01] P. Brezany, P. x. Czerwinski, and M. Winslett. A generic interface for parallel access to large data sets from HPF applications. Future Generation Computer Systems, 17(8):977– 985, June 2001. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Brandes:1994:ATT

[BZ94] T. Brandes and F. Zimmermann. Adaptor — a transformation tool for HPF programs. In Decker and Rehmann [DR94a], pages 91– 96. ISBN 0-8176-5090-3. LCCN QA76.58 .P767 1994.

Benkner:1999:CHP

[BZ99] S. Benkner and H. Zima. Compiling High Performance Fortran for distributed-memory architectures. Parallel Computing, 25(13):1785–1825, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Comeau:1990:AFP

[CA90] Raymond A. Comeau and Samon Ando. AMAD: a FOR-TRAN program to calculate added mass and damping of two-dimensional floating bodies by the boundary integral equation method. Technical communication 90/302, Defence Research Establishment Atlantic, Dartmouth, NS, Canada, 1990. v + 20 pp.

Cooper:1992:FVI

[CA92] R. K. Cooper and R. J. Allan. Fortnet (3L) v1.0: Implementation and extensions of a message-passing harness for transputers using 3L Parallel Fortran. Computer Physics Communications, 70(3):521– 543, July 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Coelho:1996:OCH

[CA96] Fabien Coelho and Corinne Ancourt. Optimal compilation of HPF remappings. Journal of Parallel and Distributed Computing, 38(2):229– 236, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jpdc.1996.0143/production;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0143/production/pdf.

Cahill:1990:HUM

[Cah90] Kevin Cahill. How to use MACSYMA to write long FORTRAN codes for noncompact simulations of gauge theories. Computers in physics, 4(2):159–165, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Campbell:2013:WES

[Cam13] Scott Campbell. ‘Wat For Ever’ student-oriented computing at the University of Waterloo. IEEE Annals of the History of Computing, 35(1):11– 22, January/March 2013. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Cann:1991:RFD

[Can91] D. Cann. Retire FORTRAN?

A debate rekindled. In IEEE [IEE91], pages 264–272. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Cann:1992:RFA

[Can92a] D. C. Cann. Retire Fortran? A Debate Rekindled. Comm. ACM, 35(8):81–89, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cann:1992:RFD

[Can92b]  David Cann.  Retire For-

tran? A debate rekindled. Comm. ACM, 35(8):81– 89, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-
0782/135231.html.

Cap:1998:NNM

[Cap98]  F. F. Cap.  A new nu-

merical method to calculate the vibrations of clamped Kirchhoff-plates of arbitrary form. Computing, 61(2): 181–188, 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

                                        URL http://www.springer.
                                        at/journals/pdf/3570.pdf;
                                         http://www.springer.at/
                                        journals/pdf/3816.pdf.

Carr:1990:FUM

[Car90] Alan Carr. A FORTRAN update of the MININEC electromagnetics modeling program. R.F. design, 13(2):59– ??, February 1990. CODEN RFDEDG. ISSN 0163-321X.

Carlson:1991:TEI

[Car91a] B. C. Carlson. A table of elliptic integrals: One quadratic factor. Mathematics of Computation, 56(193):267–280, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1991:UHP

[Car91b]  David A. Carlson.  Ultra

high performance FFTs for the CRAY-2 and CRAY-YMP supercomputers. Technical report SRC-TR-91-053, Supercomputing Research Center: IDA, Lanham, MD, USA, December 1991. 14 pp.

Carlson:1992:UPF

[Car92] David A. Carlson. Ultrahighperformance FFTS for the Cray-2 and Cray Y-MP supercomputers. The Journal of supercomputing, 6(2):

107–116, June 1992. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=6&issue=2&spage=
107.

DeCaritat:1993:HFC

[Car93] Patrice De Caritat. HEAT: a FORTRAN computer program for calculating 1-D conductive and advective heat transport in geological formations. Computers and Geosciences, 19(5): 673–??, May 1993. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Cash:1989:ABF

[Cas89a]  J. R. Cash.  Algorithm

669: BRKF45: A FORTRAN subroutine for solving firstorder systems of nonstiff initial value problems for ordinary differential equations. ACM Transactions on Mathematical Software, 15(1):29– 30, March 1989. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0098-
3500/214388.html. See also

[Hig91].

Casimir:1989:FGP

[Cas89b] Rommert J. Casimir. Fourth generation problems. ACM SIGPLAN Notices, 24(5):83– 86, May 1989. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cass:2014:TPL

[Cas14] Stephen Cass. The top 10 programming languages: Spectrum’s 2014 ranking [dataflow]. IEEE Spectrum, 51(7):68, July 2014. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Calloni:1994:IPB

[CB94] Ben A. Calloni and Donald J. Bagert. ICONIC programming in BACCII vs. textual programming: which is a better learning environment? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26 (1):188–192, March 1994. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DeCaritat:1995:TFP

[CB95] P. De Caritat and J. Braun. TWYTT: a FORTRAN program for calculating the twoway travel time of stratigraphic or other markers on seismic reflection profiles of layered geological media. Computers and Geosciences, 21(2):345–??, 1995. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Compagner:1997:RER

[CBTL97] A. Compagner, A. S. Berdnikov, S. B. Turtia, and A. Larionov. Rounding errors in random number generators. Computer Physics Communications, 106(3):207–218, November 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465597000702.

Chivers:1990:IF

[CC90] I. D. (Ian David) Chivers and M. W. (Malcolm W.) Clark. Interactive Fortran 77, volume 28 of Ellis Horwood series in computers and their applications. Ellis Horwood, New York, NY, USA, second edition, August 1990. ISBN 0-470-20101-0. 231 pp. US$37.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0470201010.

Cash:1992:AMF

[CC92a] J. R. Cash and S. Considine. Algorithm 703: MEBDF: A FORTRAN subroutine for solving first-order systems of stiff initial value problems for ordinary differential equations. ACM Transactions on Mathematical Software, 18(2):156– 158, June 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
toms/1992-18-2/p156-cash/.

Chen:1992:PFC

[CC92b] Marina Chen and James Cowie. Prototyping Fortran-90 compilers for massively parallel machines. ACM SIG-PLAN Notices, 27(7):94–105, July 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/proceedings/
                                        pldi/143095/p94-chen/.

Corey:1993:ASO

[CC93] P. D. Corey and J. R. Clymer. Approaches to speeding up old Fortran programs: Applying multiple queues. International Journal of Modelling and Simulation, 13(3):123–133, 1993. CODEN IMSIEK. ISSN 0228-6203 (print), 1925-7082 (electronic).

Chen:1994:CEC

[CC94] Dong-Yuan Chen and M. C. Chen. CPAR-language extensions to C for irregular and adaptive parallel computations. In IEEE [IEE94d], pages 501–508. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994.

Chang:1995:PMH

[CC95a] Yu-Chung Chang and Tony F. Chan. Performance modeling for high-order finite difference methods on the Connection Machine CM-2. The International Journal of Supercomputer Applications and High Performance Computing, 9(1): 40–57, Spring 1995. CODEN IJSAE9. ISSN 0890-2720.

Chapra:1995:CES

[CC95b] Steven C. Chapra and David E.

Clough. Computing for engineers and scientists: Fortran. College custom series. McGraw-Hill, New York, NY, USA, preliminary edition, 1995. ISBN 0-07-011648-2. various pp.

Cherki:1998:MFP

[CC98] S. Cherki and C. Choppy. A method for FORTRAN programs reverse engineering using algebraic specifications. Lecture Notes in Computer Science, 1376:156–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743.

Ching:1993:PBS

[CCJ93] Wai-Mee Ching, Paul Carini, and Dz-Ching Ju. A primitivebased strategy for producing efficient code for very high level programs. Computer Languages, 19(1):41–50, January 1993. CODEN COLADA. ISSN 0096-0551.

Callahan:1990:IRA

[CCK90] David Callahan, Steve Carr, and Ken Kennedy. Improving register allocation for subscripted variables. ACM SIG-PLAN Notices, 25(6):53–65, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/93542/p53-callahan/.

Callahan:1986:ICP

[CCKT86] David Callahan, Keith D. Cooper, Ken Kennedy, and Linda Torczon. Interprocedural constant propagation (compilers). ACM SIGPLAN Notices, 21(7):152–161, July 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/proceedings/
                                        plan/12276/p152-callahan/.

Chang:2001:PSS

[CCL01] R-G. Chang, T-R. Chuang, and J. K. Lee. Parallel sparse supports for array intrinsic functions of Fortran 90. The Journal of supercomputing, 18 (3):305–339, March 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chang:2004:SOP

[CCL04] Rong-Guey Chang, Tyng-Ruey Chuang, and Jenq Kuen Lee. Support and optimization for parallel sparse programs with array intrinsics of Fortran 90. Parallel Computing, 30(4):527– 550, April 2004. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2004:EPI

[CCW04] Min-Bin Chen, Tyng-Ruey Chuang, and Jan-Jan Wu. Efficient parallel implementations of near Delaunay triangulation with High Performance Fortran. Concurrency and Computation: Practice and Experience, 16(12):1143–1159, October 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Caringal:1992:FIQ

[CD92] Rizaldo B. Caringal and Phan Minh Dung. A FOR-TRAN IV to QuickBASIC translator. ACM SIGPLAN Notices, 27(2):75–87, February 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coarfa:2003:ECA

[CD03] Cristian Coarfa and Yuri Dotsenko. An emerging coarray Fortran compiler (citation only). ACM SIGPLAN Notices, page ??, 2003. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chrisochoides:1993:MDH

[CDF^{+}93] N. Chrisochoides, K. Droegemeier, G. Fox, K. Mills, and Ming Xue. A methodology for developing high performance computing models: storm-scale weather prediction. In Tentner [Ten93], pages 82–89. ISBN 1-56555-052-8. LCCN ????

Conforti:1996:PIA

[CDGM96] D. Conforti, L. De Luca, L. Grandinetti, and R. Musmanno. A parallel implementation of automatic differentiation for partially separable functions using PVM. Parallel Computing, 22(5):643– 656, August 8, 1996. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1996&
volume=22&issue=5&aid=1065.

Cownie:1994:PPP

[CDH^{+}94] J. Cownie, A. Dunlop, S. Hellberg, A. J. G. Hey, and D. Pritchard. Portable parallel programming environments — the ESPRIT PPPE project. In Dekker et al. [DSZ94], pages 135–142. ISBN 0-444-81784-0. LCCN QA76.58.E98 1994.

Coarfa:2006:ESI

[CDMC06] Cristian Coarfa, Yuri Dotsenko, and John Mellor-Crummey. Experiences with Sweep3D implementations in Co-array Fortran. The Journal of supercomputing, 36(2): 101–121, May 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=36&issue=2&spage=
101.

Clemencon:1995:AEP

[CEF^{+}95] C. Clemencon, A. Endo, J. Fritscher, A. Muller, R. Ruhl, and B. J. N. Wylie. The ’annai’ environment for portable distributed parallel programming. In El-Rewini and Shriver [ERS95], pages 242–251 (vol. 2). ISBN 0-8186-6935-7. LCCN ????

Celmaster:1996:MFR

[Cel96] William N. Celmaster. Modern Fortran revived as the language of scientific parallel computing. Digital Technical Journal, 8(3):39– 45, December 1996. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v8n3/Modern_Fortran_Revived_
as_the_07jan1997DTJN03P8.
ps;   http://www.digital.
com:80/info/DTJN03/DTJN03AH.
HTM;   http://www.digital.
com:80/info/DTJN03/DTJN03HM.
HTM;   http://www.digital.
com:80/info/DTJN03/DTJN03P8.
PS;   http://www.digital.
com:80/info/DTJN03/DTJN03PF.
PDF;   http://www.digital.
com:80/info/DTJN03/DTJN03SC.
TXT.

CCI:1991:TTA

[Cen91] Century Computing Incorporated. TAE + transportable applications environment plus. Technical Report ????, National Aeronautics and Space Administration, Goddard Space Flight Center, Greenbelt, MD, USA, 1991. ???? pp.

Cann:1990:SVF

[CF90] D. Cann and J. Feo. SISAL versus Fortran: a comparison using the Livermore Loops.

In IEEE [IEE90a], pages 626– 636. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Chandy:1995:NDC

[CF95] K. Mani Chandy and Ian Foster. A notation for deterministic cooperating processes. IEEE Transactions on Parallel and Distributed Systems, 6 (8):863–871, August 1995. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://www.
computer.org/tpds/td1995/
l0863abs.htm.

Choi:1994:SSL

[CFGG94] J. Choi, E. Fernandez-Gaucherand, and D. Gerhart. S02YSCODE: a software laboratory for stochastic systems control and decision algorithms, FOR-TRAN and MATLAB versions. In IEEE CSSS CACSD’94 [IEE94g], pages 2528–2533. ISBN ????

Choudhary:1993:UCF

[CFH^{+}93] Alok Choudhary, Geoffrey Fox, Seema Hiranandani, Ken Kennedy, Charles Koelbel, Sanjay Ranka, and Chau-Wen Tseng. Unified compilation of Fortran 77D and 90D. ACM Letters on Programming Languages and Systems, 2(1–4):95– 114, March–December 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

Chandy:1994:IST

[CFK^{+}94] Mani Chandy, Ian Foster, Ken Kennedy, Charles Koelbel, and Chau-Wen Tseng. Integrated support for task and data parallelism. Int. J. Supercomputer Appl., 8(2):80–98, Summer 1994. CODEN IJSAE9. ISSN 0890-2720.

Clemencon:1995:IRD

[CFMR95] C. Clemencon, J. Fritscher, M. J. Meehan, and R. Ruhl. An implementation of race detection and deterministic replay with MPI. In Haridi et al. [HAM95b], pages 155– 166. ISBN 3-540-60247-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I553 1995.

Corbett:1994:UEP

[CFPS94] P. F. Corbett, D. G. Feitelson, J.-P. Prost, and M. Snir. User-friendly and efficient parallel I/Os using the Vesta parallel file system. In Becker et al. [BLT94], pages 23–38. ISBN 90-5199-179-7 (IOS Press), 4-274-90003-7 (Ohmsha). LCCN ????

Capolsini:1996:MMC

[CG96] P. Capolsini and C. Gomez.

MacroC and Macrofort: C and Fortran code generation within Maple. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v3n1.html.

Special Issue featuring Engineering Applications.

Chatterjee:1993:GLA

[CGL^{+}93] Siddhartha Chatterjee, John R. Gilbert, Fred J. E. Long, Robert Schreiber, and Shang-Hua Teng. Generating local addresses and communication sets for data-parallel programs. ACM SIGPLAN Notices, 28 (7):149–158, July 1993. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Calder:1995:CSB

[CGL^{+}95a] Brad Calder, Dirk Grunwald, Donald Lindsay, James Martin, Michael Mozer, and Benjamin G. Zorn. Corpusbased static branch prediction. ACM SIGPLAN Notices, 30(6):79–92, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/207110/p79-calder/.

Chatterjee:1995:GLA

[CGL^{+}95b] Siddhartha Chatterjee, John R. Gilbert, Fred J. E. Long, Robert Schreiber, and Shang-Hua H. Teng. Generating local addresses and communication sets for data-parallel programs. Journal of Parallel and Distributed Computing, 26(1): 72–84, April 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1995.1049/production;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1995.1049/production/pdf.

Chatterjee:1993:MRA

[CGS93] Siddhartha Chatterjee, John R.

Gilbert, and Robert Schreiber. Mobile and replicated alignment of arrays in data-parallel programs. In IEEE [IEE93d], pages 420–429. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Cosnard:1994:PAC

[CGS94] Michel Cosnard, Guang R.

Gao, and Gabriel M. Silberman, editors. Parallel architectures and compilation techniques: proceedings of the IFIP WG10.3 Working Conference on Parallel Architectures and Compilation Techniques, PACT ’94, Montreal, Canada, 24–26 August, 1994, volume A-50 of IFIP Transactions. A. Computer Science and Technology. North-Holland Publishing Co., Amsterdam, The Netherlands, 1994. CO-DEN ITATEC. ISBN 0-444-81926-6. ISSN 0926-5473. LCCN QA76.58 .I46 1994.

Chatterjee:1994:ADH

[CGSS94] S. Chatterjee, J. R. Gilbert, R. Schreiber, and T. J. Sheffler. Automatic distribution in HPF. In Dongarra and Tourancheau [DT94], pages 11–18. ISBN 0-89871-343-9. LCCN QA76.58.I568 1994.

Conn:1992:SLF

[CGT92a] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. LANCELOT: a Fortran package for largescale nonlinear optimization (Release A). Number 17 in Springer Series in Computational Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55470-X.

Conn:1992:LFP

[CGT92b] Andrew R. Conn, Nick I. M. Gould, and Philippe L. Toint. Lancelot: a Fortran Package for Large-Scale Nonlinear Optimization (Release A). Number 17 in Springer series in computational mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., September 1992. ISBN 3-540-55470-X (Berlin, Heidelberg), 0-387-55470-X (New York). xviii + 330 pp. LCCN QA402.5 .C648 1992. US$89.00; US$118.95. URL http:

                                        //www.cbooks.com/sqlnut/

SP/search/gtsumt?source=&
isbn=038755470X.

Cheng:1994:PDP

[CH94] Doreen Cheng and Robert Hood. A portable debugger for parallel and distributed programs. In IEEE [IEE94f], pages 723–732. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Colet:1996:WWI

[CH96] Marc Colet and Robert Herzog. WWW2GCG, a Web interface to the GCG biological sequences analysis software. Computers and Graphics, 20 (3):445–450, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http://

www.elsevier.com/cgi-bin/
cas/tree/store/cag/cas_sub/
browse/browse.cgi?year=1996&
volume=20&issue=3&aid=9600014.

Chow:1998:OFB

[CH98] Edmond Chow and Michael A.

Heroux. An object-oriented framework for block preconditioning. ACM Transactions on Mathematical Software, 24 (2):159–183, June 1998. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Chatterjee:1993:AAA

[Cha93] S. Chatterjee. Automatic array alignment in data-parallel programs. In ACM SIGPLAN POPL ’93 [ACM93c], pages 16–28. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993. ACM order number 549930.

Chabot:1994:PCA

[Cha94a] Eric Chabot. Parallelisme et communications dans les applications scientifiques Fortran. Canadian theses = Thèses canadiennes. National Library of Canada = Bibliothèque nationale du Canada, Ottawa, Ontario, Canada, 1994. ISBN 0-315-86595-4. 3 microfiches. pp.

Chapman:1994:DDD

[Cha94b] Barbara Chapman. Dynamic data distributions in Vienna Fortran. NASA contractor report NASA CR-191575; ICASE report 93-92, National Aeronautics and Space Administration, Langley Research Center; National Technical Information Service, distributor, Hampton, VA, USA, 1994. ???? pp.

Chapman:1994:FES

[Cha94c] Stephen J. Chapman. Fortran 77 for Engineering and Science: With an Introduction to Fortran 90. HarperCollins College Publishers, New York, NY, USA, October 1994. ISBN 0-06-500068-4. xxi + 666 pp. LCCN QA76.73.F25C42 1995. US$65.89. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0065000684.

Chamberland:1995:FRG

[Cha95a] Luc Chamberland. Fortran 90: a Reference Guide. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, September 1, 1995. ISBN 0-13-397332-8 (paper). xiv + 438 pp. LCCN QA 76.73 F25 C39 1995. US$49.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0133973328.

Chapman:1995:FES

[Cha95b] Stephen J. Chapman. FOR-TRAN 77 for engineers and scientists: with an introduction to FORTRAN 90. Harper-Collins, New York, NY, USA, 1995. ISBN 0-06-500068-4. xxi + 666 + 119 pp. LCCN QA 76.73 F25 C42 1995.

Chapman:1997:FSE

[Cha97a] Stephen J. Chapman. Fortran 90/95 for Scientists and Engineers. McGraw-Hill, New York, NY, USA, August 1997. ISBN 0-07-011938-4. xiv + 874 pp. LCCN QA 76.73 F25 C425 1998. US$20.01. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0070119384.

Chapman:1997:IF

[Cha97b] Stephen J. Chapman. Introduction to Fortran. Wm. C. Brown Publishers, Dubuque, IA, USA, August 1997. ISBN 0-07-011969-4.

Charlet:2009:GGA

[Cha09] Arnaud Charlet. Gem #59: generating Ada bindings for C headers. ACM SIGADA Ada Letters, 29(2):56–60, August 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Chen:1990:CTS

[Che90] Qi Chen. CTRAN: transforming scientific FORTRAN programs to UNIX based computing environments. Thesis (M.S.), Western Michigan University, Kalamazoo, MI, USA, 1990. vi + 124 pp.

Chen:1991:ACS

[Che91] Haobo Chen. Automated conversion system for development of SQL database management system from a FORTRANbased source. Thesis (M.S.), Department of Computer Science, University of California, Davis, Davis, CA, USA, 1991. 143 pp.

Chernyaev:1992:PSV

[Che92] A. P. Chernyaev. Programming systems for vectorization and parallelization of Fortran programs for some vectorpipeline computers. Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 17(2):100–111, January 1992. CODEN PCSODA. ISSN 0361-7688 (print), 1608-3261 (electronic).

Cheng:1995:ECF

[Che95] H. H. Cheng. Extending C and FORTRAN for design automation. Journal of Mechanical Design, 117(3):390– ??, 1995. CODEN JMDEEC. ISSN 1050-0472.

Calkin:1994:PPP

[CHHW94] R. Calkin, R. Hempel, H.- C. Hoppe, and P. Wypior. Portable programming with the PARMACS message-passing library. Parallel Computing, 20(4):615– 632, March 31, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1994&
volume=20&issue=4&aid=859.

Chisman:1991:ISM

[Chi91] James Chisman. Introduction to Simulation Modeling Using GPSS/PC. Prentice-Hall, Upper Saddle River, NJ 07458, USA, July 1, 1991. ISBN 0-13-497645-2. 241 pp. US$39.80.

Cypher:1993:ARP

[CHKM93] R. Cypher, A. Ho, S. Konstantinidou, and P. Messina. Architectural requirements of parallel scientific applications with explicit communication. In Proceedings of the 20th Annual International Symposium on Computer Architecture, pages 2–13. ????, ????, 1993.

Constantine:1994:FPA

[CHL94] C. C. Constantine, R. P. Hobbs, and A. J. Lymbery. FORTRAN programs for analyzing population structure from multilocus genotypic data. The Journal of Heredity, 85(4):336–??, July 1994. CO-DEN JOHEA8. ISSN 0022-1503.

Chen:1991:FCC

[CHM91] Li Jen Chen, Chin Kun Hu, and Kit Sing Mak. FOR-TRAN code for the cluster Monte Carlo study of the q-state Potts model on Ddimensional hypercubic lattices. Computer Physics Communications, 66(2–3):377– 382, September/October 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/001046559190083W.

Choi:1991:EPF

[Cho91] Won Sik Choi. Effect of Pascal and Fortran programming instruction on the problemsolving cognitive ability in formal operational stage students. Thesis (Ed.D.), Texas Tech University, Lubbock, TX, USA, 1991. x + 160 pp.

Choudhary:1992:CFM

[Cho92] Alok Nidhi Choudhary. Compiling Fortran 77D and 90D for MIMD distributed-memory machines. Computer science technical report COMP TR92-178, Rice University, Dept. of Computer Science, Houston, TX, USA, March 1992. 10 pp.

Cooper:1992:USE

[CHT92] Keith D. Cooper, Mary W.

Hall, and Linda Torczon. Unexpected side effects of inline substitution: a case study. ACM Letters on Programming Languages and Systems, 1(1): 22–32, March 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/1057-
4514/130619.html.

Creusillet:1996:IAR

[CI96] Beatrice Creusillet and François Irigoin. Interprocedural array region analyses. International Journal of Parallel Programming, 24(6):513–546, December 1996. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Creusillet:1998:IAF

[CI98] B. Creusillet and F. Irigoin.

Interprocedural analyses of Fortran programs. Parallel Computing, 24(3–4):629– 648, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/parco/
                                        sub/1998/24/3-4/1305.pdf.

Cipra:2000:BCE

[Cip00] Barry A. Cipra. The best of the 20th Century: Editors name top 10 algorithms. SIAM News, 33(4):1–2, May 2000. ISSN 0036-1437. URL

                                        https://archive.siam.org/
                                        pdf/news/637.pdf.

Cabay:1997:AEW

[CJL97] S. Cabay, A. R. Jones, and G. Labahn. Algorithm 766: Experiments with a weakly stable algorithm for computing Padé and simultaneous Padé approximants. ACM Transactions on Mathematical Software, 23(1):91–110, March 1997. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/toms/
                                        1997-23-1/p91-cabay/.

Colonna:1994:OTS

[CJPA94] F. Colonna, L.-H. Jolly, R. A. Poirier, and J. G. Angyan. OS-IPE — a tool for scientific programming in FORTRAN. Computer Physics Communications, 81(3):293–317, 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Chen:1986:ALE

[CK86] T. Y. Chen and S. C. Kwan. An analysis of length equation using a dynamic approach. ACM SIGPLAN Notices, 21 (4):42–47, April 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carnevali:1990:SMP

[CK90] P. Carnevali and M. Kindelan.

A simplified model to predict the performance of FORTRAN vector loops on the IBM 3090/ VF. Parallel Computing, 13 (1):35–46, January 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cox:1991:TSS

[CK91] Christopher L. Cox and James A. Knisely. A tridiagonal system solver for distributed memory parallel processors with vector nodes. Journal of Parallel and Distributed Computing, 13(3):325– 331, November 1991. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chung:1994:OPE

[CKMU94] J. S. Chung, K. Karal, H. Maeda, and Y. Ueda, editors. Offshore and polar engineering: 4th International conference — April 1994, Osaka, Japan. ISOPE, Golden, CO, USA, 1994. ISBN 1-880653-10-9 (paperback: set), 1-880653-11-7 (paperback: v. 1), 1-880653-12-5 (paperback: v. 2), 1-880653-13-3 (paperback: v.

3), 1-880653-14-1 (paperback: v. 4). LCCN TC1665 .I579 1994 v. 1-4 (1994).

Cooper:1985:IIA

[CKT85] Keith D. Cooper, Ken Kennedy, and Linda Torczon. The impact of interprocedural analysis and optimization on the design of a software development environment. ACM SIG-PLAN Notices, 20(7):107–116, July 1985. CODEN SINODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Choudhary:1993:HPF

[CKZ93] Alok Choudhary, Charles Koelbel, and Mary Zosel. High Performance Fortran: Implementor and users workshop. In IEEE [IEE93d], pages 610– 614. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Crovella:1993:SLC

[CL93] Mark E. Crovella and Thomas J. LeBlanc. The search for lost cycles: a new approach to parallel program performance evaluation. Technical Report TR 479, URCSD, ????, December 1993.

Crovella:1994:PPP

[CL94] Mark E. Crovella and Thomas J. LeBlanc. Parallel performance prediction using lost cycles analysis. In IEEE [IEE94f], pages 600–609. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Chandra:1997:OCH

[CL97] Satish Chandra and James R.

Larus. Optimizing communication in HPF programs on finegrain distributed shared memory. ACM SIGPLAN Notices, 32(7):100–111, July 1997. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cai:2002:TDE

[CLiN^{+}02] DongSheng Cai, Yaoting Li, Ken ichi Nishikawa, Chiejie Xiao, and Xiaoyan Yan. Threedimensional electromagnetic particle-in-cell code using High Performance Fortran on PC cluster. Lecture Notes in Computer Science, 2327:515– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2327/23270515.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270515.
pdf.

Collins:1991:DIF

[CM91] W. R. Collins and K. W.

Miller. Defining and implementing Fortran generic abstract data types. Information and Software Technology, 33(4):281–??, May 1991. CO-DEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

Campbell:1992:CFP

[CM92] Ian D. Campbell and John H. McAndrews. CANPLOT: a FORTRAN-77 program for plotting stratigraphic data on a PostScript device. Computers and Geosciences, 18(2): 309–336, March 1992. CO-DEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Campbell:1994:PGN

[CM94] S. L. Campbell and E. Moore. Progress on a general numerical method for nonlinear higher index DAEs. II. Circuits, Systems, and Signal Processing, 13 (2-3):123–138, 1994. CODEN CSSPEH.

Chapman:1998:OHI

[CM98] B. Chapman and P. Mehrotra. OpenMP and HPF: Integrating two paradigms. Lecture Notes in Computer Science, 1470:650–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cahir:2000:PMM

[CMK00] Margaret Cahir, Robert Moench, and Alice E. Koniges. Programming models and methods. In Koniges [Kon00], chapter 3, pages 27–54. ISBN 1-55860-540-1. LCCN QA76.58 .I483 2000. Discusses PVM, MPI, SHMEM, High-Performance Fortran, and POSIX threads.

Chauhan:2003:ATD

[CMKH03] Arun Chauhan, Cheryl Mc-Cosh, Ken Kennedy, and Richard Hanson. Automatic type-driven library generation for telescoping languages. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.
sc-conference.org/sc2003/
inter_cal/inter_cal_detail.
php?eventid=10692#1; http:
//www.sc-conference.org/
sc2003/paperpdfs/pap296.
pdf.

Chapman:1993:DDD

[CMMZ93] Barbara Chapman, Piyush Mehrotra, Hans Moritsch, and Hans Zima. Dynamic data distributions in Vienna Fortran. In IEEE [IEE93d], pages 284–293. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Cesare:2002:FPS

[CMP02] L. De Cesare, D. E. Myers, and D. Posa. FORTRAN programs for space-time modeling. Computers and Geosciences, 28(2): 205–212, March 2002. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Calzarossa:2001:PIH

[CMT01] M. Calzarossa, L. Massari, and D. Tessera. Performance issues of an HPF-like compiler. Future Generation Computer Systems, 18(1):147–156, September 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Caliari:2009:APL

[CMV09] Marco Caliari, Stefanode Marchi, and Marco Vianello. Algorithm 886: Padua2D — Lagrange interpolation at Padua points on bivariate domains. ACM Transactions on Mathematical Software, 35(3): 21:1–21:11, October 2009. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Chapman:1994:SAM

[CMVZ94] B. Chapman, P. Mehrotra, J. Van Rosendale, and H. Zima. A software architecture for multidisciplinary applications: integrating task and data parallelism. In Buchberger and Volkert [BV94], pages 664– 676. ISBN 3-540-58430-7 (Berlin), 0-387-58430-7 (New York). LCCN QA76.58 .J65 1994.

Chapman:1991:VFF

[CMZ91] B. Chapman, P. Mehrotra, and H. Zima. Vienna Fortran: a Fortran language extension for distributed memory multiprocessors. Technical Report 91-72, ICASE, Hampton, VA, USA, September 1991.

Chapman:1992:PVFa

[CMZ92a] Barbara Chapman, Piyush Mehrotra, and Hans Zima. Programming in Vienna Fortran. Technical Report ACPC/TR 92-3, Austrian Center for Parallel Computation, Vienna, Austria?, 1992. i + 39 pp.

Chapman:1992:PVFb

[CMZ92b] Barbara Chapman, Piyush Mehrotra, and Hans Zima. Programming in Vienna Fortran. NASA contractor report NASA CR-189623. ICASE report 92-9, National Aeronautics and Space Administration, Langley Research Center, Hampton, VA, USA, 1992. ???? pp.

Chapman:1993:HPFb

[CMZ93a] Barbara M. Chapman, Piyush Mehrotra, and Hans P. Zima. High performance Fortran without templates: an alternative model for distribution and alignment. ACM SIG-PLAN Notices, 28(7):92–101, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chapman:1993:UDM

[CMZ93b] Barbara M. Chapman, Piyush Mehrotra, and Hans P. Zima.

User defined mappings in Vienna Fortran. ACM SIGPLAN Notices, 28(1):72–75, January 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chapman:1994:HPF

[CMZ94a] B. Chapman, P. Mehrotra, and H. Zima. High performance Fortran languages: advanced applications and their implementation. In Gentzsch and Harms [GH94a], pages 407–416. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Chapman:1994:EHAa

[CMZ94b] Barbara Chapman, Piyush Mehrotra, and Hans Zima. Extending HPF for advanced data parallel applications. NASA contractor report NASA CR-194913; ICASE report no. 94-34, Institute for Computer Applications in Science and Engineering, NASA Langley Research Center, Hampton, VA, USA, 1994.

Chapman:1995:HPF

[CMZ95] Barbara Chapman, Piyush Mehrotra, and Hans Zima. High Performance Fortran languages: Advanced applications and their implementation. Future Generation Computer Systems, 11(4–5):401–407, August 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chesshire:1994:EPD

[CN94] G. Chesshire and V. K. Naik.

An environment for parallel and distributed computation with application to overlapping grids. IBM Journal of Research and Development, 38 (3):285–300, May 1994. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
almaden.ibm.com/journal/
rd38-3.html.

Cabitza:1996:EHS

[CNBB96] G. Cabitza, C. Nardone, C. Bagaini, and A. Balzano. Experiences with HPF for scientific applications. Lecture Notes in Computer Science, 1067:290–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chunduru:1991:RFP

[CNP91] Raghu K. Chunduru, R. Nagendra, and N. S. Patangay. RESDYK — a FORTRAN program for computing apparent resistivity over an infinitely deep out-cropping vertical dike. Computers and Geosciences, 17(10):1395–1408, 1991. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Cochran:2003:NVR

[Coc03] Shannon Cochran. News and views: The robots are coming; Fortran 2000 drafted; touch relayed across Atlantic; Kramnik versus Deep Fritz; 2002 ICFP Programming Content concludes. Dr. Dobb’s Journal of Software Tools, 28(1):14, January 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7758/ddj0301n/.

Cody:1990:ETR

[Cod90a]  W. J. Cody.  ELEFUNT

test results under AST Fortran V1.8.0 on the Sequent Symmetry. Technical Report MCS-TM-138, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, July 1990. ?? pp.

Cody:1990:PEP

[Cod90b]  W. J. Cody, Jr.  Per-

formance evaluation of programs for the error and complementary error functions. ACM Transactions on Mathematical Software, 16(1):29– 37, March 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
toms/1990-16-1/p29-cody/;
 http://www.acm.org/pubs/
toc/Abstracts/0098-3500/
77628.html.

Cody:1993:ASE

[Cod93a] W. J. Cody. Algorithm 715: SPECFUN—A portable FOR-TRAN package of special function routines and test drivers.

ACM Transactions on Mathematical Software, 19(1):22– 32, March 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0098-
3500/151273.html.

Cody:1993:ASP

[Cod93b] W. J. Cody. Algorithm 715:

SPECFUN—A portable FOR-TRAN package of special function routines and test drivers. ACM Transactions on Mathematical Software, 19(1):22–32, March 1993. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Coelho:1994:CIC

[Coe94a] F. Coelho. Compilation of I/O communications for HPF. In IEEE [IEE94a], pages 102–109. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Coelho:1994:EHC

[Coe94b] F. Coelho. Experiments with HPF compilation for a network of workstations. In Gentzsch and Harms [GH94a], pages 423–428. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Coelho:1996:DHD

[Coe96] F. Coelho. Discussing HPF design issues. Lecture Notes in Computer Science, 1123: 571–??, ???? 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coffee:1993:PCR

[Cof93] Peter Coffee. Peter Coffee reports that some organizations continue to depend on FOR-TRAN. PC Week, 10(50):46– ??, December 1993. ISSN 0740-1604.

Cohen:1990:GFP

[Coh90] D. R. Cohen. GOLDCALC — a FORTRAN program for estimating the number and size of gold particles in geological samples. Computers and Geosciences, 16(2):153–??, 1990. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Coker:1991:PEP

[Cok91] A. K. Coker. Program evaluates pressure drop for single phase fluids. Hydrocarbon processing (International edition), 70(2):53–??, February 1991. CODEN IHPRBS. ISSN 0018-8190.

Coker:1993:CPE

[Cok93a] A. K. Coker. Computer program enhances guidelines for gas-liquid separator designs. The Oil and Gas Journal, 91 (19):55–??, May 1993. CODEN OIGJAV. ISSN 0030-1388.

Coker:1993:PCZ

[Cok93b] A. K. Coker. Program calculates Z-factor for natural gas. The Oil and Gas Journal, 91 (7):74–??, February 1993. CO-DEN OIGJAV. ISSN 0030-1388.

Coker:1995:FPC

[Cok95] A. Kayode Coker. Fortran Programs for Chemical Process Design, Analysis, and Simulation. Gulf Publishing, Houston, TX, USA, February 1995. ISBN 0-88415-280-4. x + 854 pp. LCCN TP155.7.C65 1995 Reserves. US$125.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0884152804.

CSEP:1991:FCS

[Com91] Computational Science Education Project. Fortran 90 and Computational Science. ????, ????, 1991. ISBN ???? ???? pp. LCCN

???? URL ftp://compsci.
cas.vanderbilt.edu/CSEP/
BOOK_PS/pl.tar.Z;   http:
//csep1.phy.ornl.gov/pl/
pl.html.

Convex:1990:CFO

[Con90] Convex Computer Corporation. CONVEX FORTRAN optimization guide. CON-VEX Computer Corporation, Richardson, TX, USA, second edition, 1990. various pp.

CDC:1991:FPG

[Con91] Control Data Corporation. FORTRAN programmer’s guide and language reference manual. Control Data Corporation, Minneapolis, MN, USA, 1991. 2 pts. in 1 pp.

Conley:1992:UMA

[Con92]  Joseph L. Conley.  User’s

manual for AeroFcn a FOR-TRAN program to compute aerodynamic parameters. NASA technical memorandum 104237, National Aeronautics and Space Administration, Ames Research Center, Dryden Flight Research Facility, Edwards, CA, USA, 1992. iii + 20 pp. For sale by the National Technical Information Service.

Cooper:1994:VAA

[Coo94] Leonard Y. Cooper. VENTCF2: an algorithm and associated FORTRAN 77 subroutine for calculating flow through a horizontal ceiling/floor vent in a zone-type compartment fire model. NISTIR 5470, U.S. Dept. of Commerce, National Institute of Standards and Technology, Gaithersburg, MD, USA, 1994. iv + 20 + 22 + 25 pp.

Cooperman:1995:SBP

[Coo95] G. Cooperman. STAR/MPI: binding a parallel library to interactive symbolic algebra systems. In Levelt [Lev95a], pages 126–132. ISBN 0-89791-699-9. LCCN A 76.95 I59 1995.

Cornell:1992:B

[Cor92]  Gary Cornell.  Basics for

DOS. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, January 1, 1992. ISBN 0-8306-2200-4, 0-8306-2199-7 (paperback). xiv + 432 pp. LCCN QA76.73.B3 C678 1992. US$21.60, US$31.95 (paperback).

Costantini:1997:APC

[Cos97a]  P. Costantini.  Algorithm

770: BVSPIS—A package for computing boundary-valued shape-preserving interpolating splines. ACM Transactions on Mathematical Software, 23(2):252–254, June 1997. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1997-23-2/p252-costantini/

.

Costantini:1997:BVS

[Cos97b]  P. Costantini.  Boundary-

valued shape-preserving interpolating splines. ACM Transactions on Mathematical Software, 23(2):229–251, June 1997. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1997-23-2/p229-costantini/

.

Counihan:1991:F

[Cou91] Martin Counihan. Fortran 90.

Pitman Publishing Ltd., London, UK, 1991. ISBN 0-273-03073-6. 309 pp. LCCN QA76.73.F25 C68 1991. See [Ame92].

Counihan:1997:FIF

[Cou97] Martin Counihan. Fortran 95: Including Fortran 90, Details of High Performance Fortran (HPF), and the Fortran Module for Variable-Length Character Strings. UCL Press, January 1997. ISBN 1-85728-367-8. US$37.95. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=1857283678.

Cai:1993:TIP

[CP93] Jiazhen Cai and Robert Paige. Towards increased productivity of algorithm implementation. In Proceedings of the ACM SIGSOFT ’93 Symposium on the Foundations of Software Engineering, pages 71–78. ????, ????, December 1993.

Crooks:1994:ADD

[CP94] P. Crooks and R. H. Perrott. Automatic data distribution. In Gentzsch and Harms [GH94a], pages 463–468. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Cray:1990:CCS

[Cra90] Cray Research, Inc. CF77 compiling system. Number SR-3071, SR-3072, SR-3074 in Publication. Cray Research, Inc., Minneapolis, MN, release 4.0 edition, 1990. ISBN ???? ???? pp. LCCN ????

Cray:1991:CCS

[Cra91a] Cray Research, Inc. CF77 compiling system. Number SR-3071, SR-3072, SR-3073, SR-3074 in Publication. Cray Research, Inc., Minneapolis, MN, release 5.0 edition, 1991. ISBN ???? ???? pp. LCCN ????

Cray:1991:UFL

[Cra91b] Cray Research, Inc. UNI-COS Fortran library reference manual. Number SR-2079 6.0 in Publication. Cray Research, Inc., Minneapolis, MN, revised edition, 1991. ISBN ???? xiv + 362 pp. LCCN ????

Cray:1992:UFL

[Cra92] Cray Research, Inc. UNICOS Fortran library reference manual. Number SR-2079 in Publication. Cray Research, Inc., Minneapolis, MN, release 7.0 edition, 1992. ISBN ???? various pp. LCCN ????

Cray:1993:CFL

[Cra93] Cray Research, Inc. CF77 Fortran language reference manual: SR-3772. Number SR-3772 in Publication. Cray Research, Inc., Minneapolis, MN, release 6.0 edition, 1993. ISBN ???? xvii + 373 pp. LCCN ????

Craddock:1995:FSL

[Cra95] David M. Craddock. A FOR-TRAN 77 simulation of a low temperature storage freezer utilizing a non-azeotropic refrigerant blend. Thesis (M.S.), Ohio University, Athens, OH, USA, March 1995. xiii + 359 pp.

Contrastin:2016:UMC

[CRDO16] M. Contrastin, A. Rice, M. Danish, and D. Orchard. Units-of-measure correctness in Fortran programs. Computing in Science and Engineering, 18(1):102–107, January/February 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Crenshaw:1990:FFH

[Cre90a] Jack Crenshaw. FORTRAN fever: High-powered compilers. Computer Language Magazine, 7(5):113–??, May 1990. CO-DEN COMLEF. ISSN 0749-2839.

CrespodaSilva:1990:USC

[Cre90b] M. R. M. Crespo da Silva. On the use of symbolic computation for automating the analysis of problems in dynamics. In Kinzel et al. [KRB^{+}90], pages 593–600. ISBN 0-7918-0515-8. LCCN TA 345 A86 1990a. Two volumes.

Creak:2003:EFO

[Cre03] Alan Creak. Everything is Fortran, in its own way. ACM SIGPLAN Notices, 38(4):7–12, April 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crockford:1990:FCA

[Cro90] William Welby Crockford.

FORTRAN code for 1986 AASHTO guide equations. Journal of transportation engineering, 116(3):396–403, May/ June 1990. CODEN JTPEDI. ISSN 0733-947X.

Crosier:1991:FPT

[Cro91]  Ronald Crosier.  FOR-

TRAN programming techniques. Journal of Quality Technology, 23(4):348–354, October 1991. CODEN JQUTAU. ISSN 0022-4065.

DuCroz:1992:BLF

[Cro92] Jeremy J. Du Croz. Building libraries with Fortran 90. Fortran Journal, 4(4):??, September/October 1992. ISSN 1060-0221.

Crouch:2007:OF

[Cro07]  Peter Crouch.  The ori-

gins of Fortran. Resurrection: The Computer Conservation Society Journal, ?? (41):??, Autumn 2007. ISSN 0958-7403. URL https://

computerconservationsociety.
org/resurrection/res41.htm#
e.

Cash:1990:FPN

[CRS90] J. R. Cash, A. D. Raptis, and T. E. Simos. A Fortran program for the numerical integration of the onedimensional Schrödinger equation using exponential and Bessel fitting methods. Computer Physics Communications, 56(3):391–407, January 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
001046559090022S.

Chivers:1990:IFH

[CS90a] Ian David Chivers and Jane Sleightholme. Interactive Fortran 77: a Hands-On Approach. Ellis Horwood Series in Computers and Their Applications. Ellis Horwood, New York, NY, USA, second edition, August 1990. ISBN 0-13-466764-6. 224 pp. LCCN QA76.73.F25 C485 1990. US$72.00; US$59.95.

URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0-13-466764-
6.

Colbrook:1990:FSD

[CS90b] A. Colbrook and C. Smythe. Formal specification of data abstraction in FORTRAN 77: Abstract arrays. Software Engineering Journal, 5(3):151–??, May 1990. CODEN SEJOED. ISSN 0268-6961.

Coschi:1990:WFL

[CS90c] Geno Coschi and Jack Bernard Schueler. WATCOM FOR-TRAN 77: language reference. WATCOM Publications, Waterloo, Ontario, Canada, 1990. ISBN 0-921329-71-7. xiii + 306 pp.

Coschi:1991:WFL

[CS91] Geno Coschi and Jack Bernard Schueler. WATCOM FOR-TRAN 77: language reference. WATCOM Publications, Waterloo, Ontario, Canada, second edition, 1991. ISBN 1-55094-027-9. xiv + 316 pp.

Chivers:1995:IF

[CS95] Ian David Chivers and Jane Sleightholme. Introducing Fortran 90. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., October 1, 1995. ISBN 3-540-19940-3 (paperback). 375 pp. LCCN QA76.73.F25 C487 1995. US$34.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=3540199403.

Chivers:2000:IF

[CS00] Ian David Chivers and Jane Sleightholme. Introducing Fortran 95. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 1-85233-276-X. xvii + 455 pp. LCCN QA76.73.F29 C48 2000. US$39.95. URL

http://www.kcl.ac.uk/kis/
support/cit/fortran/f90home.
html.

Choi:2014:AMQ

[CS14] Sou-Cheng T. Choi and Michael A. Saunders. Algorithm 937: MINRES-QLP for symmetric and Hermitian linear equations and least-squares problems. ACM Transactions on Mathematical Software, 40 (2):16:1–16:12, February 2014. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cary:1997:CCF

[CSC^{+}97] John R. Cary, Svetlana G. Shasharina, Julian C. Cummings, John V. W. Reynders, and Paul J. Hinker. Comparison of C++ and Fortran 90 for object-oriented scientific programming. Computer Physics Communications, 105(1):20–36, September 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S001046559700043X.

Csendes:1999:DRC

[Cse99] Tibor Csendes, editor. Developments in Reliable Computing: Papers presented at the International Symposium on Scientific Computing, Computer Arithmetic, and Validated Numerics, SCAN-98, in Szeged, Hungary, volume 5(3) of Reliable Computing = Nadezhnye vychisleniia. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-6057-5. LCCN QA76.9.E94 D48 1999.

Coschi:1990:WFOa

[CSS90a] Geno Coschi, Jack Bernard Schueler, and Anthony F. Scian. WATCOM FORTRAN 77: optimizing compiler and tools user’s guide. WATCOM Publications, Waterloo, Ontario, Canada, 1990. ISBN 0-921329-98-9. xvii + 345 pp.

Coschi:1990:WFOb

[CSS90b] Geno Coschi, Jack Bernard Schueler, and Anthony F. Scian. WATCOM FORTRAN 77/386: optimizing compiler and tools user’s guide. WAT-COM Publications, Waterloo, Ontario, Canada, 1990. ISBN 0-921329-70-9. xvi + 317 pp.

Coschi:1991:WFO

[CSS91] Geno Coschi, Jack Bernard Schueler, and Anthony F. Scian. WATCOM FORTRAN 77/386: optimizing compiler and tools user’s guide. WAT-COM Publications, Waterloo, Ontario, Canada, second edition, 1991. ISBN 1-55094-025-2.

Cowell:1990:TAD

[CT90] Wayne R. Cowell and Christopher P. Thompson. Tools to aid in discovering parallelism and localizing arithmetic in Fortran programs. Software —Practice and Experience, 20 (1):25–47, January 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Carpaneto:1995:ACS

[CT95] M. Dell’Amico G. Carpaneto and P. Toth. Algorithm 750: CDT: a subroutine for the exact solution of large-scale asymmetric travelling salesman problems. ACM Transactions on Mathematical Software, 21 (4):410–415, December 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cooper:2011:COC

[CT11] Keith D. Cooper and Linda Torczon. Classic optimizing compilers: IBM’s Fortran H compiler. Lecture slides

(25) for Comp 512 course at Rice University, Houston, TX, USA., Spring 2011. URL http:

//booksite.elsevier.com/
9780120884780/Graduate_Lecture_
Slides/Core_Lectures/02FortranH.

ppt. From slide 12: “[The IBM Fortran H] compiler was just 27,415 lines of Fortran + 16,721 lines of asm [assembly code]”.

Chang:1996:DIF

[CTS96] C-Y Chang, J-Y Tzeng, and J-P Sheu. Design and implementation of a Fortran assistant tool for vector compilers. International Journal of High Speed Computing, 8(1):13–46, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Cumbest:1990:AFP

[Cum90] R. J. Cumbest. AMPHAX: a FORTRAN program for calculating the lower hemisphere stereographic projections of the crystallographic axes of clinoamphibole from universal stage measurements. Computers and Geosciences, 16 (3):371–??, 1990. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Curlett:1994:NGG

[Cur94] Brian P. Curlett. NLEdit, a generic graphical user interface for Fortran programs. NASA technical memorandum 4543, National Aeronautics and Space Administration, Office of Management, Scientific and Technical Information Program; National Technical Information Service, distributor, Washington, DC, USA, 1994. ???? pp.

Calamai:1994:AFS

[CV94] Paul H. Calamai and Luis N.

Vicente. Algorithm 728: FOR-TRAN subroutines for generating quadratic bilevel programming test problems. ACM Transactions on Mathematical Software, 20(1):120–123, March 1994. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://dial.uwaterloo.ca/
~phcalama/cpp.html.

Clint:1992:CTF

[CWB92] M. Clint, J. S. Weston, and C. W. Bleakney. A comparison of two Fortran dialects for expressing parallel solutions for a problem in linear algebra. Parallel Computing, 18(12):1325– 1333, December 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Clint:1994:CPF

[CWB94] M. Clint, J. S. Weston, and C. W. Bleakney. Comparison of parallel Fortran environments on the AMT DAP510 for a linear algebra application. Concurrency, practice and experience, 6(3):193–204, May 1994. CODEN CPEXEI. ISSN 1040-3108.

Cybenko:1991:PCPa

[Cyb91]  George Cybenko.  Parallel

computing and the Perfect Benchmarks. Technical Report CSRD 1191, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1991. 12 pp.

Chiou:1990:UGQ

[CZ90] Hung-Wen Chiou and Evanghelos Zafiriou. User’s guide for QDMC version 1.0: a set of Fortran programs for constrained quadratic dynamic matrix control simulation and stability/performance study. Technical research report TR 90-29, University of Maryland, Systems Research Center, College Park, MD, USA, 1990. 23 pp.

Celledoni:2010:AFF

[CZ10] Elena Celledoni and Antonella Zanna. Algorithm 903: FRB — Fortran routines for the exact computation of free rigid body motions. ACM Transactions on Mathematical Software, 37(2): 23:1–23:24, April 2010. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Chapman:1993:HDD

[CZM93a] B. Chapman, H. Zima, and P. Mehrotra. Handling distributed data in Vienna Fortran procedures. Lecture Notes in Computer Science, 757:248– 263, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chapman:1993:HPFa

[CZM93b] Barbara Chapman, Hans Zima, and Piyush Mehrotra. High Performance Fortran without templates an alternative model for distribution and alignment. NASA contractor report NASA CR-191451; ICASE report 93-17, National Aeronautics and Space Administration, Langley Research Center; National Technical Information Service, distributor, Hampton, VA, USA, 1993. ???? pp.

Chapman:1994:HFL

[CZM94a] B. Chapman, H. Zima, and P. Mehrotra. High-Performance Fortran languages: Advanced applications and their implementation. In Gentzsch and Harms [GH94b], pages 407– 416. ISBN 3-540-57981-8 (Berlin: vol. 2: paperback), 0-387-57981-8 (New York: vol. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1– 2 (c1994). DM96.00. Two volumes.

Chapman:1994:EHAb

[CZM94b] Barbara Chapman, Hans Zima, and Piyush Mehrotra. Extending HPF for advanced dataparallel applications. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):59–70, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Danpitakkul:1990:DFP

[Dan90] Dejchai Danpitakkul. Development of a Fortran program to calculate supersonic flow using the method of characteristics. Thesis (M.S.), California State University, Northridge, Northridge, CA, USA, 1990. vii + 133 pp.

Das:2006:OSO

[Das06] Dibyendu Das. Optimizing subroutines with optional parameters in F90 via function cloning. ACM SIGPLAN Notices, 41(8):21–28, August 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dubois:1993:PFS

[DB93a] Paul F. Dubois and Lee Busby. Portability and the Fortran standard. Computers in physics, 7(2):162–165, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Dubois:1993:PPF

[DB93b] Paul F. Dubois and Lee Busby. Portable, powerful Fortran programs. Computers in physics, 7(1):38–44, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.4823145.

Dubois:1993:SP

[DB93c] Paul F. Dubois and Lee Busby.

Scientific programming. Computers in physics, 7(1):38– 44, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Cunha:1994:PPI

[dCH94] R. D. da Cunha and Tim Hopkins. PIM 1.0: The parallel iterative methods package for systems of linear equations. user’s guide (Fortran 77 version). Technical Report 2-94*, Computing Laboratory, University of Kent, Canterbury, University of Kent, Canterbury, UK, January 1994.

URL ftp://unix.hensa.ac.
uk/pub/misc/ukc.reports/
comp.sci/reports/2-94.ps.
Z.

Dongarra:19xx:ESF

[DCHH88a] Jack J. Dongarra, Jeremy J. Du Croz, Sven J. Hammarling, and Richard J. Hanson. An extended set of Fortran basic linear algebra subprograms. ACM Transactions on Mathematical Software, 14 (1):1–17, 1988. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dongarra:19xx:AES

[DCHH88b] Jack J. Dongarra, Jeremy J. Du Croz, Sven J. Hammarling, and Richard J. Hanson. Algorithm 656. An extended set of Fortran basic linear algebra subprograms: Model implementation and test programs. ACM Transactions on Mathematical Software, 14(1): 18–32, 1988. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

DiMartino:1999:HPF

[DCR99a] B. Di Martino, M. Celino, and V. Rosato. A[n] High Performance Fortran implementation of a tight-binding molecular dynamics simulation. Computer Physics Communications, 120(2–3):255–268, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

DiMartino:1999:HPM

[DCR99b] B. Di Martino, M. Celino, and V. Rosato. HPF parallelization of a molecular dynamics code: Strategies and performances. Lecture Notes in Computer Science, 1593:535– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dwarkadas:1996:ICT

[DCZ96] Sandhya Dwarkadas, Alan L.

Cox, and Willy Zwaenepoel. An integrated compile-time/ run-time software distributed shared memory system. ACM SIGPLAN Notices, 31(9):186– 197, September 1996. CODEN SINODQ. ISBN 0-89791-767-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/asplos/237090/
p186-dwarkadas/. Co-published

as SIGOPS Operating Systems Review 30(5), December 1996, and as SIGARCH Computer Architecture News, 24(special issue), October 1996.

Doolin:1997:JCL

[DD97] David M. Doolin and Jack Dongarra. JLAPACK — compiling LAPACK Fortran to Java, phase 1. Technical report CS-97-367, University of Tennessee, Knoxville, Knoxville, TN 37996, USA,

1997.  URL http://www.
netlib.org/utk/people/JackDongarra/
PAPERS/f2jpaper.ps;   http:
//www.netlib.org/utk/people/
JackDongarra/pdf/f2jpaper.
pdf.

Dayde:1999:RBB

[DD99] Michel J. Daydé and Iain S. Duff. The RISC BLAS: a blocked implementation of Level 3 BLAS for RISC processors. ACM Transactions on Mathematical Software, 25(3): 316–340, September 1999. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Darte:1996:TRT

[DDcMR96] A. Darte, F. Desprez, J. c. Mignot, and Y. Robert. TransTOOL: a restructuring tool for the parallelization of applications using High Performance Fortran. Journal of the Brazilian Computer Society, 3(2): 5–12, ???? 1996. ISSN 0104-6500 (print), 1678-4804 (electronic). URL http:/

                                        /www.cs.utk.edu/~yrobert/
                                        journals/transtool.ps.gz.

DAmbra:2010:MPP

[DDF10] Pasqua D’Ambra, Daniela Di Serafino, and Salvatore Filippone. MLD2P4: a package of parallel algebraic multilevel domain decomposition preconditioners in Fortran 95. ACM Transactions on Mathematical Software, 37(3):30:1– 30:23, September 2010. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dongarra:1995:PFI

[DDH^{+}95] J. J. Dongarra, J. Du Croz, S. Hammarling, J. Wasniewski, and A. Zemla. A proposal for a Fortran 90 interface for LAPACK. LAPACK Working Note 101, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, July 1995. URL http:

//www.netlib.org/lapack/
lawns/lawn101.ps;   http:
//www.netlib.org/lapack/
lawnspdf/lawn101.pdf. UT-

CS-95-295, July 1995.

Dongarra:1996:LF

[DDH^{+}96] J. J. Dongarra, J. Du Croz, S. Hammarling, J. Waśniewski, and A. Zemla. LAPACK for Fortran 90. Applied Mathematics and Computer Science, 6 (2):375–382, ???? 1996. CO-DEN AMCPE9. ISSN 0867-

857X.  URL http://www.
netlib.org/utk/people/JackDongarra/
PAPERS/lapack90.ps.

Denner:2017:CFB

[DDH17] Ansgar Denner, Stefan Dittmaier, and Lars Hofer. Collier: a Fortran-based complex one-loop library in extended regularizations. Computer Physics Communications, 212(??):220–238, March 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465516303320.

Dongarra:1990:ASL

[DDHD90] Jack J. Dongarra, Jeremy Du Croz, Sven Hammarling, and Iain Duff. Algorithm 679: a set of Level 3 Basic Linear Algebra Subprograms: Model implementation and test programs. ACM Transactions on Mathematical Software, 16(1):18–28, March 1990. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [Hig90b, DH92, DDP94].

Dongarra:1996:LFC

[DDHW96a] J. J. Dongarra, J. Du Croz, S. Hammarling, and J. Wasniewski. LAPACK for Fortran90 compiler. Lecture Notes in Computer Science, 1067:826–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1996:PFI

[DDHW96b] J. J. Dongarra, J. Du Croz, S. Hammarling, and J. Wasniewski. A proposal for a Fortran 90 interface for LAPACK. Lecture Notes in Computer Science, 1041:158–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dayde:1994:PBI

[DDP94] Michael J. Daydé, Iain S. Duff, and Antoine Petitet. A parallel block implementation of level-3 BLAS for MIMD vector processors. ACM Transactions on Mathematical Software, 20 (2):178–193, June 1994. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [DDHD90, Hig90b, DH92].

Doolin:1999:JCL

[DDS99] David M. Doolin, Jack Dongarra, and Keith Seymour. JLAPACK — compiling LA-PACK FORTRAN to Java. Scientific Programming, 7 (2):111–138, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.com/
app/home/contribution.asp%
3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C4%2C8%3Bjournal%
2C7%2C9%3Blinkingpublicationresults%
2C1%2C1;   http://www.netlib.
org/utk/people/JackDongarra/
PAPERS/f2jrep~1.pdf. The

software is available on the World-Wide Web at http://

www.cs.utk.edu/f2j/.

Decyk:1993:HWN

[Dec93] Viktor K. Decyk. How to write (nearly) portable Fortran programs for parallel computers. Computers in physics, 7 (4):418–425, July 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Delannoy:1993:PFG

[Del93] Claude Delannoy. Programmer en Fortran 90: guide complet. Eyrolles, Paris, France, 1993. ISBN 2-212-08723-3. xix + 413 pp.

Delves:1998:HPL

[Del98] Mike Delves. HPF: Programming Linux clusters the easy way. Linux Journal, 45:??, January 1998. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL

                                        ftp://ftp.ssc.com/pub/lj/
                                        listings/issue45/2432.tgz.

Demetriou:1995:ALF

[Dem95] I. C. Demetriou. Algorithm 742: L2CXFT: A Fortran subroutine for least squares data fitting with nonnegative second divided differences. ACM Transactions on Mathematical Software, 21(1):98–110, March 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Demetriou:1997:CFS

[Dem97] I. C. Demetriou. CXFTV2: a Fortran subroutine for the discrete least squares convex approximation. Computer Physics Communications, 100(3):297–310, March 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0010465596001580.

Demetriou:2003:LFP

[Dem03] I. C. Demetriou. L1PMA: a Fortran 77 package for best L_{1} piecewise monotonic data smoothing. Computer Physics Communications, 151(3):315–338, April 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465502007397.

Demetriou:2006:LFP

[Dem06] I. C. Demetriou. L2CXCV:

a Fortran 77 package for least squares convex/concave data smoothing. Computer Physics Communications, 174(8):643–668, April 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465506000075.

Demetriou:2007:ALF

[Dem07] Ioannis C. Demetriou. Algorithm 863: L2WPMA, a Fortran 77 package for weighted least-squares piecewise monotonic data approximation. ACM Transactions on Mathematical Software, 33(1):6:1– 6:19, March 2007. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

DeTar:1990:FAP

[DeT90] DeLos F. DeTar. FORTRAN and the art of PC programming, by Tim Ward and Eddie Bromhead. Computers and Chemistry, 14(3):259–??, 1990. CODEN COCHDK. ISSN 0097-8485.

Das:2012:NFC

[DET12] Debottam Das, Ulrich Ellwanger, and Ana M. Teixeira. NMSDECAY: a Fortran code for supersymmetric particle decays in the Next-to-Minimal Supersymmetric Standard Model. Computer Physics Communications, 183(3):774–779, March 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0010465511003833.

Deutsch:1990:FSD

[Deu90] Clayton Deutsch. A FOR-TRAN 77 subroutine for determining the fractional area of rectangular grid blocks within a polygon. Computers and Geosciences, 16(3):379– ??, 1990. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

DeVries:1994:FCC

[DeV94] Paul L. DeVries. A First Course in Computational Physics. John Wiley and Sons, New York, London, Sydney, 1994. ISBN 0-471-54869-3. xii + 424 pp. LCCN QC 52 D48 1994. US$70.95. System requirements for computer disk: IBM-compatible PC; DOS; Microsoft FORTRAN compiler; hard disk drive.

Devloo:1992:CIP

[DFL92] R. B. Devloo, L. Fezoui, and S. Lacire. A C++ interface for programming on the Connection Machine. Technical Report RR1698, INRIA-SOPHIA, Sophia-Antipolis, France, May 1992.

Dongarra:1991:PLT

[DFRR91] J. Dongarra, M. Furtney, S. Reinhardt, and J. Russell. Parallel loops — a test suite for parallelizing compilers: description and example results. Parallel Computing, 17(10–11): 1247–1255, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Doi:1995:FSL

[DFS95] Shun Doi, Hideo Fujio, and Kouta Sugihara. FEEL: a simulation language for finite element analysis. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/
proceedings/doi/doi1.html.

Detert:1994:TTS

[DG94] Ulrich Detert and Michael Gerndt. TOP 2 — tool suite for the development and testing of parallel applications. Technical Report KFA-ZAM-IB-9415, KFA Research Centre, Jülich, Jülich, Germany, 1994. 10

pp. URL ftp://ftp.zam.kfa-
juelich.de/pub/zamdoc/ib/
ib-94/ib-9415.ps.

DiDonato:1999:TFC

[DG99] Armido R. DiDonato and Russ Gnoffo. Testing a Fortran 90 compiler using the NSWC Fortran 77 Mathematics Library. Technical Report NSWCDD/TR-98/75, Naval Surface Warfare Center, Dahlgren, VA 22448-5100, USA, February 1999. v + 64 pp. URL https:

                                        //apps.dtic.mil/sti/pdfs/
                                        ADA360604.pdf.

Decyk:2008:OOD

[DG08] Viktor K. Decyk and Henry J. Gardner. Object-oriented design patterns in Fortran 90/95: mazev1, mazev2 and mazev3. Computer Physics Communications, 178(8):611–620, April 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0010465507004857.

Dodson:19xx:SEF

[DGL91a] D. S. Dodson, R. G. Grimes, and J. G. Lewis. Sparse extensions to the FORTRAN basic linear algebra subprograms. ACM Transactions on Mathematical Software, 17(2):253– 263, 1991. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dodson:1991:AMI

[DGL91b] David S. Dodson, Roger G. Grimes, and John G. Lewis.

Algorithm 692: Model implementation and test package for the sparse Basic Linear Algebra Subroutines. ACM Transactions on Mathematical Software, 17(2):264–272, June 1991. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1991-17-2/p264-dodson/.

Dodson:1991:SEF

[DGL91c] David S. Dodson, Roger G.

Grimes, and John G. Lewis. Sparse extensions to the FORTRAN Basic Linear Algebra Subroutines. ACM Transactions on Mathematical Software, 17(2):253–263, June 1991. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0098-3500/
108577.html.

Dongarra:1990:FSC

[DGR90] J. J. Dongarra, George A.

Geist, and C. H. Romine. Fortran subroutines for computing the eigenvalues and eigenvectors of a general matrix by reduction to general tridiagonal form. Technical report CS-90-116, University of Tennessee, Computer Science Dept., Knoxville, TN, USA, 1990. 13 pp.

Dongarra:1992:AFS

[DGR92] J. J. Dongarra, G. A. Geist, and C. H. Romine. Algorithm 710: FORTRAN subroutines for computing the eigenvalues and eigenvectors of a general matrix by reduction to general tridiagonal form. ACM Transactions on Mathematical Software, 18(4):392– 400, December 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

                                        URL http://www.acm.org/
                                        pubs/toc/Abstracts/0098-
                                        3500/138352.html.

Degrassi:2008:SFC

[DGS08] G. Degrassi, P. Gambino, and P. Slavich. SusyBSG: a Fortran code for BR[B → Xsy] in the MSSM with Minimal Flavor Violation. Computer Physics Communications, 179 (10):759–771, November 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0010465508002312.

Dongarra:1984:CPL

[DH84] J. J. Dongarra and R. E. Hiromoto. A collection of parallel linear equations routines for the Denelcor HEP. Parallel Computing, 1(2):133–142, December 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Demmel:1992:SBA

[DH92] James W. Demmel and Nicholas J. Higham. Stability of block algorithms with fast level-3 BLAS. ACM Transactions on Mathematical Software, 18(3): 274–291, September 1992. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [DDHD90, Hig90b, DDP94].

Doytsher:1995:FPC

[DH95] Yerahmiel Doytsher and John K.

Hall. FORTRAN programs for coordinate resection using an oblique photograph and highresolution DTM. Computers and Geosciences, 21(7):895– ??, 1995. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Dubrau:2012:TM

[DH12] Anton Willy Dubrau and Laurie Jane Hendren. Taming MATLAB. ACM SIGPLAN Notices, 47(10):503–522, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Duff:2002:OSB

[DHP02] Iain S. Duff, Michael A. Heroux, and Roldan Pozo. An overview of the Sparse Basic Linear Algebra Subprograms: The new standard from the BLAS Technical Forum. ACM Transactions on Mathematical Software, 28(2):239–267, June 2002. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dumitrascu:1990:GFP

[DI90] Liviu Dumitraşcu and Alexandru Ioachim. Generating FOR-TRAN programs from decision tables. Editura Academiei, Bucureşti, Romania, 1990. ISBN 973-27-0231-1. viii + 312 pp. LCCN T57.6 .D85613 1990. Translation of: Tehnici de construire a programelor cu structuri alternative.

DEC:1990:VFH

[Dig90a] Digital Equipment Corporation. VAX FORTRAN HPO installation guide /release notes. Digital Equipment Corporation, Maynard, MA, USA, 1990. various pp.

DEC:1990:VFP

[Dig90b] Digital Equipment Corporation. VAX FORTRAN performance guide. Digital Equipment Corporation, Maynard, MA, USA, 1990. various pp.

DEC:1992:DFL

[Dig92] Digital Equipment Corporation. DEC Fortran: language reference manual. Digital Equipment Corporation, Maynard, MA, USA, 1992. various pp.

DEC:1993:DFI

[Dig93a] Digital Equipment Corporation. DEC Fortran: installation guide for open VMS VAX systems. Digital Equipment Corporation, Maynard, MA, USA, 1993. various pp.

DEC:1993:DFP

[Dig93b] Digital Equipment Corporation. DEC Fortran: performance guide for open VMS VAX systems. Digital Equipment Corporation, Maynard, MA, USA, 1993. ISBN ???? various pp. LCCN ????

DEC:1993:DFU

[Dig93c] Digital Equipment Corporation. DEC Fortran: user manual for open VMS VAX systems. Digital Equipment Corporation, Maynard, MA, USA, 1993. various pp.

Ding:1999:HPF

[Din99] Chris H. Q. Ding. High Performance Fortran for practical scientific algorithms: an up-to-date evaluation. Future Generation Computer Systems, 15 (3):343–??, ???? 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dehbonei:1992:SIA

[DJ92] Babak Dehbonei and Pierre Jouvelot. Semantical interprocedural analysis by partial symbolic evaluation. In PEPM [PEP92], pages 14–20. ISBN ???? LCCN ????

Djouadi:2007:SFC

[DKM07] Abdelhak Djouadi, Jean-Loïc Kneur, and Gilbert Moultaka. SuSpect: a Fortran code for the supersymmetric and Higgs particle spectrum in the MSSM.

Computer Physics Communications, 176(6):426–455, March 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S001046550600419X.

Dongarra:1991:GBP

[DKMS91] J. J. Dongarra, A. Karp, K. Miura, and H. D. Simon. Gordon Bell Prize lectures (supercomputer applications). In IEEE [IEE91], pages 328–337. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

DeSturler:1997:IIS

[DL97a] E. De Sturler and D. Loher. Implementing iterative solvers for irregular sparse matrix problems in High Performance Fortran. Lecture Notes in Computer Science, 1336:293–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSturler:1997:PSI

[DL97b] E. De Sturler and D. Loher. Parallel solution of irregular, sparse matrix problems using High Performance Fortran. Lecture Notes in Computer Science, 1225:360–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dekeyser:1997:HBV

[DL97c] J.-L. Dekeyser and C. Lefebvre. HPF-Builder: a visual environment to transform Fortran 90 codes to HPF. The International Journal of Supercomputer Applications and High Performance Computing, 11(2):95–102, Summer 1997. CODEN IJSCFG. ISSN 1078-3482.

deLeeuw:2012:BRM

[dL12] Jan de Leeuw. Book review:

Modern Fortran: Style and Usage. Journal of Statistical Software, 47(BR-1):??, April 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v47/b01.

Lima:1995:PFP

[dLJEB95] B. S. L. P. de Lima, B. P. Jacob, N. F. F. Ebecken, and A. C. Benjamin. Portable FORTRAN programming tools in the development of a structural analysis program. Computers and structures, 57(6): 1109–??, 1995. CODEN CM-STCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Dasgupta:1996:QSF

[DLLR96] Indranil Dasgupta, Andrea Ruben Levi, Vittorio Lubicz, and Claudio Rebbi. QCDF90: a set of Fortran 90 modules for a high-level, efficient implementation of QCD simulations.

Computer Physics Communications, 98(3):365–397, November 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0010465596001038.

DAmore:1999:IFS

[DLM99a] Luisa D’Amore, Giuliano Laccetti, and Almerico Murli. An implementation of a Fourier series method for the numerical inversion of the Laplace transform. ACM Transactions on Mathematical Software, 25(3):279–305, September 1999. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
articles/journals/toms/1999-
25-3/p279-d_amore/p279-d_
amore.pdf;   http://www.
acm.org/pubs/citations/journals/
toms/1999-25-3/p279-d_amore/
;   http://www.acm.org/pubs/
citations/journals/toms/
1999-25-3/p279-d_amore/#
abstract;   http://www.acm.
org/pubs/citations/journals/
toms/1999-25-3/p279-d_amore/
#indterms.

DAmore:1999:AFS

[DLM99b] Luisa D’Amore, Guiliano Laccetti, and Almerico Murli. Algorithm 796: a Fortran software package for the numerical inversion of the Laplace transform based on a Fourier series method. ACM Transactions on Mathematical Software, 25(3):

306–315, September 1999. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dobmann:1995:APF

[DLS95] M. Dobmann, M. Liepelt, and K. Schittkowski. Algorithm 746: PCOMP: A Fortran code for automatic differentiation. ACM Transactions on Mathematical Software, 21(3):233– 266, September 1995. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dong:2018:TFL

[DLW^{+}18] Shao-Jun Dong, Wen-Yuan Liu, Chao Wang, Yongjian Han, G.-C. Guo, and Lixin He. TNSPackage: A Fortran2003 library designed for tensor network state methods. Computer Physics Communications, 228(??):163–177, July 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S001046551830078X.

Drouffe:1990:FPF

[DM90] J.-M. Drouffe and K. J. M.

Moriarty. FORTRAN program for the four-dimensional Ising model. International Journal of High Speed Computing, 2(2): 133–142, June 1990. CODEN IHSCEZ. ISSN 0129-0533.

Decyk:2004:SMI

[DN04] Viktor K. Decyk and Charles D. Norton. A simplified method for implementing run-time polymorphism in Fortran95. Scientific Programming, 12(1): 45–55, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

DAlberto:2009:AWM

[DN09] Paolo D’Alberto and Alexandru Nicolau. Adaptive Winograd’s matrix multiplications. ACM Transactions on Mathematical Software, 36(1):3:1– 3:23, March 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Decyk:2007:WF

[DNG07] Viktor K. Decyk, Charles D. Norton, and Henry J. Gardner. Why Fortran? Computing in Science and Engineering, 9(4): 68–71, July/August 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Decyk:1997:HEC

[DNS97] Viktor K. Decyk, Charles D. Norton, and Boleslaw K. Szymanski. How to express C++ concepts in Fortran90. Scientific Programming, 6(4):363–390, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.com/
                                        app/home/contribution.asp%

3Fwasp=a5tkxhqqy9eefb7hwkak%
26referrer=parent%26backto=
issue%2C3%2C7%3Bjournal%
2C9%2C9%3Blinkingpublicationresults%
2C1%2C1.

Decyk:1998:HSI

[DNS98] Viktor K. Decyk, Charles D.

Norton, and Boleslaw K. Szymanski. How to support inheritance and run-time polymorphism in Fortran 90. Computer Physics Communications, 115(1):9–17, December 1, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465598001015.

Dongarra:1990:PVCa

[Don90] J. J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. ?? CS-89-85, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, ?? 1990.

Dongarra:1991:PVC

[Don91] J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. ?? CS-89-85, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, ?? 1991.

Dongarra:1995:HPC

[Don95] J. J. Dongarra, editor. High performance computing: technology, methods and applications: Advanced workshop — June 1994, Cetraro, Italy, volume 10 of Advances in Parallel Computing — Amsterdam. Elsevier, Amsterdam, The Netherlands, 1995. ISBN 0-444-82163-5. ISSN 0927-5452. LCCN QA76.88 .H54 1995.

Dietz:1992:F

[DOP^{+}92] H. Dietz, M. O’Keefe, T. Parr, T. Varghese, and P. Woodward. Fortran-P. Presentation for Army High Performance Computing Research Center Research Advisory Committee, University of Minnesota, March 1992.

Dongarra:1996:MPS

[DOSW96] Jack J. Dongarra, Steve W. Otto, Marc Snir, and David Walker. A message passing standard for MPP and workstations. Comm. ACM, 39(7): 84–90, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/cacm/
234000.html.

Dotson:1993:MAP

[Dot93] Don W. Dotson. Microcomputer analysis of plate bending by the finite element method in FORTRAN, BASIC, PAS-CAL, and C languages using rectangular elements. Thesis (M.E.), Tennessee State University, Nashville, TN, USA, 1993. x + 201 pp.

Dowd:1993:HPC

[Dow93] Kevin Dowd. High Performance Computing. RISC architectures, optimization and benchmarks; A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-032-5. xxv + 371 pp. LCCN QA76.88 .D6 1993; QA76.9.A73 D68 1993. US$25.95. URL http:/

/www.oreilly.com/catalog/
9781565920323.

DeRoeck:1994:CFP

[DP94] Y. H. De Roeck and R. E.

Plessix. Combining F90 and PVM to construct synthetic seismograms by ray-tracing. In IEEE [IEE94c], pages II–653– II–658. ISBN 0-7803-2057-3, 0-7803-2056-5, 0-7803-2058-1. ISSN 0197-7385. LCCN TC 1505 O33197 1994. Three volumes. IEEE catalog no. 94CH3472-8.

DeRose:1996:MFT

[DP96] L. De Rose and D. Padua. A MATLAB to Fortran 90 translator and its effectiveness. In ACM [ACM96a], pages 309– 316. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

DeRose:1999:TTM

[DP99] Luiz De Rose and David Padua. Techniques for the translation of MATLAB programs into Fortran 90. ACM Transactions on Programming Languages and Systems, 21(2): 286–323, March 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
toplas/1999-21-2/p286-de_
rose/.

Dion:1994:PCW

[DPR94] M. Dion, J.-L. Philippe, and Y. Robert. Parallelizing compilers: what can be achieved? In Gentzsch and Harms [GH94a], pages 447–456. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Diele:2002:FRS

[DPS02] F. Diele, T. Politi, and I. Sgura. A Fortran90 routine for the solution of orthogonal differential problems. Lecture Notes in Computer Science, 2331:449– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2331/23310449.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2331/23310449.
pdf.

Dou:1997:ISV

[DPZ97] Yong Dou, Zhengbing Pang, and Xingming Zhou. Implementing a software virtual shared memory on PVM. In IEEE [IEE97], page ?? ISBN 0-8186-7876-3 (paperback and case), 0-8186-7878-X (microfiche). LCCN QA76.58 .A4 1997.

Duff:1993:MAF

[DR93a] I. S. Duff and J. K. Reid.

MA48, A Fortran code for direct solution of sparse unsymmetric linear systems of equations. Report RAL-93-072, Rutherford Appleton Laboratory, Chilton, Didcot, Berks, UK, 1993.

Duff:1993:MFC

[DR93b] I. S. Duff and J. K. Reid.

MA48, A Fortran code for direct solution of sparse unsymmetric linear systems of equations. Report RAL-93-072, Rutherford Appleton Laboratory, Chilton, Didcot, Berks, UK, 1993.

Decker:1994:PEM

[DR94a] K. M. Decker and R. M.

Rehmann, editors. Programming environments for massively parallel distributed systems: working conference of the IFIP WG10.3, April 25–29, 1994, Ascona, Italy. Birkhäuser Boston Inc., Cambridge, MA, USA, 1994. ISBN 0-8176-5090-3. LCCN QA76.58 .P767 1994.

Duff:1994:MFC

[DR94b] Iain S. Duff and J. K. Reid.

MA47, a Fortran code for direct solution of sparse symmetric indefinite structured systems of linear equations. Technical Report (to appear), Rutherford Appleton Laboratory, Chilton, Didcot, Berks, UK, 1994.

Duff:1995:MFC

[DR95a] I. S. Duff and J. K. Reid. MA47, a Fortran code for direct solution of indefinite sparse symmetric linear systems. Report RAL 95-001, Rutherford Appleton Laboratory, Chilton, Didcot, Berks, UK, ???? 1995. ???? pp.

dure95

[DR95b] Iain S. Duff and John K. Reid. MA47, a Fortran code for direct solution of indefinite sparse symmetric linear systems. Report RAL-95-001, Rutherford Appleton Laboratory, Chilton, Didcot, Berks, UK, January 1995. 63 pp.

Drezner:1992:CMN

[Dre92] Zvi Drezner. Computation of the multivariate normal integral. ACM Transactions on Mathematical Software, 18(4):470–480, December 1992. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/toms/
                                        1992-18-4/p470-drezner/.

See also [Dre93].

Drezner:1993:CAC

[Dre93]  Zvi Drezner.  Corrigen-

dum: “Algorithm 725. computation of the multivariate normal integral”. ACM Transactions on Mathematical Software, 19(4):546, December 1993. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1993-19-4/p546-drezner/.

See [Dre92].

Diaz:2003:DIP

[DRST03] Manuel Díaz, Bartolomé Rubio, Enrique Soler, and José M. Troya. Domain interaction patterns to coordinate HPF tasks. Parallel Computing, 29(7):925– 951, July 2003. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Douglas:1994:VMM

[DS94] Craig C. Douglas and Gordon Slishman. Variants of matrix-matrix multiplication for Fortran-90. ACM SIGNUM Newsletter, 29(2):4–6, April 1994. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

DeSturler:1997:SPH

[DS97] Eric De Sturler and Volker Strumpen. Scientific programming with High Performance Fortran: a case study using the xHPF compiler. Scientific Programming, 6(1):127– 152, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Denissen:2001:EDL

[DS01] Will Denissen and Henk J. Sips. Efficient dynamic local enumeration for HPF. Lecture Notes in Computer Science, 2017:355–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/2017/20170355.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2017/20170355.
pdf.

Denissen:2002:FPB

[DS02] Will Denissen and Henk J. Sips. Finding performance bugs with the TNO HPF benchmark suite. Concurrency and Computation: Practice and Experience, 14(8–9): 691–712, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/95016124/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016124{\&}PLACEBO=IE.
pdf.

deSturler:1998:PIS

[dSL98] E. de Sturler and D. Loher. Parallel iterative solvers for irregular sparse matrices in High Performance Fortran. Future Generation Computer Systems, 13(4):315–325, March 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Das:1994:SAI

[DSv94] R. Das, J. Saltz, and R. von Hanxleden. Slicing analysis and indirect accesses to distributed arrays. In Banerjee et al. [BGNP94], pages 152–168. ISBN 3-540-57659-2 (Berlin), 0-387-57659-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1993. DM122.00.

Dekker:1994:MPP

[DSZ94] L. (Leendert) Dekker, W. Smit, and J. C. Zuidervaart, editors. Massively parallel processing applications and development: proceedings of the 1994 EUROSIM Conference on Massively Parallel Processing Applications and Development, Delft, The Netherlands, 21–23 June 1994. Elsevier, Amsterdam, The Netherlands, 1994. ISBN 0-444-81784-0. LCCN QA76.58.E98 1994.

AlvesdaSilva:1992:FCO

[dSZP92] Alexandre P. Alves da Silva, Victor Hugo Quintana Z., and G. K. H. Pang. Fortran codes for the optimal estimate training 2, and probabilistic associative memory algorithms. Internal report UWE and CE-92-02, Dept. of Electrical and Computer Engineering, University of Waterloo, Waterloo, Ontario, Canada, 1992. 81 pp.

Dehnert:1993:CC

[DT93] James C. Dehnert and Ross A. Towle. Compiling for the Cydra 5. The Journal of supercomputing, 7(1–2): 181–227, May 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://www.
                                        springerlink.com/openurl.
                                        asp?genre=article&issn=0920-
                                        8542&volume=7&issue=1&spage=
                                        181.

Dongarra:1994:ETP

[DT94] J. J. Dongarra and B. Tourancheau, editors. Environments and tools for parallel scientific computing: 2nd Workshop — May 1994, Townsend, TN, Proceedings of the Workshop on Environments and Tools for Parallel Scientific Computing. SIAM Press, Philadelphia, PA, USA, 1994. ISBN 0-89871-343-9. LCCN QA76.58.I568 1994.

DuCroz:1997:DBN

[Du 97] J. Du Croz. Designing and building a new numerical library in Fortran 90. In Boisvert [Boi97], pages 197– 209. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Dubois:1997:BRM

[Dub97] Paul F. Dubois. Book review: Mark G. Gray and Randy M. Roberts, Object-Based Programming in Fortran 90. Computers in physics, 11(4):355–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/
10.1063/1.4822571.

Duff:2004:MCS

[Duf04] Iain S. Duff. MA57—a code for the solution of sparse symmetric definite and indefinite systems. ACM Transactions on Mathematical Software, 30 (2):118–144, June 2004. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dutta:1994:TFP

[Dut94] Dhruba J. Dutta. TRANS4: a FORTRAN program for computing apparent resistivity departure curves for an infinitely thick bed with transitional invaded zone in borehold geophysics. Computers and Geosciences, 20(3):293–??, April 1994. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Duval:1992:TPP

[Duv92] D. Duval. Trends in parallel programming models for high performance computers. In Ferenczi [Fer92], page 33. ISBN ???? LCCN ???? KFKI preprint/report series: KFKI-1992-34/M,N.

Dubesset:1991:FLN

[DV91] Michel Dubesset and Jean Vignes. Fortran: le langage normalisé. Collection informatique. Editions Technip, Paris, France, 1991. ISBN 2-7108-0611-8. xix + 236 pp.

Dongarra:1992:PVCa

[DV92] J. J. Dongarra and H. A. Van der Vorst. Performance of various computers using standard linear equations software in a Fortran environment. Supercomputer, 9(5):17–30, September 1992. CODEN SPCOEL. ISSN 0168-7875.

Dubesset:1993:SDF

[DV93] Michel Dubesset and Jean Vignes. Les specificités du Fortran 90. Collection informatique. Editions Technip, Paris, France, 1993. ISBN 2-7108-0652-5. xxiii + 367 pp.

Duff:1998:LLB

[DV98] I. S. Duff and C. Vömel. Level 2 and Level 3 Basic Linear Algebra Subprograms for sparse matrices: a Fortran 95 instantiation. Report TR/PA/00/18, CERFACS, Toulouse, France, ???? 1998. ???? pp.

DeTisi:2000:RAS

[DV00] Flavia De Tisi and Alba Valtulina. Remark on Algorithm 761: scattered-data surface fitting that has the accuracy of a cubic polynomial. ACM Transactions on Mathematical Software, 26(1):46–48, March 2000. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
2000-26-1/p46-de_tisi/;
http://www.acm.org/pubs/
citations/journals/toms/
2000-26-1/p46-de_tisi/p46-
de_tisi.pdf.  See [Aki96,

RB98].

Duff:2001:ISB

[DV01] I. S. Duff and C. Vömel.

The implementation of the Sparse BLAS in Fortran 95. Report TR/PA/01/27, CER-FACS, Toulouse, France, ???? 2001. ???? pp.

Duff:2002:ARM

[DV02a] Iain S. Duff and Christof Vömel. Algorithm 818: a reference model implementation of the Sparse BLAS in Fortran 95. ACM Transactions on Mathematical Software, 28(2):268– 283, June 2002. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Duff:2002:AXR

[DV02b] Iain S. Duff and Christof Vömel. Algorithm xxx: a reference model implementation of the Sparse BLAS in Fortran 95. Report RAL-TR-2002-018, Rutherford Appleton Laboratory, Chilton, Didcot, Berks, UK, 2002. ???? pp. URL

http://epubs.cclrc.ac.uk/
bitstream/289/raltr-2002018.
pdf.

deVivo:1997:PUO

[dVdVI97] Gabriela O. de Vivo, Marco de Vivo, and Germinal Isern. E pluribus unum: OOPL selection. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29 (2):17–ff., June 1997. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Duff:2000:ISB

[DVY00] I. S. Duff, C. Vömel, and M. Youan. Implementing the Sparse BLAS in Fortran 95. Report TR/PA/00/82, CER-FACS, Toulouse, France, ???? 2000. ???? pp.

Dongarra:1994:PSC

[DW94] Jack Dongarra and Jerzy Wasniewski, editors. Parallel scientific computing: First International Workshop, PARA ’94, Lyngby, Denmark, June 20– 23, 1994: proceedings, volume 879 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00.

Dwyer:2003:PRN

[DW03] Gerald P. Dwyer, Jr. and K. B. Williams. Portable random number generators. Journal of Economic Dynamics and Control, 27(4):645–650, February 2003. CODEN JED-CDH. ISSN 0165-1889 (print), 1879-1743 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0165188901000653.

DaSilva:2008:I

[DW08] Dilma Da Silva and Robert W.

Wisniewski. Introduction. Operating Systems Review, 42(1): 1, January 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dubois:1999:SPE

[DY99] Paul F. Dubois and T.-Y.

Yang. Scientific programming: Extending Python with Fortran. Computing in Science and Engineering, 1(5):66–73, September/October 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

URL http://dlib.computer.
org/cs/books/cs1999/pdf/
c5066.pdf;   http://www.
computer.org/cse/cs1999/
c5066abs.htm.

Dyadkin:1994:MP

[Dya94] Lev J. Dyadkin. Multibox parsers. ACM SIGPLAN Notices, 29(7):54–60, July 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dyadkin:1995:MPN

[Dya95] Lev J. Dyadkin. Multibox parsers. no more handwritten lexical analyzers. IEEE Software, 12(5):61–67, September 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Delves:1998:HPF

[DZ98] M. Delves and H. Zima. High Performance Fortran: a status report or: Are we ready to give up MPI? Lecture Notes in Computer Science, 1497:161– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evripidou:1998:MFP

[EB98] Paraskevas Evripidou and Robert Barry. Mapping Fortran programs to single assignment semantics for efficient parallelization. Parallel Processing Letters, 8(3):407– ??, September 1998. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Erricolo:2013:AFS

[EC13] Danilo Erricolo and Giuseppe Carluccio. Algorithm 934: Fortran 90 subroutines to compute Mathieu functions for complex values of the parameter. ACM Transactions on Mathematical Software, 40(1):8:1– 8:19, September 2013. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Einarsson:1996:FT

[ECS96] Bo Einarsson, Leonid Chubarov, and Yurij Shokin. Fortran 95 Tutorial. ????, ????, 1996. ISBN ???? ???? pp. LCCN ???? English translation-in-progress of [Ein94].

Etzel:1999:DVF

[ED99] Michael Etzel and Karen Dickinson. Digital Visual Fortran Programmer’s Guide. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1999. ISBN 1-55558-218-4. 755 (est.) pp. LCCN ???? US$49.95. URL http:

//www.compaq.com/fortran/
visual/pg-book.html.

Eichenberger:1996:MRR

[EDA96] Alexandre E. Eichenberger, Edward S. Davidson, and Santosh G. Abraham. Minimizing register requirements of a modulo schedule via optimum stage scheduling. International Journal of Parallel Programming, 24(2):103–132, April 1996. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Edelson:1990:NMF

[Ede90] David Edelson. Numerical methods with FORTRAN 77: a practical introduction, by L. V. Atkinson et al. Computers and Chemistry, 14(2):183– ??, 1990. CODEN COCHDK. ISSN 0097-8485.

Edgar:1992:FPS

[Edg92] Stacey L. Edgar. Fortran for the ’90s: Problem Solving for Scientists and Engineers. Computer Science Press, Inc., 11 Taft Court, Rockville, MD 20850, USA, January 1, 1992.

ISBN 0-7167-8247-2. xxx + 658 pp. LCCN QA76.73.F25 E35 1992. US$49.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0716782472.

Enenkel:2005:CMF

[EFG^{+}05] R. F. Enenkel, B. G. Fitch, R. S. Germain, F. G. Gustavson, A. Martin, M. Mendell, J. W. Pitera, M. C. Pitman, A. Rayshubskiy, F. Suits, W. C. Swope, and T. J. C. Ward. Custom math functions for molecular dynamics. IBM Journal of Research and Development, 49 (2/3):465–474, ???? 2005. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
rd/492/enenkel.pdf.

Erwig:2007:PFP

[EFP07] Martin Erwig, Zhe Fu, and Ben Pflaum. Parametric Fortran: program generation in scientific computing. Journal of Software Maintenance and Evolution: Research and Practice, 19 (3):155–182, May 2007. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Ehold:1999:HNL

[EGKU99] H. J. Ehold, W. N. Gansterer, D. F. Kvasnicka, and C. W. Ueberhuber. HPF and numerical libraries. Lecture Notes in Computer Science, 1557:140– 152, 1999. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehold:2002:OLP

[EGKU02] H. J. Ehold, W. N. Gansterer, D. F. Kvasnicka, and C. W. Ueberhuber. Optimizing local performance in HPF. Parallel Computing, 28(3):415–432, March 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ellwanger:2007:NFC

[EH07a] Ulrich Ellwanger and Cyril Hugonie. NMSPEC: a Fortran code for the sparticle and Higgs masses in the NMSSM with GUT scale boundary conditions. Computer Physics Communications, 177(4):399–407, August 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465507002263.

Enright:2007:RRD

[EH07b] W. H. Enright and Wayne B.

Hayes. Robust and reliable defect control for Runge–Kutta methods. ACM Transactions on Mathematical Software, 33 (1):1:1–1:19, March 2007. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Eigenmann:1991:RFP

[EHJ^{+}91] R. Eigenmann, J. Hoeflinger, G. Jaxon, Z. Li, and D. Padua.

Restructuring Fortran programs for Cedar. In Proceedings of the 1991 International Conference on Parallel Processing, volume I, Architecture, pages I–57–I–66. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, August 1991.

Eigenmann:1993:RFP

[EHJ^{+}93] R. Eigenmann, J. Hoeflinger, G. Jaxon, Z. Li, and D. Padua. Restructuring Fortran programs for Cedar. Concurrency, practice and experience, 5(7): 553–574, October 1993. CO-DEN CPEXEI. ISSN 1040-3108.

Eigenmann:1990:CFC

[Eig90a] R. Eigenmann. Cedar Fortran and its compiler. Technical Report CSRD 966, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1990. 12 pp.

Eigenmann:1990:CFR

[Eig90b] R. Eigenmann. Cedar Fortran and its restructuring compiler. Technical Report CSRD 1041, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1990. 22 pp.

Einarsson:1991:WCP

[Ein91] B. Einarsson, editor. Working Conference on Programming Environments for High– Level Scientific Problem Solving (Karlsruhe, Germany). Internation Federation for Information Processing, IBM Corporation, New York, NY, USA, September 23–27, 1991. ISBN ???? LCCN ????

Einarsson:1994:LF

[Ein94] Bo Einarsson. Lärobok i Fortran 90/95. Nationellt Superdatorcentrum, Linköpings universitet, Linköping, Sweden, 1994. ISBN ???? ???? pp. LCCN ???? Revised 1997, 1999. An English translation is in progress [ECS96].

Einarsson:1995:MLP

[Ein95] Bo Einarsson. Mixed language programming. Part 4, mixing ANSI-C with Fortran 77 or Fortran 90. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/
proceedings/einarsson/einarsson.
html.

Einarsson:1996:SET

[Ein96] Bo Einarsson. Some experiences from teaching Fortran 90. Fortran Journal, 8(1):2, 4–6, January/February 1996. ISSN 1060-0221.

Evans:1997:ACG

[EJLC97] E. W. Evans, S. P. Johnson, P. F. Leggett, and M. Cross. Automatic code generation of overlapped communications in a parallelisation tool. Parallel Computing, 23(10):1493– 1523, November 7, 1997. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1997&
volume=23&issue=10&aid=1227.

Englezos:2001:APE

[EK01] Peter Englezos and Nicolas Kalogerakis. Applied parameter estimation for chemical engineers, volume 81 of Chemical industries. Marcel Dekker, New York, NY, USA, 2001. ISBN 0-8247-9561-X. xvii + 434 pp. LCCN TP155.2.M35 E54 2001. Includes CD-ROM.

El-Khouri:1992:MFP

[EKB92] Bassam El-Khouri and Lars R. Bergman. M-prep: a Fortran 77 program for the preparatory analysis of multivariate data. Report 751, Department of Psychology, Stockholm University, Stockholm, Sweden, March 1992. 7 pp.

Ewer:1995:CSI

[EKC95] J. Ewer, B. Knight, and D. Cowell. Case study: An incremental approach to reengineering a legacy FOR-TRAN computational fluid dynamics code in C++. Advances in engineering software, 22(3):153–??, 1995. CODEN AESODT. ISSN 0965-9978 (print), 0141-1195 (electronic).

Engstler:1997:MEM

[EL97] Ch. Engstler and Ch. Lubich. Multirate extrapolation methods for differential equations with different time scales. Computing, 58(2):173–185, 1997. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.
springer.at/springer.py?
Page=10&Key=362&cat=300607/
tocs/springer.py?Page=47&
Key=340&cat=3&id_abstract=
1221&id_volume=93&id_journal=
8.

Elisseev:1998:PTD

[Eli98] V. V. Elisseev. Parallelization of three-dimensional spectral laser–plasma interaction code using High Performance Fortran. Computers in physics, 12(2):173–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.168630.

Elliott:1981:FSD

[Ell81] D. G. Elliott. FORTRAN 77 and structured design. ACM SIGPLAN Notices, 16(12):7–9, December 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ellis:1990:FPI

[Ell90] T. M. R. Ellis. Fortran 77 Programming: With an Introduction to Fortran 90 Standard. International Computer Science Series. Addison-Wes-ley, Reading, MA, USA, second edition, 1990. ISBN 0-201-41638-7 (paperback). xxi + 641 pp. LCCN QA76.73.F25 E43 1990. US$38.75. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0-201-41638-7. Rev. ed.

of: A structured approach to FORTRAN 77 programming.

Emerson:1994:BRH

[Eme94] D. R. Emerson. Book review: The high performance Fortran handbook, Scientific and engineering computation series: Charles H. Koelbel, David B. Loveman, Robert S. Schreiber, Guy L. Steele Jr. and Mary E. Zosel, M.I.T. Press, Fitzroy House, 11, Chenies Street, London WC1E 7ET, England, ISBN 0-262-61094-9, January 1994, Price: $24.95 (Paperback). Journal of Computational and Applied Mathematics, 54(3):N4, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042794902631.

Engeln-Mullges:1993:NFG

[EMR93] Gisela Engeln-Müllges and Fritz Reutter. Numerik-Algorithmen mit FORTRAN 77-Programmen. (German) [Numerical algorithms with FORTRAN 77-programs]. Bibliographisches Institut, Mannheim, 1993. ISBN 3-411-15117-X. xxvi + 1245 pp. Seventh edition. With an appendix by Jürgen Dietel et al. With a separately available computer disk.

Engeln-Mullges:1996:NAF

[EMU96] Gisela Engeln-Mullges and Frank Uhlig. Numerical Algorithms With Fortran. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., November 1996. ISBN 3-540-60529-0. xxii + 602 pp. LCCN QA297 .E56213 1996b Permanent Reserve. US$49.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=3540605290.

Engeln-Mullges:1998:BRB

[EMUP98] Gisela Engeln-Mullges, Frank Uhlig, and John D. Pryce. Book reviews: Numerical Algorithms with C and Numerical Algorithms with Fortran. SIAM Review, 40(1):171–??, March 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Erkal:1996:TTS

[EN96] A. Erkal and D. T. Numbere. Tracking thermal saturation fronts by a high level PC programming language. Proceedings — Petroleum Computer Conference, pages 83–89, 1996. CODEN PPCMEG.

Enright:1995:REC

[Enr95] Wayne Enright. Reliable error control for ODE methods. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.nsc.liu.se/~boein/
                                        ifip/kyoto/workshop-info/
                                        proceedings/enright/enright0.
                                        html.

Escaig:1991:ATM

[EO91] Y. Escaig and W. Oed. Analysis tools for micro- and autotasking programs on CRAY multiprocessor systems. Parallel Computing, 17(12):1425– 1433, December 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Espelid:1994:DAAa

[EO94] T. O. Espelid and K. J. Overholt. DQAINF: an algorithm for automatic integration of infinite oscillating tails. Numerical Algorithms, 8(1):83–101, ???? 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Enright:19xx:TFP

[EP87] W. H. Enright and J. D. Pryce. Two FORTRAN packages for assessing initial value methods. ACM Transactions on Mathematical Software, 13(1): 1–27, 1987. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Erhel:1992:DTC

[EP92] J. Erhel and B. Philippe.

Design of a toolbox to control arithmetic reliability. In Atanassova and Herzberger [AH92], pages 99– 108. ISBN 0-444-89834-4. LCCN QA76.9.C62I559 1992.

Ellis:1994:FP

[EPL94a] T. M. R. Ellis, Ivor R. Philips, and Thomas M. Lahey. Fortran 90 Programming. International computer science series. Addison-Wesley, Reading, MA, USA, May 1994. ISBN 0-201-54446-6. 825 pp. US$40.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0201544466.

Ellis:1994:FPI

[EPL94b] T. M. R. Ellis, Ivor R. Philips, and Thomas M. Lahey. Fortran 90 Programming (International Computer Science Series). Addison-Wesley, Reading, MA, USA, June 1, 1994. ISBN 0-201-54446-6. US$45.25.

Ellis:1995:FP

[EPL95] T. M. R. Ellis, Ivor R. Philips, and Thomas M. Lahey. Fortran 90 programming. International computer science series. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-54446-6. xviii + 825 pp. LCCN QA 76.73 F25 E435 1994.

Epstein:1994:CCa

[Eps94a] David Epstein. Conditional compilation. Fortran Journal, 6(??):??, May/June 1994. ISSN 1060-0221.

Epstein:1994:CCb

[Eps94b] David Epstein. Conditional compilation. Fortran Journal, 6(??):??, September/October 1994. ISSN 1060-0221.

Epstein:1996:CC

[Eps96] David Epstein. Conditional compilation. Dr. Dobb’s Journal of Software Tools, 21(5):??, May 1996. CODEN DDJTEQ. ISSN 1044-789X.

Erricolo:2006:AFS

[Err06] Danilo Erricolo. Algorithm 861: Fortran 90 subroutines for computing the expansion coefficients of Mathieu functions using Blanch’s algorithm. ACM Transactions on Mathematical Software, 32(4):622– 634, December 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

El-Rewini:1995:PTH

[ERS95] H. El-Rewini and B. D. Shriver, editors. Proceedings of the Twenty-Eighth Hawaii International Conference on System Sciences. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-6935-7. LCCN ????

Einarsson:1993:FFP

[ES93a] Bo Einarsson and Yurij Shokin.

Fortran 90 for the Fortran 77 Programmer. ????, ????, 1993. ???? pp. Also available in Russian [ES93b].

Einarsson:1993:FKD

[ES93b] Bo Einarsson and Yurij Shokin.

FORTRAN-90, Kniga dlja programmiruyushchikh na yazyke Fortran-77. Izdatel’stvo Sibirskogo Otdeleniya Rossijskaya Akademiya Nauk (Siberian Division of the Russian Academy of Sciences), Novosibirsk, Russia, 1993. ISBN 5-85826-013-6. ???? pp. LCCN

???? URL http://www.nsc.
liu.se/~boein/edu/ryska.

html. Also available in English [ES93a].

Espelid:1998:RAD

[Esp98] Terje O. Espelid. Remark on Algorithm 706: DCUTRI — an algorithm for adaptive cubature over a collection of triangles. ACM Transactions on Mathematical Software, 24(3):334–335, September 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
toms/1998-24-3/p334-espelid/

. See [BE92].

Etter:1990:SFE

[Ett90] D. M. (Delores M.) Etter.

Structured Fortran 77 for Engineers and Scientists. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, third edition, 1990. ISBN 0-8053-0051-1. xxi + 567 + [4] pp. LCCN QA76.73.F25 E85 1990. US$15.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0805300511.

Etter:1992:FNM

[Ett92] D. M. Etter. Fortran 77: With Numerical Methods for Engineers and Scientists. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, February 1992. ISBN 0-8053-1770-8. xxv + 776 pp. LCCN QA76.73.F25 E847 1992. US$56.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0805317708. System re-

quirements for computer disk (ASCII files): PC; FORTRAN 77 compiler.

Etter:1993:SFE

[Ett93]  Delores M. Etter.  Struc-

tured Fortran 77 for Engineers and Scientists. Benjamin/ Cummings Pub. Co., Menlo Park, CA, USA, fourth edition, 1993. ISBN 0-8053-1775-9. xxiii + 616 pp. LCCN QA76.73.F25 E85 1993.

Etter:1996:SFE

[Ett96] Delores M. Etter. Structured Fortran 77 for Engineers and Scientists. Addison-Wesley, Reading, MA, USA, November 1996. ISBN 0-201-49854-5. US$46.25.

Etter:1997:SFE

[Ett97]  D. M. Etter.  Structured

Fortran 77 for Engineers and Scientists. Addison-Wesley, Reading, MA, USA, fifth edition, February 1997. ISBN 0-201-49854-5. 612 pp. US$42.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0201498545.

Fabijonas:2004:AAF

[Fab04] B. R. Fabijonas. Algorithm 838: Airy functions. ACM Transactions on Mathematical Software, 30(4):491–501, December 2004. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fahringer:1994:UPG

[Fah94] T. Fahringer. Using the P03T to guide the parallelization and optimization effort under the Vienna Fortran Compilation System. In IEEE [IEE94d], pages 437–444. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994.

Fahoome:2002:JRF

[Fah02] Gail F. Fahoome. JMASM1:

RANGEN 2.0 (Fortran 90/95). Journal of Modern Applied Statistical Methods, 1(1):182–190, Winter 2002. CODEN ???? ISSN 1538-9472. URL http:/

/tbf.coe.wayne.edu/jmasm/.

Fatoohi:1994:ANS

[Fat94] R. A. Fatoohi. Adapting a Navier–Stokes solver for three parallel machines. The Journal of supercomputing, 8(2): 91–115, June 1994. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=8&issue=2&spage=
91.

Filippone:2012:OOT

[FB12] Salvatore Filippone and Alfredo Buttari. Object-oriented techniques for sparse matrix computations in Fortran 2003. ACM Transactions on Mathematical Software, 38(4):23:1– 23:20, August 2012. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Frazier:1996:CAF

[FBC96] Thomas P. Frazier, John W. Bailey, and Melissa L. Corso. Comparing Ada and FOR-TRAN lines of code: Some experimental results. Empirical Software Engineering, 1 (1):45–59, ???? 1996. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/BF00125811.

Fateman:1995:FFP

[FBWR95] Richard J. Fateman, Kevin A. Broughan, Diane K. Willcock, and Duane Rettig. Fast floating point processing in Common Lisp. ACM Transactions on Mathematical Software, 21 (1):26–62, March 1995. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fahringer:1992:APP

[FBZ92] T. Fahringer, R. Blasko, and H. P. Zima. Automatic performance prediction to support parallelization of Fortran programs for massively parallel systems. In 6th ACM International Conference on Supercomputing, pages 347–356. ????, Washington, DC, USA, July 1992.

Foster:1992:FAL

[FC92] Ian T. Foster and K. Mani Chandy. Fortran M: a language for modular parallel programming. Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, IL, June 1992.

Foster:1995:FML

[FC95] Ian T. Foster and K. Mani Chandy. FORTRAN M: a language for modular parallel programming. Journal of Parallel and Distributed Computing, 26 (1):24–35, April 1, 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1995.1044/production;
 http://www.idealibrary.

                                        com/links/doi/10.1006/jpdc.
                                        1995.1044/production/pdf.

Feng:2009:JDA

[FC09] Du Feng and Normal Clif. JMASM 29: Dominance analysis of independent data (Fortran). Journal of Modern Applied Statistical Methods, 8 (2):??, ???? 2009. CO-DEN ???? ISSN 1538-9472.

                                        URL https://jmasm.com/
                                        index.php/jmasm/article/
                                        view/459.

Fenton:2002:RTC

[FCHE02] F. H. Fenton, E. M. Cherry, H. M. Hastings, and S. J. Evans. Real-time computer simulations of excitable media: Java as a scientific language and as a wrapper for C and Fortran programs. Biosystems (A6E), 64(1):73–96, January 2002. CODEN BSYMBO. ISSN 0303-2647.

Feibus:1994:SP

[Fei94] Andy Feibus. Standards in progress. Open Systems Today, ??(163):20–??, November 1994. ISSN 1061-0839.

Fenner:1996:FEM

[Fen96] Roger T. Fenner. Finite Element Methods for Engineers. Imperial College Press and distributed by World, March 1996. ISBN 1-86094-016-1. US$18.00.

Ferenczi:1992:AWT

[Fer92] S. Ferenczi, editor. Proceedings of the 1st Austrian-Hungarian Workshop on Transputer Applications: October 8–10, 1992, Sopron, Hungary. Hungarian Academy of Sciences Central Research Institute for Physics, Budapest, Hungary, 1992. ISBN ???? LCCN ???? KFKI preprint/report series: KFKI-1992-34/M,N.

Fritzson:2005:MSC

[FES05] Peter Fritzson, Vadim Engelson, and Krishnamurthy Sheshadri. MathCode: a system for C++ or Fortran code generation from Mathematica. Mathematica Journal, 10(4):??, ???? 2005. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Fitzsimons:1993:PGD

[FG93] C. J. Fitzsimons and C. Greenough. A programming guide for the development of engineering applications software in Fortran. Technical Report RAL 93-055, Rutherford Appleton Laboratory, Chilton, Didcot, Berks, UK, 1993. ii + 27 pp.

Fanfarillo:2019:RCA

[FGBN19] Alessandro Fanfarillo, Sudip Kumar Garain, Dinshaw Balsara, and Daniel Nagle. Resilient computational applications using Coarray Fortran. Parallel Computing, 81(??):58–67, January 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0167819118303867.

Fernandez-Gaucherand:1994:SSS

[FGCG94] E. Fernández-Gaucherand, J. Choi, and D. Gerhart. S^{2}YSCODE: Stochastic systems control and decision algorithms software laboratory and FORTRAN and Matlab versions. In IEEE CSSS CACSD’94 [IEE94g], pages 179–186. ISBN ????

Fraysse:2009:ASF

[FGG09] Valérie Frayssé, Luc Giraud, and Serge Gratton. Algorithm 881: a set of flexible GMRES routines for real and complex arithmetics on highperformance computers. ACM Transactions on Mathematical Software, 35(2):13:1–13:12, July 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fraysse:2005:ASG

[FGGL05] Valŕie Frayssé, Luc Giraud, Serge Gratton, and Julien Langou. Algorithm 842: a set of GMRES routines for real and complex arithmetics on high performance computers. ACM Transactions on Mathematical Software, 31(2):228– 238, June 2005. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

FroeseFischer:2019:GFV

[FGJB19] C. Froese Fischer, G. Gaigalas, P. Jönsson, and J. Bieroń. GRASP2018 — a Fortran 95 version of the General Relativistic Atomic Structure Package. Computer Physics Communications, 237(??):184–187, April 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465518303928.

Faber:2001:IAG

[FGL01] Peter Faber, Martin Griebl, and Christian Lengauer. Issues of the automatic generation of HPF loop programs. Lecture Notes in Computer Science, 2017:359– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2017/20170359.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2017/20170359.
pdf.

Feldman:1990:FCCa

[FGMS90a] S. I. Feldman, D. M. Gay, M. W. Maimone, and N. L. Schryer. A Fortran-to-C converter. Computing Science Technical Report No. 149, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1990.

URL ftp://ftp.netlib.com/
netlib/f2c/f2c.ps.Z.

Feldman:1990:FCCb

[FGMS90b] S. I. Feldman, D. M. Gay, M. W. Maimone, and N. L.

Schryer. A Fortran to C converter. ACM SIGPLAN FOR-TRAN Forum, 9(2):21–22, October 1990. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Feld90a

[FGMS90c] S. I. Feldman, David M. Gay, M. W. Maimone, and N. L. Schryer. A Fortran-to-C converter. Computing Science Technical Report No. 149, AT&T Bell Laboratories, Murray Hill, NJ 07974, 1990.

Feldman:1990:FC

[FGMS90d] S. I. Feldman, David M. Gay, M. W. Maimone, and N. L. Schryer. A Fortran-to-C converter. Computing Science Technical Report 149, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1990.

Feldman:1993:FC

[FGMS93] S. I. Feldman, David M. Gay, Mark W. Maimone, and N. L. Schryer. A Fortran-to-C converter. Computer Science Technical Report 149, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1993.

Feldman:1995:FCC

[FGMS95] S. I. Feldman, David M. Gay, Mark W. Maimone, and N. L. Schryer. A Fortran-to-C converter. Computing Science Technical Report 149, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 22, 1995. i + 25 pp. Revision of edition of 18 May 1990.

Fahringer:2000:PMH

[FGRT00] Thomas Fahringer, Michael Gerndt, Graham Riley, and Jesper Larsson Träff. On performance modeling for HPF applications with ASL. Lecture Notes in Computer Science, 1940:191–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/1940/19400191.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1940/19400191.
pdf.

Forsley:1990:RFC

[FH90] Lawrence P. Forsley and Thomas Hess, editors. 1990 Rochester FORTH Conference: embedded systems, June 12–16th, 1990, University of Rochester. Institute for Applied FORTH Research, Rochester, NY, USA, 1990. ISBN 0-914593-10-2. LCCN QA76.73.F24 R59 1990. US$35.00. Spine title: Embedded systems. Includes abstracts and bibliographical references.

Flanders:1992:PMC

[FH92] Bob Flanders and Michael Holmes. PC Magazine C Lab Notes. Ziff-Davis Press, Emeryville, CA, USA, September 1, 1992. ISBN 1-56276-063-7. 312 pp. US$29.95. Includes disk.

Fournet:1995:FPS

[FHE95] F. Fournet, F. Hospital, and J. M. Elsen. A FORTRAN program to simulate the evolution of genetic variability in a small population. Computer applications in the biosciences: CABIOS, 11(5):469–475, October 1995. CODEN COABER. ISSN 0266-7061.

Fox:1990:FLS

[FHK^{+}90a] G. Fox, S. Hiranandani, K. Kennedy, C. Koelbel, U. Kremer, C. Tseng, and M. Wu. Fortran D language specification. Technical Report TR90-141, Department of Computer Science, Rice University, Houston, TX, USA, 1990.

Fox:1990:FDL

[FHK^{+}90b] G. Fox, S. Hiranandani, K. Kennedy, C. Koelbel, U. Kremer, C.-W. Tseng, and M.-Y. Wu. Fortran D language specification. Computer Science Department Report CRPC-TR90079, Rice University, Houston, TX, USA, December 1990.

Forth:2012:RAA

[FHP^{+}12] Shaun Forth, Paul Hovland, Eric Phipps, Jean Utke, and Andrea Walther, editors. Recent Advances in Algorithmic Differentiation, volume 87 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSA6. ISBN 3-642-30022-7 (print), 3-642-30023-5 (e-book). ISSN 1439-7358. LCCN ???? URL http:

//link.springer.com/book/
10.1007/978-3-642-30023-
3;   http://www.springerlink.
com/content/978-3-642-30023-

3. Proceedings of the Sixth International Conference on Automatic Differentiation (AD2012) held July 23–27, 2012, in Fort Collins, Colorado, USA.

Fox:1978:AFP

[FHS78] P. A. Fox, A. D. Hall, and N. L. Schryer. Algorithm 528: Framework for a portable library [Z]. ACM Transactions on Mathematical Software, 4 (2):177–188, June 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remarks [Fox79, GG99].

Fisher:1992:DTO

[FJ92] M. E. Fisher and L. S. Jennings. Discrete-time optimal control problems with general constraints. ACM Transactions on Mathematical Software, 18(4):401–413, December 1992. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1992-18-4/p401-fisher/.

Facchinei:1997:AFS

[FJS97] Francisco Facchinei, Joaquim Júdice, and João Soares. Algorithm 774. FORTRAN subroutine for generating box constrained optimization problems. ACM Transactions on Mathematical Software, 23(3):448–450, September 1997. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/toms/
                                        1997-23-3/p448-facchinei/.

Fosdick:1996:IHP

[FJSD96] Lloyd D. Fosdick, Elizabeth R. Jessup, Carolyn J. C. Schauble, and Gitta Domik. An Introduction to High-Performance Scientific Computing. MIT Press, Cambridge, MA, USA, 1996. ISBN 0-262-06181-3 (hardcover). xxiii + 760 pp. LCCN QA76 .A594 1996. US$70.00.

Ferenczi:1995:PAW

[FK95] Szabolcs Ferenczi and Peter Kacsuk, editors. Proceedings of the 2nd Austrian-Hungarian Workshop on Transputer Applications (KFKI-1995-2/M, N), September 29–October 1, 1994, Budapest, Hungary, volume 2 of Hungarian Academy of Sciences Central Research Institute for Physics — Publications, KFKI //M,N. Hungarian Acad. Sci, Budapest, Hungary, 1995. ISBN ???? LCCN ????

Foster:1996:DSB

[FKKC96] Ian Foster, David R. Kohr, Jr., Rakesh Krishnaiyer, and Alok Choudhary. Double standards: Bringing task parallelism to HPF via the message passing interface. In ACM [ACM96b], page ?? ISBN 0-89791-854-1. LCCN QA 76.88 S8573 1996. URL http:

//www.supercomp.org/sc96/
proceedings/SC96PROC/FOSTER2/

INDEX.HTM. ACM Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Flynn:1994:FCC

[FKL94] Kathleen M. Flynn, John L.

Kittle, and Alan M. Lumb. A Fortran coding convention for use in the U.S. Geological Survey, Water Resources Division. Open-file report 94-501, U.S. Geological Survey; USGS Earth Science Information Center Open-File Reports Section [distributor], Reston, VA, USA, 1994. iii + 64 pp.

Farshad:1991:HRF

[FL91] F. Farshad and J. L. LeBlanc.

How to run a FORTRAN or a BASIC computer program on PCs. Geobyte, 6(2):37–??, April 1991. ISSN 0885-6362.

Fox:1997:PFJ

[FLQZ97] Geoffrey Fox, Xiaoming Li, Zheng Qiang, and Wu Zhigang. A prototype of Fortranto-Java converter. Concurrency, practice and experience, 9(11):1047–1061, November 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
13832;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13832&PLACEBO=IE.pdf.

Special Issue: Java for computational science and engineering — simulation and modeling II.

Francis:1994:DPS

[FMW^{+}94] Rhys S. Francis, Ian D. Mathieson, Paul G. Whiting, Martin R. Dix, Harvey L. Davies, and Leon D. Rotstayn. A data parallel scientific modeling language. Journal of Parallel and Distributed Computing, 21(1):46–60, April 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1994.1041/production;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1994.1041/production/pdf.

Ford:1995:NNN

[For95] Brian Ford. The new NAG numerical PVM library (or A new parallel numerical library based on PVM). In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/
proceedings/ford/ford1.html.

Forsythe:1997:CCE

[For97] Chester Forsythe. Contemporary Computing for Engineers and Scientists Using Fortran 90. PWS-Kent Pub. Co., Boston, MA, USA, January 1, 1997. ISBN 0-534-93139-1. 544 pp. US$55.95.

Foster:1993:FML

[Fos93] I. Foster. Fortran M as a language for building earth system models. In Hoffmann and Kauranne [HK93a], pages 144–151. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Foster:1994:TPH

[Fos94] Ian Foster. Task parallelism and high-performance languages. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):27–36, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Foster:1995:DBP

[Fos95] Ian Foster. Designing and building parallel programs: concepts and tools for parallel software engineering. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-57594-9. xiii + 381 pp. LCCN QA76.642.F67 1995. US$.

URL http://www-unix.mcs.
anl.gov/dbpp/.

Foster:2017:QCF

[Fos17] Marcus P. Foster. Quantity correctness in Fortran programs. Computing in Science and Engineering, 19(4):83–87, July/August 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

URL https://www.computer.
org/csdl/mags/cs/2017/04/
mcs2017040083-abs.html.

Fox:1979:RFP

[Fox79] Phyllis Fox. Remark on “Algorithm 528: Framework for a portable library [Z]”. ACM Transactions on Mathematical Software, 5(4):524, December 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [FHS78].

Fox:1991:DHP

[Fox91a] G. Fox. Draft High Performance Fortran Language Specification: High Performance Fortran Forum, CRPC-TR92225, November 1992. Computer Information Technical, June 1991. ISBN 99922-46-36-7. US$50.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=9992246367.

Fox:1991:FDL

[Fox91b] Geoffrey Fox. Fortran D language specification. Technical report COMP TR90-14, Rice University, Dept. of Computer Science, Houston, TX, USA, 1991. 37 pp.

Fox:1994:APS

[Fox94] G. Fox. The application perspective for scalable data and task parallel languages HPF and HPC++. In Anonymous [Ano94d], pages 445–457. ISBN ???? LCCN ????

Festa:2001:AFS

[FPR01] Paola Festa, Panos M. Pardalos, and Mauricio G. C. Resende. Algorithm 815: FOR-TRAN subroutines for computing approximate solutions of feedback set problems using GRASP. ACM Transactions on Mathematical Software, 27(4):456–464, December 2001. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fernandez:1994:FPC

[FR94] Jose Fernandez and John B.

Rundle. FORTRAN program to compute displacement, potential, and gravity changes resulting from a magma intrusion in a multilayered Earth model. Computers and Geosciences, 20(4):461–??, May 1994. CO-DEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Freeman:1992:NNM

[Fre92] J. A. Freeman. Neural networks in Mathematica. AI Expert, 7(11):26–35, November 1992. CODEN AIEXEF. ISSN 0888-3785.

Fritzson:1994:CCI

[Fri94] Peter A. Fritzson, editor. Compiler construction: 5th International Conference, CC ’94, Edinburgh, UK, April 7–9, 1994: proceedings, volume 786 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-57877-3 (New York), 3-540-57877-3 (Berlin). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.C65 I52 1994.

Frisbie:1996:AED

[Fri96] Ralph Jay Frisbie. Across the Editor’s desk. Fortran Journal, 8(6):??, November/December 1996. ISSN 1060-0221. URL http:/

/www.fortran.com/fortran/
FJ/9611/#desk.

Fahringer:2002:SAS

[FSPC^{+}02] T. Fahringer, K. Sowa-Piekło, P. Czerwiński, P. Brezany, M. Bubak, R. Koppler, and R. Wismüller. SPiDER — An advanced symbolic debugger for Fortran 90/HPF programs. Concurrency and Computation: Practice and Experience, 14(2):103–136, February 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/91513538/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=91513538&PLACEBO=IE.pdf.

Filippone:1990:VLS

[FSV90] Salvatore Filippone, Paolo Santangelo, and Marcello Vitaletti. A vectorized longperiod shift-register random number generator. In IEEE [IEE90a], pages 676–684. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Fateman:2003:CCR

[FT03] Richard J. Fateman and Raymond Toy. Converting call-by-reference to call-by-value: Fortran and Lisp coexisting. In Senda [Sen03], pages 95–102. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Fausto:1991:NFP

[FTD91] R. Fausto and J. J. C. Teixeira-Dias. NBI: a FORTRAN program for molecular mechanics calculations on a microcomputer. Computers and Chemistry, 15(2):175–??, 1991. CO-DEN COCHDK. ISSN 0097-8485.

Forth:2004:JCG

[FTPR04] Shaun A. Forth, Mohamed Tadjouddine, John D. Pryce, and John K. Reid. Jacobian code generated by source transformation and vertex elimination can be as efficient as handcoding. ACM Transactions on Mathematical Software, 30(3): 266–299, September 2004. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fu:1990:EEF

[Fu90] Chuigang Fu. Evaluating the effectiveness of Fortran vectorizers by measuring total parallelism. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana, IL, USA, August 1990. ix + 56 pp. UILU-ENG-90-8029.

Fujino:1995:HOD

[Fuj95] Seiji Fujino. High-order difference schemes by modification of the right-hand side of 3D Poisson’s equation to parallel computations. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.nsc.liu.se/~boein/
                                        ifip/kyoto/workshop-info/
                                        proceedings/fujino/fujino1.
                                        html.

Furzer:1993:CTF

[Fur93] Ian Furzer. Computing teaching with Fortran 90. CEE. Chemical engineering education, 27(3):216–219, Summer 1993. CODEN CHEDAY. ISSN 0009-2479.

Feldman:1990:PFC

[FW90] S. I. Feldman and P. J. Weinberger. A portable Fortran 77 compiler. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990. URL http:

                                        //doc.cat-v.org/unix/v10/

10thEdMan/v2contents.html.

Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Faigin:1994:PIR

[FWH^{+}94] Keith A. Faigin, Stephen A.

Weatherford, Jay P. Hoeflinger, David A. Padua, and Paul M. Petersen. The Polaris internal representation. International Journal of Parallel Programming, 22(5):553– 586, October 1994. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Foster:1994:CSI

[FXAC94] I. Foster, Ming Xu, B. Avalani, and A. Choudhary. A compilation system that integrates High Performance Fortran and Fortran M. In IEEE [IEE94d], pages 293–300. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994.

Fernandez:1999:CCD

[FYR99] Jose Fernandez, Ting-To Yu, and John B. Rundle. Corrigenda: Corrigenda to “deformation produced by a rectangular dipping fault in a viscoelastic-gravitational layered earth model. part 1: Thrust fault, fltgrv and fltgrh FORTRAN programs” (computers and geosciences 22 (1996) 735–750). Computers and Geosciences, 25(3):301–??, ???? 1999. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Galison:1992:FPH

[Gal91]  Peter Galison.  Fortran,

physics, and human nature. In The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert [NRS92], chapter 10, pages 225–260. ISBN 94-010-5097-X, 94-011-2488-4 (ebook). ISSN 0068-0346. LCCN

????  URL https://link.
springer.com/chapter/10.
1007/978-94-011-2488-1_10.

Gao:2005:ERS

[Gao05] Xiao-Wei Gao. Evaluation of regular and singular domain integrals with boundary-only discretization-theory and Fortran code. Journal of Computational and Applied Mathematics, 175(2):265–290, March 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042704002328.

Gao:2006:NET

[Gao06]  Xiao-Wei Gao.  Numerical

evaluation of two-dimensional singular boundary integrals — theory and Fortran code. Journal of Computational and Applied Mathematics, 188 (1):44–64, April 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705001597.

Garc91a

[Gar91a] Oscar García. A system for the differentiation of Fortran code and an application to parameter estimation in forest growth models. In Andreas Griewank and George F. Corliss, editors, Automatic Differentiation of Algorithms: Theory, Implementation, and Application. SIAM, Philadelphia, PA, 1991.

Garcia:1991:SDF

[Gar91b] Oscar García. A system for the differentiation of Fortran code and an application to parameter estimation in forest growth models. In Andreas Griewank and George F. Corliss, editors, Automatic Differentiation of Algorithms: Theory, Implementation, and Application, page ?? SIAM Press, Philadelphia, PA, USA, 1991.

Garcia:1993:NRF

[Gar93] Alejandro Garcia. Numerical Recipes in FORTRAN. By William H. Press, Saul A Teukolsky, William T. Vetterling, and Brian P. Flannery. Journal of the American Chemical Society, 115(20): 9356–??, October 1993. CO-DEN JACSAT. ISSN 0002-7863.

Gutbrod:1996:SGT

[GAW96a] F. Gutbrod, N. Attig, and M. Weber. The SU(2)-Lattice Gauge Theory simulation code on the Intel Paragon supercomputer. Parallel Computing, 22 (3):443–463, March 1996. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gutbrod:1996:SLG

[GAW96b] F. Gutbrod, N. Attig, and M. Weber. The SU(2)-Lattice Gauge Theory simulation code on the Intel Paragon supercomputer. Parallel Computing, 22(3):443–463, April 29, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/parco/cas_sub/
                                        browse/browse.cgi?year=1996&
                                        volume=22&issue=3&aid=1057.

Gupta:1992:DAD

[GB92] Manish Gupta and Prithviraj Banerjee. Demonstration of automatic data partitioning techniques for parallelizing compilers on multicomputers. IEEE Transactions on Parallel and Distributed Systems, 3(2):179– 193, March 1992. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gerndt:1995:PAS

[GB95] M. Gerndt and R. Berrendorf. Parallelizing applications with SVM-Fortran. Lecture Notes in Computer Science, 919:793– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerbaud:1992:MAD

[GBC92] L. Gerbaud, J. Bigeon, and G. Champenois. Modular approach to describe electromechanical systems-using Macsyma to generate global approach simulation software. In IEEE [IEE92b], pages 1189– 1196 (vol. 2). ISBN 0-7803-0696-1 (casebound), 0-7803-0695-3 (softbound), 0-7803-0697-X (microfiche). LCCN TK 7835 I16 1992. Two volumes. IEEE catalog number 92CH3163-3.

Gladwell:1997:FSH

[GBDB97] I. Gladwell, K. Bouas-Dockery, and R. W. Brankin. A Fortran 90 separable Hamiltonian system solver. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3):207– 217, October 17, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/apnum/cas_sub/
browse/browse.cgi?year=1997&
volume=25&issue=2-3&aid=
815.

Garain:2015:CCF

[GBR15] Sudip Garain, Dinshaw S. Balsara, and John Reid. Comparing Coarray Fortran (CAF) with MPI for several structured mesh PDE applications. Journal of Computational Physics, 297(??):237–253, September 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S002199911500354X.

Genz:2003:ANC

[GC03] Alan Genz and Ronald Cools. An adaptive numerical cubature algorithm for simplices. ACM Transactions on Mathematical Software, 29(3):297– 308, September 2003. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Grotendorst:1994:SCT

[GDS94] Johannes Grotendorst, Juergen Dornseiffer, and Siegfried M. Schoberth. Symbolic-numeric computation techniques for problem-solving in physical chemistry and biochemistry. Technical Report KFA-ZAM-IB-9414, KFA Research Centre, Jülich, Jülich, Germany, 1994. 9 pp.

URL ftp://ftp.zam.kfa-
juelich.de/pub/zamdoc/ib/
ib-94/ib-9414.ps.

Gehrke:1995:FLG

[Geh95] Wilhelm Gehrke. Fortran 90 Language Guide. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., May 1, 1995. ISBN 3-540-19926-8 (Berlin), 0-387-19926-8 (New York). 500 pp.

LCCN QA 76.73 F25 G43 1995. US$49.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=3540199268.

Gehrke:1996:FLG

[Geh96] Wilhelm Gehrke. Fortran 95 Language Guide. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., October 1, 1996. ISBN 0-387-76062-8, 3-540-76062-8. 420 pp. LCCN QA76.73.F25 G433 1996. US$44.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=3540760628.

Gehrke:1997:FLG

[Geh97] Wilhelm Gehrke. The F Language Guide. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., June 1997. ISBN 3-540-76165-9. US$44.95.

Gentleman:2006:BRD

[Gen06]  Robert Gentleman.  Book

review: Developing Statistical Software in FORTRAN 95. Journal of Statistical Software, 17(BR-2):1–2, December 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v17/b02.

Gephart:1990:FFP

[Gep90] John W. Gephart. FMSI: a FORTRAN program for inverting fault/slickenside and earthquake focal mechanism data to obtain the regional stress tensor. Computers and Geosciences, 16(7):953–??, 1990. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Gerndt:1994:APC

[Ger94a] Michael Gerndt. Automatic parallelization of a crystal growth simulation program for distributed-memory systems. Technical Report KFA-ZAM-IB-9404, KFA Research Centre, Jülich, Jülich, Germany, 1994. 7 pp.

URL ftp://ftp.zam.kfa-
juelich.de/pub/zamdoc/ib/
ib-94/ib-9404.ps.

Gerndt:1994:PAE

[Ger94b] Michael Gerndt. Performance analysis environment for SVMfortran programs. Technical Report KFA-ZAM-IB-9417, KFA Research Centre, Jülich, Jülich, Germany, 1994. 41

pp. URL ftp://ftp.zam.kfa-
juelich.de/pub/zamdoc/ib/
ib-94/ib-9417.ps.

Gerndt:1998:HLP

[Ger98a] Michael Gerndt. High-level programming of massively parallel computers based on shared virtual memory. Parallel Computing, 24(3–4):383– 400, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.elsevier.
com/cas/tree/store/parco/
sub/1998/24/3-4/1295.pdf.

Gerndt:1998:HPM

[Ger98b] Michael Gerndt. High-level programming of massively parallel computers based on shared virtual memory. Parallel Computing, 24(3-4):383– 400, May 1998. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ghaleb:1995:CFP

[GF95a] Amin Ramadan Ghaleb and Norman Fry. CSTRAIN: A FORTRAN 77 program to study fry’s plots in twodimensional simulated models. Computers and Geosciences, 21(7):825–??, 1995. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Ghuloum:1995:FPI

[GF95b] Anwar M. Ghuloum and Allan L. Fisher. Flattening and parallelizing irregular, recurrent loop nests. ACM SIG-PLAN Notices, 30(8):58–67, August 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gay:1995:DRN

[GG95] David Gay and Eric Grosse.

d1mach revisited: no more uncommenting DATA statements. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/
proceedings/einarsson/d1mach.

html. Full paper published in [GG99].

Gay:1999:SAF

[GG99] David M. Gay and Eric Grosse. Self-adapting Fortran 77 machine constants: Comment on Algorithm 528. ACM Transactions on Mathematical Software, 25(1):123– 126, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://cm.bell-labs.
com/who/ehg/mach/d1mach.
ps;   http://doi.acm.org/
10.1145/305658.305711;
http://www.acm.org/pubs/
citations/journals/toms/
cgi-bin/TOMSbibget?Gay:1999:
SAF;   http://www.acm.org/
pubs/citations/journals/
toms/cgi-bin/TOMScitation?
Fox:1978:AFP;   http://www.
acm.org:80/pubs/citations/
journals/toms/1999-25-1/

p123-gay/. See [FHS78].

Gunnels:2001:FFL

[GGHvdG01] John A. Gunnels, Fred G. Gustavson, Greg M. Henry, and Robert A. van de Geijn. FLAME: Formal Linear Algebra Methods Environment. ACM Transactions on Mathematical Software, 27(4):422– 455, December 2001. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gawman:1993:PC

[GGK^{+}93] Ann Gawman, W. Morven Gentleman, E. Kidd, Per-Åke Larson, and J. Slonim, editors. Proceedings CASCON ’93: Toronto, Ontario, Canada, 24– 28 October 1993. Nat. Res. Council of Canada, Ottawa, Ont., Canada, 1993. ISBN ???? LCCN QA76.76.S64 C378 1993 v.1-2. Two volumes.

Garbow:1988:AFS

[GGLM88] B. S. Garbow, G. Giunta, J. N.

Lyness, and A. Murli. Algorithm 662: A FORTRAN software package for the numerical inversion of the Laplace transform based on Weeks’ method. ACM Transactions on Mathematical Software, 14(2): 171–176, June 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0098-
3500/214375.html. See also

[GL90].

Ganesan:1996:CSM

[GGW96] Ravikanth Ganesan, Kannan Govindarajan, and Min-You Wu. Comparing SIMD and MIMD programming modes. Journal of Parallel and Distributed Computing, 35(1):91– 96, May 25, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1996.0071/production;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0071/production/pdf.

Gentzsch:1994:HCNa

[GH94a] Wolfgang Gentzsch and Uwe Harms, editors. High-Perfor-mance computing and networking: International Conference and Exhibition, Munich, Germany, April 18–20, 1994: proceedings, volume 796 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Gentzsch:1994:HCNb

[GH94b] Wolfgang Gentzsch and Uwe Harms, editors. High-performance computing and networking: International Conference and Exhibition, Munich, Germany, April 18–20, 1994: proceedings, volume 797 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57981-8 (Berlin: vol. 2: paperback), 0-387-57981-8 (New York: vol. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1– 2 (c1994). DM96.00. Two volumes.

Gentzsch:1994:HPC

[GH94c] Wolfgang Gentzsch and Uwe Harms, editors. High-performance computing and networking: international conference and exhibition, Munich, Germany, April 18–20, 1994: proceedings, volume 797 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Gagunashvili:2018:CCG

[GH18] Nikolay D. Gagunashvili and Helgi Halldorsson. CHIWEI: A code of goodness of fit tests for weighted and unweighted histograms in Fortran-77, C++, R and Python. Computer Physics Communications, 231(??):245, October 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465518301449.

Gagunashvili:2019:CCC

[GHN19] Nikolay D. Gagunashvili, Helgi Halldorsson, and Helmut Neukirchen. CHICOM: Code for comparing weighted or unweighted histograms in Fortran-77, C++, R and Python. Computer Physics Communications, 245(??):Article 106872, December 2019. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465519302590.

Ghosh:2001:RCF

[Gho01] A. Ghosh. Reply to comment on “A FORTRAN program for fitting Weibull distribution and generating samples” by H. Tsai. Computers and Geosciences, 27(2): 261–262, March 2001. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Gupta:1994:IFF

[GHSJ94] S. K. S. Gupta, C.-H. Huang, P. Sadayappan, and R. W. Johnson. Implementing fast Fourier transforms on distributedmemory multiprocessors using data redistributions. Parallel Processing Letters, 4(4):477– 488, December 1994. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Gillett:1991:FPSb

[Gil91a] R. Gillett. A FORTRAN 77 program for sample-size determination in replication attempts when effect size is uncertain. Behavior research methods, instruments, and computers, 23(4):545, November 1991. CODEN BRMCEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Gillett:1991:FPSa

[Gil91b] Raphael Gillett. A FOR-TRAN 77 program for samplesize determination in replication attempts when effect size is uncertain. Behavior research methods, instruments, and computers, 23(3):442–446, August 1991. CODEN BRM-CEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Gillett:1994:END

[Gil94]  Raphael Gillett.  The ex-

act null distribution for radial maze statistics: A FOR-TRAN 77 program. Behavior research methods, instruments, and computers, 26(1): 70–72, February 1994. CODEN BRMCEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Gillett:2001:SSD

[Gil01] R. Gillett. Sample size determination for a t test given a t value from a previous study: a FORTRAN 77 program. Behavior research methods, instruments, and computers, 33 (4):544–548, November 1, 2001. CODEN BRMCEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Griewank:1996:AAP

[GJU96] Andreas Griewank, David Juedes, and Jean Utke. Algorithm 755: ADOL-C: a package for the automatic differentiation of algorithms written in C/C++. ACM Transactions on Mathematical Software, 22 (2):131–167, June 1996. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Giering:2006:TLA

[GK06] Ralf Giering and Thomas Kaminski. Tangent linear and adjoint versions of NASA/ GMAO’s Fortran 90 global weather forecast model. In Bücker et al. [BCH^{+}06], pages 275–284. CODEN LNCSA6. ISBN 3-540-28403-6 (print), 3-540-28438-9 (e-book). ISSN 1439-7358. LCCN QA304 .I58

2006. URL http://link.
springer.com/content/pdf/
10.1007/3-540-28438-9_24.

Proceedings of the Fourth International Conference on Automatic Differentiation, July 20–23, Chicago, Illinois.

Gupta:1992:MGD

[GKH^{+}92] S. K. S. Gupta, S. D. Kaushik, C.-H. Huang, J. R. Johnson, R. W. Johnson, and P. Sadayappan. A methodology for generating data distributions to optimize communication. In IEEE [IEE92c], pages 436–441. ISBN 0-8186-3200-3. LCCN QA 76.58 I42 1992. IEEE catalog no. 92TH0492-9.

Gupta:1993:AGD

[GKH^{+}93] S. K. S. Gupta, S. D. Kaushik, C.-H. Huang, J. R. Johnson, R. W. Johnson, and P. Sadayappan. On the automatic generation of data distributions. In Schnabel [Sch93b], page 82. CODEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7 .S54 v.28:1.

Gevorkyan:2019:SSC

[GKKL19] Migran N. Gevorkyan, Anna V.

Korolkova, Dmitry S. Kulyabov, and Konstantin P. Lovetskiy. Statistically significant comparative performance testing of Julia and Fortran languages in case of Runge -Kutta methods. In Numerical Methods and Applications, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2019.

URL http://link.springer.
com/chapter/10.1007/978-
3-030-10692-8_45.

Garbow:1990:RFS

[GL90] B. S. Garbow and J. N. Lyness.

Remark on “Algorithm 662: A FORTRAN software package for the numerical inversion of the Laplace transform based on Weeks’ method”. ACM Transactions on Mathematical Software, 16(4):405–406, December 1990. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0098-3500/

98302.html. See [GGLM88].

Gastineau:2010:TCA

[GL10] Mickaël Gastineau and Jacques Laskar. TRIP: a computer algebra system dedicated to celestial mechanics and perturbation series. ACM Communications in Computer Algebra, 44(4):194–197, December 2010. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Gladwell:1992:BRF

[Gla92a] Ian Gladwell. Book review: Fortran 90 Explained by M. Metcalf and J. Reid: (Oxford Science). ACM SIGPLAN Notices, 27(4):11, April 1992. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Glassy:1992:TNS

[Gla92b]  Lou Glassy.  Tiny-Ninety:

a subset of F90 for beginning programmers. Fortran Journal, 4(3):2–6, May/ June 1992. ISSN 1060-

                                        0221. URL http://www.cs.
                                        montana.edu/~glassy/papers/
                                        html/t90-article.html.

Gliss:1996:AHC

[Gli96] B. Gliss. An Ada 95 harness for converting legacy Fortran applications. Lecture Notes in Computer Science, 1088:413– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

GED:1991:F

[Glo91a] Global Engineering Documents. Fortran 90. Global Engineering Documents, Washington, DC, USA, 1991. ISBN ???? xix + 369 pp. LCCN ????

Glowinski:1991:PIC

[Glo91b] R. Glowinski, editor. Proceedings of the 10th International Conference on Computing Methods in Applied Sciences and Engineering. INRIA.

Nova Science, New York, NY, USA, 1991.

Germain:1997:HCS

[GLPE97] C. Germain, J. Laminie, M. Pallud, and D. Etiemble. An HPF case study of a domain-decomposition based irregular application. Lecture Notes in Computer Science, 1277:201–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graham:1993:OIA

[GLS93] Susan L. Graham, Steven Lucco, and Oliver Sharp. Orchestrating interactions among parallel computations. ACM SIGPLAN Notices, 28(6):100– 111, June 1993. CODEN SIN-ODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/pldi/155090/
p100-graham/.

Greenberg:1997:ACS

[GM97] Leon Greenberg and Marco Marletta. Algorithm 775. the code SLEUTH for solving fourth-order Sturm–Liouville problems. ACM Transactions on Mathematical Software, 23 (4):453–493, December 1997. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Garcia:1996:BRCb

[GMC96a] Alejandro Garcia, Susan McKay, and Wolfgang Christian. Book review: C. Redwine, Upgrading to Fortran 90. Computers in physics, 10(2):135–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.4822374.

Garcia:1996:BRCa

[GMC96b] Alejandro Garcia, Susan McKay, and Wolfgang Christian. Book review: I. Chivers and J. Sleightholme, Introducing Fortran 90. Computers in physics, 10(2):135–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.4822371.

Garcia:1996:BRJ

[GMC96c] Alejandro Garcia, Susan McKay, and Wolfgang Christian. Book review: J. Kerrigan, Migrating to Fortran 90. Computers in physics, 10(2):135–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.4822373.

Garcia:1996:BRM

[GMC96d] Alejandro Garcia, Susan McKay, and Wolfgang Christian. Book review: M. Metcalf and J. Reid, Fortran 90 Explained. Computers in physics, 10(2):

135–??, March 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.4822372.

Garcia:1996:BRWb

[GMC96e] Alejandro Garcia, Susan McKay, and Wolfgang Christian. Book review: W. Gehrke, Fortran 90 Language Guide. Computers in physics, 10(2):135–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.4822375.

Garcia:1996:BRWa

[GMC96f] Alejandro Garcia, Susan McKay, and Wolfgang Christian. Book review: W. S. Brainerd and C. H. Goldberg and J. C. Adams, Programmer’s Guide to Fortran 90. Computers in physics, 10(2):135–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.4822370.

Gerber:2018:DRP

[GMF18] F. Gerber, K. Mösinger, and R. Furrer. dotCall64: an R package providing an efficient interface to compiled C, C++, and Fortran code supporting long vectors. SoftwareX, 7(??):211–216, January/June 2018. CODEN ???? ISSN 2352-7110. URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S2352711018300785.

Guocheng:1992:FPT

[GMHC92] Pan Guocheng, Ken Moss, Tim Heiner, and James R. Carr. A FORTRAN program for three-dimensional cokriging with case demonstration. Computers and Geosciences, 18(5): 557–578, June 1992. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Gong:2016:NPG

[GML^{+}16] Jing Gong, Stefano Markidis, Erwin Laure, Matthew Otten, Paul Fischer, and Misun Min. Nekbone performance on GPUs with OpenACC and CUDA Fortran implementations. The Journal of supercomputing, 72(11):4160–4180, November 2016. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gottlieb:1992:HSF

[GMMM92] R. A. Gottlieb, D. J. Magenheimer, S. A. Meloy, and A. C. Meyer. HP 9000 series 700 FORTRAN optimizing preprocessor. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 43(3): 24–32, June 1992. CODEN HPJOAX. ISSN 0018-1153.

Gupta:1995:HCI

[GMS^{+}95] Manish Gupta, Sam Midkiff, Edith Schonberg, Ven Seshadri, David Shields, Ko-Yang Wang, Wai-Mee Ching, and Ton Ngo. An HPF compiler for the IBM SP2. In Karin [Kar95], page ?? ISBN 0-89791-816-9. LCCN QA 76.88 S8573 1995a.

URL http://www.supercomp.
org/sc95/proceedings/417_
SAMM/SC95.HTM. These pro-

ceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Grubel:1994:ATN

[GOBG^{+}94] G. Grübel, M. Otter, I. Bausch-Gall, H. Fischer, F. Breitenecker, A. Prinz, and G. Schuster. An ACSL-model translator to the neutral Fortran DSblock-model format. In cacsd94 [IEE94g], pages 143– 148. ISBN ????

Goda:1993:HPF

[God93] O. Goda. High performance Fortran. Joho-Shori (J. Information Processing Soc. Japan), 34(9):1179–1186, September 1993. CODEN JOSHA4. ISSN 0447-8053.

Gome90a

[Gom90a] C. Gomez. MACROFORT: a FORTRAN code generator for MAPLE. Rapports Techniques 119, INRIA, 1990.

Gomez:1990:MFC

[Gom90b]  Claude Gomez.  MACRO-

FORT: a FORTRAN code generator in MAPLE = MACRO-FORT: un générateur de code FORTRAN dans MAPLE. Rapports techniques 119, IN-RIA (Institut National de Recherche en Informatique et en Automatique), Rocquencourt, France, May 1990. 14 pp.

Goncalves:2001:CSP

[Gon01] M. A. Goncalves. Computing for scientists: principles of programming with C++ and Fortran 90 — R.J. Barlow and A.R. Barnett; Wiley, Chichester, 1998, 292pp., price US$49.95, ISBN 0-471-95596-5. Computers and Geosciences, 27(5):609–610, June 2001. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Goodman:1990:FCC

[Goo90a] Fred Goodman. FORTRAN to C: Character manipulation. The Journal of C Language Translation, 2(2):141– 151, September 1990. ISSN 1042-5721.

Goodman:JCLT-2-2-141

[Goo90b] Fred Goodman. FORTRAN to C: Character manipulation. The Journal of C Language Translation, 2(2):141– 151, September 1990. ISSN 1042-5721.

Goodman:1990:FCN

[Goo90c] Fred Goodman. FORTRAN to C: Numerical issues. The Journal of C Language Translation, 2(1):29–44, June 1990. ISSN 1042-5721.

Goodman:JCLT-2-1-29

[Goo90d] Fred Goodman. FORTRAN to C: Numerical issues. The Journal of C Language Translation, 2(1):29–44, June 1990. ISSN 1042-5721.

Goodman:1990:FCT

[Goo90e]  Fred Goodman.  A FOR-

TRAN to C translator, part II. The Journal of C Language Translation, 1(4):264– 274, March 1990. ISSN 1042-5721.

Goodman:JCLT-1-4-264

[Goo90f]  Fred Goodman.  A FOR-

TRAN to C translator, Part II. The Journal of C Language Translation, 1(4):264– 274, March 1990. ISSN 1042-5721.

Gross:1994:TPH

[GOS94] Thomas Gross, David R.

O’Hallaron, and Jaspal Subhlok. Task parallelism in a High Performance Fortran framework. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):16–26, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Gould:2003:CSC

[GOT03a] Nicholas I. M. Gould, Dominique Orban, and Philippe L. Toint. CUTEr and SifDec: a constrained and unconstrained testing environment, revisited. ACM Transactions on Mathematical Software, 29(4):373– 394, December 2003. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gould:2003:GLT

[GOT03b] Nicholas I. M. Gould, Dominique Orban, and Philippe L. Toint. GALAHAD, a library of thread-safe Fortran 90 packages for large-scale nonlinear optimization. ACM Transactions on Mathematical Software, 29(4):353–372, December 2003. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gouveia:1993:ATC

[Gou93] A. D. Gouveia. An aid to twodimensional contouring using nonuniform orthogonal grids — a FORTRAN algorithm. Computers and Geosciences, 19(8): 1071–??, September 1993. CO-DEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Girkar:1992:AEF

[GP92] Milind Girkar and Constantine D. Polychronopoulos. Automatic extraction of functional parallelism from ordinary programs. IEEE Transactions on Parallel and Distributed Systems, 3(2):166–178, March 1992. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Geurts:1994:FSC

[GP94] A. J. Geurts and C. Praagman. A Fortran subroutine for column reduction of polynomial matrices. Report EUT Report-WSK 94-01, Eindhoven University of Technology, Department of Mathematics and Computing Science, Eindhoven, The Netherlands, 1994. ii + 70 pp.

Geurts:1997:AFP

[GP97] A. J. Geurts and C. Praagman.

Algorithm 767: a Fortran 77 package for column reduction of polynomial matrices. ACM Transactions on Mathematical Software, 23(1):111–129, March 1997. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1997-23-1/p111-geurts/.

Guzzi:1990:CFO

[GPHL90] Mark D. Guzzi, David A.

Padua, Jay Hoeflinger, and Duncan H. Lawrie. Cedar Fortran and other vector and parallel Fortran dialects. The Journal of supercomputing, 4 (1):37–??, 1990. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Gockenbach:1999:CCL

[GPS99] Mark S. Gockenbach, Matthew J. Petro, and William W. Symes. C++ classes for linking optimization with complex simulations. ACM Transactions on Mathematical Software, 25(2): 191–212, June 1999. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gibson:1992:DIS

[GR92] D. H. Gibson and G. S. Rao. Design of the IBM System/390 computer family for numerically intensive applications: An overview for engineers and scientists. IBM Journal of Research and Development, 36 (4):695–711, July 1992. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Grego:1993:PFP

[Gre93] John M. Grego. PRASCH: A Fortran program for latent class polytomous response Rasch models. Applied psychological measurement, 17(3): 238–??, September 1993. ISSN 0146-6216.

Gray:1999:SPS

[GRE99] Mark G. Gray, Randy M. Roberts, and Tom M. Evans. Scientific programming: Shadowobject interface between Fortran 95 and C++. Computing in Science and Engineering, 1(2):63–70, March/ April 1999. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://dlib.computer.org/
cs/books/cs1999/pdf/c2063.
pdf.

Griesmer:1993:BIF

[Gri93]  James H. Griesmer.  A

brief introduction to FOR-MAC. ACM SIGPLAN Notices, 28(3):355–356, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/154766/p355-griesmer/

.

Grotendorst:1990:AFM

[Gro90] J. Grotendorst. Approximating functions by means of symbolic computation and a general extrapolation method. Computer Physics Communications, 59 (2):289–301, June 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Grose:1991:PFO

[Gro91] Timothy James Grose. The programming and functionality of OPS5 compared to LISP and FORTRAN in an aeronautical route planning system. Thesis (M.A.), University of Texas at Austin, Austin, TX, USA, 1991. vii + 63 pp.

Gockenbach:2002:EAI

[GRSS02] Mark S. Gockenbach, Daniel R.

Reynolds, Peng Shen, and William W. Symes. Efficient and automatic implementation of the adjoint state method. ACM Transactions on Mathematical Software, 28(1):22–44, March 2002. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gustavson:2007:AFS

[GRW07] Fred G. Gustavson, John K. Reid, and Jerzy Waśniewski. Algorithm 865: Fortran 95 subroutines for Cholesky factorization in block hybrid format. ACM Transactions on Mathematical Software, 33(1): 8:1–8:5, March 2007. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Garg:1990:FEAa

[GS90a] Satya Prakash Garg and Subhash Chandra Singhal. FOR-TRAN and engineering applications. Oxford and IBH Publishing Co., New Delhi, India, fourth edition, 1990. ISBN ???? xiv + 500 pp. LCCN ????

Garg:1990:FEAb

[GS90b] Satya Prakash Garg and Subhash Chandra Singhal. Fortran and Engineering Applications. South Asia Books, ????, fourth edition, December 1990. ISBN 81-204-0013-5. US$27.50. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=81-204-0013-5.

Gustafsson:1995:PSH

[GS95] N. Gustafsson and D. Salmond.

A parallel spectral HIRLAM with the data parallel programming model and with message passing — a comparison (weather forecasting). In Hoffmann and Kreitz [HK95], pages 32–48. ISBN 981-02-2211-4. LCCN QC866.E26 1994.

Gupta:1997:SAR

[GS97] Manish Gupta and Edith Schonberg. Static analysis to reduce synchronization costs data-parallel programs with remote memory copy. Parallel Processing Letters, 7(2): 145–156, June 1997. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Gomez:1998:MPG

[GS98] Claude Gomez and Tony Scott. Maple programs for generating efficient FORTRAN code for serial and vectorised machines. Computer Physics Communications, 115 (2–3):548–562, December 2, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465598001143.

Garg:2001:TOA

[GS01a] Rajat P. Garg and Ilya Sharapov. Techniques for Optimizing Applications: High Performance Computing. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-093476-3. xliv + 616 pp. LCCN QA76.88 .G37

2002. URL books/apt.pdf;
http://www.sun.com/books/
catalog/garg.html/index.

html. Part No. 806-6380-10 June 2001, Revision 01.

Grelck:2001:HVS

[GS01b] Clemens Grelck and Sven-Bodo Scholz. HPF vs. SAC – a case study (research note). Lecture Notes in Computer Science, 1900:620–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1900/19000620.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1900/19000620.
pdf.

Gil:2002:AAB

[GST02a] Amparo Gil, Javier Segura, and Nico M. Temme. Algorithm 819: AIZ, BIZ: two Fortran 77 routines for the computation of complex Airy functions. ACM Transactions on Mathematical Software, 28(3): 325–336, September 2002. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2002:AGH

[GST02b] Amparo Gil, Javier Segura, and Nico M. Temme. Algorithm 822: GIZ, HIZ: two Fortran 77 routines for the computation of complex Scorer functions. ACM Transactions on Mathematical Software, 28(4): 436–447, December 2002. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2004:AMB

[GST04a] Amparo Gil, Javier Segura, and Nico M. Temme. Algorithm 831: Modified Bessel functions of imaginary order and positive argument. ACM Transactions on Mathematical Software, 30(2):159–164, June 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2004:CSM

[GST04b] Amparo Gil, Javier Segura, and Nico M. Temme. Computing solutions of the modified Bessel differential equation for imaginary orders and positive arguments. ACM Transactions on Mathematical Software, 30 (2):145–158, June 2004. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2006:ARP

[GST06a] Amparo Gil, Javier Segura, and Nico M. Temme. Algorithm 850: Real parabolic cylinder functions U(a, x), V (a, x). ACM Transactions on Mathematical Software, 32 (1):102–112, March 2006. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2006:CRP

[GST06b] Amparo Gil, Javier Segura, and Nico M. Temme. Computing the real parabolic cylinder functions U(a, x), V (a, x). ACM Transactions on Mathematical Software, 32(1):70–101, March 2006. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2011:APC

[GST11] Amparo Gil, Javier Segura, and Nico M. Temme. Algorithm 914: Parabolic cylinder function W (a, x) and its derivative. ACM Transactions on Mathematical Software, 38(1):6:1–6:5, November 2011. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2012:IAF

[GST12] Amparo Gil, Javier Segura, and Nico M. Temme. An improved algorithm and a Fortran 90 module for computing the conical function p_{1/2+iτ}^{m} (x). Computer Physics Communications, 183(3):794–799, March 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0010465511003936.

Gondzio:1992:DAI

[GT92a] J. Gondzio and D. Tachat.

The design and application of the IMPLO — A FOR-TRAN library for linear optimization with interior point methods. Cahier 108, Laboratoire de Analyse et Modelisation de Systemes pour l’Aide a la Decision (LAMSADE), Universite de Paris Dauphine, F– 75775 Paris Cedex 16, France, January 1992.

int:Gondzio3

[GT92b] J. Gondzio and D. Tachat.

The design and application of the IMPLO – A FOR-TRAN library for linear optimization with interior point methods. Cahier 108, Laboratoire de Analyse et Modelisation de Systemes pour l’Aide a la Decision (LAMSADE), Universite de Paris Dauphine, F-75775 Paris Cedex 16, France, January 1992.

Gondzio:1994:DAI

[GT94] J. Gondzio and D. Tachat.

The design and application of IPMLO — a FORTRAN library for linear optimization with interior point methods. RAIRO: Recherche operationnelle. RAIRO: Operations research, 28(1):37–56, March 1994. CODEN RSROD3. ISSN 0399-0559.

Gould:2003:FFF

[GT03] N. I. M. Gould and Ph. L.

Toint. FILTRANE, a Fortran 95 filter-trust-region package for solving nonlinear feasibility problems. Technical Report RAL-TR-2003-017, Rutherford Appleton Laboratory, Chilton, England, 2003.

Gould:2007:FFF

[GT07] Nicholas I. M. Gould and Philippe L. Toint. FILTRANE, a Fortran 95 filter-trust-region package for solving nonlinear least-squares and nonlinear feasibility problems. ACM Transactions on Mathematical Software, 33(1):3:1–3:23, March 2007. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gorelik:1990:FSZ

[GU90] A. M. (Alla Moiseevna) Gorelik and V. L. (Viktoriia Lvovna) Ushkova. Fortran segodnia i zavtra. Algoritmy i algoritmicheskie iazyki Algoritmy i algoritmicheskie iazyki. Nauka, Moscow, Russia, 1990. ISBN 5-02-006799-7. 206 pp.

Guo:2001:DSH

[Guo01] M. Guo. Denotational semantics of an HPF-like dataparallel language model. Parallel Processing Letters, 11(2– 3):363–??, 2001. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Ganzha:1992:RBC

[GV92] V. G. Ganzha and E. V. Vorozhtsov. Resultant based code generation for the stability analysis of difference schemes with the aid of the RE-DUCE system. In Houstis and Rice [HR92], pages 185–192. ISBN 0-444-89703-8. LCCN Q334 .I45 1991.

Gustavson:2008:RFP

[GWDL08] Fred G. Gustavson, Jerzy Wasniewski, Jack J. Dongarra, and Julien Langou. Rectangular full packed format for Cholesky’s algorithm: Factorization, solution and inversion. LAPACK Working Note 199, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April 2008. URL http:

//www.netlib.org/lapack/
lawnspdf/lawn199.pdf.

Gustavson:2010:RFP

[GWDL10] Fred G. Gustavson, Jerzy Waśniewski, Jack J. Dongarra, and Julien Langou. Rectangular full packed format for Cholesky’s algorithm: factorization, solution, and inversion. ACM Transactions on Mathematical Software, 37(2):18:1– 18:21, April 2010. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gunteroth:2005:LEP

[GWE^{+}05] Kurt Gunteroth, Greg Wilson, Michael Erdmann, Dmitriy Vasilev, Ed Ravin, and Vivek Rao. Letters: Extensible programming, extended edition; A lesson in self-healing; Fortran lives. ACM Queue: Tomorrow’s Computing Today, 3 (2):12–14, March 2005. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Gardiner:1992:AFS

[GWL^{+}92] Judith D. Gardiner, Matthew R. Wette, Alan J. Laub, James J. Amato, and Cleve B. Moler. Algorithm 705: A FORTRAN-77 software package for solving the Sylvester matrix equation AXB^{T} + CXD^{T} = E. ACM Transactions on Mathematical Software, 18(2):232– 238, June 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0098-
3500/146930.html.

Hahn:1994:FSE

[Hah94] Brian D. Hahn. Fortran 90 for Scientists and Engineers. Edward Arnold, London, UK, November 1994. ISBN 0-340-60034-9 (paperback). x + 351 pp. LCCN QA76.73.F25 .H34 1994. US$35.50; UK£16.95.

URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0340600349.

Hall:1991:ICF

[Hal91] Mary Wolcott Hall. Interprocedural compilation of Fortran D for MIMD distributedmemory machines. Technical report COMP TR91-169, Rice University, Dept. of Computer Science, Houston, TX, USA, November 1991. 13 pp.

Hamilton:1985:RRK

[Ham85] Dennis E. Hamilton. Remark on “Algorithm 620: References and keywords for Collected Algorithms of the ACM”. ACM Transactions on Mathematical Software, 11(3):305– 307, September 1985. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [RH84, HM90a].

Hamilton:1993:DMA

[Ham93] Kenneth G. Hamilton. Direct memory access from PC Fortrans. Dr. Dobb’s Journal of Software Tools, 18(5):70, 72– 74, 102–104, May 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Hamilton:1995:UFP

[Ham95a] Kenneth G. Hamilton. Undocumented features of PC Fortran libraries. Dr. Dobb’s Journal of Software Tools, 20(1):30, 32, 34, 84–85, January 1995. CODEN DDJOEB. ISSN 1044-789X.

Haridi:1995:EPP

[HAM95b] Seif Haridi, Khayri Ali, and Peter Magnusson, editors. EURO-PAR ’95 parallel processing: First International EURO PAR Conference, Stockholm, Sweden, August 29–31, 1995: proceedings, number 966 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-60247-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I553 1995.

Hamilton:1998:AEP

[Ham98] K. G. Hamilton. Algorithm 780: Exponential pseudorandom distribution. ACM Transactions on Mathematical Software, 24(1):102–106, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hanson:1992:MMF

[Han92] R. J. Hanson. Matrix multiplication in Fortran 90 using Strassen’s algorithm. Fortran Journal, 4(3):6–7, May/ June 1992. ISSN 1060-0221.

Hansen:1998:EHP

[Han98] Per Brinch Hansen. An evaluation of High Performance Fortran. ACM SIGPLAN Notices, 33(3):57–64, March 1998. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). The author strongly criticizes HPF, and remarks “HPF has severe conceptual and practical limitations.”.

Harrison:19xx:IAA

[Harxx] W. L. Harrison. The interprocedural analysis and automatic parallelization of Scheme programs. CSRD Report 860, Center of Supercomputing Research and Development, University of Illinois, Urbana, IL, USA, 19xx.

Hasselman:2006:RAF

[Has06] Berend Hasselman. Remark on Algorithm 815: FORTRAN subroutines for computing approximate solutions of feedback set problems using GRASP. ACM Transactions on Mathematical Software, 32(1):166– 168, March 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Hatcher:1994:GEI

[Hat94] Philip J. Hatcher. Guest Editor’s introduction: The impact of High Performance Fortran. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):13–??, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Holoien:1991:FES

[HB91a] Martin O. Holoien and Ali Behforooz. Fortran 77 for Engineers and Scientists. Brooks/ Cole Publishing Co., Pacific Grove, CA, USA, second edition, May 1991. ISBN 0-534-14166-8. x + 463 pp. LCCN Q183.9 .H64 1991. US$47.95; US$28.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0534141668.

Holoien:FES91

[HB91b] Martin O. Holoien and Ali Behforooz. FORTRAN 77 for Engineers and Scientists. Brooks/ Cole Publishing Co., Pacific Grove, CA, USA, second edition, 1991. ISBN 0-534-14166-

8. x + 463 pp. LCCN Q183.9 .H64 1991. US$28.00.

Harris:1995:CHP

[HBB^{+}95] Jonathan Harris, John A. Bircsak, M. Regina Bolduc, Jill Ann Diewald, Israel Gale, Neil W. Johnson, Shin Lee, C. Alexander Nelson, and Carl D. Offner. Compiling High Performance Fortran for distributed-memory systems. Digital Technical Journal, 7(3):5–23, Fall 1995. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v7n3/Compiling_
High_Performance_For_02jan1996DTJJ01P8.
ps;   http://www.digital.
com:80/info/DTJJ01/DTJJ01AH.
HTM;   http://www.digital.
com:80/info/DTJJ01/DTJJ01P8.
PS;   http://www.digital.
com:80/info/DTJJ01/DTJJ01PF.
PDF;   http://www.digital.
com:80/info/DTJJ01/DTJJ01SC.
TXT.

Hawick:1993:PUM

[HBD^{+}93] K. A. Hawick, R. S. Bell, A. Dickinson, P. D. Surry, and B. J. N. Wylie. Parallelisation of the unified model data assimilation scheme. In Hoffmann and Kauranne [HK93b], pages 188–203. ISBN 981-02-1429-4. LCCN QA76.58 E354

Hanson:2001:UFI

[HBG01] Richard J. Hanson, Clay P.

Breshears, and Henry A. Gabb. Using a Fortran interface to POSIX threads. In Boisvert and Tang [BT01], pages 257– 272. ISBN 0-7923-7339-1. LCCN QA76.758 .I345 2000.

Hanson:2002:AFI

[HBG02] Richard J. Hanson, Clay P.

Breshears, and Henry A. Gabb. Algorithm 821: a Fortran interface to POSIX threads. ACM Transactions on Mathematical Software, 28(3):354– 371, September 2002. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Hindmarsh:2005:SSN

[HBG^{+}05] Alan C. Hindmarsh, Peter N. Brown, Keith E. Grant, Steven L. Lee, Radu Serban, Dan E. Shumaker, and Carol S. Woodward. SUNDI-ALS: Suite of nonlinear and differential/algebraic equation solvers. ACM Transactions on Mathematical Software, 31(3): 363–396, September 2005. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Herder:2006:MSP

[HBG^{+}06] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. Modular system programming in MINIX 3. ;login: the USENIX Association newsletter, 31(2):19–28, April 2006. CODEN LOGNEM. ISSN 1044-6397. URL http://

www.minix3.org/;   http:/
/www.usenix.org/publications/
login/2006-04/openpdfs/herder.
pdf.

Hansen:1992:FSG

[HC92] Per Christian Hansen and Tony F. Chan. FOR-TRAN subroutines for general Toeplitz systems. ACM Transactions on Mathematical Software, 18(3):256–273, September 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0098-
3500/131768.html. See also

[HC94].

Hansen:1994:CAF

[HC94] Per Christian Hansen and Tony F. Chan. Corrigendum: “Algorithm 729: FOR-TRAN subroutines for general Toeplitz systems”. ACM Transactions on Mathematical Software, 20(1):160, March 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1994-20-1/p160-hansen/. See

[HC92].

Huang:2008:FPM

[HC08] Jih-Woei Huang and Chih-Ping Chu. A flexible processor mapping technique toward data localization for blockcyclic data redistribution. The Journal of supercomputing, 45 (2):151–172, August 2008. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=45&issue=2&spage=
151.

Hart:1998:FPF

[HCD^{+}98] Hal Hart, Jim Caristi, Robert Dewar, Mark Gerhardt, Drew Hamilton, Christopher Haynes, and Sam Rebelsky. The future of programming-are fundamental changes in computer science programs coming? (panel). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 30 (1):370–371, March 1998. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hwang:2003:SAE

[HCLJ03] Gwan-Hwan Hwang, Cheng-Wei Chen, Jenq Kuen Lee, and Roy Dz-Ching Ju. Segmented alignment: An enhanced model to align data parallel programs of HPF. The Journal of supercomputing, 25(1):17–41, May 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//ipsapp009.kluweronline.
com/content/getfile/5189/
43/2/abstract.htm;   http:
//ipsapp009.kluweronline.

                                        com/content/getfile/5189/
                                        43/2/fulltext.pdf.

Hormann:1993:PRN

[HD93] Wolfgang Hörmann and G. Deflinger. A portable random number generator well suited for the rejection method. ACM Transactions on Mathematical Software, 19(4):489–495, December 1993. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/toms/
                                        1993-19-4/p489-hormann/.

Howell:2005:ABG

[HD05] Gary W. Howell and Nadia Diaa. Algorithm 841: BHESS: Gaussian reduction to a similar banded Hessenberg form. ACM Transactions on Mathematical Software, 31(1):166– 185, March 2005. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Hayashi:1994:AAS

[HDH^{+}94] Kenichi Hayashi, Tsunehisa Doi, Takeshi Horie, Yoichi Koyanagi, Osamu Shiraki, Nobutaka Imamura, Toshiyuki Shimizu, Hiroaki Ishihata, and Tatsuya Shindo. AP1000+: architectural support of PUT/ GET interface for parallelizing compiler. ACM SIGPLAN Notices, 29(11): 196–207, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
asplos/195473/p196-hayashi/

.

Hayashi:1995:AAS

[HDH^{+}95] K. Hayashi, T. Doi, T. Horie, Y. Koyanagi, O. Shiraki, N. Imamura, T. Shimizu, H. Ishihata, and T. Shindo. AP1000+: architectural support for parallelizing compilers. Transactions of the Information Processing Society of Japan, 36(7):1680–1690, July 1995. CODEN JSGRD5. ISSN 0387-5806.

Hutton:2003:PGD

[HDR03] Andrew J. Hutton, Stephanie Donovan, and C. Craig Ross, editors. Proceedings of the GCC Developers Summit May 25–27, 2003, Ottawa, Ontario Canada. ????, ????, 2003. ISBN ???? LCCN ???? URL

http://www.linux.org.uk/
~ajh/gcc/gccsummit-2003-
proceedings.pdf.

Hadi:2013:CFA

[HE13] Mohammed F. Hadi and Seyed A. Esmaeili. CUDA Fortran acceleration for the finite-difference time-domain method. Computer Physics Communications, 184(5):1395– 1400, May 2013. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465513000118.

Hempel:1994:API

[Hem94]  R. Hempel.  Application

programming interfaces for SUPRENUM. Parallel Computing, 20(10–11):1519–1526, November 3, 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/parco/cas_sub/
                                        browse/browse.cgi?year=1994&
                                        volume=20&issue=10-11&aid=
                                        910.

Hennecke:1994:RER

[Hen94]  Michael Hennecke.  RAN-

EXP: experimental random number generator package. Computer Physics Communications, 79(2):261–267, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0010465594900728.

Hennecke:1995:FIR

[Hen95]  M. Hennecke.  A Fortran

90 interface to random number generation. Computer Physics Communications, 90 (1):117–120, September 1, 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        001046559500065N.

Herbst:1990:SDU

[Her90] Robert Taylor Herbst. Software Design Using Fortran 77. Prentice-Hall, Upper Saddle River, NJ 07458, USA, February 1990. ISBN 0-13-824681-5. xx + 364 pp. LCCN QA76.73.F25H43 1990. US$32.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0-13-824681-5. Sys-

tem requirements for computer disk: PC; AT and T F77 or Dynix Fortran compiler.

Heuer:1990:FVF

[Heu90] Hans-Otto Heuer. A fast vectorized Fortran 77 program for the Monte Carlo simulation of the three-dimensional Ising system. Computer Physics Communications, 59 (2):387–398, June 1990. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/0010465590901865.

HP:1990:DFL

[Hew90a] Hewlett–Packard Company.

Domain FORTRAN language reference. Apollo Systems Div., a subsidiary of Hewlett– Packard, Palo Alto, CA, USA, version 10.8 edition, 1990. various pp.

HP:1990:FRH

[Hew90b] Hewlett–Packard Company.

FORTRAN/ 9000 reference: HP 9000 series 300/400. Hewlett–Packard Co., Fort Collins, CO, USA, 1990. various pp.

HP:1991:FRHa

[Hew91a] Hewlett–Packard Company. FORTRAN/ 9000 reference: HP 9000 series 700 computers. Hewlett–Packard, Corvallis, OR, USA, 1991. various pp.

HP:1991:FRHb

[Hew91b] Hewlett–Packard Company. FORTRAN/ 9000 reference: HP 9000 series 700 computers, technical addendum. Hewlett– Packard, Corvallis, OR, USA, 1991. various pp.

HP:1992:HFPa

[Hew92a] Hewlett–Packard Company. HP FORTRAN/ 9000 programmer’s guide. Hewlett Packard, Palo Alto, CA, USA, 1992. various pp.

HP:1992:HFPb

[Hew92b] Hewlett–Packard Company. HP FORTRAN/ 9000 programmer’s reference. Hewlett Packard, Palo Alto, CA, USA, 1992. various pp.

HP:2001:HFV

[Hew01] Hewlett–Packard Company, Palo Alto, CA, USA. HP Fortran v2.5 for the Itanium Processor Family Release Note, 2001. 22 pp. URL http://

                                        docs.hp.com/en/5969-7863/
                                        5969-7863.pdf. Manufactur-

ing Part Number: 5969-7863

Hey:1994:GEP

[Hey94] Tony Hey. The Genesis Esprit project — An overview. Parallel Computing, 20(10–11): 1605–1612, November 3, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1994&
volume=20&issue=10-11&aid=
916.

Hawick:1995:EHP

[HF95] K. A. Hawick and G. C. Fox.

Exploiting High Performance Fortran for computational fluid dynamics. In Hertzberger and Serazzi [HS95], pages 413– 419. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Hittner:1995:MFP

[HFMS95] James B. Hittner, Michael S.

Finger, James P. Mancuso, and N. Clayton Silver. A Microsoft Fortran 77 program for contrasting part correlations and related statistics. Educational and Psychological Measurement, 55(5):777–??, 1995. CODEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Hull:1994:ICE

[HFT94] T. E. Hull, Thomas F. Fairgrieve, and Ping Tak Peter Tang. Implementing complex elementary functions using exception handling. ACM Transactions on Mathematical Software, 20(2):215–244, June 1994. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Ano94b].

Hull:1997:ICA

[HFT97] T. E. Hull, Thomas F. Fairgrieve, and Ping Tak Peter Tang. Implementing the complex arcsine and arccosine functions using exception handling. ACM Transactions on Mathematical Software, 23(3):299– 335, September 1997. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
toms/1997-23-3/p299-hull/.

Harrison:1993:PPR

[HGG93] Warren Harrison, Cynthia Gens, and Bruce Gifford. pRETS: A parallel reverseengineering toolset for FOR-TRAN. Journal of Software Maintenance: Research and Practice, 5(1):37–57, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Hanson:2014:NCM

[HH14] Richard J. Hanson and Tim Hopkins. Numerical computing with modern Fortran. Applied mathematics. SIAM Press, Philadelphia, PA, USA, 2014. ISBN 1-61197-311-2 (paperback), 1-61197-312-0 (ebook). xv + 244 pp. LCCN QA76.73.F25 H367 2013.

Hanson:2018:RAM

[HH18] Richard J. Hanson and Tim Hopkins. Remark on Algorithm 539: A Modern Fortran Reference Implementation for Carefully Computing the Euclidean Norm. ACM Transactions on Mathematical Software, 44(3):24:1– 24:23, April 2018. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL https://dl.acm.org/
citation.cfm?id=3134441.

Han:1995:FPC

[HHCS95] Q. Z. Han, S. T. Hsieh, H. C.

Chiang, and H. Z. Sun. A FOR-TRAN program for the CFPs of a Boson system with F spin. Computer Physics Communications, 85(3):463–470, March 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Hall:1993:EUP

[HHK^{+}93] Mary W. Hall, Timothy J. Harvey, Ken Kennedy, Nathaniel McIntosh, Kathryn S. McKinley, Jeffrey D. Oldham, Michael H. Paleczny, and Gerald Roth. Experiences using the ParaScope Editor: an interactive parallel programming tool. ACM SIGPLAN Notices, 28 (7):33–43, July 1993. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Horiguchi:1994:ISP

[HHK94] S. Horiguchi, D. F. Hsu, and M. Kimura, editors. International Symposium on Parallel Architectures, Algorithms, and Networks (ISPAN): proceedings of the 1994, December 14–16, 1994, Kanazawa, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6507-6. LCCN QA76.58 .I5673 1994. IEEE catalog no. 94TH0697-3.

Hall:1992:ICF

[HHKT92] Mary W. Hall, Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Interprocedural compilation of Fortran D for MIMD distributed memory machines. In IEEE [IEE92d], pages 522–534. ISBN 0-8186-2630-5. LCCN QA76.5 .S894 1992. IEEE catalog no. 92CH3216-9.

Hall:1996:ICF

[HHKT96] Mary W. Hall, Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Interprocedural compilation of Fortran D. Journal of Parallel and Distributed Computing, 38(2):114– 129, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jpdc.1996.0135/production;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0135/production/pdf.

Henderson:1990:UDD

[HHLS90] Leslie Ann Henderson, Robert E.

Hiromoto, Olaf M. Lubeck, and Margaret L. Simmons. On the use of diagnostic dependenceanalysis tools in parallel programming: Experiences using PTOOL. The Journal of supercomputing, 4(1): 83–96, March 1990. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=4&issue=1&spage=
83.

Higham:1990:EMN

[Hig90a] Nicholas J. Higham. Experience with a matrix norm estimator. SIAM Journal on Scientific and Statistical Computing, 11(4):804–809, July 1990. CODEN SIJCD4. ISSN 0196-5204.

Higham:1990:EFM

[Hig90b] Nicholas J. Higham. Exploiting fast matrix multiplication within the level 3 BLAS. ACM Transactions on Mathematical Software, 16(4):352–368, December 1990. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1990-16-4/p352-higham/;
http://www.acm.org/pubs/
toc/Abstracts/0098-3500/

98290.html. Describes algorithms based on Strassen’s method which are asymptotically faster than the standard N^{3} algorithm, and in practice, faster for N ≈ 100, and examines their numerical stability. See [DDHD90, DH92, DDP94].

Higham:1991:RBF

[Hig91] Desmond J. Higham. Remark on “Algorithm 669: BRKF45: A FORTRAN subroutine for solving first-order systems of nonstiff initial value problems for ordinary differential equations”. ACM Transactions on Mathematical Software, 17(3):424–426, September 1991. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0098-3500/

116814.html. See [Cas89a].

HPFF:1992:HPF

[Hig92] High Performance Fortran Forum, ???? High Performance Fortran Language Specification, Version 0.4, 1992.

Higginbotham:1993:ISR

[Hig93a] T. F. Higginbotham. The integer square root of N via a binary search. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):41–45, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Forum:1993:HPF

[Hig93b] High Performance Fortran Forum. High Performance Fortran language specification. Scientific Programming, 2(1–2):1–170, Spring–Summer 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

HPPF:1994:HPF

[Hig94a] High Performance Fortran Forum. High Performance Fortran language specification (part III), Version 1.0, May 1993. ACM SIGPLAN FOR-TRAN Forum, 13(3):22–55, September 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

HPPF:1994:SIHa

[Hig94b] High Performance Fortran Forum. Special issue: High Performance Fortran language specification, Version 1.0 (May 3, 1993): Part II. 5. intrinsic and library procedures. ACM SIGPLAN FORTRAN Forum, 13(2):87–138, June 1994. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

HPPF:1994:SIHb

[Hig94c] High Performance Fortran Forum. Special issue: High Performance Fortran language specification, Version 1.0 (May 3, 1993): Part II. 6. extrinsic procedures. ACM SIGPLAN FORTRAN Forum, 13(2):139– 142, June 1994. CODEN ????

ISSN 1061-7264 (print), 1931-1311 (electronic).

HPPF:1994:SIHc

[Hig94d] High Performance Fortran Forum. Special issue: High Performance Fortran language specification, Version 1.0 (May 3, 1993): Part II. 7. storage and sequence association. ACM SIGPLAN FORTRAN Forum, 13(2):143–150, June 1994. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Hnatowicz:1990:GFC

[HIK90] Vladimír Hnatowicz, Valery Ilyushchenko, and Peter Kozma. GSAP: Fortran code for gamma-spectrum analysis. Computer Physics Communications, 60(1):111–125, August 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        001046559090081B.

Hildebrand:1991:CE

[Hil91] J. D. Hildebrand. A competitive edge. UNIX review, 9 (7):22–??, July 1991. CODEN UNRED5. ISSN 0742-3136.

Hollingsworth:1991:IAS

[HIM91] Jeffrey K. Hollingsworth, R. Bruce Irvin, and Barton P. Miller. The integration of application and system based metrics in a parallel program performance tool. Technical Report TR 996, Computer Sciences Department, University of Wisconsin — Madison, Madison, WI, USA, January 1991.

Hinich:2006:BRB

[Hin06] Melvin J. Hinich. Book review:

Developing Statistical Software in Fortran 95 by David R. Lemmon; Joseph L. Schafer. Technometrics, 48(4):571–572, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/25471254.

Hiranandani:1991:OFDa

[Hir91]  Seema Hiranandani.  An

overview of the Fortran D programming system. Technical report COMP TR91-154, Rice University, Dept. of Computer Science, Houston, TX, USA, March 1991. 19 pp.

Hirayama:1991:SFP

[HIS91] H. Hirayama, M. Ikeda, and N. Sagawa. Solution functions of PDEQSOL (partial differential equation solver language) for fluid problems. In IEEE [IEE91], pages 218–227. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Hamilton:1997:AR

[HJ97] Kenneth G. Hamilton and F. James. Acceleration of RANLUX. Computer Physics Communications, 101 (3):241–248, May 1, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465597000027.

Hu:2000:HHP

[HJJ^{+}00] Y. Charlie Hu, Guohua Jin, S. Lennart Johnsson, Dimitris Kehagias, and Nadia Shalaby. HPFBench: a High Performance Fortran benchmark suite. ACM Transactions on Mathematical Software, 26(1): 99–149, March 2000. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
toms/2000-26-1/p99-hu/;
http://www.acm.org/pubs/
citations/journals/toms/
2000-26-1/p99-hu/p99-hu.
pdf.

Hu:1997:HPF

[HJT97] Y. Charlie Hu, S. Lennart Johnsson, and Shang-Hua Teng. High Performance Fortran for highly irregular problems. ACM SIGPLAN Notices, 32(7):13–24, July 1997. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Havlak:1990:EIA

[HK90] P. Havlak and K. Kennedy.

Experience with interprocedural analysis of array side effects. In IEEE [IEE90a], pages 952–961. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Honda:1991:CGP

[HK91] Hiroki Honda and Hironori Kasahara. Coarse grain parallelism detection scheme of a Fortran program. Systems and computers in Japan, 22(12):24– 36, 1991. CODEN SCJAEP. ISSN 0882-1666 (print), 1520-684X (electronic).

Hall:1992:ECG

[HK92] M. W. Hall and K. Kennedy.

Efficient call graph analysis. ACM Letters on Programming Languages and Systems, 1(3):227–42, 1992. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

Hoffmann:1993:PSA

[HK93a] Geerd-R. Hoffmann and Tuomo Kauranne, editors. Parallel supercomputing in atmospheric science: proceedings of the fifth ECMWF Workshop on the Use of Parallel Processors in Meteorology, Reading UK, November 23–27, 1992. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Hoffmann:1993:PFE

[HK93b] Geerd-R. Hoffmann and Tuomo Kauranne, editors. Proceedings of the Fifth ECMWF Workshop on the Use of Parallel Processors in Meteorology. Parallel Supercomputing in Atmospheric Science. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Hu:1993:BRS

[HK93c] C. Hu and R. B. Kearfott. On bounding the range of some elementary functions in FOR-TRAN 77. Interval Computations, 1993(3):29–39, 1993. ISSN 0135-4868.

Hoffmann:1995:CAP

[HK95] Geerd-R. Hoffmann and Norbert Kreitz, editors. Coming of age: proceedings of the Sixth ECMWF Workshop on the Use of Parallel Processors in Meteorology, Reading, UK, November 21–25, 1994. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1995. ISBN 981-02-2211-4. LCCN QC866.E26 1994.

Hirananandani:1991:OFD

[HKK^{+}91a] S. Hirananandani, K. Kennedy, C. Koelbel, U. Kremer, and C-W. Tseng. An Overview of Fortran D Programming System. Technical Report TR91121, CRPC, Rice University, Houston, TX, USA, March 1991.

Hiranandani:1991:OFDb

[HKK^{+}91b] S. Hiranandani, K. Kennedy, C. Koelbel, U. Kremer, and C.- W. Tseng. An overview of the Fortran D programming system. Lecture Notes in Computer Science, 589:18–34, 1991. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hiranandani:1992:OFD

[HKK^{+}92] S. Hiranandani, K. Kennedy, C. Koelbel, U. Kremer, and C.- W. Tseng. An overview of the Fortran D programming system. Lecture Notes in Computer Science, 589:18–34, 1992. CODEN LNCSD9. ISSN 0302-9743.

Hayder:1998:CPL

[HKM98] M. E. Hayder, D. E. Keyes, and P. Mehrotra. A comparison of PETSc library and HPF implementations of an archetypal PDS computation. Advances in engineering software, 29(3):415–423, April 1998. CO-DEN AESODT. ISSN 0045-7949 (print), 1879-2243 (electronic).

Hood:1990:PPD

[HKMC90] R. Hood, K. Kennedy, and J. Mellor-Crummey. Parallel program debugging with on-the-fly anomaly detection. In IEEE [IEE90a], pages 74– 81. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Helgason:1991:SUG

[HKS91] Richard V. Helgason, Jeffery L. Kennington, and Douglas Stewart. S12 user’s guide. Technical Report 91-CSE-6, Southern Methodist University, Dallas, TX, USA, 1991. prize (\ $1.00).

Hanxleden:1994:VDF

[HKS94] R. V. Hanxleden, K. Kennedy, and J. Saltz. Value-based distributions in Fortran D. In Gentzsch and Harms [GH94b], pages 434–440. ISBN 3-540-57981-8 (Berlin: vol. 2: paperback), 0-387-57981-8 (New York: vol. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1– 2 (c1994). DM96.00. Two volumes.

Hobza:1997:PEP

[HKŠ^{+}97] Pavel Hobza, Martin Kabeláč, Jiří Šponer, Petr Mejzlík, and Jiří Vondrášek. Performance of empirical potentials (AMBER, CFF95, CVFF, CHARMM, OPLS, POLTEV), semiempirical quantum chemical methods (AM1, MNDO/M, PM3), and ab initio Hartree–Fock method for interaction of DNA bases: Comparison with nonempirical beyond Hartree–Fock results. Journal of Computational Chemistry, 18(9):1136– 1150, July 15, 1997. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hiranandani:1991:COFb

[HKT91a] S. Hiranandani, K. Kennedy, and C.-W. Tseng. Compiler optimizations for FortranD on MIMD distributedmemory machines. In IEEE [IEE91], pages 86–100. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Hiranandani:1991:COFa

[HKT91b] Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Compiler optimizations for Fortran D on MIMD distributed-memory machines. Technical report COMP TR91-15, Rice University, Dept. of Computer Science, Houston, TX, USA, 1991. 15 pp.

Hiranandani:1991:CSM

[HKT91c] Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Compiler support for machine-independent parallel programming in Fortran D. Technical report COMP TR91-149, Rice University, Dept. of Computer Science, Houston, TX, USA, 1991. 23 pp.

Hiranandani:1991:ECO

[HKT91d] Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Evaluation of compiler optimizations for Fortran D on MIMD distributed-memory machines. Technical report COMP TR91-170, Rice University, Dept. of Computer Science, Houston, TX, USA, 1991. 14 pp.

Hiranandani:1992:CSM

[HKT92a] S. Hiranandani, K. Kennedy, and C. Tseng. Compiler support for machine-independent parallel programming in Fortran D. In Saltz and Mehrotra [SM92], page ?? ISBN 0-444-88712-1. LCCN QA76.9.D5 L36 1992.

Hiranandani:1992:ECO

[HKT92b] S. Hiranandani, K. Kennedy, and C.-W. Tseng. Evaluation of compiler optimizations for Fortran D on MIMD distributed-memory machines. In 6th ACM International Conference on Supercomputing, pages 1–14. ????, Washington, DC, USA, July 1992.

Hiranandani:1992:CFD

[HKT92c] Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Compiling Fortran D for MIMD distributed-memory machines. Comm. ACM, 35(8): 66–80, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-
0782/135230.html.

Hiranandani:1993:PEFa

[HKT93a] Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Preliminary experiences with the Fortran D compiler. Technical report COMP TR93-203, Rice University, Dept. of Computer Science, Houston, TX, USA, April 1993. 13 pp.

Hiranandani:1993:PEFb

[HKT93b] Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Preliminary experiences with the Fortran D compiler. In IEEE [IEE93d], pages 338–350. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Hiranandani:1994:ECO

[HKT94] Seema Hiranandani, Ken Kennedy, and Chau-Wen Tseng. Evaluating compiler optimizations for Fortran D. Journal of Parallel and Distributed Computing, 21 (1):27–45, April 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1994.1040/production;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1994.1040/production/pdf.

Hiranandani:1994:DEN

[HKTW94] Seema Hiranandani, Ken Kennedy, Chau-Wen Tseng, and Scott Warren. The D editor: a new interactive parallel programming tool. In IEEE [IEE94f], pages 733–742, 817. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894

1994. URL http://sc94.
ameslab.gov/AP/contents.

html. IEEE catalog number 94CH34819.

Hanson:1994:BLO

[HL94] R. Hanson and T. Leite. Basic linear operators in Fortran

90. In Gentzsch and Harms [GH94b], pages 441–444. ISBN 3-540-57981-8 (Berlin: vol. 2: paperback), 0-387-57981-8 (New York: vol. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1– 2 (c1994). DM96.00. Two volumes.

Herlihy:2008:DCM

[HL08] Maurice Herlihy and Victor Luchangco. Distributed computing and the multicore revolution. ACM SIGACT News, 39(1):62–72, March 2008. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Hwang:1995:AOS

[HLJ95] Gwan-Hwan Hwang, Jenq Kuen Lee, and Dz-Ching Ju. An array operation synthesis scheme to optimize Fortran 90 programs. ACM SIGPLAN Notices, 30(8):112–122, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hwang:1998:FCA

[HLJ98] Gwan-Hwan Hwang, Jenq Kuen Lee, and Roy Dz-Ching Ju. A function-composition approach to synthesize Fortran 90 array operations. Journal of Parallel and Distributed Computing, 54(1):1–47, October 10, 1998. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1998.1481/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1998.1481/production/pdf;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1998.1481/production/ref.

Hwang:2001:AOS

[HLJ01] G-H. Hwang, J. K. Lee, and R. D. C. Ju. Array operation synthesis to optimize HPF programs on distributed memory machines. Journal of Parallel and Distributed Computing, 61 (4):467–500, April 2001. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Hopkins:1990:RRK

[HM90a] Tim Hopkins and David Morse. Remark on “Algorithm 620: References and keywords for Collected Algorithms of the ACM”. ACM Transactions on Mathematical Software, 16(4): 401–403, December 1990. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [RH84, Ham85].

Hume:1990:URS

[HM90b] Andrew G. Hume and M. Douglas McIlroy, editors. UNIX Research System. Computing Science Research Center, AT&T Bell Laboratories, Murray Hill, NJ, USA, tenth edition, 1990. ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). 702 (volume 1), vi + 630 (volume 2) pp. LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/

. Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Hahn:1992:IAE

[HM92] W. Hahn and K. Mohr. An interactive approach for extending IEEE arithmetic: auxiliary processor AP87. APL Quote Quad, 23(1):88–93, July 1992. CODEN APLQD9. ISSN 0163-6006. International Conference on APL.

Hinojosa:1993:FBF

[HM93] Juan Homero Hinojosa and Kevin L. Mickus. FORELAND BASIN — a FORTRAN program to model the formation of foreland basins resulting from the flexural deflection of the lithosphere caused by a timevarying distributed load. Computers and Geosciences, 19(9): 1321–??, October 1993. CO-DEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Hackstadt:1996:DAQ

[HM96] S. T. Hackstadt and A. D. Malony. Distributed array query and visualization for High Performance Fortran. Lecture Notes in Computer Science, 1123:55–??, ???? 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heltemes:2012:BVF

[HM12] T. A. Heltemes and G. A.

Moses. BADGER v1.0: a Fortran equation of state library. Computer Physics Communications, 183(12):2629– 2646, December 2012. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S001046551200238X.

Honda:1991:PPS

[HMKN91] Hiroki Honda, Satoshi Mizuno, Hironori Kasahara, and Seinosuke Narita. Parallel processing scheme of a basic block in a Fortran program on OS-CAR. Systems and computers in Japan, 22(11):1–??, 1991. CODEN SCJAEP. ISSN 0882-1666 (print), 1520-684X (electronic).

Halatsis:1994:PPA

[HMPT94] C. Halatsis, D. Maritsas, G. Philokyprou, and S. Theodoridis, editors. PARLE ’94: parallel architectures and languages Europe: 6th International PARLE Conference, Athens, Greece, July 4–8, 1994: proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58184-7. LCCN QA76.58 .I564 1994.

Haveraaen:2015:HPD

[HMR^{+}15] Magne Haveraaen, Karla Morris, Damian Rouson, Hari Radhakrishnan, and Clayton Carson. High-performance design patterns for modern Fortran. Scientific Programming, 2015 (??):942059:1–942059:14, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https://

                                        www.hindawi.com/journals/
                                        sp/2015/942059/.

Hwang:1995:RLS

[HMS^{+}95] Yuan-Shin Hwang, Bongki Moon, S. D. Sharma, R. Ponnusamy, R. Das, and J. H. Saltz. Runtime and language support for compiling adaptive irregular programs on distributed-memory machines. Software—Practice and Experience, 25(6):597–621, June 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hurly:1990:ARA

[HMT90] J. J. Hurly, G. T. McConville, and W. L. Taylor. Additions and revisions to the algorithms and Fortran programs to calculate quantum collision integrals for realistic intermolecular potentials. Report MLM-3635, EG and G Mound Applied Technologies, Miamisburg, OH, USA, 1990. 77 pp.

Helmbold:1991:DPE

[HMW91] D. P. Helmbold, C. E. Mcdowell, and J-Z. Wang. Determining possible event orders by analyzing sequential traces. ucsc-crl-91-36, University of California, Santa Cruz, Santa Cruz, CA, USA, September 1991. 28 pp. URL ftp://ftp.

cse.ucsc.edu/pub/tr/ucsc-
crl-91-36.ps.Z;   mailto:
:rnalib@ftp.cs.ucsc.edu.

prize (\ $4.00).

Helmbold:1993:DPE

[HMW93] David P. Helmbold, Charles E.

McDowell, and Jian-Zhong Wang. Determining possible event orders by analyzing sequential traces. IEEE Transactions on Parallel and Distributed Systems, 4(7):827–840, July 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Holmes:1990:SC

[Hol90] Bill Holmes. SPLASH cornerturn. Technical report SRC-TR-91-029, Supercomputing Research Center: IDA, Lanham, MD, USA, March 22, 1990. various pp.

Holzner:1994:BCW

[Hol94] Steve Holzner. Borland C++ for Windows Programming. Brady, February 1, 1994. ISBN 1-56686-119-5. US$40.00.

ACM:1993:ASH

[HOP93] ACM SIGPLAN HOPL-II. 2nd ACM SIGPLAN History of Programming Languages Conference (Preprints), volume 28(3). ACM Press, New York, NY 10036, USA, March 1993. CODEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN ????

Hopkins:1997:BRB

[Hop97] Tim Hopkins. Book review: Numerical algorithms with Fortran by G. Engeln–Műllges and F. Uhlig (Springer, Berlin, 1996), ISBN 3-540-60529-0: Numerical algorithms with C by G. Engeln–Műllges and F. Uhlig (Springer, Berlin, 1996), ISBN 3-540-60530-4.

Computer Physics Communications, 103(1):100–101, June 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465597000246.

Hopkins:1998:CAF

[Hop98] Tim Hopkins. Certification of Algorithm 734: a Fortran 90 code for unconstrained nonlinear minimization. ACM Transactions on Mathematical Software, 24(3):351–354, September 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Hopkins:2002:RAF

[Hop02] Tim Hopkins. Remark on Algorithm 705: a Fortran-77 software package for solving the Sylvester matrix equation AXB^{T} + CXD^{T} = E. ACM Transactions on Mathematical Software, 28(3):372– 375, September 2002. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [GWL^{+}92].

Hopkins:2003:RAF

[Hop03] Tim Hopkins. Remark on Algorithm 769: Fortran subroutines for approximate solution of sparse quadratic assignment problems using GRASP. ACM Transactions on Mathematical Software, 29(3):349– 351, September 2003. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Horw91a

[Hor91a] Jim E. Horwedel. GRESS: a preprocessor for sensitivity studies on Fortran programs. In Andreas Griewank and George F. Corliss, editors, Automatic Differentiation of Algorithms: Theory, Implementation, and Application. SIAM, Philadelphia, PA, 1991.

Horwedel:1991:GAP

[Hor91b] Jim E. Horwedel. GRESS: a preprocessor for sensitivity studies on Fortran programs. In Andreas Griewank and George F. Corliss, editors, Automatic Differentiation of Algorithms: Theory, Implementation, and Application, page ?? SIAM Press, Philadelphia, PA, USA, 1991.

Horwedel:1992:RAD

[Hor92]  Jim E. Horwedel.  Re-

verse automatic differentiation of modular Fortran programs. Technical Memorandum ORNL/TM 12050, Computing and Telecommunications Division, Oak Ridge National Laboratory, Oak Ridge, TN, USA, March 1992.

Horstmann:1996:CCL

[Hor96] C. S. Horstmann. C++ class libraries for numerical programming. C++ Report, 8(1):61– 64, 66, January 1996. CODEN CRPTE7. ISSN 1040-6042.

Horvat:2009:EFP

[Hor09] Vladimir Horvat. ERCS08: a FORTRAN program equipped with a Windows graphics user interface that calculates ECPSSR cross sections for the removal of atomic electrons. Computer Physics Communications, 180(6):995–1003, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465508004517.

Horvat:2023:FFP

[Hor23]  Vladimir Horvat.  FITEVT:

a FORTRAN program for arrival-time analysis of nucleardecay events. Computer Physics Communications, 284 (??):Article 108600, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465522003198.

Howard:1991:PIE

[How91] Kipp E. Howard. The power induced effects module: a FOR-TRAN code which estimates lift increments due to power induced effects for V/STOL flight. Thesis (M.S.), California State Polytechnic University, Pomona, CA, USA, 1991. various pp.

Holden:1995:NFP

[HP95a] Ronald R. Holden and Michael Parent. NORMUL: A FOR-TRAN program for testing multivariate normality. Behavior research methods, instruments, and computers, 27(3): 400–??, 1995. CODEN BRM-CEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Hu:1995:PMC

[HP95b] Hong Hu and Jada M. Paysour. Panel method computational performance on CM-5 and Cray-YMP. Boundary Elements Communications, 6(2): 51–54, March 1995. CODEN BECOFU. ISSN 1353-825X.

Houstis:1992:AIE

[HR92] E. N. (Elias N.) Houstis and J. R. (John Rischard) Rice, editors. Artificial Intelligence, Expert Systems and Symbolic Computing. Selected and Revised Papers from the IMACS 13th World Congress, 22–26 July 1991, Dublin, Ireland. North-Holland Publishing Co., Amsterdam, The Netherlands, 1992. ISBN 0-444-89703-8. LCCN Q334 .I45 1991.

Houstis:1998:PPS

[HRW^{+}98] E. N. Houstis, J. R. Rice, S. Weerawarana, A. C. Catlin, P. Papachiou, K.-Y. Wang, and M. Gaitatzes. PELL-PACK: a problem solving environment for PDE based applications on multicomputer platforms. ACM Transactions on Mathematical Software, 24(1): 30–73, March 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hesham:1994:PTS

[HS94a] E.-R. Hesham and B. D.

Shriver, editors. Proceedings of the Twenty-Seventh Hawaii International Conference on System Sciences. Vol. II: Software Technology, January 4– 7, 1994, Wailea, HI, USA, volume 27. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5060-5. ISSN 1060-3425. LCCN ???? IEEE catalog no. 94TH0607-2.

Hesham:1994:PTH

[HS94b] E.-R. Hesham and B. D.

Shriver, editors. Proceedings of the Twenty-Seventh Hawaii International Conference on System Sciences. Vol.II: Software Technology, January 4– 7, 1994, Wailea, HI, USA, volume 27. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5060-5. ISSN 1060-3425. LCCN ???? IEEE catalog no. 94TH0607-2.

Hertzberger:1995:HCN

[HS95] Bob Hertzberger and Giuseppe Serazzi, editors. Highperformance computing and networking: International Conference and Exhibition, Milan, Italy, May 3–5, 1995: proceedings, volume 919 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Hogg:2010:FRM

[HS10] J. D. Hogg and J. A. Scott. A fast and robust mixedprecision solver for the solution of sparse symmetric linear systems. ACM Transactions on Mathematical Software, 37(2): 17:1–17:24, April 2010. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Hatziargyriou:1991:GEF

[HT91] N. Hatziargyriou and A. Tzortzis. Graphical enhancement of FORTRAN under MS-DOS for computer-aided instruction of electric power engineering. International Journal of Electrical Engineering Education, 28(2):107–??, 1991. CODEN IJEEAF. ISSN 0020-7209.

Huang:1996:LCP

[Hua96] C.-H. Huang, editor. Languages and compilers for parallel computing: 8th international workshop, LCPC ’95, Columbus, Ohio, USA, August 10–12, 1995: proceedings, number 1033 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60765-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1995.

Huddleston:1991:IC

[Hud91a] John V. Huddleston. Introduction to Computers. Exchange Publishing Division, Buffalo, NY, NY, June 1991. ISBN 0-945261-01-2. vi + 170 pp. LCCN QA76.6.H817 1988.

Huddleston:1991:ICV

[Hud91b] John V. Huddleston. Introduction to computers — FOR-TRAN version. Exchange Pub. Division, Buffalo, NY, USA, second edition, 1991. ISBN 0-945261-04-7. 216 pp.

Huddleston:1991:ICF

[Hud91c] John V. Huddleston. Introduction to Computers: Fortran Version. Exchange Publishing Division, Buffalo, NY, NY, second edition, June 1991. ISBN 0-945261-04-7. US$24.00; US$27.50. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0945261047.

Huddleston:1996:F

[Hud96] John V. Huddleston. Fortran 90. Exchange Publishing Division, Buffalo, NY, NY, June 1996. ISBN 0-945261-07-1. 140 pp. US$42.01; US$25.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0945261071.

Huff:1993:LMS

[Huf93]  R. A. Huff.  Lifetime-

sensitive modulo scheduling. In ACM SIGPLAN PLDI ’93 [ACM93b], pages 258–267. CO-DEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7.A152 v. 28, no. 6.

Hughes:1996:FPT

[Hug96] Steven Hughes. Fortran programming tools under Linux. Linux Journal, 25:??, May 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Hunter:2000:EPG

[Hun00] Craig A. Hunter. An evaluation of PowerMac G4 systems for FORTRAN-based scientific computing with application to computational fluid dynamics simulation. Technical report, NASA Langley Research Center, Configuration Aerodynamics Branch, Hampton, VA, USA, July 2000. URL http:

//ad-www.larc.nasa.gov/~
cah/NASA_G4_Study.pdf.

Hillis:1991:WFG

[HW91] William Hillis and James William Welch. WATCOM FORTRAN 77: graphics library reference. WATCOM Publications, Waterloo, Ontario, Canada, 1991. ISBN 1-55094-019-8. vi + 132 pp.

Haering:1995:FPA

[HW95] Edward A. Haering and Stephen A. Whitmore. FOR-TRAN program for analyzing ground-based radar data: usage and derivations. NASA technical paper 3430, National Aeronautics and Space Administration, Office of Management, Scientific and Technical Information Program; Available from the National Technical Information Service, Washington, DC, USA, 1995. iii + 65 pp.

He:2009:AVS

[HWS09] Jian He, Layne T. Watson, and Masha Sosonkina. Algorithm 897: VTDIRECT95: Serial and parallel codes for the global optimization algorithm direct. ACM Transactions on Mathematical Software, 36(3): 17:1–17:24, July 2009. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [SWH15].

Hellberg:1994:PPP

[HZ94] S. A. Hellberg and E. Zaluska. A portable parallel programming environment based around PCTE. Information and Software Technology, 36 (7):419–425, July 1994. CO-DEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

Hempel:1999:AMP

[HZ99] Rolf Hempel and Falk Zimmermann. Automatic migration from PARMACS to MPI in parallel Fortran applications. Scientific Programming, 7(1):39–46, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.com/
                                        app/home/contribution.asp%
                                        3Fwasp=64cr5a4mg33tuhcbdr02%
                                        26referrer=parent%26backto=
                                        issue%2C3%2C7%3Bjournal%
                                        2C8%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

IBM:1991:VFVa

[IBM91a] IBM Corporation. VS FOR-TRAN Version 2 installation and customization for CMS, Release 5. IBM Corporation, New York, NY, USA, fifth edition, 1991. xi + 88 pp.

IBM:1991:VFVb

[IBM91b] IBM Corporation. VS FOR-TRAN Version 2 installation and customization for MVS, Release 5. IBM Corporation, New York, NY, USA, fifth edition, 1991. xi + 97 pp.

IBM:1991:VFVd

[IBM91c] IBM Corporation. VS FOR-TRAN version 2: master index and glossary. IBM Corporation, San Jose, CA, USA, fourth edition, 1991. xi + 64 pp.

IBM:1991:VFVe

[IBM91d] IBM Corporation. VS FOR-TRAN version 2: migration from the parallel FORTRAN PRPQ release 5. IBM Corporation, San Jose, CA, USA, second edition, 1991. 9 pp.

IBM:1991:VFVc

[IBM91e] IBM Corporation. VS FOR-TRAN Version 2 reference summary, Release 5. IBM Corporation, New York, NY, USA, seventh edition, 1991. iv + 77 pp.

IBM:1993:AVR

[IBM93] IBM Corporation, New York, NY, USA. AIX Version 3.2 for RISC System/6000: Optimization and Tuning Guide for Fortran, C, and C++, December 1993. viii + 305 pp. Publication No. SC09-1705-00.

Ixaru:1997:EFP

[IDVV97] L. Gr. Ixaru, H. De Meyer, G. Vanden Berghe, and M. Van Daele. EXPFIT4 — a FORTRAN program for the numerical solution of systems of nonlinear second-order initial-value problems. Computer Physics Communications, 100(1–2):71–80, February 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465596001464.

IEC:1990:III

[IEC90] IEC. ISO/IEC 9593-1 (1990-12): Information processing systems — Computer graphics — Programmer’s Hierarchical Interactive Graphics System (PHIGS) language bindings — Part 1: FORTRAN. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail:

info@iec.ch. URL: http://

www.iec.ch, 1990. 213 pp. US$230.00. URL http://www.

iec.ch/cgi-bin/procgi.pl/
www/iecwww.p?wwwlang=E&wwwprog=
cat-det.p&wartnum=018250.

IEC:1994:IIIb

[IEC94] IEC. ISO/IEC 1539-2 (1994-12) Information technology

— Programming languages — Fortran — Part 2: Varying length character strings. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail:

info@iec.ch. URL: http://

www.iec.ch, 1994. 66 pp. US$110.00. URL http://www.

iec.ch/cgi-bin/procgi.pl/
www/iecwww.p?wwwlang=E&wwwprog=
cat-det.p&wartnum=017866.

IEC:1997:III

[IEC97] IEC. ISO/IEC 1539-1 (1997-12): Information technology

— Programming languages — Fortran — Part 1: Base language. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail: info@

iec.ch. URL: http://www.

iec.ch, 1997. 346 pp. US$249.00. URL http://www.

iec.ch/cgi-bin/procgi.pl/
www/iecwww.p?wwwlang=E&wwwprog=
cat-det.p&wartnum=022616.

IEC:1998:IITa

[IEC98a] IEC. ISO/IEC/TR2 15580 (1998-12): Information technology — Programming languages — Fortran — Floatingpoint exception handling. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail:

info@iec.ch. URL: http://

www.iec.ch, 1998. 27 pp. US$56.00. URL http://www.

iec.ch/cgi-bin/procgi.pl/
www/iecwww.p?wwwlang=E&wwwprog=
cat-det.p&wartnum=023454.

IEC:1998:IITb

[IEC98b] IEC. ISO/IEC/TR2 15581 (1998-12): Information technology — Programming languages — Fortran — Enhanced data type facilities. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail:

info@iec.ch. URL: http://

www.iec.ch, 1998. 13 pp. US$33.00. URL http://www.

                                        iec.ch/cgi-bin/procgi.pl/
                                        www/iecwww.p?wwwlang=E&wwwprog=
                                        cat-det.p&wartnum=023455.

IEC:1999:III

[IEC99] IEC. ISO/IEC 1539-3 (1999-02): Information technology

— Programming languages — Fortran — Part 3: Conditional compilation. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail:

                                        info@iec.ch. URL: http://

www.iec.ch, 1999. 21 pp. US$47.00. URL http://www.

                                        iec.ch/cgi-bin/procgi.pl/
                                        www/iecwww.p?wwwlang=E&wwwprog=
                                        cat-det.p&wartnum=023629.

IEEE:1990:PSN

[IEE90a] IEEE, editor. Proceedings, Supercomputing ’90: November 12–16, 1990, New York Hilton at Rockefeller Center, New York, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

IEE:1990:POS

[IEE90b] IEEE Computer Society.Technical Committee on Operating Systems. Portable operating system interface for computer environments: FORTRAN 77 bindings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN ???? 170 pp. LCCN ????

IEEE:1991:PSA

[IEE91] IEEE, editor. Proceedings, Supercomputing ’91: Albuquerque, New Mexico, November 18–22, 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

IEEE:1992:RIS

[IEE92a] IEEE. 1003.9-1992 (R1998)

IEEE Standard for Information Technology — POSIX\begin{matrix}R \\ ◯ \end{matrix} FORTRAN 77 Language Interfaces — Part 1: Binding for System Application Program Interface (API) [Description]. IEEE, New York, NY, USA, 1992. ISBN 1-55937-230-3 (print), 0-7381-0635-6 (electronic). 186 pp. US$98.00. URL http://

standards.ieee.org/reading/

                                        ieee/std_public/description/
                                        posix/1003.9-1992_desc.html.

IEEE:1992:PRA

[IEE92b]  IEEE, editor.  PESC ’92

Record. 23rd Annual IEEE Power Electronics Specialists Conference, 29 June–3 July 1992, Toledo, Spain. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-7803-0696-1 (casebound), 0-7803-0695-3 (softbound), 0-7803-0697-X (microfiche). LCCN TK 7835 I16 1992. Two volumes. IEEE catalog number 92CH3163-3.

IEEE:1992:PFI

[IEE92c]  IEEE, editor.  Proceedings

of the Fourth IEEE Symposium on Parallel and Distributed Processing, December 1–4, 1992, Arlington, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-3200-3. LCCN QA 76.58 I42 1992. IEEE catalog no. 92TH0492-9.

IEEE:1992:PSM

[IEE92d] IEEE, editor. Proceedings, Supercomputing ’92: Minneapolis, Minnesota, November 16– 20, 1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2630-5. LCCN QA76.5 .S894 1992. IEEE catalog no. 92CH3216-9.

IEEE:1993:DPC

[IEE93a] IEEE, editor. Digest of papers: Compcon spring ’93, San Francisco, California, February 22–26, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3400-6. LCCN QA75.5.C58 1993. IEEE catalog no. 93CH3251-6.

IEEE:1993:IITa

[IEE93b] IEEE. IEEE 1003.9-1993:

Information Technology — POSIX FORTRAN 77 Language Interfaces — Part 1: Binding for System API. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN ???? ???? pp. LCCN ???? US$72. URL http://

www.ansi.org/.

IEEE:1993:PFI

[IEE93c] IEEE, editor. Proceedings of the Fifth IEEE Symposium on Parallel and Distributed Processing: December 1–4, 1993, Dallas, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4222-X. LCCN QA 76.58 I42 1993. IEEE catalog no. 93TH0584-3.

IEEE:1993:PSP

[IEE93d] IEEE, editor. Proceedings, Supercomputing ’93: Portland, Oregon, November 15–19, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

IEEE:1994:FSF

[IEE94a] IEEE, editor. Frontiers’95, the 5th Symposium on the Frontiers of Massively Parallel Computation: proceedings, February 6–9, 1995, McLean, Virginia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

IEEE:1994:IPN

[IEE94b]  IEEE, editor.  ICIP ’94:

proceedings, November 13– 16, 1994, Austin Convention Center, Austin, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6952-7 (casebound), 0-8186-6950-0 (paperback), 0-8186-6951-9 (microfiche). LCCN TA1637.I25 1994. Three volumes. IEEE catalog no. 94CH35708.

IEEE:1994:OOE

[IEE94c] IEEE, editor. Oceans 94: Oceans engineering for today’s technology and tomorrow’s preservation: proceedings, 13–16 September 13–16, 1994, Brest, France, Oceans. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-2057-3, 0-7803-2056-5, 0-7803-2058-1. ISSN 0197-7385. LCCN TC 1505 O33197 1994. Three volumes. IEEE catalog no. 94CH3472-8.

IEEE:1994:PSH

[IEE94d] IEEE, editor. Proceedings of the Scalable High-Performance Computing Conference, May 23–25, 1994, Knoxville, Tennessee, USA, Scalable High Performance Computing Conference. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994.

IEEE:1994:PSP

[IEE94e]  IEEE, editor.  Proceedings

of the Scalable Parallel Libraries Conference , October 6–8, 1993, Mississippi State, Mississippi. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-4980-1. LCCN QA76.58.S34 1993.

IEEE:1994:PSW

[IEE94f] IEEE, editor. Proceedings, Supercomputing ’94: Washington, DC, November 14–18, 1994, Supercomputing. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

IEEE:1994:PIC

[IEE94g] Proceedings of the IEEE Control Systems Society Symposium on Computer-Aided Control System Design, 7–9 March, Tucson, AZ, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN ????

IEEE:1995:IIP

[IEE95a] IEEE, editor. IPPS ’95: 9th International parallel processing symposium — April 25–28, 1995, Santa Barbara, CA, International Parallel Processing Symposium. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

IEEE:1995:PSP

[IEE95b] IEEE, editor. Proceedings of the 1994 Scalable Parallel Libraries Conference: October 12–14, 1994, Mississippi State University, Mississippi. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

IEEE:1996:PII

[IEE96]  IEEE, editor.  Proceedings

of IPPS ’96. The 10th International Parallel Processing Symposium: Honolulu, HI, USA, 15–19 April 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7255-2. LCCN QA76.58 .I565 1996. IEEE catalog number 96TB100038. IEEE Computer Society Press order number PR07255.

IEEE:1997:APD

[IEE97] IEEE, editor. Advances in parallel and distributed computing: March 19–21, 1997, Shanghai, China: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7876-3 (paperback and case), 0-8186-7878-X (microfiche). LCCN QA76.58 .A4 1997.

IEEE:2002:STI

[IEE02] IEEE, editor. SC2002: From Terabytes to Insight. Proceedings of the IEEE ACM SC 2002 Conference, November 16–22, 2002, Baltimore, MD, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1524-X. LCCN ????

IFIP:1993:ECE

[IFI93] IFIP Working Group 2.5. The enable construct for exception handling in Fortran 90. ACM SIGNUM Newsletter, 28(4):7– 16, October 1993. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

IFIP:1995:KWC

[IFI95] IFIP Working Group 2.5, editor. Kyoto Workshop 1995: Current Directions in Numerical Software and High Performance Computing, 19–20 October 1995, Kyoto, Japan. ????, ????, 1995. ISBN ???? LCCN

???? URL http://www.nsc.
liu.se/~boein/ifip/kyoto/
kyoto.html#reid;   http:
//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/
proceedings/.

Iyengar:1994:EBR

[IGHG^{+}94] Arun K. Iyengar, Thaddeus S. Grzeski, Valerie J. Ho-Gibson, Tracy A. Hoover, and John R. Vasta. An event-based, retargetable debugger. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Pack-ard Company, 45(6):33–43, December 1994. CODEN HPJOAX. ISSN 0018-1153.

URL http://www.hp.com/
hpj/94dec/dec94_33.pdf;
http://www.hp.com/hpj/toc-
12-94.html.

Ihaka:2006:BRB

[Iha06] Ross Ihaka. Book review: Developing Statistical Software in Fortran 95, by David R. Lemmon; Joseph L. Schafer.

Journal of the American Statistical Association, 101(474): 857, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/
stable/27590759.

Iwashita:2002:TLH

[IHKvW02] Hidetoshi Iwashita, Kohichiro Hotta, Sachio Kamiya, and Matthijs van Waveren. Towards a lightweight HPF compiler. Lecture Notes in Computer Science, 2327:526–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2327/23270526.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270526.
pdf.

ISO:1990:IPS

[II90]  International  Organization

for Standardization and International Electrotechnical Commission. Systemes de traitement de l’information– infographie–interfaces langage entre un programme d’application et son support graphique. International standard; ISO/IEC 9593-1. International Organization for Standardization, Geneva, Switzerland, 1990. ISBN ???? v + 213 pp. LCCN ????

ISO:1991:ISI

[II91] International Organization for Standardization and International Electrotechnical Commission. International standard: information, technology, programming languages, Fortran. International Organization for Standardization, Geneva, Switzerland, second edition, 1991. ISBN ???? xvii + 369 pp. LCCN ????

Ierotheou:1996:CAP

[IJCL96] C. S. Ierotheou, S. P. Johnson, M. Cross, and P. F. Leggett. Computer aided parallelisation tools (CAPTools)

— conceptual overview and performance on the parallelisation of structured mesh codes. Parallel Computing, 22(2):163– 195, April 5, 1996. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/parco/cas_sub/
                                        browse/browse.cgi?year=1996&
                                        volume=22&issue=2&aid=1055.

Ishizaki:1996:LPA

[IK96] K. Ishizaki and H. Komatsu. A loop parallelization algorithm for HPF compilers. Lecture Notes in Computer Science, 1033:176–190, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

IMSL:1990:UMIa

[IMS90a]  IMSL, Inc.  User’s man-

ual: IMSL SFUN/LIBRARY: FORTRAN subroutines for evaluating special functions. IMSL, Houston, TX, USA, 1990. v + 60 + [23] pp.

IMSL:1990:UMIb

[IMS90b] IMSL, Inc. Users manual:

IMSL STAT/LIBRARY: FOR-TRAN subroutines for statistical analysis. IMSL, Houston, TX, USA, 1990. ???? pp.

IMSL:1991:QR

[IMS91a] IMSL, Inc. Quick reference.

IMSL, Houston, TX, USA, 1991. ???? pp.

IMSL:1991:QRFb

[IMS91b]  IMSL, Inc.  Quick refer-

ence: FORTRAN subroutines for mathematical applications. IMSL, Houston, TX, USA, 1991. 33 pp.

IMSL:1991:QRFa

[IMS91c] IMSL, Inc. Quick reference, FORTRAN subroutines for statistical analysis. IMSL, Houston, TX, USA, 1991. 123 pp.

IMSL:1991:UMFb

[IMS91d]  IMSL, Inc.  User’s man-

ual, FORTRAN subroutines for mathematical applications MATH/LIBRARY. IMSL, Houston, TX, USA, version 2.0 edition, 1991. various pp.

IMSL:1991:UMFc

[IMS91e]  IMSL, Inc.  User’s man-

ual, FORTRAN subroutines for mathematical applications MATH/LIBRARY. IMSL, Houston, TX, USA, 1991. vi + 1372 pp. Three volumes.

IMSL:1991:UMFa

[IMS91f]  IMSL, Inc.  User’s man-

ual, FORTRAN subroutines for mathematical applications MATH/LIBRARY special functions. IMSL, Houston, TX, USA, version 2.0 edition, 1991. various pp.

IMSL:1991:UMFd

[IMS91g]  IMSL, Inc.  User’s man-

ual, FORTRAN subroutines for statistical analysis: STAT/ LIBRARY. IMSL, Houston, TX, USA, version 2.0 edition, 1991. various pp.

IMSL:1991:UMFe

[IMS91h]  IMSL, Inc.  User’s man-

ual, FORTRAN subroutines for statistical analysis: STAT/ LIBRARY. IMSL, Houston, TX, USA, 1991. vi + 1579 pp. Three volumes.

Ingres:1990:IECa

[Ing90a] Ingres. INGRES/EQUEL companion guide for FORTRAN: for the VMS operating system, release 6.3/03, December 1990; INGRES/EQUEL companion guide for COBOL: for the VMS operating system, release 6.3/03, December 1990. Ingres Corp., Alameda, CA, USA, 1990. various pp.

Ingres:1990:IECb

[Ing90b] Ingres. INGRES/ESQL companion guide for FORTRAN: for the VMS operating system, release 6.3/03, December 1990. Ingres Corp., Alameda, CA, USA, 1990. various pp.

Electrical:1991:SIT

[Ins91a] Institute of Electrical and Electronics Engineers. Standard for Information Technology: Posix Fortran 77 Language Interfaces. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, December 1991. ISBN 1-55937-140-4. US$72.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=1-55937-140-4.

IEEE:1991:SIT

[Ins91b] Institute of Electrical and Electronics Engineers and IEEE Computer Society. Technical Committee on Operating Systems. Standard for information technology — POSIX Fortran 77 language interfaces. Institute of Electrical and Electronic Engineers, New York, NY, USA, 1991. ISBN 1-55937-140-4. various pp.

IEEE:1992:ISIb

[Ins92] Institute of Electrical and Electronics Engineers and IEEE Computer Society. Technical Committee on Operating Systems and IEEE Standards Board. IEEE standard for information technology: POSIX FORTRAN 77 language interfaces — Part 1: Binding for system application program interface (API). Std 1003.9-1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 18, 1992. ISBN 1-55937-230-3. xiii + 181 pp. LCCN QA76.76.O63I454 1992. Std 1003.9-1992.

IBM:1990:AXFa

[Int90a] International Business Machines Corporation. AIX XL FORTRAN compiler/6000: language reference. Technical report, IBM Canada Ltd. Laboratory, North York, Ontario, Canada, 1990. xiv + 253 pp.

IBM:1990:AXFb

[Int90b] International Business Machines Corporation. AIX XL FORTRAN compiler/6000: user’s guide. Technical report, IBM Canada Ltd. Laboratory, New York, Ontario, Canada, 1990. xiv + 280 pp.

IBM:1990:IAXa

[Int90c] International Business Machines Corporation. IBM AIX XL FORTRAN compiler/6000: language reference. IBM Corporation, New York, NY, USA, 1990. viii + 180 pp.

IBM:1990:IAXb

[Int90d] International Business Machines Corporation. IBM AIX XL FORTRAN compiler/6000: user’s guide. IBM Corporation, New York, NY, USA, 1990. viii + 96 pp.

IBM:1990:VVF

[Int90e] International Business Machines Corporation. VAST-2 for VS FORTRAN: user’s guide: program offering. IBM Corporation, San Jose, CA, USA, second edition, 1990. vii + 164 pp.

IBM:1990:VFLa

[Int90f] International Business Machines Corporation. VS FOR-TRAN language and library reference version 2. IBM Corporation, New York, NY, USA, seventh edition, 1990. xx + 487 pp.

IBM:1990:VFLb

[Int90g] International Business Machines Corporation. VS FOR-TRAN language and library reference version 2. IBM Corporation, New York, NY, USA, sixth edition, 1990. xx + 474 pp.

IBM:1990:VFP

[Int90h] International Business Machines Corporation. VS FOR-TRAN programming guide for CMS and MVS version 2. IBM Corporation, New York, NY, USA, sixth edition, 1990. xx + 460 pp.

IBM:1990:VFVb

[Int90i] International Business Machines Corporation. VS FOR-TRAN version 2: general information, release 5. IBM Corporation, New York, NY, USA, 7th (March 1990) edition, 1990. xvi + 54 pp.

IBM:1990:VFVc

[Int90j] International Business Machines Corporation. VS FOR-TRAN version 2: general information, release 5. IBM Corporation, New York, NY, USA, eighth edition, 1990. xii + 50 pp.

IBM:1990:VFVd

[Int90k] International Business Machines Corporation. VS FOR-TRAN version 2: language and library reference. IBM Corporation, San Jose, CA, USA, sixth edition, 1990. xx + 474 pp.

IBM:1990:VFVa

[Int90l] International Business Machines Corporation. VS FOR-TRAN version 2 language and library reference, Release 5. IBM Corporation, New York, NY, USA, seventh edition, 1990. xx + 486 pp.

IBM:1990:VFVe

[Int90m] International Business Machines Corporation. VS FOR-TRAN version 2: programming guide for AIX/ 370. IBM Corporation, San Jose, CA, USA, 1990. xiv + 240 pp.

IBM:1990:VFVf

[Int90n] International Business Machines Corporation. VS FOR-TRAN version 2: programming guide for CMS and MVS. IBM Corporation, San Jose, CA, USA, sixth edition, 1990. xviii + 529 pp.

IBM:1991:VFL

[Int91a] International Business Machines Corporation. VS FOR-TRAN language and library reference version 2. IBM Corporation, New York, NY, USA, eighth edition, 1991. xx + 487 pp.

IBM:1991:VFP

[Int91b] International Business Machines Corporation. VS FOR-TRAN programming guide version 2. IBM Corporation, New York, NY, USA, seventh edition, 1991. xviii + 559 pp.

IBM:1991:VFVg

[Int91c] International Business Machines Corporation. VS FOR-TRAN version 2: general information. IBM Corporation, San Jose, CA, USA, tenth edition, 1991. xi + 51 pp.

IBM:1991:VFVh

[Int91d] International Business Machines Corporation. VS FOR-TRAN version 2: language and library reference. IBM Corporation, San Jose, CA, USA, eighth edition, 1991. xix + 482 pp.

IBM:1991:VFVf

[Int91e] International Business Machines Corporation. VS FOR-TRAN Version 2 language and library reference, Release 5.

IBM Corporation, New York, NY, USA, eighth edition, 1991. xix + 481 pp.

IBM:1991:VFVi

[Int91f] International Business Machines Corporation. VS FOR-TRAN version 2: programming guide for CMS and MVS. IBM Corporation, San Jose, CA, USA, seventh edition, 1991. xviii + 559 pp.

IBM:1992:OTG

[Int92] International Business Machines Corporation. Optimization and tuning guide for the XL FORTRAN and XL C compilers. Technical report, IBM Canada Ltd. Laboratory, North York, Ontario, Canada, 1992. x + 234 pp.

ISO:1997:IIIa

[Int97a]  International  Organization

for Standardization. ISO/ IEC 1539-1:1997: Information technology — Programming languages — Fortran — Part 1: Base language. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? ???? pp. LCCN ???? US$195.00. URL

http://webstore.ansi.org/
ansidocstore/product.asp?
sku=ISO%2FIEC+1539%2D1%3A1997.

ISO:1997:IIIg

[Int97b]  International  Organization

for Standardization. ISO/ IEC 1539-1:1997: Information technology — Programming languages — Fortran — Part 1: Base language. International Organization for Standardization, Geneva, Switzerland, 1997. ISBN ???? 346 pp. LCCN ???? CHF 288. URL http:

//www.ansi.org/;   http:/
/www.fortran.com/fortran/
Books/f95std.html;   http:
//www.iso.ch/cate/d26933.

html. ISO/IEC JTC 1/SC 22/WG 5. This is the Fortran 95 Standard. Available in English only.

ISO:1998:IITc

[Int98a] International Organization for Standardization. ISO/IEC TR 15580:1998: Information technology — Programming languages — Fortran — Floating-point exception handling. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 27 pp. LCCN ???? CHF 104; US$72.00. URL

http://webstore.ansi.org/
ansidocstore/product.asp?
sku=ISO%2FIEC+TR+15580%3A1998;
http://www.iso.ch/cate/d28230.

html. Available in English only.

ISO:1998:IITd

[Int98b] International Organization for Standardization. ISO/IEC TR 15581:1998: Information technology — Programming languages — Fortran

— Enhanced data type facilities. International Organization for Standardization, Geneva, Switzerland, 1998. ISBN ???? 13 pp. LCCN ???? CHF 68; US$46.00. URL

                                        http://webstore.ansi.org/
                                        ansidocstore/product.asp?
                                        sku=ISO%2FIEC+TR+15581%3A1998;
                                        http://www.iso.ch/cate/d28231.

html. Available in English only.

ISO:1999:IIIe

[Int99]  International  Organization

for Standardization. ISO/ IEC 1539-3:1999: Information technology — Programming languages — Fortran — Part 3: Conditional compilation. International Organization for Standardization, Geneva, Switzerland, 1999. ISBN ???? 21 pp. LCCN ???? CHF 92; US$62.00. URL

                                        http://webstore.ansi.org/
                                        ansidocstore/product.asp?
                                        sku=ISO%2FIEC+1539%2D3%3A1999;
                                        http://www.iso.ch/cate/d29926.

html. Available in English only.

IBM:19xx:FM

[Intxx] International Business Machines Corporation. FOR-TRAN manual. IBM Corporation, New York, NY, USA, 19xx. 97 pp.

ISO:2000:IIIe

[Int00]  International  Organization

for Standardization. ISO/ IEC 1539-2:2000: Information technology – Programming languages – Fortran – Part 2: Varying length character strings. International Organization for Standardization, Geneva, Switzerland, 2000. ISBN ???? ???? pp. LCCN ???? US$58.00. URL

http://webstore.ansi.org/
ansidocstore/product.asp?
sku=ISO%2FIEC+1539%2D2%3A2000.

Irvine:1991:FVC

[Irv91] Ron Irvine. Fortran vs. C as language for numerical applications. C Users Journal, 9(11): 132–??, November 1991. ISSN 0898-9788.

Iwashita:2002:VFD

[ISKvW02] Hidetoshi Iwashita, Naoki Sueyasu, Sachio Kamiya, and Matthijs van Waveren. VPP Fortran and the design of HPF/JA extensions. Concurrency and Computation: Practice and Experience, 14(8–9): 575–588, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/95016131/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016131{\&}PLACEBO=IE.
pdf.

ISO:1990:IIIa

[ISO90] ISO. ISO/IEC 9593-1:1990:

Information processing systems — Computer graphics

— Programmer’s Hierarchical Interactive Graphics System (PHIGS) language bindings — Part 1: FORTRAN. International Organization for Standardization, Geneva, Switzerland, 1990. ISBN ???? 213 pp. LCCN ???? CHF 236.

                                        URL http://www.ansi.org/;
                                         http://www.iso.ch/cate/

d17366.html. Available in English only. ISO/IEC JTC 1/SC 24/WG 4.

ISO:1994:IIIe

[ISO94] ISO. ISO/IEC 1539-2:1994: Information technology — Programming languages — FOR-TRAN — Part 2: Varying length character strings. International Organization for Standardization, Geneva, Switzerland, 1994. ISBN ???? 66 pp. LCCN ???? CHF 152.

                                        URL http://www.ansi.org/;
                                         http://www.iso.ch/cate/

d6129.html. Available in English only.

ISO:2000:FSI

[ISO00] ISO/IEC JTC1/SC22/WG5. Information technology — programming languages — Fortran – floating-point exception handing (draft). World-Wide Web document., January 19, 2000. URL ftp://ftp.

                                        nag.co.uk/sc22wg5/N1351-
                                        N1400/N1378.pdf.

ISO:2004:DIS

[ISO04a] ISO. Draft International Standard ISO/IEC 1539-1:2004(E): Information technology — Programming languages — Fortran Part 1: Base Language. International Organization for Standardization, Geneva, Switzerland, May 2004. xiv + 569 pp. URL ftp:

//ftp.nag.co.uk/sc22wg5/
N1601-N1650/N1601.pdf.gz.

ISO:2004:IIIa

[ISO04b] ISO. ISO/IEC 1539-1:2004 Information technology — Programming languages — Fortran – Part 1: Base language. International Organization for Standardization, Geneva, Switzerland, 2004. xiv + 569 pp. URL http:

//www.dkuug.dk/jtc1/sc22/
open/n3661.pdf.

ISO:2010:IIIb

[ISO10] ISO. ISO/IEC 1539-1:2010 Information technology — Programming languages — Fortran — Part 1: Base language. International Organization for Standardization, Geneva, Switzerland, June 2010. xviii + 603 pp. URL ftp:

//ftp.nag.co.uk/sc22wg5/
N1801-N1850/N1830.pdf.

Satake:2012:OGA

[iSYS12] Shin ichi Satake, Hajime Yoshimori, and Takayuki Suzuki. Optimizations of a GPU accelerated heat conduction equation by a programming of CUDA Fortran from an analysis of a PTX file. Computer Physics Communications, 183 (11):2376–2385, November 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0010465512002068.

Jung:1992:HET

[JA92] C. C.-D. Jung and D. Advani. HiPPI exploitation in TCP/IP environment. In IEEE [IEE92d], pages 610–615. ISBN 0-8186-2630-5. LCCN QA76.5 .S894 1992. IEEE catalog no. 92CH3216-9.

James:1990:RPN

[Jam90] F. James. A review of pseudorandom number generators. Computer Physics Communications, 60(3):329–344, October 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

James:1994:RFI

[Jam94] F. James. RANLUX: a Fortran implementation of the highquality pseudorandom number generator of Lüscher. Computer Physics Communications, 79(1):111–114, February 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        001046559490233X.

James:1996:ERF

[Jam96] F. James. Erratum: RAN-LUX: A Fortran implementation of the high-quality pseudorandom number generator of Lüscher [Comput. Phys. Commun. 79 (1994) 111–114]. Computer Physics Communications, 97(3):357, September 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0010465596000653.  See

[Jam94].

Joisha:2001:ECO

[JB01a] Pramod G. Joisha and Prithviraj Banerjee. The efficient computation of ownership sets in HPF. IEEE Transactions on Parallel and Distributed Systems, 12(8):769– 788, August 2001. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td2001/pdf/
l0769.pdf;   http://www.
computer.org/tpds/td2001/
l0769abs.htm.

Joisha:2001:EOS

[JB01b] Pramod G. Joisha and Prithviraj Banerjee. Exploiting ownership sets in HPF. Lecture Notes in Computer Science, 2017:259–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/2017/20170259.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2017/20170259.
pdf.

Jablonowski:1993:VVA

[JBBH93] David J. Jablonowski, John D.

Bruner, Brian Bliss, and Robert B. Haber. VASE: The visualization and application steering environment. In IEEE [IEE93d], pages 560–569. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Jin:1993:OFP

[JC93] Guo Hua Jin and Fu Jie Chen. Optimizing FORTRAN programs for hierarchical memory parallel processing systems. J. Comput. Sci. Tech. (English Ed.), 8(3):209–220, 1993.

Jezequel:2010:NVC

[JCL10] Fabienne Jézéquel, Jean-Marie Chesneaux, and Jean-Luc Lamotte. A new version of the CADNA library for estimating round-off error propagation in Fortran programs. Computer Physics Communications, 181 (11):1927–1928, November 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465510002419.

Jezequel:1993:TPT

[Jéz93] Jean-Marc Jézéquel. Transparent parallelisation through reuse: Between a compiler and a library approach. In O. Nierstrasz, editor, Proceedings ECOOP ’93, LNCS 707, pages 384–405. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., July 1993.

Jazayeri:1986:OCH

[JH86] M. Jazayeri and M. Haden.

Optimizing compilers are here (mostly). ACM SIGPLAN Notices, 21(5):61–63, May 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1993:RMI

[JL93] Dale A. Johnson and James E.

Labarre. Retraining management information systems for its role in computer integrated manufacturing. Journal of systems management, 44(9):18– ??, September 1993. CODEN JSYMA9. ISSN 0022-4839.

Johnson:1994:CM

[JM94] Stephen C. Johnson and Cleve Moler. Compiling Matlab. In USENIX Association [USE94], pages 119–127. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

Jones:1992:LAF

[Jon92a] H. W. Jones. Löwdin alphafunction, overlap integral, and computer algebra. International Journal of Quantum Chemistry, 41(5):749–754, March 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1992:LFO

[Jon92b] Herbert W. Jones. Löwdin α-function, overlap integral, and computer algebra. International Journal of Quantum Chemistry, 41(5):749–754, March 5, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jonas:1993:TPL

[Jon93] Myrtle I. Jonas. Techniques in Programming Logic: With an Introduction to Basic and Fortran. Capitol Press, August 1993. ISBN 0-9635298-0-3. US$48.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0963529803.

Jonasson:2009:ADF

[Jon09] Kristjan Jonasson. Algorithm 892: DISPMODULE, a Fortran 95 module for prettyprinting matrices. ACM Transactions on Mathematical Software, 36(1):6:1–6:7, March 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Jordan:1990:FUMa

[Jor90a] H. F. Jordan. Force user’s manual a portable, parallel FOR-TRAN. NASA contractor report NASA CR-4265, National Aeronautics and Space Administration, Office of Management, Scientific and Technical Information Division, Washington, DC, USA, 1990. ???? pp. For sale by the National Technical Information Service.

Jordan:1990:FUMb

[Jor90b] Harry F. (Harry Frederick) Jordan. Force user’s manual: a portable, parallel FORTRAN. NASA contractor report NASA CR-426, National Aeronautics and Space Administration, Office of Management, Scientific and Technical Information Division, Washington, DC, USA, 1990. iii + 33 pp. For sale by the National Technical Information Service.

Joubert:1995:FAH

[Jou95] A. Joubert. Financial applications and HPF. In Dongarra [Don95], pages 317–336. ISBN 0-444-82163-5. ISSN 0927-5452. LCCN QA76.88 .H54 1995.

Joyner:1992:FPC

[Joy92] William B. Joyner. A FOR-TRAN program for calculating nonlinear seismic ground response. Open-file report 77-671., U.S. Geological Survey, Reston, VA, USA, 1992. 50 pp.

Jones:1995:AFS

[JP95] Mark T. Jones and Paul E.

Plassmann. Algorithm 740: Fortran subroutines to compute improved incomplete Cholesky factorizations. ACM Transactions on Mathematical Software, 21(1):18–19, March 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Johnson:2020:AFC

[JPE20] S. R. Johnson, A. Prokopenko, and K. J. Evans. Automated Fortran C++ bindings for large-scale scientific applications. Computing in Science and Engineering, 22(5):84–94, 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

James:1993:ANM

[JSW93] M. L. James, G. M. Smith, and J. C. Wolford. Applied Numerical Methods for Digital Computation. HarperCollins College Publishers, New York, NY, USA, fourth edition, 1993. ISBN 0-06-500494-9. x + 709 pp. LCCN QA297.J3 1993. US$87.41.

Jonasson:2020:AFS

[JSY^{+}20] Kristjan Jonasson, Sven Sigurdsson, Hordur Freyr Yngvason, Petur Orri Ragnarsson, and Pall Melsted. Algorithm 1005: Fortran subroutines for reverse mode algorithmic differentiation of BLAS matrix operations. ACM Transactions on Mathematical Software, 46 (1):9:1–9:20, March 2020. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3382191.

Jenks:1994:HMA

[JT94] Richard D. Jenks and Barry M. Trager. How to make AXIOM into a Scratchpad. In ACM [ACM94c], pages 32–40. ISBN 0-89791-638-7. LCCN ????

URL http://www.acm.org:
80/pubs/citations/proceedings/
issac/190347/p32-jenks/.

Justice:1992:FFR

[Jus92] Harold R. Justice. FORTREX:

a Fortran runtime executive. Dr. Dobb’s Journal of Software Tools, 17(6):82–??, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Kilian:1995:CLE

[KA95] Jens Kilian and Heinz-Peter Arndt. Common Lisp as an embedded extension language. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(5):69–73, October 1995. CODEN HPJOAX. ISSN 0018-1153. URL http://

www.hp.com/hpj/toc-10-95.
html.

Kedward:2022:SF

[KAČ^{+}22] Laurence J. Kedward, Bálint Aradi, Ondřej Čertík, Milan Curcic, Sebastian Ehlert, Philipp Engel, Rohit Goswami, Michael Hirsch, Asdrubal Lozada-Blanco, Vincent Magnin, Arjen Markus, Emanuele Pagone, Ivan Pribec, Brad Richardson, Harris Snyder, John Urban, and Jérémie Vandenplas. The state of Fortran. Computing in Science and Engineering, 24(2):63–72, March/ April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kahan:2001:SFP

[Kah01] W. Kahan. SRTEST: a Fortran program to test any SRT divider’s logic for quotientdigit selection. World-Wide Web document, August 6,

                                        2001.  URL http://www.
                                        cs.berkeley.edu/~wkahan/
                                        srtest/.

Kammler:2000:FCF

[Kam00] David W. Kammler. A First Course in Fourier Analysis. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-578782-3. xv + 790 pp. LCCN QA403.5 .K36 2000.

Kapinos:2010:PPP

[KaM10] Paul Kapinos and Dieter an Mey. Productivity and performance portability of the OpenMP 3.0 tasking concept when applied to an engineering code written in Fortran 95. International Journal of Parallel Programming, 38(5– 6):379–395, October 2010. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

                                        tronic). URL http://www.
                                        springerlink.com/openurl.
                                        asp?genre=article&issn=0885-
                                        7458&volume=38&issue=5&spage=
                                        379.

Karin:1995:PAI

[Kar95] Sidney Karin, editor. Proceedings of the 1995 ACM/ IEEE Supercomputing Conference, December 3–8, 1995, San Diego Convention Center, San Diego, CA, USA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-89791-816-9. LCCN QA 76.88 S8573 1995a. URL http:

//www.supercomp.org/sc95/

proceedings/. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Karp:1996:BRU

[Kar96] Alan H. Karp. Bit reversal on uniprocessors. SIAM Review, 38(1):1–26, March 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.
siam.org/25698.htm.

Kasahara:1993:SSP

[Kas93] H. Kasahara. System software for parallel processing. Joho-Shori (J. Information Processing Soc. Japan), 34(9):1134– 1138, September 1993. CO-DEN JOSHA4. ISSN 0447-8053.

Kaylen:1990:SFS

[Kay90] Michael S. Kaylen. STS1:

a Fortran subroutine for estimating univariate structural time series models. Aewp; 1990-8 agricultural economics working paper; 1990-8., University of Missouri-Columbia, Dept. of Agricultural Economics, Columbia, MO, USA, 1990. 18 pp.

Kohn:1994:RPP

[KB94] S. R. Kohn and S. B. Baden. A robust parallel programming model for dynamic nonuniform scientific computations. In IEEE [IEE94d], pages 509–517. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994.

Kennedy:1994:IIF

[KBKT94] J. G. Kennedy, M. Behr, V. Kalro, and T. E. Tezduyar. Implementation of implicit finite element methods for incompressible flows on the CM-5. In Anonymous [Ano94p], pages 95–111. CODEN CM-MECC. ISBN ???? ISSN 0045-7825, 0374-2830. LCCN ????

Klinker:1994:PPV

[KC94] Gudrun J. Klinker and I-Yu Chen. PerfVisS: a performance visualizer for high performance Fortran programs on workstation farms. Technical report series / Cambridge Research Laboratory CRL 94/4, Digital Equipment Corp., Cambridge Research Laboratory, Cambridge, MA, USA, June 30, 1994. 18. pp.

Kearfott:1994:AIP

[KDDH94] R. B. Kearfott, M. Dawande, K. Du, and C. Hu. Algorithm 737: INTLIB: a portable Fortran-77 elementary function library. ACM Transactions on Mathematical Software, 20(4): 447–459, December 1994. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kyriakidis:1999:CNS

[KDG99] Phaedon C. Kyriakidis, Clayton V. Deutsch, and Marshall L. Grant. Calculation of the normal scores variogram used for truncated Gaussian lithofacies simulation: Theory and FORTRAN code. Computers and Geosciences, 25(2): 161–??, ???? 1999. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Kearfott:1992:IPF

[KDKSH92] R. B. Kearfott, M. Dawande, Du K.-S., and C. Hu. INTLIB: a portable FORTRAN 77 elementary function library. Interval Computations, 3(5):96– 105, 1992. ISSN 0135-4868.

Kerns:1993:BSI

[KE93] Daniel R. Kerns and Susan J.

Eggers. Balanced scheduling: instruction scheduling when memory latency is uncertain. ACM SIGPLAN Notices, 28(6):278–289, June 1993. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/155090/p278-kerns/.

Keady:1992:FSP

[Kea92] G. Keady. Fortran subroutines produced from computer algebra systems: using GEN-TRANS from REDUCE and from MACSYMA. In Noye et al. [NBC92], pages 265–272. ISBN 0-86396-172-X. LCCN ????

Kearfott:1995:IFM

[Kea95a] R. B. Kearfott. INTLIB ARITHMETIC: A Fortran 90 module for an interval data type, 1995. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kearfott:1995:FERa

[Kea95b] R. Baker Kearfott. A Fortran 90 environment for research and prototyping of enclosure algorithms for constrained and unconstrained nonlinear equations. ACM Transactions on Mathematical Software, 21(1): 63–78, March 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kearfott:1996:AIF

[Kea96a] R. Baker Kearfott. Algorithm 763: INTERVAL ARITHMETIC: A Fortran 90 module for an interval data type. ACM Transactions on Mathematical Software, 22(4):385–392, December 1996. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/toms/

1996-22-4/p385-kearfott/.

See [KDDH94].

Kearfott:1996:IFM

[Kea96b] R. Baker Kearfott. Algorithm 763: INTERVAL ARITHMETIC: A Fortran 90 module for an interval data type. ACM Transactions on Mathematical Software, 22(4):385–392, December 1996. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [KDDH94].

Keffer:1992:WCW

[Kef92] Thomas Keffer. Why C++ will replace Fortran. Dr. Dobb’s Journal of Software Tools, 17 (12 (special supplement)):39s, 40s, 42s–47s, December 1992. CODEN DDJOEB. ISSN 1044-789X.

KSR:1991:KFP

[Ken91] Kendall Square Research Corporation, Waltham, MA. KSR Fortran Programming, 1991. ???? pp.

KSR:1992:KFP

[Ken92a] Kendall Square Research Corporation. KSR Fortran programming. Kendall Square Research Corp., Waltham, MA, USA, 1992. various pp.

Kennedy:1992:SSF

[Ken92b] Ken Kennedy. Software for supercomputers of the future. The Journal of supercomputing, 5(4):251–262, February 1992. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL

http://www.springerlink.
com/openurl.asp?genre=article&
issn=0920-8542&volume=5&
issue=4&spage=251.

Kennedy:1994:CTM

[Ken94a] K. Kennedy. Compiler technology for machine-independent parallel programming. International Journal of Parallel Programming, 22(1):79–98, February 1994. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Kennedy:1994:PPS

[Ken94b] K. Kennedy. Parallel programming support in Fortran D and High Performance Fortran. In Anonymous [Ano94d], pages 427–432. ISBN ???? LCCN ????

Kerr:1990:FSP

[Ker90] Clayton P. Kerr. Fortran subroutines for process design. ??, ??, 1990. ISBN ???? 223 + [8] pp. LCCN ????

Kerrigan:1991:FCa

[Ker91a] James F. Kerrigan. From Fortran to C. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, June 1991. ISBN 0-8306-7661-9. US$24.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0830676619.

Kerrigan:1991:FCb

[Ker91b] James F. Kerrigan. From Fortran to C. Windcrest/ McGraw-Hill, Blue Ridge Summit, PA, USA, 1991. ISBN 0-8306-8661-4 (hardcover), 0-8306-7661-9 (paperback). xiv + 298 pp. LCCN QA76.73.C15 K48 1991. US$34.95 (hardcover), US$24.95 (paperback).

URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0830686614.

Kerrigan:1993:MF

[Ker93a] James F. Kerrigan. Migrating to Fortran 90. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1993. ISBN 1-56592-049-X. xxvi + 361 pp. LCCN QA76.73.F28 K47 1993. US$27.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=156592049X.

Kerrigan:MF93

[Ker93b] James F. Kerrigan. Migrating to Fortran 90. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1993. ISBN 1-56592-049-X. xxvi + 361 pp. LCCN QA76.73.F28 K47 1993. US$24.95. URL http://www.

loc.gov/catdir/enhancements/
fy0715/94174237-d.html.

Kerrigan:1993:MFP

[Ker93c] Jim Kerrigan. Migrating to Fortran 90 (Programming Languages). O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1993. ISBN 1-56592-049-X.

Kessell:1992:FDS

[Kes92] Stephen R. Kessell. FORTRAN 77, documentation and style. University revision notes series. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-55611-1 (paperback). xi + 140 pp.

Koffman:1990:PSS

[KF90] Elliot B. Koffman and Frank L. Friedman. Problem solving and structured programming in FORTRAN 77. Addison-Wes-ley, Reading, MA, USA, fourth edition, 1990. ISBN 0-201-51216-5. xvi + 652 pp. LCCN QA76.73.F25 K64 1990.

Koffman:1992:FWE

[KF92a] Elliot B. Koffman and Frank Friedman. FORTRAN, W/Engineering Applns ed. Addison-Wes-ley, Reading, MA, USA, fifth edition, November 1992. ISBN 0-201-55875-0. 675 pp. US$30.95. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=0201558750.

Koffman:1992:F

[KF92b] Elliot B. Koffman and Frank L. Friedman. Fortran. Addison-Wesley, Reading, MA, USA, 5th with engineering applications edition, 1992. ISBN 0-201-55875-0. ix + 664 + [53] pp.

Koffman:1992:FEA

[KF92c] Elliot B. Koffman and Frank L.

Friedman. Fortran: With Engineering Applications. Addison-Wesley, Reading, MA, USA, November 1, 1992. ISBN 0-201-55875-0. US$44.25.

Koffman:1992:PSS

[KF92d] Elliot B. Koffman and Frank L.

Friedman. Problem solving and structured programming in FORTRAN 77. Addison-Wesley, Reading, MA, USA, fifth edition, November 1992. ISBN 0-201-51216-5. LCCN QA76.73.F25K64 1990. US$46.25. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0201512165.

Koffman:1993:F

[KF93a] Elliot B. Koffman and Frank L.

Friedman. Fortran. Addison-Wesley, Reading, MA, USA, 5th ed. with engineering applications edition, 1993. ISBN 0-201-55875-0. ix + 664 + [53] pp. LCCN QA76.73.F25 F63 1993.

Koffman:1993:FEA

[KF93b] Elliot B. Koffman and Frank L.

Friedman. Fortran: with engineering applications. Addison-Wesley, Reading, MA, USA, fifth edition, 1993. ISBN 0-201-55875-0 (paperback). xviii + 664 pp.

Koffman:1996:F

[KF96] Elliot B. Koffman and Frank L. Friedman. Fortran. Addison-Wesley, Reading, MA, USA, fifth edition, October 1996. ISBN 0-201-59062-X. 690 pp. US$30.25. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=020159062X.

Koffman:1997:F

[KF97] Elliot B. Koffman and Frank L. Friedman. Fortran. Addison-Wesley, Reading, MA, USA, 5th ed. updated edition, 1997. ISBN 0-201-59062-X. various pp. LCCN QA76.73.F25 K637 1997. Rev. and updated ed. of: Fortran with engineering applications. 5th ed. c1993. Includes index.

Kirkup:1999:BRB

[KG99] S. Kirkup and G. C. Gaunaurd. Book reviews: Boundary Element Method in Acoustics: A Development in Fortran. Applied mechanics reviews, 52(3): B22, ???? 1999. CODEN AM-READ. ISSN 0003-6900.

Koppler:1997:VDD

[KGV97] Rainer Koppler, Siegfried Grabner, and Jens Volkert. Visualization of distributed data structures for High Performance Fortran-like languages. Scientific Programming, 6(1): 115–126, Spring 1997. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kolte:1993:LRA

[KH93] P. Kolte and M. J. Harrold. Load/store range analysis for global register allocation. In ACM SIGPLAN PLDI ’93 [ACM93b], pages 268–277. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7.A152 v. 28, no. 6.

Kuiper:2013:FPG

[KH13] Rebecca M. Kuiper and Herbert Hoijtink. A Fortran 90 program for the generalized order-restricted information criterion. Journal of Statistical Software, 54(8):??, September 2013. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v54/i08.

Khan:1992:OHO

[KHC92] B. Khan, L. Hayes, and A. P.

Cracknell. The optimisation of higher order resampling methods in a multiprocessor environment. Parallel Computing, 18(12):1335–1347, December 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kaushik:1994:ACD

[KHJS94] S. D. Kaushik, C.-H. Huang, R. W. Johnson, and P. Sadayappan. An approach to communication-efficient data redistribution. In ACM [ACM94a], pages 364–373.

ISBN 0-89791-665-4. LCCN

????

Kaushik:1995:MAR

[KHRS95] S. D. Kaushik, C.-H. Huang, J. Ramanujam, and P. Sadayappan. Multi-phase array redistribution: modeling and evaluation. In IEEE [IEE95a], pages 441–445. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

Kaushik:1995:IGI

[KHS95] S. D. Kaushik, C.-H. Huang, and P. Sadayappan. Incremental generation of index sets for array statement execution on distributed-memory machines. In Pingali et al. [PBG^{+}95], pages 251–265. ISBN 3-540-58868-X. LCCN QA76.58 .W656 1994.

Kaushik:1996:EIS

[KHS96] S. D. Kaushik, C.-H. Huang, and P. Sadayappan. Efficient index set generation for compiling HPF array statements on distributed-memory machines. Journal of Parallel and Distributed Computing, 38(2):237– 247, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jpdc.1996.0144/production;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0144/production/pdf.

Krogh:2017:RAF

[KHS17] Fred T. Krogh, Richard J. Hanson, and Philip W. Sharp. Remark on Algorithm 936: a Fortran Message Processor. ACM Transactions on Mathematical Software, 43(4):39:1, March 2017. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Kro14].

Kikuchi:1993:PAS

[Kik93] S. Kikuchi. Parallelization assist system. Joho-Shori (J. Information Processing Soc. Japan), 34(9):1158–1169, September 1993. CODEN JOSHA4. ISSN 0447-8053.

King:1992:APF

[Kin92] Roger Duane King. Automated parallelization of FORTRAN code with dynamic load balancing. Thesis (M.S.), George Mason University, Fairfax, VA, USA, 1992. iv + 103 pp.

King:1993:HPL

[Kin93] K. N. King. The history of programming languages. Dr. Dobb’s Journal of Software Tools, 18(8):18–??, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Kirkup:1993:FCE

[Kir93] S. M. Kirkup. Fortran codes for the evaluation of the discrete Helmholtz integral operators. Advances in boundary element techniques, pages 251– 281, 1993.

Kirkup:1998:FCC

[Kir98]  S. M. Kirkup.  Fortran

codes for computing the discrete Helmholtz integral operators. Advances in Computational Mathematics, 9(3–4): 391–409, November 1998. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A:1018953910353.

Kirby:2002:FSC

[Kir02] J. F. Kirby. A FORTRAN 90 subroutine to calculate array sizes prior to a mixed-radix fast Fourier transform. Computers and Geosciences, 28(8):999– 1001, October 2002. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Kouremenos:1990:TNV

[KK90] D. A. Kouremenos and R. N. Krikkis. Technical note — A VMS/Fortran 77 program to compute the normal shock waves of refrigerant R22. Advances in engineering software, 12(3):139–??, July 1990. CO-DEN AESODT. ISSN 0141-1195, 0965-9978.

Kornkven:1994:EIH

[KK94] E. Kornkven and L. V. Kale. Efficient implementation of High Performance Fortran via adaptive scheduling — an overview. In Kumar [Kum94], pages 136–141. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Kennedy:1995:ADL

[KK95a] Ken Kennedy and Ulrich Kremer. Automatic data layout for High Performance Fortran. In Karin [Kar95], page ?? ISBN 0-89791-816-9. LCCN QA 76.88 S8573 1995a.

URL http://www.supercomp.
org/sc95/proceedings/580_
UKRE/SC95.HTM. These pro-

ceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Konda:1995:SFD

[KK95b] Venkat Konda and Anup Kumar. A systematic framework for the dependence cycle removal in practical loops. Journal of Parallel and Distributed Computing, 27(2): 157–171, June 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1995.1079/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1995.1079/production/pdf.

Kennedy:1998:ADL

[KK98] Ken Kennedy and Ulrich Kremer. Automatic data layout for distributed-memory machines. ACM Transactions on Programming Languages and Systems, 20(4): 869–916, July 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
toplas/1998-20-4/p869-kennedy/

.

Kennedy:2001:CHP

[KK01] Ken Kennedy and Charles Koelbel. Chapter 1. High Performance Fortran 2.0. Lecture Notes in Computer Science, 1808:3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/1808/18080003.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1808/18080003.
pdf.

Kuiper:2010:FPC

[KKH10] Rebecca M. Kuiper, Irene Klugkist, and Herbert Hoijtink. A Fortran 90 program for confirmatory analysis of variance. Journal of Statistical Software, 34(8):??, May 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v34/i08.

Kobayashi:1995:FPN

[KKK95] R. Kobayashi, M. Konuma, and S. Kumano. FORTRAN program for a numerical solution of the nonsinglet Altarelli-Parisi equation. Computer Physics Communications, 86 (3):264–278, May 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Konovalav:1995:FDL

[KKMP95a] N. A. Konovalav, V. A.

Kryukov, S. N. Mikhailov, and A. A. Pogrebtsov. Fortran DVM: a language for portable parallel program development. Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 21 (1):35–38, January–February 1995. CODEN PROGD3. ISSN 0132-3474, 0361-7688.

Konovalov:1995:FDL

[KKMP95b] N. A. Konovalov, V. A.

Kryukov, S. N. Mikhailov, and A. A. Pogrebtsov. Fortran DVM: a language for portable parallel program development. Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 21 (1):35–??, 1995. CODEN PCSODA. ISSN 0361-7688 (print), 1608-3261 (electronic).

Kamachi:1995:HCP

[KKS^{+}95] Tsunehiko Kamachi, Kazuhiro Kusano, Kenji Suehiro, Yoshiki Seo, Masanori Tamura, Shoichi Sakon, Yukimitsu Watanabe, and Yukimasa Shiroto. HPF compiler for parallel computers: implementation and performance evaluation on Cenju-3. Nippon Electric Company research and development, 36(2): 325–334, April 1995. CODEN NECRAU. ISSN 0048-0436.

Kohler:1999:FCS

[KKY99] H. S. Köhler, N. H. Kwong, and Hashim A. Yousif. A Fortran code for solving the kadanoff-baym equations for a homogeneous fermion system. Computer Physics Communications, 123(1–3): 123–142, December 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S001046559900260X.

Kennedy:2011:RFH

[KKZ11] Ken Kennedy, Charles Koelbel, and Hans Zima. The rise and fall of High Performance Fortran. Comm. ACM, 54(11): 74–82, November 2011. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kleinrubatscher:1994:FPS

[KKZG94] Paul Kleinrubatscher, Albert Kriegshaber, Robert Zöchling, and Robert Glück. Fortran program specialization. In U. Meyer and G. Snelting, editors, Workshop Semantikgestützte Analyse, Entwicklung und Generierung von Programmen, pages 45–54. Justus-Liebig-Universität, Giessen, Germany, 1994. Report No. 9402.

Kleinrubatscher:1995:FPS

[KKZG95] Paul Kleinrubatscher, Albert Kriegshaber, Robert Zöchling, and Robert Glück. Fortran program specialization. ACM SIGPLAN Notices, 30(4):61– 70, April 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kumano:1992:FPN

[KL92] S. Kumano and J. T. Londergan. A FORTRAN program for numerical solution of the Altarelli-Parisi equations by the Laguerre method. Computer Physics Communications, 69(2–3):373–396, March/April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Koo:1995:CMO

[KLA95] H.-J. Koo, K.-S. Lee, and W.- S. Ahn. Crystal molecular orbital calculation of the lanthanum nickel oxide by means of the Microsoft Fortran. Taehan hwahakhoe chi, 39(9):685– ??, 1995. CODEN DHWHAB. ISSN 0418-2472.

Klein:1993:NRM

[Kle93] W. U. Klein. Numerical reliability of MHD flow calculations. In Adams and Kulisch [AK93], pages 397–421. ISBN 0-12-044210-8. LCCN QA76.S368 1993.

Karlovsky:1991:ANF

[KLM91] S. Karlovsky, D. Leibfritz, and D. Malon. The Argonne–NAG Fortran 90 initiative. Fortran Journal, ??(??):??, ???? 1991. ISSN 1060-0221.

Kim:2000:OOC

[KLM00] Y.-H. Kim, I.-H. Lee, and R. M. Martin. Object-oriented construction of a multigrid electronic-structure code with Fortran 90. Computer Physics Communications, 131(1):10– 25, September 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Kumar:2019:FOP

[KLM^{+}19] Ramavarmaraja Kishor Kumar, Vladimir Loncar, Paulsamy Muruganandam, Sadhan K. Adhikari, and Antun Balaz. C and Fortran OpenMP programs for rotating Bose– Einstein condensates. Computer Physics Communications, 240(??):74–82, July 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465519300827.

Knecht:1990:PQDb

[KLN90] Siegfried Knecht, Edwin Laermann, and Wolfgang E. Nagel. Parallelizing QCD with dynamical fermions on a Cray multiprocessor system. Parallel Computing, 15(1–3):3–20, September 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Koelbel:1994:HPF

[KLS^{+}94a] Charles H. Koelbel, David B.

Loveman, Robert S. Schreiber, Guy L. Steele Jr., and Mary E. Zosel. The High Performance Fortran handbook. Scientific and engineering computation. MIT Press, Cambridge, MA, USA, January 1994. ISBN 0-262-11185-3, 0-262-61094-9 (paperback). xiv + 329 pp. LCCN QA76.73.F25 H53 1994. US$32.50. URL http:

//www.mitpress.com/book-
home.tcl?isbn=0262610949.

Kuhn:1994:KPD

[KLS94b] Robert H. Kuhn, Bruce Leasure, and Sanjiv M. Shah. The KAP parallelizer for DEC Fortran and DEC C programs. Digital Technical Journal, 6(3): 57–70, Summer 1994. CODEN DTJOEL. ISSN 0898-901X.

URL ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v6n3/The_KAP_Parallelizer_
for_DEC_01jan1995DTJF05P8.
ps;   http://www.digital.
com:80/info/DTJF05/DTJF05SC.
TXT.

Kaagstrom:1998:GLB

[KLV98] Bo Kågström, Per Ling, and Charles Van Loan. GEMMbased level 3 BLAS: highperformance model implementations and performance evaluation benchmark. ACM Transactions on Mathematical Software, 24(3):268–302, September 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Knobe:1993:OTS

[KLW93] K. Knobe, J. D. Lukas, and M. Weiss. Optimization techniques for SIMD Fortran compilers. Concurrency, practice and experience, 5(7):527– 552, October 1993. CODEN CPEXEI. ISSN 1040-3108.

Koonin:1990:CPF

[KM90] Steven E. Koonin and Dawn C. Meredith. Computational Physics: Fortran Version. Addison-Wesley, Reading, MA, USA, January 1, 1990. ISBN 0-201-12779-2. xvi + 639 pp. LCCN QC20.7.E4 K66 1990. US$59.25; US$72.25.

                                        URL http://www.cbooks.
                                        com/sqlnut/SP/search/gtsumt?
                                        source=&isbn=0-201-12779-
                                        2.

Kelsey:1997:PSE

[KM97] Charles T. Kelsey, IV and Gary L. Masters. Porting scientific and engineering programs to Linux. Linux Journal, 39: ??, July 1997. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL

                                        ftp://ftp.ssc.com/pub/lj/
                                        listings/issue39/2177.tgz.

Kees:1999:CIN

[KM99] Christopher E. Kees and Cass T. Miller. C++ implementations of numerical methods for solving differentialalgebraic equations: design and optimization considerations. ACM Transactions on Mathematical Software, 25(4): 377–403, December 1999. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kim:1996:PSS

[KMBK96] Taegeun Kim, Kyeongdeok Moon, Nanjoo Ban, and Jungkwon Kim. PPTran: Source to source translator for High Performance Fortran. Parallel Algorithms and Applications, 9(3–4):213–225, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://www.

informaworld.com/smpp/content~
content=a778707980.

Karp:2023:LSD

[KMJ^{+}23] Martin Karp, Daniele Massaro, Niclas Jansson, Alistair Hart, Jacob Wahlgren, Philipp Schlatter, and Stefano Markidis. Large-scale direct numerical simulations of turbulence using GPUs and modern Fortran. The International Journal of High Performance Computing Applications, 37(5): 487–502, September 1, 2023. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/10943420231158616.

Koonin:1992:CM

[KMmYsK92] Steven E. Koonin, Dawn Meredith, Chong man Yang, and Chae sam Kim. Chonsan mullihak. Naeil ul yonun chaek. Ihwa Yoja Taehakkyo Chulpanbu, Soul Tukpyolsi, Korea, 1992. ISBN ???? xix + 643 pp. LCCN ????

Kennedy:1996:OFS

[KMR96] K. Kennedy, J. Mellor-Crummey, and G. Roth. Optimizing Fortran 90 shift operations on distributed-memory multicomputers. Lecture Notes in Computer Science, 1033:161– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamachi:1997:KPH

[KMR^{+}97] T. Kamachi, A. Muller, R. Ruhl, Y. Seo, K. Suehiro, and M. Tamura. Kemari: a portable High Performance Fortran system for distributed memory parallel processors. Scientific Programming, 6(1):41–58, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kimelman:1995:VEH

[KMS^{+}95] D. Kimelman, P. Mittal, E. Schonberg, P. F. Sweeney, Ko-Yang Wang, and D. Zernik. Visualizing the execution of High Performance Fortran (HPF) programs. In IEEE [IEE95a], pages 750–759. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

Kennedy:1991:IPP

[KMT91] Ken Kennedy, Kathryn S.

McKinley, and Chau-Wen Tseng. Interactive parallel programming using the ParaScope Editor. IEEE Transactions on Parallel and Distributed Systems, 2(3):329–341, July 1991. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kearfott:1994:FSS

[KN94] R. B. Kearfott and S. Ning.

FORTDIFF: a set of subroutines for Fortran-to-Fortran differentiation of programs, 1994.

Kalns:1995:DPD

[KN95] E. T. Kalns and L. M.

Ni. DaReL: a portable data redistribution library for distributed-memory machines. In IEEE [IEE95b], pages 78– 87. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Kearfott:2004:LTI

[KNOR04] R. Baker Kearfott, Markus Neher, Shin’ichi Oishi, and Fabien Rico. Libraries, tools, and interactive systems for verified computations four case studies. In Alt et al. [AFKL04]. ISBN 3-540-21260-4 (paperback).

Kennedy:1995:LAC

[KNS95a] K. Kennedy, N. Nedeljkovic, and A. Sethi. A lineartime algorithm for computing the memory access sequence in data-parallel programs. In Anonymous [Ano95c], pages 102–111. CODEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN ????

Kennedy:1995:LTA

[KNS95b] Ken Kennedy, Nenad Nedeljkovic, and Ajay Sethi. A lineartime algorithm for computing the memory access sequence in data-parallel programs. ACM SIGPLAN Notices, 30(8):102–111, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Knuth:2003:SPC

[Knu03] Donald E. Knuth. Selected Papers on Computer Languages. CSLI Publications, Stanford, CA, USA, 2003. ISBN 1-57586-381-2 (hardback), 1-57586-382-0 (paperback). xvi + 594 pp. LCCN QA76.7 .K63 2002. US$75.00 (hardback), US$35.00 (paperback).

Kincaid:1990:RVP

[KO90] David R. Kincaid and Thomas C. Oppe. Recent vectorization and parallelization of IT-PACKV. In O. Axelsson and L. Y. Kolotilina, editors, Preconditioned conjugate gradient methods (Nijmegen, 1989), volume 1457 of Lecture Notes in Mathematics, pages 58–78 (of 196). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-53515-2. LCCN QA3 .L35 v.1457.

King:1991:FLS

[KO91] K. N. King and A. Jefferson Offutt. A Fortran language system for mutation-based software testing. Software—Practice and Experience, 21(7):685– 718, July 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Khajah:1994:UHP

[KO94] H. G. Khajah and E. L.

Ortiz. Ultra-high precision computations. Computers and Mathematics with Applications, 27(7):41–57, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0898122194901481.

Kodama:2008:ASP

[Kod08] Masao Kodama. Algorithm 877: a subroutine package for cylindrical functions of complex order and nonnegative argument. ACM Transactions on Mathematical Software, 34(4): 22:1–22:21, July 2008. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kodama:2011:AMC

[Kod11] Masao Kodama. Algorithm 912: a module for calculating cylindrical functions of complex order and complex argument. ACM Transactions on Mathematical Software, 37 (4):47:1–47:25, February 2011. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Koelbel:1992:OHP

[Koe92] Charles Koelbel. An overview of High Performance Fortran. ACM SIGPLAN FORTRAN Forum, 11(4):9–??, December 1, 1992. ISSN 1061-7264 (print), 1931-1311 (electronic).

Koikari:2009:ABS

[Koi09] Souji Koikari. Algorithm 894: On a block Schur–Parlett algorithm for φ-functions based on the sep-inverse estimate. ACM Transactions on Mathematical Software, 36(2):12:1– 12:20, March 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Knies:1993:HPF

[KOM93] A. Knies, Michael (Michael A.) O’Keefe, and T. MacDonald. High Performance Fortran: a practical analysis. AHPCRC preprint 93-107, Army High Performance Computing Research Center, Minneapolis, MN, USA, 1993. ii + 18 pp.

Knies:1994:HPF

[KOM94] Allan Knies, Matthew O’Keefe, and Tom MacDonald. High Performance Fortran: a practical analysis. Scientific Programming, 3(3):187–199, Fall 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kondapaneni:1992:VTF

[Kon92] Prasanna Kumar Kondapaneni. A visual tool for Fortran D programming. Thesis (M.S.), Auburn University, Auburn, AL, USA, 1992. ix + 61 pp.

Konigsberg:1994:NRF

[Kon94] Lyle W. Konigsberg. Numerical Recipes in Fortran: The Art of Scientific Computing, by W. H. Press et al., and Numerical Recipes Example Book (Fortran), by W. T. Vetterling et al. Human biology, 66(2): 349–350, April 1994. CODEN HUBIAA. ISSN 0018-7143.

Koniges:2000:ISP

[Kon00] Alice E. Koniges, editor. Industrial Strength Parallel Computing. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-540-1. xxv + 597 pp. LCCN QA76.58 .I483 2000.

Koopman:1990:WSW

[Koo90] Adri Koopman. Who’s on the system and what’s he doing? The VAX professional, 12(3): 34–38, June 1990. CODEN VAXPEN. ISSN 8750-9628.

Kornbluh:1999:MSS

[Kor99] Ken Kornbluh. Math and science software [technology 1999 analysis and forecast]. IEEE Spectrum, 36(1):88–91, January 1999. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Kumar:1991:PTF

[KP91] S. P. Kumar and I. R. Philips. Portable tools for Fortran parallel programming. Concurrency, practice and experience, 3(6):559–572, December 1991. CODEN CPEXEI. ISSN 1040-3108.

Karanovic:1992:FPC

[KP92] Lj. Karanovic and D. Poleti. A FORTRAN program for conversion of PC-APD data files into ASCII files. Powder Diffraction, 7(3):179, September 1992. CODEN PODIE2. ISSN 0885-7156.

Krishnamurthy:1993:DPE

[KP93] E. V. Krishnamurthy and Chen Pin. Data parallel evaluationinterpolation algorithm for polynomial matrix inversion. Parallel Computing, 19(5):577– 589, May 1993. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kremer:1994:COR

[KR94] Ulrich Kremer and Marcelo Ramé. Compositional oil reservoir simulation in Fortran D: A feasibility study on Intel iPSC/860. Int. J. Supercomputer Appl., 8(2):119– 128, Summer 1994. CODEN IJSAE9. ISSN 0890-2720. See erratum [KR95].

Kremer:1995:ECO

[KR95] U. Kremer and M. Ramé. Erratum: Compositional Oil Reservoir Simulation in Fortran D: A Feasibility Study on Intel iPSC/860. The International Journal of Supercomputer Applications and High Performance Computing, 9(1): 71, Spring 1995. CODEN IJSAE9. ISSN 0890-2720. See [KR94].

Kraft:1994:ATF

[Kra94] Dieter Kraft. Algorithm 733:

TOMP — Fortran modules for optimal control calculations. ACM Transactions on Mathematical Software, 20(3):262– 281, September 1994. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kinzel:1990:CEP

[KRB^{+}90] G. L. Kinzel, S. M. Rohde, D. W. Bennett, et al., editors. Computers in engineering, 1990: proceedings of the 1990 ASME International Computers in Engineering Conference and Exposition, August 5–9, Boston, Massachusetts. American Society of Mechanical Engineers, 345 E. 47th St., New York, NY 10017, USA, 1990. ISBN 0-7918-0515-8. LCCN TA 345 A86 1990a. Two volumes.

Kaur:2021:FFP

[KRG21] Pardeep Kaur, Arko Roy, and Sandeep Gautam. FORTRESS:

FORTRAN programs for solving coupled Gross–Pitaevskii equations for spin-orbit coupled spin-1 Bose–Einstein condensate. Computer Physics Communications, 259(??):Article 107671, February 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S001046552030326X.

Krishnamoorthy:1986:BRB

[Kri86] Mukkai S. Krishnamoorthy. Book review: FORTRAN-Scientific Subroutine Library (by Peerless Engineering Service). SIAM Review, 28(1): 111–112, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Krommes:1990:KVA

[Kro90] John Krommes. FWEB (Krommes) vs. FWEB (Avenarius and Oppermann). TEXhax, 90(19), February 1990.

Krogh:2014:AFM

[Kro14] Fred T. Krogh. Algorithm 936: a Fortran message processor. ACM Transactions on Mathematical Software, 40(2): 15:1–15:4, February 2014. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [KHS17].

Kruessel:1990:EID

[Kru90a] Manfred Kruessel. Entwurf und Implementierung eines Dialogsystems fuer das Erstellen von Wagenauslauflisten. Studienarbeit, 1990.

Kruger:1990:EFP

[Kru90b] Anton Kruger. Efficient Fortran Programming. Wiley professional computing. John Wiley and Sons, New York, London, Sydney, September 1990. ISBN 0-471-52894-3. x + 184 pp. LCCN QA76.73.F25 K78 1990. US$46.95; US$34.95.

URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0-471-52894-
3.

Kouremenos:1990:TNF

[KRY90] D. A. Kouremenos, C. D.

Rakopoulos, and E. A. Yfantis. Technical note — A FOR-TRAN program for calculating the thermodynamic and transport properties of diesel fuel. Advances in engineering software, 12(4):190–??, October 1990. CODEN AESODT. ISSN 0141-1195, 0965-9978.

Krysl:1994:FFL

[Kry94] P. Krysl. FLIPP: Fortran library for interactive persistent programming. Engineering Computations, 11(3):213– 226, June 1994. CODEN EN-COEN. ISSN 0264-4401.

Klieme:1990:EFP

[KS90] Eckhard Klieme and Heinrich Stumpf. EVAL: A FORTRAN program for the evaluation of selection procedures. Educational and Psychological Measurement, ??(1):127–??, Spring 1990. CODEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Kennedy:2002:SIH

[KS02] K. Kennedy and Y. Seo. Special issue: High Performance Fortran. Concurrency and Computation: Practice and Experience, 14(8–9): 551–553, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/95016134/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016134{\&}PLACEBO=IE.
pdf.

Kondayya:2012:FHF

[KS12] Gundra Kondayya and Alok Shukla. A Fortran 90 Hartree–Fock program for one-dimensional periodic πconjugated systems using Pariser–Parr–Pople model. Computer Physics Communications, 183(3):677–689, March 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465511003687.

Krumbein:1995:CCT

[KSM95] William C. Krumbein, Wolfgang Scherer, and Daniel F. Merriam. CORSURF: a covariance-matrix trend-analysis FORTRAN IV computer program. Computers and Geosciences, 21(9):1065–??, 1995. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Kusters:1993:PJI

[KSW93] H. Kusters, E. Stein, and W. Werner, editors. Proceedings of the Joint International Conference on Mathematical Methods and Supercomputing in Nuclear Applications: M & C + SNA ’93, April 19–23, 1993, Congress and Exhibition Centre, Karlsruhe, Germany. Kernforschungzentrum Karlsruhe, Karlsruhe, Germany, 1993. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Kholmurodov:2000:HVL

[KSYE00] Kholmirzo Kholmurodov, William Smith, Kenji Yasuoka, and Toshikazu Ebisuzaki. A highly vectorised “link-cell” FOR-TRAN code for the DL POLY molecular dynamics simulation package. Computer Physics Communications, 125 (1–3):167–192, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465599004853.

Kamel:1990:LSC

[KSZ90] A. Kamel, P. Sguazzero, and V. Zecca. Large scale computing on clustered vector multiprocessors. In IEEE [IEE90a], pages 418–427. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Kennedy:1994:CSM

[KT94] Ken Kennedy and Kevin Timson. Centers of supercomputing — making parallel computing truly usable: research, education, and knowledge transfer at the Center for Research on Parallel Computation. The International Journal of Supercomputer Applications and High Performance Computing, 8(2):73–79, Summer 1994. CO-DEN IJSAE9. ISSN 0890-2720.

Keppens:2000:UHP

[KT00] R. Keppens and G. Toth. Using high performance Fortran for magnetohydrodynamic simulations. Parallel Computing, 26(6):705–722, May 2000. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kaufman:2002:AFP

[KTMB02] E. H. Kaufman, G. D. Taylor, P. W. Mielke, and K. J. Berry. An algorithm and Fortran program for multivariate LAD ℓ (ℓ 1 of ℓ 2) regression. Computing, 68(3):275–287, 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

URL http://link.springer.
de/link/service/journals/
00607/bibs/2068003/20680275.
htm;   http://link.springer.
de/link/service/journals/
00607/papers/2068003/20680275.
pdf.

Kang:2024:IBM

[KTP^{+}24] Hyun-Gyu Kang, Raymond S.

Tuminaro, Andrey Prokopenko, Seth R. Johnson, Andrew G. Salinger, and Katherine J. Evans. An implicit barotropic mode solver for MPAS-ocean using a modern Fortran solver interface. The International Journal of High Performance Computing Applications, 38 (3):175–191, May 1, 2024. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https://

journals.sagepub.com/doi/
abs/10.1177/10943420231205601.

Kubo91a

[Kub91a] Koichi Kubota. PADRE2, A FORTRAN precompiler yielding error estimates and second derivatives. In Andreas Griewank and George F. Corliss, editors, Automatic Differentiation of Algorithms: Theory, Implementation, and Application. SIAM, Philadelphia, PA, 1991.

Kubota:1991:PAF

[Kub91b] Koichi Kubota. PADRE2, A FORTRAN precompiler yielding error estimates and second derivatives. In Andreas Griewank and George F. Corliss, editors, Automatic Differentiation of Algorithms: Theory, Implementation, and Application, pages 251–262. SIAM Press, Philadelphia, PA, USA, 1991.

Kubota:1991:PFP

[Kub91c] Koichi Kubota. PADRE2, A FORTRAN precompiler yielding error estimates and second derivatives. In Andreas Griewank and George F. Corliss, editors, Automatic Differentiation of Algorithms: Theory, Implementation, and Application, page ?? SIAM Press, Philadelphia, PA, USA, 1991.

Kugendran:1992:ISP

[Kug92] Thurai Kugendran. Introduction to scientific programming with FORTRAN. Wm. C. Brown Publishers, Dubuque, IA, USA, 1992. ISBN 0-697-16820-4. iii + [152] pp.

Kulisch:1995:NVA

[Kul95] Ulrich Kulisch. A new vector arithmetic coprocessor chip for the PC. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.nsc.liu.se/~boein/
                                        ifip/kyoto/workshop-info/
                                        proceedings/kulisch/kulisch1.

html. The chip is the world’s first hardware implementation f the “GAMM/IMACS Proposal for Accurate Fl;ating-Point Vector Arithmetic”. It runs on any PC with a PCI bus.

Kumar:1994:PPI

[Kum94] V. K. Prasanna Kumar, editor. Parallel processing: 1st IWWP: proceedings of the First International Workshop on Parallel Processing (IWPP-94), December 26–31, 1994, Bangalore, India. Tata McGraw-Hill Pub. Co, New Delhi, India, 1994. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Kutasov:1992:PPR

[Kut92] I. M. Kutasov. Program predicts reservoir temperature and geothermal gradient. The Oil and Gas Journal, 90(22):85–86, June 1992. CODEN OIGJAV. ISSN 0030-1388.

Krishchuk:1992:IOF

[KVK92] Vladimir N. Krishchuk, Nikolay M. Vasilega, and Galina L. Kozina. Interval operations and functions library for FOR-TRAN 77 programming system and its practice using. Interval ’92 (Moscow, 1992). Interval Computations, 4:2–8, 1992. ISSN 0135-4868.

Kim:1994:CAM

[KW94] I. Kim and M. Wolfe. Communication analysis for multicomputer compilers. In Cosnard et al. [CGS94], pages 101–109. CODEN ITATEC. ISBN 0-444-81926-6. ISSN 0926-5473. LCCN QA76.58 .I46 1994.

Kim:1994:TPS

[KY94] Dongseung Kim and Byung-Guoen Yi. A two-pass scheduling algorithm for parallel programs. Parallel Computing, 20 (6):869–885, June 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kasahara:1998:DCS

[KY98a] Hironori Kasahara and Akimasa Yoshida. Data-localization compilation scheme using partial-static task assignment for Fortran coarse-grain parallel processing. Parallel Computing, 24(3-4):579–596, May 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kasahara:1998:DLC

[KY98b] Hironori Kasahara and Akimasa Yoshida. A datalocalization compilation scheme using partial-static task assignment for Fortran coarsegrain parallel processing. Parallel Computing, 24(3–4):579– 596, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

                                        URL http://www.elsevier.
                                        com/cas/tree/store/parco/
                                        sub/1998/24/3-4/1303.pdf.

Kumar:2015:FPT

[KYSV^{+}15] R. Kishor Kumar, Luis E. Young-S., Dusan Vudragović, Antun Balaz, Paulsamy Muruganandam, and S. K. Adhikari. Fortran and C programs for the time-dependent dipolar Gross–Pitaevskii equation in an anisotropic trap. Computer Physics Communications, 195(??):117–128, October 2015. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465515001344.

Kuester:1994:IFF

[KZ94a] U. Kuester and M. Zuern. Influence of Fortran 90 features on performance on Cray vector computer systems. In Gentzsch and Harms [GH94b], pages 475–483. ISBN 3-540-57981-8 (Berlin: vol. 2: paperback), 0-387-57981-8 (New York: vol. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1– 2 (c1994). DM96.00. Two volumes.

Kuster:1994:IFF

[KZ94b] U. Kuster and M. Zurn. Influence of Fortran 90 features on performance on Cray vector computer systems. Lecture Notes in Computer Science, 797:475–483, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lahey:1990:LPF

[Lah90] Lahey Computer Systems, Inc.

Lahey personal Fortran 77 reference manual. Lahey Computer Systems, Incline Village, NV, USA, revision A, June 1990 edition, 1990. xvi + x + 390 pp.

Lai:1992:FSB

[Lai92a] Ming Jun Lai. Fortran subroutines for B-nets of box splines on three- and four-directional meshes. Numerical Algorithms, 2(1):33–38, 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lai:1992:FSN

[Lai92b] Ming Jun Lai. Fortran subroutines for B-nets of box splines on three- and fourdirectional meshes. Numerical Algorithms, 2(1):33–38, February 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Langhorne:1990:RIA

[Lan90a] David B. Langhorne. The retargeting of image algebra Fortran to special-purpose architectures. Thesis (M.Eng.), University of Florida, Gainesville, FL, USA, 1990. viii + 104 pp.

LPI:1990:L

[Lan90b] Language Processors, Inc. LPI-Fortran. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-329392-0 (paperback), 0-13-329384-X. various pp.

LPI:1990:LF

[Lan90c] Language Processors, Inc. LPI-Fortran. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-329384-X.

LPI:1990:LFL

[Lan90d] Language Processors, Inc.

LPI-Fortran: Language Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, January 1990. ISBN 0-13-329392-0 (paperback), 0-13-329384-X. various pp. LCCN QA76.73.F25L66 1990. US$45.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0-13-329392-0.

LPI:1990:LLR

[Lan90e] Language Processors, Inc. LPI-FORTRAN language reference manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-329392-0 (paperback), 0-13-329384-X (LPI). various pp. LCCN QA76.73.F25L65 1990.

Lanahan:1993:GIP

[Lan93a] Robert J. Lanahan, Jr. Graphics interface (with PHIGS) for FORTRAN programmers on DEC workstations. Thesis (M.S.), San Jose State University, San Jose, CA, USA, 1993. 90 pp.

Langer:1993:PF

[Lan93b] Erasmus Langer. Programmieren in Fortran. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., March 1993. ISBN 0-387-82446-4. US$30.00. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=0387824464.

Lang:2001:SCC

[Lan01] Duncan Temple Lang. In search of C/C++ & FOR-TRAN routines. R News: the Newsletter of the R Project, 1(3):20–23, September 2001. CODEN ???? ISSN 1609-

                                        3631. URL http://CRAN.R-
                                        project.org/doc/Rnews/.

Laplante:1996:GPC

[Lap96] Phillip Laplante, editor. Great Papers in Computer Science. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-314-06365-X (paperback), 0-7803-1112-4 (hardcover). iv + 717 pp. LCCN QA76 .G686 1996. URL http:

                                        //bit.csc.lsu.edu/~chen/
                                        GreatPapers.html. Contains

a reprint of the first published paper on Fortran [BBB^{+}57].

Larsen:1993:GFP

[Lar93]  Rune B. Larsen.  “GE-

OFLUID”: a FORTRAN 77 program to compute chemical properties of gas species in C-O-H fluids. Computers and Geosciences, 19(9):1295– ??, October 1993. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Lasecki:1997:PFF

[Las97] Jennifer L. Lasecki. S-Plus functions and Fortran subroutines for multiple comparison critical points. Thesis (M.S.), University of South Carolina, Columbia, SC, USA, 1997. v + 68 pp.

Lavaud:1991:ASE

[Lav91] M. Lavaud. A^{S} TEX: a software environment on PC adapted to scientific research. In Glowinski [Glo91b], pages 779–788.

Lawrence:2001:CVF

[Law01]  Norman Lawrence.  Com-

paq Visual Fortran: a guide to creating Windows applications. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2001. ISBN 1-55558-249-4. 488 (est.) pp. LCCN QA76.73.F25 L375 2001. US$49.95. URL http:

//www.compaq.com/fortran/
visual/lawrence-book.html.

Luecke:2003:MCT

[LCC^{+}03] Glenn Luecke, Hua Chen, James Coyle, Jim Hoekstra, Marina Kraeva, and Yan Zou. MPI-CHECK: a tool for checking Fortran 90 MPI programs. Concurrency and Computation: Practice and Experience, 15(2):93–100, February 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Levine:1991:CSAb

[LCD91] D. Levine, D. Callahan, and J. Dongarra. A comparative study of automatic vectorizing compilers. Parallel Computing, 17(10–11):1223– 1244, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.netlib.
org/utk/people/JackDongarra/
PAPERS/Comparative-Study-
of-Automatic-Vectorizing-
Compilers.pdf.

LePera:1987:FVO

[LD87] T. A. LePera and H. E. Donley. FORTRAN: a viable option for the introductory computer science course. ACM SIGPLAN Notices, 22(3):102–110, March 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liska:1990:FRP

[LD90] R. Liska and L. Drska. FIDE: A REDUCE package for automation of FInite difference method for solving pDE. In Watanabe and Nagata [WN90], pages 169–176. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

URL http://www.acm.org:
80/pubs/citations/proceedings/
issac/96877/p169-liska/.

Lachanas:1998:ECG

[LE98] A. Lachanas and P. Evripidou. Exploiting course grain parallelism from FORTRAN by mapping it to IF1. Lecture Notes in Computer Science, 1470:463–??, 1998. CODEN LNCSD9. ISSN 0302-9743.

Lee:1990:HSF

[Lee90] Jinun-Chyi Lee. A high speed Fortran to C translator. Thesis (M.S. in computer science), Southern Methodist University, Dallas, TX, USA, 1990. x + 210 pp.

Lee:1997:CFF

[Lee97] J. A. N. Lee. Celebrating Fortran’s fortieth. Computer, 30 (4):8, April 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leffelaar:1993:SAS

[Lef93] P. A. Leffelaar, editor. On systems analysis and simulation of ecological processes: with examples in CSMP and FORTRAN, volume 1 of Current Issues in Production Ecology. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, August 1993. ISBN 0-7923-2434-X (hardcover). xiii + 294 pp. LCCN QH541.15.S5O43 1993. US$118.50. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=079232434X.

Lemay:1993:CPFa

[Lem93a] Gerald J. Lemay. C for Pascal and Fortran Programmers. P.S. Melvil Press, P.O. Box 973, Portsmouth 02871, RI, USA, January 1993. ISBN 1-883496-00-4. vi + 114 pp. LCCN ???? US$25.00. URL http:

//www.cbooks.com/sqlnut/

                                        SP/search/gtsumt?source=&
                                        isbn=1883496004.

Lemay:1993:CPFb

[Lem93b] Gerald J. Lemay. C for Pascal and Fortran Programmers. P.S. Melvil Press, P.O. Box 973, Portsmouth 02871, RI, USA, January 1993. ISBN 1-883496-01-2. US$35.00. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=1883496012.

Lemay:1993:CPFd

[Lem93c] Gerald J. Lemay. C for Pascal and Fortran Programmers. P.S. Melvil Press, P.O. Box 973, Portsmouth 02871, RI, USA, January 1993. ISBN 1-883496-02-0. US$15.00. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=1883496020.

Lemay:1993:CPFc

[Lem93d] Gerald J. Lemay. C for Pascal and Fortran programmers: workbook. P.S. Melvil Press, P.O. Box 973, Portsmouth 02871, RI, USA, 1993. ISBN 1-883496-02-0 (paperback). iii + 130 pp.

Leonard:1991:FF

[Leo91] Simon Leonard. Fortran five. Malice Aforethought, ????, 1991. ISBN 1-871197-02-3. 150 pp.

Leva:1992:FNR

[Lev92] Joseph L. Leva. A fast normal random number generator.

ACM Transactions on Mathematical Software, 18(4):449– 453, December 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
toms/1992-18-4/p449-leva/.

Levesque:1994:APR

[Lev94] J. M. Levesque. Applied Parallel Research’s xHPF system. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):71, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Levelt:1995:IPI

[Lev95a] A. H. M. (Antonius Henricus Maria) Levelt, editor. IS-SAC’95: proceedings of the 1995 International Symposium on Symbolic and Algebraic Computation, July 10– 12, 1995, Montréal, Canada, ISSAC — Proceedings. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-699-9. LCCN A 76.95 I59 1995.

Levy:1995:IOF

[Lev95b] G. F. Levy. Improving the output of the FORTRAN to C translator, f2c. Software— Practice and Experience, 25(2): 217–227, February 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Levy:1997:USH

[Lev97] G. F. Levy. Using <stdio.h> in the output of the FORTRAN to C translator, f2c. Software —Practice and Experience, 27(12):1369–1384, December 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        7276;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=7276&PLACEBO=IE.pdf.

Levin:1998:BRN

[Lev98] Mikhail P. Levin. Book review: Numerical Recipes In Fortran 90: The Art Of Parallel Scientific Computing. IEEE Concurrency, 6(4):79, October/December 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1998/pdf/
                                        p4078.pdf.

Lewin:1994:FDR

[Lew94] David I. Lewin. Fortran’s developer receives top engineering award. Computers in physics, 8(1):6–??, January/February 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

                                        aip.scitation.org/doi/10.
                                        1063/1.4823261.

Lujan:2000:OOO

[LFG00] Mikel Luján, T. L. Freeman, and John R. Gurd. OoLALA: an object oriented analysis and design of numerical linear algebra. ACM SIGPLAN Notices, 35(10):229–252, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

http://www.acm.org/pubs/
citations/proceedings/oops/
353171/p229-lujan/.

Lowney:1993:MTS

[LFK^{+}93] P. Geoffrey Lowney, Stefan M.

Freudenberger, Thomas J. Karzes, W. D. Lichtenstein, Robert P. Nix, John S. O’Donnell, and John C. Ruttenberg. The Multiflow Trace Scheduling compiler. The Journal of supercomputing, 7(1– 2):51–142, May 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=7&issue=1&spage=
51.

Lake:1993:POM

[LG93] Brian Lake and Chris Gray.

Parallel I/O for MIMD machines. In Proceedings of SS’93: High Performance Computing, pages 301–308. ????, Calgary, Alberta, Canada, June 1993.

Longman:1992:PDF

[LH92] R. Stewart Longman and Ronald R. Holden. PERM: a double-precision FORTRAN routine for obtaining parallel analysis eigenvalues from permuted data. Behavior research methods, instruments, and computers, 24(3):493, August 1992. CODEN BRM-CEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Luecke:1991:EFV

[LHH^{+}91] Glenn Luecke, Wagar Haque, James Hoekstra, Howard Jespersen, and James Coyle. Evaluation of Fortran vector compilers and preprocessors. Software—Practice and Experience, 21(9):891–905, September 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Luecke:1991:CPE

[LHHJ91] G. Luecke, W. Haque, J. Hoekstra, and H. Jespersen. I/O considerations for performance enhancement under the MVS operating system. Supercomputer, 8(5):41–50, September 1991. CODEN SPCOEL. ISSN 0168-7875.

Luff:2001:RFF

[LHW01] R. Luff, M. Haeckel, and K. Wallmann. Robust and fast FORTRAN and MAT-LAB(R) libraries to calculate pH distributions in marine systems. Computers and Geosciences, 27(2):157–169, March 2001. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

IMSL:1990:QRR

[Lib90a] International Mathematical and Inc. Statistical Libraries. Quick reference: raising FOR-TRAN graphics to a new power. IMSL, Houston, TX, USA, 1990. 62 pp.

IMSL:1990:UMR

[Lib90b] International Mathematical and Inc. Statistical Libraries. User’s manual: raising FOR-TRAN graphics to a new power. IMSL, Houston, TX, USA, 1990. viii + 680 pp.

Lieh:1994:SEM

[Lie94a] Junghsen Lieh. Separated-form equations of motion of controlled flexible multibody systems. Journal of Dynamic Systems, Measurement, and Control, 116(4):702–712, December 1994. CODEN JDSMAA. ISSN 0022-0434.

Lieh:1994:SFE

[Lie94b] Junghsen Lieh. Separated-form equations of motion of controlled flexible multibody systems. Journal of Dynamic Systems, Measurement, and Control, 116(4):702–712, December 1994. CODEN JDSMAA. ISSN 0022-0434.

Lignelet:1991:F

[Lig91a] Patrice Lignelet. Fortran 77.

Masson, Masson, France, 4e edition, 2e tirage edition, 1991. ISBN 2-903607-69-9. 192 pp.

Lignelet:1991:PDF

[Lig91b] Patrice Lignelet. La pratique du Fortran 77. Masson; E.S.I., Paris, France, 2e, 3e tirage edition, 1991. ISBN 2-903607-46-X (E.S.I.), 2-225-82480-0 (Masson). 231 pp.

Lignelet:1993:FAP

[Lig93] Patrice Lignelet. Fortran 90: approche par la pratique. Serie informatique. [Studio image], Menton, France, 1993. ISBN 2-909615-01-4. xii + 228 pp.

Link:1990:FPP

[Lin90] Y. Y. Link. A Fortran program for processing low speed wind tunnel test data for the Jindivik auxiliary intake. Flight mechanics technical memorandum; 426, Dept. of Defence, Defence Science and Technology Organisation Aeronautical Research Laboratory, Melbourne, Victoria, Australia, 1990. ISBN 0-646-03124-4 (not printed on item). iv + 17 + [1] pp.

Ling:1993:SHP

[Lin93] Per Ling. A set of highperformance level 3 BLAS structured and tuned for the IBM 3090 VF and implemented in Fortran 77. The Journal of supercomputing, 7(3):323– 355, September 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=7&issue=3&spage=
323.

Livadas:1991:CT

[Liv91] Panos Livadas. The C-Ghinsu tool. Technical Report SERC-TR-49-F, Software Engineering Research Centre, Utrecht, The Netherlands (??), March 1991.

Larson:2005:MCT

[LJO05] Jay Larson, Robert Jacob, and Everest Ong. The Model Coupling Toolkit: a new Fortran90 toolkit for building multiphysics parallel coupled models. The International Journal of High Performance Computing Applications, 19 (3):277–292, Fall 2005. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/19/3/
277.full.pdf+html.

Laifer:1993:DAT

[LK93a] R. Laifer and A. Knocke. Distributed applications and their use by FORTRAN. In Kusters et al. [KSW93], pages 788–?? ISBN 3-923704-11-9. LCCN ???? Two volumes.

Laifer:1993:FTU

[LK93b] R. Laifer and A. Knocke. fidl — a tool for using DCE from Fortran. Lecture Notes in Computer Science, 731:78–88, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1990:DIPa

[LM90a] Hans Lee and Paul Munsell.

The design and implementation of programs in FORTRAN 77. Prentice-Hall International, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-200510-7 (paperback). xii + 340 pp.

Lee:1990:DIPb

[LM90b] Hans Lee and Paul Munsell. The Design and Implementation of Programs in Fortran 77. Prentice-Hall, Upper Saddle River, NJ 07458, USA, January 1990. ISBN 0-13-199993-1. xii + 340 pp. LCCN QA76.73.F25 L42 1990. US$41.00; US$22.50.

URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0-13-199993-
1.

Loeliger:1994:DIO

[LM94] Jon Loeliger and R. Metzger. Developing an interprocedural optimizing compiler. ACM SIGPLAN Notices, 29(4):41– 48, April 1994. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Langla:1995:GMO

[LMG95] Joel Langla and Françoise Mazat-Gastarriet. GKS: mise en oeuvre en Fortran: une approche par l’exemple. Teknea, Toulouse (France), 1995. ISBN 2-87717-006-3. 377 pp.

Leggett:1996:IUK

[LMJC96] P. F. Leggett, A. T. J. Marsh, S. P. Johnson, and M. Cross. Integrating user knowledge with information from parallelisation tools to facilitate the automatic generation of efficient parallel FORTRAN code. Parallel Computing, 22(2):259– 288, April 5, 1996. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1996&
volume=22&issue=2&aid=1054.

Lumb:1994:UME

[LMK94] Alan M. Lumb, Richard B. Mc-Cammon, and John L. Kittle. Users manual for an expert system (HSPEXP) for calibration of the Hydrological Simulation Program — Fortran. Waterresources investigations report 94-4168, U.S. Geological Survey; USGS Earth Science Information Center Open-File Reports Section [distributor], Reston, VA, USA, 1994. iv + 102 pp.

Lorenzo:1996:HPF

[LMMW96] P. A. R. Lorenzo, A. Mueller, Y. Murakami, and B. J. N. Wylie. High Performance Fortran interfacing to ScaLA-PACK. Lecture Notes in Computer Science, 1184:457– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luksch:1997:SSE

[LMR^{+}97] Peter Luksch, Ursula Maier, Sabine Rathmayer, Matthias Weidmann, and Friedemann Unger. Sempa: Software engineering for parallel scientific computing. IEEE Concurrency, 5(3):64–72, July/ September 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

                                        URL http://dlib.computer.
                                        org/pd/books/pd1997/pdf/
                                        p3064.pdf;   http://www.
                                        computer.org/concurrency/
                                        pd1997/p3064abs.htm.

Luksan:2009:ALA

[LMV09] Ladislav Lukšan, Ctirad Matonoha, and Jan Vlček. Algorithm 896: LSA: Algorithms for large-scale optimization. ACM Transactions on Mathematical Software, 36(3):16:1–16:29, July 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Legler:1991:VFP

[LN91] David M. Legler and I. M. Navon. VARIATM — a FOR-TRAN program for objective analysis of pseudostress wind fields using large-scale conjugate-gradient minimization. Computers and Geosciences, 17(1):1–??, 1991. CO-DEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Lohr:2007:JWB

[Loh07] Steve Lohr. John W. Backus, 82, Fortran developer, dies. New York Times obituary., March 19, 2007. URL http:/

/www.nytimes.com/2007/03/
19/obituaries/20cnd-backus.
html?ex=3D1332043200&en=
3Dadde3ee5a1875330&ei=3D5124&
partner=3Dpermalink&exprod=
3Dpermalink.

Loh:2010:IHP

[Loh10]  Eugene Loh.  The ideal

HPC programming language. Comm. ACM, 53(7):42–47, July 2010. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lopez:1990:FPA

[Lop90] Hugo Rainier Ballina Lopez.

FORTRAN program for automatic terrain correction of gravity measurements. Computers and Geosciences, 16 (2):237–??, 1990. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Lorenzo:2019:AMH

[Lor19] Mark Jones Lorenzo. Abstracting Away the Machine: The History of the Fortran Programming Language (FORmula TRANslation). SE Books, 5307 West Tyson Street, Philadelphia, PA 19107, USA, 2019. ISBN 1-08-239594-3. 325 pp. LCCN QA76.73.F25 BT22 2019.

Loukides:1990:UFP

[Lou90] Michael Kosta Loukides. UNIX for FORTRAN Programmers. Nutshell handbooks Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-51-X. xviii + 244 pp. LCCN QA76.76.O63 L67 1990. US$24.95.

Lovely:1992:LAT

[Lov92] R. Lovely. Loft: a tool for automatic parallelisation of Fortran programs. In M. Valero, E. Onate, M. Jane, J. L. Larriba, and B. Suarez, editors, Parallel Computing and Transputer Applications, pages 277– 286. IOS Press, Amsterdam, The Netherlands, 1992.

Loveman:1993:HPF

[Lov93] D. B. Loveman. High performance Fortran. IEEE Parallel and Distributed Technology: Systems and Applications, 1 (1):25–42, February 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Loveman:1994:DHP

[Lov94] D. B. Loveman. The DEC High Performance Fortran 90 compiler front end. In IEEE [IEE94a], pages 46–53. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Lozier:1998:BRN

[Loz98] Daniel W. Lozier. Book reviews: Numerical recipes in Fortran 90: The art of parallel scientific computing, Volume 2 of Fortan numerical recipes, second edition, by William H. Press, Saul A. Teukolsky, William T. Vetterling and Brian P. Flannery. Mathematics of Computation, 67(221):??, January 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-
5718-98-00891-6&u=/mcom/
1998-67-221/.

Luppi:1990:SCA

[LP90] Jussi Luppi and Petri Pajunen.

Symbolic computation and automatic FORTRAN code generation for eigenvalue determination by phase integral method. Journal of Computational Physics, 88(1):15–30, May 1990. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0021999190902402.

Li:1992:ANL

[LP92] Wei Li and Keshav Pingali. Access normalization: loop restructuring for NUMA compilers. ACM SIGPLAN Notices, 27(9): 285–295, September 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
asplos/143365/p285-li/.

Li:1993:ANL

[LP93] Wei Li and K. Pingali. Access normalization: loop restructuring for NUMA computers. ACM Transactions on Computer Systems, 11(4):353–375, November 1993. CODEN AC-SYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Lin:1998:APS

[LP98a] Y. Lin and D. Padua. On the automatic parallelization of sparse and irregular Fortran programs. Lecture Notes in Computer Science, 1511:41–??, 1998. CODEN LNCSD9. ISSN 0302-9743.

Lucquin:1998:ISC

[LP98b] B. Lucquin and Olivier Pironneau. Introduction to Scientific Computing. John Wiley and Sons, New York, London, Sydney, January 1998. ISBN 0-471-97266-5. ???? pp. LCCN QC20.7.D5L8313 1998. US$64.95.

Lin:1999:APS

[LP99] Yuan Lin and David Padua. On the automatic parallelization of sparse and irregular Fortran programs. Scientific Programming, 7(3–4):231–246, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

                                        //iospress.metapress.com/
                                        app/home/contribution.asp%
                                        3Fwasp=53f7mftrrm4r73yyrqau%
                                        26referrer=parent%26backto=
                                        issue%2C4%2C12%3Bjournal%
                                        2C6%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Landau:2005:FCS

[LP05] Rubin H. Landau and M. J.

Páez. A first course in scientific computing: symbolic, graphical, and numeric problem solving using Maple, Java, Mathematica, and Fortran. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-12183-4 (hardcover). xxiv + 481 pp. LCCN Q183.9 .L36 2005.

LucambioPerez:2019:WLS

[LP19] L. R. Lucambio Pérez and L. F. Prudente. A Wolfe line search algorithm for vector optimization. ACM Transactions on Mathematical Software, 45 (4):37:1–37:23, December 2019. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:/

/dl.acm.org/citation.cfm?
id=3342104.

LeFur:1995:APA

[LPA95] M. Le Fur, J.-L. Pazat, and F. Andre. An array partitioning analysis for parallel loop distribution. In Haridi et al. [HAM95b], pages 351–364. ISBN 3-540-60247-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I553 1995.

Landi:1991:PAP

[LR91] William Landi and Barbara G.

Ryder. Pointer-induced aliasing: a problem taxonomy. In ACM [ACM91], pages 93–103. ISBN ???? LCCN ????

URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/99583/p93-landi/.

Lee:1994:EEP

[LR94] Yong-Fong Lee and Barbara G. Ryder. Effectively exploiting parallelism in data flow analysis. The Journal of supercomputing, 8(3):233– 262, November 1994. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=8&issue=3&spage=
233.

Liep90a

[LS90a] M. Liepel and K. Schittkowski. PCOMP: a FORTRAN code for automatic differentiation. Report No. 254, DFG Schwerpunktprogramm Anwendungsbezogene Optimierung und Optimale Steuerung, Mathematisches Institut, Universität Bayreuth, D-8580 Bayreuth, Germany, 1990.

Liepel:1990:PAF

[LS90b] M. Liepel and K. Schittkowski. PCOMP: A FORTRAN code for automatic differentiation. Report No. 254, DFG Schwerpunktprogramm Anwendungsbezogene Optimierung und Optimale Steuerung, Mathematisches Institut, Universität Bayreuth, D-8580 Bayreuth, Germany, 1990.

Lu:1990:IS

[LS90c] Cheng Chang Lu and Yong Ho Shin. Introducing Starbase. Technical Report CS-9003-09, Kent State University, Kent, OH, USA, 1990. URL mailto:

:sommers@mcs.kent.edu.

Liepelt:2000:RAN

[LS00] Michael Liepelt and Klaus Schittkowski. Remark on algorithm 746: new features of PCOMP: a Fortran code for automatic differentiation. ACM Transactions on Mathematical Software, 26(3):352– 362, September 2000. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lee:2004:OPD

[LS04] H. J. (Hyun Jin) Lee and W. E.

Schiesser. Ordinary and partial differential equation routines in C, C++, Fortran, Java^{(}R), Maple^{(}R), and MATLAB^{(}R). Chapman and Hall/CRC, Boca Raton, FL, USA, 2004. ISBN 1-58488-423-1. xiii + 519 pp. LCCN QA371.5.D37 L44 2004.

Lemmon:2005:DSS

[LS05] David R. Lemmon and J. L.

(Joseph L.) Schafer. Developing statistical software in Fortran 95. Statistics and computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-28123-1, 0-387-23817-4 (paperback). xv + 323 pp. LCCN QA76.5 .L453

                                        2005. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0662/2004061447-d.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0818/2004061447-
                                        t.html.

Lee:2009:AET

[LS09] Che-Rung Lee and G. W. Stewart. Algorithm 879: EIGEN-TEST — a test matrix generator for large-scale eigenproblems. ACM Transactions on Mathematical Software, 35(1): 7:1–7:11, July 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lubeck:1992:PRM

[LSW92] O. M. Lubeck, M. L. Simmons, and H. J. Wasserman. The performance realities of massively parallel processors: a case study. In IEEE [IEE92d], pages 403–412. ISBN 0-8186-2630-

5. LCCN QA76.5 .S894 1992. IEEE catalog no. 92CH3216-9.

Liu:1992:NFP

[LSZ92] Yu Xin Liu, Hong Zhou Sun, and En Guang Zhao. A new FORTRAN program for the CFP’s of a system with identical bosons. Computer Physics Communications, 70(1):154–166, May 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        001046559290098J.

Lucco:1992:DSM

[Luc92] Steven Lucco. A dynamic scheduling method for irregular parallel programs. ACM SIG-PLAN Notices, 27(7):200–211, July 1992. CODEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/143095/p200-lucco/.

Luksan:2001:ANA

[LV01] Ladislav Lukšan and Jan Vlček. Algorithm 811: NDA: algorithms for nondifferentiable optimization. ACM Transactions on Mathematical Software, 27(2):193–213, June 2001. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Levesque:1989:GFS

[LW89] J. Levesque and J. Williamson.

A Guidebook to Fortran on Supercomputers. Academic Press, San Diego, CA, 1989.

Li:1995:CPP

[LW95a] Liwei Li and Paul S. Wang.

The CL–PVM package. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(3– 4):2–8, December 1995. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Li:1995:CP

[LW95b] Liwei Li and Paul S. Wang.

The CL-PVM package. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(3/4): 2–8, December 1995. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lorton:2007:ABL

[LW07] K. Patrick Lorton and David S. Wise. Analyzing block locality in Morton-order and Morton-hybrid matrices. ACM SIGARCH Computer Architecture News, 35(4):6–12, September 2007. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Li:1990:EDD

[LYZ90] Zhiyuan Li, Pen-Chung Yew, and Chuan-Qi Zhu. Efficient data dependence analysis for parallelizing compilers. IEEE Transactions on Parallel and Distributed Systems, 1(1):26– 34, January 1990. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:1997:EHC

[LZ97] Konming Gary Li and Nabil M. Zamel. An evaluation of HPF compilers and the implementation of a parallel linear equation solver using HPF and MPI. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL

http://www.supercomp.org/
sc97/proceedings/TECH/LI/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Li:2011:FPC

[LZL11] Zi Li, Xu Zhang, and Gang Lu. A Fortran program for calculating electron or hole mobility in disordered semiconductors from first-principles. Computer Physics Communications, 182(12):2632–2637, December 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465511002633.

Mojena:1990:F

[MA90] Richard Mojena and Roy Ageloff. Fortran 77. Wadsworth, Pacific Grove, CA, USA, 1990. ISBN 0-534-11742-2. xx + 599 pp.

Muruganandam:2009:FPT

[MA09] P. Muruganandam and S. K.

Adhikari. Fortran programs for the time-dependent Gross– Pitaevskii equation in a fully anisotropic trap. Computer Physics Communications, 180(10):1888–1912, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S001046550900126X.

Muller:2018:NHP

[MA18] Michel Müller and Takayuki Aoki. New high performance GPGPU code transformation framework applied to large production weather prediction code. ACM Transactions on Parallel Computing (TOPC), 5(2):7:1–7:??, January 2018. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Maany:19xx:FAD

[Maaxx] Z. A. Maany. FORTRAN automatic differentiation package for the optimization of functions of many variables, 19xx. To appear.

Macarthur:1990:VFC

[Mac90] Alan Macarthur. A VAX FOR-TRAN command dispatcher. The VAX professional, 12(5): 26–31, October 1990. CODEN VAXPEN. ISSN 8750-9628.

MacDonald:1991:CNC

[Mac91a]  Tom MacDonald.  C for

numerical computing. The Journal of supercomputing, 5 (1):31–48, June 1991. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://www.
                                        springerlink.com/openurl.
                                        asp?genre=article&issn=0920-
                                        8542&volume=5&issue=1&spage=
                                        31.

MacDonald:1991:CCF

[Mac91b] Tom MacDonald. Cray C and Fortran interlanguage communication. The Journal of C Language Translation, 2(4): 305–317, March 1991. ISSN 1042-5721.

MacDonald:JCLT-2-4-305

[Mac91c] Tom MacDonald. Cray C and Fortran interlanguage communication. The Journal of C Language Translation, 2(4): 305–317, March 1991. ISSN 1042-5721.

Macleod:1996:AMS

[Mac96a] Allan J. Macleod. Algorithm 757: MISCFUN, a software package to compute uncommon special functions. ACM Transactions on Mathematical Software, 22(3):288–301, September 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

MacLeod:1996:RAS

[Mac96b] Allan J. MacLeod. Rational approximations, software and test methods for sine and cosine integrals. Numerical Algorithms, 12(3-4):259–272, July 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

MacLeod:1998:AFD

[Mac98] Allan J. MacLeod. Algorithm 779: Fermi–Dirac functions of order −1/2, 1/2, 3/2, 5/2. ACM Transactions on Mathematical Software, 24(1):1–12, March 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Murai:2002:IEH

[MAH^{+}02] Hitoshi Murai, Takuya Araki, Yasuharu Hayashi, Kenji Suehiro, and Yoshiki Seo. Implementation and evaluation of HPF/SX V2. Concurrency and Computation: Practice and Experience, 14(8–9): 603–629, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/95016132/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016132{\&}PLACEBO=IE.
pdf.

Mainprice:1990:FPC

[Mai90] David Mainprice. A FOR-TRAN program to calculate seismic anisotropy from the lattice preferred orientation of minerals. Computers and Geosciences, 16(3):385–??, 1990. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Maine:1991:RNF

[Mai91] R. Maine. Review of NAG Fortran 90 translator. Fortran Journal, 3(6):??, November/ December 1991. ISSN 1060-0221.

Malyshev:1991:VVV

[Mal91] A. N. Malyshev. Vvedenie v vychislitel′nuyu lineĭnuyu algebru (s prilozheniem algoritmov na FORTRANe). (Russian) [Introduction to computational linear algebra (with the application of algorithms in FORTRAN)]. Edited and with a preface and an appendix by S. K. Godunov. “Nauka” Sibirsk. Otdel., Novosibirsk, Siberia, 1991. ISBN 5-02-029682-1. 229 pp.

MNHPCTEC:19xx:FCC

[Manxx] Manchester and North High Performance Computing Training & Education Centre. Fortran 90: a conversion course for Fortran 77 programmers. World-Wide Web document,

19xx.  URL http://www.
hpctec.mcc.ac.uk/hpctec/
courses/Fortran90/F90course.
html.

Marani:1990:TFC

[Mar90] Avi Marani. Two FORTRAN compilers for microcomputers: Ryan-McFarland and Microsoft. Journal of Chemical Information and Computer Sciences, 30(1):96–??, February 1990. CODEN JCISD8. ISSN 0095-2338.

Marshall:1992:ATS

[Mar92] A. C. Marshall. The automatic translation of the supernode MK II occam library into portable Fortran code. In M. Valero, E. Onate, M. Jane, J. L. Larriba, and B. Suarez, editors, Parallel Computing and Transputer Applications, pages 307–316. IOS Press, Amsterdam, The Netherlands, 1992.

Marquet:1993:LED

[Mar93] P. Marquet. Languages and expressions of data parallelism.

Technique et Science Informatiques, 12(6):685–714, 1993. CODEN TTSIDJ. ISSN 0752-4072, 0264-7419.

Margenov:1998:BNR

[Mar98] Svetozar Margenov. Book news & reviews: Introduction to Fortran 90 for Engineers and Scientists. IEEE Computational Science & Engineering, 5(3): 87, July/September 1998. CO-DEN ISCEE4. ISSN 1063-6552 (print), 1558-1861 (electronic).

                                        URL http://dlib.computer.
                                        org/cs/books/cs1998/pdf/
                                        c3087.pdf.

Martens:2007:FFP

[Mar07] China Martens. Father of Fortran programming language dies. ComputerWorld online news story., March 20, 2007. URL http:/

                                        /cwflyris.computerworld.
                                        com/t/1369934/419952/55837/
                                        2/.

Mashaw:1992:PBB

[Mas92a] Bijan Mashaw. Programming byte by byte: structured FOR-TRAN. American Computer Press, Livermore, CA, USA, third edition, 1992. ISBN 0-934433-08-9. xvii + 541 pp.

Maslov:1992:DEW

[Mas92b] Vadim Maslov. Delinearization: an efficient way to break multiloop dependence equations. ACM SIGPLAN Notices, 27(7):152–161, July 1992. CODEN SINODQ.

ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/143095/p152-maslov/.

Mashaw:1993:PBB

[Mas93a]  Bijan Mashaw.  Program-

ming Byte by Byte: Structured Fortran. America Computer Press, Livermore, CA, USA, third edition, June 1993. ISBN 0-934433-08-

9. US$34.45. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0934433089.

Maslov:1993:LAD

[Mas93b] Vadim Maslov. Lazy array data-flow dependence analysis. Technical report, University of Maryland, College Park, College Park, MD, USA, 1993.

URL ftp://ftp.cs.umd.edu/
pub/papers/TRs/3110.ps.Z.

Maslov:1994:LAD

[Mas94] Vadim Maslov. Lazy array data-flow dependence analysis. In ACM [ACM94b], pages 311–325. ISBN 0-89791-636-

0. LCCN QA76.7 .A15 1994.

URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/174675/p311-maslov/.

Mathews:1990:UCS

[Mat90] J. Mathews. Using computer symbolic algebra in applied engineering mathematics to teach Fourier series. CoED, 10(1):47– 52, January–March 1990. CO-DEN CWLJDP. ISSN 0736-8607.

Mackey:1992:RFP

[MB92] Scudder D. Mackey and John S. Bridge. A revised FORTRAN program to simulate alluvial stratigraphy. Computers and Geosciences, 18(2): 119–182, March 1992. CO-DEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Machacek:1995:HPF

[MB95] M. Machacek and E. Bertschinger. A High Performance Fortran PM code for cosmological Nbody simulations. In Anonymous [Ano95a], page 2.07. ISBN ???? ISSN 0002-7537. LCCN ????

Merlin:1999:MDP

[MBFC99] John Merlin, Scott Baden, Stephen Fink, and Barbara Chapman. Multiple data parallelism with HPF and KeLP. Future Generation Computer Systems, 15(3):393–405, April 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Mickevicius:2011:FPH

[MBGK11] S. Mickevicius, E. Brazauskas, D. Germanas, and R. K. Kalinauskas. The four-particle harmonic-oscillator brackets: Compact expressions and updated Fortran program. Computer Physics Communications, 182(6):1377–1381, June 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465511000774.

Mayo:1991:FW

[MC91] William E. Mayo and Martin Cwiakala. The FORTRAN 90 workbook. McGraw-Hill, New York, NY, USA, 1991. ISBN ???? vi + 176 pp. LCCN ????

Mckenzie:1992:CFP

[MC92] Dean P. Mckenzie and David M.

Clarke. CUTOFF: A FOR-TRAN program for establishing thresholds for screening indices. Educational and Psychological Measurement, 52(4): 891–894, Winter 1992. CO-DEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Mayo:1994:SOT

[MC94] William E. Mayo and Martin Cwiakala. Schaum’s Outline of Theory and Problems of Programming With Fortran 77. Schaum’s Outline Series. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-041155-7. 346 pp. US$13.95; US$14.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0070411557.

Mayo:1995:SOTa

[MC95a] William E. Mayo and Martin Cwiakala. Schaum’s outline of theory and problems of programming with Fortran 77. Schaum’s outline series. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-041155-7. vi + 346 pp. LCCN QA76.73.F25 M3945 1995.

Mayo:1995:SOTb

[MC95b] William E. Mayo and Martin Cwiakala. Schaum’s Outline of Theory and Problems of Programming With Fortran 90. Schaum’s Outline Series. McGraw-Hill, New York, NY, USA, May 1995. ISBN 0-07-041156-5. viii + 440 pp. LCCN QA76.73.F25 M3947 1995. US$14.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0070411565.

Morandi:1996:PAC

[MC96] R. Morandi and C. Conti. A parallel algorithm to construct a curve network interpolating non-gridded functional data in R^{3}. Supercomputer, 12(4):33– 45, December 1996. CODEN SPCOEL. ISSN 0168-7875.

Mani:2017:RPR

[MCA17] B. K. Mani, S. Chattopadhyay, and D. Angom. RC-CPAC: a parallel relativistic coupled-cluster program for closed-shell and one-valence atoms and ions in FORTRAN. Computer Physics Communications, 213(??):136–154, April 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465516303629.

Mellor-Crummey:2002:AOS

[MCAB^{+}02] J. Mellor-Crummey, V. Adve, B. Broom, D. Chavarría-Miranda, R. Fowler, G. Jin, K. Kennedy, and Q. Yi. Advanced optimization strategies in the Rice dHPF compiler. Concurrency and Computation: Practice and Experience, 14(8–9):741–767, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/95016130/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016130{\&}PLACEBO=IE.
pdf.

McBryan:1991:CII

[McB91] O. McBryan. A comparison of the Intel iPSC/860 and the Suprenum-1 parallel computers. Supercomputer, 8(1):6–17, January 1991. CODEN SP-COEL. ISSN 0168-7875.

McBane:2006:PCD

[McB06] George C. McBane. Programs to compute distribution functions and critical values for extreme value ratios for outlier detection. Journal of Statistical Software, 16(3):1–9, May 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v16/i03.

McCalpin:1995:CPB

[McC95] John D. McCalpin. A comment on “Performance Benchmark Results for Selected Fortran 90 Compilers” by Prentice and Ameko. Fortran Journal, 7(2):16–19, ???? 1995. ISSN 1060-0221. URL ftp:/

                                        /perelandra.cms.udel.edu/
                                        pub/mccalpin/fortran90.ps.

See [McC95].

McCalpin:1996:CSS

[McC96]  John D. McCalpin.  A

case study of some issues in the optimization of Fortran 90 array notation. Scientific Programming, 5(3):219– 237, Fall 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL http://reality.sgi.
                                        com/employees/mccalpin/papers/
                                        f90.ps.

McDonald:1993:CLF

[McD93] T. M. McDonald. Converting legacy Fortran applications to distributed applications. Lecture Notes in Computer Science, 731:89–103, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McGrath:1991:ZCF

[McG91] Peter H. McGrath. ZERO CROSSOVER — a FORTRAN program to determine the dip and extent of geological boundary using horizontal derivatives of upward-continued gravity data. Computers and Geosciences, 17(7):1017–1032, 1991. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Merlin:1996:SSH

[MCH96] J. H. Merlin, D. B. Carpenter, and A. J. G. Hey. SHPF: a Subset High Performance Fortran compilation system. Fortran Journal, 8(2):2–6, March/ April 1996. ISSN 1060-0221.

McJones:2017:HFF

[McJ17a] Paul McJones. History of FORTRAN and FORTRAN II. Computer History Museum Web site document., March

7, 2017. URL http://www.
softwarepreservation.org/
projects/FORTRAN/. The site

is a treasure trove of early Fortran history, with numerous links to bibliographies, documents, emulators, films, interviews, memoirs, photographs, and source code.

McJones:2017:SOF

[McJ17b] Paul McJones. In search of the original Fortran compiler. IEEE Annals of the History of Computing, 39(2):81–88, ???? 2017. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http://

www.softwarepreservation.
org/projects/FORTRAN/.

McJones:2017:RJB

[McJ17c] Paul McJones. Remembering John Backus. Web site., April

1, 2017. URL http://www.

mcjones.org/dustydecks/archives/
2007/04/01/60/.

Machiels:1997:FEO

[MD97] L. Machiels and M. O. Deville. Fortran 90: An entry to object-oriented programming for solution of partial differential equations. ACM Transactions on Mathematical Software, 23(1):32–49, March 1997. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1997-23-1/p32-machiels/.

Martorana:1994:KPW

[MDD94] A. Martorana, G. Deganello, and D. Duca. K17: a program written in vectorial Fortran for the simulation of powder diffraction patterns. Journal of Applied Crystallography, 27(??):74–79, February 1994. CODEN JACGAR. ISSN 0021-8898.

Muhlleitner:2005:SFC

[MDM05] M. Mühlleitner, A. Djouadi, and Y. Mambrini. SDE-CAY: a Fortran code for the decays of the supersymmetric particles in the MSSM. Computer Physics Communications, 168(1):46–70, May 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465505000822.

Mohr:2007:FPA

[MDV07] Bernd Mohr, Luiz DeRose, and Jeffrey Vetter. A framework for performance analysis of Co-Array Fortran. Concurrency and Computation: Practice and Experience, 19(17):2207– 2218, December 10, 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mehrotra:1993:DPP

[Meh93a] P. Mehrotra. Data parallel programming: the promises and limitations of High Performance Fortran. In Volkert [Vol93], page 114. ISBN 0-387-57314-3 (U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I5 1993. DM58.00.

Mehrotra:1993:ILD

[Meh93b] P. Mehrotra. Invited lecture: Data parallel programming: The promises and limitations of High Performance Fortran. In Volkert [Vol93], pages 114–?? ISBN 0-387-57314-3 (U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I5 1993. DM58.00.

Mehrotra:1994:HPF

[Meh94] P. Mehrotra. High Performance FORTRAN. In Wieseman [Wie94], pages 546–561. ISBN ???? ISSN 0191-7811. LCCN ????

Meissner:1995:F

[Mei95] Loren P. Meissner. Fortran 90. PWS Pub. Co., Boston, MA, USA, 1995. ISBN 0-534-93372-6. xii + 679 pp. LCCN QA 76.73 F25 M435 1995. US$50.95.

Meissner:1996:POT

[Mei96]  Loren Meissner.  Personal

observations teaching with F. Fortran Journal, 8(6): ??, November/December 1996. ISSN 1060-0221. URL http:/

                                        /www.fortran.com/fortran/
                                        FJ/9611/#teaching.

GarciaMerayo:1991:PF

[Mer91] Felix Garcia Merayo. Programacion en FORTRAN 77. Paraninfo, Madrid, Spain, 3a, corr. y ampliada edition, 1991. ISBN 84-283-1818-2. 399 pp.

Meredith:1992:NPF

[Mer92a] Roger W. Meredith. Numeric precision in FORTRAN computing. Computers in physics, 6(5):506–512, September 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Merlin:1992:AFA

[Mer92b] J. H. Merlin. Adapting Fortran 90 array programs for distributed memory architectures. Lecture Notes in Computer Science, 591(??):184–200, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Metcalf:1992:FEF

[Met92a] M. Metcalf. A first encounter with Fortran 90. Fortran Journal, 4(1):2–7, January/ February 1992. ISSN 1060-0221.

Metcalf:1992:SPA

[Met92b] Michael Metcalf. Still programming after all these years. New scientist, 135(1838):30– ??, September 1992. CODEN NWSCAL. ISSN 0262-4079, 0028-6664.

Metcalf:1995:HPF

[Met95] M. Metcalf. High Performance Fortran. In Vandoni [Van95], pages 193–197. ISBN 92-9083-076-X. ISSN 0007-8328. LCCN QC770 .E83 v.95, no. 5.

Metcalf:1999:FH

[Met99a] Michael Metcalf. Fortran 90/ 95/HPF. ACM SIGPLAN Notices, 34(12):24–29, December 1999. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Metcalf:1999:RFI

[Met99b] Michael Metcalf. Resources:

Fortran information file. ACM SIGPLAN Notices, 34(12):24– 29, December 1999. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

http://ftp.informatik.rwth-
aachen.de/dblp/db/indices/
a-tree/m/Metcalf:Michael.
html.

Metcalf:1999:IFCa

[Met99c] Mike Metcalf. Information file on compilers, tools, books, courses, tutorials, and the standard for the Fortran language and its derivatives: Version of 20 May 1999 (the penultimate year of the millennium). Scientific Programming, 7(1):327– 333, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Metcalf:1999:IFCb

[Met99d] Mike Metcalf. Information file on compilers, tools, books, courses, tutorials, and the standard for the Fortran language and its derivatives: Version of 20 May 1999 (the penultimate year of the millennium). Scientific Programming, 7(3–4):327– 333, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.com/
app/home/contribution.asp%
3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C11%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Meyers:2000:NCIb

[Mey00] Randy Meyers. The new C: It all began with FORTRAN. C/C++ Users Journal, 18(11): 80–??, November 2000. CO-DEN CCUJEX. ISSN 1075-2838.

Meyers:2001:NCW

[Mey01] Randy Meyers. The new C:

Why variable length arrays? C/C++ Users Journal, 19(10): 46–??, October 2001. CODEN CCUJEX. ISSN 1075-2838.

Miyoshi:1994:DAN

[MFI^{+}94] H. Miyoshi, M. Fukuda, T. Iwamiya, T. Nakamura, M. Tuchiya, M. Yoshida, K. Yamamoto, Y. Yamamoto, S. Ogawa, Y. Matsuo, T. Yamane, M. Takamura, M. Ikeda, S. Okada, Y. Sakamoto, et al. Development and achievement of NAL Numerical Wind Tunnel (NWT) for CFD computations. In IEEE [IEE94f], pages 685–692. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Meerbergen:2009:CBE

[MFK09] Karl Meerbergen, Kresimir Fresl, and Toon Knapen. C++ bindings to external software libraries with examples from BLAS, LAPACK, UMFPACK, and MUMPS. ACM Transactions on Mathematical Software, 36(4):22:1–22:23, August 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

More:1981:AFS

[MGH81] Jorge J. Moré, Burton S. Garbow, and Kenneth E. Hillstrom. Algorithm 566: FOR-TRAN subroutines for testing unconstrained optimization software [C5 [E4]]. ACM Transactions on Mathematical Software, 7(1):136–140, March 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [AFS94].

MacLaren:1991:FPS

[MH91] L. D. MacLaren and S. D. Hill. A Fortran program for spectral analysis using the fast Fourier transform. Flight mechanics technical memorandum; 432, Dept. of Defence, Defence Science and Technology Organisation Aeronautical Research Laboratory, Melbourne, Victoria, Australia, 1991. ISBN 0-646-05178-4 (not printed on item). ii + 16 + [3] pp.

Merlin:1995:IHP

[MH95] John Merlin and Anthony Hey. An introduction to High Performance Fortran. Scientific Programming, 4(2):87– 113, Summer 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Mulder:2012:BFP

[MHdL12] Joris Mulder, Herbert Hoijtink, and Christiaan de Leeuw. BIEMS: a Fortran 90 program for calculating Bayes factors for inequality and equality constrained models. Journal of Statistical Software, 46 (2):??, January 2012. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v46/i02.

Matsumoto:1996:AAP

[MHT96] Aki Matsumoto, D. S. Han, and Takao Tsuda. Alias analysis of pointers in Pascal and Fortran 90: Dependence analysis between pointer references. Acta Informatica, 33(2):99–130, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer-ny.
com/link/service/journals/
00236/bibs/6033002/60330099.
htm.

Microsoft:1991:MF

[Mic91] Microsoft. Microsoft FOR-TRAN, 1991.

Microsoft:1993:MFPb

[Mic93a] Microsoft Corporation. Microsoft FORTRAN powerStation, 1993. 5 computer disks + 1 language guide (xiv + 557 p.) + 1 getting started guide (iv + 3) + 1 error messages guide (vii + 102 p.).

Microsoft:1993:MFPa

[Mic93b] Microsoft Corporation. Microsoft FORTRAN powerStation 32 for Windows NT development system for Windows NT, 1993.

Michalakes:1997:MSP

[Mic97] J. Michalakes. MM90: a scalable parallel implementation of the Penn State/NCAR Mesoscale Model (MM5). Parallel Computing, 23(14):2173– 2186, December 17, 1997. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1997&
volume=23&issue=14&aid=1265.

Miller:1991:MST

[Mil91] Phillip C. Miller. Middleman; a software tool to aid in parallelizing dusty-deck Fortran. Thesis (M.S.), Oregon Graduate Institute of Science and Technology, Beaverton, OR, USA, 1991. vii + 86 pp.

Millar:1992:CFM

[Mil92]  G. T. Millar.  CIRCLE:

a FORTRAN monohull ship turning circle prediction program. Technical communication 92/302, Defence Research Establishment Atlantic, Dartmouth, NS, Canada, 1992. v + 26 pp.

Millman:1993:AP

[Mil93] Joel Millman. Artistic programming. Forbes, 151(7): 94–??, March 1993. CODEN FORBA5. ISSN 0015-6914.

Miller:2004:AMF

[Mil04a] Alan Miller. Alan Miller’s Fortran software. Web site, February 4, 2004. URL https://

jblevins.org/mirror/amiller/

. From the Web site: All code written by Alan Miller is released into the public domain.

Miller:2004:CMS

[Mil04b] Alan J. Miller. A collection of mathematical and statistical routines in FOR-TRAN 90. Journal of Statistical Software, 9(3):??, February 25, 2004. CODEN JS-SOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/
v09/i03/F90lib.pdf.

Mochizuki:1995:WML

[MIN^{+}95] M. Mochizuki, Y. Itoh, M. Nii, T. Takagi, and Y. Mitsui. A wideband monolithic lossy match power amplifier having an LPF/HPF-combined interstage network. IEICE Transactions on Electronics, E78-C(9): 1252–1254, September 1995. CODEN IELEEJ. ISSN 0916-8524.

Miranda:1990:FCP

[Mir90] Manuel J. Miranda. Fortran 77 computer program for analysis of multicell thin walled section under combined loading. Thesis (M.S.), California State University, Northridge, Northridge, CA, USA, 1990. xi + 108 pp.

Mitchell:1992:SBC

[Mit92] Ed Mitchell. Secrets of the Borland C++ Masters. Howard W. Sams, Indianapolis, IN 46268, USA, November 1, 1992. ISBN 0-672-30137-7. Includes disk.

Mitra:1993:FPP

[Mit93] Mitra. A Fortran: Program for problems in beam bending. CoED, 3(4):35–40, October 1993. CODEN CWLJDP. ISSN 0736-8607.

Mitchell:1997:SMP

[Mit97] William F. Mitchell. Stop-Watch: a module for portable measurement of execution time. Fortran Journal, 9(1): ??, 1997. ISSN 1060-0221.

URL http://math.nist.gov/
~WMitchell/papers/stopwatch_
fj.ps.gz;   http://www.fortran.
com/fortran/FJ/9701/mitchell/
mitchell.html.

Mitchell:2002:DPA

[Mit02] W. F. Mitchell. The design of a parallel adaptive multi-level code in Fortran

90. Lecture Notes in Computer Science, 2331:672–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2331/23310672.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2331/23310672.
pdf.

Marazzi:1993:ARF

[MJR93] Alfio Marazzi, Johann Joss, and Alex Randriamiharisoa. Algorithms, Routines, and S Functions for Robust Statistics: The FORTRAN Library ROBETH with an Interface to S-PLUS. Wadsworth and Brooks/Cole Publishing Co., Pacific Grove, CA, USA and Pacific Grove, CA, USA, 1993.

ISBN 0-534-19698-5. xii + 436 pp. LCCN QA276.4 .M267 1993. US$69.50. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0534196985.

Murthy:1992:SAR

[MKC92] V. K. Murthy, E. V. Krishnamurthy, and Pin Chen. Systolic algorithm for rational interpolation and Padé approximation. Parallel Computing, 18 (1):75–83, January 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

McShan:1995:AIP

[MKF95] D. L. McShan, M. L. Kessler, and B. A. Fraass. Advanced interactive planning techniques for conformal therapy: high level beam descriptions and volumetric mapping techniques. International Journal of Radiation Oncology, Biology, Physics, 33(5):1061– 1072, December 1, 1995. CO-DEN IOBPD3. ISSN 0360-3016.

Majaess:1992:AAA

[MKFB92] Fouad Majaess, Patrick Keast, Graeme Fairweather, and Karin R. Bennett. Algorithm 704: ABDPACK and ABBPACK—FORTRAN programs for the solution of almost block diagonal linear systems arising in spline collocation at Gaussian points with monomial basis functions.

ACM Transactions on Mathematical Software, 18(2):205– 210, June 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0098-
3500/146927.html.

Moore:1994:MPP

[MKS94] M. S. Moore, G. Karsai, and J. Sztipanovits. Model-based programming for parallel image processing. In IEEE [IEE94b], pages 811–815 (vol. 3). ISBN 0-8186-6952-7 (casebound), 0-8186-6950-0 (paperback), 0-8186-6951-9 (microfiche). LCCN TA1637.I25 1994. Three volumes. IEEE catalog no. 94CH35708.

Maley:1996:FSA

[MKS^{+}96] D. Maley, P. L. Kilpatrick, E. W. Schreiner, N. S. Scott, and G. H. F. Diercksen. The formal specification of abstract data types and their implementation in Fortran 90: implementation issues concerning the use of pointers. Computer Physics Communications, 98(1–2):167–180, October 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0010465596000938.

Meadows:1994:MCF

[MM94] L. Meadows and D. Miles. Migrating CM Fortran applications to HPF. In IEEE [IEE94a], pages 37–40. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Moreira:1998:FCC

[MM98] José E. Moreira and Samuel P.

Midkiff. Fortran 90 in CSE: a case study. IEEE Computational Science & Engineering, 5(2):39–49, April/ June 1998. CODEN IS-CEE4. ISSN 1063-6552 (print), 1558-1861 (electronic).

URL http://dlib.computer.
org/cs/books/cs1998/pdf/
c2039.pdf;   http://www.
computer.org/cse/cs1998/
c2039abs.htm.

Mullick:2002:FPC

[MM02] M. Mullick and R. K. Majumdar. A Fortran program for computing the mise-a-la-masse response over a dykelike body. Computers and Geosciences, 28(9):1119–1126, November 2002. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Meinke:2008:SVS

[MMEH08] Jan H. Meinke, Sandipan Mohanty, Frank Eisenmenger, and Ulrich H. E. Hansmann. SMMP v. 3.0 — simulating proteins and protein interactions in Python and Fortran. Computer Physics Communications, 178(6):459–470, March 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0010465507004614.

Moreira:1998:CFC

[MMG98] J. E. Moreira, S. P. Midikiff, and M. Gupta. A comparison of Fortran, C/C++, and Java for numerical computing. IEEE Antennas and Propagation Magazine, 40(5):102–105, October 1998. CODEN IAP-MEZ. ISSN 1045-9243.

Moreira:2000:FMJ

[MMG00] José E. Moreira, Samuel P. Midkiff, and Manish Gupta. From flop to megaflops: Java for technical computing. ACM Transactions on Programming Languages and Systems, 22(2): 265–295, March 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL http://www.acm.org/
                                        pubs/citations/journals/
                                        toplas/2000-22-2/p265-moreira/

.

Milligan:1992:FED

[MMRS92] P. Milligan, R. K. McConnell, S. A. Rea, and P. P. Sage. Fort-Port: An environment for the development of parallel Fortran programs. Microprocessing and Microprogramming, 34(??):73– 76, 1992. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

Martins:2009:POO

[MMT09] Joaquim R. R. A. Martins, Christopher Marriage, and Nathan Tedford. pyMDO: An object-oriented framework for multidisciplinary design optimization. ACM Transactions on Mathematical Software, 36 (4):20:1–20:25, August 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Middleton:1995:EDS

[MMV95] D. Middleton, P. Mehrotra, and J. Van Rosendale. Expressing direct simulation Monte Carlo methods in High Performance Fortran. In Bailey et al. [BBG^{+}95], pages 698–703. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Meadows:1995:PRS

[MMY95a] L. F. Meadows, D. Miles, and M. Young. Performance results of several High Performance Fortran benchmarks. In IEEE [IEE95a], pages 516–517. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

Miles:1995:PRS

[MMY95b] D. Miles, L. Meadows, and M. Young. Performance results of several High Performance Fortran benchmarks. In IEEE [IEE95a], pages 516–517. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

Morales:2001:APF

[MN01] José Luis Morales and Jorge Nocedal. Algorithm 809: PREQN: Fortran 77 subroutines for preconditioning the conjugate gradient method. ACM Transactions on Mathematical Software, 27(1):83–91, March 2001. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Morales:2011:RAB

[MN11] José Luis Morales and Jorge Nocedal. Remark on “Algorithm 778: L-BFGS-B: Fortran subroutines for large-scale bound constrained optimization”. ACM Transactions on Mathematical Software, 38(1): 7:1–7:4, November 2011. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [ZBLN97].

Marsaglia:1990:DBR

[MNZ90] George Marsaglia, B. Narasimhan, and Arif Zaman. The distance between random points in rectangles. Communications in Statistics: Theory and Methods, 19(11):4199–4212, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moldenhauer:2012:FIS

[Mol12] Jacob Moldenhauer. Fortranning it with style [review of Modern Fortran: Style and Usage (Clerman, N.S. and Spector, W.; 2012)]. Computing in Science and Engineering, 14(5):5–6, September/October 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Moore:1995:OFA

[Moo95a] David L. Moore. Objectoriented facilities in Ada 95. Dr. Dobb’s Journal of Software Tools, 20(10):??, October 1995. CODEN DDJTEQ. ISSN 1044-789X.

Moore:1995:OOF

[Moo95b] David L. Moore. Objectoriented facilities in Ada 95. Dr. Dobb’s Journal of Software Tools, 20(10):28, 30, 32, 34– 35, October 1995. CODEN DDJOEB. ISSN 1044-789X.

Morris:1981:CAR

[Mor81] Alfred H. Morris, Jr. Can Ada replace FORTRAN for numerical computation? ACM SIG-PLAN Notices, 16(12):10–13, December 1981. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morris:2015:EMI

[Mor15] Karla Morris. Emulating multiple inheritance in Fortran 2003/2008. Scientific Programming, 2015(??):126069:1– 126069:7, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2015/126069/.

Mosli:1995:CCF

[Mös95] Bernd Mösli. A comparison of C++, FORTRAN 90 and Oberon-2 for scientific programming. In GISI 95, pages 740–748. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-642-79958-2. ISSN 1431-472X.

URL http://link.springer.
com/chapter/10.1007/978-
3-642-79958-7_103.

Maslov:1993:SPC

[MP93] Vadim Maslov and William Pugh. Simplifying polynomial constraints over integers to make dependence analysis more precise. Technical report, University of Maryland, College Park, College Park, MD, USA, 1993. URL

ftp://ftp.cs.umd.edu/pub/
papers/TRs/3109.ps.Z. This

is revised version of the CS-TR-3109 report that appeared July, 1993.

Metcalf:1987:FE

[MR87] Michael Metcalf and John Ker Reid. Fortran 8x Explained. Oxford University Press and Clarendon Press, Walton Street, Oxford OX2 6DP, UK and Oxford, UK, 1987. ISBN 0-19-853751-4 (hardcover), 0-19-853731-X (paperback). xiv + 262 pp. LCCN QA76.73.F26 M48 1987. UK£12.50 (US$25.00 U.S.), 0198537514. See also [Ame87, Ame90b].

Metcalf:1990:FEa

[MR90a] Michael Metcalf and John Ker Reid. Fortran 8x explained. Clarendon Press; Oxford University Press, Oxford, UK, revised, repr. (with corrections) edition, 1990. ISBN 0-19-853824-3. xv + 315 pp.

Metcalf:1990:FE

[MR90b] Michael Metcalf and John Ker Reid. Fortran 90 Explained. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853772-7 (paperback). xiv + 294 pp. LCCN QA76.73.F28 M48 1990. US$22.45. See [Ame92].

Metcalf:1991:FE

[MR91] Michael Metcalf and John Reid. Fortran 90 Explained. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, February 1991. ISBN 0-19-853772-7. US$40.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0198537727.

Metcalf:1992:FE

[MR92] Michael Metcalf and John Ker Reid. Fortran 90 explained. Oxford science publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1992. ISBN 0-19-853772-7 (paperback). xiv + 294 pp.

Metcalf:1993:FE

[MR93a] Michael Metcalf and John Ker Reid. Fortran 90 explained. Oxford science publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-853772-7 (paperback). xiv + 306 pp.

Miminis:1993:AFS

[MR93b] George Miminis and Michael Reid. Algorithm 718: A FOR-TRAN subroutine to solve the eigenvalues allocation problem for single-input systems. ACM Transactions on Mathematical Software, 19(2):224– 232, June 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0098-
3500/152618.html.

Metcalf:1994:FE

[MR94] Michael Metcalf and John Ker Reid. Fortran 90 explained. Oxford science publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, reprinted with corrections edition, 1994. ISBN 0-19-853772-7. xiv + 306 pp.

Miminis:1995:AFS

[MR95a] George Miminis and Helmut Roth. Algorithm 747: A Fortran subroutine to solve the eigenvalue assignment problem for multiinput systems using state feedback. ACM Transactions on Mathematical Software, 21(3):299–326, September 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Mueller:1995:EHP

[MR95b] A. Mueller and R. Ruehl. Extending High Performance Fortran for the support of unstructured computations. In ACM [ACM95a], pages 127– 136. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995.

Metcalf:1996:FE

[MR96a] Michael Metcalf and John Ker Reid. Fortran 90/95 Explained. Oxford science publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, May 1996. ISBN 0-19-851888-9. xv + 345 pp. LCCN QA76.73.F28 M48 1996. US$29.95; US$26.95.

URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0198518889.

Muller:1996:CDI

[MR96b] A. Muller and R. Ruhl.

Communication-buffers for data-parallel, irregular computations. In Szymanski and Sinharoy [SS96], pages 295–298. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Metcalf:1999:FE

[MR99] Michael Metcalf and John K.

Reid. Fortran 90/95 Explained. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 1999. ISBN 0-19-850558-2. 358 pp.

Mraz:1994:RVP

[Mra94] Ronald Mraz. Reducing the variance of point to point transfers in the IBM 9076 parallel computer. In IEEE [IEE94f], pages 620–629. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Metcalf:2004:FE

[MRC04] Michael Metcalf, John Ker Reid, and Malcolm Cohen. Fortran 95/2003 Explained. Numerical mathematics and scientific computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-852692-X (hardcover), 0-19-852693-8 (paperback). xviii + 416 pp. LCCN QA76.73.F235 M48 2004.

Metcalf:2011:MFE

[MRC11] Michael Metcalf, John Ker Reid, and Malcolm Cohen. Modern Fortran explained. Numerical mathematics and scientific computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, fourth edition, 2011. ISBN 0-19-960142-9 (paperback), 0-19-960141-0. xx + 488 pp. LCCN QA76.73.F235 M48

2011. URL http://www.loc.
gov/catdir/enhancements/
fy1312/2012360954-b.html;
 http://www.loc.gov/catdir/
enhancements/fy1312/2012360954-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy1312/2012360954-t.html.

Metcalf:2018:MFE

[MRC18] Michael Metcalf, John Ker Reid, and Malcolm Cohen. Modern Fortran explained: incorporating Fortran 2018. Numerical mathematics and scientific computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, fifth edition, 2018. ISBN 0-19-185002-0 (e-book), 0-19-881188-8 (paperback), 0-19-881189-6 (hardcover). xxii + 522 pp. LCCN QA76.73.F25.

Metcalf:2023:MFE

[MRCB23] Michael Metcalf, John Reid, Malcolm Cohen, and Reinhold Bader. Modern Fortran Explained: Incorporating Fortran 2023. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, sixth edition, December 2023. ISBN 0-19-887657-2. LCCN QA76.73.F25 M4 2023.

Metcalf:1993:FCF

[MRG^{+}93] Michael Metcalf, John Reid, Denis Girou, Michel Caillat, and Bernard Pichon. Fortran 90: les concepts fondamentaux. AFNOR technique. AFNOR, Paris, France, 1993. ISBN 2-12-486513-7. xxii + 376 pp.

Metzger:1993:ICP

[MS93a] Robert Metzger and Sean Stroud. Interprocedural constant propagation: an empirical study. ACM Letters on Programming Languages and Systems, 2(4):213– 232, March 1993. CODEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

                                        URL http://www.acm.org/
                                        pubs/toc/Abstracts/1057-
                                        4514/176526.html.

Morgan:1993:PF

[MS93b] J. S. (J Steve) Morgan and J. L. (J Lawrie) Schonfelder. Programming in Fortran 90. Computer science texts. Alfred Waller, Henley-on-Thames, UK, 1993. ISBN 1-872474-06-3 (paperback). xiv + 350 pp.

Mudge:1994:PTS

[MS94] T. N. Mudge and B. D. Shriver, editors. Proceedings of the Twenty-Seventh Hawaii Internation Conference on System Sciences Vol. I: Architecture. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5050-8. LCCN ???? IEEE catalog no. 94TH0607-2.

Mascagni:2000:ASS

[MS00a] Michael Mascagni and Ashok Srinivasan. Algorithm 806: SPRNG: a scalable library for pseudorandom number generation. ACM Transactions on Mathematical Software, 26 (3):436–461, September 2000. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See correction [MS00b].

Mascagni:2000:CAS

[MS00b] Michael Mascagni and Ashok Srinivasan. Corrigendum: Algorithm 806: SPRNG: a scalable library for pseudorandom number generation. ACM Transactions on Mathematical Software, 26(4):618–619, December 2000. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [MS00a].

Martins:2003:CSD

[MSA03] Joaquim R. R. A. Martins, Peter Sturdza, and Juan J. Alonso. The complexstep derivative approximation. ACM Transactions on Mathematical Software, 29(3):245– 262, September 2003. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Merodio:1992:FPC

[MSB92] Julio C. Merodio, Luis A. Spalleti, and Luis M. Bertone. A FORTRAN program for the calculation of normative composition of clay minerals and pelitic rocks. Computers and Geosciences, 18(1):47–62, January 1992. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

McLay:1996:MSM

[MSC96] R. T. McLay, S. Swift, and G. F. Carey. Maximizing sparse matrix-vector product performance on RISC based MIMD computers. Journal of Parallel and Distributed Computing, 37(2):146– 158, September 15, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jpdc.1996.0115/production;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0115/production/pdf.

Mamidi:2022:PAG

[MSP^{+}22] Nischay Ram Mamidi, Dhruv Saxena, Kumar Prasun, Anil Nemili, Bharatkumar Sharma, and S. M. Deshpande. Performance analysis of GPU accelerated meshfree q-LSKUM solvers in Fortran, C, Python, and Julia. In IEEE, editor, 2022 IEEE 29th International Conference on High Performance Computing, Data, and Analytics (HiPC), pages 156–

165. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2022.

Morel-Seytoux:1990:UMK

[MSZ90] Hubert J. Morel-Seytoux and Chuan Mian Zhang. User’s manual for Kergen-Kersim Package: Fortran 77 programs for simulation of a streamaquifer system, interim report. Hydrology Days Publications, Fort Collins, CO, USA, 1990. ISBN ???? ii + 37 + 38 pp. LCCN ????

Martello:1990:KPA

[MT90] Silvano Martello and Paolo Toth. Knapsack Problems: Algorithms and Computer Implementations. John Wiley and Sons, New York, London, Sydney, 1990. ISBN ???? LCCN ????

McKevitt:2025:AFC

[MVK25] James McKevitt, Eduard I.

Vorobyov, and Igor Kulikov. Accelerating Fortran codes: a method for integrating Coarray Fortran with CUDA Fortran and OpenMP. Journal of Parallel and Distributed Computing, 195(??):??, January 2025. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0743731524001412.

Mehrotra:1998:HPFb

[MVZ98a] P. Mehrotra, J. Van Rosendale, and H. Zima. High Performance Fortran: Status and prospects. Lecture Notes in Computer Science, 1541:345– 356, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehrotra:1998:HPFa

[MVZ98b] Piyush Mehrotra, John Van Rosendale, and Hans Zima. High Performance Fortran: History, status and future. Parallel Computing, 24(3–4): 325–354, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.elsevier.
com/cas/tree/store/parco/
sub/1998/24/3-4/1293.pdf.

Marsh:1990:UMP

[MWM90] B. D. Marsh, T. M. Williams, and G. L. Mathieson. The use of mixed Prolog/Fortran for battle simulation. OR; the journal of the Operational Research Society, 41(4):311– 318, April 1990. CODEN OPRQAK. ISSN 0160-5682 (print), 1476-9360 (electronic).

Morton:1995:LLP

[MWO95] Don Morton, Kefei Wang, and David O. Ogbe. Lessons learned in porting Fortran/ PVM code to the Cray T3D. IEEE Parallel and Distributed Technology: Systems and Applications, 3(1):4–11, 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Mehrotra:2000:HPF

[MZ00] Piyush Mehrotra and Hans Zima. High Performance Fortran for aerospace applications. NASA contractor report NASA/CR-2000-210321, National Technical Information Service, Washington, DC, USA, 2000. ???? pp. Shipping list number 2001-0362-M.

Mehrotra:2001:HPF

[MZ01] Piyush Mehrotra and Hans Zima. High Performance Fortran for aerospace applications. Parallel Computing, 27(4):477– 501, March 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/35/21/47/28/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/35/21/47/28/30/article.
pdf.

Marsaglia:1994:REI

[MZM94] George Marsaglia, Arif Zaman, and John C. W. Marsaglia. Rapid evaluation of the inverse of the normal distribution function. Statistics & Probability Letters, 19(4):259–266, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Marsaglia:1990:TUR

[MZT90] George Marsaglia, Arif Zaman, and Wai Wan Tsang. Toward a universal random number generator. Statistics & Probability Letters, 9(1):35–39, 1990. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Nagel:1990:EAC

[Nag90] W. E. Nagel. Exploiting autotasking on a CRAY Y-MP: an improved software interface to multitasking. Parallel Computing, 13(2):225–233, February 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nagai:1995:BFI

[Nag95] T. Nagai. Benchmarking Fortran intrinsic functions. International Journal of High Speed Computing, 7(2):217–??, 1995.

CODEN IHSCEZ. ISSN 0129-0533.

Nag:19xx:NFM

[Nagxx] Nag. NAg FORTRAN mini manual. NAg, Oxford, UK, 19xx. Distributed by NAg (USA).

Nagle:2001:MFV

[Nag01] Dan Nagle. Multithreading, Fthreads, and Visual Fortran. Dr. Dobb’s Journal of Software Tools, 26(7):36, 38, 40, July 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2001/2001_07/
                                        fthreads.zip.

Nagle:2002:FS

[Nag02] Dan Nagle. The Fortran 2000 Standard. Dr. Dobb’s Journal of Software Tools, 27(2):48, 50– 52, February 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2002/2002_02/fortn2k.
                                        txt.

Naito:2017:BBT

[Nai17] O. Naito. A browser-based tool for conversion between Fortran NAMELIST and XML/ HTML. SoftwareX, 6(??):19– 24, ???? 2017. CODEN ???? ISSN 2352-7110. URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S235271101630036X.

Nakao:1990:SAN

[Nak90] M. Nakao. State of the art for numerical computations with guaranteed accuracy. Journal of Information Processing, 31 (9):1177–1190, 1990. CODEN JIPRDE. ISSN 0387-6101.

Nakao:1995:GEB

[Nak95a] Mitsuhiro Nakao. Guaranteed error bounds for the finite element solutions of elliptic partial differential equations. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/
proceedings/nakao/nakao1.
html.

Nakatani:1995:SIHb

[Nak95b] T. Nakatani. Status and issues of HPF compilers. Transactions of the Institute of Electronics, Information and Communication Engineers D-I, J78D-I(2):142–148, February 1995. CODEN DTRDES.

Nakatani:1995:SIHa

[Nak95c] Toshio Nakatani. Status and issues of HPF compilers. Systems and computers in Japan, 26(12):1–??, ???? 1995. CO-DEN SCJAEP. ISSN 0882-1666 (print), 1520-684X (electronic).

Nance:1993:HDE

[Nan93a] Richard E. Nance. A history of discrete event simulation programming languages. ACM SIGPLAN Notices, 28(3):149–175, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/proceedings/
                                        plan/154766/p149-nance/.

Nandakumar:1993:FPD

[Nan93b] Ratna Nandakumar. A FOR-TRAN 77 program for detecting differential item functioning through the Mantel– Haenszel statistic. Educational and Psychological Measurement, 53(3):679–684, Fall 1993. CODEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Nandakumar:1993:FPA

[Nan93c] Ratna Nandakumar. A FOR-TRAN program for assessing unidimensionality of binary data using Holland and Rosenbaum’s methodology. Multivariate behavioral research, 28 (1):63–66, January 1993. CO-DEN MVBRAV. ISSN 0027-3171.

Nardelli:1995:PUP

[Nar95] M. Nardelli. PARST95 — an update to PARST: a system of Fortran routines for calculating molecular structure parameters from the results of crystal structure analyses. Journal of Applied Crystallography, 28(5): 659–??, 1995. CODEN JAC-GAR. ISSN 0021-8898.

Nataf:1992:ASN

[Nat92] J.-M. Nataf. Algorithm of simplification of nonlinear equations systems. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 26(3): 9–16, August 1992. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

NASA:2000:DCB

[Nat00] National Aeronautics and Space Administration. Developing CORBA-based distributed scientific applications from legacy Fortran programs. Technical Report NASA/TM-2000-209950, National Aeronautics and Space Administration, Washington, DC, USA, December 6, 2000.

Noye:1992:CTA

[NBC92] John Noye, Basil Benjamin, and Len Colgan, editors. Computational techniques and applications: proceedings of 5th International Computational Techniques and Applications Conference, held at The University of Adelaide, 14–17 July, 1991, Computational Techniques and Applications. Australian Mathematics Society, Adelaide, South Australia, Australia, 1992. ISBN 0-86396-172-X. LCCN ????

Nanthaamornphong:2015:EUC

[NCMF15] Aziz Nanthaamornphong, Jeffrey Carver, Karla Morris, and Salvatore Filippone. Extracting UML class diagrams from object-oriented Fortran: ForUML. Scientific Programming, 2015(??):421816:1– 421816:15, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2015/421816/.

Norton:1996:POO

[NDS96] Charles D. Norton, Viktor K. Decyk, and Boleslaw K. Szymanski. On parallel object oriented programming in Fortran 90. ACM SIGAPP Applied Computing Review, 4(1):27– 31, April 1996. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/240732.240742.

Norton:2007:TAM

[NDSG07] Charles D. Norton, Viktor K. Decyk, Boleslaw K. Szymanski, and Henry Gardner. The transition and adoption to modern programming concepts for scientific computing in Fortran. Scientific Programming, 15(1): 27–44, ???? 2007. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Neumiller:2001:HSP

[Neu01] Kurt Neumiller. Hydrological simulation program — FOR-TRAN (HSPF) calibration for Mica Creek, Idaho. Special report 01-01, National Council for Air and Stream Improvement, Research Triangle Park, NC, USA, January 2001. various pp.

Neshyba:1993:ILC

[NG93] S. P. Neshyba and R. R.

Gamache. Improved linebroadening coefficients for asymmetric rotor molecules with application to ozone lines broadened by nitrogen. Journal of Quantitative Spectroscopy and Radiative Transfer, 50(5): 443–453, November 1993. CO-DEN JQSRAE. ISSN 0022-4073.

Nguyen:1991:UMT

[Ngu91] T. X. Nguyen. User manual for two simple PostScript output Fortran plotting routines interim report. NASA contractor report NASA CR-187590, National Aeronautics and Space Administration, Langley Research Center; National Technical Information Service, distributor, Hampton, VA, USA, 1991. ???? pp.

Naumann:2009:OVE

[NH09] Uwe Naumann and Yuxiao Hu. Optimal vertex elimination in single-expression-use graphs. ACM Transactions on Mathematical Software, 35(1): 2:1–2:20, July 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Nguyen:2003:AVF

[NI03] T. V. N. Nguyen and F. Irigoin.

Alias verification for Fortran code optimization. J.UCS: Journal of Universal Computer Science, 9(3):270–??, March 28, 2003. CODEN ???? ISSN 0948-6968. URL http://www.

                                        jucs.org/jucs_9_3/alias_
                                        verification_for_fortran.

Nicolau:1991:ALC

[Nic91] Alexandru Nicolau. Advances in languages and compilers for parallel processing. Research monographs in parallel and distributed computing. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-64028-7. ISSN 0953-7767. 467 pp. LCCN QA76.7 .A38 1991.

Nielsen:1992:BFP

[Nie92] Roger L. Nielsen. BIGD.FOR: a FORTRAN program to calculate trace-element partition coefficients for natural mafic and intermediate composition magmas. Computers and Geosciences, 18(7):773–??, August 1992. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Nikhil:1993:PPL

[Nik93] R. S. Nikhil. The parallel programming language Id and its compilation for parallel machines. International Journal of High Speed Computing, 5(2): 171–223, 1993. CODEN IH-SCEZ. ISSN 0129-0533.

Nishida:1995:BPE

[Nis95]  Takaaki Nishida.  Bifurca-

tion problems for equations of fluid dynamics and computer assisted proof. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ????

LCCN ???? URL http:

//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/
proceedings/nishida/nishida.
ps.

Nakamura:1994:EPV

[NIY^{+}94] H. Nakamura, H. Imori, Y. Yamashita, K. Nakazawa, T. Boku, H. Li, and I. Nakata. Evaluation of pseudo vector processor based on slidewindowed registers. In Mudge and Shriver [MS94], pages 368– 377. ISBN 0-8186-5050-8. LCCN ???? IEEE catalog no. 94TH0607-2.

Nesbitt:1994:FPG

[NJ94a] Lloyd Nesbitt and Michael R.

Jones. A FORTRAN program to generate comparison data to test tomography inversion programs. NOAA technical memorandum ERL ETL 245, U.S. Dept. of Commerce, National Oceanic and Atmospheric Administration, Environmental Research Laboratories, Environmental Technology Laboratory, Boulder, CO, USA, 1994. iv + 63 pp.

Nesbitt:1994:FPP

[NJ94b] Lloyd Nesbitt and R. Michael Jones. A FORTRAN program for performing nonperturbative ocean acoustic tomography inversion. Noaa technical memorandum erl etl; 243, U.S. Dept. of Commerce, National Oceanic and Atmospheric Administration, Environmental Research Laboratories, Environmental Technology Laboratory, Boulder, CO, USA, 1994. iv + 165 pp.

Niewels:1994:SDA

[NJ94c] J. Niewels and W. Jorden. Systematical development of an autonomous HPF driven and controlled inspection robot. In Chung et al. [CKMU94], pages 324–329. ISBN 1-880653-10-9 (paperback: set), 1-880653-11-7 (paperback: v. 1), 1-880653-12-5 (paperback: v. 2), 1-880653-13-3 (paperback: v. 3), 1-880653-14-1 (paperback: v. 4). LCCN TC1665 .I579 1994 v. 1-4 (1994).

Neelakantan:1994:IIT

[NK94] K. Neelakantan and D. K. Kumar. Is it time to retire FOR-TRAN? Current science, 66(5): 334–??, 1994. CODEN CUS-CAM. ISSN 0011-3891.

Nyhoff:1992:FES

[NL92] Larry R. Nyhoff and Sanford Leestma. Fortran 77 for Engineers and Scientists. Macmillan; Maxwell Macmillan Canada; Maxwell Macmillan International, New York, NY, USA, third edition, January 1992. ISBN 0-02-388655-2. xii + 630 pp. LCCN QA76.73.F25N9 1992. US$58.33. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=0023886552.

Nyhoff:1995:FNMa

[NL95a] Larry Nyhoff and Sanford Leestma. Fortran 77 and Numerical Methods for Engineers and Scientists. Prentice-Hall, Upper Saddle River, NJ 07458, USA, April 1995. ISBN 0-02-388741-9 (paperback). xviii + 764 pp. LCCN QA 76.73 F25 N89 1995. US$65.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0023887419. Includes

disk.

Nyhoff:1995:FNMb

[NL95b] Larry R. Nyhoff and Sanford Leestma. FORTRAN 77 and numerical methods for engineers and scientists. Macmillan College Pub.: Maxwell Macmillan International; Maxwell Macmillan Canada, New York, NY, USA, 1995. ISBN 0-02-388741-9. xviii + 764 pp. LCCN QA 76.73 F25 N89 1995.

Nyhoff:1996:IFE

[NL96] Larry R. Nyhoff and Sanford C. Leestma. Introduction to Fortran 90 for Engineers and Scientists. Prentice Hall Modular Series for Engineering. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-505215-7. 360 pp. US$31.60. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0135052157.

Nyhoff:1997:FES

[NL97a] Larry R. Nyhoff and Sanford Leestma. Fortran 90 for Engineers and Scientists. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-519729-5. xxvii + 119 pp. LCCN QA76.73.F25 N92 1997. US$53.00. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=0135197295.

Nyhoff:1997:IFE

[NL97b] Larry R. Nyhoff and Sanford Leestma. Introduction to Fortran 90 for engineers and scientists. Prentice Hall modular series for engineering. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-505215-7. xiv + 359 + A37 + I5 pp. LCCN QA76.73.F25 N925 1997. US$31.60.

Nanthaamornphong:2019:EFA

[NL19] Aziz Nanthaamornphong and Anawat Leatongkam. Extended ForUML for automatic generation of UML sequence diagrams from objectoriented Fortran. Scientific Programming, 2019(1): 2542686:1–2542686:??, 2019. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https:/

                                        /onlinelibrary.wiley.com/
                                        doi/epdf/10.1155/2019/2542686.

Norman:2023:PCC

[NLBB23] Matthew Norman, Isaac Lyngaas, Abhishek Bagusetty, and Mark Berrill. Portable C++ code that can look and feel like Fortran code with Yet Another Kernel Launcher (YAKL). International Journal of Parallel Programming, 51(4-5): 209–230, October 2023. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10766-022-00739-0.

Namashivayam:2020:MFI

[NLE^{+}20] Naveen Namashivayam, Bill Long, Deepak Eachempati, Bob Cernohous, and Mark Pagel. A modern Fortran interface in OpenSHMEM need for interoperability with Parallel Fortran using coarrays. ACM Transactions on Parallel Computing (TOPC), 7(4): 24:1–24:25, December 2020. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

URL https://dl.acm.org/
doi/10.1145/3418084.

Nyhoff:1996:FESa

[NLN96] Larry R. Nyhoff, Sanford Leestma, and Larry Nyhoff. Fortran 77 for Engineers and Scientists: With an Introduction to Fortran 90. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, January 1996. ISBN 0-13-363003-X. xxv + 884 pp. LCCN QA76.73.F25 N9 1996. US$53.00; US$52.00.

URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=013363003X.

Norman:2015:CSC

[NLVE15] Matthew Norman, Jeffrey Larkin, Aaron Vose, and Katherine Evans. A case study of CUDA FORTRAN and OpenACC for an atmospheric climate kernel. Journal of Computational Science, 9:1–6, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1877750315000605.

Nishitani:2002:TCI

[NNON02] Yasunori Nishitani, Kiyoshi Negishi, Hiroshi Ohta, and Eiji Nunohiro. Techniques for compiling and implementing all NAS parallel benchmarks in HPF. Concurrency and Computation: Practice and Experience, 14(8–9): 769–787, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/95016133/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=95016133{\&}PLACEBO=IE.
                                        pdf.

Noble:1990:FTF

[Nob90] J. V. Noble. A FORmula TRANslator for Forth. In RFC:1990 [RFC90], pages 110–

113. ISBN ???? LCCN ????

Nucciarone:1997:NLD

[NOL97] Jeffrey J. Nucciarone, Yusuf Ozyoruk, and Lyle N. Long.

New life in dusty decks: Results of porting a CM Fortranbased aeroacoustic model to High Performance Fortran. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

//www.supercomp.org/sc97/
proceedings/TECH/NUCCIARO/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Norwood:1991:MWP

[Nor91]  John Norwood.  Mixed-

language Windows programming. Dr. Dobb’s Journal of Software Tools, 16(10):104– 108, 130, October 1991. CO-DEN DDJOEB. ISSN 1044-789X.

Nardin:1992:ACN

[NPB92] Mark Nardin, W. F. Perger, and Atul Bhalla. Algorithm 707: CONHYP: a numerical evaluator of the confluent hypergeometric function for complex arguments of large magnitudes. ACM Transactions on Mathematical Software, 18(3):345–349, September 1992. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1992-18-3/p345-nardin/.

Numrich:1998:CAF

[NR98a] Robert W. Numrich and John Reid. Co-Array Fortran for parallel programming. ACM SIGPLAN FORTRAN Forum, 17(2):1–31, August 1998. ISSN 1061-7264 (print), 1931-1311 (electronic).

Numrich:1998:SRC

[NR98b] Robert W. Numrich and John Reid. Special report: Co-Array Fortran for parallel programming. ACM SIGPLAN FORTRAN Forum, 17(2):1–31, August 1998. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Naumann:2005:DEF

[NR05] Uwe Naumann and Jan Riehme. A differentiation-enabled Fortran 95 compiler. ACM Transactions on Mathematical Software, 31(4):458–474, December 2005. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Naumann:2006:CAN

[NR06] Uwe Naumann and Jan Riehme. Computing adjoints with the NAGWare Fortran 95 compiler. In Bücker et al. [BCH^{+}06], pages 159–169. CO-DEN LNCSA6. ISBN 3-540-28403-6 (print), 3-540-28438-9 (e-book). ISSN 1439-7358. LCCN QA304 .I58

                                        2006. URL http://link.
                                        springer.com/content/pdf/
                                        10.1007/3-540-28438-9_14.

Proceedings of the Fourth International Conference on Automatic Differentiation, July 20–23, Chicago, Illinois.

Numrich:1998:WMS

[NRK98] R. W. Numrich, J. Reid, and K. Kim. Writing a multigrid solver using co-array Fortran. Lecture Notes in Computer Science, 1541:390–399, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nye:1992:IPS

[NRS92] Mary Jo Nye, Joan L.

Richards, and Roger H. Stuewer. The Invention of Physical Science: Intersections of Mathematics, Theology and Natural Philosophy Since the Seventeenth Century: Essays in Honor of Erwin N. Hiebert, volume 139 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 94-010-5097-X, 94-011-2488-4 (e-book). ISSN 0068-0346. xxxiv + 278 pp. LCCN ???? URL https:

//link.springer.com/book/
10.1007/978-94-011-2488-1.

Nash:1992:ABS

[NS92] Stephen G. Nash and Ariela Sofer. Algorithm 711: BTN: Software for parallel unconstrained optimization. ACM Transactions on Mathematical Software, 18(4):414–448, December 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/

                                        pubs/citations/journals/
                                        toms/1992-18-4/p414-nash/.

Nesterenko:2011:QFV

[NS11] A. V. Nesterenko and C. Simolo. QCDMAPT F: Fortran version of QCDMAPT package. Computer Physics Communications, 182(10):2303–2304, October 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0010465511002116.

Numrich:1998:DEF

[NSJD98] R. W. Numrich, J. L. Steidel, B. H. Johnson, and B. Dupont de Dinechin. Definition of the F_{−−} extension to Fortran 90. Lecture Notes in Computer Science, 1366:292–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743.

Naterop:2020:HRN

[NSU20] L. Naterop, A. Signer, and Y. Ulrich. handyG — rapid numerical evaluation of generalised polylogarithms in Fortran. Computer Physics Communications, 253(??):Article 107165, August 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0010465520300230.

Noid:1990:MDS

[NSWP90] D. W. Noid, B. G. Sumpter, B. Wunderlich, and G. A. Pfeffer. Molecular dynamics simulations of polymers: Methods for optimal Fortran programming. Journal of Computational Chemistry, 11(2):236– 241, March 1990. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

NAG:1990:NFLa

[Num90a] Numerical Algorithms Group.

NAG Fortran library introductory guide: Mark 14. NAG Inc., Downers Grove, IL, USA, 1990. various pp.

NAG:1990:NFLb

[Num90b] Numerical Algorithms Group.

NAG Fortran library manual, mark 14. Numerical Algorithms Group, Oxford, UK, 1990. ???? pp.

NAG:1990:NFLc

[Num90c] Numerical Algorithms Group.

NAG Fortran library, Mark 14: maintenance bulletin FR14C. Numerical Algorithms Group, Oxford, UK, 1990. 58 pp.

NAG:1991:HNF

[Num91a] Numerical Algorithms Group.

Handbook for the NAG Fortran workstation library, Version 1. Numerical Algorithms Group, Inc., Oxford, UK, 1991. various pp.

NAG:1991:NFLa

[Num91b] Numerical Algorithms Group.

NAG FORTRAN library introductory guide: Mark 15. NAG, Oxford, UK, 1991. various pp.

NAG:1991:NFLb

[Num91c] Numerical Algorithms Group. NAG Fortran library manual, mark 15. NAG, Oxford, UK, 1991. ???? pp.

NRS:1992:NRF

[Num92] Numerical Recipes Software. Numerical Recipes in Fortran. Cambridge University Press, Cambridge, UK, October 1992. ISBN 0-521-35467-

6. US$34.50. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=0521354676;   http:/
                                        /www.nr.com/nronline_switcher.
                                        html.

NAG:1993:NFLa

[Num93a] Numerical Algorithms Group. NAG Fortran library introductory guide: Mark 16. NAG, Oxford, UK, 1993. various pp.

NAG:1993:NFLb

[Num93b] Numerical Algorithms Group. NAG Fortran library manual, mark 16. NAG, Oxford, UK, 1993. ???? pp.

NAG:19xx:NFL

[Numxx] Numerical Algorithms Group. NAg FORTRAN library manual. Numerical Algorithms Group, Oxford, United Kingdom, 19xx. various pp.

Numrich:2005:PNA

[Num05] Robert W. Numrich. Parallel numerical algorithms based on tensor notation and Co-Array Fortran syntax. Parallel Computing, 31(6):588–607, June 2005. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Norwood:1994:SMP

[NV94] John Norwood and Shankar Vaidyanathan. Symmetric multiprocessing for PCs. Dr. Dobb’s Journal of Software Tools, 19(1):80, 82–85, 98– 99, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Ngo:1996:FCS

[NVC96] Dat Ngo, Gorden Videen, and Petr Chýlek. A FORTRAN code for the scattering of EM waves by a sphere with a nonconcentric spherical inclusion. Computer Physics Communications, 99(1):94–112, December 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465596001099.

Nieto-Vesperinas:1993:FRE

[NVFNP93] M. Nieto-Vesperinas, F. J.

Fuentes, R. Navarro, and M. J. Perez-Ilzarbe. A FORTRAN routine to estimate a function of two variables from its autocorrelation. Computer Physics Communications, 78 (1):211–217, December 1993. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Ning:1994:ADE

[NVG94] Qi Ning, V. Van Dongen, and G. R. Gao. Automatic decomposition in EPPP compiler. In Botsford et al. [BGG^{+}94], pages 283–291. ISBN ???? LCCN ????

Yi:1990:SIC

[nY90] Yong no Yi. Soputuweo immun: chobo tangye e imhan haksup kyojae: chonja kyesan kaeron, Fortran, Cobol, Basic. Hangmundang, Soul Tukpyolsi, Korea, 1990. ISBN ???? 340 pp. LCCN ????

Navon:1991:ETE

[NY91] I. M. Navon and Jian Yu. EXSHALL: a Turkel-Zwas explicit large time-step FOR-TRAN program for solving the shallow-water equations in spherical coordinates. Computers and Geosciences, 17(9): 1311–1344, 1991. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Okuda:2002:OEE

[OA02] Hiroshi Okuda and Norihisa Anan. Optimization of element-by-element FEM in HPF 1.1. Concurrency and Computation: Practice and Experience, 14(8–9): 647–663, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/95016128/
                                        START;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=95016128{\&}PLACEBO=IE.
pdf.

OBoyle:1993:DPA

[O’B93] Michael O’Boyle. A data partioning algorithm for distributed memory compilation. Technical Report UMCS-93-7-1, University of Manchester, Department of Computer Science, Manchester, UK, July 1993. URL mailto:

:techreports@cs.man.ac.uk.

Olmos:1994:FFL

[OC94] J. Olmos and K. Chan. FLAN:

Fortran language database utility for SONGS 2/3 simulator. Simulation series, 26(3): 41–??, 1994. CODEN SMC-PAX. ISSN 0735-9276.

Ochem:2009:GIA

[Och09] Quentin Ochem. Gem #55:

introduction to Ada /Java interfacing. ACM SIGADA Ada Letters, 29(2):43–45, August 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ottenstein:1992:ECF

[OE92] Karl J. Ottenstein and Steven J.

Ellcey. Experience compiling Fortran to program dependence graphs. Software— Practice and Experience, 22 (1):41–62, January 1992. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Oed:1993:CRM

[Oed93] Wilfried Oed. The Cray Research massively parallel processor system CRAY T3D. Technical report, Cray Research GmbH, München, Germany, November 15, 1993.

Offner:1994:DSM

[Off94] C. D. Offner. A data structure for managing parallel operations. In Hesham and Shriver [HS94b], pages 33–42. ISBN 0-8186-5060-5. ISSN 1060-3425. LCCN ???? IEEE catalog no. 94TH0607-2.

Offner:1998:PBH

[Off98] Carl D. Offner. Per Brinch Hansen’s concerns about High Performance Fortran. ACM SIGPLAN Notices, 33(8):34– 39, August 1998. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ogino:2002:TDG

[Ogi02]  Tatsuki Ogino.  Three-

dimensional global MHD simulation code for the Earth’s magnetosphere using HPF/JA. Concurrency and Computation: Practice and Experience, 14(8–9):631–646, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/95016136/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?

ID=95016136{\&}PLACEBO=IE.
pdf.

Okawa:1990:LAP

[OH90] Y. Okawa and N. Haraguchi.

A linear array of processors with partially shared memory for parallel solution of PDE. In IEEE [IEE90a], pages 41– 48. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Overbey:2009:RLR

[OJ09] Jeffrey L. Overbey and Ralph E.

Johnson. Regrowing a language: refactoring tools allow programming languages to evolve. ACM SIGPLAN Notices, 44(10):493–502, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

OKeefe:1993:FTA

[O’K93]  Matthew O’Keefe.  The

Fortran-P translator: automatic translation of Fortran 77 programs for massively parallel processors. AHPCRC preprint 93-021, Army High Performance Computing Research Center, Minneapolis, MN, USA, 1993. iii + 34 pp.

Okumura:1995:AIA

[Oku95] Kohshi Okumura. On the applications of interval arithmetic electrical network analysis. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.nsc.liu.se/~boein/
                                        ifip/kyoto/workshop-info/
                                        proceedings/okumura/okumura1.
                                        html.

Olagnon:1992:ENF

[Ola92] M. Olagnon. Experience with NagWare f90. Fortran Journal, 4(6):2–5, November/December 1992. ISSN 1060-0221.

Olagnon:1993:FFF

[Ola93] Michel Olagnon. F90 FAN’s: Fortran 90 frequently asked about news. December 7, 1993. Available by anonymous

                                        ftp from molene.ifremer.fr,

1993.

Olagnon:1995:FFP

[Ola95] M. Olagnon. f90ppr: a Fortran 90 pre-processor; A Fortran 90 pretty-printer. Fortran Journal, 7(2):8–14, March/April 1995. ISSN 1060-0221.

Olagnon:1996:LGN

[Ola96] Michel Olagnon. Little giants — the new Fortran subset. Fortran Journal, 8(6): ??, November/December 1996. ISSN 1060-0221. URL http:/

                                        /www.fortran.com/fortran/
                                        FJ/9611/#ficiency.

Olver:1991:RUB

[Olv91] F. W. J. Olver. Book review: Mathematical Function Library for Microsoft-FORTRAN, by United Laboratories, Inc.: Wiley, New York, 1989. Mathematics of Computation, 56(194):879– 885, April 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/1991-56-194/
S0025-5718-91-99751-6?active=
current.

Oldehoeft:1990:MAI

[OM90] R. R. Oldehoeft and J. R. Mc-Graw. Mixed applicative and imperative programs. Parallel Computing, 13(2):175–191, February 1990. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Orkwis:1992:NMS

[OM92] Paul D. Orkwis and D. Scott McRae. Newton’s method solver for high-speed viscous separated flowfields. American Institute of Aeronautics and Astronautics Journal, 30 (1):78–85, January 1992. CO-DEN AIAJAH. ISSN 0001-1452.

Okabe:1995:NFA

[ONT95] Y. Okabe, M. Nakamura, and T. Tsuda. New fast algorithms for first-order linear recurrences on vector computers. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/
proceedings/okabe/okabe.
ps.

Orlando:1998:CLE

[OP98a] S. Orlando and R. Perego. A coordination layer for exploiting task parallelism with HPF. Lecture Notes in Computer Science, 1511:386–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orlando:1998:MRS

[OP98b] S. Orlando and R. Perego. An MPI-based run-time support to coordinate HPF tasks. Lecture Notes in Computer Science, 1497:289–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ott:2020:FKD

[OPB^{+}20] Jordan Ott, Mike Pritchard, Natalie Best, Erik Linstead, Milan Curcic, and Pierre Baldi. A Fortran–Keras deep learning bridge for scientific computing. Scientific Programming, 2020(1):8888811:1–8888811:??, 2020. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https:/

                                        /onlinelibrary.wiley.com/
                                        doi/epdf/10.1155/2020/8888811.

OKeefe:1995:FPT

[OPE^{+}95] Matthew O’Keefe, Terence Parr, B. Kevin Edgar, Steve Anderson, Paul Woodward, and Hank Dietz. Fortran-P translator: towards automatic translation of Fortran 77 programs for massively parallel processors. Scientific Programming, 4(1):1–21, Spring 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Orlando:2000:MDT

[OPP00] S. Orlando, P. Palmerini, and R. Perego. Mixed data and task parallelism with HPF and PVM. Cluster Computing, 3 (3):201–213, 2000. CODEN ???? ISSN 1386-7857.

Orchard:2015:EFT

[ORO15] Dominic Orchard, Andrew Rice, and Oleg Oshmyan. Evolving Fortran types with inferred units-of-measure. Journal of Computational Science, 9:156–162, July 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S1877750315000563.

Ortega:1994:IFSb

[Ort94a] James M. Ortega. An Introduction to Fortran 90 for Scientific Computing. Saunders College Pub., Fort Worth, TX, USA, August 1994. ISBN 0-03-010198-0. vi + 234 pp. LCCN QA 76.73 F25 O75 1994b. US$30.00; US$20.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0030101980.

Ortega:1994:IFSa

[Ort94b] James M. Ortega. An Introduction to Fortran for Scientific Computing. Saunders College Pub., Fort Worth, TX, USA, January 1994. ISBN 0-03-003128-1. viii + 204 pp. LCCN QA76.73.F25 O75 1994. US$30.00. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=0030031281.

Osyczka:1992:CAM

[Osy92] Andrzej Osyczka. Computer Aided multicriterion Optimization System (CAMOS): software package in FORTRAN. International Software Publishers, Krakow, Poland, 1992. ISBN ???? ???? pp. LCCN ????

Ong:1993:COE

[OT93] C. L. Ong and W. T. Tsai. Class and object extraction from imperative code. Journal of Object-Oriented Programming, 6(1):58–??, March 1993. CODEN JOOPEC. ISSN 0896-8438.

Overill:1991:EVG

[Ove91] R. Overill. On the efficient vectorization of the general firstorder linear recurrence relation. Supercomputer, 8(2):31– 36, March 1991. CODEN SP-COEL. ISSN 0168-7875.

Prentice:1994:PBR

[PA94] J. K. Prentice and Agbeli K. Ameko. Performance benchmark results for selected Fortran 90 compilers. Fortran Journal, 6(6):??, November/ December 1994. ISSN 1060-0221. See comment [McC95].

Padua:2000:FC

[Pad00]  David Padua.  The For-

tran I compiler. Computing in Science and Engineering, 2(1):70–75, January/ February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

URL http://dlib.computer.
org/cs/books/cs2000/pdf/
c1070.pdf;   http://www.
computer.org/cse/cs1999/
c1070abs.htm.

Page:1995:PPG

[Pag95] Clive G. Page. The professional programmers guide to Fortran 77. Pitman Publishing Ltd., London, UK, 1995. ISBN 0-273-02856-1. ix + 180 pp. LCCN QA 76.73 F25 P34 1988. US$.

URL ftp://ftp.star.le.ac.
uk/pub/fortran/prof77.ps.
gz.

Pajunen:1990:GME

[Paj90] P. Pajunen. Generation of more efficient Fortran code via symbolic software. Supercomputer, 7(1):6–41, January 1990. CODEN SPCOEL. ISSN 0168-7875.

Pfeiffer:1990:BOS

[PAK^{+}90] Wayne Pfeiffer, Arnold Alagar, Anke Kamrath, Robert H. Leary, and Jack Rogers. Benchmarking and optimization of scientific codes on the CRAY X-MP, CRAY-2, and SCS-40 vector computers. The Journal of supercomputing, 4(2): 131–152, June 1990. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=4&issue=2&spage=
131.

Pao:1999:EAI

[Pao99] Yen-Ching Pao. Engineering Analysis: Interactive Methods and Programs with FOR-TRAN, QuickBASIC, MAT-LAB, and Mathematica. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1999. ISBN 0-8493-2016-X. 360 pp. LCCN TA345.P36 1999. US$79.95.

Pao:2001:EAI

[Pao01]  Y. C. Pao.  Engineering

analysis: interactive methods and programs with FORTRAN, QuickBASIC, MATLAB, and Mathematica. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001.

Papadimitriou:1993:KNA

[Pap93]  Pythagoras  Papadimitriou.

The KSR1—A numerical analyst’s perspective. Numerical Analysis Report 242, University of Manchester, Manchester, UK, December 1993.

URL ftp://vtx.ma.man.ac.
uk/pub/narep/narep242.ps.
Z.

Parker:1986:SFC

[Par86] J. R. Parker. A subset FOR-TRAN compiler for a modified Harvard architecture. ACM SIGPLAN Notices, 21(9):57– 62, September 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parker:1994:BFL

[Par94] Andrew M. Parker. Binding FORTRAN layered earth modeling routines with Windows graphical user interface. Computers and Geosciences, 20(2): 221–??, March 1994. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Pasewark:1995:MWV

[Pas95] Pasewark. Microsoft Works 3.0 DOS Version; Tutorial and Applications (Dt-FORTRAN). South Western, January 1995. ISBN 0-538-63435-9. US$28.95; US$26.05. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0538634359.

Paul:1993:FCP

[Pau93] Christopher A. H. Paul. A FORTRAN 77 code for plotting stability regions. Numerical Analysis Report 230, University of Manchester, Manchester, UK, May 1993.

URL ftp://vtx.ma.man.ac.
uk/pub/narep/narep230.ps.
Z.

Pazat:1996:THP

[Paz96] J.-L Pazat. Tools for High Performance Fortran: a survey. Lecture Notes in Computer Science, 1132:134–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pingali:1995:LCP

[PBG^{+}95] K. Pingali, U. Banerjee, D. Gelernter, A. Nicolau, and D. Padua, editors. Languages and compilers for parallel computing: 7th International Workshop, Ithaca, NY, USA, August 8–10, 1994: proceedings, volume 892 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-58868-X. LCCN QA76.58 .W656 1994.

Pasquarell:1995:PFC

[PBU95] G. C. Pasquarell, D. G. Boyer, and J. B. Urban. PACKER — A FORTRAN 77 code for collection, analysis, and display of interval pressure injection test data. Computers and Geosciences, 21(4):481–??, 1995. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Prins:1998:EIC

[PCS98] J. F. Prins, S. Chatterjee, and M. Simons. Expressing irregular computations in modern Fortran dialects. Lecture Notes in Computer Science, 1511:1– ??, 1998. CODEN LNCSD9. ISSN 0302-9743.

Prins:1999:ICF

[PCS99] Jan F. Prins, Siddhartha Chatterjee, and Martin Simons. Irregular computations in Fortran — expression and implementation strategies. Scientific Programming, 7(3–4):313– 326, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.com/
app/home/contribution.asp%
3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C10%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Perrin:1996:DPP

[PD96] Guy-Rene Perrin and Alain Darte. The data parallel programming model: foundations, HPF realization, and scientific applications, volume 1132 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNCSD9. ISBN 3-540-61736-1. ISSN 0302-9743 (print), 1611-3349 (electronic). xv + 284 pp. LCCN QA76.642 .D38 1996.

Ponnusamy:1993:UCD

[PDS^{+}93] R. Ponnusamy, R. Das, J. Saltz, D. Mavriplis, and A. Choudhary. Unstructured computations and the Dybbuk runtime system. In Tentner [Ten93], pages 170–175. ISBN 1-56555-052-8. LCCN ????

Pelliccia:1993:DFH

[Pel93] Anthony Pelliccia. Development of a FORTRAN hydrologic model and its application to a Caribbean mangrove wetland. Thesis (M.S.), University of North Carolina at Charlotte, Charlotte, NC, USA, 1993. x + 127 pp.

ACM:1992:PAW

[PEP92] Proceedings of the 1992 ACM Workshop on Partial Evaluation and Semantics-Based Program Manipulation. ACM Press, New York, NY 10036, USA, June 1992. ISBN ???? LCCN ????

Per:1993:SHL

[Per93] Ling Per. A set of highperformance level 3 BLAS structured and tuned for the IBM 3090 VF and implemented in Fortran 77. The Journal of supercomputing, 7(3):323–356, September 1993. CODEN JO-SUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Perrin:1994:SSA

[Per94] G.-R. Perrin. Some synthesis aspects for data parallel programming. In Gentzsch and Harms [GH94a], pages 469–474. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Peters:1991:SMV

[Pet91] A. Peters. Sparse matrix vector multiplication techniques on the IBM 3090 VF. Parallel Computing, 17(12):1409–1424, December 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pais:2004:UHP

[PFS^{+}04] V. A. Pais, N. Fournier, M. A.

Sutton, K. J. Weston, and U. Dragosits. Using High Performance Fortran to parallelise a multi-layer atmospheric transport model. Parallel Computing, 30(1):21–33, January 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Peng:2010:AFC

[PG10] Liang-You Peng and Qihuang Gong. An accurate Fortran code for computing hydrogenic continuum wave functions at a wide range of parameters. Computer Physics Communications, 181(12):2098– 2101, December 2010. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465510003413.

Polychronopoulos:1990:SPA

[PGH^{+}90] C. Polychronopoulos, M. Girkar, M. Haghighat, C. Lee, B. Leung, and D. Schouten. The structure of Parafrase-2: An advanced parallelizing compiler for C and Fortran. In D. Gelernter, A. Nicolau, and D. Padua, editors, Proceedings of the Third (??) Workshop on Languages and Compilers for Parallel Computing. MIT Press, Cambridge, MA, USA, 1990.

Parashar:1996:CTP

[PH96] Manish Parashar and Salim Hariri. Compile-time performance prediction of HPF/ Fortran 90D. IEEE Parallel and Distributed Technology: Systems and Applications, 4(1):57–73, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL

http://dlib.computer.org/
pd/books/pd1996/pdf/p1057.
pdf;   http://www.computer.
org/concurrency/pd1996/p1057abs.
htm.

Pascual:2006:ETT

[PH06] Valéerie Pascual and Laurent Hascoëet. Extension of TAPE-NADE toward Fortran 95. In Bücker et al. [BCH^{+}06], pages 171–179. CODEN LNCSA6. ISBN 3-540-28403-6 (print), 3-540-28438-9 (e-book). ISSN 1439-7358. LCCN QA304 .I58

2006. URL http://link.
springer.com/content/pdf/
10.1007/3-540-28438-9_15.

Proceedings of the Fourth International Conference on Automatic Differentiation, July 20–23, Chicago, Illinois.

Ponnusamy:1995:SID

[PHD^{+}95] Ravi Ponnusamy, Yuan-Shin Hwang, Raja Das, Joel H. Saltz, Alok Choudhary, and Geoffrey Fox. Supporting irregular distributions using dataparallel languages. IEEE Parallel and Distributed Technology: Systems and Applications, 3(1):12–24, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL

http://dlib.computer.org/
pd/books/pd1995/pdf/h10012.
pdf;   http://www.computer.
org/concurrency/pd1995/p1012abs.
htm.

Parashar:1994:DAD

[PHHF94a] M. Parashar, S. Hariri, T. Haupt, and G. C. Fox. Design of an application development toolkit for HPF/Fortran 90D. In Kumar [Kum94], pages 482–487. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Parashar:1994:IPH

[PHHF94b] Manish Parashar, Salim Hariri, Tomasz Haupt, and Geoffrey C. Fox. Interpreting the performance of HPF/Fortran 90D. In IEEE [IEE94f], pages 743–752. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Phillips:1991:PBL

[Phi91a] C. Phillips. The performance of the BLAS and LAPACK on a shared memory scalar multiprocessor. Parallel Computing, 17(6–7):751–761, September 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Phillips:1991:TTP

[Phi91b] Jeffrey D. Phillips. TER-RACE a terracing procedure for gridded data, with Fortran programs, and VAX command procedure, Unix C-shell, and DOS batch file implementations. Open-file report 92-5-A, U.S. Geological Survey: [Books and Open-File Reports Section distributor], Denver, CO, USA, 1991. iii + 20 pp.

Phillips:1992:TTP

[Phi92] Jeffrey D. Phillips. TER-RACE a terracing procedure for gridded data, with Fortran programs, and VAX command procedure, Unix C-shell, and DOS batch file implementations, 1992.

Picard:1994:PDF

[Pic94] J. Picard. Portage d’applications Fortran sous Windows. Travaux scientifiques des chercheurs du Service de santé des armées durant l’année, 15(??):285–??, 1994. ISSN 0243-7473.

Pifer:1996:WDO

[Pif96] Joe Pifer. Are we doing our students a favor by making them learn Fortran? Computers in physics, 10(6):509–??, November 1996. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/
10.1063/1.4822490.

Plesinger:1993:FIT

[Ple93]  A. Plesinger.  Fast inver-

sion of transient calibration responses of arbitrary digital seismograph systems. Computers and Geosciences, 19 (2):189–201, 1993. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Picano:1993:PSA

[PMBH93] Silvio Picano, David G. Meyer, Eugene D. Brooks, III, and Joseph E. Hoag. Performance and scalability aspects of directory-based cache coherence in shared-memory multiprocessors. In Proceedings of the 1993 International Conference on Parallel Processing, volume II — Software, pages II–103–II–106. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, August 1993.

Pan:1992:FPT

[PMHC92] Guocheng Pan, Ken Moss, Tim Heiner, and James R. Carr. A FORTRAN program for three-dimensional cokriging with case demonstration. Computers and Geosciences, 18(5): 557–??, June 1992. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Pase:1993:MFP

[PMM93] Douglas M. Pase, Tom Mac-Donald, and Andrew Meltzer. MPP Fortran programming model. Technical report, Cray Research, Inc., Minneapolis, MN, October 11, 1993.

                                        URL ftp://ftp.cray.com/
                                        product-info/program_env/
                                        program_model.html.

Pase:1994:CFP

[PMM94] Douglas M. Pase, Tom Mac-Donald, and Andrew Meltzer. CRAFT Fortran programming model. Scientific Programming, 3(3):227–253, Fall 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pletzer:2008:EFD

[PMM^{+}08] Alexander Pletzer, Douglas McCune, Stefan Muszala, Srinath Vadlamani, and Scott Kruger. Exposing Fortran derived types to C and other languages. Computing in Science and Engineering, 10(4): 86–92, July/August 2008. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pohl:1997:CFP

[Poh97] Ira Pohl. C++ for Fortran Programmers. Addison-Wes-ley, Reading, MA, USA, July 1, 1997. ISBN 0-201-92483-8. 582 pp. US$39.76. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=0201924838.

Ponnusamy:1994:SIDa

[Pon94a] Ravi Ponnusamy. Supporting irregular distributions in FORTRAN 90D/HPF compilers. Computer science technical report series CS-TR-3268, University of Maryland, College Park, MD, USA, May 1994. 20 pp.

Ponnusamy:1994:SIDb

[Pon94b] Ravi Ponnusamy. Supporting irregular distributions in FORTRAN 90D/HPF compilers. Computer science technical report series CS-TR-3268.1, University of Maryland, College Park, MD, USA, November 1994. 25 pp.

Porscha:1990:MES

[Por90] W. Porscha. Maskenorientierte Eingabe zur Steuerung von geologischen Programmen. Studienarbeit, 1990.

ACM:1993:FAS

[PPP93] Fourth ACM SIGPLAN Symposium on Principles and Practice of Parallel Programming, volume 28(7). ACM Press, New York, NY 10036, USA, July 1993. CODEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN ????

Pardalos:1997:AFS

[PPR97] Panos M. Pardalos, Leonidas S.

Pitsolulis, and Mauricio G. C. Resende. Algorithm 769: Fortran subroutines for approximate solution of sparse quadratic assignment problems using GRASP. ACM Transactions on Mathematical Software, 23(2):196–208, June 1997. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1997-23-2/p196-pardalos/.

Palmer:1994:WND

[PPW94] D. W. Palmer, J. F. Prins, and S. Westfold. Work-efficient nested data-parallelism. In IEEE [IEE94a], pages 186–193. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Parsons:1994:RRT

[PQ94] R. Parsons and D. Quinlan. Run-time recognition of task parallelism within the P++ parallel array class library. In IEEE [IEE94e], pages 77–86. ISBN 0-8186-4980-1. LCCN QA76.58.S34 1993.

Padberg:1991:BCA

[PR91] Manfred Padberg and Giovanni Rinaldi. A branch-and-cut algorithm for the resolution of large-scale symmetric traveling salesman problems. SIAM Review, 33(1):60–100, March 1991. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Prasad:1990:IUO

[Pra90] B. S. V. Prasad. An input utility for older Fortran programs. Computers in industry, 13(4):361–366, 1990. CO-DEN CINUD4. ISSN 0166-3615 (print), 1872-6194 (electronic).

Press:1992:NRFb

[Pre92a] William H. Press. Numerical recipes in FORTRAN, 1992. ISBN 0-521-43719-9.

Press:1992:NRFc

[Pre92b] William H. Press. Numerical recipes in FORTRAN, 1992. ISBN 0-521-43717-2.

Press:1992:NRFd

[Pre92c] William H. Press. Numerical recipes in FORTRAN, 1992. ISBN 0-521-43719-9.

Press:1992:NRFe

[Pre92d] William H. Press. Numerical recipes in FORTRAN, 1992.

Prentice:1993:ATF

[Pre93a]  John K. Prentice.  Auto-

matic translation of Fortran 77 to Fortran 90 using VAST-90. Fortran Journal, 5(4): ??, July/August 1993. ISSN 1060-0221. URL http:/

/www.google.com/search?q=
cache:www.unics.uni-hannover.
de/rrzn/gehrke/vast_paper.
html+%22Fortran+Journal%
22&hl=en;   http://www.unics.
uni-hannover.de/rrzn/gehrke/
vast_paper.html.

Prentice:1993:PSV

[Pre93b] John K. Prentice. Pacific-Sierra VAST/77to90. Fortran Journal, 5(3):??, May/ June 1993. ISSN 1060-0221.

URL http://www.psrv.com/
vast77to90.html.

Prentice:1993:PBS

[Pre93c] John K. Prentice. A performance benchmark study of Fortran 90 compilers. Fortran Journal, 5(3):2–7, May/June 1993. ISSN 1060-0221.

Press:1993:NRFc

[Pre93d] William H. Press. Numerical Recipes: Fortran Set. Cambridge University Press, Cambridge, UK, second edition, January 1993. ISBN 0-521-44610-4. US$95.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0521446104;   http:/
/www.nr.com/nronline_switcher.

html. Includes 3.5in floppy disk.

Press:1993:NRFa

[Pre93e] William H. Press. Numerical recipes in FORTRAN. Cambridge University Press, Cambridge, UK, 2nd version 2.0 edition, 1993. ISBN 0-521-43716-4 (diskette). 1 computer disk pp. LCCN QA76.73.C15 N865 1992.

Press:1993:NRFb

[Pre93f] William H. Press. Numerical recipes in FORTRAN. Cambridge University Press, Cambridge, UK, 2nd, v2.0 edition, 1993. ISBN 0-521-43716-4. 1 computer disk pp. LCCN QA297.N866 1992.

Pressman:1993:FTT

[Pre93g] Michael H. Pressman. Fortran for today and tomorrow. Wm. C. Brown Publishers, Dubuque, IA, USA, 1993. ISBN 0-697-04483-1. xxii + 485 pp. LCCN QA76.73.F25 P717 1993.

Prentice:1994:U

[Pre94a]  John K. Prentice.  Un-

known. Fortran Journal, 6(??): ??, November/December 1994. ISSN 1060-0221.

Press:1994:NRF

[Pre94b] William H. Press. Numerical recipes in FORTRAN: the art of scientific computing. Cambridge University Press, Cambridge, UK, 2d, corrected to software version 2.04 edition, 1994. ISBN 0-521-43064-X (book), 0-521-43719-9 (FORTRAN diskette IBM 3.5, 720K). xxvi + 963 pp. LCCN QA76.73.C15 N865 1992.

Preston:1999:NTS

[Pre99] Nicky Williams Preston. New type signatures for legacy Fortran subroutines. ACM SIG-SOFT Software Engineering Notes, 24(5):76–85, September 1999. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pria:1993:MPI

[Pri93] A. Pria. ‘Mathematica’, a powerful informatics environment for the solution of engineering problems. Revista española de electronica, June 1993. CO-DEN RVEEBT. ISSN 0482-6396.

Priestley:2010:SOM

[Pri10] Mark Priestley. A Science of Operations: Machines, Logic and the Invention of Programming. History of computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-84882-554-4 (hardcover), 1-84882-555-2 (e-book). ix + 341 pp. LCCN QA76.6 .P737 2010.

Papamichael:1999:CMF

[PRS99] N. (Nicolas) Papamichael, Stephan Ruscheweyh, and E. B. Saff, editors. Computational methods and function theory 1997: proceedings of the Third CMFT Conference, 13–17 October 1997, Nicosia, Cyprus, volume 11 of Series in approximations and decompositions. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1999. ISBN 981-02-3626-3. LCCN QA297 .I473 1997.

Pryce:1999:TPS

[Pry99] J. D. Pryce. A test package for Sturm–Liouville solvers. ACM Transactions on Mathematical Software, 25(1):21–57, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Preppernau:1996:FPP

[PS96] Bryan L. Preppernau and Isaac Shokair. Fortran Powerstation 4.0 ports Fortran code smoothly to the PC. Computers in physics, 10(3):275–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/
10.1063/1.4822402.

Pal:2008:FPS

[PS08] Hridis Kumar Pal and Alok Shukla. A Fortran 90 program to solve the Hartree– Fock equations for interacting spin-1/2 fermions confined in harmonic potentials. Computer Physics Communications, 179(4):267–274, August 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465508001446.

Ponnusamy:1993:DRS

[PSC^{+}93a] R. Ponnusamy, J. Saltz, A. Choudhary, R. Das, and D. Mavriplis. The Dybbuk runtime system. In IEEE [IEE93a], pages 205–212. ISBN 0-8186-3400-6. LCCN QA75.5.C58 1993. IEEE catalog no. 93CH3251-6.

Ponnusamy:1993:RCT

[PSC93b] Ravi Ponnusamy, Joel Saltz, and Alok Choudhary. Runtime compilation techniques for data partitioning and communication schedule reuse. In IEEE [IEE93d], pages 361–370. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Ponnusamy:1995:RSC

[PSC^{+}95] R. Ponnusamy, J. Saltz, A. Choudhary, Yuan-Shin Hwang, and G. Fox. Runtime support and compilation methods for user-specified irregular data distributions. IEEE Transactions on Parallel and Distributed Systems, 6(8):815– 831, August 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pan:2003:SHI

[PSG03] Yi Pan, Joseph J. S. Shang, and Minyi Guo. A scalable HPF implementation of a finite-volume computational electromagnetics application on a CRAY T3E parallel system. Concurrency and Computation: Practice and Experience, 15(6):607–621, May 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Parthasarathy:1994:SSF

[PSPE94] S. Parthasarathy, K. Sekar, M. N. Ponnuswamy, and N. Elango. STATCW and STATCN: Fortran programs for conducting statistical tests for centrosymmetry. Journal of Applied Crystallography, 27(4): 653–??, 1994. CODEN JAC-GAR. ISSN 0021-8898.

Papazachos:1993:FPC

[PT93] C. B. Papazachos and G. N.

Tsokas. A FORTRAN program for the computation of 2-dimensional inverse filters in magnetic prospecting. Computers and Geosciences, 19(5): 705–??, May 1993. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Press:1996:NRFa

[PTM96] William H. Press, Saul A.

Teukolsky, and Michael Metcalf. Numerical Recipes in Fortran 90: The Art of Parallel Scientific Computing, volume 2 of Fortran Numerical Recipes. Cambridge University Press, Cambridge, UK, second edition, September 1, 1996. ISBN 0-521-57439-0 (hardcover). 200 pp. LCCN QA76.73.F25 N85 1996. US$44.95; US$40.45.

URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0521574390;
http://www.nr.com/nronline_
switcher.html.

Patzay:1992:NSM

[PTS92] Gyorgy Patzay, B. Toth, and Istvan Szabo. Numerical simulation of multicomponent ion exchange co-current and counter-current elution for porous exchangers. Hungarian journal of industrial chemistry, 20(3):179–182, 1992. CODEN HJICAI. ISSN 0133-0276.

Press:1996:NRFb

[PTV96] William H. Press, Saul A.

Teukolsky, and William T. Vetterling. Numerical Recipes in Fortran 77 and Fortran 90: The Art of Scientific and Parallel Computing. Cambridge University Press, Cambridge, UK, September 1996. ISBN 0-521-57440-

4. US$39.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0521574404;   http:/
/www.nr.com/; http://www.
nr.com/nronline_switcher.
html.

Press:1992:NRFa

[PTVF92] William H. Press, Saul A. Teukolsky, William T. Vetterling, and Brian P. Flannery. Numerical recipes in FORTRAN: the art of scientific computing. Cambridge University Press, Cambridge, UK, second edition, October 1, 1992. ISBN 0-521-43064-X (book), 0-521-43721-0 (example book) 0-521-43717-2 (diskette), 0-521-43719-9 (diskette), 0-521-43716-4 (diskette). xxvi + 963 pp. LCCN QA76.73.C15 N865 1992. US$39.95; US$54.95.

URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=052143064X;
 http://www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0521437164;   http:
//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0521437172;   http:
//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&

isbn=0521437199. Includes

MacIntosh disk.

Pugh:1990:EFR

[Pug90] Kenneth Pugh. Emulating Fortran’s ‘repeat’ format construct. C Users Journal, 8(3): 36–??, March 1990. ISSN 0898-9788.

Pugh:1994:MFC

[Pug94] Kenneth Pugh. Moving from FORTRAN to C. C Users Journal, 12(6):75–??, June 1994. ISSN 0898-9788.

Parkinson:1984:CAG

[PW84] D. Parkinson and M. Wunderlich. A compact algorithm for Gaussian elimination over GF(2) implemented on highly parallel computers. Parallel Computing, 1(1):65–73, August 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Patel:1993:FPS

[PW93] M. K. Patel and J. P. Waterhouse. A FORTRAN program for the statistical analysis of incomplete time series data sets by a method of partition. Computers in Biology and Medicine, 23(2):155–165, March 1993. CODEN CBM-DAW. ISSN 0010-4825.

Procassini:1993:PGO

[PWD93] Richard J. Procassini, Scott R.

Whitman, and William P. Dannevik. Porting a global ocean model onto a shared-memory multiprocessor: Observations and guidelines. The Journal of supercomputing, 7(3):287– 321, September 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=7&issue=3&spage=
287.

Perrott:1993:LPD

[PZA93] R. H. Perrott and Adib Zarea-Aliabadi. Languages for programming distributed memory systems. Computing and Control Engineering Journal, 4(6): 269–276, December 1993. CO-DEN CCEJEL. ISSN 0956-3385 (print), 1741-0460 (electronic).

Pachucki:2016:HFS

[PZY16] K. Pachucki, M. Zientkiewicz, and V. A. Yerokhin. H2SOLV: Fortran solver for diatomic molecules in explicitly correlated exponential basis. Computer Physics Communications, 208(??):162–168, November 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465516302120.

Qiang:2000:FIO

[QRH00] J. Qiang, R. D. Ryne, and S. Habib. Fortran implementation of object-oriented design in parallel beam dynamics simulations. Computer Physics Communications, 133 (1):18–33, December 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Queisser:2000:CRW

[Que00] Andrew Queisser. Classes for reading and writing parameter blocks. C/C++ Users Journal, 18(10):22–??, October 2000. CODEN CCUJEX. ISSN 1075-2838.

Ruhl:1990:PFC

[RA90] Roland Ruhl and Marco Annaratone. Parallelization of FORTRAN code on distributed-memory parallel processors. ACM SIGARCH Computer Architecture News, 18(3):342–353, September 1990. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Raghavachari:1995:BRH

[Rag95] Mukund Raghavachari. Book review: The High Performance Fortran Handbook by Charles Koelbel, David Loveman, Robert Schreiber, Guy Steele Jr., and Mary Zosel. ACM SIGPLAN Notices, 30 (7):5, July 1995. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rajendran:1995:FPC

[Raj95] S. Rajendran. A FORTRAN program for computing the geometric properties of plane lamina and axi-symmetric bodies. Computers and structures, 54(5):859–??, 1995. CODEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Ramsay:1990:MFS

[Ram90] James O. Ramsay. MATFIT: A FORTRAN subroutine for comparing two matrices in a subspace. Psychometrika, ?? (3):551–555, September 1990. CODEN PSMIAX. ISSN 0033-3123.

Rappoldt:1990:RMF

[Rap90] C. Rappoldt. Reference manual of the FORTRAN utility library TTUTIL with applications. Simulation reports cabott; nr. 20, Centre for Agrobiological Research (CABO-DLO) and Dept. of Theoretical Production Ecology (TPE), Agricultural University, Wageningen, The Netherlands, 1990. 122 pp.

Raportirenko:1994:GPS

[Rap94] A. M. Raportirenko. GSL: a portable standard Lisp interpreter. In Becks and Perret-Gallix [BPG94], pages 645– 650. ISBN 981-02-1699-8. LCCN QC793.47.E4I58 1993.

Ratzer:1995:FA

[Rat95] Gerald F. G. Ratzer. Fortran 90 and algorithms. Kendall/Hunt Pub., Dubuque, IA, USA, 1995. ISBN 0-7872-1566-X. 307 pp. LCCN QA76.73.F25R36

Rouson:2010:DPM

[RAX10] Damian W. I. Rouson, Helgi Adalsteinsson, and Jim Xia. Design patterns for multiphysics modeling in Fortran 2003 and C++. ACM Transactions on Mathematical Software, 37(1):3:1–3:30, January 2010. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Renka:1998:RA

[RB98] Robert J. Renka and Ron Brown. Remark on Algorithm

761. ACM Transactions on Mathematical Software, 24(4): 383–385, December 1998. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Aki96, DV00].

Renka:1999:AAT

[RB99] R. J. Renka and Ron Brown.

Algorithm 792: Accuracy tests of ACM algorithms for interpolation of scattered data in the plane. ACM Transactions on Mathematical Software, 25(1): 78–94, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL ftp://netlib.bell-
labs.com/netlib/toms/792.
gz;   http://doi.acm.org/
10.1145/305658.305745;
http://phase.etl.go.jp/netlib/
toms/792;   http://www.acm.
org/pubs/citations/journals/
toms/cgi-bin/TOMSbibget?
Renka:1999:AAT; http://www.
acm.org:80/pubs/citations/
journals/toms/1999-25-1/

                                        p78-renka/;   http://www.
                                        hensa.ac.uk/netlib/toms/
                                        792.gz;   http://www.netlib.
                                        no/netlib/toms/792; http:
                                        //www.netlib.org/toms/792.

Rao:2010:AGM

[RBD^{+}10] Anil V. Rao, David A. Benson, Christopher Darby, Michael A. Patterson, Camila Francolin, Ilyssa Sanders, and Geoffrey T. Huntington. Algorithm 902: GPOPS, a MATLAB software [sic] for solving multiplephase optimal control problems using the Gauss pseudospectral method. ACM Transactions on Mathematical Software, 37(2):22:1–22:39, April 2010. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See corrigendum [RBD^{+}11].

Rao:2011:CAG

[RBD^{+}11] Anil V. Rao, David A. Benson, Christopher Darby, Michael A. Patterson, Camila Francolin, Ilyssa Sanders, and Geoffrey T. Huntington. Corrigendum: Algorithm 902: GPOPS, a MATLAB software for solving multiple-phase optimal control problems using the Gauss pseudospectral method. ACM Transactions on Mathematical Software, 38(1):9:1–9:2, November 2011. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [RBD^{+}10].

Reale:1992:FPI

[RBS92] F. Reale, M. Barbera, and S. Sciortino. Fast parallel implementation of multidimensional data-domain FORTRAN codes on distributed-memory processor arrays. Computer Physics Communications, 72 (2):129–143, November 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Reale:1993:PDH

[RBS93a] F. Reale, M. Barbera, and S. Sciortino. A parallel 2-d hydrodynamic FORTRAN code for astrophysical applications on a Meiko computing surface. Future Generation Computer Systems, 9(1):19–24, May 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Reale:1993:PHF

[RBS93b] F. Reale, M. Barbera, and S. Sciortino. A parallel 2-D hydrodynamic FORTRAN code for astrophysical applications on a Meiko computing surface. Future Generation Computer Systems, 9(1):19–24, May 1993. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rostaing:1991:ATA

[RD91] Nicole Rostaing and Stephane Dalmas. Analyse et transformation automatiques de programmes Fortran a l’aide d’un langage fonctionnel type = automatic analysis and transformation of Fortran programs using a typed functional language. Rapports de recherche 1518, Institut National de Recherche en Informatique et en Automatique, Le Chesnay, France, 1991. 10 pp.

Rowlingson:1992:SSP

[RD92] B. S. Rowlingson and P. J. Diggle. Splancs: Spatial point pattern analysis code in S-Plus. In P. Bresnahan, E. Corwin, and D. Cowen, editors, Proceedings: 5th International Symposium on Spatial Data Handling: August 3–7, 1992, Charleston, SC, USA, volume 2, pages 713–722. Humanities and Social Sciences Computing Lab, University of South Carolina, Charleston, SC, USA, 1992. ISBN 0-9633532-0-9 (vol. 1), 0-9633532-1-7 (vol. 2), 0-9633532-2-5 (set). Two volumes.

Redwine:1995:UF

[Red95] Cooper Redwine. Upgrading to Fortran 90. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., September 1, 1995. ISBN 0-387-97995-6. xiv + 501 pp. LCCN QA 76.73 F25 R43 1995. US$44.95.

Reid:1992:AFV

[Rei92a] J. Reid. The advantages of Fortran 90. Die Vorzuge von Fortran 90. Computing, 48 (??):219–238, 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Reid:1992:VFG

[Rei92b] John Reid. The advantages of Fortran 90. Computing, 48(3): 219–238, September 1992. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Reid:1992:AF

[Rei92c] John K. Reid. The advantages of Fortran 90. Computing, 48 (3-4):219–238, 1992. CODEN COSPDM. ISSN 0010-485X (print), 1436-5057 (electronic).

Reichelt:1993:IFO

[Rei93] Eric R. Reichelt. Implementing Fortran ODE solver LSODE using MatLab. Thesis (M.S.), San Diego State University, San Diego, CA, USA, 1993. vii + 65 pp.

Reid:1995:EHFa

[Rei95a] John Reid. Exception handling in Fortran. ACM SIGPLAN FORTRAN Forum, 14(3):9–15, September 1995. ISSN 1061-7264 (print), 1931-1311 (electronic).

Reid:1995:EHFb

[Rei95b] John Reid. Exception handling in Fortran. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/

proceedings/reid/reid.txt.

Presented by Bo Einarsson.

Reid:1995:PME

[Rei95c] John Reid. Procedures in a module for exception handling. Technical report, ????, ????, November 1, 1995.

Reid:1996:RFF

[Rei96a] J. K. Reid. Remark on “Fast Floating-Point Processing in Common Lisp”. ACM Transactions on Mathematical Software, 22(4):496–497, December 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Reid:1996:AF

[Rei96b] John Reid. An appreciation of F. Fortran Journal, 8(6):3– 9, November/December 1996. ISSN 1060-0221. URL http:/

/www.fortran.com/fortran/
FJ/9611/#appreciation.

Reid:1997:TAE

[Rei97] J. K. Reid. Two approaches to exception handling in Fortran. In Boisvert [Boi97], pages 210–223. ISBN 0-412-80530-8. LCCN QA297.I35 1996. US$146.50.

Reid:2002:CAF

[Rei02] John Reid. Co-array Fortran for full and sparse matrices. Lecture Notes in Computer Science, 2367:61– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2367/23670061.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2367/23670061.
pdf.

Reid:2003:FF

[Rei03]  John Reid.  The future

of Fortran. Computing in Science and Engineering, 5(4):59–67, July/August 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2003/04/c4059.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/04/c4059.
pdf.

Reid:2004:OF

[Rei04] John Reid. An overview of Fortran 2003. ACM SIGPLAN Notices, 39(8):31–38, August 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Renka:1996:ATC

[Ren96a]  R. J. Renka.  Algorithm

751: TRIPACK: a constrained two-dimensional Delaunay triangulation package. ACM Transactions on Mathematical Software, 22(1):1–8, March 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [Ren99a].

Renka:1996:ASS

[Ren96b] R. J. Renka. Algorithm 752: SRFPACK: software for scattered data fitting with a constrained surface under tension. ACM Transactions on Mathematical Software, 22(1):9–17, March 1996. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [Ren99b].

Renka:1997:ASD

[Ren97a]  Robert J. Renka.  Algo-

rithm 772. STRIPACK: Delaunay triangulation and Voronoi diagram on the surface of a sphere. ACM Transactions on Mathematical Software, 23(3):416–434, September 1997. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1997-23-3/p416-renka/.

Renka:1997:ASI

[Ren97b]  Robert J. Renka.  Al-

gorithm 773. SSRFPACK: Interpolation of scattered data on the surface of a sphere with a surface under tension. ACM Transactions on Mathematical Software, 23(3):435–442, September 1997. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1997-23-3/p435-renka/.

Renka:1999:RAa

[Ren99a] R. J. Renka. Remark on Algorithm 751. ACM Transactions on Mathematical Software, 25 (1):97–98, March 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Ren96a].

Renka:1999:RAb

[Ren99b] R. J. Renka. Remark on Algorithm 752. ACM Transactions on Mathematical Software, 25 (1):99–100, March 1999. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Ren96b].

Renka:2003:ADD

[Ren03] Robert J. Renka. Algorithm 828: DNSPLIN1: discrete nonlinear spline interpolation. ACM Transactions on Mathematical Software, 29(4):458– 468, December 2003. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Renka:2004:ACI

[Ren04]  Robert J. Renka.  Algo-

rithm 833: CSRFPACK— interpolation of scattered data with a C^{1} convexity-preserving surface. ACM Transactions on Mathematical Software, 30(2): 200–211, June 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Renka:2009:ATT

[Ren09]  Robert J. Renka.  Algo-

rithm 893: TSPACK: tension spline package for curve design and data fitting. ACM Transactions on Mathematical Software, 36(1):7:1–7:8, March 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

RFC:1990:ESR

[RFC90]  Embedded Systems:  1990

Rochester Forth Conference. University of Rochester, Rochester, NY, USA, June 1990. ISBN ???? LCCN ????

Russ:1996:HAT

[RFRH96] S. H. Russ, B. Flachs, J. Robinson, and B. Heckel. Hector: automated task allocation for MPI. In IEEE [IEE96], pages 344–348. ISBN 0-8186-7255-2. LCCN QA76.58 .I565 1996. IEEE catalog number 96TB100038. IEEE Computer Society Press order number PR07255.

Resende:1998:AFS

[RFS98] Mauricio G. C. Resende, Thomas A. Feo, and Stuart H. Smith. Algorithm 787: Fortran subroutines for approximate solution of maximum independent set problems using GRASP. ACM Transactions on Mathematical Software, 24(4):386–394, December 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Reichel:1990:AFS

[RG90a] L. Reichel and W. B. Gragg. Algorithm 686: FORTRAN subroutines for updating the QR decomposition. ACM Transactions on Mathematical Software, 16(4):369–377, December 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0098-
3500/98291.html.  Cited

in Åke Björck’s bibliography on least squares, which is available by anonymous

ftp from math.liu.se in
pub/references.

Reichel:1990:FSU

[RG90b] L. Reichel and W. B. Gragg.

FORTRAN subroutines for updating the QR decomposition. ACM Transactions on Mathematical Software, 16:369–377, 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Rice:1984:ARK

[RH84] John R. Rice and Richard J.

Hanson. Algorithm 620: References and keywords for Collected Algorithms of the ACM. ACM Transactions on Mathematical Software, 10(4):359– 360, December 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [Ham85, HM90a].

Rogers:1994:MFP

[RH94] H. J. Rogers and R. K. Hambleton. MH: A FORTRAN 77 program to compute the Mantel– Haenszel statistic for detecting differential item functioning. Educational and Psychological Measurement, 54(1): 101–??, Spring 1994. CODEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Rhee:1993:FSE

[Rhe93] Kijong Rhee. A FORTRAN solution for evaluating the coefficients of the power method for nonnormal transformation. Educational and Psychological Measurement, 53(1):107– 110, Spring 1993. CODEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Rodriguez:1996:POW

[RHH96] Bernardo Rodriguez, Leslie Hart, and Tom Henderson. Parallelizing operational weather forecast models for portable and fast execution. Journal of Parallel and Distributed Computing, 37(2):159– 170, September 15, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jpdc.1996.0116/production;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0116/production/pdf.

Ribar:1992:FPW

[Rib92]  L. John Ribar.  FOR-

TRAN Programming for Windows. Osborne/McGraw-Hill, Berkeley, CA, USA, December 1992. ISBN 0-07-881908-3. US$29.95; US$26.95. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0078819083.

Ribeiro:2002:FCN

[Rib02] M. Ribeiro. A Fortran code for null geodesic solutions in the lemaître–tolman–bondi spacetime. Computer Physics Communications, 148(2):236–241, October 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Rice:1995:PSE

[Ric95] John Rice. Problem solving environments for scientific computing. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/
proceedings/rice/rice1.html.

Richardson:2006:TCP

[Ric06]  William Richardson.  A

treatment of computational precision, number representation, and large integers in an introductory Fortran course. Journal of Computers in Mathematics and Science Teaching, 25(3):213–237, 2006. CODEN JCMTDV. ISSN 0731-9258. URL http:

//www.editlib.org/index.
cfm?CFID=21448942&CFTOKEN=
41513643&fuseaction=Reader.
ViewAbstract&paper_id=6298.

Ritland:1990:SFC

[Rit90] K. Ritland. A series of FOR-TRAN computer programs for estimating plant mating systems. The Journal of Heredity, 81(3):235–??, May 1990. CO-DEN JOHEA8. ISSN 0022-1503.

Reddy:1992:MFP

[RKMJ92] P. J. Reddy, D. Krishna, U. Suryanarayana Murty, and K. Jamil. A microcomputer FORTRAN program for rapid determination of lethal concentrations of biocides in mosquito control. Computer applications in the biosciences: CABIOS, 8 (3):209–213, June 1992. CO-DEN COABER. ISSN 0266-7061.

Reese:1991:OOF

[RL91] Donna S. Reese and Ed Luke. Object oriented Fortran for development of portable parallel programs. In Proceedings of the 3rd IEEE Symposium on Parallel Distributed Processing, pages 608–615. ????, Dallas, TX, USA, December 1991.

Reid:2020:HCS

[RLS20] John Reid, Bill Long, and Jon Steidel. History of coarrays and SPMD parallelism in Fortran. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):72:1– 72:30, June 2020. URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3386322.

Rafelski:1990:PFP

[RM90] H. E. Rafelski and B. Müller.

PC-Fortran programs for muon reactivation calculations in muon-catalyzed fusion. Computer Physics Communications, 59(3):521–526, July 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
001046559090093G.

Roth:1997:CSH

[RMCKB97] Gerald Roth, John Mellor-Crummey, Ken Kennedy, and R. Gregg Brickner. Compiling stencils in High Performance Fortran. In ACM [ACM97], page ?? ISBN 0-89791-985-8. LCCN ???? URL http:

//www.supercomp.org/sc97/
proceedings/TECH/ROTH/INDEX.

HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Rouson:2005:DMA

[RMX05] Damian W. I. Rouson, Karla Morris, and Xiaofeng Xu. Dynamic memory de-allocation in Fortran 95/2003 derived type calculus. Scientific Programming, 13(3):189–203, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Rouson:2012:IYP

[RMX12] Damian Rouson, Karla Morris, and Jim Xia. This isn’t your parents’ Fortran: Managing C++ objects with modern Fortran. Computing in Science and Engineering, 14(2): 46–54, March/April 2012. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Reid:2007:CAN

[RN07] John Reid and Robert W. Numrich. Co-arrays in the next Fortran Standard. Scientific Programming, 15(1):9–26, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Robinet:19xx:EDL

[Robxx] Bernard Robinet. Etude d’un langage algorithmique: FOR-TRAN. Institut de programmation, Paris, France, 19xx. various pp.

DiRoccaferrera:19xx:NFI

[Rocxx] Giuseppe Maria Ferrero Di Roccaferrera. Notes on Fortran IV language. Technical report, Quantitative Methods Dept., College of Business Administration Syracuse University, Syracuse, NY, USA, 19xx. 118 pp.

Rodriguez:1990:VTP

[Rod90] Brad Rodriguez. VECTOR-FORTH — programming an array processor in Forth. In Forsley and Hess [FH90], pages 128–129. ISBN 0-914593-10-2. LCCN QA76.73.F24 R59 1990.

US$35.00. Spine title: Embedded systems. Includes abstracts and bibliographical references.

Ross:1993:CCF

[Ros93] John W. Ross. Calling C functions with variably dimensioned arrays. Dr. Dobb’s Journal of Software Tools, 18(8):52, 54, 56, August 1993. CODEN DDJOEB. ISSN 1044-789X.

Roth:1993:OFP

[Rot93] Gerald Roth. Optimizing Fortran 90D programs for SIMD execution. Technical report COMP TR93-213, Rice University, Dept. of Computer Science, Houston, TX, USA, April 1993. 19 pp.

Roth:19xx:SMO

[Rotxx]  Peter N. Roth.  Storage

management in object based Fortran77. Fortran Journal, ??(??):??, ???? 19xx. ISSN 1060-0221. URL http:

//www.engineeringobjects.
com/zips/smoopftn.zip.

Rouse:1990:IF

[Rou90] Rouse. Introduction to Fortran 77. Holt, Reinhart, and Winston, New York, NY, USA, October 1990. ISBN 0-03-069394-2. US$26.53. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0-03-069394-2.

Robinson:1993:CFP

[RP93] Stuart Robinson and Martin Prime. C for Fortran programmers. Rutherford Appleton Laboratory, Chilton, Didcot, Berks, UK, 1993. ISBN 1-85889-051-9. ???? pp.

Rauchwerger:1995:LTS

[RP95] Lawrence Rauchwerger and David Padua. The LRPD test: speculative run-time parallelization of loops with privatization and reduction parallelization. ACM SIGPLAN Notices, 30(6):218–232, June 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/207110/p218-rauchwerger/

.

Radul:2012:AFI

[RP12] Alexey Radul and Barak A. Pearlmutter. AD in Fortran: Implementation via prepreprocessor. In Forth et al. [FHP^{+}12], pages 273–284. CO-DEN LNCSA6. ISBN 3-642-30022-7 (print), 3-642-30023-5 (e-book). ISSN 1439-7358. LCCN ???? URL

http://link.springer.com/
content/pdf/10.1007/978-
3-642-30023-3_25. Proceed-

ings of the Sixth International Conference on Automatic Differentiation (AD2012) held July 23–27, 2012, in Fort Collins, Colorado, USA.

Roy:2020:FFT

[RPG^{+}20] Arko Roy, Sukla Pal, S. Gautam, D. Angom, and P. Muruganandam. FACt: FOR-TRAN toolbox for calculating fluctuations in atomic condensates. Computer Physics Communications, 256(??):Article 107288, November 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465520300989.

Resende:1996:AFS

[RPL96] Mauricio G. C. Resende, Panos M. Pardalos, and Yong Li. Algorithm 754: Fortran subroutines for approximate solution of dense quadratic assignment problems using GRASP. ACM Transactions on Mathematical Software, 22(1):104–118, March 1996. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0098-3500/
225553.html.

Reed:1992:CWK

[RR92] J. F. Reed, 3d and J. J. Reed.

Cohen’s weighted kappa with Turbo Pascal (FORTRAN). Computer Methods and Programs in Biomedicine, 38(2– 3):153–165, July 1992. CO-DEN CMPBEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Ribar:1993:FPW

[RR93] L. John Ribar and John Ribar. Fortran Programming for Windows. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, January 1, 1993. ISBN 0-07-881908-3. xiii + 238 pp. LCCN QA76.73.F25R49 1993. US$29.95.

Ribeiro:1999:AFS

[RR99] Celso C. Ribeiro and Mauricio G. C. Resende. Algorithm 797: Fortran subroutines for approximate solution of graph planarization problems using GRASP. ACM Transactions on Mathematical Software, 25(3):341–352, September 1999. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Radhakrishnan:2015:UCP

[RRM^{+}15] Hari Radhakrishnan, Damian W. I. Rouson, Karla Morris, Sameer Shende, and Stavros C. Kassinos. Using coarrays to parallelize legacy Fortran applications: Strategy and case study. Scientific Programming, 2015(??):904983:1– 904983:12, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2015/904983/.

Rouson:2008:GFA

[RRX^{+}08] Damian W. I. Rouson, Robert Rosenberg, Xiaofeng Xu, Irene Moulitsas, and Stavros C. Kassinos. A grid-free abstraction of the Navier–Stokes equations in Fortran 95/2003. ACM Transactions on Mathematical Software, 34(1):2:1–2:33, January 2008. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Rao:1992:NMPa

[RS92a] S. Balachandra Rao and C. K.

Shantha. Numerical Methods With Programs in Basic, Fortran and Pascal. Stosius Inc/Advent Books Division, London, UK, December 1992. ISBN 0-86311-370-2 (paperback). xiii + 446 pp. US$20.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0863113702.

Rao:1992:NMPb

[RS92b] S. Balachandra Rao and C. K.

Shantha. Numerical methods: with programs in BASIC, FORTRAN and Pascal. Universities Press (India), Hyderabad, India, 1992. ISBN 0-86311-370-2. xiii + 446 pp.

Rabenseifner:1993:CDR

[RS93] R. Rabenseifner and A. Schuch.

Comparison of DCE RPC, DFN-RPC, ONC and PVM. In Schill [Sch93a], pages 39– 46. ISBN 3-540-57306-2, 0-387-57306-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C55I58 1993.

Reid:2009:AFV

[RS09a] John K. Reid and Jennifer A.

Scott. Algorithm 891: a Fortran virtual memory system.

ACM Transactions on Mathematical Software, 36(1):5:1– 5:12, March 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Reid:2009:CSC

[RS09b] John K. Reid and Jennifer A. Scott. An out-of-core sparse Cholesky solver. ACM Transactions on Mathematical Software, 36(2):9:1–9:33, March 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Ramaswamy:1997:FET

[RSB97] S. Ramaswamy, S. Sapatnekar, and P. Banerjee. A framework for exploiting task and data parallelism on distributed memory multicomputers. IEEE Transactions on Parallel and Distributed Systems, 8(11):1098– 1116, November 1997. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1997/pdf/
l1098.pdf;   http://www.
computer.org/tpds/td1997/
l1098abs.htm.

Ralston:1990:FPS

[RTY90] B. Ralston, F. Thomas, and F. Yeung. Flood prediction — a study in Fortran optimization and connectivity. Supercomputer, 7(4):48–51, July 1990. CODEN SPCOEL. ISSN 0168-7875.

Ruby:1993:FEM

[Rub93] Lawrence Ruby. Fortran 90 Explained by Michael Metcalf and John Reid and FOR-TRAN 77 for Engineers and Scientists: Includes Preview of FORTRAN 90, 3rd ed. by Larry Nyhoff and Sanford Leestma. American Journal of Physics, 61(4):383–??, April 1993. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Renes:1992:MGC

[RVV^{+}92] W. A. Renes, M. Vanbegin, P. Van Dooren, J. W. J. Beckers, and J. The MATLAB gateway compiler. A tool for automatic linking of FORTRAN routines to MATLAB. In Barker [Bar92], pages 95–100. ISBN 0-08-041269-6. LCCN TJ213 .C57 1992.

Rosing:1999:PPP

[RY99] Matthew Rosing and Steve B.

Yabusaki. A programmable preprocessor for parallelizing Fortran90. In ACM [ACM99], page ??

Ryskin:1995:BFB

[Rys95] Alexander Ryskin. Bookshelf:

Fortran becomes a language for the nineties: review of Migrating to Fortran 90. IEEE Software, 12(4):110–111, July 1995. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Reddy:1994:F

[RZ94a] Rama N. Reddy and Carol A. Ziegler. Fortran 77 With 90. West Information Publishing Group, ????, second edition, January 1994. ISBN 0-314-48135-4. LCCN QA76.73.F25R42 1989. US$51.75.

URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0314481354.

Reddy:1994:FAS

[RZ94b] Rama N. Reddy and Carol A. Ziegler. Fortran 77 With 90: Applications for Scientists and Engineers. West Publishing Company, St. Paul, MN, USA, second edition, January 1994. ISBN 0-314-02861-7. xvi + 640 pp. LCCN QA76.73.F25R42 1994. US$53.25. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0314028617.

Sabot:1992:OCF

[Sab92] Gary Sabot. Optimized CM Fortran compiler for the Connection Machine computer. Technical Report TR-219, Thinking Machines Corporation, 245 First Street, Cambridge, MA 01241, 1992. (Proceedings of the 25th Hawaii International conference on system sciences, pp.162-172 (Kauai Hawaii: January 7-10, 1992)).

Sabot:1994:OCF

[Sab94] G. W. Sabot. Optimizing CM Fortran compiler for Connection Machine computers. Journal of Parallel and Distributed Computing, 23(2):224– 238, November 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1994.1134/production;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1994.1134/production/pdf.

Sabot:1995:HPC

[Sab95] Gary W. Sabot. High Performance Computing: Problem Solving With Parallel and Vector Architectures. Addison-Wesley, Reading, MA, USA, May 1995. ISBN 0-201-59830-

2. xvi + 246 pp. LCCN QA76.5 .H4817 1995. US$46.21.

Stevenson:1992:VCF

[SAC^{+}92] D. E. Stevenson, L. K.

Ammons, W. G. Crosmun, A. Jackson, and G. L. Raj. A vector C and Fortran compiler for the FPS T-series: Experiences with compiling to occam I. Software—Practice and Experience, 22(5):371–390, May 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Saini:1995:NEP

[Sai95] Subhash Saini. NAS experiences of porting CM Fortran codes to HPF on IBM SP2 and SGI Power Challenge. Technical report NAS-95-010, NASA-CR 203382, National Technical Information Service, Washington, DC, USA, 1995. ?? pp.

Salemi:1992:LPC

[Sal92] Joe Salemi. Locating PC-based compilers for PL/1, FOR-TRAN, and COBOL; improving performance on slower or older machines. PC Magazine, 11(13):509–516, July 1992. CODEN PCMGEP. ISSN 0888-8507.

SS:1995:KTG

[Sal95] Salford Software. KB0013: Taming the GUI dragon. Fortran Journal, 7(1):??, January/February 1995. ISSN 1060-0221. URL http:/

                                        /www.suns.salford.ac.uk/
                                        support/compilers/kb/skb0013.
                                        shtml.

Sala:2006:OOF

[Sal06]  Marzio Sala.  An object-

oriented framework for the development of scalable parallel multilevel preconditioners. ACM Transactions on Mathematical Software, 32(3):396– 416, September 2006. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Santavicca:1992:FMT

[San92] Jeffery W. Santavicca. Fluid mechanics tutorials in GKS supported FORTRAN. Thesis (M.S.), Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1992. xi + 211 pp.

Sarkar:1991:PIP

[Sar91] V. Sarkar. PTRAN — the IBM parallel translation system. In B. K. Szymanski, editor, Parallel Functional Languages and Compilers, pages 309–392. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-52243-8.

Sarafin:1994:WT

[Sar94] Gregory C. Sarafin. The Wrapi toolkit. Dr. Dobb’s Journal of Software Tools, 19(3):26, 28, 30, 32, 34, March 1994. CO-DEN DDJOEB. ISSN 1044-789X.

Sarkar:1997:ASH

[Sar97] V. Sarkar. Automatic selection of high-order transformations in the IBM XL FOR-TRAN compilers. IBM Journal of Research and Development, 41(3):233–??, ???? 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
almaden.ibm.com/journal/
rd/413/sarkar.html.

Sarkadi:2000:FPC

[Sar00] L. Sarkadi. A Fortran program to calculate the matrix elements of the Coulomb interaction involving hydrogenic wave functions. Computer Physics Communications, 133(1):119– 127, December 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Sarkadi:2017:FPC

[Sar17] L. Sarkadi. A Fortran program to calculate the matrix elements of the Coulomb interaction involving hydrogenic wave functions. Computer Physics Communications, 212(??):280–282, March 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S001046551630337X.

Smith:1990:IDA

[SAS90] Kevin Smith, Bill Appelbe, and Kurt Stirewalt. Incremental dependence analysis for interactive parallelization. ACM SIGARCH Computer Architecture News, 18(3):330–341, September 1990. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). Proceedings 1990 International Conference on Supercomputing, ACM SIGARCH Computer Architecture News.

Sattin:1997:LFP

[Sat97] Fabio Sattin. Lyap: a Fortran 90 program to compute the Lyapunov exponents of a dynamical system from a time series. Computer Physics Communications, 107 (1–3):253–257, December 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465597001240.

Savage:1995:SFO

[Sav95] M. Savage. SEEK: a Fortran optimization program using a feasible directions gradient search. NASA contractor report 4641, Army Research Laboratory technical report ARL-TR-645, National Aeronautics and Space Administration, Office of Management, Scientific and Technical Information Program; National Technical Information Service, distributor, Washington, DC, USA, 1995. ???? pp.

Silver:1991:FPT

[SB91] N. Clayton Silver and Robert T.

Burkey. A FORTRAN 77 program for testing the differences among independent correlations. Educational and Psychological Measurement, 51(3): 641–??, Fall 1991. CODEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Smith:1992:GSF

[SB92] George K. Smith and Charles R.

Bauer. Getting started with Fortran 77. Four Pi, Norridge, IL, USA, 1992. ISBN ???? 181 pp. LCCN ????

Somerville:2001:FSI

[SB01] Paul N. Somerville and Frank Bretz. FORTRAN 90 and SAS-IML programs for computation of critical values for multiple testing and simultaneous confidence intervals. Journal of Statistical Software, 6(5):1– 14, 2001. CODEN JSSOBK.

ISSN ???? URL http://www.
jstatsoft.org/v06/i05;
http://www.jstatsoft.org/
v06/i05/EXAMPLE/;   http:
//www.jstatsoft.org/v06/
i05/FORTRAN/;   http://www.
jstatsoft.org/v06/i05/jss2r.
pdf;   http://www.jstatsoft.
org/v06/i05/SAS/;   http:
//www.jstatsoft.org/v06/
i05/updates.

Shterenlikht:2019:MVF

[SC19] Anton Shterenlikht and Luis Cebamanos. MPI vs Fortran coarrays beyond 100k cores: 3D cellular automata. Parallel Computing, 84(??):37– 49, May 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167819118303181.

Schlichting:1990:NFLb

[Sch90] J. J. F. M. Schlichting. NUMVEC FORTRAN library manual. Report. Centrum voor Wiskunde en Informatica NM-9017, Stichting Mathematisch Centrum, Amsterdam, The Netherlands, 1990. 3 + 6 pp.

Schildt:1991:ACE

[Sch91a] Herbert Schildt. The art of C: elegant programming solutions. Osborne/McGraw-Hill, Berkeley, CA, USA, September 1, 1991. ISBN 0-07-881691-2. xvi + 459 pp. LCCN QA 76.73 C15 S3423 1991. US$39.95. System requirements for computer disk: PC; virtually any C compiler including Turbo C, Turbo C++, Microsoft C, and others that are ANSI C-compatible.

Schneck:1991:BRO

[Sch91b]  Paul B. Schneck.  Book

review: Optimizing FOR-TRAN Programs by C. F. Schofield (Halstead Press, 1989). ACM SIGARCH Computer Architecture News, 19(1): 153–154, March 1991. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Schill:1993:DOD

[Sch93a] Alexander Schill, editor. DCE — the OSF distributed computing environment: client/server model and beyond: International DCE Workshop, Karlsruhe, Germany, October 7– 8, 1993: proceedings, number 731 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57306-2, 0-387-57306-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.C55I58 1993.

Schnabel:1993:WLC

[Sch93b] Robert B. Schnabel, editor.

Workshop on Languages, Compilers, and Run-Time Environments for Distributed Memory Multiprocessors: Boulder, Colorado, September 30–October 2, 1992: extended abstracts of talks and poster presentations, volume 28(1) of ACM SIG-PLAN Notices. ACM Press, New York, NY 10036, USA, January 1993. CODEN SIN-ODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7 .S54 v.28:1.

Schonfelder:1993:FAO

[Sch93c] Lawrie Schonfelder. Fortran 90 is already object oriented (well almost). Fortran Journal, 5(2): 4–6, ???? 1993. ISSN 1060-0221.

Schuster:1994:PPG

[Sch94] V. J. Schuster. PGHPF from the Portland Group. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):72, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Schreiber:1996:SIC

[Sch96a] R. Schreiber. Support for irregular computation in High Performance Fortran. Lecture Notes in Computer Science, 1117:285–??, ???? 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreiber:1996:IH

[Sch96b] R. S. Schreiber. An introduction to HPF. Lecture Notes in Computer Science, 1132:27– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreiber:1997:HPF

[Sch97] Robert Schreiber. High Performance Fortran, Version 2. Parallel Processing Letters, 7(4): 437–??, December 1997. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Schonfelder:1999:VPA

[Sch99] J. L. Schonfelder. Variable precision arithmetic: Version 1.1: a Fortran 95 module. World-Wide Web document., Decem-

ber 1999. URL http://pcwww.
liv.ac.uk/~jls/vpa20.f95;
http://pcwww.liv.ac.uk/~
jls/vpa20.htm.

Schurdak:19xx:AUC

[Schxx] John J. Schurdak. An Approach to the Use of Computers in the Instructional Process and Evaluation. ????, ????, 19xx. ISBN ???? 42 pp. LCCN ????

Schonfelder:2003:VPA

[Sch03] J. L. Schonfelder. Variable precision arithmetic: a Fortran 95 module. Scientific Programming, 11(1):67–76, 2003. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Schlittgen:2007:BRD

[Sch07] Rainer Schlittgen. Book review: David Lemmon and Joseph Schafer: Developing Statistical Software in Fortran 95. AStA. Advances in Statistical Analysis, 91(4):449– 450, December 2007. CO-DEN ???? ISSN 1863-8171 (print), 1863-818X (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s10182-007-
0050-1.

NewScientist:1992:T

[Sci92] New Scientist. Technology. World press review, 39(12):47– ??, December 1992. ISSN 0195-8895.

SCAI:1993:FRM

[Sci93] Scientific Computing Associates, Inc., New Haven, CT, USA. FORTRAN-Linda Reference Manual, 1993.

STI:19xx:UF

[Scixxa] Scientific Toolworks, Inc. Understand for Fortran. World-Wide Web document., 19xx.

URL http://www.scitools.

com/uf.html. From the vendor Web site: “Understand for FORTRAN is an interactive development environment (IDE) tool providing reverse engineering, automatic documentation, metrics and cross referencing of FORTRAN source code. It supports FOR-TRAN 77 (F77) and FOR-TRAN 90 (F9X) language standards, with common VAX and Cray extensions.”.

STI:19xx:UFU

[Scixxb] Scientific Toolworks, Inc., 1579 Broad Brook Rd, South Royalton,VT 05068, Tel: (802)

763-2995, FAX: (802)-763-3066. Understand for Fortran: User Guide and Reference Manual, Version 1.4, 19xx.

URL http://www.scitools.
com/manuals/latest/html/
understand_f/. From the

vendor Web site: “Understand for FORTRAN is an interactive development environment (IDE) tool providing reverse engineering, automatic documentation, metrics and cross referencing of FOR-TRAN source code. It supports FORTRAN 77 (F77) and FORTRAN 90 (F9X) language standards, with common VAX and Cray extensions.”.

Scott:1993:VBD

[Sco93] D. F. Scott. Visual Basic for DOS Developer’s Guide. Howard W. Sams, Indianapolis, IN 46268, USA, May 1, 1993. ISBN 0-672-30321-3. Includes disk.

Silver:1990:FIP

[SD90] N. Clayton Silver and William P.

Dunlap. A FORTRAN IV program for testing the significance of correlation matrices. Educational and Psychological Measurement, ??(3):597–??, Fall 1990. CODEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Stearns:1992:SPA

[SD92] Samuel D. Stearns and Ruth A.

David. Signal Processing Algorithms Using Fortran and C. Prentice-Hall Signal Processing Series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, November 1992. ISBN 0-13-816307-3, 0-13-812694-1. xiii + 331 pp. LCCN TK5102.5.S6993 1993. US$62.00; US$56.00; US$45.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0138126941;   http:
//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0138163073. Includes

3.5in disk.

Stearns:1993:SPA

[SD93] Samuel D. Stearns and Ruth A. David. Signal processing algorithms using Fortran and C. Prentice-Hall signal processing series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-816307-3, 0-13-812694-1. xiii + 331 pp. LCCN TK5102.5.S6993 1993 Reserves. System requirements for computer disk: IBM PC or compatible; MS-DOS; VGA graphics. Cover title: Signal processing algorithms in Fortran and C.

Silber:1999:NLT

[SD99] G.-A. Silber and A. Darte. The Nestor library: a tool for implementing Fortran source to source transformations. Lecture Notes in Computer Science, ??(1593):653–??, ???? 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seymour:2001:ATF

[SD01] Keith Seymour and Jack Dongarra. Automatic translation of Fortran to JVM bytecode. In ACM [ACM01], pages 126–133. ISBN 1-58113-359-6. LCCN QA76.9.O35 A26 2001.

URL http://www.netlib.
org/utk/people/JackDongarra/
PAPERS/f2jreport.pdf;   http:
//www.philippsen.com/JGI2001/
camerareadyabstracts/51.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500126.ps.

Seymour:2003:ATF

[SD03] Keith Seymour and Jack Dongarra. Automatic translation of Fortran to JVM bytecode. Concurrency and Computation: Practice and Experience, 15(3–5):207–222, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sips:1998:ALE

[SDv98] Henk J. Sips, Will Denissen, and Kees van Reeuwijk. Analysis of local enumeration and storage schemes in HPF. Parallel Computing, 24(3–4):355– 382, May 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.elsevier.
com/cas/tree/store/parco/
sub/1998/24/3-4/1294.pdf.

Seeley:2004:HWF

[See04] Donn Seeley. How not to write Fortran in any language. ACM Queue: Tomorrow’s Computing Today, 2(9):58–65, December/January 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Senda:2003:IPI

[Sen03] J. Rafael Senda, editor. IS-SAC 2003: Proceedings of the 2003 International Symposium on Symbolic and Algebraic Computation, August 3–6, 2003, Drexel University, Philadelphia, PA, USA. ACM Press, New York, NY 10036, USA, 2003. ISBN 1-58113-641-2. LCCN QA76.95. ACM order number 505030.

Schlick:1992:TETa

[SF92] Tamar Schlick and Aaron Fogelson. TNPACK—A truncated Newton minimization package for large-scale problems: I. algorithm and usage. ACM Transactions on Mathematical Software, 18(1):46–70, March 1992. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1992-18-1/p46-schlick/.

Stein:1993:DAO

[SF93] J. Stein and G. C. Fox. Dependence analysis for outer loop parallelization of existing Fortran-77 programs. Concurrency, practice and experience, 5(8):659–674, December 1993. CODEN CPEXEI. ISSN 1040-3108.

Sakagami:2002:PEJ

[SF02] Hitoshi Sakagami and Shingo Furubayashi. Performance evaluation for Japanese HPF compilers with special benchmark suite. Lecture Notes in Computer Science, 2327:491– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2327/23270491.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270491.
pdf.

Stamatiadis:2010:ATA

[SF10] S. Stamatiadis and S. C. Farantos. auto deriv: Tool for automatic differentiation of a Fortran code. Computer Physics Communications, 181 (10):1818–1819, October 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Silver:1992:FPT

[SFB92] N. Clayton Silver, Michael S.

Finger, and Robert T. Burkey. A FORTRAN 77 program for testing trends among independent correlations. Educational and Psychological Measurement, 52(1):109–112, Spring 1992. CODEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Sang:2002:DCB

[SFKL02] J. Sang, G. Follen, C. Kim, and I. Lopez. Development of CORBA-based engineering applications from legacy Fortran programs. Information and Software Technology, 44 (3):175–184, March 15, 2002. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

DeSilva:1993:CPPa

[SG93a] Clarence De Silva and Lalith Gamage. C-language programming part 11–some useful features of C. M and C, Measurements and Control, 27(3): 86–??, June 1993. ISSN 0148-0057.

DeSilva:1993:CPPb

[SG93b] Clarence De Silva and Lalith Gamage. C-language programming part 12–programming language comparison. M and C, Measurements and Control, 27(4):118–??, September 1993. ISSN 0148-0057.

DeSilva:1993:CPPd

[SG93c] Clarence De Silva and Lalith Gamage. C-language programming part 12–programming language comparison. Medical Electronics, 24(4):50–??, September 1993. ISSN 0149-9734.

DeSilva:1993:CPPc

[SG93d] Clarence De Silva and Lalith Ganage. C-language programming part 11 — some useful features of C. Medical Electronics, 24(3):52–??, June 1993. ISSN 0149-9734.

Sreedhar:1995:LTA

[SG95] Vugranam C. Sreedhar and Guang R. Gao. A linear time algorithm for placing [phi]- nodes. In ACM [ACM95b], pages 62–73. ISBN 0-89791-692-1. LCCN QA 76.7 A11

1995. URL http://www.acm.
org:80/pubs/citations/proceedings/
plan/199448/p62-sreedhar/.

ACM order number: 549950.

Scott:1997:GOF

[SGMS97] T. C. Scott, I. P. Grant, M. B.

Monagan, and V. R. Saunders. Generation of optimized FORTRAN code for molecular integrals of Gaussian-type functions. Maple Technical Newsletter, 4(2):15–24, 1997. ISSN 1061-5733.

Sandlin:1991:PIE

[SH91] Doral R. Sandlin and Kipp E.

Howard. The power induced effects module a FORTRAN code which estimates lift increments due to power induced effects for V/STOL flight. [NASA contractor report]; NASA CR-188081, Cal Poly State University; National Aeronautics and Space Administration, San Luis Obispo, CA, USA, 1991. ???? pp.

Sun:1997:FCP

[SH97] Yang Sun and Kenji Hara. Fortran code of the Projected Shell Model: feasible shell model calculations for heavy nuclei. Computer Physics Communications, 104(1–3):245–258, August 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465597000647.

Sheng:2007:JAI

[SH07] Yanyan Sheng and Todd C. Headrick. JMASM 27: an algorithm for implementing Gibbs sampling for 2PNO IRT models (Fortran). Journal of Modern Applied Statistical Methods, 6(1):??, ???? 2007. CO-DEN ???? ISSN 1538-9472.

URL https://jmasm.com/
index.php/jmasm/article/
view/308.

Sheng:2009:JGS

[SH09] Yanyan Sheng and Todd C. Headrick. JMASM 28: Gibbs sampling for 2PNO multi-unidimensional item response theory models (Fortran). Journal of Modern Applied Statistical Methods, 8 (2):??, ???? 2009. CO-DEN ???? ISSN 1538-9472.

URL https://jmasm.com/
index.php/jmasm/article/
view/458.

Shah:1994:FSE

[Sha94] Lacky V. Shah. Fortran specifications for ensemble: research project. Master of science, plan ii, Dept. of Electrical; Engineering and Computer Sciences, University of California, Berkeley, Berkeley, CA, USA, 1994. 31 pp.

Sharp:1995:AAM

[Sha95] Bill Sharp. The Alpha Alternative to Mainframes. Datamation, 41(8):61–62, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Stewart:1991:ADF

[SHCP91] A. Stewart, C. Holt, M. Clint, and R. Perrott. Algorithm 128: a DAP FOR-TRAN subroutine for thinning digitised pictures. The Computer Journal, 34(4):377– 379, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_34/Issue_04/tiff/
377.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_04/tiff/
378.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_04/tiff/
379.tif.

Sherrill-Lubinski:1991:SVF

[She91] Sherrill-Lubinski Corporation.

SL-GMS VAX Fortran interface reference: object-oriented graphical modeling system, version 4.0. Technical report, Sherrill-Lubinski Corporation, ??, ??, USA, January 21, 1991. 113 pp.

Shepard:1992:PGB

[She92]  R. Shepard.  A proposal

for generic BLAS, LINPACK, and LAPACK: a step towards portability. Fortran Journal, 4(2):6–12, ???? 1992. ISSN 1060-0221.

Shiau:1993:OOP

[Shi93a]  Lie June Shiau.  Object-

oriented programming for numerical methods. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):5–8, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shirts:1993:AMM

[Shi93b] Randall B. Shirts. Algorithm 721: MTIEU1 and MTIEU2: Two subroutines to compute eigenvalues and solutions to Mathieu’s differential equation for noninteger and integer order. ACM Transactions on Mathematical Software, 19(3): 391–406, September 1993. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Shirer:1998:FSO

[Shi98] Donald L. Shirer. Fortran 90 subsets offer simplified environment for first-time users. Computers in physics, 12(2): 166–??, March 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

                                        aip.scitation.org/doi/10.
                                        1063/1.168632.

Shterenlikht:2019:QIF

[Sht19] A. Shterenlikht. On quality of implementation of Fortran 2008 complex intrinsic functions on branch cuts. ACM Transactions on Mathematical Software, 45(1):11:1– 11:9, March 2019. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL https://dl.acm.org/
citation.cfm?id=3301318.

Shindo:1995:HCA

[SIDH95] T. Shindo, H. Iwashita, T. Doi, and J. Hagiwara. HPF compiler for the AP1000. In ACM [ACM95a], pages 190– 194. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995.

Siegal:1994:PEI

[Sie94a] Howard Jay Siegal, editor.

Proceedings / Eighth International Parallel Processing Symposium, April 26–29, 1994, Cancun, Mexico. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5602-6. LCCN QA76.58.I58 1994. IEEE catalog no. 94CH34819.

Siegel:1994:PEI

[Sie94b] Howard Jay Siegel, editor. Proceedings / Eighth International Parallel Processing Symposium, April 26–29, 1994, Cancun, Mexico. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5602-6, 0-8186-5601-8. ISSN 1063-7133. LCCN QA 76.58 I56 1994.

SGI:1992:FLR

[Sil92a] Silicon Graphics Computer Systems. Fortran 77 language reference manual. Silicon Graphics, Mountain View, CA, USA, 1992. various pp.

SGI:1992:FPG

[Sil92b] Silicon Graphics Computer Systems. Fortran 77 programmer’s guide. Silicon Graphics, Mountain View, CA, USA, 1992. various pp.

Silbar:1993:INI

[Sil93] Richard R. Silbar. An interactive NeXTSTEP interface to a Fortran code for solving coupled differential equations. Computers in physics, 7(5):592–598, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Silver:2001:DFP

[Sil01] N. C. Silver. DIFALPHA: a FORTRAN 77 program for testing the difference between independent alpha coefficients with different test lengths. Applied psychological measurement, 25(1):68(–67), March 2001. CODEN ???? ISSN 0146-6216.

Seo:2002:HJE

[SIOS02] Yoshiki Seo, Hidetoshi Iwashita, Hiroshi Ohta, and Hitoshi Sakagami. HPF/JA: extensions of High Performance Fortran for accelerating realworld applications. Concurrency and Computation: Practice and Experience, 14(8–9): 555–573, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/95016127/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016127{\&}PLACEBO=IE.
pdf.

Schieber:1994:RRT

[SJ94] Colleen D. Schieber and Eric E.

Johnson. RATCHET: realtime address trace compression hardware for extended traces. ACM SIGMETRICS Performance Evaluation Review, 21 (3–4):22–32, April 1994. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Scott:1994:FSA

[SKM94] N. S. Scott, P. L. Kilpatrick, and D. Maley. The formal specification of abstract data types and their implementation in Fortran 90. Computer Physics Communications, 84(1–3):201– 225, November 1994. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Sehr:1991:FTF

[SKP91] David C. Sehr, Laxmikant V.

Kale, and David A. Padua.

Fortran-style transformations for functional programs (extended abstract). In Proceedings of the 1991 International Conference on Parallel Processing, volume II, Software, pages II–282–II–283. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, August 1991.

Slaton:1996:FC

[Sla96] Jonathan Slaton. F World is cool. Fortran Journal, 8(6): ??, November/December 1996. ISSN 1060-0221. URL http:/

                                        /www.fortran.com/fortran/
                                        FJ/9611/#cool.

Shen:1990:ESFa

[SLY90a] Zhiyu Shen, Zhiyuan Li, and Pen-Chung Yew. An empirical study of Fortran programs for parallelizing compilers. Report CSRD 983, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL, USA, 1990. 37 pp.

Shen:1990:ESFb

[SLY90b] Zhiyu Shen, Zhiyuan Li, and Pen-Chung Yew. Empirical study of Fortran programs for parallelizing compilers. IEEE Transactions on Parallel and Distributed Systems, 1(3):356– 364, July 1990. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Schonfelder:1990:DSF

[SM90] J. L. Schonfelder and J. S. Morgan. Dynamic strings in Fortran 90. Software—Practice and Experience, 20(12):1259– 1271, December 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Saltz:1992:LCR

[SM92] Joel Saltz and Piyush Mehrotra, editors. Languages, compilers, and run-time environments for distributed memory machines, volume 3 of Advances in parallel computing. Elsevier, Amsterdam, The Netherlands, 1992. ISBN 0-444-88712-1. viii + 314 pp. LCCN QA76.9.D5 L36 1992.

Sass:1994:EUT

[SM94] Ron Sass and Matt Mutka.

Enabling unimodular transformations. In IEEE [IEE94f], pages 753–762. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

URL http://sc94.ameslab.
gov/AP/contents.html. IEEE

catalog number 94CH34819.

Sherlock:1995:AFD

[SM95] Barry G. Sherlock and Donald M. Monro. Algorithm 749: Fast discrete cosine transform. ACM Transactions on Mathematical Software, 21(4):372– 378, December 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sakagami:2002:CCP

[SM02a] H. Sakagami and T. Mizuno. Compatibility comparison and performance evaluation for Japanese HPF compilers using scientific applications. Concurrency and Computation: Practice and Experience, 14(8–9): 679–689, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/95016123/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016123{\&}PLACEBO=IE.
pdf.

Shires:2002:EHM

[SM02b] D. Shires and R. Mohan. An evaluation of HPF and MPI approaches and performance in unstructured finite element simulations. Journal of Mathematical Modelling and Algorithms, 1(3):153–167, 2002. CODEN ???? ISSN 1570-1166.

Shires:2003:OPF

[SM03] Dale Shires and Ram Mohan. Optimization and performance of a Fortran 90 MPI-based unstructured code on large-scale parallel systems. The Journal of supercomputing, 25(2):131–141, June 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//ipsapp009.kluweronline.
com/content/getfile/5189/
44/4/abstract.htm;   http:
//ipsapp009.kluweronline.
com/content/getfile/5189/
44/4/fulltext.pdf.

Sharma:2017:MCF

[SM17] Anuj Sharma and Irene Moulitsas. MPI to Coarray Fortran: Experiences with a CFD solver for unstructured meshes. Scientific Programming, 2017(1): 3409647:1–3409647:??, 2017. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https:/

/onlinelibrary.wiley.com/
doi/epdf/10.1155/2017/3409647.

Seabaugh:1990:EIF

[SMB90] Alan Carter Seabaugh, John J.

Mathias, and Michael I. Bell. EPROP, an interactive FOR-TRAN program for computing selected electronic properties of gallium arsenide and silicon. NIST special publication. Semiconductor measurement technology 400-85, U.S. Dept. of Commerce, National Institute of Standards and Technology, Gaithersburg, MD, USA, May 1990. v + 117 pp. For sale by the Supt. of Docs., U.S. G.P.O.

Sabot:1991:CFO

[SMG91] Gary Sabot, Janet Marantz, and David Gingold. CM Fortran optimization notes: slicewise model. Technical report series TMC-184, Thinking Machines Corp., Cambridge, MA, USA, March 1991. vii + 35 pp.

Strout:1991:ECS

[SMH91] Robert E. Strout, II, James R. McGraw, and Alan C. Hindmarsh. An examination of the conversion of software to multiprocessors. Journal of Parallel and Distributed Computing, 13(1):1–16, September 1, 1991. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Smith:1991:AFP

[Smi91] David M. Smith. Algorithm 693: A FORTRAN package for floating-point multipleprecision arithmetic. ACM Transactions on Mathematical Software, 17(2):273–283, June 1991. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0098-3500/
                                        108585.html.

Smith:1992:OFT

[Smi92] Glen Smith. Object-oriented Fortran tutorial. Technical report, Engineering Research Center, Mississippi State University, Mississippi State, MS, USA, 1992.

Smith:1993:OOF

[Smi93a] Glen D. Smith. Object oriented FORTRAN: tutorial. MSSU-EIRS-ERC 93-17, Engineering Research Center for Computational Field Simulation, Mississippi State, MS, USA, 1993. 18 + [33] pp.

Smith:1993:RFP

[Smi93b] R. I. Smith. Ricicle: A Fortran program to refine unit cell parameters of incommensurate structures. Powder Diffraction, 8(3):168–172, September 1993. CODEN PODIE2. ISSN 0885-7156.

Smith:1994:PFF

[Smi94]  I. M. Smith.  Program-

ming in FORTRAN 90: a First Course for Engineers and Scientists. John Wiley and Sons, New York, London, Sydney, September 1994. ISBN 0-471-94185-9. 209 pp. US$45.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0471941859.

Smith:1995:WHA

[Smi95a] Brian T. Smith. Workshop:

HPF: a user’s perspective. In Karin [Kar95], page ?? ISBN 0-89791-816-9. LCCN QA 76.88 S8573 1995a. URL http:

//www.supercomp.org/sc95/
proceedings/MISC/WORKSHOP.

HTM#HPF. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Smith:1995:PFF

[Smi95b] I. M. (Ian Moffat) Smith. Programming in Fortran 90: a First Course for Engineers and Scientists. John Wiley and Sons, New York, London, Sydney, May 1995. ISBN 0-471-94185-9 (paperback). viii + 209 pp. LCCN QA76.73.F25 S55 1995. US$45.00.

Smith:1998:AMP

[Smi98] David M. Smith. Algorithm 786: Multiple-precision complex arithmetic and functions. ACM Transactions on Mathematical Software, 24(4):359– 367, December 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [Bai95, Bre78, Bre79, BHY80].

Smith:2000:SPF

[Smi00] Brian Smith. Sneak peek at Fortran 2002. Research & Development, 42(3):39, March 2000. CODEN REDEEA. ISSN 0746-9179.

Smith:2001:AFS

[Smi01] David M. Smith. Algorithm 814: Fortran 90 software for floating-point multiple precision arithmetic, gamma and related functions. ACM Transactions on Mathematical Software, 27(4):377–387, December 2001. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Smith:2011:AMP

[Smi11] David M. Smith. Algorithm 911: Multiple-precision exponential integral and related functions. ACM Transactions on Mathematical Software, 37 (4):46:1–46:16, February 2011. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Smolarski:1994:EF

[Smo94] Dennis Chester Smolarski. The ESSENTIALS of FORTRAN. Research and Education Association, New York, NY, USA, revised printing 1994 edition, 1994. ISBN ???? viii + 120 pp. LCCN ????

Su:2006:APP

[SMSW06] Hai-Jun Su, J. Michael Mc-Carthy, Masha Sosonkina, and Layne T. Watson. Algorithm 857: POLSYS GLP—a parallel general linear product homotopy code for solving polynomial systems of equations. ACM Transactions on Mathematical Software, 32(4):561– 579, December 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sakagami:2002:TTD

[SMSY02] Hitoshi Sakagami, Hitoshi Murai, Yoshiki Seo, and Mitsuo Yokokawa. 14.9 TFLOPS three-dimensional fluid simulation for fusion science with HPF on the Earth Simulator. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.
                                        org/paperpdfs/pap.pap147.
                                        pdf.

Sumana:1994:PRA

[SN94] S. Sumana and U. Nagaraj Shenoy. Porting realistic applications to HPF — a case study. In Kumar [Kum94], pages 370–375. ISBN 0-07-462332-X. LCCN QA 76.58 I587 1994.

Snir:1992:PI

[Sni92] Marc Snir. Proposal for IO. Posted to HPFF I/O Forum, August 31, 1992. Second Draft.

Stout:1992:CPE

[SNJ^{+}92] William Stout, Ratna Nandakumar, Brian Junker, Hua-Hua Chang, and Duane Steidinger. Computer program exchange dimtest. Applied psychological measurement, 16(3): 236–??, September 1992. ISSN 0146-6216.

Satoh:2006:ATL

[SNK06] Makoto Satoh, Kiyoshi Negishi, and Atsushi Kobayashi. Analysis of two-level data mapping in an HPF compiler for distributed-memory machines. Parallel Computing, 32(4):280– 300, April 2006. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schuster:1993:RHP

[SNMC93] V. J. Schuster, S. Nakamoto, L. Meadows, and A. Choudhary. Retargetable High Performance Fortran compilers. In Anonymous [Ano93q], pages 111–123. ISBN ???? LCCN ????

Snyder:2007:SPF

[Sny07] W. Van Snyder. Scientific programming in Fortran. Scientific Programming, 15(1):3–8, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

SSVG:1993:FCV

[Sof93] Software Standards Validation Group. 1978 Fortran compiler validation system user’s guide. Technical Report NI-STIR 5287, U.S. Dept. of Commerce, National Institute of Standards and Technology, Gaithersburg, MD, USA, 1993. various pp. For sale by the National Technical Information Services.

Stichnoth:1994:GCA

[SOG94] James M. Stichnoth, David O’Hallaron, and Thomas R. Gross. Generating communication for array statements: Design, implementation, and evaluation. Journal of Parallel and Distributed Computing, 21 (1):150–159, April 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1994.1048/production;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1994.1048/production/pdf.

Somerville:1998:FPE

[Som98] Paul N. Somerville. A Fortran 90 program for evaluation of multivariate normal and multivariate t integrals over convex regions. Journal of Statistical Software, 3(4):1– 10, 1998. CODEN JSSOBK.

                                        ISSN ???? URL http://www.
                                        jstatsoft.org/v03/i04;
                                        http://www.jstatsoft.org/
                                        v03/i04/mvi3.for;   http:
                                        //www.jstatsoft.org/v03/
                                        i04/somerville.pdf;   http:
                                        //www.jstatsoft.org/v03/
                                        i04/somerville.ps;   http:
                                        //www.jstatsoft.org/v03/
                                        i04/updates;   http://www.
                                        jstatsoft.org/v03/i04/winmvi3.
                                        exe.

Somerville:2007:CCV

[Som07] Paul N. Somerville. Calculation of critical values for Somerville’s FDR procedures. Journal of Statistical Software, 21(6):1–14, October 2007. CO-DEN JSSOBK. ISSN 1548-

                                        7660.  URL http://www.
                                        jstatsoft.org/v21/i06.

Sawdey:1993:IFC

[SOP93] Aaron Sawdey, Matthew O’Keefe, and Terence Parr. Implementing a Fortran 77 to CM Fortran translator using the SORCERER source-to-source translator generator. AH-PCRC preprint 93-102, Army High Performance Computing Research Center, Minneapolis, MN, USA, 1993. ii + 15 pp.

Soul91a

[Sou91a] Edgar J. Soulié. User’s experience with Fortran precompilers for least squares optimization problems. In Andreas Griewank and George F. Corliss, editors, Automatic Differentiation of Algorithms: Theory, Implementation, and Application. SIAM, Philadelphia, PA, 1991.

Soulie:1991:UEF

[Sou91b] Edgar J. Soulié. User’s experience with Fortran precompilers for least squares optimization problems. In Andreas Griewank and George F. Corliss, editors, Automatic Differentiation of Algorithms: Theory, Implementation, and Application, page ?? SIAM Press, Philadelphia, PA, USA, 1991.

Step91a

[SP91a] Bruce R. Stephens and John D.

Pryce. DAPRE: a differentiation arithmetic system for FORTRAN. Technical Report ACM–91–3, Royal Military College of Science, Shrivenham, UK, 1991.

Stephens:1991:DAD

[SP91b] Bruce R. Stephens and John D.

Pryce. DAPRE: a differentiation arithmetic system for FORTRAN. Technical Report ACM–91–3, Royal Military College of Science, Shrivenham, UK, 1991.

Spencer:1993:RNR

[Spe93] Henry Spencer. Review: Numerical Recipes in FORTRAN. ;login: the USENIX Association newsletter, 18(3):36–37, May/June 1993. CODEN LOGNEM. ISSN 1044-6397.

Spearing:1994:PFP

[Spe94] Dane R. Spearing. ptchg: a FORTRAN program for pointcharge calculations of electric field gradients (EFGs). Computers and Geosciences, 20(4): 615–??, May 1994. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

SSC:1996:FR

[Spe96a] Specialized Systems Consultants, Inc.Staff. Fortran 77 Reference. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, December 1996. ISBN 0-916151-08-5. US$2.50.

Staff:1996:FR

[Spe96b] Specialized Systems Consultants, Inc.Staff. Fortran 77 Reference. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, December 1996. ISBN 0-916151-08-5. US$2.50. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=0916151085.

Sperry:19xx:FII

[Spexx] Sperry Rand Corporation.Univac Division. FORTRAN IV: instructor’s guide. Sperry Rand Corp., St. Paul, MN, USA, 19xx. ???? pp.

Stamatiadis:2000:ATA

[SPF00] S. Stamatiadis, R. Prosmiti, and S. C. Farantos. auto deriv: Tool for automatic differentiation of a Fortran code. Computer Physics Communications, 127(2):343–355, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Sharma:1994:RCS

[SPM^{+}94] S. D. Sharma, R. Ponnusamy, B. Moon, Yuan-Shin Hwang, R. Das, and J. Saltz. Run-time and compile-time support for adaptive irregular problems. In IEEE [IEE94f], pages 97–106. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Spoerl:1994:SHC

[Spo94] J. R. Spoerl. Softech’s HPF compiler. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):75, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Sahulka:1991:FCI

[SPS^{+}91] R. J. Sahulka, E. C. Plachy, L. J. Scarborough, R. G. Scarborough, and S. W. White. FORTRAN for clusters of IBM ES/ 3090 multiprocessors. IBM Systems Journal, 30(3):296– 311, 1991. CODEN IBMSA7. ISSN 0018-8670.

Sivaraman:1995:PSP

[SR95] H. Sivaraman and C. S. Raghavendra. Parallelizing sequential programs to a cluster of workstations. In Agrawal [Agr95], pages 38–41. ISBN 0-8493-2618-4. LCCN QA76.58.I34 1995.

Souravlas:2004:PTD

[SR04] Stavros Souravlas and Manos Roumeliotis. A pipeline technique for dynamic data transfer on a multiprocessor Grid. International Journal of Parallel Programming, 32(5):361– 388, October 2004. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (elec-

                                        tronic). URL http://www.
                                        springerlink.com/openurl.
                                        asp?genre=article&issn=0885-
                                        7458&volume=32&issue=5&spage=
                                        361.

Sreenath:1992:HCE

[Sre92] N. Sreenath. A hybrid computation environment for multibody simulation. Mathematics and Computers in Simulation, 34(2):121–140, August 1992. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Sagiv:1996:PID

[SRH96] Mooly Sagiv, Thomas Reps, and Susan Horwitz. Precise interprocedural dataflow analysis with applications to constant propagation. Theoretical Computer Science, 167(1– 2):131–170, October 30, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/tcs/cas_sub/browse/
browse.cgi?year=1996&volume=
167&issue=1-2&aid=2274.

Sondergard:1990:FOF

[SRM90] M. A. Sondergard, J. E. Robinson, and D. F. Merriam. FILT-PC, a one-dimensional Fourier transform program in FOR-TRAN for the PC. Computers and Geology, 6(??):251– ??, 1990. CODEN COMGDX. ISSN 0271-0420.

Sathe:1990:FPC

[SS90] P. V. Sathe and Shubha Sathyendranath. FORTRAN programs for computation of optical properties of the sea from radiation data collected by in situ spectrometers. Computers and Geosciences, 16 (8):1085–??, 1990. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Schuette:1993:ILE

[SS93] Michael A. Schuette and John P. Shen. Instructionlevel experimental evaluation of the Multiflow TRACE 14/ 300 VLIW computer. The Journal of supercomputing, 7 (1–2):249–271, May 1993. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://www.
                                        springerlink.com/openurl.
                                        asp?genre=article&issn=0920-
                                        8542&volume=7&issue=1&spage=
                                        249.

Schick:1994:FEC

[SS94] William Schick and Gordon Silverman. FORTRAN 90 and Engineering Computation. John Wiley and Sons, New York, London, Sydney, November 1994. ISBN 0-471-58512-2. 425 pp. US$65.95. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=0471585122.

Schick:1995:FEC

[SS95] William Schick and Gordon Silverman. Fortran 90 and Engineering Computation. John Wiley and Sons, New York, London, Sydney, January 1995. ISBN 0-471-58512-2 (paper). xvi + 425 pp. LCCN QA76.73.F25 S333 1995. US$65.95.

Szymanski:1996:LCR

[SS96] Boleslaw K. Szymanski and Balaram Sinharoy, editors. Languages, Compilers and Run-Time Systems for Scalable Computers, 22–24 May 1995, Troy, NY, USA. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Stenger:1999:CMS

[SS99] Frank Stenger and Ross Schmidtlein. Conformal maps via Sinc methods. In Papamichael et al. [PRS99], pages 505–549. ISBN 981-02-3626-3. LCCN QA297 .I473 1997.

Subhlok:2000:APM

[SS00] J. Subhlok and P. Steenkiste.

Airshed pollution modeling in an HPF style environment. Journal of Parallel and Distributed Computing, 60(6):690– 715, June 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sahu:2009:FIH

[SS09] Sridhar Sahu and Alok Shukla.

Fortran 90 implementation of the Hartree–Fock approach within the CNDO/2 and INDO models. Computer Physics Communications, 180(5):724–734, May 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465508003834.

Sony:2010:GPF

[SS10] Priya Sony and Alok Shukla.

A general purpose Fortran 90 electronic structure program for conjugated systems using Pariser–Parr–Pople model. Computer Physics Communications, 181(4):821–830, April 2010. CODEN CPHCBZ.

ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S001046550900407X.

Shih:2000:EAG

[SSC00] K-P. Shih, J-P. Sheu, and C-Y. Chang. Efficient address generation for affine subscripts in data-parallel programs. The Journal of supercomputing, 17(2):205–227, January 2000. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Sussman:1993:BIL

[SSG93] G. J. Sussman, G. L. Steele, Jr., and R. P. Gabriel. A brief introduction to Lisp. ACM SIGPLAN Notices, 28 (3):361–362, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/proceedings/
                                        plan/154766/p361-lee/.

Suzuoka:1994:PDB

[SSG94] T. Suzuoka, J. Subhlok, and T. Gross. Performance debugging based on scalability analysis. In IEEE [IEE94a], pages 406–413. ISBN 0-8186-6965-

9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Suzuoka:1997:PDT

[SSG97] Takashi Suzuoka, Jaspal Subhlok, and Thomas Gross. A performance debugging tool for high performance Fortran programs. Concurrency, practice and experience, 9(10):927–945, October 1997. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
13809;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13809&PLACEBO=IE.pdf.

Schneider:2010:NFP

[SSG^{+}10] Barry I. Schneider, Javier Segura, Amparo Gil, Xiaoxu Guan, and Klaus Bartschat. A new Fortran 90 program to compute regular and irregular associated Legendre functions. Computer Physics Communications, 181(12):2091– 2097, December 2010. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465510003450.

Schneider:2018:NFP

[SSG^{+}18] Barry I. Schneider, Javier Segura, Amparo Gil, Xiaoxu Guan, and Klaus Bartschat. A new Fortran 90 program to compute regular and irregular associated Legendre functions (new version announcement). Computer Physics Communications, 225(??):192–193, April 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465517304186.

Sala:2008:PHP

[SSH08] Marzio Sala, W. F. Spotz, and M. A. Heroux. PyTrilinos: High-performance distributedmemory solvers for Python. ACM Transactions on Mathematical Software, 34(2):7:1– 7:33, March 2008. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sherriff:1991:CFP

[SSLG91] Barbara L. Sherriff, Vijay Singh, Jianjie Liang, and H. Douglas Grundy. CHEMSHIFT: a FORTRAN program to calculate ^{29}Si MAS NMR chemical shift of silicate minerals. Computers and Geosciences, 17(7):967–972, 1991. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Subhlok:1993:ETD

[SSOG93] Jaspal Subhlok, James M. Stichnoth, David R. O’Hallaron, and Thomas Gross. Exploiting task and data parallelism on a multicomputer. ACM SIGPLAN Notices, 28(7):13– 22, July 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stringfellow:1999:GSS

[SSS99] N. D. Stringfellow, R. N. L. Smith, and V. V. S. S. Sastry. A generic structure for scalar and vector boundary element codes using Fortran 90. Advances in engineering software, 30(5):313–??, ???? 1999. CO-DEN AESODT. ISSN 0965-9978 (print), 0141-1195 (electronic).

Scarborough:1991:CIE

[SSW91] L. J. Scarborough, R. G. Scarborough, and S. W. White. Clustering IBM Enterprise System/3090 computers for parallel execution of FORTRAN programs. IBM Journal of Research and Development, 35(5/6):667–679, September/ November 1991. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Snelting:1990:PTS

[ST90] G. Snelting and C. Thies.

Programming tools for the Suprenum supercomputer. In Toulouse ’90. Third International Workshop. Software Engineering and its Applications Proceedings, pages 951– 963 (vol. 2). EC2, Nanterre, France, December 1990. 2 vol.

Skillicorn:1995:PLP

[ST95] David B. Skillicorn and Domenico Talia. Programming languages for parallel processing. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-6502-5 (case). x + 399 pp. LCCN QA76.642 .P76 1995.

Stanford:1994:PQC

[Sta94] N. Stanford. Portable QCD codes for massively parallel processors. In Anonymous [Ano94i], pages 817–819. CO-DEN NPBSE7. ISBN ???? ISSN 0920-5632 (print), 1873-3832 (electronic). LCCN ????

Steppeler:1990:FFE

[Ste90] J. Steppeler. FE2DY: a finite element FORTRAN program for the solution of the shallow-water equations with energy conservation. Computers and Geosciences, 16 (5):645–??, 1990. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Sternberg:1991:PFP

[Ste91] M. J. Sternberg. PROMOT: a FORTRAN program to scan protein sequences against a library of known motifs. Computer applications in the biosciences: CABIOS, 7(2):257– 260, April 1991. CODEN COABER. ISSN 0266-7061.

Steele:1993:HPF

[Ste93] Guy L. Steele, Jr. High Performance Fortran: Status report. ACM SIGPLAN Notices, 28(1):1–4, January 1993. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stetter:1995:UAA

[Ste95a] Hans Stetter. The use of algebraic algorithms in scientific computing. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ????

LCCN ???? URL http:

//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/
proceedings/stetter/stetter1.
html.

Stewart:1995:RAD

[Ste95b] A. Stewart. Reasoning about data-parallel array assignment. Journal of Parallel and Distributed Computing, 27(1): 79–85, May 1995. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1995.1073/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1995.1073/production/pdf.

Stol:1993:FVV

[Sto93] W. Stol. The FORTRAN version of the van Keulen-Seligman CSMP-Spring wheat model. Simulation reports cabo-tt; no. 30, Centre for Agrobiological Research, Wageningen Agricultural University, Wageningen, The Netherlands, May 1993. 112 + I20 pp.

Straka:2005:ATA

[Str05] Christian W. Straka. ADF95:

Tool for automatic differentiation of a FORTRAN code designed for large numbers of independent variables. Computer Physics Communications, 168(2):123–139, June 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465505000810.

Sabot:1991:CPE

[STVS91] G. Sabot, L. Tennies, A. Vasilevsky, and R. Shapiro. Compiler parallelization of an elliptic grid generator for 1990 Gordon Bell Prize. In IEEE [IEE91], pages 338–346. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Shabaev:2015:QFP

[STY15] V. M. Shabaev, I. I. Tupitsyn, and V. A. Yerokhin. QEDMOD: Fortran program for calculating the model Lamb-shift operator. Computer Physics Communications, pages 175–181, ???? 2015. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465514004081.

Shabaev:2018:QFP

[STY18] V. M. Shabaev, I. I. Tupitsyn, and V. A. Yerokhin. QEDMOD: Fortran program for calculating the model Lamb-shift operator. Computer Physics Communications, 223(??):69, February 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465517303478.

Sugihara:1995:CAN

[Sug95] Kokichi Sugihara. Combinatorial abstraction — a new paradigm for robust geometric computation. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/
proceedings/sugihara/sugihara1.
html.

Sullivan:1991:VPI

[Sul91] Steve Sullivan. Vector and parallel implementations of the wavelet transform. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1991. viii + 64 pp.

Sun:1992:SFU

[Sun92a] SPARCompiler FORTRAN 2.0.1: User’s Guide. Sun Microsystems, Inc., Mountain View, CA, USA, October 1992. Part No. 800-6552-11, Revision A.

Sun:1992:SFN

[Sun92b] SPARCompiler FORTRAN:

Numerical Computation Guide. Sun Microsystems, Inc., Mountain View, CA, USA, October 1992. ???? pp. Part No. 800-7097-11, Revision A.

SunSoft:1993:SAD

[Sun93] SunSoft. Solaris application developer’s guide. SunSoft, Mountain Views, Calif., 1993. ISBN 0-13-205097-8. xvi + 102 + 6 pp. LCCN QA76.76.S95 S65 1993.

SunSoft:1994:FRM

[Sun94] SunSoft. FORTRAN 3.0.1 reference manual. SunSoft, Mountain View, CA, USA, 1994. xxvi + 478 pp.

Sun:2005:FIA

[Sun05] Sun Microsystems, Inc. Fortran 95 Interval Arithmetic Programming Reference Manual: Sun Studio 10. Sun Microsystems, Inc., Santa Clara, CA 95054, USA, January 2005. xiv + 141 pp. URL

                                        http://docs-pdf.sun.com/
                                        819-0503/819-0503.pdf;
                                        http://docs.sun.com/app/
                                        docs/doc/819-0503?q=819-

0503-10. Part no. 819-0503-10. Schlichting:1990:NFLa

[Sv90] J. J. F. M. Schlichting and H. A. van der Vorst. NUMVEC FORTRAN library manual. Report. Centrum voor Wiskunde en Informatica NM-9018, Stichting Mathematisch Centrum, Amsterdam, The Netherlands, 1990. 11 pp.

Subhlok:1995:OMS

[SV95] Jaspal Subhlok and Gary Vondran. Optimal mapping of sequences of data parallel tasks. ACM SIGPLAN Notices, 30 (8):134–143, August 1995. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sips:1996:ALE

[SVD96] H. J. Sips, K. Van Reeuwijk, and W. Denissen. Analysis of local enumeration and storage schemes in HPF. In ACM [ACM96a], pages 10–17. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Slape:1991:AMS

[SW91] J. K. Slape and P. J. L. Wallis.

Algorithm 127: a modification of Sale’s algorithm to accommodate FORTRAN 77. The Computer Journal, 34(4):373– 376, August 1991. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_34/Issue_04/tiff/
373.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_04/tiff/
374.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_04/tiff/
375.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_34/Issue_04/tiff/
376.tif.

Seamons:1994:EAI

[SW94] Kent E. Seamons and Marianne Winslett. An efficient abstract interface for multidimensional array I/O. In IEEE [IEE94f], pages 650–659. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994.

                                        URL http://sc94.ameslab.
                                        gov/AP/contents.html. IEEE

catalog number 94CH34819.

Swarztrauber:1984:FAV

[Swa84] P. N. Swarztrauber. FFT algorithms for vector computers. Parallel Computing, 1(1):45– 63, August 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sabot:1993:PEF

[SWBO93] Gary Sabot, Skef Wholey, Jonas Berlin, and Paul Oppenheimer. Parallel execution of a Fortran 77 weather prediction model. In IEEE [IEE93d], pages 538–545. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Sosonkina:2015:RAV

[SWH15] Masha Sosonkina, Layne T. Watson, and Jian He. Remark on algorithm 897: VT-DIRECT95: Serial and parallel codes for the global optimization algorithm DIRECT. ACM Transactions on Mathematical Software, 41(3):22:1– 22:2, June 2015. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [HWS09].

Silver:1995:MFP

[SWM95] N. Clayton Silver, Diane L.

Wadiak, and Catherine J. Massey. A Microsoft Fortran 77 program for testing the differences among independent first-order partial correlations. Educational and Psychological Measurement, 55(2):245– ??, 1995. CODEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Srinivasan:1992:IFP

[SWO92] A. R. Srinivasan, John D.

Westbrook, and Wilma K. Olson. An interactive FOR-TRAN program for threedimensional molecular visualization. Computers and Chemistry, 16(3):265–266, 1992. CO-DEN COCHDK. ISSN 0097-8485.

Sawdayi:1990:MFD

[SWW90] R. Sawdayi, G. Wagenbreth, and J. Williamson. MIMDizer: Functional and data decomposition; creating parallel programs from scratch, transforming existing Fortran programs to parallel. In ICASE Workshop. ????, ????, May 1990.

Schneider:1990:FPP

[SZ90] Klaus Schneider and Ronald F.

Zernicke. A Fortran package for the planar analysis of limb intersegmental dynamics from spatial coordinate-time data.

Advances in engineering software, 12(3):123–??, July 1990. CODEN AESODT. ISSN 0141-1195, 0965-9978.

Szelenyi:1991:VPE

[SZ91] F. Szelenyi and V. Zecca. Visualizing parallel execution of FORTRAN programs. IBM Journal of Research and Development, 35(1/2):270–282, January/March 1991. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Schulte:1997:AIS

[SZAB97] M. J. Schulte, V. Zelov, A. Akkas, and J. C. Burley. Adding interval support to the GNU Fortran compiler. Technical report, Lehigh University, Bethlehem, PA, USA, September 1997.

                                        URL http://home.ku.edu.
                                        tr/~ahakkas/publications/
                                        Add_Intv_Sup_Comp.pdf.

Schulte:1998:SAP

[SZAB98] M. J. Schulte, V. Zelov, A. Akkas, and J. C. Burley. Summary of accomplishments and plans for future research on the interval enhanced GNU Fortran compiler. Technical report, Lehigh University, Bethlehem, PA, USA, June 1998.

                                        URL http://home.ku.edu.
                                        tr/~ahakkas/publications/
                                        Summary.pdf.

Schulte:1999:IEG

[SZAB99] Michael J. Schulte, Vitaly Zelov, Ahmet Akkas, and James Craig Burley. The interval-enhanced GNU Fortran compiler. In Csendes [Cse99], pages 311–322. ISBN 0-7923-6057-5. LCCN QA76.9.E94

D48 1999. URL http://mesa.
ece.wisc.edu/publications/
cp_1999-12.pdf.

Szelenyi:1990:APF

[Sze90] F. Szelenyi. Analyzing parallel FORTRAN programs with an execution profiler. Technical Report ICE-VS06, IBM European Center for Scientific and Engineering Computing, Rome, Italy, May 1990.

Schulz-Ziemer:1995:HIP

[SZG95] G. Schulz-Ziemer and A. Geiger.

HPF on Intel Paragon and CRAFT on CRAY T3D: basic performance measurements and experiments with a block-sparse CG-algorithm. In Hertzberger and Serazzi [HS95], pages 618–625. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Sarma:1998:UHP

[SZM98] G. Sarma, T. Zacharia, and D. Miles. Using High Performance Fortran for parallel programming. Computers and Mathematics with Applications, 35(12):41–57, June 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0898122198000959.

Szymanski:2007:FPL

[Szy07] Boleslaw K. Szymanski. Fortran programming language and scientific programming: 50 years of mutual growth. Scientific Programming, 15(1):1–2, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Takeda:2001:AME

[TAH^{+}01] K. Takeda, N. K. Allsopp, J. C. Hardwick, P. C. Macey, D. A. Nicole, S. J. Cox, and D. J. Lancaster. An assessment of MPI environments for Windows NT. The Journal of supercomputing, 19(3): 315–323, July 2001. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://www.
                                        wkap.nl/oasis.htm/338207.

Tallin:1991:PTD

[Tal91] Andrew G. Tallin. Programming techniques: Dynamic memory allocation in FOR-TRAN. Journal of Computing in Civil Engineering, 5(1):122– ??, January 1991. CODEN JC-CEE5. ISSN 0887-3801.

Talay:1994:PSP

[Tal94] D. Talay. Presto: a software package for the simulation of diffusion processes. Statistics and Computing, 4(4):247– 252, December 1994. CO-DEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

Tamasanis:1995:MMW

[Tam95] Doug Tamasanis. Mathematica meets Warp. BYTE Magazine, 20(5), May 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Taqi:2016:VFP

[Taq16] Ali H. Taqi. A visual Fortran 90 program for the twoparticle or two-hole excitations of nuclei: the PPRPA program. SoftwareX, 5(??):44– 50, ???? 2016. CODEN ???? ISSN 2352-7110. URL

http://www.sciencedirect.
com/science/article/pii/
S2352711016300048.

Taylor:1986:VIP

[Tay86] D. Taylor. Visionaries: an interview with Professor Michael O’Leary. ACM SIGPLAN Notices, 21(6):41–54, June 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taylor:1997:STS

[Tay97]  Howard L. Taylor.  Some

thoughts on sub-languages of Fortran 90/95. Fortran Journal, 9(2):??, Fall 1997. ISSN 1060-0221.

Taylor:1999:BRC

[Tay99] Peter R. Taylor. Book review: Y. C. Pao, Engineering analysis: Interactive methods and programs with FORTRAN, QuickBasic, MATLAB, and Mathematica (1999) CRC Press,Bristol 0-8493-2016-X. Computer Physics Communications, 120(2–3):271–272, August 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465599002374.

Thakur:1994:COD

[TBC94a] R. Thakur, R. Bordawekar, and A. Choudhary. Compilation of out-of-core data, parallel programs for distributed memory machines. ACM SIGARCH Computer Architecture News, 22(4):23– 28, September 1994. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Thakur:1994:CRS

[TBC94b] R. Thakur, R. Bordawekar, and A. Choudhury (or Choudhary??). Compiler and runtime support for out-of-core HPF programs. In ACM [ACM94a], pages 382–391. ISBN 0-89791-665-4. LCCN ????

Tian:2002:IOC

[TBG^{+}02] Xinmin Tian, Aart Bik, Milind Girkar, Paul Grey, Hideki Saito, and Ernesto Su. Intel(R) OpenMP C++/Fortran compiler for hyper-threading technology: Implementation and performance. Intel Technology Journal, 6(1):36– 46, February 2002. ISSN 1535-766X. URL http://

developer.intel.com/technology/
itj/2002/volume06issue01/
vol6iss1_hyper_threading_
technology.pdf.

Thakur:1994:RAR

[TCF94] R. Thakur, A. Choudhary, and G. Fox. Runtime array redistribution in HPF programs. In IEEE [IEE94d], pages 309–

316. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994.

Templon:1997:BRG

[TDMC97] J. A. Templon, W. M. Dennis, Susan R. McKay, and Wolfgang Christian. Book review: Gisela Engeln-Müllges and Frank Uhlig, Numerical Algorithms with Fortran. Numerical Algorithms with C. Computers in physics, 11 (6):594–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.4822607.

Teague:1994:PPR

[Tea94]  Neal Teague.  Program

projects realistic solution-gas-drive GOR. The Oil and Gas Journal, 92(8):51–??, February 1994. CODEN OIGJAV. ISSN 0030-1388.

Teeter:1990:TFC

[Tee90] Eric Teeter. Translating Fortran to C. C Users Journal, 8(10):91–??, October 1990. ISSN 0898-9788.

Templeman:1996:AFP

[Tem96] Templeman. Advanced Fortran Programming for Windows. John Wiley and Sons, New York, London, Sydney, December 1996. ISBN 0-471-95685-6. US$39.95. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=0471956856.

Tentner:1993:HPC

[Ten93] Adrien Michel Tentner, editor. High Performance Computing Symposium 1993. Grand Challenges in Computer Simulation. Proceedings of the 1993 Simulation Multiconference on the High Performance Computing Symposium. SCS, San Diego, CA, USA, 1993. ISBN 1-56555-052-8. LCCN ????

Teodor:2001:WSD

[Teo01] Dan Teodor. Web servers and dynamic content. Linux Journal, 82:172, 174, 176, 178– 179, February 2001. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http://

                                        noframes.linuxjournal.com/
                                        lj-issues/issue82/4386.html.

Thacker:1993:NRF

[Tha93] B. H. Thacker. Numerical Recipes in Fortran; Second edition. William H Press et al. Applied mechanics reviews, 46(9): B121, September 1993. CO-DEN AMREAD. ISSN 0003-6900.

TMC:1991:GSC

[Thi91] Thinking Machines Corporation. Getting started in CM Fortran. Thinking Machines Corp., Cambridge, MA, USA, 1991. v + 46 pp.

Thornburg:1986:KIC

[Tho86] J. Thornburg. Keyword input for C. ACM SIGPLAN Notices, 21(6):55–58, June 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thompson:1990:MFP

[Tho90] Bruce Thompson. MULTI-NOR: A FORTRAN program that assists in evaluating multivariate normality. Educational and Psychological Measurement, ??(4):845–??, Winter 1990. CODEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Thole:1993:HPF

[Tho93] C. A. Thole. High Performance Fortran. In Anonymous [Ano93n], pages 885–892 (or 885–891??). ISBN ???? ISSN 0254-6213. LCCN ????

Thompson:1997:ACMb

[Tho97a] William J. Thompson. Atlas for Computing Mathematical Functions: An Illustrated Guidebook for Practitioners With Programs in Fortran 90 and Mathematica. Wiley-Interscience, New York, NY, USA, June 1997.

ISBN 0-471-18171-4. xiv + 888 pp. LCCN QA331.T386 1997. US$99.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0471181714.

Thompson:1997:TFS

[Tho97b] William J. Thompson. Three Fortran systems for desktop computers. Computers in physics, 11(4):370–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.4822573.

Thompson:2013:AIG

[Tho13] Ian Thompson. Algorithm 926: Incomplete Gamma functions with negative arguments. ACM Transactions on Mathematical Software, 39(2):14:1– 14:9, February 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Tipper:1991:FPC

[Tip91] John C. Tipper. FORTRAN programs to construct the planar Voronoi diagram. Computers and Geosciences, 17 (5):597–??, 1991. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Tanaka:1990:CTF

[TIUG90] Yoshikazu Tanaka, Kyouko Iwasawa, Yukio Umetani, and Shizuo Gotou. Compiling techniques for first-order linear recurrences on a vector computer. The Journal of supercomputing, 4(1): 63–82, March 1990. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=4&issue=1&spage=
63.

Tu:1990:CCP

[TJ90] Jho-Ju Tu and John R. Johnson. Can computer programming improve problem-solving ability? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 22(2):30–33, June 1990. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tirumalai:1990:PLE

[TLS90] P. Tirumalai, M. Lee, and M. Schlansker. Parallelization of loops with exits on pipelined architectures. In IEEE [IEE90a], pages 200– 212. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Tirumalai:1991:PWL

[TLS91] Parthasarathy P. Tirumalai, Meng Lee, and Michael S. Schlansker. Parallelization of WHILE loops on pipelined architectures. The Journal of supercomputing, 5(2–3):119– 136, October 1991. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=5&issue=2&spage=
119.

Tinetti:2013:RFL

[TMD13] Fernando G. Tinetti, Mariano Méndez, and Armando De Giusti. Restructuring Fortran legacy applications for parallel computing in multiprocessors. The Journal of supercomputing, 64(2):638–659, May 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11227-012-0863-x.

Tracy:2018:CMC

[TOC18] Fred Thomas Tracy, Thomas C. Oppe, and Maureen K. Corcoran. A comparison of MPI and co-array FORTRAN for large finite element variably saturated flow simulations. Scalable Computing: Practice and Experience, 19(4):423–432, ???? 2018. CODEN ???? ISSN 1895-1767. URL https:

//www.scpe.org/index.php/
scpe/article/view/1468.

Tomayko:1999:DAF

[Tom99] James E. Tomayko. Departments: Anecdotes: Fortran. IEEE Annals of the History of Computing, 21(2):65–??, ???? 1999. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Thomas:2004:LLF

[TOML04] James W. Thomas, Jon P.

Okada, Peter Markstein, and Ren-Cang Li. The Libm library and floating-point arithmetic in HP-UX for Itaniumbased systems: Updated for HP-UX 11i v2. Technical report, Hewlett–Packard Corporation, Palo Alto, CA, USA, December 3, 2004. 26

pp. URL http://h21007.
www2.hp.com/dspp/ddl/ddl_
Download_File_TRX/1,1249,
942,00.pdf.

Torby:1991:FE

[Tor91]  Bruce J. Torby.  Fortran

77 for Engineers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, January 1991. ISBN 0-13-326745-8. x + 462 pp. US$60.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0133267458.

Torres:2010:ADT

[Tor10] Germán A. Torres. Algorithm 900: a discrete time Kalman filter package for large scale problems. ACM Transactions on Mathematical Software, 37(1): 11:1–11:16, January 2010. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Touzeau:1984:FCF

[Tou84] Roy F. Touzeau. A Fortran compiler for the FPS-164 scientific computer. ACM SIGPLAN Notices, 19(6):48– 57, June 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thirumalai:1996:ECA

[TR96] Ashwath Thirumalai and J. Ramanujam. Efficient computation of address sequences in data parallel programs using closed forms for basis vectors. Journal of Parallel and Distributed Computing, 38(2):188– 203, November 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

                                        tronic). URL http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jpdc.1996.0140/production;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0140/production/pdf.

Treharne:1991:RFS

[Tre91] James T. Treharne. RSSP: a Fortran simulation package for use in teaching Response Surface Methodology. Thesis (M.S.), Auburn University, Auburn, AL, USA, 1991. viii + 97 pp.

Tremblay:1995:PF

[Tre95] Jean-Paul Tremblay. Programming in Fortran 77. Mc-Graw-Hill, New York, NY, USA, February 1995. ISBN 0-07-065179-5. US$19.95.

URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0070651795.

Treggiari:1997:DFM

[Tre97] L. P. Treggiari. Development of the Fortran Module Wizard within DIGITAL Visual Fortran. Digital Technical Journal, 9(1):13–??, ???? 1997. CODEN DTJOEL. ISSN 0898-901X. URL http:/

/www.digital.com:80/info/
DTJP02/DTJP02AH.HTM;   http:
//www.digital.com:80/info/
DTJP02/DTJP02HM.HTM;   http:
//www.digital.com:80/info/
DTJP02/DTJP02P8.PS;   http:
//www.digital.com:80/info/
DTJP02/DTJP02PF.PDF;   http:
//www.digital.com:80/info/
DTJP02/DTJP02SC.TXT.

Trouve:1990:RAP

[Tro90] Jean-Bernard Trouvé. Restructuration automatisée de programmes Fortran en pseudocode schématique. Canadian theses = Thèses canadiennes. Bibliothèque nationale du Canada, Ottawa, Ontario, Canada, 1990. ISBN 0-315-52714-5. 3 microfiches (221 im.) pp.

Taylor:1991:NMF

[TRS91] P. Taylor, P. Rosenberg, and M. G. Samsonova. A new method for finding long consensus patterns in nucleic acid sequences. Computer applications in the biosciences: CABIOS, 7:495–500, 1991.

CODEN COABER. ISSN 0266-7061.

Thirumalai:1996:CGO

[TRV96] A. Thirumalai, J. Ramanujam, and A. Venkatachar. Communication generation and optimization for HPF. In Szymanski and Sinharoy [SS96], pages 311–316. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Thompson:2006:FFD

[TS06a] S. Thompson and L. F. Shampine. A friendly Fortran DDE solver. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4):503–516, March/April 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tiwari:2006:BSB

[TS06b] Rakesh Prabhat Tiwari and Alok Shukla. A basis-set based Fortran program to solve the Gross–Pitaevskii equation for dilute Bose gases in harmonic and anharmonic traps. Computer Physics Communications, 174(12):966–982, June 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0010465506000993.

Trim:2025:AEF

[TS25] Sean J. Trim and Raymond J. Spiteri. Algorithm 1054: ellipFor, a Fortran software library for Legendre elliptic integrals and Jacobi elliptic functions with generalized input arguments. ACM Transactions on Mathematical Software, 51 (1):6:1–6:??, March 2025. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Tsai:2001:CFP

[Tsa01] H. Tsai. Comment on “A FORTRAN program for fitting Weibull distribution and generating samples” by A. Ghosh. Computers and Geosciences, 27(2):257–259, March 2001. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Tseng:1993:OFD

[Tse93] Chau-Wen Tseng. An optimizing Fortran D compiler for MIMD distributed-memory machines. Thesis (Ph.D.), Rice University, Dept. of Computer Science, Houston, TX, USA, 1993. xiii + 193 pp.

Tseng:1997:PPF

[Tse97] Chau-Wen Tseng. Performance of the prototype Fortran D compiler. Software—Practice and Experience, 27(7):763–796, July 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
7317;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=7317&PLACEBO=IE.pdf.

Toint:1992:LFS

[TT92] Ph. L. Toint and D. Tuyttens. LSNNO, A FORTRAN subroutine for solving largescale nonlinear network optimization problems. ACM Transactions on Mathematical Software, 18(3):308–328, September 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

                                        URL http://www.acm.org/
                                        pubs/toc/Abstracts/0098-
                                        3500/131771.html.

Tobochnik:1993:FCP

[TT93] Neal Tobochnik and Jan Tobochnik. Fortran compilers for personal computers. Computers in physics, 7(6):672– 677, November 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Turgut:1993:TFP

[Tur93] Altan Turgut. TOPOVEL: a FORTRAN program for tomographic inversion of porosity and velocity structure in the subsurface. Computers and Geosciences, 19(4):511– ??, April 1993. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Ting:1992:VWP

[TYJ92] Julian J.-L. Ting, J. M. Yuan, and T.-F. Jiang. Vectorizable wave propagation FOR-TRAN code for calculations of multiphoton dissociation.

Computer Physics Communications, 70(2):417–425, June 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
001046559290204C.

Thacker:2010:AMS

[TZW^{+}10] William I. Thacker, Jingwei Zhang, Laynet Watson, Jeffrey B. Birch, Manjula A. Iyer, and Michael W. Berry. Algorithm 905: Modified Shepard algorithm for interpolation of scattered multivariate data. ACM Transactions on Mathematical Software, 37(3):34:1– 34:20, September 2010. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Udegbunam:1991:FPI

[Ude91] Emmanuel O. Udegbunam. A FORTRAN program for interpretation of relative permeability from unsteady-state displacements with capillary pressure included. Computers and Geosciences, 17(10): 1351–1358, 1991. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Ueberhuber:1997:NCM

[Ueb97] Christoph W. Ueberhuber. Numerical Computation: Methods, Software, and Analysis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-62058-3 (vol. 1: softcover), 3-540-62057-5 (vol. 2: softcover), 3-642-59118-3 (ebook). xvi + 474 (vol. 1), xvi + 495 (vol. 2) pp. LCCN QA297 .U2413 1997. US$44.95 (vol. 1), US$49.95 (vol. 2).

Utter-Honig:1991:GAP

[UHP91] S. Utter-Honig and C. M. Pancake. Graphical animation of parallel Fortran programs. In IEEE [IEE91], pages 491–500. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Uberhuber:1993:SEF

[UM93] Christoph Uberhuber and Peter Meditz. Software-Entwicklung in Fortran 90. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., March 1993. ISBN 0-387-82450-2. US$40.00. URL http:

                                        //www.cbooks.com/sqlnut/
                                        SP/search/gtsumt?source=&
                                        isbn=0387824502.

Umemura:1991:FNL

[Ume91] K. Umemura. Floating-point number LISP. Software—Practice and Experience, 21(10): 1015–1026, October 1991. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ulberg:1994:BRC

[UMM94] Dima E. Ulberg, A. John Mallinckrodt, and Susan McKay. Book review: Charles H. Koelbel, David B. Loveman, Robert S. Schreiber, Guy L. Steele, Jr., and Mary E. Zosel, The High Performance Fortran Handbook. Computers in physics, 8(4):428–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.4823319.

Utke:2008:OFM

[UNF^{+}08] Jean Utke, Uwe Naumann, Mike Fagan, Nathan Tallent, Michelle Strout, Patrick Heimbach, Chris Hill, and Carl Wunsch. OpenAD/F: a modular open-source tool for automatic differentiation of Fortran codes. ACM Transactions on Mathematical Software, 34 (4):18:1–18:36, July 2008. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

USEPA:1993:HSP

[Uni93] United States.Environmental Protection Agency. Hydrological simulation program — FORTRAN: user’s manual for release 10. Technical report, Environmental Research Laboratory, Office of Research and Development, U.S. Environmental Protection Agency, Athens, GA, USA, 1993. v + 660 pp.

UCOCD:19xx:FCP

[Unixx] United States.Office of Civil Defense. A Fortran computer program for radiation shielding analysis of structures. Praeger-Kavanagh-Waterbury, Engineers-Architects, ????, 19xx. 135 pp.

Strathclyde:1992:GAF

[Uni2 ] University of Strathclyde. General advice for FORTRAN programmers. Technical report, University of Strathclyde, Glasgow, Scotland, 1992 (??). 2 pp.

USNRC:2001:TMFa

[U.S01a] U.S.Nuclear Regulatory Commission. Trac-M/Fortran 90 (Version 3.0) programmer’s manual. Technical Report NUREG/CR-6725, U.S. Nuclear Regulatory Commission, Washington, DC, USA, 2001.

USNRC:2001:TMFb

[U.S01b] U.S.Nuclear Regulatory Commission. Trac-M/Fortran 90 (Version 3.0) theory manual. Technical Report NUREG/CR-6724, U.S. Nuclear Regulatory Commission, Washington, DC, USA, 2001.

USNRC:2001:TMFc

[U.S01c] U.S.Nuclear Regulatory Commission. Trac-M/Fortran 90 (Version 3.0) user’s manual.

Technical Report NUREG/CR-6722, U.S. Nuclear Regulatory Commission, Washington, DC, USA, 2001.

USENIX:1994:PUSb

[USE94] USENIX Association, editor.

Proceedings of the USENIX Symposium on Very High Level Languages (VHLL): October 26–28, 1994, Santa Fe, New Mexico, USA. USENIX Association, Berkeley, CA, USA, 1994. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

Utter:1990:VSP

[Utt90] Paula Sue Utter. A visualization system for parallel Fortran programs. Thesis (M.S.), Auburn University, Auburn, AL, USA, 1990. xiii + 163 pp.

Ujaldon:1995:NDL

[UZCZ95] M. Ujaldon, E. L. Zapata, B. M. Chapman, and H. P. Zima. New data-parallel language features for sparse matrix computations. In IEEE [IEE95a], pages 742–749. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

Ujaldon:1996:DLF

[UZCZ96] M. Ujaldon, E. L. Zapata, B. M. Chapman, and H. P. Zima. Data-parallel language features for sparse codes. In Szymanski and Sinharoy [SS96], pages 253–264. ISBN 0-7923-9635-9. LCCN QA76.58.L37 1996.

Ujaldon:1997:VFH

[UZCZ97] M. Ujaldon, E. L. Zapata, B. M. Chapman, and H. P. Zima. Vienna-Fortran/HPF extensions for sparse and irregular problems and their compilation. IEEE Transactions on Parallel and Distributed Systems, 8(10):1068– 1083, October 1997. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.
org/td/books/td1997/pdf/
l1068.pdf;   http://www.
computer.org/tpds/td1997/
l1068abs.htm.

Vetterling:1993:NREa

[V^{+}93] William T. Vetterling et al. Numerical Recipes Example Book. Cambridge University Press, Cambridge, UK, March 1993. ISBN 0-521-31330-9. viii + 179 pp. LCCN QA76.73.F25N84

1985.  URL http://www.
nr.com/nronline_switcher.
html.

Vaidyanathan:1993:MFW

[Vai93] Shankar Vaidyanathan. Multitasking Fortran and Windows NT. Dr. Dobb’s Journal of Software Tools, 18(10): 21–??, October 1993. CO-DEN DDJOEB. ISSN 1044-789X. Special Issue: Windows Sourcebook.

Vajapeyam:1992:ILC

[Vaj92] Sriram Vajapeyam. Instruction level characterization of the Cray Y-MP processor. Technical Report TR 1086, Computer Sciences Department, University of Wisconsin, Madison, WI, USA, May 1992.

VanTuyl:1984:EF

[Van84] Robert R. Van Tuyl. On evolution of FORTRAN. ACM SIG-PLAN Notices, 19(11):52–55, November 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanKraalingen:1990:FVC

[van90a] D. W. G. van Kraalingen. The FORTRAN version of CSMP MACROS (Modules for Annual CRop Simulation). Simulation report cabo-tt; nr. 21, Centre for Agrobiological Research (CABO) and Dept. of Theoretical Production Ecology (TPE), Agricultural University, Wageningen, The Netherlands, 1990. 145 pp.

vanMechelen:1990:FPD

[van90b] Iven van Mechelen. A FOR-TRAN program for the detection of logical relations between a set of predictors and a criterion variable. Multivariate behavioral research, 25(2): 207–??, April 1990. CODEN MVBRAV. ISSN 0027-3171.

vanWaveren:1994:HPF

[van94a] G. M. van Waveren. High Performance Fortran. In Gentzsch and Harms [GH94a], pages 429–433. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994).

Vanderlip:1994:PSV

[Van94b] J. Vanderlip. Pacific Sierra’s VAST-HPF and VAST/77toHPF. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):74, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Vandoni:1995:SCA

[Van95] C. E. Vandoni, editor. 1995 CERN School of Computing: Arles, France, 20 August – 2 September 1995: proceedings, number 5 in CERN European Organization for Nuclear Research — Reports. CERN, Geneva, Switzerland, 1995. ISBN 92-9083-076-X. ISSN 0007-8328. LCCN QC770 .E83 v.95, no. 5.

Vanderbauwhede:2022:MLF

[Van22] Wim Vanderbauwhede. Making legacy Fortran code type safe through automated program transformation. The Journal of supercomputing, 78 (2):2988–3028, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s11227-021-03839-9.

Vardi:1995:ISC

[Var95] Ilan Vardi. Introduction to Symbolic Computation. TE-LOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1995. ISBN 0-387-94147-9. 320 pp. LCCN ???? US$49.95. Covers C, Fortran, Maple, and Mathematica.

Varga:1997:CMF

[Var97] Kálmán Varga. A combined Mathematica–Fortran program package for analytical calculation of the matrix elements of the microscopic cluster model. Computer Physics Communications, 104(1–3):259–274, August 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465597000313.

Vouk:1995:EEL

[VBA95] M. A. Vouk, R. Balay, and J. Amborosiano. EDSS — an environment for largescale numerical computing and decision making. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/
proceedings/vouk/vouk1.html.

Villaverde:2018:PTI

[VBB18] Alejandro F. Villaverde, Kolja Becker, and Julio R. Banga. PREMER: a tool to infer biological networks. IEEE/ ACM Transactions on Computational Biology and Bioinformatics, 15(4):1193–1202, July 2018. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Verschaeren:1997:NPF

[VCV97a] Dennis Verschaeren, Annie Cuyt, and Brigitte Verdonk. On the need for predictable floating-point arithmetic in the programming languages Fortran 90 and C/C++. ACM SIGPLAN Notices, 32(3):57– 64, March 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vershaeren:1997:NPF

[VCV97b] Dennis Vershaeren, Annie Cuyt, and Brigitte Verdonk. On the need for predictable floating-point arithmetic in the programming languages Fortran 90 and C/C++. ACM SIGPLAN Notices, 32(3):57– ??, March 1997. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanReeuwijk:1996:IFH

[vDSP96] Kees van Reeuwijk, Will Denissen, Henk J. Sips, and Edwin M. R. M. Paalvast. An implementation framework for HPF distributed arrays on message-passing parallel computer systems. IEEE Transactions on Parallel and Distributed Systems, 7(9):897–914, September 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://www.computer.
org/tpds/td1996/l0897abs.
htm.

Veen:1994:PHP

[Vee94] A. H. Veen. The Prepare HPF programming environment. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):73, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Veldhuizen:1997:SCC

[Vel97] Todd Veldhuizen. Scientific computing: C++ versus Fortran: C++ has more than caught up. Dr. Dobb’s Journal of Software Tools, 22(11): 34, 36–38, 91, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Vesely:1991:FCS

[Ves91] Vítězslav Veselý. Fast cellstructured algorithm for digit reversal of arbitrary length. SIAM Journal on Scientific and Statistical Computing, 12 (2):298–310, March 1991. CO-DEN SIJCD4. ISSN 0196-5204.

Vetterling:1993:NREb

[Vet93] William T. Vetterling. Numerical recipes example book (Fortran). Cambridge University Press, Cambridge, UK, second edition, 1993. ISBN 0-521-43721-0 (paperback). viii + 245 pp. LCCN QA76.76.F25 N85

vonHippel:2006:TAO

[vH06] G. M. von Hippel. TaylUR, an arbitrary-order diagonal automatic differentiation package for Fortran 95. Computer Physics Communications, 174(7):569–576, April 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465506000166.

vonHippel:2007:NVA

[vH07] G. M. von Hippel. New version announcement for TaylUR, an arbitrary-order diagonal automatic differentiation package for Fortran 95. Computer Physics Communications, 176(11–12):710–711, June 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S001046550700197X.

vonHippel:2010:TMA

[vH10] G. M. von Hippel. TaylUR 3, a multivariate arbitraryorder automatic differentiation package for Fortran 95. Computer Physics Communications, 181(3):705–706, March 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465509003464.

vonHanxleden:2000:BCP

[vHK00] Reinhard von Hanxleden and Ken Kennedy. A balanced code placement framework. ACM Transactions on Programming Languages and Systems, 22(5):816–860, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL

http://www.acm.org/pubs/
articles/journals/toplas/
2000-22-5/p816-von_hanxleden/
p816-von_hanxleden.pdf;
http://www.acm.org/pubs/
citations/journals/toplas/
2000-22-5/p816-von_hanxleden/

.

Hanxleden:1994:VBD

[vHKS94a] R. v. Hanxleden, K. Kennedy, and J. Saltz. Value-based distributions in Fortran D. Lecture Notes in Computer Science, 797:434–440, 1994. CO-DEN LNCSD9. ISSN 0302-9743.

vonHanxleden:1994:VDF

[vHKS94b] R. von Hanxleden, K. Kennedy, and J. Saltz. Value-based distributions in Fortran D. In Gentzsch and Harms [GH94c], pages 434–440. ISBN 0-387-57981-8 (New York), 3-540-57981-8 (Berlin). LCCN QA76.88.I57 1994. DM96.00. Two volumes.

Vignes:1993:SAR

[Vig93] J. Vignes. A stochastic arithmetic for reliable scientific computation. Mathematics and Computers in Simulation, 35 (3):233–261, September 1993. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Vilot:1994:IST

[Vil94] M. J. Vilot. An introduction to the Standard Template Library. C++ Report, 6(8):22– 29, 35, October 1994. CODEN CRPTE7. ISSN 1040-6042.

Violett:1990:DAE

[Vio90] Duane L. Violett. Dynamic array expansion using VAX FOR-TRAN. The VAX professional, 12(6):17–??, December 1990. CODEN VAXPEN. ISSN 8750-9628.

Veldhuizen:1997:WCB

[VJ97a] T. L. Veldhuizen and M. E. Jernigan. Will C++ be faster than Fortran? Lecture Notes in Computer Science, 1343:49– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veldhuizen:1997:WCF

[VJ97b] Todd L. Veldhuizen and M. E. Jernigan. Will C++ be faster than Fortran? In Proceedings of the 1st International Scientific Computing in Object-Oriented Parallel Environments (ISCOPE’97), Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc.,

                                        1997. URL http://extreme.

indiana.edu/~tveldhui/papers/

.

vonHanxleden:1994:GTB

[vK94] Reinhard von Hanxleden and Ken Kennedy. GIVE-N-TAKE

— a balanced code placement framework. ACM SIG-PLAN Notices, 29(6):107–120, June 1994. CODEN SINODQ. ISBN 0-89791-598-4. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/178243/p107-von_hanxleden/

.

Vielhaber:1993:NVI

[VKB93] J. P. Vielhaber, J. V. Kuhlman, and J. S. Barrett. NONMEM version III implementation on a VAX 9000: a DCL procedure for single-step execution and the unrealized advantage of a vectorizing FORTRAN compiler. Computer Methods and Programs in Biomedicine, 40(2):103–115, June 1993. CO-DEN CMPBEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

vonHanxleden:1992:CAIa

[vKK92] Reinhard von Hanxleden, Ken Kennedy, and Charles Koelbel. Compiler analysis for irregular problems in Fortran D. Technical report COMP TR92-198, Rice University, Dept. of Computer Science, Houston, TX, USA, 1992. 15 pp.

vonHanxleden:1993:CAI

[vKK^{+}93] R. von Hanxleden, K. Kennedy, C. Koelbel, R. Das, and J. Saltz. Compiler analysis for irregular problems in Fortran D. Lecture Notes in Computer Science, 757:97–111, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vonHanxleden:1994:VDA

[vKS94] R. von Hanxleden, K. Kennedy, and J. Saltz. Value-based distributions and alignments in Fortran D. Journal of Programming Languages, 2(3): 259–282, September 1994. CO-DEN JPLAER. ISSN 0963-9306.

Vesier:1992:TCM

[VLLY92] C. C. Vesier, J. D. Lemmon, Jr., R. A. Levine, and A. P. Yoganathan. A three-dimensional computational model of a thinwalled left ventricle. In IEEE [IEE92d], pages 73–82. ISBN 0-8186-2630-5. LCCN QA76.5 .S894 1992. IEEE catalog no. 92CH3216-9.

Volkert:1993:PCS

[Vol93] Jens Volkert, editor. Parallel computation: Second International ACPC Conference, Gmunden, Austria, October 4– 6, 1993: proceedings, volume 734 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-57314-3 (U.S.). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I5 1993. DM58.00.

vonHanxleden:1992:CAIb

[von92] Reinhard von Hanxleden.

Compiler analysis for irregular problems in FORTRAN D. NASA contractor report NASA CR 189664; ICASE report 92-22, National Aeronautics and Space Administration, Langley Research Center; National Technical Information Service, distributor, Hampton, VA, USA, 1992. 19 pp.

vonLaszewski:1992:PBL

[vPMF92] G. von Laszewski, M. Parashar, A. G. Mohamed, and G. C. Fox. On the parallelization of blocked LU factorization algorithms on distributed memory architectures. In IEEE [IEE92d], pages 170–179. ISBN 0-8186-2630-5. LCCN QA76.5 .S894 1992. IEEE catalog no. 92CH3216-9.

Venkatachar:1997:CGB

[VRT97] A. Venkatachar, J. Ramanujam, and A. Thirumalai. Communication generation for block-cyclic distributions. Parallel Processing Letters, 7(2): 195–202, June 1997. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Vajapeyam:1991:ESC

[VSH91] Sriram Vajapeyam, Gurindar S.

Sohi, and Wei-Chung Hsu. An empirical study of the CRAY Y-MP processor using the Perfect Club benchmarks. ACM SIGARCH Computer Architecture News, 19(3):170–179, May 1991. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic). IEEE catalog number 91CH2995-9.

Vetterling:1992:NRE

[VTP92] William T. Vetterling, Saul A. Teukolsky, and William H. Press. Numerical recipes: example book (FORTRAN). Cambridge University Press, Cambridge, UK, second edition, October 1992. ISBN 0-521-43721-0. viii + 245 pp. LCCN QA76.76.F25 N85 1992. UK£19.95; US$29.95.

URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0521437210.

vanGaans:1990:MLR

[vV90] P. F. M. van Gaans and S. P. Vriend. Multiple linear regression with correlations among the predictor variables. theory and computer algorithm RIDGE (FORTRAN 77). Computers and Geosciences, 16(7):933–??, 1990. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

vanWaveren:2002:CGH

[vWAH^{+}02] Matthijs van Waveren, Cliff Addison, Peter Harrison, Dave Orange, Norman Brown, and Hidetoshi Iwashita. Code generator for the HPF Library and Fortran 95 transformational functions. Concurrency and Computation: Practice and Experience, 14(8–9): 589–602, July/August 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/95016135/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=95016135{\&}PLACEBO=IE.
pdf.

Wagener:1995:SIF

[W^{+}95] J. L. Wagener et al. Special issue: Fortran 95 Committee Draft, May 1995 (X3J3/95-007R1, SC22/WG56-N1122). ACM SIGPLAN FORTRAN Forum, 14(2):xviii + 354, June 1995. ISSN 1061-7264 (print), 1931-1311 (electronic).

Wagenbreth:1994:AAH

[Wag94] G. Wagenbreth. APR’s approach to High Performance Fortran for distributed memory multiprocessor systems. In IEEE [IEE94a], pages 41–45. ISBN 0-8186-6965-9. LCCN QA76.58.S95 1994. IEEE catalog no. 95TH8024.

Wasniewski:1998:RFC

[WAG98] J. Wasniewski, B. S. Andersen, and F. Gustavson. Recursive formulation of Cholesky algorithm in Fortran 90. Lecture Notes in Computer Science, 1541:574–578, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walsh:1990:LEQ

[Wal90] John F. Walsh. Locating errors in questionnaire style data: a FORTRAN program. Perceptual and Motor Skills, 71(2): 537–??, October 1990. CODEN PMOSAZ. ISSN 0031-5125.

Wallich:1991:FF

[Wal91a] Paul Wallich. Fortran forever. Scientific American, 265 (1):112, July 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

                                        URL http://www.nature.
                                        com/scientificamerican/journal/
                                        v265/n1/pdf/scientificamerican0791-
                                        112.pdf.

Walsh:1991:MFR

[Wal91b] John F. Walsh. Microsoft’s FORTRAN 5.0 random number generator: be aware. Perceptual and Motor Skills, 72(1): 257–??, February 1991. CO-DEN PMOSAZ. ISSN 0031-5125.

Walsh:1992:SPG

[Wal92] John F. Walsh. A simple procedure for generating nonnormal data sets: A FORTRAN program. Teaching of Psychology, 19(4):243–244, December 1992. ISSN 0098-6283.

Walter:1993:AXF

[Wal93a] W. V. Walter. ACRITH-XSC a Fortran-like language for verified scientific computing. In Adams and Kulisch [AK93], pages 45–70. ISBN 0-12-044210-8. LCCN QA76.S368 1993.

Walter:1993:FXP

[Wal93b] W. V. Walter. FORTRAN-XSC. A portable Fortran 90 module library for accurate and reliable scientific computing. In Albrecht et al. [AAS93], pages 265–285. CO-DEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Walster:2000:UII

[Wal00] G. William Walster. The use and implementation of interval data types in Fortran. ACM SIGPLAN FORTRAN Forum, 19(2):2–15, August 2000. ISSN 1061-7264 (print), 1931-1311 (electronic). Describes the interval arithmetic support in the Sun Microsystems Forte Developer 6 Fortran 95 (f95) compiler.

Walster:2001:IAF

[Wal01] G. William Walster. Interval angles and the Fortran ATAN2 intrinsic function. Report, Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, April 14, 2001. 17 pp.

Wallcraft:2002:CCA

[Wal02a] Alan J. Wallcraft. A comparison of Co-Array Fortran and OpenMP Fortran for SPMD programming. The Journal of supercomputing, 22(3):231–250, July 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//ipsapp008.kluweronline.
com/content/getfile/5189/
36/1/abstract.htm;   http:
//ipsapp008.kluweronline.
com/content/getfile/5189/
36/1/fulltext.pdf.

Walster:2002:IAF

[Wal02b] G. William Walster. Interval angles and the Fortran ATAN2 intrinsic function. Report, Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, May 17, 2002. 17

pp. URL http://developers.
sun.com/solaris/articles/
int_angles/interval-angles.
pdf.

Wampler:1990:OOP

[Wam90a]  K. Dean Wampler.  The

object-oriented programming paradigm (OOPP) and FOR-TRAN programs. Computers in physics, 4(4):385–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.168371.

Wampler:1990:OPP

[Wam90b]  K. Dean Wampler.  The

object-oriented programming paradigm (OOPP) and FOR-TRAN programs. Computers in physics, 4(4):385–??, 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Waligora:1997:IAO

[WBS97] Sharon Waligora, John Bailey, and Mike Stark. The impact of Ada and object-oriented design in NASA Goddard’s Flight Dynamics Division. ACM SIGADA Ada Letters, 17(3): 67–86, May/June 1997. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). Extensive study of a decade of large system software development in Ada, Fortran, C, and C++, with the finding that development is moving away from Ada to C and C++ on two main grounds: lack of adequate software development environments for Ada, and high cost (3 to 8 times per seat).

Winer:1992:PMB

[WC92] Ethan Winer and Phil Cramer.

PC Magazine Basic: Lab Notes. Ziff-Davis Press, Emeryville, CA, USA, November 1, 1992. ISBN 1-56276-068-8. Includes disk.

Williams:1992:TFP

[WCN92] C. J. Williams, J. C. Christian, and J. A. Norton, Jr. TWINAN90: a FOR-TRAN program for conducting ANOVA-based and likelihoodbased analyses of twin data. Computer Methods and Programs in Biomedicine, 38(2– 3):167–176, July 1992. CO-DEN CMPBEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Wasniewski:1998:HPLb

[WD98] J. Waśniewski and J. Dongarra. High performance linear algebra package for FOR-TRAN 90. Lecture Notes in Computer Science, 1541:579– 581, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weatherford:1994:HPE

[Wea94] Stephen Andrew Weatherford. High-level pattern-matching extensions to C++ for Fortran program manipulation in Polaris. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana, IL, USA, 1994. viii + 104 pp.

Weinman:1991:VFa

[Wei91a] David G. Weinman. VAX Fortran. The Boyd & Fraser programming language series. PWS-Kent Pub. Co., Boston, MA, USA, May 1991. ISBN 0-87835-172-8. xiii + 450 pp. LCCN QA76.8.V37W45 1986.

Weinman:1991:VFb

[Wei91b] David G. Weinman. VAX FORTRAN. PWS-Kent series in engineering and computer science. PWS-Kent Pub. Co., Boston, MA, USA, second edition, 1991. ISBN 0-534-92565-0. xiii + 479 pp. LCCN QA76.8.V37W45 1991.

Weinman:1991:VFP

[Wei91c] David G. Weinman. Vax Fortran. Pws-Kent Series in Engineering and Computer Science. PWS-Kent Pub. Co., Boston, MA, USA, May 1991. ISBN 0-534-92565-0. US$44.95.

Weinman:1993:VF

[Wei93] David G. Weinman. VAX FORTRAN. PWS-Kent series in engineering and computer science. PWS Publishing Co., Boston, MA, USA, second edition, 1993. ISBN 0-534-92565-0 (paperback). xiii + 479 pp. LCCN QA76.8.V37W45 1991.

Wei:1994:BRI

[Wei94] Sha Xin Wei. Book review:

InterCall 2.1: The tortoise and the hare. Mathematica Journal, 4(1):??, Winter 1994. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (elec-

tronic). URL http://www.
mathematica-journal.com/
issue/v4i1/reviews/wei/22-
25_xinwei.mj.pdf;   http:
//www.mathematica-journal.
com/issue/v4i1/reviews/wei/
index.html.

Weisfeld:1995:PSH

[Wei95] Matt Weisfeld. Portable screen handling. Dr. Dobb’s Journal of Software Tools, 20(5):142– 143, 146, 148–149, May 1995. CODEN DDJOEB. ISSN 1044-789X.

Weste:1996:WFM

[Wes96] Michael Weste. Writing a Fortran 90 Makefile maker in Perl. Fortran Journal, 8(3): 8–12, May/June 1996. ISSN 1060-0221.

Wilson:1994:SIR

[WFW^{+}94] Robert P. Wilson, Robert S. French, Christopher S. Wilson, Saman P. Amarasinghe, Jennifer M. Anderson, Steve W. K. Tjiang, Shih-Wei Liao, Chau-Wen Tseng, Mary W. Hall, Monica S. Lam, and John L. Hennessy. SUIF: an infrastructure for research on parallelizing and optimizing compilers. ACM SIGPLAN Notices, 29(12):31–37, December 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weideman:1992:UGRa

[WHL92a] Mark H. Weideman, Vince H. Hammond, and Alfred C. Loos. User’s guide to resin infusion simulation program in the Fortran language. Vpi-e; 92-04 ccms; 92-03 interim report / nasa-virginia tech composites program; 88 ccms (series); 92-03. interim report (nasavirginia tech composites program); 88., College of Engineering, Virginia Polytechnic Institute and State University, Blacksburg, VA, USA, 1992. i + 82 pp.

Weideman:1992:UGRb

[WHL92b] Mark H. Weideman, Vince H.

Hammond, and Alfred C. Loos. User’s guide to resin infusion simulation program in the Fortran language. [NASA contractor report]; NASA CR-189928, Virginia Polytechnic Institute and State University; National Aeronautics and Space Administration; National Technical Information Service, distributor, Blacksburg, VA, USA, 1992. ???? pp.

Wang:1995:NFP

[WHL95] J.-J. Wang, Q.-Z. Han, and Y.- X. Liu. A new Fortran program for CFPs of an identical fermion system. Computer Physics Communications, 85 (1):99–109, January 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Wolfe:1994:AAA

[WI94] M. Wolfe and M. Ikei. Automatic array alignment for distributed memory multicomputers. In Hesham and Shriver [HS94b], pages 23–32. ISBN 0-8186-5060-5. ISSN 1060-3425. LCCN ???? IEEE catalog no. 94TH0607-2.

Wichmann:1989:SPI

[Wic89] B. A. Wichmann. Scientific processing in ISO-Pascal: a proposal to get the benefits of mixed precision floating-point. ACM SIGPLAN Notices, 24 (6):20–22, June 1989. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wicker:1999:SSW

[Wic99] Louis J. Wicker. Simulating severe weather. Dr. Dobb’s Journal of Software Tools, 24(3):18– 20, 22, 24, 26–28, March 1999. CODEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/1999/9902/9902toc.htm.

Wieseman:1994:RCR

[Wie94] C. D. Wieseman, editor. Role of computers in research and development at Langley Research Center: Workshop entitled “The role of computers in LARC R and D” — June 1994, Hampton, VA, number 10159 in NASA Conference Publication. National Aeronautics and Space Administration, Washington, DC, USA, 1994. ISBN ???? ISSN 0191-7811. LCCN ????

Wieder:1999:ANH

[Wie99] Thomas Wieder. Algorithm 794: Numerical Hankel transform by the Fortran program HANKEL. ACM Transactions on Mathematical Software, 25 (2):240–250, June 1999. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

                                        tronic). URL ftp://netlib.
                                        bell-labs.com/netlib/toms/
                                        794.gz;   http://doi.acm.
                                        org/10.1145/317275.317284;
                                         http://phase.etl.go.jp/
                                        netlib/toms/794;   http:/
                                        /www.acm.org/pubs/citations/
                                        journals/toms/1999-25-2/
                                        p240-wieder/;   http://www.
                                        acm.org/pubs/citations/journals/

toms/cgi-bin/TOMSbibget?
Wieder:1999:ANH; http://
www.hensa.ac.uk/netlib/toms/
794.gz;   http://www.netlib.
no/netlib/toms/794; http:
//www.netlib.org/toms/794.

Wilkes:1993:CPF

[Wil93] Maurice V. Wilkes. Computing perspectives: From Fortran and Algol to object-oriented languages. Comm. ACM, 36 (7):21–23, July 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-
0782/159553.html.

Wille:1995:ASFa

[Wil95a] David R. Wille. Advanced Scientific Fortran. John Wiley and Sons, New York, London, Sydney, 1995. ISBN 0-471-95383-0. xvii + 234 pp. LCCN QA76.73.F25 W55 1995. US$62.00; US$49.95.

URL http://www.cbooks.
com/sqlnut/SP/search/gtsumt?
source=&isbn=0471953830.

Wille:1995:ASFb

[Wil95b] David R. Wille. Advanced scientific Fortran. John Wiley and Sons, New York, London, Sydney, 1995. ISBN 0-471-95383-0. xvii + 234 pp. LCCN QA76.73.F25 W55 1995.

Weickmann:1994:FPP

[WJ94] Ann Weickmann and R. Michael Jones. A FORTRAN program for performing Abel transforms. Noaa technical memorandum erl etl; 244, U.S. Dept. of Commerce, National Oceanic and Atmospheric Administration, Environmental Research Laboratories, Environmental Technology Laboratory, Boulder, CO, USA, 1994. iv + 158 pp.

Wang:2004:BBS

[WKM04] R. Wang, P. Keast, and P. Muir. BACOL: B-spline adaptive collocation software for 1-D parabolic PDEs. ACM Transactions on Mathematical Software, 30(4):454–470, December 2004. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Walker:1997:NCF

[WMCU97] Edward Walker, Gary Morgan, Bruce Cass, and Zygmunt Ulanowski. A note on compiling FORTRAN loop kernels onto a dataflow architecture. Parallel Computing, 22(11): 1545–1557, January 26, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1997&
volume=22&issue=11&aid=1103.

Watson:1997:ASF

[WMMW97] Layne T. Watson, Robert C. Melville, Alexander P. Morgan, and Homer F. Walker. Algorithm 777. HOMPACK90: a suite of Fortran 90 codes for globally convergent homotopy algorithms. ACM Transactions on Mathematical Software, 23(4):514–549, December 1997. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Watanabe:1990:IPI

[WN90] Shunro Watanabe and Morio Nagata, editors. ISSAC ’90: proceedings of the International Symposium on Symbolic and Algebraic Computation: August 20–24, 1990, Tokyo, Japan. ACM Press and Addison-Wesley, New York, NY 10036, USA and Reading, MA, USA, 1990. ISBN 0-89791-401-5 (ACM), 0-201-54892-5 (Addison-Wesley). LCCN QA76.95 .I57 1990.

Watanabe:1994:MSP

[WNO94] T. (Tsutomu) Watanabe, Makoto Natori, and Tsutomu Oguni. Mathematical Software for the P.C. and Work Stations: a Collection of Fortran 77 Programs. North-Holland Publishing Co., Amsterdam, The Netherlands, June 1994. ISBN 0-444-82000-0. xiv + 387 pp. LCCN QA 76.73 F25 F6813 1994. US$178.50. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0444820000. Translation

of: FORTRAN 77 ni yoru suchi keisan sofutowea.

Walker:1996:RBC

[WO96] D. W. Walker and S. W. Otto.

Redistribution of block-cyclic data distributions using MPI. Concurrency, practice and experience, 8(9):707–728, November 1996. CODEN CPEXEI. ISSN 1040-3108. URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        23305.

Wolf:1991:FSC

[Wol91]  Gert W. Wolf.  A FOR-

TRAN subroutine for cartographic generalization. Computers and Geosciences, 17 (10):1359–1382, 1991. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Wollan:1992:PRN

[Wol92] Peter C. Wollan. A portable random number generator for parallel computers. Communications in Statistics: Simulation and Computation, 21(4): 1247–1254, ???? 1992. CO-DEN CSSCDB. ISSN 0361-0918.

Won:1993:EPF

[WR93] Sik Choi Won and Judi Repman. Effects of Pascal and FORTRAN programming on the problem-solving abilities of college students. Journal of Research on Computing in Education, 25(3):290–302, Spring 1993. CODEN JRCEE8. ISSN 0888-6504.

Wright:1990:OPM

[Wri90a] E. Alaphia Wright. Open pit mine design models: an introduction with FORTRAN/ 77 programs. Series on mining engineering; v. 8. Trans Tech Publications, Clausthal-Zellerfeld, Federal Republic of Germany, 1990. ISBN ???? xi + 187 pp. LCCN ????

Wright:1990:NCU

[Wri90b] F. Wright. The ‘effective number of codons’ used in a gene. Gene, 87(1):23–29, March 1, 1990. CODEN GENED6. ISSN 0378-1119.

Wright:1991:OPM

[Wri91] E. Alphia Wright. Open Pit Mine Design Models: An Introduction With Fortran 77 Programs, volume 8 of Series on Mining Engineering. Trans Tech Publications, Ltd., Clausthal, Germany; Rockport, MA, USA, July 1991. ISBN 0-87849-083-3. US$75.00. URL http:

//www.cbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=0878490833.

Wright:1999:PSL

[Wri99] Ed Wright. Porting the SPICE library. Dr. Dobb’s Journal of Software Tools, 24(12):94, 96–99, December 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_12/
doc2c.zip;   http://www.
ddj.com/ftp/1999/1999_12/
spice.txt.

Wagemann:1990:CLC

[WRL90] R. Wagemann, G. Regehr, and R. Lypka. Compiler listing for chemical speciation program: MACS80.VERS.4.1 for Microsoft: Fortran version 5 for MS-DOS. Canadian data report of fisheries and aquatic sciences 818, Dept. of Fisheries and Oceans, Central and Arctic Region, Winnipeg, Manitoba, Canada, December 1990. x + 190 pp.

Wang:1994:MFP

[WS94] M. C. Wang and N. C. Silver. A Microsoft FORTRAN 77 program for determining the confidence interval around the estimate of the population correlation coefficient for the vote-counting method. Educational and Psychological Measurement, 54(1):105–??, Spring 1994. CODEN EPMEAJ. ISSN 0013-1644 (print), 1552-3888 (electronic).

Wholey:1994:TMP

[WSL94] S. Wholey, R. Shapiro, and D. Loshin. Thinking Machines’ plans for HPF. IEEE Parallel and Distributed Technology: Systems and Applications, 2(3):76, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Wise:2000:APP

[WSW00] Steven M. Wise, Andrew J. Sommese, and Layne T. Watson. Algorithm 801: POL-SYS PLP: a partitioned linear product homotopy code for solving polynomial systems of equations. ACM Transactions on Mathematical Software, 26(1):176–200, March 2000. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
2000-26-1/p176-wise/;   http:
//www.acm.org/pubs/citations/
journals/toms/2000-26-1/
p176-wise/p176-wise.pdf.

White:1990:PCA

[WTW90] S. W. White, D. C. Torney, and C. C. Whittaker. A parallel computational approach using a cluster of IBM ES/ 3090 600Js for physical mapping of chromosomes. In IEEE [IEE90a], pages 112– 121. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Wu:1993:NRF

[Wu93] J. Wu. Numerical Recipes in Fortran; Second Edition. William H Press et al. Applied mechanics reviews, 46(7):B97, July 1993. CODEN AMREAD. ISSN 0003-6900.

Williams:1990:ISN

[WW90] M. A. Williams and D. G.

Wilson. Iterative solution of a nonlinear system arising in phase-change problems. SIAM Journal on Scientific and Statistical Computing, 11(6):1087– 1101, November 1990. CODEN SIJCD4. ISSN 0196-5204.

Weerawarana:1992:PCG

[WW92] Sanjiva Weerawarana and Paul S. Wang. A portable code generator for CRAY FORTRAN. ACM Transactions on Mathematical Software, 18(3):241–255, September 1992. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0098-3500/
                                        131767.html.

Wolfram:1993:STN

[WW93] Stephen Wolfram and Wolfram Research, Inc. Selected Tutorial Notes. Wolfram Research, Inc., 100 Trade Center Drive, Champaign, IL 61820-7237, USA, 1993. ISBN 1-880083-07-8. 405 pp. LCCN ????

Wakatani:1994:NAA

[WW94] A. Wakatani and M. Wolfe. A new approach to array redistribution: strip mining redistribution. In Halatsis et al. [HMPT94], pages 323– 335. ISBN 3-540-58184-7. LCCN QA76.58 .I564 1994.

Wakatani:1995:OAR

[WW95] Akiyoshi Wakatani and Michael Wolfe. Optimization of array redistribution for distributed memory multicomputers. Parallel Computing, 21(9):1485– 1490, September 12, 1995. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1995&
volume=21&issue=9&aid=1006.

Wan:2014:FFP

[WW14] Lu-Ping Wan and Jian-Xiong Wang. FDCHQHP: a Fortran package for heavy quarkonium hadroproduction. Computer Physics Communications, 185 (11):2939–2949, November 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465514002276.

Waheed:1999:PNB

[WYJ99] Abdul Waheed, Jerry Yan, and Haoqiang Jin. Parallelization of NAS benchmarks for shared memory multiprocessors. Future Generation Computer Systems, 15(3):353–??, ???? 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Xu:1990:MLC

[XH90] Z. Xu and K. Hwang. Molecule:

a language construct for development of parallel programs. IEEE Transactions on Software Engineering, SE-16(5): 587–599, May 1990. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Xiong:2024:GSH

[XHY^{+}24] Qiyang Xiong, Shiyong Huang, Zhigang Yuan, Bharatkumar Sharma, Lvlin Kuang, Kui Jiang, and Lin Yu. GPIC: a set of high-efficiency CUDA Fortran code using GPU for particle-in-cell simulation in space physics. Computer Physics Communications, 295 (??):Article 108994, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0010465523003399.

Xu:1993:AFP

[Xu93] Y. Xu. ARLOSS: a FOR-TRAN program for modeling the effects of initial ^{40}Ar losses on ^{40}Ar,^{39}Ar dating. Computers and Geosciences, 19(4): 533–??, April 1993. CODEN CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Xiang:1995:FSM

[XWK95] Y. Xiang, C. M. Wang, and S. Kitipornchai. FORTRAN subroutines for mathematical operations on polynomial functions. Computers and structures, 56(4):541–??, 1995. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Yamamoto:1995:NSL

[Yam95]  Tetsuro Yamamoto.  Non-

linear SOR-like methods and their applications. In IFIP Working Group 2.5 [IFI95], page ?? ISBN ???? LCCN ???? URL http:

//www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/
proceedings/yamamoto/yamamoto1.
html.

Yan:1994:PTA

[Yan94a] J. C. Yan. Performance tuning with AIMS — an Automated Instrumentation and Monitoring System for multicomputers. In Hesham and Shriver [HS94a], pages 625–633. ISBN 0-8186-5060-5. ISSN 1060-3425. LCCN ???? IEEE catalog no. 94TH0607-2.

Yanik:1994:BRB

[Yan94b] Elizabeth Greenwell Yanik.

Book review: Numerical Recipes in FORTRAN-The Art of Scientific Computing 2nd Ed. (W. H. Press, W. T. Vetterling, S. A. Teukolsky and B. P. Flannery). SIAM Review, 36(1):149, March 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Yang:1995:RMV

[Yan95] Ying Yang. Random-effects model validation: a program package using a FORTRAN program (REML) and S-plus. Pittsburgh, PA, USA, 1995. v + 126 pp.

Yousif:1992:FCS

[YB92] H. A. Yousif and E. Boutros. A FORTRAN code for the scattering of EM plane waves by an infinitely long cylinder at oblique incidence. Computer Physics Communications, 69 (2–3):406–414, March/April 1992. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Yu:2013:DST

[YB13] Wenbin Yu and Maxwell Blair. DNAD, a simple tool for automatic differentiation of Fortran codes using dual numbers. Computer Physics Communications, 184(5):1446–1452, May 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465513000027.

Yang:2014:PMI

[YBMCB14] Chaoran Yang, Wesley Bland, John Mellor-Crummey, and Pavan Balaji. Portable, MPIinteroperable Coarray Fortran. ACM SIGPLAN Notices, 49 (8):81–92, August 2014. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yau:1997:EHP

[YFH97] H. W. Yau, G. C. Fox, and K. A. Hawick. Evaluation of High Performance Fortran through application kernels. Lecture Notes in Computer Science, 1225:772–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1994:HPF

[YGS^{+}94] S. X. Yang, D. Gannon, S. Srinivas, F. Bodin, and P. Bode. High Performance Fortran interface to the parallel C++. In IEEE [IEE94d], pages 301–308. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994.

Yi:1993:AGM

[YH93] Kwangkeun Yi and Williams Ludwell Harrison, III. Automatic generation and management of interprocedural program analyses. In ACM SIG-PLAN POPL ’93 [ACM93c], pages 246–259. ISBN 0-89791-560-7 (soft cover), 0-89791-561-5 (series hard cover). LCCN QA76.7 .A15 1993.

URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/158511/p246-yi/. ACM

order number 549930.

Yip:1990:FCG

[Yip90] Simon P. Yip. Fortran 77, a concise guide. Forward Information, Melbourne, Victoria, Australia, 1990. ISBN 1-875493-00-X. 106 pp.

Yousif:1990:FCS

[YK90] Hashim A. Yousif and Sigurd Köhler. A Fortran code for the scattering of EM plane waves by two cylinders at normal incidence. Computer Physics Communications, 59(2):371–385, June 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0010465590901854.

Kuo:19xx:FSF

[yKxx] Fu yin Kuo. FORTRAN suan fa hui pien. Kuo fang kung yeh chu pan she: Hsin hua shu tien Pei-ching fa hsing so fa hsing, Peking, China, ti 1 pan edition, 19xx. ISBN ???? various pp. LCCN ????

Yoshida:1996:DFM

[YKK96] A. Yoshida, K. Koshizuka, and H. Kasahara. Datalocalization for Fortran macrodataflow computation using partial static task assignment. In ACM [ACM96a], pages 61–68. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Yang:1995:PCM

[YO95] Bwolen Yang and D. R. O’Hallaron. Procedure call models for distributed parameters in data parallel programs. In IEEE [IEE95b], pages 157– 164. ISBN 0-8186-6895-4. LCCN QA76.58 .S34 1994.

Yu:2002:CDP

[YRF02] T.-T. Yu, J. B. Rundle, and J. Fernandez. Corrigendum to “Deformation produced by a rectangular dipping fault in a viscoelastic-gravitational layered earth model. Part II: strike-slip fault-STRGRV and STRGRH FORTRAN programs” [Computers and Geosciences 22 (1996) 751– 764]. Computers and Geosciences, 28(1):89–91, February 2002. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Young-S:2017:OGI

[YSMA^{+}17] Luis E. Young-S., Paulsamy Muruganandam, Sadhan K. Adhikari, Vladimir Loncar, Dusan Vudragović, and Antun Balaz. OpenMP GNU and Intel Fortran programs for solving the time-dependent Gross– Pitaevskii equation. Computer Physics Communications, 220 (??):503–506, November 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465517302321.

Young-S:2023:OFP

[YSMBA23] Luis E. Young-S., Paulsamy Muruganandam, Antun Balaz, and Sadhan K. Adhikari. OpenMP Fortran programs for solving the time-dependent dipolar Gross–Pitaevskii equation. Computer Physics Communications, 286(??):Article 108669, May 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0010465523000140.

Young-S:2016:OFP

[YSVM^{+}16] Luis E. Young-S., Dusan Vudragović, Paulsamy Muruganandam, Sadhan K. Adhikari, and Antun Balaz. OpenMP Fortran and C programs for solving the timedependent Gross–Pitaevskii equation in an anisotropic trap. Computer Physics Communications, 204(??):209–213, July 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S001046551630073X.

Yu:2001:SII

[Yu01] Z. W. Yu. Surface interpolation from irregularly distributed points using surface splines, with Fortran program. Computers and Geosciences, 27(7): 877–882, August 2001. CO-DEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Yang:1994:DIP

[YWS^{+}94] B. Yang, J. Webb, J. M. Stichnoth, D. R. O’Hallaron, and T. Gross. Do and Merge: integrating parallel loops and reductions. In Banerjee et al. [BGNP94], pages 169–183. ISBN 3-540-57659-2 (Berlin), 0-387-57659-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .W656 1993. DM122.00.

Yasue:1993:FCM

[YYM93] T. Yasue, H. Yamana, and Y. Muraoka. A FORTRAN compiling method for dataflow machines and its prototype compiler for the parallel processing system-harray. Lecture Notes in Computer Science, 757:482–496, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:2007:BSP

[YYX^{+}07] Xuejun Yang, Xiaobo Yan, Zuocheng Xing, Yu Deng, Jiang Jiang, and Ying Zhang. A 64-bit stream processor architecture for scientific applications. ACM SIGARCH Computer Architecture News, 35 (2):210–219, May 2007. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Zarea-Aliabadi:1993:LPD

[ZA93] Adib Zarea-Aliabadi. Languages for programming distributed memory systems. Computing and Control Engineering Journal, 4(6):269– ??, December 1, 1993. CO-DEN CCEJEL. ISSN 0956-3385 (print), 1741-0460 (electronic).

Zaghloul:2011:ACF

[ZA11] Mofreh R. Zaghloul and Ahmed N. Ali. Algorithm 916: Computing the Faddeyeva and Voigt functions. ACM Transactions on Mathematical Software, 38(2):15:1–15:22, December 2011. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [Zag16].

Zaghloul:2016:RAC

[Zag16] Mofreh R. Zaghloul. Remark on “Algorithm 916: Computing the Faddeyeva and Voigt Functions”: Efficiency improvements and Fortran translation. ACM Transactions on Mathematical Software, 42(3): 26:1–26:9, May 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [ZA11].

Zahn:1992:FPR

[Zah92] John J. Zahn. FORTRAN programs for reliability analysis. General technical report FPL GTR-72, U.S. Dept. of Agriculture, Forest Service, Forest Products Laboratory, Madison, WI, USA, 1992. 25 pp.

Zirkel:1994:UF

[ZB94a] Gene Zirkel and Eli Berlinger. Understanding FORTRAN 77 and 90. PWS Pub. Co., Boston, MA, USA, 1994. ISBN 0-534-93447-1. xvii + 670 pp. LCCN QA76.73.F25Z57 1994.

Zirkel:1994:UFR

[ZB94b] Gene Zirkel and Eli Berlinger. Understanding Fortran 77 and 90: Reference Guide Understanding Fortran 77 & 90. PWS-Kent Pub. Co., Boston, MA, USA, January 1994. ISBN 0-534-93447-1. US$66.95.

Zima:1992:VFLa

[ZBC^{+}92] Hans Zima, Peter Brezany, Barbara Chapman, Piyush Mahrota, and Andreas Schwald. Vienna Fortran: a language specification, version 1.1. Technical Report Series ACPC/TR 92-4, Austrian Center for Parallel Computation, [1992?], Vienna, Austria?, 1992. 86 pp.

Zima:1994:SVF

[ZBC94] Hans P. Zima, Peter Brezany, and Barbara M. Chapman. SU-PERB and Vienna Fortran. Parallel Computing, 20(10–11): 1487–1517, November 3, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/
browse/browse.cgi?year=1994&
volume=20&issue=10-11&aid=
909.

Zhu:1997:ALF

[ZBLN97] Ciyou Zhu, Richard H. Byrd, Peihuang Lu, and Jorge Nocedal. Algorithm 778. L-BFGS-B: Fortran subroutines for Large-Scale bound constrained optimization. ACM Transactions on Mathematical Software, 23(4):550–560, December 1997. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [MN11].

Zwolak:2007:AOW

[ZBW07] Jason W. Zwolak, Paul T.

Boggs, and Layne T. Watson. Algorithm 869: ODR-PACK95: a weighted orthogonal distance regression code with bound constraints. ACM Transactions on Mathematical Software, 33(4):27:1–27:12, August 2007. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Zima:1993:DDD

[ZCMM93] H. Zima, B. Chapman, H. Moritsch, and P. Mehrotra. Dynamic data distributions in Vienna Fortran. In IEEE [IEE93d], pages 284–295. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Zima:1995:CTS

[ZCP95] H. P. Zima, B. M. Chapman, and M. Pantan. Compiler technology for scalable parallel architectures — a short overview. In Hoffmann and Kreitz [HK95], pages 240–249. ISBN 981-02-2211-4. LCCN QC866.E26 1994.

Zhang:1992:FPD

[ZE92] Wenchao Zhang and K. E. Evans. A FORTRAN program for the design of laminates with required mechanical properties. Computers and structures, 45(5):919–??, December 1992. CODEN CM-STCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Zeichick:1992:WGI

[Zei92] Alan L. Zeichick. Watcom Group Inc.’s FORTRAN 77/ 386 v. 8.5. Computer Language Magazine, 9(3):17–22, March 1992. CODEN COMLEF. ISSN 0749-2839.

Zima:1992:VFLb

[Zim92]  Hans Zima.  Vienna For-

tran, a language specification. NASA contractor report NASA CR-189629, National Aeronautics and Space Administration, Langley Research Center; National Technical Information Service, distributor, Hampton, VA, USA, 1992. 86 pp.

Zima:1999:IHP

[Zim99] H. P. Zima. An introduction to HPF+ project. Lecture Notes in Computer Science, 1615:9– ??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zima:2002:HPF

[Zim02] Hans P. Zima. High Performance Fortran — history, status and future (HiWEP invited talk). Lecture Notes in Computer Science, 2327:490– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2327/23270490.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270490.
pdf.

Zima:2007:FLA

[Zim07] Hans P. Zima. From FOR-TRAN 77 to locality-aware high productivity languages for peta-scale computing. Scientific Programming, 15(1):45– 65, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Zachary:1995:ECC

[ZJEP95a] J. L. Zachary, C. R. Johnson, E. N. Eide, and K. W. Parker. An entry-level course in computational engineering and science. In Anonymous [Ano95b], pages 209–213. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zachary:1995:ELC

[ZJEP95b] Joseph L. Zachary, Christopher R. Johnson, Eric N. Eide, and Kenneth W. Parker. An entry-level course in computational engineering and science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27 (1):209–213, March 1995. CO-DEN SIGSD3. ISBN 0-89791-693-X. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

ftp://ftp.math.utah.edu/
pub/mirrors/ftp.ira.uka.
de/bibliography/Compiler/
fortran.bib;   ftp://ftp.
math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/
Compiler/FORTRAN/fortran3.
bib;   ftp://ftp.math.utah.
edu/pub/mirrors/ftp.ira.

uka.de/bibliography/Misc/
DBLP/1995.bib.

Zhou:1991:CAR

[ZMR^{+}91] Susan Y. J. Zhou, Sati Mazumdar, Carol K. Redmond, Michael H. Dong, and Joseph P. Costantino. Computations of adjusted rates and lifetime risks from occupational cohort data: a program package using FOR-TRAN and GLIM. Computers and Biomedical Research, 24 (1):29–46, February 1991. CO-DEN CBMRB7. ISSN 0010-4809 (print), 1090-2368 (electronic).

Zosel:1993:HPF

[Zos93] M. E. Zosel. High performance Fortran: an overview. In IEEE [IEE93a], pages 132–136. ISBN 0-8186-3400-6. LCCN QA75.5.C58 1993. IEEE catalog no. 93CH3251-6.

Zhou:1990:UGF

[ZT90] Jian Zhou and Andre Leon Tits. User’s guide for FSQP version 2.0: a Fortran code for solving optimization problems, possibly minimax, with general inequality constraints and linear equality constraints, generating feasible iterates. Technical research report TR 90-60r1, University of Maryland, Systems Research Center, College Park, MD, USA, 1990. 29 pp.

Zhu:1994:LFP

[ZZN94] Keyun Zhu, Xiaolei Zou, and I. M. Navon. LADFEUDX — a FORTRAN program for variational data assimilation with a finite-element shallow-water equations model. Technical Report FSU-SCRI-94-06, Supercomputer Computations Research Institute, Tallahassee, FL, USA, January 1994. 26 pp.

Zhang:2019:PFC

[ZZSW19] Z. Zhang, D.-B. Zhang, T. Sun, and R. M. Wentzcovitch. phq: a Fortran code to compute phonon quasiparticle properties and dispersions. Computer Physics Communications, 243(??):110–120, October 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0010465519301523.